A Complete Bibliography of ACM SIGMETRICS Performance Evaluation Review

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 September 2024 Version 1.58

Title word cross-reference

+ [2080, 2169]. 1/N [2716]. 3 [3420, 1970, 3331, 2849, 936, 2706]. 4 [2060]. 5 [2941]. $9.81 [3143]. α [1860]. c [1870]. d [3335, 2884, 2636, 2650, 3063, 1192, 1955, 2651, 2494]. Δ [2010]. G/G/∞ [2023]. k [1577, 3447, 3343, 3373, 2066, 3122, 3315]. λ(n)/C_{k}/1/N [189]. M [2921, 3446]. n × n [3095, 3065].

-Accurate [2716]. -dimensional [1955]. -Graphs [2010]. -optimal [1870]. -th [2921]. -TLB [1192]. -weighted [1860].

.NET [1821].

1} [1720, 1592, 1915, 1691, 2298, 1170, 671, 198, 2997, 807, 1230, 3343, 3373, 3062, 450, 1128, 320, 3445, 3024, 2640, 1169, 2289, 1792, 3389, 1439]. 1-FB [1348]. 1-type [1188, 1267]. 11 [267]. 11/780 [267]. 1100’ [419, 257]. 1100-of [419]. 1100/42 [257]. 145 [86]. 1992 [706].

2 [1963, 575]. 2-dimensional [2088]. 2-Level [1429]. 2.7 [3223]. 2000 [929, 666]. 2000-user [11]. 2010 [1916]. 2011 [2015]. 2016 [2693]. 2024 [3454]. 23-3 [310]. 256 [3123].

3-phase [3384]. 3-tiered [1332]. 3/23 [310]. 3/23-3/24/82 [310]. 3/24/82 [310]. 3000 [266]. 360 [24]. 370 [86]. 370/145 [86]. 3G [2286, 2608]. 3G/4G [2608].

42 [257]. 43XX [363]. 4G [2608].

5000 [738]. 52779 [127]. 5890 [555].

6 [3329]. 60 [38, 2475].

780 [267].

802.11 [1900, 1298, 1919]. 802.11-operated [1298]. 802.11e [1385]. 802.11s [1993]. 802.16e [1691].

9/30/91 [660].

a.k.a [2884]. abandonment [2009]. abandonments [2356]. ABR [932, 874]. Absorbing [3217]. Abstract [2566, 38, 2446, 3196, 2716, 3191, 2575, 2728, 206, 599, 2893, 946, 945, 1687, 2056, 2057, 2058, 939, 1170, 401, 691, 1695, 2052, 398, 940, 780, 781, 2059, 395, 2060, 396, 2061, 2062, 2063, 2064, 783, 2065, 1045, 1019, 1874, 2066, 2053, 2067, 2068, 784, 1044, 941, 2054, 635, 448, 2069, 1446, 942, 439, 2055, 696, 2135, 2070, 2071, 786, 2072, 801, 787, 2073, 943, 944]. abstracting [2252]. Abstraction [1374, 818]. Abstracts [87, 2015, 2016]. AC [2795]. Academic [9, 642]. ACCEL [1299]. ACCEL-RATE [1299]. Accelerated [2724, 3296]. Accelerating [2730, 3282, 2492, 2251]. acceleration [2177]. Acceptance [2065, 2450, 117]. Acceptance-rejection [2065]. Access [2785, 2924, 607, 2740, 3010, 3068, 2839, 3142, 705, 946, 1146, 2142, 1645, 1991, 2233, 2784, 1609, 1083, 3276, 1868, 694, 464, 2293, 614, 454, 1189, 558, 1049, 1921, 707, 1836, 484, 1424, 826, 1051, 1879, 690, 483, 2887, 1919, 374, 2277, 1837, 1481, 1901]. Access-Delay [3010]. Accessibility [3425]. accessing [285]. accounting [2266, 235, 246]. Accuracy [1596, 2936, 725, 1569, 3228, 1397]. Accurate [2731, 2400, 2548, 2716, 540, 759, 795, 1254, 1985, 1059, 821, 1283, 1320, 1383, 1301, 2363, 1794, 1430, 1411, 1422]. ACE [2252]. Achievability [2591]. Achievable [3128, 1889, 867, 1034]. Achievement [2808]. Achieving [2605, 3181, 1921, 1403, 2622, 3127, 2930, 1459]. acknowledgement [483]. acknowledgement-based [483]. acknowledgements [943]. ACM [3356, 645, 3346, 3066, 3176, 2865, 2899, 2892]. Acoustic [2835]. Acquisition [3158, 3397, 1143]. Across [2744, 2131, 2372, 2563, 2905, 2658, 603, 1946]. Action [3329]. Active [818, 495, 2773, 1394, 1245, 2192]. activities [1362, 749]. Activity [2446, 3392, 1756, 1884, 1466, 1202]. Ad [2964, 1212, 2320, 1325, 1115, 2010, 1084]. Ad-Hoc [2964, 1212, 1115]. ADAM [74]. adaptability [3228]. Adaptation [2846, 1196]. Adaptive [44, 2713, 231, 1213, 900, 879, 2894, 2596, 1563, 2738, 2774, 2533, 3428, 2462, 585, 513, 1163, 1306, 1997, 91, 1276, 773, 1261, 395, 1413, 396, 1809, 1778, 1815, 989, 724, 1879, 2097, 1967, 253, 46, 981, 1435]. Adaptivity [3359]. added [985]. Addison [606, 608, 645, 700]. Addison-Wesley [606, 608, 645]. addition [21]. Additional [2829, 800]. additive [1070, 2695]. Address [996, 736, 1032, 754, 843, 769, 972, 752]. adiabatic [1763]. adjustment [999]. Admission [2096, 1196, 1163, 1778, 1895]. Admit [2557]. Adoption [3413, 2916, 2440, 2167, 2027]. ADP [164, 84, 214, 2124]. adPerf [3133]. Adrian [644, 701]. Ads [3133]. adsorption [2424]. Advance [2445, 2577, 1497, 1625]. Advanced [1750, 2894, 3044, 3114, 2939]. Adversarial [2845, 3060, 3412, 2820, 3035, 2957, 3314, 2373]. adversaries [3382, 1661, 1297]. advertisement [1797]. Advertising [2675, 2984, 2444, 1899, 2329]. advisor [563]. Aeronautical [2943]. affecting [616, 574]. affects [2029, 1022]. affinities [1039]. affinity [833, 741, 2772]. after [2766]. Against [3041, 3035, 3382, 2168]. Age [3055, 3267, 3372, 2822, 1761, 2339, 1749]. Age-Based [2822]. Age-of-Information [3055]. agenda [1819]. Agent [3275, 3269, 3311, 3319, 3297]. Agents [3399, 2567, 2625, 2616, 1216]. aggregate [343, 1350, 1188]. aggregated [1162]. aggregates [1577, 1773]. Aggregation [644, 2973, 3118, 2942, 701, 263, 2130, 1654, 348, 1267]. aggregation-based [1267]. Aggregators [2655]. Aggressive [3300]. aggressiveness [2293]. Aging [2003, 3094, 887]. Agnostic [3008]. Agreement [2644, 1172]. Agreements [3142]. Ahead [2585, 2982, 2701, 2670]. AI [3134, 3400, 3251]. aid [113]. aided [1685]. aids [306, 162]. Aimbots [2633]. AIMD [3210]. Air [2941, 409]. Airborne [2962]. Airline [2667]. alchemical [219]. algebra [1929]. algebraic [763]. algebras [1017]. Algol [38]. Algol-60 [38]. Algorithm [3446, 2752, 3014, 3018, 2994, 3258, 2485, 2549, 3083, 3088, 3145, 2825, 2819, 1262, 2291, 1234, 1869, 1300, 1480, 398, 437, 1560, 359, 1261, 1535, 1252, 8, 2061, 772, 2405, 1845, 1070, 1381, 989, 1810, 1109, 1028, 261, 759, 482, 1175, 848, 444, 2213, 639, 1366, 2343, 2018]. Algorithmic [2824, 1387, 679, 1699, 2972, 2985, 1511]. algorithmization [2054]. Algorithms [3029, 3045, 2918, 3055, 2472, 2590, 3250, 3441, 2488, 2487, 2628, 2989, 3030, 3310, 3117, 3076, 2520, 2737, 644, 3149, 2479, 3146, 3091, 3189, 3245, 3453, 3409, 3410, 2623, 2881, 2745, 3312, 2766, 3147, 701, 2573, 3037, 2519, 2860, 3038, 2948, 370, 1222, 984, 1805, 1389, 178, 364, 1409, 2409, 519, 1097, 2232, 1679, 1622, 1301, 1478, 433, 570, 466, 1107, 748, 145, 1023, 554, 514, 2248, 413, 459, 419, 481, 962, 2055, 487, 2239, 585, 995, 2347, 777, 2349, 1464, 1411]. Alignment [3014]. All-or-Nothing [3188]. all-to-one [2319]. Allen [642]. Allocation [2790, 3416, 3309, 3305, 3119, 2932, 2647, 3263, 3400, 2905, 2477, 2648, 3253, 2467, 2478, 3252, 3293, 2920, 2883, 3255, 3019, 3050, 2535, 2637, 2638, 2992, 3415, 2464, 2558, 2726, 808, 31, 275, 1623, 1194, 364, 1765, 2307, 2281, 2431, 1625, 885, 756, 1948, 2885, 844, 560, 1336, 1296, 419, 1036, 2229, 515, 2390, 297, 1626, 1723, 1865]. Allocations [2828, 2372, 2532]. Allow [2450]. AlloX [2905]. Aloha [1443]. alternating [2175]. Alternative [1143, 737, 2043, 114]. alternatives [528, 194]. Ambivalence [3427]. Amdahl [555, 649, 1786]. Among [3369, 1025, 227]. amongst [1508]. Amplify [3427]. AMVA [1061]. AN1 [681]. analyses [659]. Analysi [595, 607]. Analysing [2649, 3099, 872]. Analysis [3008, 2862, 382, 461, 2630, 641, 1943, 2946, 1185, 1721, 2204, 2291, 1158, 1129, 1041, 2074, 651, 2579, 3123, 2731, 1531, 3196, 2528, 1449, 2613, 2718, 2481, 2769, 2850, 3250, 3290, 3406, 449, 3014, 1503, 3395, 1535, 1009, 597, 605, 646, 658, 2542, 647, 2647, 2496, 2917, 3349, 3370, 2581, 2934, 2969, 868, 2340, 3091, 3073, 3358, 3259, 2543, 1735, 3028, 789, 2720, 1544, 3272, 2881, 3163, 3298, 3113, 3218, 2778, 732, 1508, 1523, 653, 2658, 670, 3080, 3348, 2617, 2852, 954, 3078, 1670, 3180, 755, 3454, 204, 3009, 3034, 3445, 674, 3436, 3291, 33, 2822, 3213, 2629]. Analysis [2709, 626, 1012, 1102, 1087, 1157, 1227, 1433, 2876, 3016, 3053, 3156, 3234, 3332, 3439, 2717, 1428, 385, 418, 564, 2072, 194, 1108, 3025, 3085, 860, 202, 3142, 2536, 2675, 2639, 2841, 705, 1293, 2093, 685, 2345, 1941, 1212, 3160, 426, 2105, 1994, 338, 1316, 1159, 3057, 1570, 1737, 736, 2412, 2188, 1016, 1922, 714, 507, 1132, 687, 2346, 50, 350, 951, 1480, 1646, 1740, 1839, 1004, 1702, 1237, 1094, 834, 856, 1979, 621, 1368, 1405, 1986, 163, 2193, 398, 40, 280, 319, 473, 768, 2080, 562, 1700, 159, 160, 298, 804, 245, 3384, 1603]. analysis [3455, 37, 3059, 2236, 491, 1718, 566, 1889, 1924, 1215, 622, 1720, 2036, 1325, 2162, 1732, 2344, 496, 100, 2111, 1550, 1372, 1008, 698, 558, 920, 1845, 1065, 432, 212, 1049, 501, 1908, 1747, 669, 1290, 1264, 378, 1454, 506, 739, 1907, 2238, 557, 619, 1173, 446, 442, 156, 2254, 924, 693, 728, 612, 2192, 618, 676, 22, 577, 1380, 554, 391, 2172, 556, 934, 468, 283, 635, 1877, 448, 1343, 1878, 2017, 1753, 514, 1926, 761, 1330, 482, 2253, 711, 1294, 876, 1385, 1005, 628, 34, 1267, 299, 305, 479, 1779, 1175, 1162, 494, 572, 573]. analysis [1644, 1006, 2215, 203, 344, 242, 576, 763, 1021, 1184, 1542, 2768, 848, 3104, 141, 441, 444, 592, 451, 722, 2393, 1244, 677, 1455, 1754, 1668, 353, 2381, 708, 1481, 1792, 429, 2407, 743, 400, 550, 1055, 1304, 1558, 668, 1238, 2428, 2305, 2772, 260, 262, 369, 1081, 1092, 1504]. analyst [301]. Analytic [409, 498, 256, 145, 2089, 408, 273, 272, 149, 2085, 144, 721, 855, 1177, 420, 813, 930, 1328, 567]. Analytical [627, 3032, 276, 684, 1993, 2215, 1867, 1234, 90, 613, 164, 2385, 1485, 1066, 12, 2012, 1169, 1406, 2927]. Analytics [2973, 2351, 2885, 2342, 2339]. Analyze [2877, 1733]. Analyzer [3179, 1750, 181, 2334, 324]. Analyzing [1145, 1221, 870, 2618, 2795, 2683, 2554, 1588, 3355, 3102, 2669, 733, 1498, 2984, 343, 268, 1900, 100, 1968, 667, 1142, 2280, 1414]. Anarchy [2741, 1873]. Anatomy [665, 2368, 2385]. andlearning [3455]. Android [3069, 3003]. animation [1418]. annotating [1914]. anomalies [2150, 1333, 1885]. Anomaly [3112, 2775, 1810, 1139, 1554, 1806, 1759, 1964]. Anonymity [2810, 3322, 2723, 2221]. answered [790]. Anthony [751]. Anticipating [2556]. Antipattern [3219]. anycast [1423]. APC [2182]. aperiodic [615]. APL [322]. App [3110, 3003, 3224]. appear [1438]. Apples [1922, 5]. Apples-to-apples [1922]. Appliances [2659]. applicability [410, 836]. Application [2748, 2780, 1207, 933, 3286, 2641, 3370, 3379, 3069, 221, 2480, 497, 2580, 3258, 3410, 1028, 3362, 3235, 2783, 3147, 2940, 602, 3136, 3003, 54, 1584, 367, 820, 1237, 757, 1067, 243, 829, 164, 3384, 485, 632, 2409, 1932, 2150, 504, 1756, 1704, 1295, 1853, 2123, 1268, 327, 1484, 934, 635, 2208, 638, 266, 2155, 2070, 119, 1481, 2399, 1621, 3282]. Application-level [3379, 1237, 1484]. application-VM-physical-machine [2409]. application/file [1067]. application/file-level [1067]. Applications [2461, 3045, 3090, 3030, 642, 952, 2679, 2469, 2581, 3246, 3443, 3331, 2803, 3385, 84, 2736, 3236, 2792, 2158, 511, 2278, 1163, 1076, 994, 493, 993, 1359, 1711, 1830, 781, 805, 1617, 807, 1469, 1852, 624, 2157, 2063, 85, 415, 872, 1097, 1588, 832, 1923, 1872, 238, 1285, 3110, 1142, 579, 936, 1854, 1559, 2301, 2427, 541, 1310, 2215, 1657, 816, 1122, 1886, 1102, 715, 107, 2201, 1244, 1840, 1406, 1202, 1042, 2212, 2420, 3197, 789]. Applicationson [2524]. applicative [196]. Applied [102, 3353, 248, 1691, 231, 1638, 1092]. Applying [1504, 1734, 1821, 1965]. Apportionment [3301]. Approach [2474, 382, 2823, 2687, 2859, 2821, 2963, 2866, 2689, 2555, 2973, 2486, 2605, 2620, 3174, 2680, 2972, 2951, 2888, 2717, 2462, 3050, 3005, 2877, 2878, 2759, 3116, 946, 2746, 1842, 3232, 2074, 1619, 616, 1645, 2422, 223, 275, 401, 1434, 713, 1405, 138, 1025, 1276, 15, 219, 2160, 2307, 2019, 1594, 1546, 753, 2949, 525, 865, 1372, 1640, 45, 1068, 1736, 1437, 2262, 553, 2410, 2167, 679, 1302, 309, 1553, 315, 1858, 283, 560, 1918, 1502, 1440, 1005, 462, 474, 1404, 770, 220, 985, 588, 742, 59, 1228, 1800, 2930, 831]. Approaches [2758, 3032, 2497, 1216, 125, 423, 98, 136, 1699, 1758, 1824]. Approximate [589, 830, 2550, 440, 1603, 3091, 273, 558, 2977, 3381, 3150, 761, 2593, 743, 370, 1141, 342, 856, 1560, 562, 652, 1215, 516, 1495, 1060, 514, 1175, 414, 708, 369, 1081, 1344]. approximately [3166]. Approximating [1613, 543]. Approximation [3216, 2548, 2716, 2853, 2886, 2917, 3146, 3318, 3084, 2610, 383, 1093, 2079, 32, 2109, 1897, 2009, 728, 618, 542, 1896, 1105, 1533, 2416]. Approximations [1793, 2804, 2808, 3387, 3054, 3130, 520, 2715, 1475, 453, 1123, 2303, 1996, 2046, 1605, 2302, 1173, 1903, 2289, 1358]. Apps [3300, 3124, 2333]. AQM [1123, 1311]. arbiters [273, 432]. Arbitrage [3431]. Arbitrary [3188, 2698, 3267, 3372, 1791, 1869, 1678]. Architectural [3280, 2581, 1504, 993, 2274, 1912, 1030, 564]. Architecture [869, 1071, 2977, 896, 2096, 1632, 1279, 692, 888, 1277, 1178, 408, 582, 386, 1249, 518, 1283, 1581, 1932, 2219, 1547, 666, 2231, 852, 1908, 2138, 506, 1796, 1655, 2337, 1285, 476, 1484, 1804, 1945, 1800, 2251, 1493]. Architecture-Aware [2977]. architecture-level [1932]. Architectures [1080, 3189, 3070, 3190, 1222, 2156, 1924, 430, 458, 1827, 1814, 1923, 666, 778, 1559, 890, 1031, 456]. archives [707]. Area [658, 304, 306, 870, 1011, 1076, 1004, 352, 1932, 522, 1201, 249, 1454, 309, 276, 381, 533, 1127, 2147, 1715, 482, 2135, 427, 1574]. Areas [2872, 1980]. Argus [1963]. Argus-2 [1963]. arising [1662]. ARM [2245]. ARM-based [2245]. ARMA [445]. armed [2082, 2829, 2906]. arms [6]. ARMv8 [3189]. Arnold [642]. ARPANET [293]. array [1096, 810, 630, 2249]. arrays [555, 866, 662, 721, 1460]. Arrival [762, 3163, 3267, 3372, 2074, 1985, 2090, 1514, 1546, 2065, 2153, 3062, 1095, 2087, 968]. arrivals [1489, 1612, 1252, 1790, 1544, 2086, 1595, 295, 2199, 1102, 2919]. arrivals/departures [1595]. Art [641, 595]. Article [588]. artificial [365]. artistry [2]. as-level [1663, 1217, 1739, 1410]. ASIP [3374]. asked [3209]. Aspects [203, 2094, 1387, 347]. assemble [1173]. assemble-to-order [1173]. assembling [1729]. assembly [1355, 843]. Assessing [1365]. Assessment [2697, 2972, 2094]. asset [2255]. Assets [3040]. Assignment [2592, 2759, 1721, 940, 1776, 2317, 523]. assist [1977]. assistance [2068]. Assisted [2753, 2561, 2384, 2533, 1671, 1858]. Association [2914, 2535]. associative [931, 1852, 634]. associativity [1828]. assumption [1900]. assumptions [217]. Assurance [99, 324, 129, 115, 120, 106, 130, 102, 107, 128]. Asymmetric [3344, 834, 920, 2257]. Asymptotic [760, 2730, 2826, 1889, 2432, 2431, 3376, 3259, 3129, 1165, 2789, 3173, 921, 494, 2503, 3326, 3127, 1533, 1358, 1645, 1986, 1354, 1095, 1105, 2788, 1256]. Asymptotically [2864, 3162, 3038, 3079, 1262]. Asymptotics [2612, 2304, 2911, 2085, 2066, 2434, 1058, 1474]. Asynchronous [847, 2724, 739, 3289, 1587, 2205, 519, 2110]. At-Scale [3406]. ATM [891, 805, 1009, 557, 953, 822, 874, 909]. attached [893, 855]. Attack [2739, 2702, 2337, 3105, 2271]. Attacker [2695]. Attacks [3412, 2761, 2835, 2897, 3056, 2958, 3041, 3167, 3187, 2908, 2497, 2766, 2893, 3035, 3236, 3231, 2767, 1806, 2112]. Attention [2675]. Attraction [2995]. Auction [2924, 2438, 2558, 2352, 2354]. Auctions [2450, 2470, 2328, 1515]. Augmentation [3429]. Augmented [3360, 3409, 3277, 3164, 3157]. Augmenting [3260]. AURORA [2660]. Authors [307]. Auto [2178, 2409, 3179]. Auto-generation [2178]. auto-scaling [2409]. Auto-SDA [3179]. autocorrelated [1624]. Autocorrelation [1979, 1602]. Automata [3289, 1373]. automate [1327]. Automated [3179, 3400, 2493, 2619, 1046, 2393, 2259, 242, 2124]. Automatic [1713, 1484, 1703, 3227, 1391, 1228, 116, 1324, 1862].

Automatically [1820]. Automation [1131]. automatization [1712]. Autonomous [3044, 3135]. Autonomy [3399]. autoregressive [2336]. Autoscaling [3273, 2396]. autotuning [2175]. Availability [973, 2662, 861, 1274, 397, 1200, 1835, 1788, 1813, 1415, 1513, 526]. available [1811, 1398, 1496, 767, 1322]. Average [1159, 2737, 3342, 1269, 1892]. Average-Cost [2737]. averaging [1941, 1881]. Avoidance [3101, 2939, 1053]. avoiding [2044, 2256]. Award [3356, 2728]. Aware [3294, 3203, 2682, 3386, 2977, 3229, 2563, 3410, 2620, 3192, 3023, 3157, 2839, 3390, 2474, 2096, 1686, 1950, 1676, 2796, 2149, 2186, 2103, 2405, 1305, 1704, 1622, 2115, 1936, 1284, 1520, 2594, 1652, 1804, 2539, 1891, 2516, 1208, 1176, 2272, 2558]. Awareness [3362]. Away [3438]. aways [3115]. AWS [2554]. Axiomatic [3185]. Axiomatizing [2991]. Azure [2793].

B [640, 594, 596, 663, 3090]. B-Matching [3090]. B.V [646]. Back [3413, 2621, 2622, 355]. Back-off [2622]. back-up [355]. Backbone [1555, 2663, 1209, 581]. backbones [1683]. Backend [3400, 1831]. backfilling [1179]. background [1295, 1769]. Backhaul [2841]. backlog [1991, 1600]. backoff [1784]. backtrack [2127]. backup [2195]. Bad [2580]. Bagrodia [588]. Balanced [2859, 2372, 2890, 2691, 2631, 2863, 260, 1475, 1164, 860]. Balancing [2862, 2867, 2562, 3075, 2970, 2969, 3140, 2612, 3229, 2623, 2881, 2651, 2864, 2658, 3139, 3291, 2971, 3325, 2755, 3138, 2471, 2860, 2855, 2913, 2968, 3038, 3079, 3283, 1318, 1863, 1722, 2191, 900, 1864, 2198, 1825, 517, 840, 496, 2417, 2104, 435, 434, 1953, 2031, 289, 2267, 2195, 1412, 2002, 1304, 2313, 812, 1091]. balking [792]. Band [3277, 2366]. Bandit [2986, 3174, 2829, 3116, 2906, 2082]. Bandits [2488, 3117, 2826, 3263, 3308, 3173, 3450, 3311, 2369]. Bandwidth [2523, 2506, 2637, 2800, 971, 735, 3420, 1811, 1818, 1535, 986, 1996, 2444, 1398, 1392, 2532, 1068, 1496, 1394, 1940, 1595, 2008, 1048, 2113, 903, 1757, 1322, 816, 1499, 1626]. Bandwidth-efficient [971, 1048]. bandwidth-sharing [1499, 1626]. bandwidths [1018]. banked [2121]. bankrupt [2166]. Banyan [728, 618, 684]. Barrier [3270, 3320, 174, 2683]. Base [2662, 2962, 2711, 338, 205]. Based [2823, 2713, 2696, 644, 2589, 2724, 2648, 3345, 2625, 2742, 599, 2888, 2822, 2971, 3227, 2492, 3236, 701, 2637, 2501, 2984, 3321, 2733, 2294, 2570, 1221, 1040, 992, 1355, 379, 659, 2682, 3419, 1991, 869, 1688, 274, 2531, 3108, 1004, 1767, 2850, 3423, 1399, 1086, 891, 2719, 1056, 1181, 1025, 1365, 1307, 2770, 521, 1963, 1562, 3442, 2479, 3179, 1594, 879, 882, 1059, 1089, 392, 1037, 1546, 1969, 2521, 518, 1420, 2499, 872, 2777, 122, 1235, 1675, 3246, 1622, 1748, 1710, 1115, 2973, 3258, 435, 673, 2180, 1378, 1044, 1690, 3033, 989, 1611, 873, 1003, 3110, 1053, 924, 771, 1699, 1758]. based [502, 3275, 1633, 2922, 3348, 3398, 1117, 1180, 2132, 166, 703, 1880, 1104, 1148, 1899, 1294, 1440, 1267, 3233, 561, 1806, 675, 2246, 1883, 2229, 483, 2418, 3088, 1404, 1205, 2016, 848, 2179, 3035, 836, 722, 418, 2776, 2245, 373, 787, 3254, 1920, 2239, 2995, 588, 1516, 1152, 270, 1558, 1794, 2341, 1526, 639, 1039, 1533, 1574, 2950, 2018, 625, 2968, 1587, 3189, 2943]. Basic [1784, 1739, 2157, 2268]. Basics [2709]. Batch [2988, 828, 3096, 45, 2464, 1514, 2299, 68, 78, 881, 2919]. batch-networks [2299]. Batch-Processing [3096]. Batched [1463, 1159]. Batching [3322, 1972, 432, 2012]. BATS [2396]. Battery [2662, 2936, 2309, 2759, 2226]. battery-diesel [2309]. Bayes [2491]. Bayesian [2324, 553, 234, 1380, 391, 3006]. BCMP [436]. BDAS [2361]. BE [283, 2552, 1492, 1753, 790]. Beamforming [2932, 3330]. Beams [2475]. Before [2906]. beginning [600, 1527]. Behaving [2616]. Behavior [2525, 2551, 2789, 2751, 2549, 2503, 2928, 1145, 984, 1198, 2275, 651, 1619, 689, 1093, 1123, 1981, 931, 1851, 879, 27, 1192, 524, 1065, 2367, 434, 2021, 884, 1362, 2154, 636, 827, 921, 918, 1882, 483, 930, 1121, 2545, 1754, 733, 583, 944]. Behavioral [531, 2571]. behaviors [2161, 2094]. Behaviour [2659, 184, 445, 446, 512, 402]. Behind [3126]. Benchmark [57, 650, 657, 3154, 2493, 610, 3002, 79, 2178, 662, 575, 1501, 666, 941, 1926, 603, 699, 1644, 592]. Benchmarked [24, 26]. Benchmarking [3175, 182, 750, 1925, 3355, 2180, 158, 888, 144, 1936, 1220, 1944]. Benchmarks [211, 1246, 2948, 1708, 1820, 713, 955, 1192, 593, 665, 887, 1931]. Beneficial [2741]. beneficiary [1445]. beneficiary-donor [1445]. Benefits [2560, 741, 495, 842, 1334, 2243]. Bernoulli [321, 728]. Best [3354, 2923, 2966, 265, 1167, 998, 1532, 999, 1677, 1712]. best-effort [998, 1677]. BestFit [2432]. bestseller [2333]. Better [959, 2482, 2689, 3285, 2592, 1577, 5, 1488]. between [1949, 2805, 2681, 986, 1348, 3384, 478, 1822, 2936, 2403, 3142, 2354, 2509, 1855]. Beware [3]. Beyond [1541, 3421, 2344, 3229, 2127, 3247, 3415, 117]. BG [1925]. BG/P [1925]. BGP [1314, 3413, 1315, 2649, 1770]. Bias [3292, 3318, 3158]. Bias-Variance [3158]. Biased [3308, 2625, 2418, 2143]. Biases [2571]. bibliography [358]. Bidding [2328, 3182, 2670]. Big [2461, 2460, 2514, 2463, 2338, 2501, 2733, 2345, 2337, 1491, 2342, 2339, 2361, 1799, 2341]. BigData [2343]. bigger [1457]. bike [3160]. Bilateral [2057, 2509]. Billing [2443]. Billion [2840, 2729]. Billion-Scale [2840, 2729]. Bin [3045, 2692, 2521, 3433, 3306, 2494, 2432]. Bin-Packing [3433]. binary [1869, 2061, 2349]. binomial [2919]. binpacking [1167]. biochemical [2685]. biological [1641, 1638, 1640]. biology [1639]. Bipartite [2104, 3077, 3208, 2759, 1890]. bird [311]. Birkhoff [1164]. birth [2072]. Bistro [1076, 1178]. bit [1567]. Bitcoin [3103, 3104, 2446, 3080, 2723]. Bitcoin-NG [3080]. BitTorrent [1425, 1871, 1966, 2107, 1571, 1724, 1670, 1692, 1973]. BitTorrent-like [1724, 1670, 1973]. Black [2758, 1816, 3115]. Black-box [2758, 1816]. Blackbox [1834]. BlackjackBench [2174]. BlenX [1641]. bloat [2122]. Bloated [3405]. Block [3086, 3283, 1067, 2952, 2268, 185, 377, 373]. Blockchain [2955, 3100, 3352, 2951, 3101, 3324, 2948]. Blockchains [3201, 3099]. blocked [385]. Blocking [552, 1114, 2612, 2588, 358, 147, 34, 203, 1136, 1556]. BlockSim [2955]. bloom [1852, 2599, 3395, 2506, 1569, 3396]. Blue [1501]. Bluetooth [1610]. BMAP [1439]. BMAP/MAP/1 [1439]. Boltzmann [3347]. BONES [3417]. Book [641, 750, 640, 595, 594, 596, 598, 597, 604, 607, 606, 605, 608, 751, 746, 745, 654, 653, 663, 700, 701, 705, 117]. Boolean [763]. Boosting [2662, 2738, 2774, 1153]. Bootstrapped [2811]. BoPF [2900]. Bot [3431]. both [1154]. Botnet [1753]. Bottlenecks [3426, 814, 633, 667, 2277]. Bottom [1577]. Bottom- [1577]. Bought [3143]. Bound [1646, 2997, 2705, 349, 1965, 376, 639, 260]. Bounded [2840, 2966, 2550, 1167, 2109, 2367, 2273]. Bounding [1584, 1482, 1511, 882, 526, 3168, 679, 774]. Bounds [2790, 2817, 2804, 2921, 3341, 2488, 2487, 2591, 3447, 2469, 3178, 3340, 3376, 2571, 2495, 2911, 294, 2250, 1306, 1600, 2269, 440, 807, 2198, 1969, 678, 1020, 1671, 1590, 1573, 1397, 1782, 1393]. bowl [1656]. box [2758, 1816, 2366, 3115]. BRADO [1589]. braids [1655]. branch [1979, 639]. branch-and-bound [639]. breakdowns [691]. Breaking [174, 3383]. Breathing [2597]. Bregman [3151]. Bregman-style [3151]. BrickX [2039]. Bridge [1998]. bridges [522]. Bridging [1317]. Brief [643, 642, 645, 644, 646, 657, 656, 658, 655, 647, 66]. bring [2129]. Bringing [839]. broadband [933, 1362, 2277, 1837]. broadcast [1132, 289, 924, 612, 535, 1048, 2012]. Broadcasting [3088, 740]. Broadcasting-based [3088]. brokerage [219]. Brown [2460]. Brownian [1731, 2069]. Browser [3323]. Browsers [3009, 3034]. Browsing [2916, 2154]. BSTs [1330]. BT [1931]. bucket [1034]. budget [2319, 2396]. budget-constrained [2396]. Budgeted [2664]. Budgets [2488, 3415]. Buffer [3419, 1514, 2998, 897, 3198, 2819, 760, 1762, 1389, 1388, 1067, 1612, 562, 627, 491, 1183, 2405, 2120, 864, 923, 732, 670, 1117, 1642, 287, 297]. buffered [470, 146, 2303, 1419, 619]. bufferless [1535]. Buffers [3279, 534, 584, 2273, 1386, 539, 377]. build [1245, 2221]. Building [1985, 421, 1569, 2152, 2156, 1076, 1821, 377, 2039]. built [1219]. built-in [1219]. bulk [529, 1771, 420]. Bundling [2107, 1966]. Burman [2152]. Burst [3303, 1262]. Burstable [2722]. Burstiness [2900, 1646]. Burstiness-Fairness [2900]. bursty [1052, 1706, 1820, 834, 1114, 1396, 287]. bus [938, 977, 1225, 621, 433, 381, 535, 836]. bus-based [836]. buses [285]. Business [161, 592, 1145, 160, 1351, 1362, 790]. Busy [1718, 922, 283, 1051]. Byzantine [2845, 3269]. Byzantine-robust [3269].

C [597, 419, 699]. C240 [580, 590]. CA [2530, 2622]. cable [1010]. Cables [3437]. CAC [1023]. Cache [2544, 2921, 716, 2642, 2479, 2743, 748, 884, 2742, 2990, 779, 602, 2907, 2614, 2412, 1805, 1022, 274, 814, 1389, 1388, 768, 1485, 519, 1547, 1725, 1183, 565, 1545, 2120, 702, 2173, 732, 690, 1681, 561, 833, 813, 428, 741, 733, 1312, 429, 550, 668, 2383, 2836]. cache-affinity [741]. cache-based [274]. cache-coherent [565]. CAche-conscious [2383]. cache/memory [1547]. Caches [2595, 2704, 3061, 2669, 1024, 1828, 931, 2140, 2246, 835, 930, 712, 715, 1944, 1219]. Caching [2836, 2687, 2713, 2817, 3052, 2914, 2956, 2641, 2634, 3285, 2596, 3010, 3087, 3312, 3084, 2903, 2920, 2818, 2412, 714, 957, 815, 1866, 1045, 863, 1211, 1515, 2154, 1038, 942, 2398, 2545, 878, 1030, 916, 1039]. CADENUS [1365]. Cady [594]. Calculate [2994]. Calculating [189, 819]. calculations [1514]. Calculus [2731, 2629, 3383, 2269, 1490, 1717, 2541, 560]. calibrating [914, 182]. calibration [1436]. California [1232]. call [1489]. calls [1204]. Cambridge [354]. Campaigns [2583, 1753]. campus [1084]. CAMs [1492]. Can [1887, 1219, 3457, 3193, 1492, 2293, 2592, 2169, 1753, 2363, 790, 1126, 2112]. cancel [2884]. cancel-on-start [2884]. CANIT [1234]. Canonical [3014]. Cantor [1800]. Capabilities [3002, 800]. capacities [1508]. Capacity [167, 2703, 2698, 1731, 1618, 594, 2700, 2738, 2774, 1115, 2467, 173, 3164, 2783, 3077, 3208, 2992, 3420, 1842, 2164, 1564, 1680, 2426, 42, 1207, 1237, 1986, 1469, 2431, 478, 1320, 2402, 450, 1387, 503, 192, 1606, 2007, 2283, 2011, 1896, 1030, 801, 894, 214, 2311]. Capital [3427]. CapProbe [1320]. CApRI [2383]. Capsules [3304]. Capturing [3304, 1217, 1130]. Car [2579, 2696, 3059, 2620]. car-sharing [3059]. Carathéodory [2593]. Carbon [3411, 3410, 1875, 2283]. Carbon-Aware [3410]. Carbon-Efficiency [3411]. carbon-neutral [2283]. CarbonScaler [3411]. cardinality [1695]. cards [1676, 1340]. Carlo [3013, 1522]. carrier [1609]. CART [1331]. Cartography [3437]. Cascade [2729]. cascaded [776]. Cascades [2967, 2118]. Cascading [2795, 2764, 3042, 2188, 2214, 2021, 2254]. Case [3203, 2732, 3056, 3092, 2756, 3415, 1159, 95, 2412, 888, 2921, 493, 1067, 1032, 76, 1365, 531, 782, 478, 1925, 2417, 221, 1191, 1370, 635, 1880, 1005, 372, 1006, 1919, 3236, 803, 788, 791, 1176, 59, 214]. cases [2342]. Casualties [2463]. CAT [1241]. catastrophes [2062, 2919]. Catchment [2975]. categorization [1703]. catenet [486]. CAUCHY [2000]. caught [1193]. Causal [3249, 3009, 3093, 3363, 2831]. cause [1140, 2293, 2231]. caused [1551]. causes [853, 1668]. causing [2359]. CC [794]. CC-NUMA [794]. CD [407]. CDN [2169]. cell [1319, 1680, 2384, 1640, 1385, 1044]. Cell-based [1044]. cell-biological [1640]. Cello [917]. Cellular [2517, 2662, 2619, 3183, 1960, 2841, 2032, 1057, 1114, 1113, 2276, 1836, 1959, 2106, 2225, 2135]. Censored [2814]. Center [2537, 3060, 868, 3407, 637, 2755, 2458, 2278, 2426, 2415, 2148, 2019, 11, 2413, 2256, 2355, 2283, 1835, 2255, 1984, 1701, 2313, 2317, 789]. Centers [3141, 2681, 2504, 2702, 2761, 2700, 2835, 2455, 2754, 2794, 2798, 2803, 2908, 2519, 2034, 1407, 2114, 3228, 1968, 2115, 2223, 2185, 1952, 2145]. central [141, 142, 2919]. Centralities [3037]. centrality [2117]. Centralization [3299, 2649, 1947]. Centralized [2455, 1216, 417]. centres [368]. Centric [2920, 2038, 1350, 1579]. certain [209, 179]. certainty [2308]. certification [1, 418]. CFTL [1879]. Chain [3015, 398, 2262, 1522, 618, 282, 387]. Chains [3218, 3027, 1099, 1261, 2236, 1187, 772, 1343]. Chair [3356]. Chairs [3098, 3367, 3368]. challenge [2200, 1465]. Challenges [3069, 1815, 1458, 2580, 2223, 580, 590, 2585, 2052, 1072, 1917, 2342, 1984, 2338, 1824, 2341, 1823]. Chance [2692]. Chances [3422]. Chandy [588]. Change [3428, 1213]. changes [616, 235, 246, 1551]. changing [1349]. Channel [3294, 1868, 2835, 3402, 3451, 2783, 984, 1762, 834, 408, 1431, 1470, 484, 1775, 147, 374, 894, 1723, 2927, 837]. channel-exit [484]. Channels [2474, 3058, 2376, 2282, 288, 1674, 1471, 1111, 483]. characterisation [1029]. Characterising [512]. Characteristics [357, 1915, 511, 1767, 1937, 757, 1384, 2274, 1000, 406, 1483, 854, 1045, 1211, 374, 306, 1931]. Characterization [1146, 1602, 2613, 2718, 3406, 694, 2764, 2720, 2684, 599, 1311, 1007, 2524, 477, 850, 1010, 659, 2004, 2092, 1067, 2174, 965, 781, 245, 423, 531, 478, 333, 581, 332, 1581, 161, 886, 1139, 2136, 1362, 637, 1046, 2106, 715, 2408, 1960, 2420]. Characterizations [541]. characterize [1141]. Characterizing [1079, 1198, 689, 3132, 2241, 2226, 2060, 2126, 3191, 1725, 3296, 1192, 1968, 3431, 2633, 579, 1698, 3133, 1978, 1959, 1121, 1884, 1886, 1427, 2145, 2694, 3323, 2706, 2134, 1831, 816, 2252]. Charging [68, 3147, 2189, 2746, 3165, 3184, 3384, 67, 2930]. charts [114]. Chasm [3153]. Cheap [2522]. cheat [1193]. Cheaters [2040]. Cheating [3113]. Checking [3221, 1369, 1750, 1370, 1639, 1747]. checkpointing [1512, 862, 1946]. chip [1766, 2036, 1932, 1690, 1940]. chip-multiprocessors [1940]. Chipotle [3143]. Chips [2613, 2720]. Choice [2570, 3078, 2397, 71, 29, 2485]. Choice-based [2570]. Choices [2636, 2548, 2569, 3325, 2494, 2025, 2854]. CHOKe [1256]. choose [1335]. Chord [1341]. chunk [1788]. churn [2408]. circuit [193, 778, 586]. circuit-switched [778]. Circuits [2593]. Cisco [1451]. Cities [2660]. Ck [320]. claim [315]. Class [2719, 2857, 2787, 1683, 2476, 2096, 828, 2531, 1730, 856, 731, 2236, 2281, 1059, 2230, 201, 2356, 341, 923, 2304, 809, 296, 600, 2099, 761, 1267, 515, 2239, 202, 1894, 148, 3197]. Class-of-service [1683]. class-oriented [1730]. Classes [3457, 3453, 3381, 589, 384, 1171, 436, 206]. Classification [3126, 3414, 2963, 2957, 1688, 1382, 1567, 1585, 2259, 553, 1859, 3110, 1758, 1380, 711, 402, 1289, 2338]. classifications [1495, 1537]. classifiers [1382, 1492]. Classifying [2151, 1395, 2878, 2399]. Classroom [3210]. Clean [2822]. click [2320]. Clicks [2602, 2320]. Client [3352, 1201, 1195, 1486, 761]. client-server [761]. Clients [3264, 1155, 986]. Clock [1693, 1906, 989, 1611, 670, 1653, 2405]. Clones [2739]. Close [2837, 463]. Closed [683, 2809, 2730, 3026, 2540, 1646, 398, 1864, 1889, 536, 32, 516, 1165, 544, 348, 979, 881, 148]. Closed-loop [683]. Closer [3425, 839]. Cloud [3066, 2692, 2989, 2554, 2507, 3377, 3411, 2581, 3134, 3032, 3238, 2574, 2580, 2557, 2560, 2467, 2561, 2714, 2556, 2965, 2834, 3205, 1984, 2725, 3255, 2462, 3266, 3036, 2722, 2536, 2699, 2878, 2466, 2558, 2726, 2606, 2657, 3283, 2331, 2217, 2335, 2186, 2773, 2409, 1976, 2216, 2210, 2396, 2223, 2168, 1831, 2421, 2147, 2113, 1967, 2427, 3256, 2229, 2352, 2215, 2366, 2146, 2201, 2143, 2212, 2211]. Cloud-of-Clouds [2561]. CloudBridge [1988]. Clouds [3419, 3286, 2561, 3278, 2470, 2511, 2296]. Cluster [1040, 711, 2539, 675, 1200, 989, 890, 1108, 2389]. Cluster-based [675, 1040]. Clustered [2128, 1091]. Clustering [1571, 2598, 2512, 401, 2006, 2370, 2349]. Clusters [2460, 2562, 2900, 973, 2372, 1696, 2905, 1831, 2172, 2143, 186, 2035, 988]. CME [3215]. CMG [87]. CMP [1827, 1547]. CMPs [1808, 1675, 1876, 1977]. Co [1702, 3251, 2519, 1625, 2228]. co-allocation [1625]. Co-designing [1702]. co-generation [2228]. Co-Location [2519]. Co-Simulation [3251]. Coalition [2666, 3199, 2573]. CoBF [3330]. Code [121, 2184, 327, 1850]. Coded [3276, 3259, 2415, 798, 1907]. CodeMRI [1708]. codes [1688, 2383, 2111, 2376]. CODESTAR [485]. coding [1359, 2007, 1652]. coefficient [1348]. Coflows [2705]. coherence [814, 813]. coherent [565]. coincident [2256]. collaboration [1388]. Collaborative [2579, 2564, 2604, 1039, 1693]. Collaboratively [2966]. Collapse [3388, 3073, 3028, 2913, 200]. Collected [3258]. Collecting [2348]. Collection [2976, 671, 2546, 1287, 37, 1193, 1335, 1715, 166, 2418, 2239, 2334]. collective [1872]. collision [3058, 298, 1902, 1674]. collusions [2162]. Colocation [2459, 2559]. Coloring [1854, 3382]. columns [2349]. Combination [89]. Combinatorial [3117, 3263]. Combinatorics [3037]. Combined [1686, 176, 271, 1675]. Combining [2493, 1245, 1333, 684]. Come [3326]. comes [1638]. command [355]. Comments [30, 47]. Commerce [2639, 1363, 1077, 958, 1027, 1364, 2877]. commercial [1588, 565, 674, 1174, 1238]. Commitment [2656]. Committee [3368, 3356]. commodity [2388, 1853]. common [747, 3209]. Commons [2451, 2321]. communicating [157]. Communication [2681, 1400, 2844, 640, 596, 646, 3149, 2449, 663, 3023, 2093, 321, 1254, 2050, 2178, 821, 1756, 249, 784, 851, 1698, 1160, 299, 305, 1111, 833, 483, 2418, 1404, 220, 2175, 1202, 915]. Communication-Aware [3023]. Communications [2946, 2944, 175, 2943, 10, 248, 2429, 396, 693, 612, 2407]. Community [2779, 2632, 3262, 3202, 3371]. CoMoM [1730]. Compact [2501, 843]. compaction [552, 532]. Companies [2674]. Comparative [652, 1581, 3298, 2386, 249, 22, 47]. Comparing [1525, 2758, 1036, 1992, 2135, 229, 5, 2154]. Comparison [505, 1607, 2161, 2394, 2050, 1511, 247, 395, 1826, 106, 458, 783, 1113, 1563, 1168, 435, 1472, 26, 771, 23, 775, 82, 2177, 682, 24, 1272, 727, 2163]. Comparisons [2512]. Compatibilities [3336]. Compatibility [3139, 386, 1848]. compensating [1653]. compensation [295]. competing [2676]. Competition [2473, 2667, 2740, 3369, 2331, 2949, 3346, 3176]. Competitions [2611]. Competitive [3055, 3386, 3182, 2665, 2983, 1543, 2982, 3147, 2321, 593, 2421]. competitiveness [2250, 2136, 1681]. compilation [877]. compile [668]. compile-time [668]. Compiler [2847]. Compiler-Inserted [2847]. Complementarities [2573]. complementary [792, 2220]. Complete [660, 1335, 3388, 778, 186, 944]. Completely [2751]. Completion [2705, 757, 862, 1527, 1459]. Complex [2520, 3012, 2094, 1830, 753, 2208, 2215, 1840, 1709]. Complexity [2687, 3076, 325, 2815, 2936, 2861, 2860, 2121, 2142, 247, 1203, 394, 2186, 392, 324, 2403]. component [1365]. component-based [1365]. components [2074, 397]. Composed [3217]. Composite [648, 892, 188, 742]. Compositional [2095]. Comprehensive [3350]. compressed [2076, 980]. Compression [1927, 2620, 294, 1300, 996, 1970, 752, 1764]. Compression-Aware [2620]. compressive [1982]. Compromise [3322, 2026]. Computable [2495]. Computation [3244, 2786, 2745, 889, 1355, 2000, 2216, 2041, 1599, 2414]. Computational [178, 644, 821, 2254, 701, 1247, 347, 1248]. Computationally [3015]. computations [945, 931, 540, 1104, 155, 2039]. Compute [3216, 2690, 3054, 1437, 1060, 1831]. Compute-Sync-Merge [2690]. Computer [641, 2, 138, 9, 595, 597, 606, 642, 646, 751, 131, 745, 192, 3454, 264, 602, 343, 735, 167, 44, 95, 379, 49, 616, 689, 409, 56, 31, 42, 50, 1277, 1740, 91, 210, 63, 40, 222, 265, 302, 159, 61, 43, 1064, 37, 397, 64, 390, 3209, 430, 250, 11, 272, 98, 136, 279, 122, 404, 149, 45, 66, 134, 284, 17, 193, 678, 88, 238, 72, 431, 577, 67, 526, 281, 348, 448, 41, 62, 69, 96, 681, 560, 177, 703, 704, 82, 147, 53, 12, 474, 133, 137, 1857, 344, 19, 515, 388, 97, 944, 812]. computer [513, 1751, 10, 640, 594, 596, 643, 654, 663]. computers [216, 180, 463, 396, 851, 756, 817, 717, 101, 726, 24, 263]. Computing [3066, 2946, 3067, 2767, 1240, 2554, 104, 3134, 2555, 2454, 2977, 2574, 2905, 1082, 2691, 2648, 2765, 2171, 3426, 3205, 3255, 2979, 3251, 3266, 2553, 2657, 1512, 1795, 278, 969, 1015, 1738, 1438, 1252, 839, 1815, 65, 1965, 678, 1201, 2049, 615, 2100, 2216, 156, 261, 339, 674, 1967, 2427, 2352, 2215, 1246, 1697, 2146, 2201, 270, 2387, 1243, 346, 2212]. concatenation [1134]. Concave [2357]. Concentrated [3215]. Concentrator [295]. concept [7, 934]. Concurrency [2748, 2780, 3191, 472, 444, 418, 708, 417, 203]. Concurrent [3086, 2516, 688, 436, 1360, 676, 628, 255, 727]. concurrent-read [628]. Condition [2588, 2305]. conditional [1395]. Conditions [2921, 3122, 2968, 1167, 1127]. conference [2015, 87, 2899]. Conferencing [2777]. Confidence [133, 1487, 1018, 3371]. Confident [1654]. Configuration [103, 224, 2608, 225, 2207, 223, 1795, 1023, 1364, 190]. Configurations [2842, 1488, 1532, 1557]. configuring [797, 941]. Confirmation [3102, 3103, 3104]. conflict [482, 1414, 2224]. Conflicts [2771]. Confluence [3376]. Congested [2727]. Congestion [2142, 2571, 3192, 3171, 2939, 2991, 2331, 2189, 2194, 1449, 1308, 1307, 1591, 2108, 1848, 2949, 1942, 1085, 873, 2391, 1053, 556, 1054, 1849, 1204, 2768, 995, 2112, 943, 744]. congestion-dependent [556]. conjectures [483]. Connected [3155]. Connection [2800, 1056, 1510, 1337, 2403, 566]. Connection-Level [2800]. connection-rate [1056]. Connections [2529, 2001, 1053, 1152]. Connectivity [2570, 1845, 1663, 1491, 2389]. Connie [608, 700]. Conquering [2361]. conscious [2383]. Consensus [3201, 3088, 3235, 3035, 2948, 56]. Consensus-based [3035]. Consensus-forming [3088]. Consequences [2985]. Conservatism [2445]. conserve [1467]. conserving [2360]. Considerations [455, 65, 78, 374]. considered [1471]. considering [740]. Consistency [1074, 429]. Consistent [3071, 1083]. Consolidation [3229, 1988, 2035]. Constant [3318, 3317, 3404, 1869, 2431, 1572]. Constant-stepsize [3404]. Constrained [2568, 3310, 3429, 3118, 3174, 3333, 3078, 3023, 3157, 1101, 1240, 341, 2396, 817]. Constraint [3146, 3181, 3145, 3195, 438]. constrainted [411]. Constraints [3151, 2687, 2692, 3440, 3182, 2983, 3181, 3139, 2439, 2653, 3403, 2638, 3046, 3138, 3048, 1842, 2082, 1292, 1996, 1678, 1070, 2319, 2267, 34, 2227, 385]. Constructed [2791]. Construction [2641, 57, 1555, 1522]. consumer [2324]. consumers [226]. Consumption [2579, 3123, 2618, 2734, 2455, 2458, 1610, 1684, 1064, 2184, 1629, 1875, 1918, 1199, 2245, 2252, 3374]. Contagion [2447]. containers [2144]. Contemporary [2848]. Content [2713, 2914, 1353, 2149, 2455, 2453, 2565, 2961, 2594, 3213, 2498, 3425, 1799, 3142, 2326, 1024, 736, 1805, 1966, 1772, 2350, 1813, 2380, 1051, 2323, 2398, 2242, 2243, 1634, 1960, 2327, 988]. Content-aware [2149]. contention [378, 1003, 1177, 937, 500, 1763, 141]. contents [975]. Context [1208, 2365, 1176]. Context-aware [1208, 1176]. context-sensitive [2365]. Continual [3159]. Continuous [3423, 3452, 1715, 2584, 3027, 990, 1613, 2126, 737, 797, 999, 1978, 1992, 190, 944, 971]. Continuous-Time [3452, 1613, 1978]. Continuum [3067]. Contractive [3404]. Contracts [3132, 2576, 1905]. Contractual [3142]. Contrasting [2621]. contribution [1265]. Control [3017, 2993, 2809, 2986, 1400, 2826, 2646, 3316, 352, 121, 2555, 3284, 1360, 2962, 3118, 2740, 3186, 3401, 3449, 2551, 3095, 3065, 3345, 539, 2762, 3451, 2736, 3365, 2694, 2712, 2792, 2991, 3281, 2771, 2096, 505, 683, 2003, 2189, 2194, 1163, 49, 225, 328, 409, 2142, 1645, 223, 1308, 317, 1307, 2108, 2310, 2307, 2281, 800, 1518, 1848, 1912, 1778, 351, 1821, 472, 1942, 122, 1972, 1552, 1085, 286, 784, 1196, 883, 1752, 873, 1674, 556, 1116, 2010, 1573, 318, 483, 572, 300, 444, 418, 353, 708, 585, 995, 1152, 943, 202, 744, 727, 1533, 513, 417, 3171]. Control-Based [3345]. control-plane [1573]. control-theoretic [572]. controlled [288, 913]. Controller [3285, 1821, 416]. Controlling [2773, 1332, 1003, 1125]. controls [352]. conventions [25]. Convergence [3452, 2844, 2572, 2605, 2649, 2861, 2830, 3319, 3404, 1943, 1033, 1175]. Converging [3401]. Converse [2591]. Conversion [3440, 3409, 2110]. convertible [1879]. Convex [3151, 2484, 2890, 3018, 1095, 3181, 3333, 2889, 2904, 3378, 2982, 3242, 2825, 2261, 2356, 3049, 1565, 2728, 580, 590]. conveyor [449]. Convolutional [3054, 376]. Convolutions [3189]. Conway [644, 701]. Cooling [2522, 2115, 2141]. Cooperation [2473, 3210, 1805]. Cooperative [3297, 2687, 1772, 1443, 1211, 2424, 916]. Coordinated [1154, 2438, 1876, 844, 2559, 3330]. Coordination [2325, 2575, 1069, 539]. copies [205]. Coping [1309]. coprocessors [877]. Copy [2035, 1436]. Core [2704, 2615, 2456, 2573, 2757, 2394, 2284, 902, 1963, 1924, 1756, 1644, 21, 2183, 1939]. Cornell [868]. Corrected [2289].

Correcting [2312]. Correction [547, 2384]. correctness [108]. Correlated [3207, 3014, 3263, 984, 671, 1861, 2282, 1437, 1604, 1102, 2389]. Correlation [3229, 1779]. Correlation-Aware [3229]. correlations [1271, 3197]. coscheduling [1184]. COSEM [3261]. Cosmos [3221]. Cost [2836, 2988, 2731, 3100, 2646, 2737, 3352, 2812, 2543, 2563, 1351, 3295, 3065, 3378, 51, 2776, 3266, 3036, 2466, 2519, 2152, 16, 2415, 2191, 363, 1976, 2356, 1875, 2206, 78, 1804, 1952]. Cost-Effective [3352]. Cost-Efficient [2836, 3036, 3295]. Cost-Identifiability [2776]. Cost-optimal [51]. cost-performance [363]. Costly [2593]. Costs [3409, 3095, 2556, 2452, 3084, 3019, 3242, 2470, 1407, 2307, 2391]. Count [314]. Counter [1655, 1493]. Counterfeit [3131]. Countermeasures [2583]. counters [1796]. counting [317, 1560, 1594, 326, 1707, 315, 312]. Coupled [2830, 31, 1399]. Coupling [3404]. Coupon [2515, 1376, 1911]. Cournot [2740]. Coverage [2634, 3033, 2762, 2683]. Covert [3167, 3187]. COX [357]. COX-distribution [357]. CP [74, 2323]. CP-V [74]. CPS [2770]. CPS-based [2770]. CPU [1607, 43, 1675, 2905, 3001]. CPU/GPU [2905]. CPU/link [1675]. CPU2000 [1581, 1192]. CPU2006 [1581]. CPUs [1582]. Cramér [1782]. Craps [3113]. Crawler [1073]. Crawler-friendly [1073]. Crawls [2601]. Cray [637, 2177]. CRC [658, 746]. credible [137]. Credit [3200]. Criteria [159, 212]. criterion [43]. Critical [2770, 2583, 2894, 2763, 2898, 3040, 3039, 2179, 2892, 1617, 564, 254]. Critical-Infrastructure-Targeted [2583]. Criticality [2615, 2360, 2020, 2209]. Critically [3140]. Cross [1518, 3412, 1285, 2852, 3437, 3436, 68, 1426, 1404]. Cross-architecture [1285]. Cross-Layer [3437, 3436, 1518, 2852, 1404]. Cross-Path [3412]. cross-traffic [1426]. Crossbar [2689, 1419]. Crowd [2753, 2486]. Crowd-Sourced [2753]. Crowd-Sourcing [2486]. crowded [2225]. Crowdsensing [2684]. Crowdsourcing [2485, 2230, 2428]. Crypto [3100]. Crypto-Hotwire [3100]. Cryptocurrency [2810, 3131, 2954, 3323, 2949]. Cryptocurrency-themed [3323]. Crystal [2976]. CS [3457]. CSI [2847]. CSMA [2528, 407, 352, 2262, 1883, 1992, 2530, 2622, 1846, 2927]. CSMA/CA [2530, 2622]. CSMA/CD [407]. CTIP [905]. CUDA [1926]. Cultivation [3353]. cumulative [1033, 819]. curation [2326, 2350]. curcuit [425]. Curing [2812]. Current [1149, 2580, 2585, 2422, 125]. Curtain [3126]. Curve [2534, 42, 1405, 1546, 998, 744]. custom [1153]. Customer [2703, 808, 2295, 600]. customers [671, 436, 385]. cutoff [318]. cutting [1097]. CVaR [3244]. CWS [1861]. Cyber [2894, 2898, 2806, 2497, 2766, 2893, 1905, 2339]. cyber-insurance [1905]. Cyber-Physical [2766, 2893, 2339]. Cyber-Security [2806]. cyberinfrastructure [2159]. CyberInsurance [2202]. cybersecurity [2202]. cyberspace [2166]. Cycle [439, 217]. Cyclic [3383, 2076, 453, 32, 673]. cyclic-service [453]. Cypress [1930].

D [644, 701, 2934, 2854, 3140, 3325, 2928, 3420, 1266, 2123, 3331, 2849, 936, 3277, 790, 2706, 3152]. D-Band [3277]. D-factor [2123]. D-hops [3152]. D-redundant [1266]. D.P. [93]. D2D [2429]. DAG [2950]. DAG-based [2950]. DAGsim [2950]. Daisy [906]. damage [1584]. dampening [2073]. Dancing [2623, 812]. Dandelion [2810, 2723]. Dangling [3321]. DARC [1668]. DASD [191, 812]. Data [2517, 2708, 3141, 2461, 3261, 3212, 2681, 2537, 3423, 2330, 2719, 280, 319, 2504, 2702, 2891, 3280, 2721, 3000, 3146, 3421, 3060, 1568, 100, 2761, 2700, 2835, 2455, 3429, 1441, 1301, 2973, 3258, 3284, 1478, 2754, 2794, 3158, 3282, 2256, 3361, 2798, 3300, 2803, 802, 2561, 2377, 2514, 3397, 2255, 2617, 232, 3407, 2624, 561, 2463, 3363, 3256, 2908, 2979, 2978, 3137, 2540, 2607, 2800, 2755, 3303, 3099, 2501, 2458, 2519, 2476, 1411, 2733, 2782, 945, 2278, 2345, 505, 2130, 1832, 338, 736, 1390, 760, 1254, 1319, 2426, 2034, 2415, 2050, 529, 1154, 407, 2044, 1368, 205, 1950, 627, 197, 2383, 2148, 2114, 1490]. data [37, 591, 2019, 737, 460, 3228, 1392, 2413, 1545, 1458, 134, 1965, 1381, 1771, 1968, 1565, 1810, 2115, 2355, 1774, 2337, 2283, 2223, 1835, 2185, 635, 2922, 96, 1715, 1879, 166, 690, 1229, 1333, 1036, 2342, 89, 1984, 1143, 318, 1764, 2339, 2418, 1006, 119, 2361, 2338, 1701, 51, 733, 1952, 585, 2334, 2145, 2313, 2265, 2317, 2213, 2341, 1960, 2212, 155, 1459, 1939, 1477, 92]. Data-flow [3300]. Data-intensive [3256, 945, 505, 1965]. data-locality [1390]. Data-Rich [3363]. Database [3031, 750, 657, 694, 574, 457, 458, 415, 832, 524, 378, 442, 508, 670, 204, 255, 444, 708, 403]. databases [575, 2220]. Datacenter [3279, 3090, 2500, 3192, 2705, 3281, 2833, 2831, 2253, 3233, 2252, 2315, 3171]. Datacenter-scale [2831]. Datacenters [3379, 2635, 2658, 2559, 2509, 1983, 2036, 1814, 2402, 2116, 2272]. dataflow [505, 150, 151]. Datasets [3301, 2045]. day [2670]. day-ahead [2670]. DBs [2280]. DC [2795]. DCC [3066]. DCN [2846]. DCTCP [1943]. DDoS [1509, 3167, 3187]. De-ossifying [1975]. Deadline [1418, 2698, 2929, 3403, 2699, 2558, 2203, 3165, 1268, 1697]. Deadline-aware [2558]. Deadline-Sensitive [2699]. Deadlines [3206, 2880, 600]. death [2072]. debate [1683]. debugging [234, 734, 335, 510, 876, 1946]. decay [804]. decay-usage [804]. December [706]. Decentralised [3173, 2887]. Decentralized [3269, 195, 435, 1378, 1677]. deciding [1101].

Decision

[3045, 3421, 3345, 2526, 3359, 3006, 2830, 2877, 1371, 1839, 1382, 2081, 1066, 1031]. Decision-Making [3045, 3421, 3359]. decision-support [1066]. decisions [329, 2335]. declarations [313]. DECnet [493, 373]. decomposability [186]. Decomposable [2751]. Decomposition [644, 1904, 2744, 701, 3247, 187, 1756, 2067, 503]. Deconstructing [2734]. decoupling [1900]. Decreases [2538]. DECstation [738]. DECwindows [579]. deduction [280, 319]. deduplicated [1907]. Deduplication [2561, 2922]. Deduplication-Assisted [2561]. Deep [3115, 3029, 1871, 3191, 3296, 3295, 2961, 2958, 3391, 3407, 3277, 2846]. DeepLabCut [3354]. DeepSec [2958]. defects [239]. Defending [3035]. Defense [2894]. deferrable [2310]. Deferred [2450]. Deficiencies [64]. define [609]. Defined [2681, 2505]. Defining [72]. definition [280, 319]. deflection [743]. Defragmenting [2229]. Degradable [404, 271, 346]. Degradation [2517, 2972, 2312]. Degree [2918, 2855, 1913]. Degree-Greedy [2918]. Delay [2816, 1991, 2586, 2686, 1771, 2605, 3010, 2654, 3065, 482, 1883, 1058, 2101, 1499, 2800, 2911, 2861, 3127, 2538, 2860, 2855, 2913, 2968, 1293, 1592, 461, 735, 2142, 1649, 1986, 1253, 1432, 1678, 2262, 1454, 1196, 1351, 2007, 679, 1551, 1804, 374, 2788, 2401]. Delay-Cost [3065]. delay-friendliness [1649]. Delay-Optimal [2654, 2860, 1499]. Delayed [2799, 2802, 3378, 2536, 2527, 683, 1019, 2302, 385]. Delays [2233, 2844, 3333, 2951, 2498, 718, 343, 1338, 1463, 470, 1450, 467, 1100, 390, 272, 1719, 486, 1160, 1887]. deletion [2376]. delivered [241]. Delivering [2257, 1713]. Delivery [2713, 2673, 2565, 2594, 2498, 1833, 1189, 1955, 2398, 1799, 1634]. Delta [2933]. DeltaTree [2516]. Demand [2814, 2537, 2656, 2752, 2703, 2457, 2753, 2321, 2515, 3192, 2459, 2813, 2559, 2657, 2029, 857, 1309, 1339, 1937, 796, 2256, 2355, 2283, 1284, 1402, 2380, 2616, 2113, 1670, 2229, 897, 2308, 2073, 812]. Demand-Aware [3192]. demand-based [2229]. Demand-insensitive [2321]. demand-responsive [2283]. Demands [3188, 2595, 2540, 38, 2880, 1574]. Democratizing [3394]. demonstrating [575]. Demystify [3003]. Demystifying [3012, 3086, 2463]. Denial [1379, 2112]. Denial-of-service [1379]. denied [1579]. Denis [746]. Dense [2528, 2785, 2704, 3330]. Densification [1662]. density [1194]. department [789]. Departure [3163, 3184, 1439]. Departures [3307, 1595]. dependability [1744, 1511, 1743, 623, 722]. dependable [1458]. dependence [876]. Dependencies [3383]. Dependency [3299]. Dependent [2809, 2643, 3026, 3196, 2969, 1807, 792, 333, 404, 368, 809, 556, 970, 1792, 1251, 369, 774, 3135]. deployability [1053]. deployed [1035, 2400]. deploying [2216]. Deployment [3072, 3277, 1648]. deployments [2217]. Depth [3294, 2850, 677]. derivation [498, 2336]. derivations [85]. Deriving [1782]. Descent [2845, 2890, 3007, 3046]. Description [907]. descriptor [2014]. Design [641, 2731, 2769, 3290, 891, 2752, 2572, 519, 2931, 2499, 3358, 2916, 2580, 2720, 3351, 508, 2448, 2439, 113, 848, 3050, 456, 995, 2639, 1493, 268, 1280, 649, 1382, 1738, 365, 2019, 898, 425, 1932, 1658, 2232, 1923, 1332, 1704, 1303, 1561, 2180, 1685, 1690, 2238, 238, 126, 1402, 235, 246, 1812, 635, 2248, 112, 1652, 153, 1404, 203, 220, 1935, 2381, 194, 142, 2428, 2327, 114, 2354]. Design-Induced [2720]. Designing [252, 2222, 2841, 2558, 2860, 1702, 1905]. designs [1843, 2152, 1722, 472, 1561, 1870, 228, 109]. desktop [1035, 993]. despite [1647].

Desrochers [658]. destruction [1584].

Detailed [2851]. details

[957, 2867, 2868, 2869, 2870, 2871, 2872, 2807, 2873, 2874, 2875]. Detecting [2517, 2258, 3132, 2510, 2633, 3300, 2959, 1054, 1857, 1885, 3321, 616, 633, 865]. Detection [3419, 2791, 3219, 3112, 2775, 2779, 3258, 3428, 2958, 2508, 3230, 3303, 3005, 3382, 3231, 1509, 1997, 1213, 473, 1684, 1963, 996, 2231, 1085, 1810, 1139, 1554, 1333, 2347, 961, 1964, 1862]. detector [2117]. Detectors [3167, 1806]. determinate [946]. determination [294, 8, 89]. Determining [1469, 1138, 3040, 306]. Deterministic [2731, 949, 2513, 3315, 337, 806, 1893, 2018]. Develop [41]. developing [370]. Development [703, 137, 129, 218, 243, 118, 485, 130, 1561, 510, 241]. developmental [241]. deviation [1094, 933, 1240]. Deviations [2748, 2780, 3376, 2422]. Device [3400, 2334, 181, 1331, 2387]. Devices [3123, 2718, 3069, 3296, 2006, 1037, 1825, 1386, 855, 269, 1959]. devil [957]. DHCP [2408]. Diagnosing [1550, 1591, 525, 783]. Diagnosis [2831, 1816, 2124]. Diagnostic [3298]. diagrams [1613]. DIAMOND [181]. did [3123, 1312]. Diego [637]. diesel [2309]. Different [3301, 984, 247, 317, 2047, 478, 1085, 1508, 1803]. Differential [2448, 1669, 1898, 234]. Differentials [2503]. Differentiated [3283, 1151, 991, 2097, 1075]. Differentiation [2322, 1302, 2165, 1385, 1034]. DiffForward [3283]. diffserv [1196]. diffserv-enabled [1196]. Diffusing [2631, 2863]. Diffusion [2441, 2303, 2021, 2926, 1894, 2080, 32, 1507, 1605, 2053, 2297, 2370, 2360, 1521]. Diffusions [2811]. digital [781, 858, 181]. dilemma [2953]. Dilemmas [2557]. Dimension [3347, 602, 2858]. Dimensional [3027, 3007, 2993, 3388, 209, 1286, 1955, 2068, 1859, 2380, 2088, 2885, 2263, 2913]. dimensionally [1343]. Diophantine [1643]. direct [1338]. direct-indirect [1338].

Directed [3013, 1675]. direction [2345, 1531, 1891]. direction-aware [1891]. directionality [1687]. directions [1566, 1246]. Disaggregated [3280, 3278]. Disaggregation [3379, 440, 2915]. Disaster [2896, 3041, 2410, 3043]. Disasters [2765]. discarded [1773]. Discipline [3387, 1994, 451]. disciplines [1293, 1270, 17, 1294, 572, 573, 375]. DISCO [2963]. Discount [2639, 2877]. Discovery [2526, 1958, 1409, 1854]. Discrete [3312, 3337, 337, 1100, 1239, 2063, 1640, 462, 2155, 588, 881, 2336]. discrete-event [1100, 1239]. discrete-state [2155]. discrete-time [881]. Discretize [2963]. discretizing [190]. Discriminant [2259]. discriminatory [1994, 544, 1792, 1358]. Disk [465, 1665, 826, 2492, 1224, 946, 1462, 1159, 1096, 1185, 1463, 2291, 1570, 631, 810, 630, 866, 1503, 797, 415, 416, 256, 721, 2180, 466, 6, 355, 906, 903, 147, 1803, 1822, 1046, 917, 930, 2244, 1210, 1460, 373, 523, 812, 777, 811, 1850, 1464, 1219]. Disk-Based [2492, 2180]. disk-level [1803]. Disk-tape [826]. DiskAccel [2492]. disks [1096, 938, 977, 714, 555, 902, 893]. Disparities [3153]. Dispatchers [3079]. Dispatching [2024, 3020, 3063, 3174, 3344, 3390, 1222, 2005, 29, 2103]. Disruptions [3047, 1550]. dissatisfaction [2258]. Dissection [2013]. Dissemination [2491, 1386, 2377, 1508, 1523]. Dissertation [3356]. DiST [1249]. Distance [2480, 1223, 2259, 537, 682, 469, 2433, 1682]. distance-vector [682]. distances [861]. Distancing [3179]. distinct [206]. Distinguishing [316]. Distributed [2326, 3066, 2946, 2687, 2564, 2678, 2845, 2752, 3030, 2760, 2844, 2659, 3149, 2586, 2554, 2522, 2964, 2775, 2987, 963, 1866, 2724, 2994, 1845, 2563, 3056, 2608, 3351, 3308, 3271, 2410, 3287, 2745, 2560, 2688, 2929, 2147, 1881, 2624, 3339, 3434, 1572, 300, 3007, 3035, 2842, 2717, 601, 1888, 3190, 2471, 505, 1780, 2189, 935, 1163, 2000, 143, 1869, 617, 1035, 1505, 275, 480, 900, 1818, 940, 1359, 1906, 1781, 841, 1617, 1738, 1490, 397, 782, 1361, 518, 517, 664, 614, 1097, 2316, 578, 1640, 1815, 65, 2973, 2754, 2794, 2049, 1810, 2216, 289, 2798, 489, 817, 1513, 579, 2172]. distributed [508, 2880, 936, 2885, 755, 339, 413, 2253, 1689, 479, 675, 1356, 2195, 726, 2235, 892, 1184, 1466, 1947, 803, 1202, 270, 290, 375, 509, 2272, 2950, 2343, 1865, 2018, 1031, 3238]. Distributing [2242, 1074]. Distribution [2937, 2938, 1018, 2455, 3265, 3375, 2561, 3088, 3043, 2787, 2755, 3215, 3335, 2190, 357, 830, 1805, 205, 1353, 1025, 2083, 807, 2261, 2001, 1580, 1597, 2063, 672, 2109, 1772, 2066, 450, 1301, 1381, 88, 2086, 1774, 2192, 2434, 320, 1911, 1527, 2033, 2224]. Distribution-free [1018]. Distributions [3207, 2998, 3452, 2064, 2057, 2058, 2077, 1170, 1863, 1218, 1482, 2090, 1695, 410, 1140, 2060, 2750, 1240, 840, 1187, 2299, 2065, 1605, 1165, 201, 2199, 543, 2155, 1229, 2055, 414, 1668, 1521, 1894, 819]. Dive [3391]. divergence [56, 1964]. Diverse [2827]. Diversity [2866, 2865, 524, 3393]. diving [1871]. Divisible [2925]. DL [3398]. DL-based [3398]. DLMS [3261]. DLMS/COSEM [3261]. DNA [1355]. DNA-based [1355]. DNNs [3355]. DNS [1423]. Do [3017, 1819, 6, 1971]. Doctoral [3356]. documentation [242]. documents [2109]. Does [2122, 1580, 2832, 3162, 3427, 1610, 2266]. DOG [1241]. DoKnowMe [2582]. DOLL [3238]. Domain [2582, 1314, 2133, 1686, 1987, 1692, 2649]. Domains [3033, 3321, 383]. Domestic [2659]. Dominant [3054]. donor [1445]. Don’t [3408, 2129, 2614]. DoS-resistant [3201]. double [2085]. down [2129, 2123, 1627, 1635, 2033]. Downgrading [2647]. Download [2708, 2345, 1339]. downtime [2193]. Drain [2481, 2850, 2226]. DRAM [2613, 2718, 891, 2851, 3012, 1972, 2386, 2720, 1707, 1673, 1796, 1768]. DRAM-based [891]. Draw [591]. Drift [2469, 2902, 2803, 1653]. Drive [2627, 1185, 854, 1629, 1046, 811]. Driven [2976, 3219, 2454, 2962, 2585, 2669, 3251, 2075, 2412, 714, 1861, 2662, 37, 725, 2348, 1968, 2582, 489, 1956, 35, 2725, 46, 1759, 530, 550, 2846]. Drives [2954, 1570, 1767, 1436, 1822, 930, 2239, 777]. Driving [3044, 3135]. Drop [1530, 2557]. DSL [1837]. DTN [1920]. DTNs [2676]. Dual [2345, 2940, 3046, 3317, 535]. dual-bus [535]. duality [1124]. due [235, 246, 200]. DUMP [215]. Duo [3281]. duopoly [2169]. Duplex [2478]. Durability [1689]. durable [2332]. Duration [2611, 1088]. during [2225, 509]. DV [591]. DV-Draw [591]. DVFS [3206, 1834]. Dynamic [2980, 3090, 3193, 1349, 2769, 2094, 568, 2295, 3149, 2821, 2449, 2581, 1976, 3047, 3032, 2786, 3237, 2840, 2575, 2623, 3306, 1447, 2648, 2965, 3034, 3130, 3205, 3293, 1764, 1105, 2978, 3144, 1202, 1634, 583, 3036, 3317, 2992, 2711, 2098, 340, 2550, 990, 1868, 1203, 2307, 2281, 2372, 2431, 1696, 840, 1050, 669, 1264, 1196, 739, 1107, 2009, 676, 637, 1881, 1978, 1777, 813, 2352, 1291, 1184, 1272, 1668, 1651, 2145, 1682, 2224, 988]. Dynamically [3266]. Dynamics [2566, 2923, 2732, 3401, 2625, 1849, 1313, 2518, 1314, 2092, 259, 1848, 1856, 1847, 1051, 1775, 1959, 2375, 1528, 2072, 1913]. dynamism [2390].

e-bike [3160]. e-business [1145, 1351, 1362]. E-commerce [2877, 1363, 1077, 1027, 1364, 2639]. Each [3299]. Early [2849, 1509, 868, 311, 905]. early-bird [311]. Ease [582]. ECF [2707]. ECI [2836]. ECI-Cache [2836]. Eclipse [1745]. Economic [2965, 2903, 2559, 2320, 2218]. economically [294]. Economics [2735, 2331, 1905]. economies [278]. economies-of-scale [278]. Economy [2672]. ecosystem [2949]. Ed [646, 657]. EDF [856, 3165]. Edge [2500, 2914, 3107, 3296, 2691, 3268, 3256, 3148, 3177, 1181, 2949, 2910]. edge-based [1181]. Edge-cloud [3256]. Edited [750, 751]. Edition [642, 751]. Editor [1506, 587, 1448, 1457, 1342, 1433]. editorial [1243]. Editors [956]. EDP [1004]. EDP/ERP [1004]. EDP/ERP-oriented [1004]. eDRAM [2121]. Edwin [751]. EELRU [981]. Effect [313, 2533, 2569, 3200, 2401, 1722, 405, 317, 1597, 1595, 484, 1526, 2027]. Effective [2739, 841, 3352, 2691, 2984, 294, 331, 1300, 111, 1018, 2008, 981]. Effectiveness [734, 129, 210]. Effects [146, 2621, 147, 3185, 2649, 2573, 613, 1127, 1989, 1484, 269, 1887, 2360]. efficacy [2148, 2386]. Efficiency [2944, 1609, 2923, 2583, 3411, 2704, 2740, 2620, 3329, 2756, 1999, 1922, 1630, 2419, 1972, 703, 1493]. Efficient [2836, 3215, 2130, 3261, 3419, 383, 861, 3201, 2487, 2628, 3015, 3117, 2014, 571, 731, 1409, 3280, 3285, 1809, 2987, 3273, 3453, 2230, 784, 1752, 1685, 2663, 489, 1586, 1062, 1229, 2196, 2427, 2660, 3270, 3320, 2909, 892, 715, 3264, 2462, 1626, 3036, 550, 2414, 3037, 2925, 2133, 2152, 1480, 1421, 1577, 398, 437, 1844, 551, 2045, 2108, 2419, 1299, 2316, 1183, 1070, 1301, 2104, 3295, 3287, 1109, 1836, 1048, 1116, 1763, 2155, 1143, 2363, 2033, 2181, 2315, 971, 1411, 1344, 3353]. efficiently [1711, 801]. effort [129, 998, 999, 520, 1677]. efforts [1246]. Egalitarian [2430]. EigenTrust [3258]. Eigenvectors [3054]. EJB [1130]. Elastic [2921, 3204, 2504, 2260, 1357, 1147, 1306, 1801]. Elasticity [3411, 2688, 2882]. Electric [2746, 2191, 3147, 2759, 2410, 2313]. electrical [2189]. Electricity [2754, 2794, 2798, 2655, 3043, 2576, 2701, 2519, 2509, 2283, 1918, 2670, 2915, 2354]. electronic [116, 958, 356, 325]. elements [1291]. Elevator [696]. Elevator-type [696]. Ellis [656]. Elon [1844]. Elsevier [646]. elusive [965, 1663]. email [2093]. embedded [1783, 1621]. Embedding [2682, 3300, 3232, 1479, 2128, 2987]. Embra [845]. Emergence [2659, 2279]. Emergency [2946, 2962, 2657]. Emerging [2872]. eMIVA [1617]. EmNet [1802]. empathic [1802]. Empirical [1388, 210, 2185, 2709, 403, 184, 86, 1022, 2050, 138, 280, 319, 545, 767, 2253, 1798, 831]. empiricist [13, 17]. Emulating [1295]. emulation [1716, 877, 1220, 1677, 1709]. Enabled [3261, 1196, 3366]. Enabling [2850, 2922, 3293, 3256, 3148, 1844]. encapsulated [1578]. encoding [1441]. encourage [1265]. Encrypted [3126, 3072, 3109]. Encyclopedia [751]. End [2517, 2698, 3131, 1080, 1631, 3355, 1398, 486, 2502, 3185, 3403, 1975, 1834, 649, 1032, 1421, 1600, 1490, 1469, 1914, 1454, 1351, 1633, 1054, 859, 1938, 1334, 943, 1329, 774]. End-Host [3185, 1914]. end-network [1975]. End-to-End [2517, 2698, 3355, 2502, 3403, 3131, 1631, 1398, 1834, 1421, 1600, 1490, 1469, 1454, 1351, 1054, 859, 1938, 1334, 943, 1329, 774]. Endpoint [3258]. Endurance [2836, 2241]. Energy [3151, 3029, 3261, 2734, 2481, 2850, 2661, 2108, 2455, 2563, 2754, 2794, 2515, 3271, 2620, 3329, 2798, 1836, 1936, 2353, 2756, 2684, 3001, 2660, 1199, 3157, 2116, 2536, 2458, 3353, 2284, 1922, 2034, 1610, 1407, 1676, 2796, 2388, 2266, 1064, 2423, 2186, 2378, 2419, 2184, 2316, 1704, 1815, 2402, 1875, 2043, 1968, 1921, 2228, 1559, 1467, 2144, 2041, 2033, 2245, 1952, 1920, 2315]. Energy-Aware [3157, 1936, 2796, 2186]. energy-efficiency [1922]. Energy-efficient [2108, 3353, 2315]. energy-performance [1559]. Energy-Response [2536].

Energyaware [2464]. enforcement [2257]. Eng [588]. Engagement [2960, 2375]. engine [1667]. Engineering [3249, 2804, 608, 2585, 700, 1686, 229, 1742, 1207, 1237, 933, 1131, 1315, 2149, 872, 1772, 2290, 2015, 1233, 1284, 1402, 2242, 1945, 2317]. engines [2161, 2222, 1125]. English [315]. enhance [139]. Enhanced [3352, 2615, 3417]. enhancements [1741]. Enhancing [3155, 2719, 3360, 3137, 2241]. Enough [2956, 1567, 2203, 1799]. Entanglement [3025, 3077, 3208]. enterprise [2312, 1781, 2150, 871]. Entropic [3424]. Entropy [2641, 386, 450, 1478, 2425, 400]. ENUM [1632]. envelope [2301]. envelopes [3062]. Environment [3294, 3114, 2929, 1010, 379, 225, 760, 1349, 834, 627, 694, 582, 661, 568, 782, 2773, 332, 2150, 2214, 486, 150, 78, 533, 1128, 482, 765, 1653, 1291, 417, 1751, 1281]. environment-induced [1653]. Environmental [1919, 1838]. Environmentally [2454, 1305]. Environments [3032, 2832, 3177, 3251, 3266, 2022, 939, 1512, 1825, 100, 1320, 2042, 1044, 1633, 674, 1184, 1697, 1435]. Epidemic [1672, 2967, 3260, 2551, 2789, 2751, 2508, 3159, 1386, 2119, 2118, 1521]. Epidemic-Like [2551, 2789, 2751]. Epidemics [2589, 2812, 2367, 1752]. epidemiology [3275]. EPT [3419]. EPT-based [3419]. Equalizing [1901]. equation [1900]. equations [697]. Equilibria [2626]. Equilibrium [2572, 2786, 3068, 3142, 1787, 189, 1885]. equipment [1918]. equivalence [2805, 187, 792]. equivalent [188]. Era [2941, 2815, 1786]. Erasure [3276, 3259, 2415, 1907]. Erasure-Coded [3259, 2415, 1907]. Erdos [2591, 3014]. Ergodicity [3235]. Erlang [1489]. Erol [604]. ERP-oriented [1004]. Erratum [3208]. Error [2804, 350, 1193, 2376, 107, 2610, 2384, 405, 1963, 2111, 2386, 1085, 2007, 556, 520, 1886, 1944]. error-estimating [2111]. Error-free [1193]. errors [1570, 1503, 179, 556, 240, 241, 1768]. ES-based [703]. ESS [325]. essential [3455]. establishment [109]. estimate [1000, 3275, 1291, 2788]. Estimated [2716]. estimates [687]. Estimating [1273, 1695, 2199, 2111, 2376, 1381, 1478, 620, 1214]. Estimation [3008, 2814, 2937, 3108, 3015, 2938, 3169, 1100, 2729, 2832, 3362, 3402, 3269, 2766, 2640, 2540, 2757, 3136, 1811, 3166, 1275, 1223, 1577, 1654, 1773, 76, 1818, 2023, 179, 1239, 1299, 1398, 3202, 1320, 2153, 1496, 1301, 1109, 1774, 1426, 809, 1693, 1590, 3371, 1117, 1322, 1682, 1477]. estimator [1338, 547]. Estimators [2501, 350, 241, 520, 1782]. Ethereum [3132, 3102, 3105]. Ethernet [352, 276, 579]. Ethernet-like [276]. Ethernets [406]. ethical [2219]. Euclidean [2288, 1479]. EURO [2200]. EV [3184, 3384, 2930]. evaluate [2374, 109]. Evaluating [1024, 229, 1766, 2186, 3189, 1394, 972, 1515, 1575, 2896, 2171, 1408, 1204, 1644, 712, 1136, 621, 1600, 1851, 864, 171, 59]. Evaluation [3354, 2731, 3212, 2801, 2482, 9, 3214, 643, 647, 2679, 2865, 415, 2848, 2777, 2582, 745, 3217, 654, 3456, 2807, 2456, 533, 371, 2684, 154, 1803, 740, 255, 3190, 1242, 1216, 3225, 2098, 1363, 1096, 1369, 1737, 927, 987, 95, 129, 2, 56, 904, 1002, 42, 3224, 1730, 2092, 555, 713, 1388, 891, 1341, 138, 163, 210, 364, 1382, 933, 2796, 15, 265, 197, 1746, 955, 1963, 176, 652, 455, 591, 2140, 157, 737, 1638, 821, 1283, 1017, 332, 832, 271, 98, 136, 929, 565, 362, 284, 393, 1622, 852, 1115, 1303, 1085, 249, 183, 1690, 1667, 3223, 490]. evaluation [2185, 502, 381, 47, 1023, 549, 1990, 166, 703, 704, 890, 308, 1652, 53, 12, 481, 324, 2177, 1566, 377, 2249, 97, 995, 668, 443, 346, 1422, 2158]. Evaluations [3418, 1197]. even [1947]. Event [2630, 2850, 2662, 2525, 2078, 37, 1100, 1239, 1640, 489, 1633, 462, 267, 588, 1794].

Event-based [2850, 1633, 1794]. Event-driven [2662, 37]. Events [39, 336, 1551, 2225, 2124]. Everything [3004]. Everywhere [3438]. Evictions [2463]. evidence [853, 1753]. Evolution [3221, 955, 121, 822]. evolutionary [1738]. Evolutions [2941]. Evolving [3401]. Exact [2412, 644, 2542, 2261, 2236, 2496, 2525, 3381, 3180, 1188, 701, 3085, 1521, 1480, 398, 1167, 1796, 1060, 1267, 148]. examination [1935]. examining [720]. example [130]. examples [820]. Exchange [2449, 2473, 3449, 778, 469]. exclusive [378, 628]. exclusive-write [628]. Executing [2988]. Execution [2699, 2464, 718, 116, 925, 1390, 766, 695, 1853, 151, 371, 530, 2212]. Exemplar [904]. Existence [3244, 982, 188]. existing [1035]. exit [484]. exogenous [2217]. Exoskeleton [3354]. expansion [296, 913]. Expansions [2917, 1165]. expectations [2240]. Expected [2716, 1167, 769, 679]. expediting [1431]. expense [172]. Experience [3072, 1795, 621, 2721, 1713, 120, 104, 70, 1821, 905]. experienced [129]. Experiences [846, 967, 1248, 1084]. Experiment [2499, 115, 313, 503, 802, 638, 1857]. Experimental [641, 904, 2613, 2718, 2851, 3012, 3351, 53, 946, 405, 1963, 406, 8, 2214, 2386, 17, 1561, 1023, 635, 380, 168, 241, 265]. experimentation [582, 719]. Experimenting [2457, 1084]. Experiments [858, 2492, 2159, 323, 688, 2219, 322]. Expert [426, 521, 504]. Explained [2852]. Explaining [3153]. Explicit [2622, 873, 2173]. Exploit [2441, 1276]. Exploiting [2719, 840, 2615, 2262, 1989, 1467, 1653, 2466, 889, 552, 2014, 1822, 2252]. Exploration [3393, 861, 1932, 1923, 1685, 680]. Exploring [2049, 1697, 1334, 2313]. Exponential [3207, 1600, 807, 3340, 3376, 2057, 2077, 3337, 2060, 622, 2065, 2053, 3215]. Exponentially [2538]. exposed [1855]. Exposing [2811]. Expressnet [460]. Extended [2566, 38, 2446, 1927, 3196, 2716, 3191, 2575, 206, 2893, 946, 945, 1687, 939, 1170, 401, 691, 1695, 940, 780, 781, 395, 396, 783, 1045, 1019, 1874, 784, 1044, 941, 448, 1446, 942, 752, 439, 21, 2135, 786, 801, 787, 943, 944]. Extending [2446, 3114, 1786]. extensible [764]. extension [436]. Extensions [2997, 347, 3323, 1278, 2045, 315]. Externalities [2440, 1648]. externality [2027]. Extinction [1507]. Extra [2631, 2863]. extraction [811]. Extrapolation [3318]. Extremal [2064]. Extreme [2589]. Extremely [3122]. extremescale [1923]. Eye [2603, 3360].

F [487]. fable [213]. face [61, 1823]. Facebook [2709]. Facilitating [3196, 2440, 1549]. facility [37]. Facing [3167]. factor [2123]. factorial [1594]. factorial-moment [1594]. factorial-moment/strip-integral [1594]. Factorization [2056, 3196, 2728, 2999]. Factors [574, 681, 616]. Faculty [2866]. Fading [2843]. failed [2018]. failing [397]. Failure [2895, 2901, 1303, 2833, 1997, 861, 1702, 104, 3228, 1378, 1527, 1945, 961, 1850]. Failures [2795, 2764, 2557, 2483, 2635, 3042, 1309, 2188, 1441, 2386, 2254, 2253]. Fair [2790, 3416, 3305, 2131, 2994, 3087, 1595, 2443, 1417, 2920, 2883, 2828, 2465, 1837, 2638, 2800, 2471, 1696, 1719, 1605, 1394, 1403, 998, 1347, 848]. Fairness [2472, 2859, 2932, 2473, 2931, 3047, 3440, 2824, 2900, 3293, 2498, 1537, 3427, 1943, 1843, 1141, 1110, 1475, 2924, 2203, 1630, 932, 1049, 1296, 1524, 3233, 2257, 1992, 1697, 2125, 2930, 2430]. fake [3382]. fake-post [3382]. False [3395, 2506]. Family [2479]. Fare [2667]. Farms [2592, 1765, 1601, 1776, 1091]. FaRNet [2263]. Farout [2207]. Fast [2152, 1390, 2731, 342, 843, 2744, 1495, 2605, 3295, 3287, 1148, 2861, 668, 2711, 795, 1254, 1134, 1300, 766, 534, 2024, 77, 80, 1059, 1283, 1335, 1298, 1033, 1707, 1667, 215, 1930, 1103, 2263, 2414, 845, 2336, 1800, 2343, 504, 215]. Fast-Convergence [2605, 2861]. Faster [913, 959, 1299, 1457]. fat [1574]. fat-tree [1574]. FatTree [2458]. Fault [2555, 2842, 992, 473, 898, 553, 2095, 577, 448, 801, 1043]. fault-based [992]. Fault-Tolerant [2842, 448]. faults [1504, 238]. Faulty [2529]. favoring [1474]. FB [1525, 1348]. FCFS [2294, 3120]. FDDI [439]. Feasibility [1035, 3111, 1822, 1917]. Feasible [2443, 3167]. Feature [3167, 2598, 163]. Features [2959, 959, 1750, 1648]. FEC [1261]. federal [66, 97]. Federated [3432, 3264, 3116, 3231]. FedQV [3432]. Feedback [31, 2986, 3308, 3333, 537, 2629, 2996, 3017, 197, 298, 872, 1914, 2302, 201, 113, 943]. feedback-based [872]. Feedforward [2956, 1731]. fetching [3442]. few [1958, 1567]. FFT [3189]. FFT-based [3189]. FFTs [902]. Fi [3329, 2839]. Fiat [3301]. Fibre [837]. Field [3216, 3292, 2500, 2528, 2785, 2716, 2853, 2886, 2917, 2686, 2515, 2483, 2658, 2626, 3205, 162, 2717, 2995, 2610, 2715, 3160, 1316, 2784, 1761, 2645, 3059, 1898, 1841, 2046, 1263, 311, 3130, 1896, 3291, 1768, 1909, 2239, 2928]. FIFO [562, 258]. fifteenth [754]. FIG [3220]. fight [2320]. Fighting [3105]. figure [48]. figures [55, 14, 25]. File [893, 707, 1029, 887, 1708, 268, 838, 617, 1035, 904, 1074, 1341, 768, 975, 1140, 1379, 764, 782, 926, 664, 1507, 454, 2337, 446, 1508, 1523, 767, 560, 366, 420, 674, 479, 675, 1692, 1660, 2033, 1153, 744, 523, 1422]. file-level [1067]. file-sharing [1692]. file-system [1708, 975]. files [1146]. filesystems [748]. Filter [2350, 1495]. Filtering [2604, 1880]. Filters [2599, 3395, 2506, 1852, 1569, 1452]. find [814]. Finding [1720, 1488, 1532, 2534, 1070]. findings [1755]. Fine [1363, 2850, 1704, 1961, 3304, 3254, 1067, 2148, 2244]. Fine-grained [2850, 1704, 1961, 3304, 3254, 1067, 2148, 2244]. Finger [3331]. Fingerprinting [2598, 2362]. Fingerprints [3155]. Finite [3250, 2998, 534, 2844, 2645, 1842, 1612, 2081, 491, 1718, 584, 450, 151, 923, 619, 1386, 484, 676, 287, 377, 3220]. Finite-Buffer [2998, 1612]. finite-buffered [619]. Finite-Sample [3250]. Finite-Sum [2844]. firewall [1148]. Firming [2264]. First [2761, 2053, 3329, 2459, 2908, 3326, 1190, 2106, 2225, 1583, 2336]. First-Come [3326]. first-order [2336]. first-principles [1583]. First-Served [3326]. Fisher [2499, 2111]. Fit [2832, 1167]. Fitness [2611, 1461, 1548]. Fitting [3337, 2090, 2061, 1229, 1712]. Fixed [1123, 3422, 1996, 3371, 1385, 589, 1093, 1900, 2676, 1662, 1533]. Fixed-point [1996, 1533]. Flash [2719, 2849, 2483, 2706, 2384, 1767, 2405, 2120, 1879, 2099, 2239]. Flash-based [2719, 2239]. Flexibilities [2560]. Flexibility [3193, 2307, 2237]. Flexible [1497, 3287, 2514, 2990, 2443, 2475, 2913, 563, 449, 1283, 1625, 748, 2010, 1476, 2910, 1770, 845]. Flipped [3210]. Floating [2961, 2706]. Flood [1232]. flooding [1798]. Flow [2795, 3076, 1292, 2932, 2994, 3229, 1565, 3095, 3451, 1289, 2056, 1505, 317, 2796, 197, 2131, 2261, 720, 2001, 1299, 351, 1942, 1605, 1068, 1301, 1381, 288, 784, 1856, 1774, 1655, 3300, 1595, 1847, 556, 1116, 2990, 2363, 318, 572, 1321, 188, 353, 585, 155, 1411]. Flow-Control [3451, 784]. flow-controlled [288]. Flow-level [1565, 1942, 1605, 1856, 1847, 1321]. flowcharts [114]. FLOWR [2399]. Flows [3188, 3397, 2194, 1316, 1357, 1306, 1623, 1270, 337, 1215, 1394, 1290, 431, 1624, 999, 3233, 1054, 1474]. fluctuating [1483]. fluctuations [1311]. Fluid [2696, 2531, 2796, 2679, 1728, 2786, 2302, 1494, 1257, 3095, 3065, 3205, 2041, 2110, 2056, 2075, 922, 1807, 1603, 1260, 2068, 1897, 2086, 1858, 2088, 2069, 2070, 2930, 2416, 1344]. Fluid-Flow [3095]. fluid-fluid [2075]. Fluid-Model-Based [2696]. focused [1549]. Foes [2674]. Fog [3251]. follow [1580, 2350]. Following [3041]. footprint [1875]. Footprints [428, 1656]. forecast [2187]. Forensics [2643, 2119]. forever [2062]. Foreword [3356, 3367, 1637, 1536, 1608, 1620, 1519, 1593, 1506, 1448, 1457, 1342, 1433]. Fork [2495, 2588, 2858, 860, 979, 1556, 1865, 830]. Fork-Join [2495, 860, 979]. fork/join [1556]. form [2935, 762, 2079, 178, 536, 3334, 1605, 1561, 3257, 2055, 881]. Formal [2810, 1734]. Formalizing [1346]. format [1098]. Formation [2568, 2666, 3235, 3199, 2170]. formed [829]. forming [3088]. forms [3017, 1842, 1241]. formula [1420]. formula-based [1420]. formulae [85]. Fortier [658]. FORTRAN [316, 317]. Forums [3322]. Forward [2603]. Forwarders [1216]. Forwarding [2711, 3283, 2133, 1783, 2043, 1333]. foster [2350]. Foundation [647, 3385]. Foundations [323, 365]. FPGA [1963]. FPGA-based [1963]. Fractal [602, 1800]. Fragmentation [2813, 1868, 1785]. frame [1008, 848, 513]. frame-based [848]. Framework [2955, 3250, 3063, 2454, 3295, 3345, 3304, 3082, 3437, 2847, 3006, 2558, 3135, 1250, 1867, 2935, 2156, 1076, 1922, 1254, 123, 2044, 1421, 2045, 1924, 1283, 1932, 832, 1298, 2374, 2404, 2670, 917, 2352, 1476, 2212, 1429, 1865, 2430, 2927]. frameworks [1365, 2216]. fraud [2320]. Free [1530, 2709, 801, 980, 2220, 1018, 1193, 1783]. Free-Drop [1530]. Freestyle [2623]. Frequency [3089, 2667, 2762, 2757, 2712, 793, 1115, 561]. frequency-based [561]. frequently [982]. Freshness [3361, 3161, 3365]. Frieberger [9]. friendliness [1649]. friendly [1357, 1073, 215, 999]. Friends [2674]. Frontier [2728]. FTP [707]. Full [2478, 1190, 2922, 1664, 1279, 1283, 2110]. Full-data [2922]. Full-Duplex [2478]. Full-system [1190, 1664, 1283]. Fully [3280, 1677]. FuncPipe [3295]. Function [3385, 2526, 116, 1762, 2085]. Function-as-a-Service [3385]. Functional [88, 503]. Functions [3204, 2641, 1791, 1447, 206]. fund [356]. Fundamental [3052, 2628, 805, 1483, 806, 3187, 1397, 2607, 364, 2347]. fundamentals [63, 478, 2279, 598]. Further [187, 47]. fused [1929]. Fusion [2781]. Future [2946, 3070, 1566, 836, 1462, 120, 1246, 2300, 995]. Fuzzing [3154].

G [1592, 1691, 2298, 1158, 384, 671, 198, 856, 2997, 1348, 3447, 1597, 1720, 3343, 3373, 450, 654, 3445, 1188, 1267, 3024, 3122, 1169, 2289, 3389, 2941, 1161, 800, 1258, 3343, 3373, 450, 1970]. G-networks [1161, 800, 1258]. G/G/ [3343, 3373]. G/G/1 [3343, 3373, 450]. Gacha [3290]. GAIA [1509]. Gains [2478]. Game [2451, 2769, 3290, 2821, 2162, 2667, 2668, 2515, 2571, 3161, 2695, 2320, 1912, 2949, 2167, 1302, 1233, 573]. Game-Theoretic [2821, 573]. Games [2645, 2923, 2572, 2633, 2571, 2626, 2445, 3199, 2573, 1801, 2318, 1698, 1849]. gaming [2040]. gamma [2077]. Ganesha [1816]. gang [966]. gang-scheduling [966]. GAO [66]. Gap [3015, 827]. Garbage [2976, 2546, 1287, 1193, 2239]. Gate [2706, 1896]. gate-limited [1896]. gateway [572, 573]. Gateways [1920]. gathering [134, 96]. gating [2394]. Gaze [3360]. Gazer [2976]. GB [3123, 2772]. GB-PANDAS [2772]. Gbps [3247]. GDM [2387]. Gelenbe [604]. GemDroid [2374]. Gemini [2177]. GEMM [1930]. Gene [1501]. Gene/L [1501]. General [3452, 3063, 3453, 3065, 3122, 2439, 2928, 370, 2298, 340, 2074, 1863, 3367, 1182, 691, 36, 2750, 2281, 619, 549, 348, 414, 148, 1865]. general-purpose [36]. generalised [1746]. Generalization [2643, 441]. Generalized [1171, 2589, 3073, 3028, 1252, 2434, 670, 429, 947]. Generalizing [2450]. generally [2046]. generally-timed [2046]. generated [3337, 2327]. Generating [927, 1583, 951, 1820, 360, 2065, 1247]. Generation [2946, 3429, 2493, 2518, 116, 2156, 2034, 1056, 2178, 2310, 2256, 2228, 917]. Generative [3112, 3060, 765, 2957, 259]. Generator [3220, 764, 2309, 1321, 1156]. GENI [2159]. Geo [2973, 2754, 2794, 2798, 2624, 2278, 2253, 2272]. Geo-Distributed [2624, 2973, 2754, 2794, 2798, 2253, 2272]. geo-replication [2278]. Geographic [2687, 1955, 1138]. Geographical [2031, 2658, 1953]. Geographically [2563, 2045]. geolocation [2160, 1980]. Geometric [3033, 2208, 2249]. geometry [2204]. Georganas [644, 701]. George [658]. get [1193]. Gets [2602]. Getting [279, 1017]. GHz [2475]. GI [1170, 807, 1230, 3062, 1170, 2997, 2080, 1267, 1169]. GI/G/1 [2997, 1267, 1169]. GI/G/1-type [1267]. GI/GI/1 [1170]. GI/Ph/n [2080]. Gibson [3]. GISMO [1156]. Gittins [3444, 3343, 3373, 3122, 3162]. glass [1957]. Glauber [2518]. Global [2995, 3319, 1069, 496, 921, 880, 1770]. globally [989, 916, 290]. globally-managed [916]. globally-synchronizing [989]. Gnutella [1427]. Go [3285, 3123, 1289, 1312]. Go-to-Controller [3285]. goal [965]. goes [117]. Going [3092]. gone [1829]. good [2332, 2120]. Google [2200, 1375, 1831, 2954, 2364]. Gossip [1951, 1737, 1761, 1735, 1621]. Gossiping [3116, 1734, 1732, 1733, 1736, 2424, 1468]. Got [3413]. governed [1274]. government [66]. GPGPU [2387, 2757, 3136, 2251]. GPGPUs [2420]. GPRS [1086]. GPS [1472, 1180]. GPU [3086, 2848, 2615, 2905, 2172, 1930, 2181]. GPUs [3420, 3191, 1925, 3289, 2978]. Gradient [2845, 2844, 3169, 2532, 2724, 3269, 3007, 2940, 687]. Gradient-Based [2724]. Gradients [3378]. grain [1363]. grained [2850, 1067, 2148, 1704, 1961, 3304, 2244, 3254]. Granular [3247]. granularity [2329]. GranularNF [3247]. Graph [2590, 2811, 2591, 1097, 3260, 2690, 3054, 3348, 3230, 2632, 2497, 2547, 2999, 3037, 2368, 1756, 1957, 2104, 2127, 2267, 2118, 154, 119]. graphical [2782]. graphics [61]. Graphs [2918, 3014, 2589, 2548, 2967, 3013, 2010, 3291, 2966, 3311, 3152, 1958, 60, 2126, 2344, 22, 47, 25, 1662, 1160, 2224]. Gray [750, 657]. GRE [1578]. GreatSPN [3225, 1741]. greed [573]. Greedy [2918, 1783, 2491, 3145, 2265, 2546, 1262, 2170]. Greedy-Bayes [2491]. Green [2661, 2457, 2658, 2509, 2032, 2122, 2038, 2402, 2317]. Greenhouse [3353, 2314]. Greening [2537, 1953, 1837]. GreenMetrics [1916, 2030]. Grid [3393, 2766, 2188, 2191, 1544, 2313, 1243, 1248, 2311, 1245, 1246, 1244]. GridG [1247]. Grids [2937, 2795, 2760, 2515, 2497, 2509, 2767, 2422, 2410, 1247, 2192, 2893, 2354]. grooming [1118]. groove [1734]. Ground [3154, 1663]. Ground-Truth [3154]. Group [3440, 894, 66, 1133]. Group-guaranteed [894]. grouping [1044]. groups [875, 557]. growth [238, 2208]. GSD [162]. GSO [2945]. GSPN [951]. GSPNs [953]. GSTF [2291]. Gt [2640]. GuaNary [3419]. Guaranteed [2502, 1955, 894, 1042]. Guaranteeing [2163]. Guarantees [2698, 2810, 2504, 2596, 2840, 3396, 1535, 1546, 1694, 806, 1068, 1351, 2035, 774, 1459]. Guardrails [2970]. guessing [1557]. Guest [1243, 1506, 956, 1448, 1457, 1342, 1433]. guide [1484, 419]. guided [861]. guideline [98, 136]. guidelines [66].

H [597, 605, 646, 647, 653]. Hadoop [2721, 2042, 2468]. Half [2534]. Half-Latency [2534]. Halfin [2078, 2881, 3325]. Hall [640, 594, 596, 605, 655, 653, 663]. Halstead [244, 547, 322, 315, 520]. HAM [348]. HANA [2584]. hand [1057]. hand-off [1057]. Handbook [750, 657]. handling [2310]. handover [2943]. Hands [2801]. Hands-on [2801]. Happen [3395]. haptic [1143]. Hard [1629, 1268, 2778, 1504, 1292, 615, 438, 1436]. Hard-sphere [2778]. Hardness [3101]. Hardware [3399, 3355, 190, 2394, 730, 74, 2174, 2388, 18, 238, 1940, 23, 1148, 1408, 813, 89, 752, 1988, 2177, 1977, 10, 19, 1935, 2013, 2124]. hardware-based [1148]. Hardware-Software [3355, 1988]. harmful [1471]. Harmonic [2442, 2937]. Harnessing [3067]. Harpoon [1321]. HARRIS [666]. HARTS [740]. harvested [1920]. Harvesting [3151, 3300, 2314, 2423, 2378]. Hash [2614, 2049, 759]. Hashing [2922, 1569]. Hastings [2127]. haul [3438, 3328]. having [151]. hazard [3442, 1019]. Head [3360]. Head-Mounted [3360]. headers [1126]. HEAL [3407]. heartbeat [1997]. heat [2314]. Heavy [1994, 3452, 2595, 2970, 3327, 3091, 2902, 3073, 3265, 3376, 3375, 3028, 2549, 2587, 2803, 3061, 3450, 3064, 1893, 2503, 3194, 1792, 2800, 3344, 3390, 3240, 2715, 2858, 2860, 2855, 2913, 2968, 1357, 2285, 1475, 1995, 1512, 2287, 1438, 1019, 1892, 2279, 2289, 2305, 2772]. Heavy-Tailed [2595, 1357, 2289]. heavy-tails [1512, 1438, 2279]. Heavy-Traffic [2549, 2587, 2800, 3344, 3390, 2860, 3073, 3028, 1893, 1792, 2855, 2913, 2968, 1892, 2772]. Hegemony [3153]. held [2240]. Help [2592, 2363]. HEMI [39]. Hershel [2362]. heSRPT [3022, 3074]. Heterogeneity [2498, 2466, 2417, 2273, 2027]. Heterogeneous [3141, 3276, 2562, 3020, 3075, 3063, 3259, 2707, 2815, 2756, 3256, 2653, 2883, 2637, 2471, 3338, 3380, 3079, 3231, 1171, 2019, 2036, 2103, 1866, 2302, 2104, 1936, 438, 2048, 494, 1977, 1946, 1697, 2327]. HeteroScouts [1977]. Heuristic [2682, 2531, 1167, 261, 348]. heuristically [45]. Heuristics [2097, 434]. Hewlett [904]. HFTraC [2712]. Hidden [1111, 2915, 1261, 1614, 1855]. Hierarchical [329, 2932, 3183, 1651, 988, 2193, 627, 855, 906, 388, 564, 302]. hierarchies [331, 349, 183, 376]. hierarchy [951, 1391]. Hieroglyph [2690]. High [2836, 3420, 2944, 1619, 1194, 2923, 3000, 2100, 690, 2171, 599, 2244, 3007, 1030, 2712, 3281, 735, 2121, 692, 1503, 969, 1061, 360, 291, 1569, 1776, 1665, 678, 2262, 976, 1053, 2380, 1633, 1137, 2263, 1174, 20]. High-bandwidth [3420]. High-capacity [3420, 1030]. High-density [1194]. high-dimensional [2380]. high-end [1633]. High-Endurance [2836]. High-Frequency [2712]. High-level [1619, 599, 360]. high-order [2262]. High-Performance [3000, 2100, 692, 969]. high-reliability [1503, 1665]. high-speed [678]. High-Throughput [3281, 2244, 2121]. high-variability [1776]. high-volume [1174]. HighEnd [2049]. higher [1597, 441, 1395]. Highleyman [605, 653]. Highlight [2936]. highly [994, 250, 634, 767]. highly-associative [634]. Hill [745]. Hints [2580]. histograms [1289]. history [1420, 312]. history-based [1420]. Hit [2544, 2742, 2907]. HMTT [1664]. Hoc [2964, 1212, 1325, 1115, 2010, 1084]. holding [2356]. Holland [647]. home [2134, 1802]. homogeneity [198]. homogeneous [2022, 350, 437, 622, 289, 296]. Homophilous [3153]. Hop [3438, 2784, 2698, 1469, 1631, 2316, 1113, 2153, 678, 2676, 1764, 2530, 1199, 1846, 1845]. hop-by-hop [1631, 2316]. HOP-TERRAIN [1845]. Hopping [2456, 1115, 2247]. hops [3152]. Horizon [3309, 3270, 3320]. Horwood [656]. Host [3185, 1450, 1274, 1914, 3246]. host-based [3246]. Hosting [3268, 3321, 1407]. Hosting-Based [3321]. Hosts [3407, 1479, 1138, 300]. HostView [1914]. Hot [2708, 2702, 1686, 2114, 500, 1313]. Hot-Data [2708]. hot-potato [1686, 1313]. hot-spot [500]. Hotwire [3100]. Hound [2831]. Hour [2701]. Hour-Ahead [2701]. Howarth [594]. HPCC [1501]. HPDC [2037]. HPDC/SIGMETRICS [2037]. HSLAN [492]. HTTP [2406, 2533]. httperf [960]. hub [1839]. Human [97, 61]. Hurst [934]. Hurt [3193]. Hybrid [735, 3420, 2976, 2758, 2673, 2905, 3443, 2945, 2509, 2003, 1338, 1254, 1276, 1658, 1050, 759, 1926, 2309, 19, 2341, 1931, 2039, 39]. Hyper [2501, 2996, 987, 3228]. Hyper-Compact [2501]. Hyper-Scalable [2996]. hyper-scale [3228]. HyperBench [3002]. hyperbolic [1783]. hyperchannel [408]. hypercube [498, 496, 623, 778, 629]. Hyperexponential [2995]. hypergraph [2344]. Hyperscale [2524]. hysteresis [882, 1059].

I/O [2836, 340, 938, 977, 422, 713, 720, 796, 1809, 100, 901, 256, 674, 918, 3426, 2280, 194, 1153, 1228, 1429]. IaaS [1976, 2470]. IBM [868, 86, 363, 310, 789, 162, 24]. Iconic [61]. ICPE [2015]. IDBUG [118]. Idealized [3085, 1965]. Identical [2934]. Identifiability [2776]. Identification [2937, 2493, 2004, 1181, 2279, 2048, 2247, 1759]. Identify [3135, 1291]. Identifying [2681, 3040, 94]. Idle [2750, 2882, 889, 945, 973, 801]. idleness [1586, 1769]. IEEE [611, 588, 1691, 1298, 2677, 1385, 1993, 1919, 2535]. If [3009, 1527, 2215]. Ignite [2564]. II [2868, 2875]. Illegal [3100]. Illuminating [3301]. Image [1701]. Images [3294]. Imbalance [2617, 2855, 3384]. Impact [1110, 2523, 3387, 2671, 2565, 2765, 2838, 2498, 2577, 2408, 2389, 2841, 2093, 1314, 2190, 1287, 1389, 1610, 1834, 2241, 1766, 2226, 485, 2001, 179, 1473, 624, 1057, 2150, 1942, 1077, 1940, 1835, 2370, 755, 918, 1408, 966, 2375, 1919, 968, 2433, 2073, 2145]. Impacts [1470, 824, 1624, 1340]. impaired [1789]. impatience [2029]. imperfect [1477]. Implementation [3258, 1690, 334, 1132, 529, 224, 1332, 989, 1930, 2132, 1926, 1652, 93, 152, 1108]. implementations [1148, 1931]. implemented [2394]. Implementing [225, 634, 916, 67]. implication [1960]. Implications [2481, 2322, 3355, 2028, 1038, 1755, 528, 2122, 1767, 1889, 1442, 1045, 2240, 2154, 2099, 1753]. Implicit [1368, 935]. Implied [2469]. imply [2122]. importance [2143]. Impossibility [3101, 2547]. Improbability [3220]. Improve [2620, 2556, 1183, 1972, 1498, 1219]. Improved [1793, 3188, 2591, 368, 3192, 2705, 2999, 1979]. Improvement [3326, 105, 42, 74, 253]. Improving [3141, 1155, 3231, 1997, 2044, 3071, 2661, 1063, 3120, 3447, 3228, 1200, 920, 2042, 3094, 2043, 2486, 1107, 2849, 1513, 2806, 3205, 1006, 2435, 3109, 1832, 940, 1078, 1545, 1770]. IMS [421]. In-degree [1913]. In-depth [2850]. Incentive [1848, 1265, 3080, 2607, 2327, 1302, 2676, 2428]. Incentives [2436, 2452, 2320, 1571]. incentivize [2024]. Incentivized [2781]. Incentivizing [2703, 1858, 3070]. including [613]. inclusion [716]. inclusions [1898]. Incompletely [2587]. Incorporating [369]. Increase [2700]. Increasing [2748, 2780, 731, 1159, 3442, 887, 731]. Increasingly [3126]. Incremental [2916, 1053]. Independent [2518, 1664, 1094, 1120, 360, 1908, 1312].

Index [2474, 2823, 307, 3111, 2356, 1895, 2001, 1774, 707, 1516, 3172]. indexing [545]. Indicators [2684, 2038]. indices [402]. indirect [1338, 1502]. individual [1802, 2530]. Indoor [2475]. Induced [2720, 2312, 1653]. Industrial [2797, 2694]. industry [1925]. inefficiencies [1965]. Inequalities [3051, 2889, 2904]. inexact [1526, 1647]. infection [2119]. infer [1867, 1918]. Inference [3249, 2601, 2730, 3400, 3284, 2888, 3363, 3364, 2512, 3037, 1014, 1449, 2365, 1453, 1410, 1412, 1782, 1329]. Inferring [3072, 1872, 1195, 2975, 1754]. infinite [2294, 208, 1343, 954, 2919]. infinite-server [2919]. infinite-state [954]. Influence [2442, 3286, 2729, 3262, 718, 2340]. Information [2441, 3294, 2781, 3055, 2816, 2643, 2499, 2916, 2477, 3118, 3361, 3267, 3372, 2577, 2497, 3365, 2992, 2711, 683, 935, 736, 2005, 1839, 1154, 2325, 1705, 2220, 7, 783, 865, 2111, 4, 1819, 2153, 2259, 1561, 1386, 1265, 1142, 1905, 2329, 2300, 220, 2685, 1526, 1647, 1521, 1477]. Information-based [2499]. Information-update [3361]. Informed [878]. informing [1683]. Infrastructure [3438, 2769, 2770, 2583, 2894, 2700, 2763, 2898, 3039, 2892, 2939, 869, 1275, 1351, 1244, 1960]. Infrastructures [3100, 2897, 2768]. Infusing [3134]. ingress [1403]. Inherent [2824, 1937]. inherently [1048]. inhomogeneous [1691]. initiated [395, 775]. inline [551]. Innovation [2671, 2565]. Input [3149, 2689, 3091, 3265, 2654, 3095, 3065, 3083, 2999, 3136, 2298, 1482, 629, 141, 1251]. Input-Dynamic [3149]. Input-Queued [2689, 3091, 3265, 2654, 3095, 3065, 3083, 2999]. input/output [629]. InputQueued [2205]. inputs [1094, 1120]. Insensitive [1318, 2612, 1392, 2321, 1393]. Insensitivity [1601, 3129, 2800, 3336]. Inserted [2847]. Insertion [2921, 2376, 2433]. insights [1831, 1084]. instabilities [2359]. Instability [2608, 1260, 2293, 2530]. installations [95]. Instances [3238, 2722, 2773]. Institute [310]. Instruction [314, 386, 1250, 766, 73, 668]. instruction-set [766]. Instructional [3211]. Instrument [39]. Instrumentation [645, 632, 2847, 2016, 1617, 510, 180]. Instrumentation-based [2016]. insularity [455]. Insurance [2806, 1170, 2168, 1905, 2202]. integral [1594]. Integrated [863, 2942, 815, 407, 425, 508, 459, 1988, 318, 1751]. Integrating [851, 153]. Integration [1357, 2584, 1306, 235, 246, 2069, 1794]. Intel [3419, 888, 629, 3004, 2852]. Intelligent [3288, 3148, 3177, 18]. Intensity [2611, 2640]. Intensive [3282, 945, 505, 1965, 3256]. Inter [932, 2502, 2649, 2133, 1686, 1987, 1183, 827, 2253, 1692, 1939]. inter-core [1939]. inter-datacenter [2253]. Inter-domain [2649, 2133, 1686, 1987, 1692]. Inter-IXP [2502]. Inter-receiver [932]. inter-reference [1183, 827]. interacting [1316, 186]. Interaction [1676, 1442, 825, 1862]. Interaction-aware [1676]. Interactions [3012, 2296, 1498, 2313]. Interactive [2988, 3331, 2464, 182, 994, 75, 79, 91, 1063, 485, 335, 283, 402, 2403, 158, 1042]. Interchange [2305, 1894]. Interchangeable [2883]. intercloud [2042, 2044]. Interconnect [2171, 586, 2177]. Interconnected [3042, 381]. interconnection [470, 584, 501, 619, 693, 500, 867, 469, 743, 567]. Intercontinental [3438]. Interdependencies [2763, 366]. Interdependency [2768]. Interdependent [2897, 3217]. interest [2348, 1588]. interest-driven [2348]. Interface [2941, 1340]. Interference [2944, 1622, 2468, 1869, 54, 2214, 779]. Interference-aware [1622]. Interferences [2511]. interleaved [1796]. Intermittent [2701, 525, 2228, 2353]. internal [1338, 1014]. International [3066, 3454, 2892, 2015]. Internet [1156, 1975, 3413, 2674, 2194, 2275, 1209, 1584, 1110, 1132, 1218, 1112, 1217, 2322, 1591, 991, 1000, 1826, 1069, 1253, 2151, 2378, 1580, 2522, 872, 1008, 1396, 1050, 1072, 1149, 1383, 1345, 1771, 1875, 1479, 1648, 2167, 1921, 2453, 1401, 2735, 3092, 2008, 1053, 2815, 1380, 1877, 1663, 822, 1138, 1635, 2208, 638, 2626, 3436, 1757, 3070, 3369, 2975, 1959, 1703, 3068, 1289, 1030, 1406, 995, 1153, 961, 1549]. Internet-like [1635]. Internet-of-Things [2815]. Internet-Scale [2522]. Internet-wide [3413]. internets [869]. internetworks [909]. Interoperability [3352, 2211]. Interposed [1288, 1429]. interpretation [802]. interpreting [689]. Interpretive [40]. interruptions [338, 412]. Interrupts [3408]. Interscience [595]. Intersection [2733]. Interval [1611, 1130, 1721, 1688, 1013, 774]. intervals [1997, 1018, 133, 1487]. intra [1314, 1686, 2050, 1503, 1665, 2253]. intra- [1686, 2253]. intra-disk [1503, 1665]. intra-domain [1314]. intra-node [2050]. Intrapersonal [2771]. intrinsic [1975, 1767, 202]. Introducing [1414, 2169]. Introduction [2159, 948, 956, 2801, 745, 2069, 264]. IntroPerf [2365]. Intrusion [3230, 1684, 2338]. intrusive [3379, 2366]. invariant [68]. invariants [850]. Inventory [2646, 3182, 2983, 2736, 2792, 3048, 2067, 1899, 2073]. Inverse [3143, 3215, 3243]. inversive [1169]. Invertible [3396]. Investigating [2045, 2502, 2027, 1129]. Investigation [2693, 2954, 210, 545, 380, 132]. investment [172]. IO [1825]. IoT [2941, 3331, 2660]. IP [1086, 1056, 1009, 2160, 738, 1257, 1139, 1103, 3437, 1333, 1980, 1126, 913, 1313, 1043, 1208, 1176, 1683]. IP-based [1086].

IP/ATM [1009]. IPFS [3425]. iPSC [629]. iPSC/2 [629]. IPTV [1775]. IPv6 [1739, 3301]. IQ [1819]. IRIS [731]. Ironmodel [1666]. Irregular [2600, 2383, 931, 2420]. ISA [2815]. islands [213, 1876]. isolated [381]. Isolating [1340]. ISP [1915, 1975, 1009, 1772, 2169, 2323, 2443, 1005, 2242, 1583]. ISPs [1382, 2169]. Issue [2461, 3170, 3246, 901, 1082, 2899, 895, 907, 964, 2876, 3016, 3053, 3156, 3234, 3332, 3439, 1007, 1916, 1740, 3214, 950, 1012, 1087, 1157, 1227, 1433, 1243]. Issues [56, 1082, 456, 991, 969, 423, 898, 1072, 871, 997, 896, 2235, 427]. Itanium(R) [1581]. Item [3433]. Iteration [3319, 809]. Iterative [2485, 2438, 3083, 1059, 347, 2403, 1587]. itself [2075, 312]. IXP [2502].

J [640, 594, 596, 658, 789, 663, 588]. J2EE [1275]. Jackson [2434]. Jain [641, 595]. Java [2122, 2050, 2801, 1064, 1065, 1121]. JavaScript [3304]. Jetson [3296]. Jim [750, 657]. JINSIL [892]. jMarkov [2091]. JMS [1500]. JMT [1742]. Job [2856, 3448, 3204, 3349, 3370, 1088, 11, 2467, 3381, 3064, 2995, 2928, 967, 1766, 3334, 3327, 1597, 362, 1544, 3257, 1785, 308, 206, 966, 968, 1647, 260]. job-size [1647]. Jobs [3022, 3203, 3178, 2563, 3381, 2747, 2911, 3121, 3127, 2699, 3239, 1230, 817, 1474]. jobscheduling [1186]. John [641, 604, 746, 419]. Join [2884, 2750, 2882, 2495, 2588, 2858, 759, 860, 979, 1304, 1556, 1865, 830]. Join-Idle-Queue [2750, 2882]. Join-shortest-work [2884]. joins [826]. Joint [2415, 2914, 2506, 3316, 3375, 2292, 1402, 2624, 2497, 2535, 2317, 1723, 2409, 1945, 2243, 2430, 2436]. Jointly [2349]. Journaling [2922]. JS [3304]. JSQ [3452, 1601, 2066, 2996]. jump [2053]. junctions [2768]. June [588].

Kalman [1452]. Kant [745]. Karhunen [2336]. Kaufmann [750, 657]. Keeping [2970]. Kelly [1166, 2165, 419]. Kernel [3408, 787, 1152, 1389, 2336]. Kernel-based [787, 1152]. Kernelized [3308]. Kernels [3086, 1929, 510]. Key [3000, 2105, 1133]. key-value [2105]. keynote [1032, 754, 3099]. kinetic [2378, 2685]. King [640, 596, 663]. Kingman [2997]. Kiviat [60, 14, 22, 47, 25]. KL [1964]. Knapsack [3146, 3147, 3307, 3145]. knapsacks [1903]. Knee [2534]. Know [3143, 3004, 1819]. Knowledge [518, 3352, 2582, 2797, 2328, 1350]. Knowledge-based [518]. Knowledge-driven [2582]. Known [3022, 3027, 1310]. Koskela [645]. KPC [2090, 1712]. KPC-Toolbox [1712, 2090]. KPI [3108]. Kraken [2504]. Krishna [745]. KSR1 [719].

L [588, 1501]. L2 [1944]. Labeling [3014, 2230]. Labelling [3227]. Laboratories [2797]. laboratory [229]. Labs [3298]. lack [2685]. Lagrangian [2760, 2521]. Lagrangian-based [2521]. Lambda [1026]. LambdaGrids [1497]. LAN [1198, 1166, 481]. landmark [1682]. Language [7, 599, 111, 174, 360, 1851, 554, 908, 152, 155, 907]. languages [124, 185, 2177]. LANs [1298]. LaPalice [672]. Laplace [3215, 3243]. Laplacian [2433]. laptops [2154]. Large [2630, 3029, 2980, 2748, 2780, 2918, 1094, 3406, 3414, 2750, 2749, 2891, 3140, 3433, 3376, 3237, 3096, 2680, 3217, 2483, 2688, 2959, 2448, 3339, 3434, 2498, 2653, 2717, 2926, 2553, 2494, 3037, 3079, 2098, 1262, 2105, 129, 2188, 2422, 1616, 869, 1480, 1702, 1795, 1382, 933, 1818, 2303, 1409, 975, 1828, 575, 1240, 1841, 782, 847, 1230, 1187, 1114, 415, 100, 2340, 504, 2276, 130, 1165, 1970, 1044, 2238, 1257, 1897, 2377, 2392, 1989, 1854, 2222, 1775, 1870, 675, 269, 1768, 2106, 494, 1476, 2041, 2227, 2146, 2201, 1377, 1460, 1931, 1913, 849, 513, 625, 1219, 1855]. Large-Scale [2630, 3029, 2749, 3140, 3096, 2483, 2688, 2959, 3339, 3434, 2494, 2750, 2498, 2653, 2926, 2553, 3079, 2098, 2105, 129, 2188, 1616, 1702, 1818, 2303, 1409, 975, 1841, 782, 2340, 1044, 2238, 1257, 1897, 2377, 2222, 675, 1768, 1476, 2227, 1931, 1913, 625]. Larger [3422, 3275]. Largescale [2442, 2334]. largest [1291]. LargeWeb [2907]. LAS-based [1294]. lasting [1753]. Latency [2682, 2613, 3071, 2523, 2496, 3444, 2505, 2543, 2720, 3329, 2594, 2534, 3324, 2553, 2415, 405, 1961, 2138, 1424, 1757, 1337, 2363, 2244, 2016, 1668, 915]. Latency-Cost [2543]. lateness [883]. Latent [3270, 3320, 1570]. Latte [3071]. Law [3152, 1274, 1887, 649, 1786]. Laws [3187]. Layer [3437, 3436, 1518, 1872, 1295, 2365, 983, 2852, 1879, 1404, 2221]. Layered [1688, 1033]. layout [1022]. layouts [1568]. lazy [200]. LDAP [1055]. leaders [2552]. lean [2122]. Leaping [3222]. learned [1375, 3111]. Learning [2782, 3297, 3115, 3029, 2986, 3108, 2758, 2845, 3250, 3432, 2487, 3316, 3170, 3191, 2869, 2964, 3060, 2967, 3296, 3237, 3409, 2962, 2486, 3295, 3401, 3345, 2827, 2961, 2958, 2514, 2118, 3348, 3347, 3102, 2526, 3277, 2951, 3164, 3252, 3270, 3320, 2875, 3183, 3105, 2966, 3007, 3035, 3157, 2828, 3264, 2462, 3171, 3266, 3036, 2846, 2960, 2830, 2825, 3319, 3405, 2831, 3197, 3366, 2397, 104, 2358, 2324, 2259, 1553, 1880, 3172, 2012, 2338, 585, 2341, 2399, 1228, 2343, 2158, 3231, 3238]. Learning-Augmented [3409, 3157]. learning-based [2341]. Learning-enabled [3366]. lease [2545]. Least [2854, 982]. Leave [3155]. Lecture [2808]. ledger [2950]. legacy [1918]. Length [3265, 3375, 2840, 2820, 2587, 2540, 547, 1694, 450, 1119, 2199, 320]. Length-Bounded [2840]. Lengths [3192, 2928, 1167, 3197]. lens [2111]. LEO [3394, 2945]. less [1048, 2026]. Lessons [3115, 2460, 2851]. let [2129]. Leung [597]. Level [3408, 2656, 2932, 2475, 2800, 2907, 1429, 1293, 1915, 1619, 1207, 1237, 1217, 810, 1388, 1067, 754, 410, 1739, 360, 3379, 1728, 7, 351, 772, 1932, 1942, 2240, 1605, 497, 1470, 784, 1172, 1565, 1856, 871, 1410, 1847, 1484, 1663, 1803, 599, 712, 1321, 1460, 803, 1153, 155]. levels [1669, 172, 1889, 478]. Leveraging [2937, 3432, 1773, 3411, 3331, 1859, 3262]. LFU [982]. Liberty [1281]. library [3216, 3224, 781, 122]. Licensed [2924]. life [217, 1638, 277]. life-cycle [217]. Lifetime [3119, 2719, 3094, 2849, 1766, 2683, 840, 769, 1378, 1440]. lifetime-based [1378]. Light [3352, 2417, 2715, 3389, 1621, 1475, 1995]. Light-Tailed [3389]. Light-weight [1621]. Lightened [3288]. Lighting [2660]. Lightweight [2601, 2810, 3397, 2160, 1853]. Like [2702, 2835, 2551, 2789, 2751, 2114, 1580, 1724, 276, 145, 1635, 1670, 1481, 1973]. Likelihood [2540, 687, 1181, 1774, 1782]. Limit [2858, 2750, 1898, 2360, 2110, 2919, 1894]. Limitation [2855]. Limitations [3422, 3418, 992, 903, 538]. Limited [2762, 2787, 685, 757, 1560, 495, 614, 2328, 673, 1896, 677, 2237, 881]. Limits [3052, 2785, 2628, 2645, 2520, 2686, 2902, 1201, 842, 2789, 2803, 2607, 49, 2784, 1807, 806, 1893, 2347, 1894, 2305]. Line [2760, 2547, 3247, 498, 2286, 731, 1611, 356, 2407, 1248, 811]. Linear [3143, 3318, 3182, 3095, 2439, 3048, 1013, 30, 575, 1098, 1354, 1929, 1727]. lineariser [368]. Lingering [2235]. Linguistics [311]. Link [2775, 2477, 2945, 2475, 1686, 1584, 1982, 1450, 1000, 1675, 1470, 2379, 1085, 983, 276, 1572, 682, 2110, 1326]. Link-Level [2475, 1470]. link-state [682]. linkability [2220]. linked [88]. LinkedIn [2675]. linking [16]. Links [3437, 920]. Linux [2693]. LIRS [1183]. List [2479].

List-based [2479]. Listing [3396]. lists [1354]. little [1947]. Live [2630, 1672, 1910, 1671]. lived [1316, 1413]. LiveLab [1909]. lmbench [888]. Load [3297, 2862, 3119, 2618, 2734, 3026, 3196, 1164, 1025, 3076, 2867, 2562, 1864, 3075, 2970, 3447, 517, 614, 2969, 3140, 2612, 3229, 3410, 434, 2623, 2881, 289, 1139, 2651, 2864, 2658, 2617, 2762, 3009, 3139, 3291, 2971, 2002, 3325, 2183, 3138, 1091, 2631, 2863, 2471, 2860, 2855, 2913, 2968, 3038, 3079, 1318, 1863, 75, 1722, 1213, 900, 940, 395, 495, 2310, 2198, 1825, 1483, 840, 2417, 435, 1953, 2031, 2410, 2267, 2195, 740, 880, 2313, 744, 812, 2915, 369]. Load-Balanced [2631, 2863]. Load-Balancing [2562, 2623, 434]. Load-Dependent [3026, 3196]. Load-Optimization [3076]. Load-side [2762]. Loaded [2854, 3140]. loader [74]. loading [16]. Loads [2631, 2863, 2104]. Local [658, 2779, 1454, 2745, 2896, 2966, 3281, 1722, 2052, 2325, 1864, 352, 496, 416, 522, 1845, 1787, 249, 2256, 309, 276, 381, 533, 482, 361, 299, 305, 427, 304, 306]. local-rule [2052]. localities [1276]. Locality [2346, 1446, 2978, 3137, 2516, 3138, 552, 1390, 367, 1871, 1045, 2405, 707, 446, 827, 1312, 403, 2213]. Locality-aware [2516]. Localization [2895, 2901, 2000, 1845, 2404]. Localized [2508, 3319]. Localizing [2517, 2547]. Locally [2690, 339]. Locally-Sufficient [2690]. Locating [2926, 1326, 1216]. Location [3300, 2619, 2984, 2519, 1226, 1138]. Location-Based [2984]. lock [378]. locking [1074, 204, 677]. Loeve [2336]. Log [3227, 268, 1779]. logging [1484, 2244]. logic [1368, 1567]. LoGPC [937]. LogStamp [3227]. Long [3438, 1705, 3293, 3270, 3320, 1844, 1413, 1253, 1239, 1719, 1753, 970, 1229, 1251]. Long-haul [3438]. long-lived [1413]. long-range [1251]. long-range-dependent [970]. long-run [1239]. long-tail [1253]. long-tailed [1229]. Long-term [3293, 1844]. Longest [3111]. Longevity [2719]. Look [3126, 2761, 3329, 2459, 2908, 2982, 3425, 463, 2106, 2225]. Look-Ahead [2982]. lookahead [1224]. Looking [2603]. Lookup [3287, 3396, 2265]. lookups [1103, 913]. loop [683, 659, 996, 489]. loop-based [659]. loops [299, 305]. Loses [3097]. Loss [1793, 3045, 2998, 2525, 2740, 2849, 2529, 1652, 2494, 3336, 1273, 1564, 1014, 1578, 1050, 1679, 1165, 1961, 923, 1606, 1727, 1877, 1048, 1933, 532, 1614, 2201, 1455, 744]. Loss-aware [1652]. loss-less [1048]. loss-load [744]. Losses [3049, 1052, 984, 1528]. Loud [2835]. Lounge [2652]. love [1375]. Low [2687, 2604, 2970, 2605, 3010, 3270, 3320, 2828, 1377, 3006, 2939, 2861, 2860, 2121, 1185, 3166, 2076, 2191, 1828, 986, 1183, 1295, 1955, 2403, 2244, 944]. low-associativity [1828]. low-bandwidth [986]. Low-Complexity [2687, 2861, 2860, 2403]. low-cost [2191]. Low-Delay [2605, 2861]. low-latency [2244]. low-priority [1295]. Low-Rank [3270, 3320, 3166]. Low-Regret [3006]. Low-Voltage [2939]. Lower [2790, 2488, 2487, 3447, 2076, 1965]. lower-bound [1965]. Loyalty [2666]. LPM [3285]. LRFU [1276]. LRU [562, 2634, 1725, 2743, 982, 2990, 2818]. LTE [2429, 2837, 2839]. LTE-Aware [2839]. LTERadar [2839]. LU [292]. LU-LU [292]. lumping [1099]. Lux [3301]. Lyapunov [3250, 2641, 3027].

M [461, 338, 1269, 2295, 598, 1904, 2236, 286, 1128, 588, 3315, 1592, 461, 1691, 2298, 338, 1158, 1269, 384, 671, 198, 856, 2295, 1348, 1904, 2236, 3447, 1597, 1720, 1230, 2066, 286, 1128, 320, 3445, 1188, 3024, 3122, 1792, 3315, 3389]. M-JSQ [2066]. M/Ck/1 [320]. M/G/ [3447, 3122]. M/G/1} [1720, 1592, 1691, 2298, 671, 198, 1348, 3445, 1188, 3024, 3389]. M/G/1-FB [1348].

M/G/1-type [1188]. M/G/1/./EDF [856]. M/G/1/SRPT [1158]. M/G/m [384]. M/G/s} [1597]. M/GI/1 [1230]. M/M [2066]. M/M-JSQ [2066]. M/M/ [3315]. M/M/1 [1128]. M/M/1/SRPT [1269]. M/M/k [1904]. M/M/k/N [2295]. M/M/k/setup [2236]. M/M/m [461, 338, 286]. M/PH/1 [1792]. M68020 [666]. MAC [1900, 1443, 2677, 2407]. MACH [511, 729]. Mache [532]. Machine [3119, 3108, 2758, 2845, 3060, 3102, 3164, 599, 3105, 2960, 3405, 360, 2409, 877, 5, 476, 1513, 1553, 135, 372, 2106, 2338, 1824, 270, 2317, 2341, 845, 758, 155, 1228, 2158, 1064, 314, 566]. machine-based [270]. machine-learning [1228]. machine-to-machine [2106]. Machines [2681, 3347, 2468, 497, 2095, 2183]. Machuel [746]. Macmillan [598]. macro [2155]. macroscopic [2368]. Magma [3154]. magnifying [1957]. mainframe [458]. maintaining [2138]. maintenance [1796, 2097]. Majority [2625]. Majorization [2064, 1141]. make [527]. Making [3045, 3421, 572, 573, 3359, 3006, 1839, 1477]. MalAlert [2959]. Malcolm [3297]. Malicious [3186, 3323, 2878]. Malware [3414, 2583, 2959, 2259]. MAM [2089]. MAMA [1012, 1087, 1157, 1227, 1433, 2876, 3016, 3053, 3156, 3234, 3332, 3439]. Mambo [1279]. Manage [2707, 1402]. managed [916]. Management [3297, 2461, 2944, 103, 2661, 594, 2868, 3377, 2555, 3182, 2608, 2961, 2873, 2797, 2594, 2990, 2725, 3251, 3254, 3036, 2819, 3048, 528, 1024, 1040, 2284, 331, 1839, 1067, 1766, 1676, 773, 2266, 2114, 980, 574, 2186, 175, 1983, 1833, 1827, 66, 1673, 2115, 466, 873, 748, 1836, 1027, 1586, 1876, 1284, 2147, 1486, 204, 1899, 690, 561, 1938, 1701, 749, 1952, 1498, 2387, 2272, 2013, 1133, 988, 224]. Manager [3350, 906]. Managing [49, 1407, 1875, 2815, 1424, 1559, 1692, 880, 1458]. MANET [1432]. MANETs [1867, 2070]. Manipulation [2655]. Mankind [2597]. manufacturing [449, 503, 1653]. Many [2816, 2204, 2552, 1924, 2302, 2086, 1474, 2527]. many-core [1924]. many-flows [1474]. many-server [2302, 2086, 2527]. Many-Sources [2816]. manycore [2290]. MAP [2087, 1439, 2467, 2213, 2056, 1482, 2289]. MAP-modulated [2056]. MAP/G/1 [2289]. Mapping [3391]. MapReduce [2044, 2045, 2372, 2679, 2563, 2292, 1816, 1777, 2101, 2360, 2213, 2464, 2466]. MAPs [1602]. MARC [3370]. March [710]. Margaret [645]. marginal [2308]. Marker [3094]. Market [2700, 2454, 2916, 2560, 2624, 3068, 2701, 2925]. Market-Driven [2454]. Marketing [2489]. Marketplace [3278, 2333]. Marketplaces [2981, 2626]. Markets [3029, 3422, 3237, 2754, 2794, 2740, 2798, 2714, 2806, 2965, 2655, 3068, 3144, 2670, 2354]. marking [1034]. Markov [1371, 992, 2682, 1099, 1613, 3015, 2081, 2287, 1261, 2826, 209, 2236, 1187, 772, 3453, 208, 2262, 1020, 3345, 1522, 3218, 728, 618, 1060, 2526, 1343, 723, 2155, 1188, 1111, 1614, 3027, 2041, 708, 2830, 3005, 2915, 2877, 2919, 2416, 680]. Markov-modulated [2287, 2826, 728, 2919]. Markov-regulated [2416]. Markovian [3224, 1985, 2090, 2234, 1612, 807, 952, 2061, 2062, 3318, 3013, 2087, 1128, 724, 3222, 188, 823, 656]. MaRS [682, 3279]. MART [1685]. MART-aided [1685]. martingale [2301]. martingale-envelope [2301]. martingales [2157]. mass [855]. massive [2128]. massively [632]. Match [3051, 3137]. Matching [3090, 3193, 2591, 3091, 3358, 3237, 3144, 2664, 3317, 2759, 3152, 2294, 1890, 2550, 1709, 3111]. MAthematical [1012, 1087, 1157, 1227, 2876, 3016, 3053, 3156, 3234, 3332, 1237, 1096, 132, 1433, 3439]. mathematics [2054]. MATLAB [2090]. matrices [3166, 1412, 2349, 1411]. Matrix [3008, 3215, 2728, 3338, 3380, 1915, 2089, 2057, 2077, 3337, 1613, 2060, 2065, 1736, 2173, 1291, 1452, 1169, 2425, 1782, 1477]. matrix-analytic [2089]. matrix-exponential [2057, 2060]. matter [2266]. Matters [3093, 1491]. mature [1017]. Max [2891, 3091, 2994]. Max-Min [2994]. Max-Weight [3091]. Maximization [3310, 2574, 2742, 2439, 3145, 2470, 1659, 1399, 2340, 1350, 1899, 2011]. Maximizing [2544, 1468, 3262, 2907, 2468, 1172, 2370]. Maximum [3055, 1181, 1774, 2425, 1256, 2540, 450, 400, 1782]. MaxWeight [2549, 2503]. MCCAbased [1993]. McCahn [746]. McGraw [745]. McGraw-Hill [745]. McKerrow [606]. mCRL [1734]. MDC [2047]. MDP [3342]. MDPs [3244, 2737]. Mean [1592, 3216, 3292, 3160, 3074, 2528, 2784, 2785, 2645, 1898, 2716, 2853, 2886, 2917, 2686, 2046, 3140, 2515, 2658, 2626, 3130, 3291, 3205, 3168, 2717, 2995, 979, 2928, 2610, 2715, 1293, 1316, 350, 1761, 621, 3059, 1252, 1841, 958, 506, 679, 1263, 1596, 514, 761, 1896, 1175, 441, 2239, 743, 262, 369]. Mean-Field [2528, 2785, 2716, 2686, 3205, 2717, 2610, 2715, 2784, 2645, 2046, 3130, 3291, 2928, 1316]. mean-value [506]. MeanField [1737]. Measure [2771, 367, 394, 386, 392, 932, 1296, 1524, 722]. measured [1596]. Measurement [641, 3123, 606, 3069, 3430, 3394, 3428, 3298, 446, 177, 1055, 3124, 2733, 248, 2, 1132, 50, 1450, 123, 480, 124, 1654, 389, 36, 1456, 629, 325, 1018, 2232, 1008, 393, 1710, 1582, 1655, 1633, 236, 327, 266, 1005, 1054, 2106, 2363, 1476, 228, 722, 742, 2364, 799, 1391, 1411, 300]. measurement-based [1710, 722]. Measurements [38, 2931, 2505, 58, 3304, 636, 1007, 184, 216, 1207, 1237, 1181, 2131, 1009, 1253, 291, 928, 1914, 17, 1961, 2138, 1245, 1877, 1117, 914, 277, 82, 1397, 1870, 1412, 2016, 1549, 1329]. Measures [2791, 1221, 507, 1511, 1613, 247, 882, 457, 1437, 1060, 361, 790, 109]. Measuring [1011, 2134, 2184, 336, 454, 2413, 3300, 3392, 1297, 447, 171, 19, 3321, 125, 1979, 1713, 1578, 322, 960, 638, 1446, 1909, 509, 3413]. Mechanism [3261, 2931, 2448, 2439, 3050, 2657, 1040, 1083, 1299, 2165, 1350, 2026, 943, 744, 2428]. Mechanisms [2718, 3191, 2437, 2575, 2720, 2607, 2639, 2332, 2924, 991, 1001, 954, 1311]. Media [1156, 1522, 1146, 990, 797, 1580, 1049, 2350, 999, 1670, 1502, 971]. Mediator [2925]. medium [2142, 1645, 1902, 2887]. Meeting [2203, 69, 1938]. Meets [2500, 2846, 1369, 2142, 2958]. Membership [2599]. Memories [2976, 2843, 2384, 1852, 634, 2099]. Memory [511, 3420, 1927, 794, 2719, 2523, 464, 2586, 3418, 2480, 1065, 2620, 2849, 3278, 2483, 2797, 3304, 2966, 2757, 2992, 945, 973, 685, 528, 44, 1664, 992, 38, 54, 28, 146, 1767, 891, 2241, 1560, 2014, 773, 2388, 551, 980, 531, 471, 879, 633, 536, 2385, 519, 1299, 2176, 1547, 578, 1191, 565, 974, 1929, 341, 818, 1673, 1003, 667, 817, 2172, 1424, 844, 755, 1933, 876, 168, 462, 726, 966, 1121, 770, 1122, 626, 2182, 836, 741, 1136, 787, 880, 916, 2387, 2125, 2013, 1391, 569]. Memory-reference [511]. memory-stalls [2176]. memory-system [818]. MEMS [1037]. MEMS-based [1037]. MemSpy [667]. Memtrade [3278]. Merge [2542, 2690, 3101]. Mergesorts [727]. merging [857]. merit [55, 25, 48]. Mesa [335]. mesh [498, 1552, 1611, 1812, 1599, 1798]. Message [3368, 1432, 3098, 3106, 3226, 843, 756, 771, 937, 740, 588]. message-based [588]. message-passing [756, 771, 937]. Messaging [2586, 1268]. Metaverse [3392]. MetaVRadar [3392]. METE [1938]. Metering [3261, 2939, 1814, 1917]. Method [2862, 2804, 2760, 2902, 2803, 2940, 2715, 343, 3215, 2152, 1979, 1738, 1249, 633, 872, 516, 272, 2085, 1522, 347, 1902, 809, 348, 253, 1267, 94, 550, 148, 1214]. methodologies [176]. Methodology [184, 2848, 2582, 2439, 248, 1716, 379, 218, 1469, 2444, 929, 1065, 284, 972, 508, 96, 102, 1006, 128, 788, 791]. Methods [2730, 2844, 2778, 321, 1185, 2089, 1734, 652, 106, 1638, 536, 624, 1239, 1420, 2065, 208, 212, 368, 703, 154, 89, 763, 2175, 1888]. metric [1826, 1783, 1579, 1882, 2182]. Metrics [2815, 2972, 2250, 244, 229, 317, 865, 227, 313, 1908, 126, 1936, 1838, 230, 1587]. metro [2135]. metro-area [2135]. Metropolis [2127]. metropolitan [1980]. MH [59]. MH-TSS [59]. mHTTP [2406]. Micro [852, 2155, 216, 458, 447, 94, 1800]. Micro-architecture [852, 1800]. micro-processor [94]. micro-processors [447]. Microarchitectural [2420]. Microarchitecture [3398, 1203, 1685, 1430]. microbenchmarks [1340]. microcode [510]. Microgrids [2753, 2228, 2354]. microprocessor [1504, 1963]. microprocessors [1002, 1062]. microscopic [1212]. microsensors [1135]. Microservice [3273]. Microservices [3377, 3350]. Middle [3003]. middleware [1030]. might [2114]. migrating [495]. migration [2191, 1910, 1690, 626, 1459]. migration-based [1690]. MIL [127]. MIL-S-52779 [127]. Millimeter [3277]. Millimeter-Wave [3277]. MIMD [632, 344]. Min [2891, 2994, 639]. MIN-based [639]. Min-Max [2891]. Mine [2953]. Mini [3124, 222, 122]. Mini-Apps [3124]. mini-computer [122]. minicomputer [79, 166]. minicomputer-based [166]. Minimal [3047, 2880, 2929, 3192]. Minimal-Variance [2929, 2880]. Minimization [3239, 1351]. Minimize [3074, 1976]. Minimizing [3055, 862, 2103, 1077, 2820, 2990, 3084, 3267, 3372, 2458, 2468, 2227]. minimum [679]. Mining [3100, 2949, 2343]. minutes [1430]. Miracle [3430]. Mirage [3430]. Mirror [3046]. mirrored [631]. MIS [172]. misconceptions [747]. Misinformation [3005]. Misra [588]. Miss [2633, 2545, 602, 1022, 1725, 715, 620]. Misses [2990]. missing [7]. mission [576, 763]. Mitigate [3427, 1887]. Mitigating [2900, 3426]. Mitigation [2739, 2799, 2505, 2802, 2386, 2768]. Mix [3137, 3]. Mixed [2988, 1270, 2397, 2053, 68]. mixed-exponential [2053]. Mixing [3101, 2233]. mixtures [3337]. ML [3284, 3110]. ML-based [3110]. MLC [2384]. mmWave [3331, 3330]. Mobile [3123, 2618, 2734, 2661, 3107, 2454, 1082, 2608, 3300, 3304, 2837, 3177, 2536, 1216, 1986, 1910, 1325, 1114, 2333, 1866, 1874, 2120, 1386, 2100, 2374, 2255, 1990, 2026, 2375, 1726, 374, 2334, 1960, 2399, 2212, 2684]. mobile-application [2399]. Mobility [2608, 2962, 2813, 3232, 1531, 911, 2389]. Mobility-on-Demand [2813]. Möbius [1743, 3114]. modal [3355]. mode [3202, 3371]. Model [3008, 1369, 3221, 2696, 2590, 1004, 2850, 3406, 2523, 3219, 2673, 3327, 2592, 3400, 3129, 2691, 2453, 2585, 3381, 1117, 2838, 1956, 3205, 3267, 3372, 2717, 3254, 2494, 2476, 2294, 3335, 2056, 191, 1166, 182, 2075, 90, 2164, 760, 140, 2207, 409, 617, 1489, 28, 1505, 2531, 1890, 463, 1194, 1531, 1861, 1669, 621, 408, 1750, 1934, 613, 2796, 30, 257, 2059, 1315, 259, 3059, 720, 1841, 2385, 1485, 792, 32, 351, 1445, 2061, 1431, 522, 1923, 1605, 1066, 1733, 702, 1643, 1787, 2053, 1370, 1639, 1747, 721, 864, 1724, 2123, 234, 238, 2086, 1124, 1401, 1130, 2173]. model [771, 855, 2424, 78, 1060, 549, 2088, 2885, 356, 420, 827, 1440, 154, 34, 12, 1882, 89, 930, 1993, 1726, 2249, 1406, 2239, 1312, 567, 620]. Model-based [1004, 1117, 3254, 771]. Model-driven [1956, 1861]. Modeling [3457, 285, 345, 641, 1687, 729, 2429, 399, 340, 938, 977, 2579, 422, 834, 1937, 2758, 837, 2200, 2770, 3214, 302, 598, 658, 655, 931, 1253, 1037, 1825, 664, 1080, 1598, 2176, 2493, 1545, 1235, 2763, 2533, 2139, 1386, 3361, 431, 476, 953, 3114, 1788, 1813, 1548, 586, 2677, 1775, 3001, 2329, 785, 2936, 1236, 1255, 2627, 2339, 1528, 1012, 1087, 1157, 1184, 1227, 1433, 2876, 3016, 3053, 3156, 3234, 3332, 3439, 3159, 1824, 915, 2694, 3109, 1259, 3211, 2428, 2878, 155, 1250, 1096, 1209, 379, 1619, 1742, 994, 1254, 1147, 401, 747, 1740, 2091, 1444, 2226]. modeling [1781, 1743, 424, 390, 425, 854, 1789, 1191, 1332, 497, 1748, 1494, 144, 2180, 1544, 1685, 2238, 1582, 3223, 309, 295, 1177, 2192, 579, 937, 2172, 235, 246, 1103, 540, 759, 908, 308, 2395, 719, 1111, 474, 94, 1135, 1959, 1884, 1660, 2215, 255, 1412, 287, 1021, 2393, 1244, 388, 733, 708, 799, 1238, 1228, 1797, 1865, 1092, 1751]. Modelling [1163, 2217, 2801, 1641, 3031, 640, 596, 2667, 354, 949, 947, 502, 339, 663, 417, 3225, 95, 143, 1434, 627, 661, 652, 430, 1925, 1372, 131, 1736, 256, 145, 2012, 107, 742, 492]. Models [3292, 3175, 3072, 2795, 2322, 2851, 3112, 3334, 2931, 2581, 3260, 3296, 2820, 3295, 3095, 3065, 2815, 3217, 156, 2778, 3205, 2995, 2800, 2960, 2610, 1439, 2948, 477, 1744, 2935, 2156, 54, 334, 187, 342, 383, 2550, 861, 1730, 1706, 1985, 1613, 1056, 437, 2080, 1307, 1830, 164, 108, 209, 800, 179, 445, 2046, 273, 272, 149, 1345, 1965, 1257, 192, 1443, 2218, 2391, 276, 1285, 1615, 684, 1461, 3275, 281, 261, 348, 2054, 366, 909, 413, 912, 82, 765, 723, 2155, 101, 137, 264, 2246, 1310, 200, 1614, 1657, 2070, 84, 1174, 2041, 786, 1666, 1840, 263]. models [132, 1628, 1331, 2915, 346, 1329, 262, 774, 1344, 680, 385]. Modem [2850, 1010, 1106]. Moderate [3447]. Modern [3399, 2696, 2613, 2718, 2720, 3255, 3283, 2394, 1974, 1062, 447, 777]. Modified [3272, 3145]. Modular [296]. Modulated [2826, 2056, 1020, 728, 2919, 2287]. modulating [2202]. module [275]. modules [834]. Molloy [598]. Moment [2058, 3041, 885, 441, 2305]. moment/strip [1594]. Moments [3027, 1602, 1597, 1969, 1447, 441, 1395]. moments-based [1969]. Monetization [2489]. Money [2735]. Monit [475]. Monitor [2909, 601, 74, 18, 10, 19, 267]. Monitoring [1450, 3108, 3302, 3039, 2584, 39, 2776, 3365, 3109, 509, 1656, 1780, 1664, 2286, 1702, 1421, 1661, 181, 475, 1453, 1970, 1245, 851, 1715, 1573, 1657, 2366, 1621]. monitors [490, 62, 23]. monopolist [2332]. Monotone [3145, 1999, 525]. Monotonic [2308]. Monotonicity [2764]. Monte [3013, 1522]. Moon [2633]. Moore [2433]. Morgan [750, 657]. Moshe [607, 705]. most [3209, 279]. Motion [3331, 2069]. motivated [45]. Mounted [3360]. move [6]. Movement [3280, 881]. Movers [2378]. Moving [2749]. MP [637]. MPEG [925, 798, 886]. MPEG-coded [798]. MPI [1926, 2177, 1931]. MPI/CUDA [1926]. MPI/OpenMP [1931]. MPLS [1134]. MPTCP [2707]. MPX [2852]. MQTT [2942]. mRSC [2993]. MSR [2018]. MSR-based [2018]. Mt [2640]. Mt/Gt/1 [2640]. MTOOL [633]. MTS [52]. Much [2956, 2675, 2293, 2116]. Multi [3297, 2993, 2472, 2531, 3388, 2537, 3302, 3276, 2599, 2698, 2504, 2634, 1231, 3178, 1286, 3355, 2761, 2700, 2835, 3400, 2704, 2153, 2406, 2754, 2794, 2900, 466, 2798, 983, 2438, 2756, 2990, 3167, 3269, 2908, 3068, 3027, 2465, 2971, 1460, 2535, 3311, 2638, 2911, 3121, 3317, 2829, 3319, 2511, 2913, 2906, 2096, 2382, 2121, 2397, 1222, 2284, 2078, 1762, 1319, 1680, 79, 2784, 834, 856, 1388, 2082, 2024, 1781, 1469, 1851, 1059, 2063, 458, 2949, 772, 1113, 2230, 2365, 885, 678, 201, 2356, 2123, 2297, 1351, 1859, 809, 1624, 1343, 2263, 761, 147, 1417, 1896, 2196, 1764, 2530, 1199, 1364, 1105, 878, 1406]. multi [2183, 1723, 1846, 417, 2927, 3197]. Multi-access [3276]. Multi-Agent [3269, 3311, 3319, 3297]. Multi-armed [2829, 2906, 2082]. multi-banked [2121]. multi-cell [1319, 1680]. multi-channel [1762, 1723, 2927]. Multi-class [2531, 2096, 856, 1059, 2230, 201, 2356, 809, 761, 3197]. multi-computer [147]. Multi-Core [2704, 2284]. multi-cryptocurrency [2949]. Multi-Dimensional [3027, 2993, 3388, 1286, 2913, 1859, 2263]. multi-dimensionally [1343]. Multi-disk [466]. Multi-Feature [3167]. Multi-flow [2990]. Multi-hop [2698, 2153, 2784, 1469, 1113, 678, 1764, 2530, 1199, 1846]. Multi-layer [983, 2365]. Multi-level [1460, 1388, 772]. Multi-LRU [2634]. multi-micro [458]. Multi-modal [3355]. Multi-Model [3400]. multi-nomial [2397]. Multi-path [1231, 1105]. Multi-Pipe [3302]. multi-process [878]. Multi-Processor [2756]. multi-queue [1417]. Multi-Rate [2535]. multi-regional [2196]. Multi-Resource [2472, 2900, 3068, 2638, 2465, 2123]. Multi-Server [3178, 2971, 3121, 2078, 2024, 1417, 1896]. multi-service [885]. Multi-Set [2599]. multi-shop [2382]. Multi-source [2406]. multi-state [2297]. multi-system [417]. Multi-Task [2911]. Multi-Tenant [2537, 2761, 2700, 2908, 2504, 2835, 2511]. multi-tier [1222, 1781, 1851, 1406]. multi-tiered [1351, 1624, 1364]. Multi-timescale [2754, 2794, 2798]. Multi-Unit [2438]. multi-user [79]. multi-variate [2063]. Multi-way [3317]. multiaccess [298]. multicarrier [2419]. Multicast [1339, 1623, 1032, 1359, 534, 1578, 932, 1189, 875, 1068, 1033, 1044, 1812, 775, 910, 985, 1197, 1043]. Multicasting [2047]. multichain [516]. multichannel [557, 1856]. Multiclass [2786, 411, 2465, 2805, 1016, 1480, 1730, 1349, 978, 414]. Multicommodity [3188, 1505]. multicomputer [795]. multicomputers [784, 530]. multicore [2156, 2290, 1908, 1673, 2257, 2048, 2144, 2245, 2181, 1931, 1786]. multicores [1938, 1886, 1939]. Multicut [2840]. MultiDefender [2318]. Multidimensional [3073, 3028, 176, 1903]. Multigraph [2502]. multihomed [1990]. Multihop [2477, 3403, 1867, 1811, 1623, 821, 1416, 2357]. Multilayer [2857]. multilevel [3089]. Multimedia [895, 1328, 2274, 737, 839, 899, 936, 896, 1036, 1122, 892, 894]. multiparty [1698]. Multipath [2628, 2710, 2248, 3082, 1942, 2406, 501, 2273, 1887]. multipathing [191]. multiphase [778]. multiplayer [1698]. Multiple [2779, 2450, 1179, 341, 2485, 3272, 2747, 2547, 3147, 3195, 2695, 3079, 705, 1221, 938, 977, 180, 408, 1773, 398, 440, 662, 1015, 3384, 2372, 875, 212, 1168, 1020, 1903, 489, 2392, 282, 387, 1644, 2435, 515, 2073, 2347, 758, 1477, 607]. multiple-chain [398]. Multiple-choice [2485]. multiple-loop [489]. Multiple-Processor [3272]. Multiple-queue [1179]. multiple-server [440, 1015]. multiple-set [1773]. multiple-subscripted [662]. multiplexers [805, 776, 318]. multiplexing [1094, 1120, 438, 971]. Multiprocessor [691, 511, 507, 31, 146, 814, 716, 473, 551, 531, 536, 196, 336, 565, 669, 1853, 506, 570, 431, 1003, 502, 78, 586, 1559, 867, 835, 626, 733, 429, 849, 604]. multiprocessor-based [1003]. multiprocessors [528, 274, 794, 621, 1766, 804, 725, 2036, 972, 1690, 1940, 490, 844, 836, 741, 1136, 2125, 639, 569, 625]. Multiprogrammed [964, 939, 32, 1335, 570, 499, 82, 53, 726, 263]. Multiprogramming [27, 342, 383, 31, 613, 571, 410, 207, 58]. multiprogramming-multiprocessor [31]. multirate [821]. Multiresource [3448]. Multiscale [3292]. Multiserver [2879, 2912, 3334, 3327, 3357, 3349, 3370, 3381, 3064, 3168, 1534, 436, 1542, 2009]. Multiserver-Job [3349, 3370, 3334, 3327]. multistability [1385]. multistage [470, 1654, 584, 501, 619, 567]. Multitask [3272]. multitasking [702]. multithreaded [928, 1559, 1886, 2393]. multithreading [842, 1408, 1186]. multitier [1820, 1834]. MultiTrack [1804]. multiuser [574]. Multivariate [3112, 2077]. multiversion [443]. must [1527]. mutual [1705]. MVA [589]. MVS [167, 521, 165, 170]. MVS/XA [521]. My [3413, 3123, 5, 3110]. Myrinet [989]. Myrinet-based [989]. myth [1337]. Myths [1287].

N [2295, 487, 2080]. N/1/F [487]. Nagle [1028]. name [2265]. Named [2721, 2265]. NAND [2384, 2405, 3094, 2849, 2706]. Nano [2455]. NAS [1931]. NASLU [1926]. Nassi [114]. national [581]. native [1390]. Natural [3041]. naughty [934]. Nautilus [3437]. NBS [248]. NDT7 [3298]. Neal [592]. Near [3279, 3433, 3024, 2979, 3403, 186, 2513, 3417, 575]. Near-complete [186]. near-linear [575]. Near-Optimal [3279, 3433, 3024, 3403, 2513, 3417]. nearest [496]. Nearly [2751, 3122, 1393, 1893]. Necessary [2588, 2968]. need [3017, 2186, 355]. Negative [3395]. negatives [2129]. negotiating [2444]. Nehalem [1998]. Neighbor [2384, 1770, 496, 1973]. Neighbor-cell [2384]. Neighbor-specific [1770]. Nelson [592]. Nesterovian [2605]. nesting [392]. Net [3097, 829, 3223]. NetBSD [888]. NetDiffusion [3429]. NetEcon [2436]. NetFlow [1451]. NetQuest [1476]. NetReplay [1817]. Nets [948, 2679, 949, 3177, 697, 762, 837, 952, 950, 954, 823, 429, 1746, 1748, 947]. netted [379]. Network [2447, 2568, 3052, 2682, 2731, 1982, 1182, 2489, 2769, 480, 1399, 3302, 3071, 2322, 3286, 2572, 2732, 2554, 2749, 2666, 3112, 2931, 2499, 2775, 3069, 3394, 3429, 3328, 3246, 2541, 2533, 2697, 2840, 3258, 1648, 2820, 3428, 2605, 2453, 2609, 2943, 3392, 3298, 2958, 2119, 2972, 2959, 2619, 1763, 2569, 2888, 3252, 2598, 2629, 2909, 3125, 3200, 3183, 287, 1945, 3148, 3236, 303, 2892, 385, 2209, 2776, 2381, 3262, 3081, 2723, 2540, 2755, 2846, 2861, 3109, 1682, 799, 1007, 2675, 2501, 3037, 2841, 3281, 3427, 2733, 3383, 3353, 248, 2093, 2159, 1975, 1338, 1817, 1755, 1716, 1309, 1040]. network [825, 1198, 870, 959, 927, 1534, 2040, 2004, 1540, 846, 187, 301, 275, 1093, 1123, 1014, 1004, 1449, 1223, 1695, 408, 1421, 1405, 2269, 1181, 437, 1676, 1359, 768, 661, 1684, 2131, 1009, 1585, 652, 1315, 1270, 1490, 1603, 37, 396, 893, 2001, 179, 2170, 2949, 1001, 1932, 1431, 1550, 2358, 2219, 522, 2232, 1717, 1772, 2276, 1756, 1872, 1942, 2413, 1441, 1845, 2153, 149, 1068, 193, 1290, 1478, 2139, 2319, 1454, 1211, 1970, 1810, 192, 1245, 309, 276, 3257, 855, 544, 1284, 1402, 2434, 1847, 937, 533, 535, 281, 884, 1989, 261, 348, 681, 1340, 2885, 2263, 1491, 482]. network [2253, 82, 1551, 1116, 1294, 361, 1440, 2395, 1652, 34, 905, 2427, 1397, 1337, 1408, 1111, 1779, 2225, 2375, 2055, 1657, 1321, 2070, 1476, 1174, 985, 2338, 2201, 2910, 263, 1891, 2020, 1084, 2243, 1998, 2408, 270, 1709, 1794, 2341, 1960, 2027, 1533, 1366, 1894, 2163, 1683, 262, 1326, 1031, 304, 306]. Network-assisted [2533]. network-attached [893, 855]. Network-Caching [3052]. Network-design [2381]. network-on-chip [1932]. network-performance [2232]. network-wide [1550, 1756, 1810]. Networked [2686, 2740, 3314, 3319, 3366]. Networking [3408, 2810, 2721, 2793, 2871, 746, 3456, 2475, 849, 3011, 738, 1710, 2265]. networkload [765]. Networks [2442, 3279, 2862, 382, 2823, 1793, 2946, 2980, 2809, 3090, 2816, 3026, 2528, 2785, 2914, 1400, 2436, 3076, 2661, 2698, 2642, 2662, 2956, 643, 644, 658, 3149, 3316, 3170, 2644, 2473, 2634, 3435, 2522, 2964, 3060, 2505, 2894, 3260, 2596, 3358, 2857, 2477, 2608, 2962, 3351, 2665, 3087, 3186, 3054, 2803, 2565, 3218, 2508, 654, 2594, 2729, 3348, 2677, 2765, 3393, 3039, 2942, 3042, 3230, 2669, 3205, 3451, 2705, 2498, 3256, 3222, 2909, 3364, 3199, 3183, 2837, 3043, 3324, 3159, 701, 2926, 2957, 3365, 3403, 2622, 2939, 2727, 3037, 2458, 2588, 2858, 2476, 2833, 3153, 3330, 370, 735, 294]. networks [2003, 2032, 1212, 1687, 1999, 919, 1842, 340, 3232, 2412, 1869, 3231, 1016, 1564, 1922, 1762, 1254, 1995, 498, 1805, 2233, 470, 1616, 2426, 2369, 1482, 1623, 1480, 1646, 2079, 2286, 2784, 1609, 178, 1349, 407, 1405, 1986, 1013, 398, 933, 1734, 1711, 440, 1713, 2134, 1731, 2148, 2683, 1844, 2400, 2287, 2320, 302, 1906, 349, 1535, 1161, 1996, 2368, 2140, 1260, 2261, 1413, 491, 1469, 584, 352, 800, 1889, 2293, 1518, 536, 821, 1325, 2162, 1510, 1258, 2299, 1114, 516, 1113, 2340, 1866, 1678, 1442, 1392, 1942, 1679, 1441, 1305, 1552, 1874, 885, 1165, 1470, 501, 1783]. networks [368, 411, 1387, 1494, 678, 288, 1264, 1303, 1802, 341, 1555, 1752, 2021, 1378, 1360, 2297, 619, 1257, 1565, 1856, 2318, 1606, 489, 1611, 1206, 1265, 1302, 1263, 1836, 976, 2254, 924, 693, 728, 381, 618, 684, 2377, 1415, 296, 1468, 1727, 1520, 1854, 724, 1990, 2370, 1715, 999, 2026, 911, 1635, 2208, 500, 514, 358, 2097, 908, 1416, 2011, 2010, 1881, 1323, 1440, 867, 1038, 459, 1978, 1870, 282, 387, 1135, 1984, 2398, 1857, 2102, 1911, 1636, 1764, 572, 573, 1909, 300, 2418, 2530, 1404, 1726, 1199, 2128, 2435, 848, 2277, 377, 469, 1377, 1313, 1428, 1455, 2209]. networks [2685, 427, 1799, 1651, 186, 860, 979, 585, 1499, 1626, 1634, 1516, 583, 538, 743, 400, 2357, 1304, 1599, 2317, 2389, 290, 2315, 567, 1823, 881, 1723, 1556, 1846, 2327, 1574, 1798, 148, 260, 1973, 1797, 1865, 2430, 2927, 1751, 1901, 1855]. Neumann [1164]. Neural [2956, 3260, 3230, 3177, 2846, 3417]. Neural-Enhanced [3417]. neutral [2161, 2283, 1920]. Neutrality [3097, 2322, 3092, 1683, 2671]. neutralize [2112]. Never [2633]. Newer [2482]. Newreno [1199]. News [2491].

Newsvendor [2164]. Next

[2946, 2156, 917]. Next-Generation [2946]. NFA [2132]. NFS [976]. NFSv4.1 [2482]. NFT [3430]. NFV [3247]. NG [3080]. NHFSSTONES [665]. nice [934]. Nicholas [644]. Nicolas [701]. No [3312, 3155, 1088, 532]. no-loss [532]. No-regret [3312]. NoC [1675]. NoC-based [1675]. Node [1760, 2838, 3232, 2050, 2423, 1378, 1415, 2434, 1304, 2027, 2018]. Nodes [2563, 3186, 2005, 2530, 1855]. Noise [3286, 2508, 2363, 1594, 1993]. Noisy [2599, 2967, 3253]. nomial [2397]. Non [3388, 1986, 3379, 2704, 3376, 2671, 3013, 3049, 2728, 3450, 2834, 2366, 2360, 823, 2695, 2825, 2022, 1489, 3224, 2241, 2161, 952, 1898, 1100, 1189, 207, 1095, 1565, 1443, 1169, 1136, 1214]. Non- [3376, 2671]. Non-additive [2695]. Non-asymptotic [1986]. non-blocking [1136]. Non-complete [3388]. Non-Convex [2825, 3049, 2728, 1565]. non-cooperative [1443]. non-homogeneous [2022]. Non-intrusive [3379, 2366]. non-inversive [1169]. Non-Markovian [3013, 823, 3224, 952]. non-neutral [2161]. non-Poisson [1489]. Non-Preemptive [2834, 207]. non-regenerative [1100]. non-responsive [1214]. non-sequential [1189]. non-smooth [1898]. Non-stationary [3450, 1095]. Non-Volatile [2704, 2241]. Non-work-conserving [2360]. nonconcave [1399]. Noncooperative [1674]. Nonexponential [400]. nonGaussian [970]. nonillion [2346]. noniterative [516]. Nonlinear [2611]. Nonlinearity [2811]. Nonnegative [3245]. nonneutral [2162]. NonNeutrality [2453]. nonpreemptive [451]. nonreentrant [345]. Nonsmooth [3404]. nonstationarity [1112]. NonStop [575]. nonuniform [834]. North [647]. North-Holland [647]. NOS [283]. NOS/BE [283]. NoSQL [3031, 2280].

Nostrand [751]. note [86, 55, 1225, 412, 363, 134, 26, 1615, 48, 20]. notes [10]. Nothing [3188]. Novel [3026, 2902, 932, 1655, 835, 1920, 1493]. ns} [912]. NSFNET [581]. NTP [3391]. NUCA [1690]. Nudge [3120, 3326, 2222, 3446]. Nudge- [3446]. NUMA [794, 669, 2179, 625]. Number [2964, 205, 2773, 2109, 1394, 2227]. numbers [1588, 240]. Numerical [2080, 208, 2087, 282, 387, 3215, 3243, 339, 414]. NVIDIA [3086, 3191]. NVMe [2524].

O [642, 656, 2836, 3416, 340, 938, 977, 422, 713, 720, 796, 1809, 100, 901, 256, 674, 918, 3426, 2280, 194, 1153, 1228, 1429]. O-RAN [3416]. OASIS [152]. Obfuscating [3155]. Obfuscation [2490, 2600]. Object [2817, 1568, 832, 899, 1031]. object-oriented [1568]. Objectives [2472, 3359, 609, 73]. Objects [2594, 892, 1156]. Oblivious [1574, 1309, 2136]. Observations [1451, 2829, 2906, 1425, 2369, 2347]. Observe [2906]. observed [1619]. Obtaining [1137, 580, 590]. OCCAM [2888]. Occasional [3159]. Occupancy [2920, 268, 1360]. Occupancy-Centric [2920]. occupied [2227]. occurs [1527]. Octave [3218]. ODE [2995]. ODE-based [2995]. off [1949, 1252, 1728, 1057, 1611, 2776, 2622, 1259, 1162]. off-line [1611]. Offering [2701]. Offloading [3107, 2536]. Offs [2824, 1672, 2924, 2014, 3158, 1336, 2936, 2877, 2349]. offset [1693]. old [1345]. On-demand [2616, 1339, 1670]. On-Device [3400]. On-line [811, 2286, 731, 356, 1248]. on-off [1728, 1162]. on/off [1252, 1259]. Once [2922]. One [2832, 2822, 1450, 1078, 2151, 1578, 1335, 2319, 1693, 2885]. one-dimensional [2885]. one-way [2151, 1578, 1693]. ones [3275]. ongoing [1246]. OnLine [3238, 2981, 3151, 3309, 3045, 3305, 3052, 3090, 3119, 3386, 2484, 2603, 3441, 2989, 3030, 1144, 1047, 2890, 3018, 2409, 2521, 3263, 2931, 3440, 1068, 3409, 3410, 3182, 3118, 2206, 2983, 3049, 2486, 2633, 3174, 3181, 2228, 3253, 2827, 2283, 2714, 2267, 3333, 3268, 3269, 3313, 3252, 3378, 2982, 3183, 3147, 3307, 3019, 3050, 3242, 3227, 2664, 3006, 2830, 3046, 2825, 2819, 3048, 3415, 3314, 2470, 2519, 2558, 2726, 1711, 1167, 1598, 2358, 2328, 1515, 1543, 1899, 2246, 2352, 1823, 1464, 2343]. Only [3093, 2056, 2057, 2058, 2052, 2059, 2060, 2061, 2062, 2063, 2064, 2065, 2066, 2053, 2067, 2068, 2054, 2069, 2055, 2070, 2016, 2071, 2072, 2073]. Ookla [3298]. OP2 [1924, 2172]. OPEDo [1744]. Open [1089, 1349, 1864]. OpenAirInterface [1716]. Opening [3115]. OpenMP [1931]. operands [316]. Operate [3061]. operated [1298]. Operating [888, 611, 601, 959, 222, 613, 624, 1001, 1605, 769, 1853, 510, 535, 1484, 447, 253, 917, 152]. Operation [2718, 1417, 509]. Operational [291, 346, 467, 2286, 1407, 280, 319, 284, 1970, 2557]. Operations [3175, 519]. Operator [2678, 3114]. operators [317]. Opinion [2625]. OPNET [2006]. opportunistic [1949, 2026, 1956, 1459]. Opportunities [2655, 1121]. OPSS [1616]. Opteron [1934]. optical [1262, 415, 924, 2110]. Optics [3057]. Optimal [3279, 1791, 3206, 2194, 1534, 1373, 3022, 2817, 1564, 328, 808, 2005, 2531, 1623, 2678, 199, 2488, 1359, 3244, 2826, 2646, 1765, 2562, 2187, 3342, 233, 3285, 3327, 3357, 1361, 3377, 3433, 1678, 1392, 1874, 2402, 288, 286, 2754, 2794, 3410, 3158, 3186, 3401, 3424, 1606, 2654, 2798, 3095, 3065, 2587, 2467, 2688, 2864, 438, 1785, 778, 3347, 2526, 3039, 600, 3084, 3024, 3064, 3122, 3162, 978, 2102, 3068, 1657, 1174, 3019, 515, 3403, 3266, 1952, 2513, 2727, 3417, 3005, 3138, 1850, 3390, 2825, 2819, 3240, 3389, 1973, 2726, 3338, 3380, 2860, 3038, 3079, 1855, 1949, 1352, 857, 1262, 1309, 1672, 1150, 1118]. optimal [2261, 2307, 2281, 1720, 29, 2111, 1441, 1472, 2410, 2391, 1905, 1870, 859, 1883, 2247, 715, 51, 2296, 1499, 1599, 2213, 971, 1493]. Optimality [1843, 2856, 3446, 2596, 3173, 2736, 2546, 2792, 2855, 2913, 2968, 1762, 2550, 1354, 2432, 2431, 1790, 1533, 2772, 1962]. Optimally [2827, 2747]. Optimisation [3031]. Optimistic [529, 3312, 910, 527, 784, 418, 2183, 1414]. Optimization [2823, 3151, 3029, 3057, 2613, 2484, 2603, 2791, 1400, 205, 2989, 3030, 3076, 3244, 2844, 2506, 2891, 3169, 2744, 2890, 3018, 2724, 3096, 3182, 2983, 3049, 2605, 3181, 2745, 1401, 2945, 2714, 3333, 3313, 2888, 3378, 2982, 2940, 2861, 3048, 1464, 1592, 1250, 1744, 44, 1632, 1686, 2415, 1731, 2295, 2358, 2238, 2206, 2292, 2216, 2396, 1878, 1956, 1777, 372, 89, 2403, 1404, 1205, 1435, 2020, 523, 812, 1366, 2430]. Optimizations [2728, 1121]. Optimize [2963, 2466, 1447, 1870]. Optimized [2921, 3243, 3412, 1436, 2143]. Optimizing [2988, 2284, 2656, 2335, 2595, 1501, 3411, 3444, 2480, 3449, 2173, 1557, 1323, 2910, 2776, 744, 2699, 938, 977, 1758]. Option [2966, 2053]. Options [2450]. Oracle [443, 1829]. oracles [2403]. oranges [5]. order [3335, 2805, 1994, 2060, 2262, 2410, 1173, 2012, 2336]. ordering [1150, 176, 1095]. organizers [3106, 3226]. organizing [1805, 2348]. oriented [2096, 1730, 1004, 621, 1568, 2231, 1031]. Origin [929]. origins [1337]. ORIS [3224]. orphan [1077]. ORSuite [3175]. OS-1100’ [419]. OSNs [2093, 2601]. OSPF [1043]. ossifying [1975]. other [2093, 200]. our [15, 1710]. out-of-band [2366]. Out-of-core [902]. out-of-the-box [2366]. Outage [2311]. Outage-capacity [2311]. outcomes [3455]. Outlier [2791]. output [629, 141, 1312]. outsourcing [1137]. Outward [2729]. Over-the-Top [2674]. overall [94]. Overallocation [193]. Overcoming [3270, 3320, 2685, 2855]. Overcommitment [2692]. Overflow [3419, 1514, 1117]. Overhead [3180, 3445, 1869, 1572, 200, 944]. overlapped [1471]. overlapping [2292]. Overlaps [2634]. overlay [1616, 1442, 1522, 1804, 1636, 1377]. overlays [1643, 2221]. Overload [2621, 1158, 730, 1147, 202, 2930]. overloading [2422]. overruns [207]. overview [1716, 2077, 650, 3223, 1030, 1599]. Overwhelming [2508]. Own [3299].

P [656, 1925]. P.A.P.A. [324]. P1400 [58]. P2P [2029, 1434, 1643, 1787, 1494, 1303, 1265, 1302, 1226, 1788, 2028, 1323, 1804, 1754, 2513, 1913]. PaaS [2580]. Pace [3299, 1107]. pack [147]. package [334, 2091, 3218, 722]. Packard [904]. Packet [1382, 1492, 2974, 459, 3267, 3372, 3403, 294, 984, 2205, 834, 1981, 1056, 1078, 1567, 1594, 1050, 1119, 2138, 1774, 1859, 1758, 502, 1847, 1424, 914, 1294, 2363, 2362, 1528, 1614, 848, 194, 2110]. packet-counting [1594]. packet-loss [1050]. packet-switched [1847, 848]. Packets [2735, 1954, 1829]. Packing [3045, 2521, 3433, 3306, 2653, 2494, 1013, 2692, 1167, 2432, 2227]. Page [3419, 2618, 2734, 2594, 277, 3009, 2907, 2614, 879, 330, 669, 981, 2390, 258, 513]. Page-aware [2594]. Page-Level [2907]. PageRank [2985]. pages [992]. paging [86, 651, 331, 30, 636, 2260, 269]. Pair [2548]. Pair-Approximation [2548]. Pairwise [2512]. PAM [516]. PANDAS [2772]. Panel [226, 3421, 822]. panelists [304]. Paper [2015]. Paradigm [2990, 2071]. Paradox [3158]. Paragon [838]. Parallel [697, 2856, 3022, 3074, 488, 2287, 963, 2493, 2268, 1929, 901, 2691, 3083, 724, 911, 2621, 462, 964, 456, 538, 2553, 2911, 3127, 3240, 2511, 889, 2294, 718, 477, 559, 1873, 825, 838, 2005, 692, 967, 1808, 757, 199, 902, 464, 773, 3058, 841, 629, 632, 548, 847, 1361, 624, 430, 578, 1640, 698, 497, 475, 2541, 863, 2066, 1179, 739, 499, 476, 851, 156, 756, 817, 580, 590, 771, 371, 468, 549, 540, 776, 1895, 755, 339, 876, 717, 918, 910, 978, 726, 835, 1220, 770, 816, 968, 1184, 1244, 786, 2183, 788, 791, 2414, 758]. parallel [727, 1931, 639, 1228]. parallel-architecture [692]. parallel-read [199]. Parallelism [2815, 2978, 1859, 413, 541, 1887]. Parallelizable [3203, 686]. parallelized [833]. ParalleX [1935]. Parameter [366, 1329, 179, 1366]. parameterize [1706]. parameterized [250, 1285]. parameters [1130, 1364, 228, 564, 811, 1391]. Parametric [2567, 309]. Pareto [1592]. PARMA [1944]. Parsimonious [2776]. Parsing [3227]. Part [647]. Partial [2601, 2775, 2025, 2130, 2005, 1099, 1441, 2153, 1946]. Partially [1471, 508]. Participant [71]. participants [1958]. particle [2306]. Partition [2691]. partitioned [1569]. Partitioning [3383, 568, 1325, 1097, 1940, 2212]. Partitions [2895, 2901]. Party [3299, 3133, 985]. Pascal [547]. Passage [1187, 2053]. passing [756, 771, 937]. passive [1453, 1245]. Past [2603, 1462, 2262, 251]. patches [1874]. Patching [2694, 1132, 1026]. Path [3386, 1661, 3412, 3284, 2707, 3300, 2017, 3070, 3185, 3125, 1867, 340, 1134, 1997, 1231, 100, 2153, 1070, 2273, 1410, 1105, 2179, 1334, 1326]. Path-Aware [3386]. path-based [2179]. Path-quality [1661]. Path-vector [2017]. pathchar [1000]. pathologies [2124]. Paths [2707, 1134, 2184]. pathways [546]. patience [2302]. patient [2071]. pattern [619, 41]. Patterns [2681, 464, 330, 454, 1756, 707, 2263, 1879, 968]. Paul [658]. Pause [3410]. PAWS [1071]. PAWS-2000 [1071]. Payment [3451]. Payments [2578]. PBS [1562]. PC [989, 1180]. PCA [1951, 1554, 1806]. PCA-based [1806]. pClock [1546]. PCMARK(R)05 [1644]. PCs [1035]. PDE [2862]. PEACH [3294]. Peak [2452, 2256, 580, 590]. peaks [2252]. PEDCAD [832]. PEEK [389]. Peer [2579, 3271, 1502, 3324, 2573, 1352, 1616, 1659, 1341, 1379, 1507, 1658, 1378, 1671, 1206, 1813, 2028, 1415, 1858, 1508, 1523, 1854, 2196, 2033, 1428, 1304, 1521, 1973]. peer-assisted [1671, 1858]. peer-supported [2196]. Peer-to-Peer [2579, 3324, 1502, 1616, 1659, 1341, 1379, 1507, 1658, 1378, 1206, 1813, 1415, 1508, 1523, 2033, 1428, 1304, 1521, 1973]. Peering [2167, 3142, 2444]. PeerPressure [1324]. PEMA [3350]. Penrose [2433]. pension [2084]. people [2220]. PEPA [1745]. Per-Flow [2994, 1299, 1655, 2363]. per-hop [2153]. Per-Link [2477]. Per-queue [1264]. per-session [678]. perceived [1486]. percentiles [958]. Perfect [3083, 2778, 1516, 1999]. PerfNA [3246]. performability [507]. Performance [2517, 685, 641, 2098, 1222, 3141, 2461, 2946, 3057, 1016, 1147, 2472, 2859, 1645, 1002, 3212, 274, 527, 3193, 2730, 2801, 1083, 631, 2758, 630, 407, 2523, 74, 3219, 473, 3286, 9, 265, 197, 3031, 2006, 1700, 1828, 1781, 349, 640, 824, 595, 594, 596, 598, 643, 604, 607, 2932, 647, 629, 2504, 2140, 1743, 471, 2963, 566, 157, 2679, 1924, 1638, 406, 460, 292, 196, 546, 622, 1080, 3000, 430, 1367, 425, 2865, 2581, 2969, 3343, 3373, 3069, 3107, 1113, 1833, 2704, 2777, 928, 1191, 698, 565, 2615, 609, 362, 3296, 2480, 3246, 1305, 1071, 3094, 501, 2580, 1085]. Performance [183, 433, 1082, 2582, 3272, 619, 949, 3428, 745, 1671, 2620, 1173, 3163, 490, 871, 2815, 693, 612, 2710, 381, 654, 3456, 676, 1624, 2807, 577, 535, 997, 2585, 1103, 468, 2556, 635, 3004, 2676, 1990, 3348, 3407, 2529, 2099, 1926, 3454, 1330, 166, 2684, 603, 663, 3034, 3133, 3001, 1294, 876, 135, 168, 308, 361, 2171, 1465, 2584, 628, 299, 305, 479, 3205, 481, 1356, 3185, 3213, 2627, 682, 3425, 344, 444, 677, 427, 456, 373, 749, 3190, 429, 2407, 788, 791, 290, 142, 443, 727, 1931, 3211, 2524, 492, 2466, 2511, 2771, 3011, 705, 2029, 1221, 648]. performance [1212, 321, 528, 559, 1363, 1744, 426, 1096, 1266, 44, 1369, 1163, 1198, 870, 195, 927, 1011, 987, 95, 49, 143, 616, 1742, 1425, 2122, 760, 1287, 409, 939, 56, 1110, 1306, 1319, 1672, 1991, 28, 869, 692, 301, 1280, 814, 42, 888, 252, 463, 1389, 688, 1610, 747, 1740, 1004, 2312, 1120, 794, 713, 1834, 2038, 1178, 493, 1384, 2044, 990, 138, 163, 210, 363, 364, 1934, 40, 940, 440, 2796, 15, 2388, 1746, 2094, 1830, 2683, 164, 969, 159, 160, 495, 1131, 805, 1063, 986, 1348, 455, 1456, 3455, 591, 720, 584, 574, 2001, 882, 291, 457]. performance [821, 1037, 2385, 1485, 624, 1597, 1283, 2046, 458, 525, 783, 865, 293, 1017, 854, 332, 1317, 1588, 2150, 2214, 250, 832, 271, 2176, 1827, 504, 2417, 2232, 98, 136, 932, 1183, 472, 4, 1942, 1914, 1332, 1845, 738, 793, 666, 475, 2042, 2365, 432, 284, 393, 1748, 2290, 852, 1908, 972, 1437, 678, 1747, 1303, 721, 1211, 2043, 506, 570, 864, 923, 1685, 2015, 842, 2100, 3223, 126, 1109, 1130, 309, 2396, 802, 1595, 2008, 1285, 976, 734, 851, 484, 335, 580, 590, 771, 502, 2028, 1500, 2377, 1513, 1028, 579, 1633, 960, 1127, 1575, 549, 283]. performance [1137, 41, 1062, 1557, 681, 1340, 1635, 1559, 704, 638, 890, 147, 81, 690, 867, 1652, 447, 253, 12, 1822, 785, 1397, 942, 419, 474, 94, 2280, 966, 2225, 1759, 1644, 1006, 1566, 712, 1122, 2135, 84, 171, 1012, 1021, 1102, 1087, 1157, 1227, 1433, 2876, 3016, 3053, 3156, 3234, 3332, 3439, 2041, 306, 162, 2182, 2277, 377, 1887, 1334, 2393, 1244, 786, 303, 1666, 1840, 1210, 1824, 1272, 801, 1136, 97, 742, 550, 1055, 1153, 1328, 1331, 961, 668, 1248, 509, 1238, 2035, 2124, 1228, 774, 1429, 1219, 1587, 2158, 200, 606, 605, 608, 645, 653, 700]. performance-prediction [771]. performance-problems [525]. performing [150]. period [922, 1718]. Periodic [1132, 2646, 689, 615, 2087, 1048, 1160]. Periodic-Review [2646]. peripheral [18]. Perishable [3305]. Permissioned [3201]. Permutations [2748, 2780]. permuted [1414]. persistence [1441]. Persistent [2894, 3044, 2529, 2733, 1498, 1152]. Personalized [2578, 2985]. Perspective [2515, 3185, 2077, 1789, 1233, 239, 3233, 1642, 1935]. perspectives [1717]. perturbation [687]. perturbed [2083]. PET [661]. Peter [640, 596, 663]. Petri [947, 697, 762, 948, 837, 1746, 952, 2679, 1748, 949, 950, 3223, 954, 823, 429]. PGAS [2177]. PH [1792, 2080, 1482, 2064]. PH-Distributions [2064]. Phase [3203, 3207, 2493, 2297, 2755, 2058, 2090, 2079, 3384, 2063, 1552, 2086, 2155, 1229, 2048, 2289, 1792]. Phase-Aware [3203]. phase-dependent [1792]. phase-type [2058, 2090, 2079, 2063, 2086, 2155, 1229, 2289]. Phased [576, 763]. Phased-mission [763]. Phases [2612, 1041, 2292, 446, 512]. phenomena [779]. phenomenon [1041]. Philips [58]. Phillip [606]. Phishing [3414]. Physical [2500, 2758, 2497, 2766, 2893, 1609, 2409, 2339, 255]. physics [1434]. pick [1282]. Picocell [3277]. picture [1491]. Pictures [1714]. piecewise [2640, 2416]. piecewise-stationary [2640]. piggyback [857]. PIM [1043]. pins [2348]. pinterest [2348]. Pipe [3302, 3223]. PIPE2 [1746]. Pipeline [1576, 477, 915]. Pipelined [3295, 372]. pipelining [295]. piracy [2323]. PIRATE [1827]. pitfalls [1028]. Placement [2678, 2642, 3086, 3287, 2838, 2624, 3177, 2638, 1194, 1088, 669, 366, 2427, 2179, 2143, 1634, 523, 1960]. Plackett [2152]. Plan [2476, 393]. Plane [3245, 3397, 793, 1350, 1573]. planet [2552]. planet-wide [2552]. Planning [2703, 172, 594, 167, 1564, 1618, 2431, 478, 2402, 503, 192, 1606, 2304, 802, 173, 214]. plans [116, 2084]. Platform [3221, 3261, 2740, 1632, 1664, 1808, 1814]. Platforms [2836, 3416, 2093, 2186, 1547, 2374]. Play [2906, 2364]. playback [858]. Players [2964]. Playing [2923, 2571]. plenty [1707]. PLM [1033]. plugin [1745]. POBO [3377]. pocket [1064]. Point [3428, 1093, 1123, 1900, 1996, 720, 1854, 1385, 1533]. Points [2839, 1282, 1919]. poisoning [1806]. Poisson [1489, 1594, 2919]. Poisson-noise [1594]. Policies [2921, 3448, 3244, 3086, 2969, 2533, 2526, 3078, 2822, 2096, 857, 2805, 1721, 2056, 2412, 714, 2005, 31, 1525, 91, 757, 1090, 768, 1738, 2187, 352, 233, 2431, 624, 1778, 1445, 1547, 2356, 435, 982, 3257, 2168, 2185, 756, 1515, 1027, 539, 936, 1895, 600, 2097, 204, 2933, 1681, 1860, 677, 1224, 1210, 1770, 1395, 1526, 727, 1474]. policing [1403]. Policy [3342, 3444, 3373, 3162, 2928, 3317, 3319, 1861, 2140, 1485, 1183, 1605, 1667, 206, 46, 1081, 3343, 3122]. politely [1078]. Polling [382, 2285, 829, 1896, 1162, 696]. pollution [1494]. Pólya [2611]. Polynomial [3117, 2737]. Polynomial-Time [3117]. Pong [1591]. Ponzi [3132]. Pooch [746]. pool [1821, 625, 3391]. Pooling [3388, 3265, 3375, 2818]. Pools [2859, 3384, 1942]. PoP [1915]. popular [1830, 1500, 1840]. Popularity [3391, 1775, 1029]. Population [2566, 2886, 1480, 411, 484, 1440, 385]. Portable [944, 2174, 1228]. Portfolio [2725]. Portfolio-driven [2725]. Posit [2160]. Positive [2506, 2440]. possession [272, 101]. possible [853, 386]. Post [2815, 3043, 3382]. Post-disaster [3043]. Post-ISA [2815]. Posted [3019, 2726]. Poster [2003, 2004, 2005, 2006, 2007, 2008, 2009, 2011, 2010, 2012, 2013, 1041, 1047, 1045, 1050, 1049, 1044, 1048, 1051, 1046, 1043, 1042]. posters [2037]. poststore [719]. potato [1686, 1313]. Potential [2923, 2502, 498, 316, 986, 3230]. Potentials [992, 2390, 538]. POTRA [2156]. Power [1280, 2811, 2618, 2795, 2850, 2938, 2760, 2732, 2636, 2589, 2548, 2851, 3379, 2764, 2895, 2901, 2854, 3140, 2761, 2700, 2835, 2704, 3063, 1582, 2663, 2651, 3393, 3040, 3078, 3042, 2144, 2497, 2766, 2893, 2940, 3325, 3195, 2757, 2755, 2494, 2701, 3152, 2025, 1691, 1791, 2190, 2394, 2188, 2156, 2122, 2767, 2422, 2034, 2953, 1766, 2261, 1765, 2264, 1518, 2036, 1932, 2358, 1629, 2417, 1704, 1972, 1552, 1917, 2410, 355, 2254, 2185, 1674, 2192, 1876, 1520, 1822, 1887, 1947, 1952, 2002, 2407, 2252, 2313, 1723, 2430]. power-aware [1704]. Power-Efficient [2663]. Power-Law [3152]. power-line [2407]. Power-of- [3063, 2651]. Power-of-d [3325]. Power-of-d-choices [2854]. Power-performance [1280]. Power-to-Water [3042]. POWER7 [1925, 2173, 2456]. powered [1305]. Powering [2033]. PowerPC [1279]. Practical [2817, 3072, 3273, 2478, 1840, 2224, 54, 2180, 523]. Practice [2714, 2373, 1539, 839, 2107, 1370, 1426, 2211]. Practices [3354, 2580, 2694]. practitioners [3209]. Pre [2906, 3442, 1436]. pre-copy [1436]. pre-fetching [3442]. Pre-Observations [2906]. Precedence [3023, 3157]. Precedence-Constrained [3023, 3157]. precision [1180]. Predict [3051]. Predictability [1106, 1979, 1524]. predictable [1636]. predicted [713]. Predicting [925, 1684, 3103, 472, 958, 1415, 240, 1635, 1614, 321, 1588, 356, 1248]. Prediction [3294, 1420, 2567, 3125, 602, 2960, 2833, 1834, 1144, 164, 3228, 1191, 1929, 183, 1685, 771, 1513, 1553, 1062, 135, 1882, 84, 2338, 788, 791, 1331]. Predictions [3119, 2484, 2603, 2989, 3049, 3010, 3306, 3253, 3333, 3164, 3313, 1285]. Predictive [1444, 2555, 3328, 2172, 2841, 2351, 1116]. PredictRoute [3125]. Preemptible [3238]. Preemption [3445, 757, 1088, 537]. preemption-distance [537]. Preemptive [3180, 2834, 384, 207, 414]. prefetch [1250]. Prefetching [3084, 905, 838, 992, 1389, 815, 1154, 1276, 986, 863, 842, 1940, 2173, 712, 878, 916, 727]. Prefix [3111, 913, 2833]. Preliminary [1935, 33, 2177]. Prelimit [3404]. premium [2444]. premiums [2202]. Prentice [640, 594, 596, 605, 655, 653, 663]. Prentice-Hall [594, 655]. Preparation [3212]. Preprototyping [877]. prerequisite [42]. prescient [1250]. Presence [3075, 1052, 1503, 1661, 1961, 2353, 1692, 1993]. present [1462]. Preservation [370]. Preserve [2557]. Press [9, 642, 645, 644, 658, 746, 701]. prevalent [1966]. Prevention [207]. preventive [2002]. Price [2322, 2165, 2452, 2655, 2813, 2445, 1873, 2331, 63, 1618, 2295, 2166, 2321, 2168, 3105, 2308]. priced [1373]. Prices [3422, 2954, 3019, 2675, 2726, 2924, 2444, 2283]. Pricing [2570, 2980, 2332, 2197, 2436, 3422, 2673, 2333, 2379, 2355, 2672, 2168, 2113, 2626, 3451, 93, 2296, 3144, 2727, 2509, 2133, 64, 65, 66, 2053, 1360, 2210, 1350, 1402, 69, 1204, 2308, 2112, 2330]. pricing-network [1402]. Primal [2940, 3046, 3317]. Primal-Dual [2940, 3046, 3317]. primitive [1817]. primitives [2050]. principle [1941, 1660]. principled [2019]. Principles [3258, 1583]. printers [1838]. priorities [1179, 1472, 537, 1186]. prioritization [1347, 1777]. prioritized [1162]. Priority [2460, 647, 3443, 3180, 2463, 1185, 919, 399, 384, 2059, 1562, 1295, 435, 2257, 1058, 318, 414, 2071, 451, 1487]. priority-based [1562, 435]. PRISM [1370, 1735, 1747]. Privacy [3155, 2568, 3360, 2581, 2351, 3158, 3034, 2569, 2448, 3101, 2607, 2219, 2218]. Privacy-Constrained [2568]. privacy-sensitive [2219]. Private [2451, 3308, 3142, 2321]. Pro [2452]. Proactive [3294, 3010, 2538, 2401]. Probabilistic [2752, 3259, 1370, 1226, 554, 2247, 3337, 808, 108, 1361, 1639, 1747, 442, 1374, 391, 535, 1327, 1311]. probabilities [1170, 189, 1117]. Probability [2506, 1510, 2742, 3037, 2422, 1240, 1507, 1165, 1736, 1902, 1521, 642]. probably [1488, 1532]. probe [1205]. probe-based [1205]. Probing [2298, 2776, 1578, 2232, 2933, 2363, 2640, 2002]. Probing-based [2776]. Problem [2790, 2682, 3143, 2840, 3410, 3147, 3307, 2382, 687, 2346, 275, 2081, 1167, 1419, 2141, 1905, 339, 1604, 1779, 1642, 1435, 523, 487]. Problems [2646, 67, 1352, 2023, 2281, 1089, 2051, 525, 783, 865, 1345, 2304, 68, 560, 2338, 1891, 2341]. ProbMela [1371]. procedure [1438, 360]. procedures [50]. Proceedings [99, 2030]. Process [1017, 3282, 3345, 2849, 3402, 2584, 3235, 2877, 2074, 2075, 828, 1784, 2241, 1013, 2306, 243, 2083, 720, 840, 2085, 434, 2297, 728, 878, 254, 1251, 2125, 2336, 1439]. process-flow [720]. Processes [2986, 2886, 2611, 3453, 2551, 2789, 2751, 2526, 3159, 2830, 1371, 1170, 1985, 2090, 2052, 1641, 495, 1718, 157, 2065, 1019, 208, 1095, 2087, 1854, 1529, 1188, 1267, 1169, 2072, 1481]. Processing [2678, 488, 750, 605, 657, 3358, 3418, 2690, 3096, 3289, 653, 2466, 1505, 694, 613, 111, 1807, 2045, 2541, 676, 67, 1424, 468, 549, 670, 755, 51, 1947, 2341, 142, 1865, 963]. Processing-in-Memory [3418]. Processor [3206, 3021, 3129, 3272, 756, 2756, 3205, 726, 2787, 2928, 375, 569, 625, 1293, 1541, 321, 2805, 498, 1171, 2005, 716, 757, 1203, 1934, 1766, 731, 197, 1249, 1473, 1588, 580, 590, 544, 2392, 94, 1186, 1677, 141, 1792, 1358]. processor-input-output [141]. Processor-pool-based [625]. processor-sharing [1293, 544]. processors [622, 151, 1109, 844, 447, 2048, 1644, 2181]. procrastinating [452]. Procurement [2754, 2794, 2798, 3068, 2353]. Producers [226, 2701]. producing [530, 158]. Product [2079, 3334, 1241, 1605, 1838, 2055, 2935, 762, 1842, 178, 2014, 1098, 536, 3257, 881, 1797, 324]. Product-form [2079, 3334, 2055, 762, 3257, 881]. Product-forms [1241, 1842]. Production [2736, 2792, 2067, 1544, 283, 2073]. Production-Inventory [2736, 2792]. production/inventory [2073]. products [359, 93]. prof [814]. PROFET [3001]. PROFGEN [360]. Profile [2976, 3206]. Profile-Driven [2976]. profilers [360]. profiles [14]. Profiling [780, 2662, 3009, 3093, 2280, 2475, 1390, 766, 695, 1704, 1888, 1202]. Profit [2470, 2169]. profitability [2952]. Profitable [2806]. Profits [2084, 2444, 1172]. Program [3368, 330, 185, 372, 1882, 2847, 324, 105, 184, 559, 1022, 367, 1937, 247, 118, 108, 389, 386, 862, 27, 445, 512, 235, 246, 371, 827, 918, 523, 314]. program/architecture [386]. programmed [490]. programming [225, 124, 1923, 739, 2216, 564]. Programs [2666, 602, 38, 814, 688, 464, 662, 316, 547, 196, 695, 1568, 928, 1191, 475, 1065, 151, 667, 771, 315, 937, 876, 2393]. Progress [2899, 2468, 364]. Project [224, 393, 520]. projects [241]. proliferation [1494]. PROLOG [521]. PROLOG-based [521]. prologue [251]. promoting [2059]. promotion [1047]. Proof [2446, 2437]. propagating [1258]. propagation [1874, 911]. proper [128]. Properties [2764, 3331, 2789, 2056, 688, 1421, 1591, 1739, 1442, 2268, 1119, 2279, 514, 439, 1175, 1860, 1846]. Property [2783, 1885]. Prophesy [1244]. Prophet [3051, 2889, 2904, 3006]. Proportional [1015, 2821, 2689, 2477, 2471, 1809, 1288, 1605, 1196, 1948]. proportional-delay [1196]. Proportional-share [1015]. Proportionally [2828, 2800]. Proposal [73, 2406, 393]. proposed [666]. Protecting [3041]. Protection [3419, 2769, 2770, 3322]. PROTEUS [692]. Protocol [3017, 3429, 1234, 1141, 480, 493, 1900, 780, 613, 773, 298, 396, 1735, 1303, 784, 1265, 612, 999, 1763, 1356, 439, 1126, 708, 492]. Protocol-Constrained [3429]. Protocols [607, 705, 1737, 987, 1339, 529, 2148, 292, 1732, 872, 1733, 1049, 1736, 1648, 1401, 1443, 484, 884, 1529, 911, 1670, 775, 1573, 910, 833, 682, 1197, 429, 583, 2950]. prototype [838, 630, 476]. prototypes [1709, 1794]. Prototyping [220, 866, 221, 717]. prover [698]. provide [1302]. Provided [2570]. Provider [2606]. Providers [2674, 3369, 3142, 2984, 2167]. Providing [1075, 1042, 2125, 774, 806, 1142]. Provision [2453]. Provisioning [2989, 2502, 2146, 2201, 2462, 1976, 2283, 1989, 2421, 2271, 1967, 1038, 2352, 2309, 2910]. proxies [986, 871, 942]. proxy [1010, 1024, 957, 1154, 1038, 1039]. PRSDs [1927]. Pruning [3338, 3380]. PS [1525, 1601]. Pseudo [2697, 475]. pseudo-parallel [475]. Pseudo-Separation [2697]. Pseudonymous [3322]. PSON [1636]. Public [3100, 2660, 2722, 1198, 707]. publications [709, 789]. Publisher [2669]. Publisher-Driven [2669]. Publishers [750, 646, 657, 656, 2028]. Pull [2968]. Pull-based [2968]. Pulls [3430]. Purchasing [2437, 2624]. purity [8]. purpose [36]. Pushing [1078]. putting [686].

Q [2089, 3172]. Q-learning [3172].

Q-MAM [2089]. QBD [2083, 1718, 2085]. QBDs [2076]. QBETS [1590]. QCN [1941]. QDSL [1669]. QoE [2630, 2823, 2286, 3108, 2960, 3109]. QoS [2096, 1025, 1546, 1827, 1547, 1372, 958, 1921, 1520, 2972, 2026, 2097, 1116, 1938, 2819]. QoS-aware [2096]. QoS-based [1025]. QPME [1748]. QPS [3083, 3083]. Quadratic [3432]. Qualitative [1860]. Quality [99, 2731, 3072, 2486, 2945, 3369, 1197, 105, 1832, 129, 115, 125, 123, 1713, 1618, 120, 1661, 226, 130, 1819, 327, 102, 1323, 230, 107, 128, 1075, 1042, 109]. Quantification [2680]. Quantify [2523]. Quantifying [2460, 1064, 2319, 1835, 1336, 867, 2978, 2243]. Quantile [76]. Quantitative [1732, 2848, 3113, 129, 3224, 768, 1638, 2444, 1325, 1113, 2123, 327, 21, 597]. Quantum [3245, 3453, 3025, 3085, 3190, 1720]. Quarter [3245]. Quartz [559]. quasi [2072]. quasi-birth-and-death [2072]. queries [1852, 758]. Query [3423, 1851, 676]. Query-based [3423]. quest [1818]. Questions [304, 3209, 1345, 790]. Queue [2816, 2998, 2023, 656, 655, 2750, 2689, 3265, 3375, 2652, 2786, 2477, 3443, 2820, 873, 2587, 2882, 2503, 2540, 2999, 2855, 1592, 1691, 2298, 338, 671, 922, 350, 1514, 449, 1307, 2295, 1252, 1167, 1720, 1694, 450, 368, 1437, 201, 1168, 1264, 2356, 1179, 286, 673, 1892, 1627, 2199, 1128, 542, 1590, 320, 1347, 1417, 1446, 1883, 2102, 2640, 1311, 2289, 1792, 583, 2919, 1439, 2416, 1081, 3197]. queue-based [1307, 1883]. Queue-Differentials [2503]. queue-length [320]. Queue-Proportional [2689, 2477]. Queue-Size [2999, 2102]. Queued [2689, 3091, 3265, 2654, 3095, 3065, 3083, 2999]. Queueing [2809, 1995, 470, 3026, 3196, 3341, 3384, 643, 642, 644, 3316, 491, 792, 452, 3178, 2902, 557, 3163, 3218, 924, 654, 448, 3348, 358, 2951, 2495, 3205, 1162, 3222, 2300, 451, 3236, 2237, 2540, 3121, 3127, 1251, 2858, 370, 1999, 2746, 2373, 1016, 2197, 187, 198, 275, 1480, 1646, 2079, 178, 1349, 1517, 398, 437, 440, 661, 2287, 349, 1101, 1260, 800, 1889, 882, 1059, 179, 536, 753, 1597, 1969, 3209, 516, 522, 149, 1165, 1748, 368, 411, 341, 883, 923, 1268, 1020, 192, 1447, 281, 261, 348, 724, 2885, 514, 1296, 34, 101, 133, 282, 387, 1058, 494, 1660]. queueing [1174, 848, 377, 263, 132, 186, 860, 1516, 538, 202, 375, 881, 1329, 1556, 148, 260, 262, 3197, 385, 647]. Queueing-based [924]. Queueing-Network [3236]. queueing-theoretic [2885]. Queues [2294, 2708, 2531, 2998, 2542, 3170, 3357, 3340, 3376, 2543, 2787, 3194, 1873, 2204, 2074, 2078, 384, 1170, 856, 1612, 2080, 2024, 1807, 2059, 3089, 807, 1348, 1603, 1473, 1361, 1483, 672, 2302, 3062, 2066, 2068, 2086, 189, 2009, 2199, 1895, 1878, 600, 537, 978, 1893, 414, 1102, 1677, 2072, 400, 3374]. Queuing [3381, 701, 1669, 1253, 1317, 934, 82, 3233, 200, 84]. Quickest [3005]. quickly [1488, 1532, 1557, 790]. QuickProbe [1496]. QuickStop [3005]. Quid [2452]. Quo [2452]. quota [1264]. QWI [3172].

R [595, 588, 864]. R-tree [864]. Rack [3297]. Radio [2661, 911]. Radiometric [3155]. raid [1665, 810, 1503, 1460]. RAIDframe [866]. Rain [3328]. Raj [641]. Ralston [751]. RAN [3416]. Random [2918, 2785, 2589, 2293, 3245, 2897, 2629, 2418, 2971, 2926, 2518, 2022, 718, 1958, 3335, 2805, 1994, 1645, 1991, 2233, 1531, 2784, 1213, 829, 2126, 2140, 1101, 1240, 1864, 2065, 1191, 1957, 2127, 1752, 2021, 484, 2119, 2199, 2370, 1978, 1036, 483, 2435, 1891, 1481, 1304, 1366, 1901]. Random-Access [2785, 2233, 2784, 484, 1901]. random-order-of-service [1994]. randomaccess [1501]. randomization [820]. Randomized [2862, 1863, 2623, 2392, 2631, 2819, 2863, 2354, 2198, 1796, 1763]. Randomly [2616]. randomness [2139]. Range [2137, 1705, 1398, 970, 2110, 1251]. Ranging [3397]. Rank [3166, 2487, 3270, 3320, 2130, 2076, 1558]. rank-based [1558]. Ranking [2590, 1364, 60, 1787, 1628]. Rankings [2827]. Rao [1782]. Raphael [607, 705]. rapid [866]. RAPL [2184]. Rare [2078, 2525]. Rare-event [2078]. RATCHET [752]. Rate [1307, 2844, 1413, 2477, 2478, 2742, 3339, 2535, 3338, 3380, 1762, 1056, 1996, 1019, 1565, 2216, 873, 2009, 999, 859, 520, 1162, 1723, 2035, 2430, 1299, 3017]. Rate-adaptive [1413]. rate-and-power [1723]. Rate-based [1307]. Rate-Matrix [3338, 3380]. RateOptimal [2205]. Rates [3395, 2923, 2994, 2622, 1273, 1022, 2234, 3442, 1240, 2087, 2392, 1403, 1414]. Rating [3097, 2428]. Ratio [602]. rational [2074, 2065]. Rationing [2976]. Ratios [2544, 2907, 687, 363, 620]. Raw [2958]. RawPower [2958]. Ray [578]. RCAT [1258]. RCP [3017]. RDP [1850]. Re [2987, 3398]. Re- [2987]. Re-Thinking [3398]. Reaching [2840]. REACT [2893]. reactions [2685]. reactive [1431, 2192]. Read [2602, 199, 355, 628]. readahead [930]. reader [558]. readings [587]. Real [2310, 3418, 666, 3385, 3150, 601, 3303, 3382, 729, 730, 631, 111, 731, 1700, 77, 80, 839, 1821, 1383, 207, 433, 883, 615, 899, 1895, 704, 1408, 752, 790, 708, 1823, 1081]. real-coloring [3382]. Real-Time [601, 2310, 666, 3385, 3150, 3303, 729, 730, 631, 111, 731, 77, 80, 839, 1383, 207, 433, 883, 615, 899, 1895, 752, 708, 1081]. real-world [790]. realistic [1708, 1247]. realities [1287]. Reality [3360, 3392]. realizing [1394]. really [1580, 1579, 1219]. realtime [2051, 2189]. RealWear [3094]. reasoning [1327]. reasons [2258]. Rebecca [645]. RECAL [398, 536]. Receipts [2974]. receive [2273, 1779]. receiver [395, 932, 775]. receiver-initiated [395, 775]. receivers [2047]. recency [1183]. recently [982]. recipes [1712]. recognition [2263]. Recommendations [2914, 2510, 66, 97]. Reconfigurable [3279, 3090, 3076, 3281, 1589]. Reconfiguration [3328]. Reconnaissance [3412]. reconnect [467]. reconstruction [2400, 2153]. Record [1946]. recording [1598, 944]. Recovery [2769, 2896, 2497, 1134, 1438, 1050, 1441, 1303, 2410, 921, 1945, 803, 1043, 1850, 2018]. Recruiting [2866]. recurrence [697, 548].

Recursive [2859, 2236, 1343, 1366, 2039]. Recycling [3395]. RED [1235]. RED-based [1235]. Redesigning [2723, 1699]. redirection [1155, 1634]. Redistribution [2764]. Reduce [2467, 2556, 2553, 1249, 2262]. Reduced [2718, 469]. Reduced-Voltage [2718]. Reducing [2496, 796, 1902, 2594, 2452, 3344, 1203, 859]. Reduction [2544, 2720, 2561, 3324, 2076, 1203, 996, 2036, 974, 1582, 2401]. reductions [829, 1191]. Redundancy [1755, 3021, 3141, 2636, 2650, 3387, 2934, 2543, 2774, 3128, 3241, 3335, 2884, 2935, 1503, 1665, 1467, 2933]. Redundancy- [2650, 3335, 2884]. Redundant [2496, 1266, 555, 2193]. reentrant [345]. Refactoring [3219]. Reference [1759, 511, 736, 1154, 764, 879, 330, 1183, 1335, 524, 707, 827, 1446, 733, 1312]. Reference-driven [1759]. references [1067]. referencing [446, 403]. Refined [3216, 2853, 2886]. Refinement [3292, 1374]. Refining [3130, 2416]. Reflection [2736, 2792]. Refresh [2620, 2121]. Regenerative [820, 76, 1100, 1162]. Regime [2479, 2881, 3325, 2078, 1762, 1761, 2380, 1474]. Regimes [2715, 1981, 1358]. Region [3077, 3208]. regional [2196]. regions [1565, 1822]. Register [2501, 2733, 463]. register-to-register [463]. registers [1567]. regression [222, 1918]. Regret [3052, 2604, 2488, 2487, 2820, 3181, 3268, 2828, 3006, 3317, 2250, 2550, 3312]. Regrets [2989]. regular [1510]. regulated [1094, 1120, 1777, 2416]. Regulating [3264]. Regulation [2762]. Regulations [3092]. Reilly [751]. Reinforcement [3250, 3170, 3345, 3270, 3320, 3171, 2843, 3319, 2012]. Reinhold [751]. Reject [2557]. rejection [2065]. rejuvenation [862, 1108]. Relating [112]. relations [843, 548]. relationship [1348]. relationships [525, 227, 2321]. Relative [991, 2641, 1461, 3093, 1681, 1548].

Relaunch [2799, 2802]. relaxation [2261]. relay [1008, 2011]. relaying [1650]. relays [2676]. release [575]. releases [699]. relevance [2373, 887]. relevant [934]. Reliability [2698, 1907, 1867, 507, 2426, 1740, 1399, 1750, 1503, 2094, 424, 271, 1665, 1545, 1345, 1747, 234, 238, 235, 236, 246, 1465, 576, 51, 139]. reliability-based [1399]. Reliable [2703, 2485, 1249, 875, 983, 775, 910, 985, 1133]. Relinquishment [2450]. Relocating [2980]. remaining [1807]. Remote [3362, 3402, 2050, 1274]. removal [1594, 2433]. rendering [1418]. reneging [1473, 792]. Renewable [2115, 2034, 2310, 1968]. renewables [2031, 2311]. Renewal [2986, 2236]. Reno [1235, 1199]. Rent [3082]. Rent-Seeking [3082]. rental [2382]. Rényi [2591, 3014, 3424]. reordering [2383, 1961, 930]. reorganization [255]. Reorganizing [3137]. Repair [3043]. repairable [526]. repairman [200]. Repairs [3043, 691]. repeat [312]. repeated [1691, 2207]. Replacement [2479, 329, 1389, 562, 2140, 879, 1485, 1183, 2405, 2120, 145, 670, 1681, 561, 981, 715, 258]. Replay [1927, 1853, 1946, 1422]. Replaying [2208]. Replenishable [3415]. replica [1634]. Replicas [2934, 1635]. replicated [1025, 508, 1689, 962, 444]. Replication [2678, 2592, 2738, 2774, 3114, 2621, 2717, 2553, 2278, 338, 990, 199, 737, 1206, 1376, 2196, 479, 675, 1428, 1414, 2414]. Report [3454, 456, 33]. reporting [303]. Reports [3414]. Representation [3207, 3347, 61]. Representations [2958]. Representative [2492, 927, 1217, 1583]. Reproducing [3153]. reprogramming [1844]. Reputation [2672, 2639, 2877]. Request [2921, 480, 1045, 1703, 930]. request-response [480]. Requests [2496, 2743, 2556, 146, 1972, 1077, 761]. Requirement [2698]. Requirements [111, 110, 120, 1106, 112, 859, 966, 816]. ReRack [2034]. RESCU [1050]. Research [3346, 3176, 310, 2947, 789, 1823, 1277, 2037, 218, 332, 1149, 1819, 1710, 150, 1566]. resequencing [461, 1164]. Reservation [2445, 2577, 1090, 3059, 1497, 673, 1899, 451]. reservation-based [1899]. reservations [1625, 1179, 2113]. reserve [2924]. reserves [1040]. reservoir [1728]. RESET [3349, 3370]. resets [1161, 1258]. residential [2190]. Resilience [2840, 2972, 3436, 3136, 1379, 1378]. Resiliency [1906, 2680]. Resilient [2841]. resistance [2433]. Resistances [2984]. resistant [3201]. resizing [2145]. resolution [2047, 339, 482, 1763, 20]. Resource [2790, 2461, 3416, 3309, 2682, 2472, 2859, 3388, 2989, 2661, 2457, 2932, 2647, 2868, 3377, 3263, 3350, 3265, 3375, 2437, 1625, 2900, 3287, 3449, 3253, 2478, 1027, 3161, 2873, 2965, 3252, 3293, 2725, 3068, 3199, 3255, 2462, 2818, 3019, 3050, 2535, 3251, 3254, 2638, 2992, 2558, 2726, 1040, 31, 1623, 2092, 1937, 1090, 1684, 1292, 2307, 2281, 2372, 2431, 1694, 1778, 1814, 272, 1719, 558, 1942, 885, 1787, 1360, 2123, 2304, 72, 1226, 2885, 1336, 1296, 2427, 419, 859, 101, 2229, 1938, 2352, 2465, 1626, 1248, 1865, 988]. resource-allocation [1296]. Resource-efficient [3287]. Resource-Latency [2682]. Resource-on-Demand [2457]. resource-sharing [1778]. Resource-Utility [3449]. Resources [2980, 3305, 2905, 3056, 2648, 2920, 2883, 2576, 3036, 2695, 2925, 1560, 333, 151, 2210, 1967, 515]. respect [1395]. Responce [2559]. Responding [127]. Response [2537, 2656, 2752, 2703, 2923, 2650, 2970, 2753, 3258, 201, 2515, 3180, 2459, 3162, 2787, 2536, 3344, 2657, 830, 216, 384, 350, 480, 2200, 1807, 1252, 1892, 2256, 2355, 679, 542, 1195, 1486, 1596, 356, 543, 206, 2308, 2296, 2414, 1395, 744, 370]. responsible [1719]. responsive [2283, 1214]. responsiveness [1078]. restart [3275, 1527]. Restless [2826]. Restoration [1134, 1309]. restore [355, 215]. Restrained [1769]. Restricted [3347]. restructuring [185]. Result [3101, 2985, 1438]. Results [2721, 3334, 2753, 3381, 184, 1096, 187, 1482, 660, 805, 1240, 1904, 1776, 593, 168, 706, 710, 403]. Resume [3410, 384]. Retail [2814, 161]. retailers [2073]. Retention [2849, 2386]. rethinking [2240]. retransmission [533]. Retransmissions [2282, 2109, 2359, 1595]. retrieval [2274, 737, 892]. Retroactive [2974]. retrofitting [1780]. Reuse [2480, 2246, 1939]. Reuse-based [2246]. Revenue [2322, 2574, 2453, 3092, 1899, 2924, 1284, 2323]. revenue-fairness [2924]. Reversal [3235]. Reverse [2590]. Review [9, 640, 2646, 654, 2807, 653, 419, 641, 750, 595, 594, 596, 598, 597, 643, 604, 607, 606, 605, 608, 642, 645, 644, 646, 657, 656, 658, 655, 751, 647, 746, 66, 745, 700, 1246, 701, 588, 705, 2865]. Reviews [663, 2324]. Revisit [2085]. Revisited [3080, 1873, 1166, 1384, 1047]. Revisiting [1271, 1679, 2240, 2608, 2398, 3145]. Reward [2682, 731, 2236, 2676, 819]. Rewards [2307]. RFID [2247]. Rice [488]. Rich [3363]. Ridesharing [2696]. Rigorous [1877]. ring [498, 291, 439, 492, 354]. Rings [382, 1118]. RIO [899]. rises [2168]. Rising [2728]. Risk [2791, 3441, 2898, 2648, 1170, 1284, 391]. risk-aware [1284]. Risk-Based [2648]. Risk-Sensitive [3441]. Risks [3321, 2084]. RLNQ [3170]. rmf [3216]. Road [2970]. ROADEF [2200]. ROADEF/EURO [2200]. Robertazzi [643, 654]. Robin [3240, 1720]. robot [1145, 2255]. Robotics [3399]. robots [1125]. Robust [3382, 2993, 2791, 2019, 3169, 3409, 1233, 1491, 3277, 2957, 3311, 1477, 2400, 1383, 3230, 3269, 1666, 2020, 943, 730]. Robustness [1445, 3104, 1843, 2319]. Role [2480, 1964]. Rom [607, 705]. Ronald [655]. Root [2231, 1668]. RoSense [3274]. Round [1453, 3240, 1720]. Round-Robin [3240, 1720]. Round-trip [1453]. roundtrips [1496]. Route [3192, 409, 1531, 1829]. Router [2842, 1317, 1104, 1321, 1642, 1214]. Routers [2663, 2194, 1235, 1921, 1859, 1148]. routes [1643]. Routing [2696, 3076, 3058, 1801, 2735, 3150, 2838, 3451, 2649, 2975, 3240, 3281, 1915, 1975, 1309, 687, 1359, 1987, 1315, 1069, 1361, 1601, 2316, 1392, 1552, 1650, 1068, 1622, 288, 1955, 679, 2392, 542, 2136, 1895, 2017, 1551, 1573, 867, 459, 1333, 1956, 682, 740, 1105, 469, 2249, 1377, 1313, 1651, 2243, 585, 1516, 743, 1599, 1770, 1574, 3197]. Row [2480]. rows [2349]. RPC [1242, 3284, 785]. RPG [326]. RPS [467, 81]. RSIO [1862]. RTBH [3413]. RTC [3110]. RTE’s [251]. RTT [1053]. RTT-based [1053]. RTX [666]. Rug [3430]. Rule [2625, 2534, 2052, 1354]. Rules [3135, 317, 315, 312]. Rumor [2490, 2600, 2665, 2117, 2547, 2347]. rumours [2288]. Run [2581, 757, 1239, 876]. Run-time [2581, 876]. run-to-completion [757]. Running [1711, 1144, 1886]. Runtime [2963, 2122, 1851, 1965].

S [127, 3216, 1597]. SA [3404]. SACK [1199]. SADPonzi [3132]. safari [1289]. Safe [3377, 2631, 2863]. said [1779]. sales [2333]. Sample [3250, 3270, 3320, 620]. Sample-Efficient [3270, 3320]. sampled [1451, 1774]. SAMPLER [266]. SAMPLER/3000 [266]. samplers [2127]. samples [1773]. Sampling [2689, 2583, 2778, 2974, 3402, 2971, 2492, 716, 1213, 1654, 1950, 2129, 993, 1292, 2131, 1957, 2127, 734, 1662, 2418, 1657, 2143, 1487, 1430, 3197]. San [637]. Sandy [1998]. SAP [2584]. SARA [154]. SATCOM [2947]. Satellite [2941, 2944, 3394, 2942, 869]. Satellite-Terrestrial [2942]. satisfaction [177]. satisfying [1802]. Saturated [3334, 2587, 3381]. Saturation [3435]. Saving [2141, 2519, 1691, 2185]. savings [2043, 1921]. SCADA [3393]. Scalability [3286, 1189, 3418, 3192, 1556, 2588, 2029, 1275, 1365, 770, 2435, 2249, 2389, 831]. Scalable [2590, 3075, 1072, 875, 1650, 1970, 1633, 3034, 2996, 1076, 1339, 1178, 768, 1987, 1249, 1578, 1910, 1589, 2344, 1563, 2138, 1667, 1245, 1003, 1226, 1990, 1323, 1870, 788, 791, 1133, 1422, 2270]. Scale [3297, 2630, 3029, 3406, 3414, 2749, 2522, 3140, 3096, 2840, 2680, 2483, 2688, 2617, 2959, 2693, 3339, 3434, 2494, 2098, 2105, 129, 2188, 1616, 1702, 1795, 278, 1818, 2303, 1409, 975, 2750, 2368, 1841, 782, 3228, 2340, 504, 130, 1783, 1044, 2238, 1257, 1897, 2223, 2377, 3275, 1989, 2729, 2222, 675, 1768, 2106, 2498, 1476, 2041, 2227, 2653, 2146, 2201, 2926, 2553, 1931, 1913, 849, 513, 2831, 625, 3079]. scale-free [1783]. Scale-out [2617]. scaled [1635]. scaled-down [1635]. scales [1221, 1396, 934, 1803]. scaleup [575]. Scaling [2816, 2426, 2587, 2815, 3187, 3164, 3339, 3434, 3081, 3195, 2757, 2999, 2858, 1791, 1843, 3337, 2531, 713, 2203, 3089, 176, 1490, 893, 2409, 2268, 1675, 1107, 941, 2102]. Scalings [2686]. Scam [3431]. Scan [382, 216]. scanning [3231]. SCAT [261]. scattered [2045]. Scenario [3135]. Scenario-dependent [3135]. Scenarios [3388, 1319]. Scheduler [3086, 2707, 729, 828, 1562, 362, 1394, 308, 46]. Schedulers [2460, 3266, 1607]. Schedules [3071, 3428, 2593]. Scheduling [2474, 3241, 935, 2285, 2856, 3022, 3074, 3203, 1539, 1762, 1540, 2234, 3446, 3448, 2760, 969, 2659, 3184, 3204, 2507, 2689, 3327, 3357, 3447, 1215, 1694, 2870, 3032, 3288, 3259, 2563, 3272, 615, 1856, 3361, 2654, 3065, 499, 817, 2929, 936, 3180, 2834, 1104, 2539, 1160, 2463, 833, 2747, 2822, 3024, 3064, 3248, 964, 2705, 2403, 2503, 3183, 3023, 3157, 3255, 3043, 3403, 686, 3326, 1647, 758, 777, 3389, 1293, 1949, 2121, 1418, 2805, 1159, 1185, 1463, 2291, 1129, 1534, 1373, 1869, 2205, 1538, 1525, 1861, 967, 1349, 91, 1722, 757, 1766, 731, 841, 1535, 804, 1015, 1738, 1270, 465, 1271, 839, 2051]. scheduling [1809, 624, 1760, 1678, 1563, 1552, 45, 2262, 1179, 433, 570, 1472, 1447, 2228, 1674, 1769, 535, 2141, 2880, 600, 380, 1543, 1294, 3233, 918, 1337, 2257, 1572, 1356, 2048, 2195, 978, 726, 966, 541, 1860, 917, 2235, 2309, 626, 968, 1174, 1977, 848, 1697, 2101, 1224, 1210, 741, 1499, 254, 1599, 2213, 1346, 1395, 1393, 2125, 2772, 569, 1464, 1429, 1862, 625, 1081]. Scheme [2836, 2438, 473, 1503, 2316, 1972, 2120, 1264, 2676, 670, 1048, 1116, 318, 1920, 812, 1459]. Schemes [2944, 3132, 2651, 2860, 399, 2205, 900, 562, 1033, 67, 2392, 479, 418, 803, 2013, 2163]. Science [3067, 3212, 642, 646, 751, 2, 323, 265, 313, 327, 240, 239, 520, 228, 317, 315]. Scientific [659, 591, 3426, 464, 1285, 540]. Score [310]. Scoring [212]. screen [2047]. scrubbing [1665]. SCSI [1225, 811]. SCTP [1997]. SDA [3179]. SDN [2943, 2649, 2909]. SDN-based [2943]. SDP [2330]. SDRAWKCAB [15]. SE [588]. SE-13 [588]. Seamless [2943, 1353]. Search [2022, 2642, 2671, 1206, 2516, 850, 1232, 231, 2161, 1864, 1663, 2435, 1125, 1428, 1558, 1366]. Searching [2632]. Second [642, 885, 2015, 2410, 712, 611]. second-order [2410]. secondary [268]. sector [1570]. Secure [2609, 893]. Securing [2946, 3007]. Security [3155, 2447, 3360, 3258, 3329, 2806, 2892, 2695, 2341, 3003, 3321, 1874, 1648, 2318, 1905, 2027]. seeding [2370]. Seeking [3082]. seen [2077]. Segmentation [2981]. Segmentation-Thickness [2981]. Segmented [258]. Select [3266]. Selected [709, 789]. selecting [865]. Selection [3386, 2854, 2619, 962, 3070, 3185, 3242, 1352, 340, 392, 2402, 1472, 2028, 89, 2071, 119, 190, 1973]. Selective [2092]. Selectively [1780]. Self [1355, 1805, 853, 2661, 926, 1778, 1001, 1729, 1967, 3264, 1846, 713, 2059, 824, 1594, 941, 934, 1573, 2399]. Self-adaptive [1778, 1967]. Self-assembling [1729]. self-configuring [941]. self-learning [2399]. self-monitoring [1573]. Self-organizing [1805]. self-promoting [2059]. Self-Regulating [3264]. self-scaling [713, 941]. self-similar [934]. Self-similarity [853, 926, 824, 1594]. Self-sufficiency [2661]. Self-synchronizing [1846]. Self-tuning [1001]. selfish [1555]. selfishness [1975]. Selftuning [2012]. sell [2953]. Semantically [1462]. Semantically-smart [1462]. semantics [610]. Semi [3117, 1585, 437, 1128, 2915, 831]. Semi-Bandits [3117]. semi-empirical [831]. semi-homogeneous [437]. semi-Markov [2915]. semi-Markovian [1128]. Semi-supervised [1585]. Semidefinite [1878]. send [1779]. send-receive [1779]. sender [395, 775]. sender-initiated [395, 775]. sensing [1982, 1609]. Sensitive [3441, 2699, 2219, 2365]. Sensitivity [2188, 507, 1554, 401, 1211, 976, 723, 2381, 1720]. Sensor [3033, 2677, 3353, 1509, 2683, 1844, 2400, 1413, 2423, 1441, 1305, 1845, 1729, 2377, 1440]. Sensors [2781]. Sentiment [2954]. Separability [3449]. separable [148, 262]. Separation [2697, 2818]. Sequence [3227, 1725]. Sequential [3175, 2643, 3202, 3359, 2576, 825, 199, 1189, 2424, 690]. sequential-write [199]. serialization [343, 390]. Series [3008, 3112, 2865, 3109, 363, 1590]. serpentine [854]. Served [3326]. Server [2531, 3435, 1080, 2592, 2854, 3178, 2557, 3163, 2971, 3121, 2470, 343, 1216, 1155, 850, 2056, 927, 2078, 617, 1995, 1194, 1407, 1150, 440, 2796, 2024, 781, 1015, 3384, 1252, 1765, 797, 893, 1601, 1776, 1283, 452, 1976, 2417, 2302, 958, 1437, 1168, 673, 1196, 2086, 1627, 544, 960, 899, 1127, 1575, 1195, 1051, 420, 761, 1543, 765, 1417, 1896, 1337, 200, 1220, 1021, 1102, 373, 1075, 880, 1152, 2919, 988, 1422, 2527]. server-limited [1896]. Serverless [3295, 3426, 1035]. Servers [3276, 2749, 3075, 1071, 2725, 2653, 3144, 3195, 3240, 2294, 3021, 1040, 2284, 2204, 959, 808, 1073, 1074, 2312, 1154, 990, 900, 1025, 1444, 1851, 1271, 796, 898, 1072, 1149, 941, 1027, 1500, 997, 1038, 1408, 1036, 962, 2144, 897, 188, 2227, 894, 2245, 186, 1153, 849, 369, 1031]. Service [3413, 2674, 2578, 2662, 2586, 3387, 3433, 2738, 2774, 2565, 2688, 2882, 2616, 3385, 3268, 3084, 3339, 3434, 3369, 2629, 2653, 2984, 2476, 2606, 3383, 1592, 2096, 3335, 2298, 1155, 1994, 338, 919, 453, 1863, 412, 1482, 2079, 1669, 1405, 1150, 2024, 2287, 1618, 1379, 1061, 2750, 1889, 1598, 1288, 2302, 1295, 2231, 885, 2067, 673, 1268, 1172, 1897, 2086, 1302, 1579, 2392, 998, 600, 2271, 1034, 1162, 1636, 572, 573, 414, 985, 2227, 1197, 2910, 2296, 1075, 2289, 2112, 961, 1042, 1251, 2163, 1683, 2401, 731]. service-level-agreement [1172]. service-oriented [2231]. Services [2517, 2692, 2449, 2473, 2574, 2502, 3010, 2813, 3101, 2462, 3148, 3254, 2907, 2440, 3283, 1132, 1151, 1086, 407, 63, 991, 1833, 66, 1077, 1875, 2169, 2100, 1226, 1858, 2421, 2147, 69, 2097, 2253, 1703, 2146, 1406]. Serving [3284, 2380, 2617, 2538, 2401]. Session [2867, 2868, 2869, 226, 2870, 2871, 2872, 2807, 2873, 2874, 2875, 1041, 1047, 292, 1045, 1050, 1049, 678, 1044, 2899, 236, 1048, 822, 1051, 1046, 1043, 1042]. sessions [1079, 1357, 932, 1652]. Set [2599, 2832, 1035, 766, 1773, 169, 259, 931, 29, 1183, 1066, 145]. Sets [2791, 2891, 2518, 563, 165, 170, 1229]. Settings [2845, 3421, 1248]. Settling [2026]. Setup [3343, 3373, 3315, 1904, 2236]. several [1041, 210, 2295, 132]. SGI [929]. SGX [3004]. Shade [766]. shakers [2378].

Shallow [3279]. Shape [60]. shapers [1119]. Shaping [2669, 2683]. Shapley [1858]. share [1015, 1809]. Shared [2924, 2659, 2437, 2920, 285, 528, 191, 1995, 383, 621, 891, 551, 536, 1546, 1694, 1191, 565, 558, 972, 378, 2123, 1902, 147, 876, 462, 1054, 770, 626, 836, 741, 1136, 2110, 2125, 569]. shared-memory [528, 551, 626, 741, 1136, 2125]. Sharing [2451, 3021, 2579, 2821, 3129, 2533, 2672, 2438, 3161, 2965, 3205, 2577, 3199, 2465, 2787, 2928, 2800, 2501, 1293, 1541, 3160, 2805, 2164, 760, 1171, 2005, 1341, 627, 1379, 395, 495, 3059, 1996, 1696, 1473, 1778, 2348, 614, 1719, 2359, 1288, 1392, 1605, 1394, 1264, 1571, 1265, 3092, 1595, 544, 2392, 2323, 269, 33, 1692, 897, 740, 1660, 1677, 1499, 1626, 1792, 2183, 2181, 59, 1723, 92, 1358]. sharing/batch [68]. Sharma [656]. Sharp [2269, 3341, 2051, 3178]. SHARPE [1749]. shedding [2410]. shelf [686]. Shift [3275]. Shifting [3410, 2411, 2256, 2527]. Shneiderman [114]. Shoot [2633]. shop [2382]. Short [2674, 2671, 2676, 2670, 2675, 1316, 2184, 1396, 1049, 1528, 1474]. short-term [1049, 1528]. shortest [2884, 1807, 542, 1081]. Should [2739, 3422, 3456, 3004, 1933, 2552, 2127]. Shrinking [3424]. Shuffle [3282, 469]. shuffle-exchange [469]. Shuffling [2878]. Side [2460, 2835, 2369, 2762]. Sided [3144, 3194, 2727]. Sidi [607, 705]. SIGMETRICS [3346, 3356, 3176, 2899, 2865, 3066, 2037, 963, 87, 69, 2728, 2892]. SIGMETRICS/CMG [87]. signal [2226]. Signals [3331]. Signatures [3040, 2404]. signed [2129]. significantly [1249]. SIMD [877]. SimFlex [1283]. similar [934]. similarity [853, 824, 1594, 926, 1558]. Simmons [645]. Simple [3448, 2523, 2691, 2881, 1177, 3024, 1800, 140, 394, 1249, 1392, 1320, 378, 809, 78, 487, 200, 981]. SimpleScalar [1278]. Simplex [2708]. simplifications [829]. simplifies [2242].

SimPoint [1282]. SIMROT [1987].

Simulated [151]. Simulating [1170, 1019, 165, 170, 3001, 3222, 735, 1335, 155]. Simulation [641, 2955, 143, 75, 2525, 2848, 3398, 970, 1928, 3251, 1281, 2950, 1999, 697, 1632, 795, 2078, 1254, 846, 1280, 2034, 688, 91, 1203, 76, 773, 571, 1641, 197, 1253, 1249, 1282, 1283, 1932, 1431, 1923, 634, 1298, 131, 150, 88, 818, 1190, 1484, 724, 776, 462, 2155, 910, 153, 137, 835, 2670, 255, 530, 715, 787, 152, 588, 1516, 538, 270, 550, 1430, 142, 139, 845, 1800, 2251]. simulations [716, 1808, 725, 548, 1640, 974, 489, 911, 133, 1105, 388, 2183, 1709, 1794]. simulator [1279, 1616, 692, 766, 154, 35]. simulators [3275]. Simultaneous [556, 101, 2250, 336, 1408, 1186, 272]. Simultaneously [3181]. Single [2531, 2965, 2632, 1995, 1090, 1015, 1252, 797, 769, 1437, 1168, 765, 1385, 2362, 200, 1644, 469, 2357, 1850, 2018]. single-address-space [769]. single-packet [2362]. single-resource [1090]. single-server [1015, 200]. single-server-multiple-queue [1168]. single-state [469]. SINR [1622]. SINR-based [1622]. SIP [1575]. SIQUEUE [661]. SIQUEUE-PET [661]. SITA [2204]. sites [1145, 1363, 1332, 1364, 1174, 1238]. Situational [3362]. Size [2474, 2917, 2729, 2832, 3233, 3267, 3372, 881, 3390, 2999, 1721, 1218, 1140, 298, 2001, 1597, 1720, 2103, 1301, 1381, 1774, 235, 246, 2102, 385, 1526, 1647]. Size- [3390]. Size-aware [2474, 2103]. Size-based [3233]. Size-limited [881]. Sizes [3022, 2817, 3433, 3064, 2995, 1480, 1215, 29, 1066]. Sizing [3198, 1558, 3136, 379, 359, 1642, 2315]. sketches [1577, 1558]. skew [694, 547]. Skewed [3435]. Skewness [2009]. ski [2382]. skill [2294]. skills [3455]. Skype [1714]. SLA [1642]. SlackSim [1808]. sleep [1729]. Slicing [3183]. Slick [1954]. Sliding [3083].

Sliding-Window [3083]. SLITS [3288]. slotted [1443, 1992, 492]. slotted-Aloha [1443]. Slow [2744, 1041, 2123, 1627]. slow-down [2123, 1627]. Slowdown [3074, 3168, 1230, 2103]. slowly [1349]. SM [814]. SM-prof [814]. Small [160, 1852, 2711, 1762, 547, 279]. small-buffer [1762]. smaller [1492, 3275, 1457]. Smart [3123, 3261, 3132, 3020, 2515, 3150, 2660, 2509, 1462, 2410, 2192, 1899, 2313, 2311, 1750, 1346, 1393, 2330]. SmartNICs [3282]. smartphone [2226, 1909, 2481]. Smartphones [2154, 2003, 1884, 2960]. Smash [3287]. SMCSolver [2089]. Smith [608, 700]. Smooth [1419, 1898]. Smoothed [2890, 3018, 3313, 687]. smoothing [859]. Smoothness [2571]. SMR [2627]. SMS [2276]. SNA [292]. Snapshots [2643]. SOAP [2822]. Social [3249, 2781, 2489, 2602, 2644, 3179, 2931, 2665, 2439, 3262, 3427, 2040, 1711, 2368, 2348, 2340, 2219, 2324, 2350, 2370, 1911, 2128, 2072, 1823, 1797]. social-driven [2348]. Soft [3206, 3092, 1944, 1057, 1106, 1895, 1886, 1081]. SOFTDOC [242]. Softw [588]. Software [99, 103, 2681, 3219, 608, 120, 424, 2505, 3355, 1675, 1, 14, 1673, 117, 311, 235, 236, 246, 2585, 239, 700, 312, 105, 3211, 3135, 345, 244, 195, 229, 129, 115, 231, 225, 323, 125, 223, 123, 359, 394, 1131, 36, 106, 233, 237, 226, 325, 333, 227, 504, 130, 313, 738, 13, 17, 2290, 234, 238, 1859, 476, 1177, 327, 240, 112, 23, 1330, 1104, 1148, 230, 813, 241, 520, 1988, 203, 344, 242, 158, 113, 119, 107, 1327, 228, 1935, 128, 1108, 742, 1794, 139, 109, 317, 224, 315]. software-based [1104, 1148]. Software-Defined [2681]. Software-directed [1675]. Sojourn [2299, 3239, 1269, 1447, 1446]. Solar [2758, 3271, 2264]. solid [1767, 2238, 269, 2239]. solid-state [2238]. Solution [536, 753, 1175, 730, 342, 209, 1098, 1059, 32, 772, 516, 282, 387, 414, 1169, 190, 523, 148, 262, 1587]. Solutions [2804, 3026, 2646, 3453, 589, 920, 1257, 723, 1188, 2342]. Solving [1612, 3143, 348, 1344]. Some [3115, 1425, 747, 63, 2542, 2702, 2835, 4, 1119, 2654, 241, 1310, 483, 10, 1210, 56, 2114, 2051, 2157, 322, 1529, 168, 282, 387, 418]. Sometimes [2482]. Sons [641, 604]. sorting [481]. Source [2490, 2600, 2926, 1728, 2406, 2402, 2007, 1615, 2117, 2347, 121]. Sourced [2753]. Sources [2816, 2547, 1045, 1020, 2228, 2353, 1857, 1162, 1477]. Sourcing [2486]. SP [1931]. SP2 [868]. Space [3388, 3073, 3028, 2920, 2411, 829, 1456, 2288, 980, 2232, 769, 972, 1685, 1060, 315, 232, 1657, 287, 1493, 2913, 680]. Space-Shared [2920]. Spaces [2438, 1783, 1955]. spacetime [2204]. spaghetti [1656]. spam [2276, 1495, 1753, 1880]. spammer [2004]. spanning [2371]. SPARC [666]. spark [2361, 2679]. Sparse [2996, 1958]. Sparsity [2937, 3288]. Sparsity-Lightened [3288]. Spatial [2785, 2634, 2525, 3291, 552, 1761, 1640, 2405, 1854]. Spatially [3078]. Spatio [1322, 1591]. Spatio-temporal [1322, 1591]. SPEC [955, 1581, 1192, 593]. Special [2461, 1916, 1740, 3214, 3170, 3246, 901, 1082, 950, 2899, 895, 907, 964, 1012, 1087, 1157, 1227, 1433, 2876, 3016, 3053, 3156, 3234, 3332, 3439, 1007, 1243]. specific [1770]. Specification [222, 2560, 156]. specifications [788, 791]. Spectral [2944, 3015, 2764, 2419, 1437, 1888]. spectroscopy [1576]. Spectrum [2451, 2438, 1868, 1826, 982, 1723, 2224]. Speculation [2539]. Speculation-aware [2539]. Speculative [2699]. speech [1615]. Speed [3206, 2531, 3093, 3434, 3081, 3195, 3247, 1791, 1843, 1567, 2203, 678, 1053, 2370]. Speed-scaling [2531]. Speeding [388, 2518, 3136]. speeds [2287]. Speedtest [3298]. speedup [575]. sphere [2778]. Spins [2644]. Splicing [3285]. Split [2542, 3177, 1232]. Split-Merge [2542]. SplitPlace [3177]. SplitRPC [3284]. Splitting [2081, 3284, 2148]. SPMD [1927]. Sponsored [2489, 2510, 2476]. Spot [3422, 2700, 500]. Spread [2441, 2665, 3262, 3303, 2733, 2395]. Spreading [2643, 2288, 3159, 1719, 2119]. spreadsheet [753]. Springer [643, 607, 654, 705]. Springer-Verlag [643, 607, 654, 705]. Spy [2490, 335]. SQ [2928]. SQL [575]. square [498]. SRM [921]. SRPT [1949, 1158, 1129, 1269, 2879, 2912, 3447, 3272, 1892]. SRPT- [3447]. SSD [2306, 3228, 2240, 2635]. SSDs [2150, 3094, 2922, 2390, 2524, 2546]. STAB [1322]. Stability [1941, 1890, 3334, 1168, 3128, 2530, 2887, 1943, 870, 1531, 1223, 1101, 1095, 1264, 1565, 1595, 2017, 374, 1770, 1962]. Stabilization [3314]. Stable [2133, 1069, 2965, 3070, 2535, 2573, 2270]. stably [2411]. Stack [3408, 3284, 2852, 2693, 2598, 666, 634, 2365]. Stage [2667, 32]. staggered [1904]. Stake [2446]. stalls [2176]. stamping [1109, 2363]. standard [2887]. standards [1710, 2211]. Star [2728]. Stardust [1466]. StarShip [3426]. start [2884, 1041, 1056, 237]. starvation [2927]. State [2937, 3294, 2809, 2816, 2804, 3388, 3452, 2938, 829, 1807, 2917, 3073, 2994, 3028, 2881, 3449, 3130, 2766, 3390, 3404, 2913, 680, 2075, 1511, 1767, 2808, 2479, 792, 333, 772, 2238, 2297, 2095, 1060, 954, 2155, 269, 682, 1657, 469, 2239, 1498, 1587]. State-Aware [3390]. state-based [1587]. State-dependent [1807, 333]. State-space [2913]. statecharts [1372]. Stateful [3247]. States [3217, 1414]. static [16, 334, 693, 637, 242]. Stationarity [1529]. Stationary [687, 2083, 337, 2743, 2929, 3335, 1482, 1613, 2750, 1240, 1605, 2066, 1095, 1060, 2434, 3450, 2640, 1894, 2416]. Stations [3026, 2662, 2962, 440, 436, 263]. Statistic [2493]. Statistical [3221, 1014, 2845, 2959, 35, 3007, 3035, 2783, 3037, 1434, 1421, 1056, 1934, 525, 98, 136, 2328, 1394, 1495, 2136, 635, 1877, 776, 474, 1882, 1202, 774, 9]. Statistically [3117]. Statistics [3212, 642, 11, 1872, 52, 1707, 1796, 1614, 1493]. Status [3161]. Stay [3155, 3422]. Steady [2804, 2808, 3452, 2479, 2917, 2881, 3404, 1511, 772]. Steady-State [2804, 3452, 2917, 2881, 3404, 2808, 2479, 1511, 772]. stealing [939, 1841, 3257]. Stealthy [1806]. Stein [2804, 2715]. steps [4]. stepsize [3404]. Stepsizes [3318]. StoCharts [1372]. Stochastic [3225, 1793, 3045, 948, 2369, 3214, 2645, 2646, 3316, 1167, 2281, 2198, 2521, 3433, 3318, 3358, 2857, 2724, 1942, 2067, 2238, 949, 238, 1020, 2605, 3449, 2551, 2789, 2751, 947, 3217, 1284, 2088, 3235, 1542, 2717, 786, 3326, 3025, 1628, 3046, 697, 762, 1534, 2075, 1564, 1511, 837, 1368, 1750, 1405, 2269, 2091, 1746, 829, 3059, 952, 1260, 2307, 2432, 1100, 1510, 2358, 1679, 1378, 1360, 1724, 950, 1606, 1903, 1402, 1727, 2054, 954, 413, 2069, 978, 1726, 1533, 774, 646, 655]. Stochastically [3120]. stop [237]. Stopping [3005, 2412]. Storage [3123, 2578, 2719, 1983, 2753, 2623, 3287, 895, 2556, 2717, 1331, 2637, 2701, 2190, 268, 1266, 2415, 331, 1388, 2306, 1503, 2187, 796, 1037, 1546, 1630, 1286, 1665, 1288, 2120, 2180, 183, 1907, 2238, 2168, 855, 906, 1461, 1548, 1137, 896, 2099, 1467, 1689, 34, 1205, 21, 1466, 894, 2116, 1850, 2035, 513, 2018]. Storage-as-a-Service [2578]. Storage-Assisted [2753]. storages [2405]. Store [3000, 2105]. stored [1670, 1502, 859, 1952]. stores [329]. Storms [2507]. story [2054]. Straggler [2739, 2799, 2802, 2553, 2831]. straight [498]. Strategic [2441, 2894, 3324, 3144, 2607, 2296]. Strategies [2457, 3182, 2797, 2896, 3039, 2838, 3084, 3068, 2701, 1309, 2394, 1280, 815, 737, 2952, 1085, 2028, 533, 556, 2353, 380, 813, 2390]. Strategy [2437, 2961, 1232, 2417, 2167, 535]. Strategy-proof [2437]. Strategyproof [3421]. stratified [1957]. Stream [2678, 1505, 1218, 2212]. Streaming [2630, 3072, 3108, 2628, 3310, 2673, 1156, 2533, 2973, 2513, 3303, 3417, 1146, 1357, 1306, 1672, 1616, 1132, 1444, 1231, 1589, 1189, 1301, 1381, 1478, 1671, 1810, 2943, 1812, 1990, 1670, 1502, 1993, 1197, 1328, 971, 1411]. Streams [2507, 736, 1950, 2129, 1045, 798, 886, 732, 903, 1323]. strength [2226]. stretch [1955]. Strict [3139, 1472, 2880]. Strided [3189]. strip-integral [1594]. Striping [810, 1036]. Strong [2572, 1173]. Strongly [2737, 3389, 186]. Structural [2697, 1290, 2789, 281, 2985, 1140, 2259, 1928]. Structure [2590, 1950, 2646, 266, 3270, 3320, 3359, 3262, 124, 64, 1663, 413]. Structure-aware [1950]. Structured [217, 3453, 225, 1613, 2506, 1643, 185, 1303, 1121]. Structures [3000, 1368, 2306, 1679, 2104, 371, 3111]. Stubborn [2625]. Student [3346, 3176, 2037]. Students [3456]. studies [993, 1594, 98, 136, 1370]. Study [3399, 2500, 3414, 2732, 3387, 2851, 3012, 3430, 2704, 2710, 2483, 2756, 81, 1726, 2830, 3124, 86, 1222, 244, 1234, 2122, 1022, 888, 401, 815, 91, 1178, 405, 493, 990, 1868, 76, 1365, 582, 2006, 975, 531, 782, 406, 478, 1925, 1113, 2214, 221, 1191, 2386, 2210, 2087, 635, 767, 637, 1005, 813, 372, 1768, 1006, 1919, 770, 21, 1122, 2364, 788, 791, 1328, 59, 214, 1798, 831]. Studying [688, 2368, 1940, 1939, 250]. style [3151]. Sub [3419, 2881, 3325, 1981]. Sub-Halfin [2881, 3325]. Sub-Page [3419]. subcube [623]. Subexponential [672, 1170, 1251]. subgraph [2433]. subject [2062, 1070]. Subjects [2607]. Sublinear [3181]. Submarine [3437]. Submodular [3310, 3146, 3145, 2593]. subnetworks [411]. subpopulation [1381]. subprograms [16].

Subscriber [2452]. subscripted [662]. Subscription [2453]. Subsequences [2748, 2780, 1159]. subset [1852]. Subsidies [2440]. subsidization [68]. subsidy [2163]. Substation [2732]. subsumes [982]. subsystem [33]. subsystems [422, 404, 256]. successes [2052]. Successful [3239, 130]. Successive [1870]. sufficiency [2661]. Sufficient [2690, 2588, 2968, 125]. SUGAR [3136]. suitability [1479]. Suite [3175, 3002, 1278, 650]. suites [603]. Sum [2844, 3195, 10]. summaries [1383]. Summarization [3146]. Summary [2753, 963, 706, 710, 938, 611, 310, 517, 738, 1802, 1582, 1884, 84, 741]. Sun [666]. Supercomputer [637, 1501]. supercomputers [1931]. superhashing [2953]. Supermarket [3129]. supermartingales [1600]. superpage [1047]. Superposition [2668, 1056, 732]. Superposition-Traffic [2668]. superscalar [1109]. Supervised [3036, 1585]. Supply [3051, 3019, 2701, 513]. Support [3419, 3280, 903, 3183, 1975, 1151, 1131, 1617, 1066, 936, 24, 849, 1031]. supported [2196]. Supporting [998, 859, 2837, 1535]. Surprising [1776]. surrogate [272]. Survey [149, 390, 2211]. Survivability [2192]. survivable [2371]. survive [2062]. susceptibility [2367]. Sustainability [2965, 2271, 1839, 1835]. Sustainable [2454, 2115]. SVR4 [828]. SW [3083]. SW-QPS [3083]. Swapping [2759, 257]. swarming [1788, 1813, 2028]. swarms [1962, 2270]. sweep [1729]. sweep-and-sleep [1729]. Switch [3073, 3265, 3028, 2549, 2587, 3025, 3085, 2315, 2833, 453, 1151, 834, 891, 837, 557, 1003, 2110]. switch-over [453]. switched [1847, 778, 586, 1294, 2102, 848, 194]. Switches [3302, 2689, 3091, 2654, 3095, 3065, 3083, 2999, 2205, 1164, 1790, 1419, 502].

Switching [325, 3328, 3409, 3083, 3333, 3180, 3378, 3077, 3208, 2631, 2863, 294, 1262, 116, 1517, 2335, 1419, 2206, 459, 1334, 2357]. SWNs [953]. Symbiotic [1186]. symmetries [1099]. Symposium [963]. Syms [26]. Sync [2690]. synchronization [1906, 1413, 624, 1001, 972, 1611, 413, 1653, 1136, 443, 2035]. Synchronized [1709]. synchronizing [989, 826, 1846]. Synchronous [2886, 248]. Synectics [10]. Synergy [2774]. Synopsis [164]. SynRGen [764]. Synthesis [2052, 2842, 2095, 2425]. Synthetic [2993, 1561, 617, 688, 421]. System [2937, 2564, 3193, 218, 3452, 2938, 176, 597, 606, 605, 608, 646, 2650, 325, 3134, 161, 3418, 3258, 745, 3163, 653, 2616, 2852, 3040, 700, 704, 3001, 2660, 2629, 601, 1708, 1221, 959, 1664, 379, 1619, 143, 90, 1742, 617, 1279, 1035, 904, 75, 31, 888, 79, 1074, 2286, 2092, 967, 794, 631, 1767, 480, 405, 649, 1032, 449, 40, 222, 940, 613, 521, 975, 164, 160, 43, 2750, 629, 720, 397, 782, 882, 1059, 457, 2385, 624, 1597, 518, 32, 1230, 1283, 336, 504, 4, 122, 2240, 1383, 1065, 2365, 134, 393, 131, 378, 88, 818, 1268, 126]. system [679, 58, 667, 442, 1190, 510, 1513, 577, 1484, 283, 767, 62, 96, 1488, 1532, 1557, 366, 356, 1330, 166, 339, 82, 1775, 147, 1524, 447, 53, 12, 153, 269, 94, 137, 264, 1058, 1938, 1660, 2309, 887, 242, 1122, 763, 892, 84, 220, 374, 2227, 19, 1935, 2788, 1466, 388, 1030, 801, 2237, 152, 194, 2073, 916, 2428, 2399, 417, 1574, 1219, 1031, 302, 465, 121, 161]. system-level [2240]. system-wide [1938]. Systematic [3399, 3135, 1005]. Systematically [1832]. Systems [3457, 3249, 641, 2955, 3029, 2696, 1927, 2678, 2730, 3276, 2436, 3360, 3031, 750, 640, 2659, 595, 594, 596, 643, 657, 647, 310, 2586, 3020, 3387, 3280, 2525, 2686, 2879, 2912, 3334, 2895, 2901, 3140, 3134, 3433, 2902, 2704, 3063, 2555, 3259, 3096, 3129, 901, 3258, 789, 2485, 3056, 3361, 3401, 2680, 2467, 2478, 2945, 654, 2688, 895, 3004, 2965, 2529, 3150, 3454, 2621, 663, 3128, 2874, 2495, 2448, 3339, 3434, 3168, 964, 2577, 306, 2653, 3035, 2717, 2940, 2971, 2928, 2694, 2877, 3336, 3314, 2471, 3405, 3338, 3380, 2855, 2968, 3079, 3366, 2029, 857, 343, 461, 477, 505, 1780]. systems [2098, 3160, 1262, 2190, 935, 1462, 167, 426, 838, 1266, 830, 340, 182, 116, 95, 2285, 616, 689, 507, 409, 1254, 453, 28, 146, 50, 252, 198, 275, 3224, 1434, 1820, 1669, 691, 1702, 91, 1388, 1659, 1795, 1834, 1368, 2241, 278, 199, 1090, 1868, 2193, 1818, 473, 768, 111, 1503, 2303, 2094, 829, 1684, 1781, 1379, 2203, 1015, 807, 652, 1098, 1490, 1438, 980, 3059, 531, 1101, 1864, 1898, 1841, 337, 574, 175, 1638, 233, 1089, 847, 926, 1546, 1630, 753, 1694, 1778, 1969, 1100, 1239, 390, 517, 664, 2419, 3209, 614, 1057, 2344, 783, 865]. systems [1001, 415, 452, 250, 271, 2103, 1665, 1658, 272, 578, 1066, 1458, 1819, 404, 149, 207, 769, 2290, 1115, 1729, 1639, 1853, 1168, 1179, 1561, 1571, 883, 923, 1724, 2238, 2123, 615, 1673, 1020, 1897, 289, 1245, 1173, 1447, 68, 2337, 431, 499, 3257, 953, 1376, 484, 809, 156, 1936, 1374, 1699, 1758, 855, 1788, 1813, 2028, 1586, 1624, 539, 2392, 508, 526, 281, 1989, 468, 261, 348, 448, 896, 1816, 2271, 560, 755, 761, 1467, 703, 204, 1543, 1689, 1417, 253, 1896, 2196, 479, 2257, 675, 133, 1310, 2339, 494, 917, 696, 200, 770, 158, 1542, 113, 162, 2179]. systems [1946, 2182, 2366, 444, 24, 2201, 515, 786, 51, 132, 97, 1108, 1481, 880, 788, 791, 202, 1153, 2341, 375, 812, 59, 346, 509, 1329, 1251, 1850, 2428, 2013, 1913, 513, 2039, 2401, 2018, 3197, 1751, 2527, 611, 789].

T [789, 216]. T-scan [216]. Table [2614, 85]. Tables [3396, 2049, 1377]. Tackling [3276]. Tactical [2342]. Tail [3446, 3444, 3340, 3376, 2066, 2434, 3450, 3162, 1474, 3389, 1170, 1253, 2001, 672, 2085, 1774, 1887]. Tail-Optimal [3389]. Tailed [2595, 3389, 1357, 1019, 1229, 2289]. Tailor [140]. Tails [1538, 2858, 1512, 1438, 2279, 2101]. Takagi [646, 647]. Take [3115, 2202]. Take-aways [3115]. tale [2250, 213]. Talisman [795]. Talk [2728, 2936, 2674, 2671, 2676, 2670, 2675]. Tandem [1627, 2299, 3374, 1482]. TANGRAMII [1751]. TAO [3398]. TAPAS [3217]. tape [854, 826]. TARDIS [2411]. Target [2622]. Targeted [2583, 2491, 3041, 1911, 2329]. Targeting [2520]. Task [940, 2592, 3107, 2738, 2911, 1721, 622, 1776, 207, 66, 1160, 626, 786, 2213, 2414]. tasking [1305]. Tasks [2799, 3322, 2802, 2747, 3023, 3157, 3137, 1185, 731, 1144, 615, 1769, 1527, 686]. Tau [3222]. TauSSA [3222]. taxonomies [2211]. TB [3062]. TB/GI/1 [3062]. TBBT [1422]. TCAM [1688, 1699, 1758, 2132]. TCAM-based [1688, 1699, 1758, 2132]. TCAMs [1567]. TCP [1273, 1052, 1212, 984, 2194, 1316, 1041, 1530, 1357, 1234, 1649, 1093, 1123, 1981, 1056, 2001, 1420, 920, 1235, 738, 1085, 2273, 1124, 983, 1053, 2710, 1028, 1553, 822, 874, 999, 2248, 3082, 912, 2693, 1034, 1236, 1255, 1199, 1126, 1328, 1259, 1208, 1176, 2008]. TCP-friendly [1357, 999]. TCP-Vegas [1259]. TCP/IP [738, 1126, 1208, 1176]. Teach [3457, 3456]. Teaching [3455, 3454, 3210, 3211]. Technical [611, 69, 3368, 95]. Technique [3349, 285, 370, 735, 1578, 1320, 22, 1343, 1870, 153, 835, 2788, 979, 1344]. Techniques [641, 551, 3370, 2505, 2832, 3102, 3105, 139, 505, 1024, 231, 83, 1449, 990, 2241, 933, 1818, 1061, 176, 821, 390, 783, 293, 130, 2386, 771, 1380, 635, 1143, 353]. technologies [1983, 1984, 1272]. Technology [3211, 249]. TED [911, 908, 909, 912]. TeksDB [3000]. telecom [1918]. telecommunication [908]. Telecommunications [746, 2193, 907]. teleconferencing [805]. Telegram [2954]. telephone [1003, 374]. Telephony [2004]. teleprocessing [90]. tell [1126]. Telling [2851]. Temperature [2114, 2272, 2422]. Temporal [1045, 2195, 2598, 2927, 1591, 2405, 2095, 827, 1322]. Temporally [946]. temporary [2068]. Tenant [2537, 2761, 2700, 2908, 2504, 2835, 2511]. tensor [1098]. terabytes [1375]. term [1844, 1049, 3293, 1528]. terminal [94, 24]. ternary [1492, 1852]. TERRAIN [1845]. Terrestrial [2942, 869]. TES [807]. TES/GI/1 [807]. Test [3298, 2619, 230, 116, 222, 119]. testbed [1132, 1910, 488]. Testing [2599, 1632, 231, 111, 233, 237, 1, 117, 232]. tests [1321]. text [987, 315]. Textbooks [3213]. texts [163]. th [2921]. their [3030, 2150, 1362, 2155, 541, 1210]. them [609]. themed [3323]. theorem [698, 1763, 2919]. Theorems [2593, 762, 2330, 1784]. Theoretic [2821, 2667, 1912, 2162, 2167, 1302, 1233, 2885, 572, 573]. Theoretical [1573, 2497, 648, 1868]. Theory [948, 643, 642, 644, 655, 868, 1426, 654, 2714, 3362, 2951, 3248, 3198, 2783, 701, 3194, 2860, 3038, 2373, 820, 1517, 63, 1359, 164, 64, 839, 7, 3209, 2157, 1821, 1317, 4, 322, 883, 1263, 2248, 1857, 119]. there [2552]. Thermal [2702, 2908, 1932, 1673, 2141]. Thickness [2981]. thin [1201]. thin-client [1201]. Things [2378, 2815]. Thinking [2744, 3398]. Third [751, 3299, 3133, 985, 3454]. Third-Party [3299, 3133, 985]. Thomas [643, 654]. Thorough [3406]. those [958]. Thoughts [2597]. thrashing [328, 727]. Thread [3086, 3288, 2048, 528, 1821, 2179]. Threat [2770]. Threats [2894, 3044, 3003]. Three [487, 1607, 227]. Threshold [1090, 2589, 882, 1059, 1089, 1445, 1970, 1533]. Threshold-Based [2589, 882, 1059, 1089]. Thresholds [3159, 2051]. Throughput [3279, 3188, 3096, 3329, 1500, 2677, 3200, 374, 2772, 2588, 2858, 3281, 2913, 2121, 589, 938, 977, 2142, 1275, 621, 1469, 1420, 2413, 920, 1470, 809, 1553, 1468, 1727, 1385, 1236, 1255, 2244, 1697, 2788, 1256, 2213]. Throughput-delay [374]. throughput-fairness [1697]. throughput-oriented [621]. Throughputs [2622, 1596, 1992, 1901]. Throw [3408]. Tie [2760]. Tie-Line [2760]. tier [1222, 1781, 1851, 1427, 1406, 1915]. Tier-1 [1915]. Tiered [2642, 1332, 1351, 1624, 1364]. tiers [2180]. Tight [1969]. Tightening [3447]. Tighter [3027, 1773]. tightest [1326]. Time [2474, 3008, 370, 382, 3117, 3452, 2650, 3342, 3112, 3140, 3433, 3181, 2551, 2789, 2751, 776, 3180, 3009, 3162, 2649, 2705, 3235, 3027, 2940, 2787, 601, 2637, 3109, 2830, 1794, 2536, 3344, 3239, 1221, 729, 1155, 830, 1269, 730, 216, 1564, 453, 1863, 2346, 350, 1482, 1997, 2234, 631, 856, 1613, 2411, 1203, 2014, 111, 1078, 731, 1144, 2400, 1807, 1061, 298, 77, 80, 807, 2126, 1252, 2310, 1413, 862, 1249, 839, 1187, 2299, 2581, 1396, 1497, 1625, 2153, 666, 1383, 207, 2067, 2379, 201, 1453, 433, 883, 615, 1892, 1565, 2086, 1109, 1447, 68, 679, 2087, 809, 1693, 899, 934]. time [542, 3385, 1895, 1590, 1195, 1486, 3150, 356, 543, 876, 1446, 1604, 1978, 1803, 73, 269, 752, 1653, 33, 439, 2363, 1527, 414, 708, 254, 3303, 668, 1395, 744, 1823, 881, 59, 1459, 1493, 1081, 685, 773, 622]. Time-Average [3342]. Time-Based [2637]. time-correlated [1604]. time-critical [254]. time-dependent [809]. time-memory [2014]. Time-parallel [776]. Time-Series [3112]. time-sharing [68, 269, 33]. time-sharing/batch [68]. time-stamping [1109]. Time-Varying [3433, 2551, 2789, 2751, 2940, 2474, 3181, 1564, 2234, 1565, 2086, 2087]. time-windows [1497, 1625]. timed [1373, 2046, 429]. Timeline [2669]. Timelines [2931, 3061, 2669]. Timely [3362, 3364]. Timepatch [835]. Timer [3442]. Timer-based [3442]. Times [2970, 3343, 3373, 3043, 3315, 1592, 718, 2298, 1355, 925, 2412, 384, 2233, 453, 1056, 2200, 1807, 3184, 2750, 3103, 622, 2302, 2053, 1596, 914, 3104, 2289, 2414, 1251]. timescale [2754, 2794, 2798]. Timescales [2744]. timesharing [93]. Timing [2835, 1750, 1190, 1180, 20]. timing-first [1190]. TIPME [1063]. TLB [1192]. Token [382, 491, 381, 1034, 439]. tolerance [801]. Tolerant [2842, 898, 1771, 2337, 577, 448]. Tolerating [2273, 2849]. Tomographic [2838]. Tomography [2499, 2775, 2776, 1982, 1182, 2153, 1454, 2137]. Too [2956, 2702]. Tool [3355, 3217, 110, 3436, 2448, 559, 1744, 1741, 1509, 814, 3224, 1839, 2091, 1746, 118, 1617, 36, 1851, 326, 250, 475, 1748, 322, 3223, 335, 960, 266, 324, 2016, 142, 1278, 3216]. Toolbox [3248, 1987, 1712, 2090]. Toolkit [3125, 3224, 1928]. Tools [1277, 2801, 1742, 2089, 83, 1740, 210, 120, 802, 734, 3214]. Top [2674]. Topic [2395]. Topics [2942]. Topological [1758, 2202]. Topologies [2864, 3192, 1182, 1217, 1826, 2413, 2237, 1583]. Topology [1958, 3200, 2846, 1014, 1181, 1409, 1739, 2319, 2010, 1427]. TOSME [3214]. Total [2646, 2705, 943, 2012]. Total-Cost [2646]. TPC [660, 699, 790]. TPC-C [699]. TPC-D [790]. Trace [3155, 714, 1927, 974, 552, 1664, 1300, 993, 996, 725, 877, 1335, 734, 532, 35, 752, 267, 550, 620, 1422]. trace-compression [1300]. trace-driven [725, 35, 550]. trace-sample [620]. traces [1341, 2178, 1193, 2365, 674, 752, 530]. Tracing [110, 551, 578, 983]. track [355]. Tracking [1656, 3131, 3331, 2255, 1452, 1466]. tractable [1599]. Trade [2824, 3158, 2936, 2776, 2877, 2509, 1949, 1672, 2924, 2014, 1336, 2349]. Trade-off [2776, 1949]. Trade-Offs [2824, 3158, 2936, 2877, 1672, 2014, 1336, 2349]. Tradeoff [2981, 2900, 2536, 2007, 2041, 1697, 2311, 1855]. Tradeoffs [2032, 2520, 2586, 3361, 898, 2351, 806, 2238, 1812, 1559, 1822, 1204]. Trading [3423, 3271, 2639]. Traffic [3126, 2500, 1680, 3072, 3108, 3276, 3452, 2970, 3327, 581, 2668, 3091, 2902, 3265, 3376, 3375, 3429, 3424, 2549, 2587, 2803, 2391, 3061, 2974, 2959, 2452, 3064, 2503, 1412, 1452, 3194, 2800, 2846, 3109, 1007, 2712, 1238, 3344, 3390, 3240, 2715, 2860, 3283, 2093, 1915, 1314, 1755, 2133, 1994, 1686, 1209, 2285, 409, 1147, 1475, 1540, 1995, 1171, 1112, 1207, 1237, 834, 1695, 2411, 933, 853, 2148, 2287, 1118, 1009, 1585, 824, 2149, 1252, 1253, 2151, 460, 1580, 1299, 1114, 872, 3073, 1974, 2417, 1008, 1396, 1772, 1756, 1872, 1552, 3062, 1801, 1383, 806, 798, 1290, 1478, 2139, 3028]. traffic [619, 1892, 1810, 1426, 1139, 1233, 1615, 1284, 1402, 1380, 438, 2147, 934, 998, 1895, 2263, 970, 1554, 1333, 1692, 1959, 2106, 1162, 2242, 1893, 1885, 1321, 1291, 287, 1021, 2640, 848, 1945, 1377, 1754, 2425, 1792, 1634, 1782, 2317, 1259, 799, 2399, 2305, 2772, 1574, 1214, 1411, 1477, 2855, 2913, 2968, 3197, 1092]. Traffic-driven [2846]. Training [3406, 3296, 3295]. trajectory [2404]. Trans [588]. Transaction [750, 605, 657, 653, 3102, 631, 694, 676, 670]. Transactional [94, 2388, 2013]. transactions [1656, 3103]. Transfer [3261, 2800, 987, 529, 1795, 1295, 356, 1920, 744]. transfers [1771, 420]. Transform [3445, 2336]. Transformation [478, 912, 3215, 3243, 1758]. Transformations [2514]. Transformer [3406]. Transient [2988, 2737, 2479, 2917, 2786, 923, 2725, 1481, 1092, 1158, 1511, 1878, 723, 1455, 819]. Transiently [3071, 2609]. Transit [3150, 2167, 1693, 914]. Transition [3449, 2297]. transitions [2076, 2046, 512]. translation [1879]. Transmission [2506, 340, 2050, 2234, 2044, 1353, 1535, 1050, 1033, 557, 1674, 483]. transmissions [1789]. Transparency [2520]. Transparent [1853, 2365, 1769]. transplant [2059]. transplay [1946]. Transport [3424, 351, 1631, 1295, 2391, 1529, 2221]. Transportation [2768]. transposes [2173]. transposition [1354]. transputer [967, 568]. Trap [2249]. TRAPEDS [530]. treatment [467]. Tree [2506, 2895, 2901, 2516, 1382, 536, 864, 371, 482, 2247, 1481, 1574]. tree-like [1481]. tree-RECAL [536]. Tree-structured [2506]. treeness [1757]. Trees [2600, 2061, 2062, 1589, 2095, 2371]. trend [245]. trends [2093, 1149, 2947, 703, 1984, 2954]. trials [2330]. triangles [1682]. Triangular [1098]. trip [1453]. Tripartite [3077, 3208]. trisection [1800]. trivial [8]. trouble [1231]. Troubleshooting [3407, 1416, 1983, 1324]. true [1561]. Truncation [3130, 3027]. Trust [2946, 3258]. Truth [3154, 1663]. Truthful [2438, 2664, 2558]. TS [2405]. TS-CLOCK [2405]. TSS [59]. TTL [2713, 2412, 900, 2595, 2973]. TTL-Based [2713, 2973]. ’Tube [1971]. Tunable [2371]. Tune [1026]. Tuning [83, 559, 1795, 521, 293, 1001, 504, 144, 1142, 510, 1327, 1621, 1031]. Tunnel [717]. tuple [1456]. turbocharging [2203]. TurboSMARTS [1430]. tussles [2166]. Tutorial [2278, 2735, 62, 426, 422, 423, 424, 425, 256, 427, 92]. TV [1598]. twenty [1749]. twin [1524, 1356]. twisting [1019]. Twitter [2368, 2602]. Two [1293, 1352, 2997, 2548, 2667, 1552, 1345, 2068, 2218, 3381, 3078, 3190, 3144, 3194, 2727, 2476, 3017, 2025, 1216, 2250, 247, 213, 209, 37, 406, 32, 2417, 920, 1496, 1070, 1836, 771, 2434, 2676, 1427, 353, 1749, 803, 1894, 2163]. Two-Class [2476, 1894]. two-computer [37]. Two-dimensional [2068, 209]. two-hop [2676]. Two-level [1293, 803]. two-node [2434]. Two-phase [1552]. Two-Sided [3144, 3194, 2727]. Two-Stage [2667, 32]. two-tier [1427]. two-way [920]. Type [3207, 1592, 2058, 2090, 2079, 2063, 2086, 1443, 2155, 1188, 1229, 1267, 696, 1169, 2289]. types [808, 2295].

U [608, 700]. U.S. [409]. ubiquitous [1465]. UBR [874]. Udo [746]. UDP [1256]. Ultra [2528, 2785, 3341]. Ultra-Dense [2528, 2785]. Ultra-Sharp [3341]. UML [1372]. unbalanced [431]. unbalancing [940]. Unbiased [3169, 2127]. Unboundedly [548]. Uncertain [3051, 3033, 3010, 3043, 2576, 3184, 1574]. Uncertainties [2703, 2310, 2187, 1130]. Uncertainty [3309, 2791, 2760, 2812, 3165, 1284, 1402, 2113]. Uncooperative [1308]. Uncoordinated [2451]. uncountable [2075]. uncountable-state [2075]. uncovered [1629]. undecidability [1240]. Under-price [3105]. Underload [2621]. underlying [2258, 2306]. understand [1965]. Understanding [2275, 3232, 2613, 2718, 1767, 2274, 2650, 1230, 2150, 2276, 3431, 1124, 1486, 1551, 2375, 2511, 3011, 2258, 1974, 1812, 2245]. undesirable [1754]. undetectable [2767]. Undirected [3311]. uneconomic [2670]. Unfairness [3427, 1129, 1446]. unicast [1181, 1652]. unidirectional [535]. unification [315]. Unified [382, 3250, 644, 2404, 3050, 701, 467, 1562, 1268, 2249, 1865]. Uniform [3045, 2109, 2549]. uniformization [724]. unifying [2935, 1842]. Unique [3040]. uniqueness [1385]. Unit [2438, 14, 72, 142]. units [68]. Univac [257]. universal [1274, 2117, 1962, 2270]. Universality [2651]. University [95, 11]. UNIX [651, 794, 454, 636, 380, 828, 366]. Unknown [3204, 3316, 2964, 3064, 3314, 1215, 1088, 2380]. Unleashing [2390]. unnecessary [2044]. Unobservable [3186]. unrecoverable [1503]. Unreliable [3107, 3402, 3164, 3313, 1512, 1438]. Unscaled [2503]. unstructured [124, 1206, 1428]. Unsupervised [2598, 1880]. unsupervised-learning-based [1880]. Unveiling [1756, 3230]. Update [3071, 278, 3361, 1604]. Updates [2609, 3161, 1074, 2129, 847]. Updating [3213, 3198, 297]. Upgrades [2619]. upload [1178, 1508, 1076]. upon [2152, 3120, 2557]. upper [294, 2076]. Uptime [2193]. urban [2314]. Urgent [3067]. URL [2957]. URLs [3414]. Urn [2611]. Usability [2916]. Usage [2038, 3112, 52, 804, 43]. Use [95, 757, 2093, 216, 1099, 1237, 74, 993, 662, 1282, 1561, 510, 1446, 1423, 2342, 541, 113, 97, 109, 1477]. used [130, 982]. useful [919, 1142]. User [2570, 3408, 2914, 70, 2786, 2535, 2960, 2538, 1146, 1079, 2258, 1198, 79, 11, 1914, 1802, 784, 1350, 1579, 1362, 177, 402, 2375, 1884, 1888, 1153, 2327, 2163, 2401, 1862]. user-centric [1350, 1579]. user-generated [2327]. User-Level [3408, 784, 1153]. User-Provided [2570]. Users [3322, 2878, 2234, 1582, 1508, 1909]. Using [945, 3354, 1141, 3419, 2484, 2603, 2791, 3360, 3100, 1000, 1910, 3060, 2835, 1923, 3238, 3094, 1639, 883, 3282, 126, 3054, 2958, 1060, 2922, 3102, 2959, 3009, 2966, 1327, 2553, 2722, 2984, 3281, 2159, 2130, 3225, 1656, 1155, 338, 697, 244, 1686, 1373, 1927, 1074, 1997, 1985, 555, 1937, 408, 1341, 1613, 1359, 1560, 197, 843, 829, 410, 1261, 1063, 237, 1852, 2036, 2184, 1569, 1200, 525, 877, 3111, 18, 1872, 3418, 1019, 875, 1625, 2365, 1068, 852, 1454, 192, 490, 1285, 1902, 2173, 1513, 1380, 1693, 2255, 724, 240, 1635, 876, 2404, 970, 561, 1408, 1777, 1882, 2048, 1911]. using [682, 2229, 2403, 1885, 1220, 763, 1452, 913, 2143, 388, 2020, 801, 1952, 429, 2939, 668, 744, 1329, 2124, 262, 3383]. Utilities [3316, 2695, 1399, 355]. Utility [1659, 3204, 3449, 2742, 2439, 889, 973, 1399, 1288, 212, 1447, 1350]. Utilization [2700, 1832, 498, 43, 2335, 2051, 1769, 1116, 19]. utilizing [1154].

V [480, 74, 465]. v1.3 [3220]. Vacation [647, 2056]. vacations [1691]. validation [729, 1002, 1280, 747, 1200, 1710, 281, 35]. validity [1900, 1239]. valleys [2252]. Value [3000, 2526, 2903, 2982, 3210, 2607, 2105, 621, 43, 506, 1858, 514, 761, 2329, 1005, 1175, 985, 441, 128, 979, 743, 2145, 262, 369]. value-added [985]. Values [2589, 2716]. VANET [2710]. Variability [3286, 2217, 1061, 1776, 1470, 859, 2073]. Variable [2817, 1552, 1119, 1162, 190]. Variance [2544, 3158, 2929, 2009, 2880, 1446]. variate [2063]. variates [2065]. Variation [2613, 2720, 2849, 2241, 1348, 1398, 235, 246, 603]. variational [1902]. variations [560, 1653]. Varying [3433, 2551, 2789, 2751, 2940, 2474, 1564, 2234, 2126, 1565, 2086, 3181, 2087]. VAX [144, 267, 254]. VAX/VMS [144, 254]. VBR [806, 798, 886, 1048, 903]. vector [463, 2014, 580, 590, 2017, 682]. vector-descriptor [2014]. vectors [852]. Vegas [1124, 1236, 1255, 1259]. Vehicle [3147, 2984, 3135, 2189, 2746, 3165]. Vehicles [2759]. Vehicular [2961, 2903, 2209]. Verification [3428, 1371, 1367]. Verlag [643, 607, 654, 705]. Versatile [2121, 1830]. version [1281, 2013]. versioning [676]. versions [2220]. Versus [3159, 1307, 1665, 2455, 666, 682, 1998]. vertices [2207]. very [415, 1460]. via [3155, 3008, 2601, 2446, 2643, 2489, 3276, 1203, 1807, 2335, 2236, 1864, 2496, 2641, 2716, 2890, 536, 548, 3334, 2895, 2901, 1258, 2775, 3412, 2738, 2774, 2690, 1650, 1957, 1453, 1970, 2256, 3424, 2745, 3300, 1522, 1468, 2972, 2806, 2012, 1054, 2257, 530, 1888, 708, 1328, 2440, 3211, 2125, 2999, 2878, 3283]. Viability [2606]. vibration [2312]. vibration-induced [2312]. Video [2630, 3072, 3108, 3179, 2673, 2777, 2533, 3417, 2029, 857, 2275, 1132, 1713, 781, 805, 858, 796, 1789, 1050, 806, 2943, 1990, 1048, 903, 1323, 859, 2375, 897, 812]. Video-based [3179]. video-on-demand [2029, 857, 796, 897, 812]. videos [2047]. View [3438, 2902, 280, 319, 720, 1912, 457, 2268, 1736, 311, 1121]. view-probability-matrix [1736]. Viewing [1789, 2275]. views [226, 2155]. VIII [87]. Violations [3181, 3135]. Viral [2489]. Virtual [2682, 2500, 3119, 2681, 1064, 1814, 3392, 3093, 270, 2501, 2468, 2606, 44, 992, 38, 28, 737, 1696, 877, 974, 193, 288, 1044, 906, 135, 168, 717, 2670, 300, 1435, 2143, 1824, 2317]. virtual-disk [906]. Virtualization [2171, 3002, 3003, 1910, 1286, 1429]. Virtualized [2836, 3416, 3419, 3004, 2832, 1825, 1983, 2214, 1967, 1701]. viruses [1857]. VirusTotal [3414]. visibility [2333]. Vision [3033]. Vision-based [3033]. visualise [814]. Visualization [2564, 645, 591, 1715, 1042]. Visualizing [793]. VM [2409, 1917, 2834]. VMS [144, 254, 2988]. VMTP [480]. VOD [898, 3424, 2196]. voice [407, 460, 1139, 1698, 318, 1204]. voice-data [318]. voice/data [407, 460]. Volatile [2704, 2241]. Volatility [2598]. Voltage [2718, 2939, 1675, 1107, 1876]. Volume [647, 173, 1174]. volunteer [1738]. Voting [3432, 554]. VPC3 [1300]. VPN [1650]. VPNs [1336]. VR [2837]. vs [2441, 1216, 2674, 3408, 2490, 1631, 2169, 2119, 2742, 2154, 3233]. Vulnerability [2897, 2697, 3040, 2693, 2188, 1504, 1886]. Vulnerable [3405].

W [655, 746]. WAIN [3106, 3226]. wait [677]. Waiting [453, 3140, 856, 807]. Waiting-time [453]. waits [1225]. wake [888]. Walk [2971, 2127, 2418]. Walking [1957]. Walks [3245, 2126, 1101, 1240, 1957, 1978, 2435, 1891]. Walter [9]. WAN [1220]. warehouses [1006]. Warp [685, 773, 622]. wars [1618]. Water [3042]. watermarks [211]. Watson [789]. WattScope [3379]. Wave [3277]. wavelength [2110]. wavelets [970]. Way [2585, 1490, 2151, 1578, 920, 1693, 3317]. WCSF [298]. WDM [1118]. Weak [1260, 2922]. Weak-Hashing-based [2922]. Weakly [2830, 1083]. Weakness [3041]. Wear [3011]. Weaving [3000]. Web [1071, 2326, 3126, 1221, 1222, 1155, 850, 1010, 1079, 1024, 927, 1011, 1073, 1074, 957, 1151, 1083, 1154, 900, 1025, 853, 1150, 1078, 2520, 986, 1271, 1080, 872, 1974, 1045, 2916, 1332, 1196, 871, 941, 960, 2594, 3304, 1195, 1051, 2154, 3009, 962, 1220, 1126, 1021, 1102, 1174, 2277, 1125, 1272, 1075, 1152, 1628, 3109, 1091, 1238, 1039, 988, 1031]. Web-based [1221]. WebRTC [2777]. WebRTC-based [2777]. Websites [2566]. Wechat [3124]. Weight [3091, 1621]. Weighted [1696, 3047, 1948, 2705, 2920, 2207, 1826, 1760, 1957, 554, 1860]. weights [1686, 1472, 1792]. Welfare [2470, 2163]. well [829, 2112]. well-formed [829]. Wesley [606, 608, 645, 700]. Westmere [1925]. What-If [3009, 2215]. Where [3123, 1312, 2116, 1829]. Which [2739, 2075, 139]. while [740, 1770]. White [2438]. Whitt [2078, 2881, 3325]. Whittle [2474, 2823, 3172]. Who [3413, 3097, 2602, 2560, 1779, 2771]. Wi [3329, 2839]. Wi-Fi [3329, 2839]. Wide [853, 1150, 1075, 3413, 870, 1011, 1076, 2552, 1004, 1550, 1756, 1201, 1810, 1127, 2147, 1715, 1938]. wide-area [870, 1076, 1201, 1127, 2147, 1715]. widely [2240]. Wiener [3402]. WiFi [3155, 2429, 2135]. Wilbur [605, 653]. Wild [2481, 3431, 2275, 1768, 1666]. Wiley [641, 595, 597, 604]. Wiley-Interscience [595]. Will [2633, 312]. Wind [717, 2187]. Window [3083, 2346, 298, 29, 482, 585, 583, 695]. windowing [954]. windows [1497, 1625]. Wins [3097]. WIP [2899]. Wire [1567]. wired [1320]. Wireless [2823, 1319, 2816, 2914, 2634, 3328, 2783, 3365, 2476, 3353, 1212, 1687, 984, 1716, 1198, 1869, 1762, 1623, 1609, 1083, 1986, 1676, 2226, 1844, 2293, 1518, 1910, 1113, 1431, 1320, 1552, 1874, 1298, 1049, 1470, 1387, 1264, 1085, 1555, 1856, 983, 1611, 1263, 1674, 2424, 1468, 1520, 1854, 1715, 911, 2097, 1416, 2011, 2010, 1440, 1652, 1135, 1764, 1909, 2530, 1404, 1199, 2020, 1084, 1651, 2389, 1846, 2430, 1855]. wirespeed [1707]. Wisconsin [717]. WiSE [2866]. withholding [2952]. within [749]. Without [3343, 3373, 2994, 3001, 2816, 1567, 1069, 1790, 3265, 3375, 1180, 2363]. WLANS [1680, 1384, 2457, 1385, 2535]. WOJ [2922]. Wolfe [655]. Work [2446, 2899, 2884, 939, 1841, 85, 68, 572, 573, 2360, 2035]. Work-in-Progress [2899]. workbench [301]. Workflows [3426, 3256, 2465]. Working [2832, 563, 169, 259, 29, 1066, 145, 165, 170]. working-set [1066]. Workload [3457, 1010, 2105, 994, 245, 423, 3012, 1790, 3411, 2969, 2987, 332, 1205, 1007, 477, 850, 659, 617, 334, 401, 1646, 1985, 405, 2191, 965, 781, 365, 782, 478, 333, 868, 88, 503, 1211, 2115, 2256, 637, 711, 2289, 2145, 2272, 1712, 2527]. Workload-DRAM [3012]. Workloads [2988, 3191, 3060, 825, 927, 1564, 1706, 1861, 2274, 421, 841, 1825, 1776, 1156, 1335, 1831, 1803, 1121]. works [140]. Workshop [3066, 611, 2436, 3214, 3170, 3246, 3454, 84, 2876, 3016, 3053, 3156, 3234, 3332, 3439, 2892, 2030, 164, 1433, 99, 1071, 1012, 1087, 1157, 1227, 456, 1916]. Workstation [611, 973, 1200]. workstations [889, 825]. World [3299, 3393, 1700, 1821, 790, 853, 1150, 1075]. worry [1933]. Worst [2921, 3056, 3236, 3415]. Worst-case [2921, 3236]. Worth [2675, 1231]. would [2062]. WPI [650]. WPIN [2436]. Write [2976, 3419, 2922, 714, 2241, 199, 1191, 355, 628]. Write-Once [2922]. Write-Rationing [2976]. writer [558]. Wroclaw [95]. Wrong [2571, 243]. WSN [1920]. WSNs [2000]. WWW [997, 1127].

x [3328, 780, 695, 637]. x-haul [3328]. X-MP [637]. x86 [888]. XA [521]. XACML [1667]. Xaminer [3436]. Xen [1607]. Xengine [1667]. XML [1242]. XML-RPC [1242]. Xprof [695]. XRAY [180].

Yes [3395]. YOUQMON [2286]. YouTube [2286, 2960].

Z [809]. Z-iteration [809]. Zephyr [3352]. Zero [3097, 3100, 3352, 3178, 3258, 3121, 3127]. Zero-Knowledge [3352]. Zero-Queueing [3178]. Zero-Rating [3097]. Zigbee [2006]. Zipf [1580]. Zipf-like [1580]. zone [1304]. zone-balancing [1304].

References

Keirstead:1972:STC

[1] Ralph E. Keirstead and Donn B. Parker. Software testing and certification. ACM SIGMETRICS Performance Evaluation Review, 1(1):3–8, March 1972. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bell:1972:CME

[2] Thomas E. Bell. Computer measurement and evaluation: artistry, or science? ACM SIGMETRICS Performance Evaluation Review, 1(2):4–10, June 1972. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Palme:1972:BGM

[3] Jacob Palme. Beware of the Gibson mix. ACM SIGMETRICS Performance Evaluation Review, 1(2):10–11, June 1972. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Johnson:1972:SST

[4] Robert R. Johnson. Some steps toward an information system performance theory. ACM SIGMET-RICS Performance Evaluation Review, 1(3):4–15, September 1972. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kernighan:1972:CAO

[5] B. W. Kernighan, P. J. Plauger, and D. J. Plauger. On comparing apples and oranges, or, my machine is better than your machine. ACM SIGMET-RICS Performance Evaluation Review, 1(3):16–20, September 1972. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lynch:1972:DDA

[6] W. C. Lynch. Do disk arms move? ACM SIGMETRICS Performance Evaluation Review, 1(4):3–16, December 1972. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Halstead:1973:LLM

[7] M. H. Halstead. Language level, a missing concept in information theory. ACM SIGMETRICS Performance Evaluation Review, 2(1):7–9, March 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Halstead:1973:EDP

[8] M. H. Halstead. An experimental determination of the ‘purity’ of a trivial algorithm. ACM SIGMETRICS Performance Evaluation Review, 2(1):10– 15, March 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Denning:1973:RSC

[9] Peter J. Denning. Review of ‘Statistical Computer Performance Evaluation’ by Walter Frieberger; Academic Press (1972). ACM SIGMETRICS Performance Evaluation Review, 2(1):16–22, March 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Svobodova:1973:CSN

[10] Liba Svobodova. Communications: Some notes on the Computer Synectics hardware monitor sum. ACM SIGMETRICS Performance Evaluation Review, 2(1):23–25, March 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ishida:1973:JSU

[11] Haruhisa Ishida and Nobumasa Takahashi. Job statistics at a 2000-user university computer center. ACM SIGMETRICS Performance Evaluation Review, 2(2):2–13, June 1973. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rice:1973:AMC

[12] Don R. Rice. An analytical model for computer system performance evaluation. ACM SIGMETRICS Performance Evaluation Review, 2(2):14–30, June 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kolence:1973:SE

[13] Kenneth W. Kolence. The software empiricist. ACM SIGMETRICS Performance Evaluation Review, 2(2):31–36, June 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kolence:1973:SUP

[14] Kenneth W. Kolence and Philip J. Kiviat. Software unit profiles & Kiviat figures. ACM SIGMETRICS Performance Evaluation Review, 2(3):2–12, September 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Denning:1973:WOA

[15] Peter J. Denning. Why our approach to performance evaluation is SDRAWKCAB. ACM SIGMETRICS Performance Evaluation Review, 2(3): 13–16, September 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Beck:1973:CSL

[16] Norman Beck and Gordon Ashby. On cost of static linking and loading of subprograms. ACM SIGMETRICS Performance Evaluation Review, 2(3): 17–20, September 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kolence:1973:SEE

[17] Ken Kolence. The software empiricist experimental disciplines & computer measurements. ACM SIGMET-RICS Performance Evaluation Review, 2(3):21–23, September 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hughes:1973:UHM

[18] James Hughes and David Cronshaw. On using a hardware monitor as an intelligent peripheral. ACM SIGMET-RICS Performance Evaluation Review, 2(4):3–19, December 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Svobodova:1973:MCS

[19] Liba Svobodova. Measuring computer system utilization with a hardware and a hybrid monitor. ACM SIGMET-RICS Performance Evaluation Review, 2(4):20–34, December 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wortman:1974:NHR

[20] David B. Wortman. A note on high resolution timing. ACM SIGMETRICS Performance Evaluation Review, 3(1): 3–9, March 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Snyder:1974:QSA

[21] Rowan Snyder. A quantitative study of the addition of extended core storage. ACM SIGMETRICS Performance Evaluation Review, 3(1):10–33, March 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Merrill:1974:TCA

[22] H. E. Barry Merrill. A technique for comparative analysis of Kiviat graphs. ACM SIGMETRICS Performance Evaluation Review, 3(1):34–39, March 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Peterson:1974:CSH

[23] Thomas G. Peterson. A comparison of software and hardware monitors. ACM SIGMETRICS Performance Evaluation Review, 3(2):2–5, June 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Syms:1974:BCT

[24] Gordon H. Syms. Benchmarked comparison of terminal support systems for IBM 360 computers. ACM SIGMET-RICS Performance Evaluation Review, 3(2):6–34, June 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Morris:1974:KGC

[25] Michael F. Morris. Kiviat graphs: conventions and ‘figures of merit’. ACM SIGMETRICS Performance Evaluation Review, 3(3):2–8, October 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lujanac:1974:NSB

[26] Paul L. Lujanac. A note on Syms’ benchmarked comparison. ACM SIG-METRICS Performance Evaluation Review, 3(3):9–10, October 1974. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Graham:1974:MPB

[27] G. Scott Graham and Peter J. Denning. Multiprogramming and program behavior. ACM SIGMETRICS Performance Evaluation Review, 3(4):1–8, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brandwain:1974:MPV

[28] A. Brandwain, J. Buzen, E. Gelenbe, and D. Potier. A model of performance for virtual memory systems. ACM SIGMETRICS Performance Evaluation Review, 3(4):9, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Henderson:1974:OCW

[29] Greg Henderson and Juan Rodriguez-Rosell. The optimal choice of window sizes for working set dispatching. ACM SIGMETRICS Performance Evaluation Review, 3(4):10–33, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Denning:1974:CLP

[30] Peter J. Denning. Comments on a linear paging model. ACM SIGMET-RICS Performance Evaluation Review, 3(4):34–48, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brice:1974:FCR

[31] Richard S. Brice and J. C. Browne. Feedback coupled resource allocation policies in the multiprogrammingmultiprocessor computer system. ACM SIGMETRICS Performance Evaluation Review, 3(4):49–53, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Halachmi:1974:CCT

[32] Baruch Halachmi and W. R. Franta. A closed, cyclic, two-stage multiprogrammed system model and its diffusion approximation solution. ACM SIGMETRICS Performance Evaluation Review, 3(4):54–64, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schwetman:1974:ATS

[33] H. D. Schwetman. Analysis of a timesharing subsystem (a preliminary report). ACM SIGMETRICS Performance Evaluation Review, 3(4):65–75, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reiser:1974:ASC

[34] M. Reiser and A. G. Konheim. The analysis of storage constraints by a queueing network model with blocking. ACM SIGMETRICS Performance Evaluation Review, 3(4):76–81, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schatzoff:1974:SVT

[35] M. Schatzoff and C. C. Tillman. Statistical validation of a trace-driven simulator. ACM SIGMETRICS Performance Evaluation Review, 3(4):82–93, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ferrari:1974:GPS

[36] Domenico Ferrari and Mark Liu. A general-purpose software measurement tool. ACM SIGMETRICS Performance Evaluation Review, 3(4):94–105, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Foley:1974:EDD

[37] James D. Foley and John W. McInroy. An event-driven data collection and analysis facility for a twocomputer network. ACM SIGMET-RICS Performance Evaluation Review, 3(4):106–120, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Batson:1974:MVM

[38] A. P. Batson and R. E. Brundage. Measurements of the virtual memory demands of Algol-60 programs (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 3(4): 121–126, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sebastian:1974:HHE

[39] Peter R. Sebastian. HEMI (Hybrid Events Monitoring Instrument). ACM SIGMETRICS Performance Evaluation Review, 3(4):127–139, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cox:1974:IAC

[40] Springer W. Cox. Interpretive analysis of computer system performance. ACM SIGMETRICS Performance Evaluation Review, 3(4):140–155, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Noe:1974:DYC

[41] J. D. Noe and N. W. Runstein. Develop your computer performance pattern. ACM SIGMETRICS Performance Evaluation Review, 3(4):156– 165, December 1974. CODEN ????

ISSN 0163-5999 (print), 1557-9484 (electronic).

Brotherton:1974:CCC

[42] D. E. Brotherton. The computer capacity curve — a prerequisite for computer performance evaluation and improvement. ACM SIGMETRICS Performance Evaluation Review, 3(4): 166–179, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Erikson:1974:VCU

[43] Warren J. Erikson. The value of CPU utilization as a criterion for computer system usage. ACM SIGMETRICS Performance Evaluation Review, 3(4): 180–187, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Badel:1974:AOP

[44] M. Badel, E. Gelenbe, J. Leroudier, D. Potier, and J. Lenfant. Adaptive optimization of the performance of a virtual memory computer. ACM SIGMETRICS Performance Evaluation Review, 3(4):188, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kimbleton:1974:BCS

[45] Stephen R. Kimbleton. Batch computer scheduling: a heuristically motivated approach. ACM SIGMETRICS Performance Evaluation Review, 3(4): 189–198, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sharp:1974:APD

[46] Joseph C. Sharp and James N. Roberts. An adaptive policy driven scheduler. ACM SIGMETRICS Performance Evaluation Review, 3(4):199– 208, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Merrill:1975:FCC

[47] H. W. Barry Merrill. Further comments on comparative evaluation of Kiviat graphs. ACM SIGMETRICS Performance Evaluation Review, 4(1):1–10, January 1975. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stevens:1975:NFM

[48] Barry A. Stevens. A note on figure of merit. ACM SIGMETRICS Performance Evaluation Review, 4(1):11–19, January 1975. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bell:1975:MCP

[49] Thomas E. Bell. Managing computer performance with control limits. ACM SIGMETRICS Performance Evaluation Review, 4(1):21–28, January 1975. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Browne:1975:AMP

[50] J. C. Browne. An analysis of measurement procedures for computer systems. ACM SIGMETRICS Performance Evaluation Review, 4(1):29–32, January 1975. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Terplan:1975:COR

[51] Kornel Terplan. Cost-optimal reliability of data processing systems. ACM SIGMETRICS Performance Evaluation Review, 4(2):1–12, April 1975. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Landwehr:1975:USM

[52] Carl E. Landwehr. Usage statistics for MTS. ACM SIGMETRICS Performance Evaluation Review, 4(2):13–23, April 1975. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reddy:1975:EEM

[53] Y. V. Reddy. Experimental evaluation of a multiprogrammed computer system. ACM SIGMETRICS Performance Evaluation Review, 4(2):24–32, April 1975. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bhandarkar:1975:PAM

[54] Dileep P. Bhandarkar. A practical application of memory interference models. ACM SIGMETRICS Performance Evaluation Review, 4(2):33–39, April 1975. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bahr:1975:NFM

[55] Dieter Bahr. A note on figures of merit. ACM SIGMETRICS Performance Evaluation Review, 4(3):1–3, July 1975. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Boehm:1975:ICP

[56] B. W. Boehm and T. E. Bell. Issues in computer performance evaluation: some consensus, some divergence. ACM SIGMETRICS Performance Evaluation Review, 4(3):4–39, July 1975. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Barber:1975:BC

[57] Eric Ole Barber, Arne Asphjell, and Arve Dispen. Benchmark construction. ACM SIGMETRICS Performance Evaluation Review, 4(4):3–14, October 1975. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marrevee:1975:MPP

[58] J. P. Marrevée. Measurements of the Philips P1400 multiprogramming system. ACM SIGMETRICS Performance Evaluation Review, 4(4):15–45, October 1975. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wright:1976:AET

[59] Linda S. Wright and William A. Burnette. An approach to evaluating time sharing systems: MH-TSS a case study. ACM SIGMETRICS Performance Evaluation Review, 5(1):8–28, January 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Calcagni:1976:SRK

[60] John M. Calcagni. Shape in ranking Kiviat graphs. ACM SIGMET-RICS Performance Evaluation Review, 5(1):35–37, January 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Eisenfeld:1976:IRH

[61] J. Eisenfeld, David R. Barker, and David J. Mishelvich. Iconic representation of the human face with computer graphics. ACM SIGMET-RICS Performance Evaluation Review, 5(1):38–39, January 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nutt:1976:TCS

[62] Gary J. Nutt. Tutorial: computer system monitors. ACM SIGMET-RICS Performance Evaluation Review, 5(1):41–51, January 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cotton:1976:SFP

[63] Ira W. Cotton. Some fundamentals of price theory for computer services. ACM SIGMETRICS Performance Evaluation Review, 5(1c):1–12, March 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Giammo:1976:DCP

[64] Thomas Giammo. Deficiencies in computer pricing structure theory. ACM SIGMETRICS Performance Evaluation Review, 5(1c):13–21, March 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kimbleton:1976:CPD

[65] Stephen R. Kimbleton. Considerations in pricing distributed computing. ACM SIGMETRICS Performance Evaluation Review, 5(1c):22–30, March 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kiviat:1976:BRG

[66] Philip J. Kiviat. A brief review of the GAO task group’s recommendations on management guidelines for pricing computer services in the federal government. ACM SIGMETRICS Performance Evaluation Review, 5(1c):71–83, March 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Morris:1976:PIP

[67] Michael F. Morris. Problems in implementing and processing computer charging schemes. ACM SIGMET-RICS Performance Evaluation Review, 5(1c):84–88, March 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Luderer:1976:CPM

[68] Gottfried W. R. Luderer. Charging problems in mixed time-sharing/batch systems: cross subsidization and invariant work units. ACM SIGMET-RICS Performance Evaluation Review, 5(1c):89–93, March 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Oatey:1976:STM

[69] David J. Oatey. SIGMETRICS technical meeting on pricing computer services. ACM SIGMETRICS Performance Evaluation Review, 5(1c):94– 102, March 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gutsche:1976:UE

[70] Richard H. Gutsche. User experience. ACM SIGMETRICS Performance Evaluation Review, 5(1c):103– 107, March 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anonymous:1976:PC

[71] Anonymous. Participant’s choice. ACM SIGMETRICS Performance Evaluation Review, 5(1c):108–122, March 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Luderer:1976:DCR

[72] Gottfried W. R. Luderer. Defining a computer resource unit. ACM SIGMETRICS Performance Evaluation Review, 5(2):5–10, April 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Roehr:1976:PIT

[73] K. Roehr and K. Niebel. Proposal for instruction time objectives. ACM SIGMETRICS Performance Evaluation Review, 5(2):11–18, April 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Collins:1976:PIC

[74] John P. Collins. Performance improvement of the CP-V loader through use of the ADAM hardware monitor. ACM SIGMETRICS Performance Evaluation Review, 5(2):63–67, April 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brandwajn:1976:SLI

[75] A. Brandwajn. Simulation of the load of an interactive system. ACM SIGMETRICS Performance Evaluation Review, 5(2):69–92, April 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coppens:1976:QER

[76] G. W. J. Coppens, M. P. F. M. van Dongen, and J. P. C. Kleijnen. Quantile estimation in regenerative simulation: a case study. ACM SIGMET-RICS Performance Evaluation Review, 5(3):5–15, Summer 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Estell:1976:HFRa

[77] Robert G. Estell. How fast is ‘realtime’? ACM SIGMETRICS Performance Evaluation Review, 5(3):16–18, Summer 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mills:1976:SMC

[78] Philip M. Mills. A simple model for cost considerations in a batch multiprocessor environment. ACM SIGMET-RICS Performance Evaluation Review, 5(3):19–27, Summer 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Buchanan:1976:IBM

[79] Irene Buchanan and David A. Duce. An interactive benchmark for a multiuser minicomputer system. ACM SIGMETRICS Performance Evaluation Review, 5(4):5–17, Fall 1976. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Estell:1976:HFRb

[80] Robert G. Estell. How fast is ‘realtime’? ACM SIGMETRICS Performance Evaluation Review, 5(4):18–20, Fall 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rafii:1976:SPR

[81] Abbas Rafii. Study of the performance of RPS. ACM SIGMETRICS Performance Evaluation Review, 5(4):21–38, Fall 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Price:1976:CQN

[82] Thomas G. Price. A comparison of queuing network models and measurements of a multiprogrammed computer system. ACM SIGMETRICS Performance Evaluation Review, 5(4):39–62, Fall 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Buzen:1976:TTT

[83] J. P. Buzen. Tuning: tools and techniques. ACM SIGMETRICS Performance Evaluation Review, 5(4):63–81, Fall 1976. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Spiegel:1977:WSA

[84] Mitchell G. Spiegel. Workshop summary: ‘applications of queuing models to ADP system performance prediction’. ACM SIGMETRICS Performance Evaluation Review, 6(1):13–33, Winter 1977. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hellerman:1977:TWF

[85] L. Hellerman. A table of work formulae with derivations and applications. ACM SIGMETRICS Performance Evaluation Review, 6(1):35–54, Winter 1977. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Allen:1977:NES

[86] R. C. Allen and S. R. Clark. A note on an empirical study of paging on an IBM 370/145. ACM SIGMETRICS Performance Evaluation Review, 6(1):55–62, Winter 1977. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Morrison:1977:ASC

[87] Robert L. Morrison. Abstracts from the 1977 SIGMETRICS/CMG VIII conference. ACM SIGMETRICS Performance Evaluation Review, 6(2):3–21, Spring 1977. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lazos:1977:FDW

[88] Constantine Lazos. Functional distribution of the workload of a linked computer system and its simulation. ACM SIGMETRICS Performance Evaluation Review, 6(3):5–14, Summer 1977. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Scheer:1977:COM

[89] A.-W. Scheer. Combination of an optimization model for hardware selection with data determination methods. ACM SIGMETRICS Performance Evaluation Review, 6(3):15–26, Summer 1977. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berinato:1977:AMT

[90] Terence Berinato. An analytical model of a teleprocessing system. ACM SIGMETRICS Performance Evaluation Review, 6(3):27–32, Summer 1977. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chanson:1977:SSA

[91] Samuel T. Chanson and Craig D. Bishop. A simulation study of adaptive scheduling policies in interactive computer systems. ACM SIGMET-RICS Performance Evaluation Review, 6(3):33–39, Summer 1977. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ziegler:1977:DST

[92] Kurt Ziegler. A data sharing tutorial. ACM SIGMETRICS Performance Evaluation Review, 6(4):3–7, Fall 1977. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Scott:1977:PDP

[93] Shirley E. Scott. Pricing D.P. products: a timesharing implementation. ACM SIGMETRICS Performance Evaluation Review, 6(4):8–12, Fall 1977. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sarzotti:1977:TTS

[94] Alain Sarzotti. Transactional terminal system on micro-processor: a method for identifying & modeling overall performance. ACM SIGMETRICS Performance Evaluation Review, 6(4):13–22, Fall 1977. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bazewicz:1977:UMP

[95] Mieczyslaw Bazewicz and Adam Peterseil. Use of modelling in performance evaluation of computer systems: a case of installations in the Technical University of Wroclaw. ACM SIGMETRICS Performance Evaluation Review, 6(4):22–26, Fall 1977. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Orchard:1977:NMC

[96] R. A. Orchard. A new methodology for computer system data gathering. ACM SIGMETRICS Performance Evaluation Review, 6(4):27–41, Fall 1977. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Underwood:1978:HPE

[97] Mark A. Underwood. Human performance evaluation in the use of federal computer systems: recommendations. ACM SIGMETRICS Performance Evaluation Review, 7(1–2):6– 14, Spring-Summer 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jain:1978:GSA

[98] Aridaman K. Jain. A guideline to statistical approaches in computer performance evaluation studies. ACM SIGMETRICS Performance Evaluation Review, 7(1–2):18–32, Spring-Summer 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anonymous:1978:PSQ

[99] Anonymous. Proceedings of the Software Quality and Assurance Workshop. ACM SIGMETRICS Performance Evaluation Review, 7(1–2):32, Spring-Summer 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Honig:1978:DPA

[100] Howard P. Honig. Data path analysis: analyzing large I/O environments. ACM SIGMETRICS Performance Evaluation Review, 7(1–2):34– 37, Spring-Summer 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sauer:1978:SRP

[101] C. H. Sauer and E. A. MacNair. Simultaneous resource possession in queueing models of computers. ACM SIGMET-RICS Performance Evaluation Review, 7(1–2):41–52, Spring-Summer 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pfau:1978:AQA

[102] Pamela R. Pfau. Applied quality assurance methodology. ACM SIGMET-RICS Performance Evaluation Review, 7(3–4):1–8, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bersoff:1978:SCM

[103] Edward H. Bersoff, Vilas D. Henderson, and Stan G. Siegel. Software configuration management. ACM SIGMETRICS Performance Evaluation Review, 7(3–4):9–17, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Glass:1978:CFL

[104] Robert L. Glass. Computing failure: a learning experience. ACM SIGMET-RICS Performance Evaluation Review, 7(3–4):18–19, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Woodmancy:1978:SQI

[105] Donald A. Woodmancy. A software quality improvement program. ACM SIGMETRICS Performance Evaluation Review, 7(3–4):20–26, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fujii:1978:CSA

[106] Marilyn S. Fujii. A comparison of software assurance methods. ACM SIGMETRICS Performance Evaluation Review, 7(3–4):27–32, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sukert:1978:EMA

[107] Alan N. Sukert and Amrit L. Goel. Error modelling applications in software quality assurance. ACM SIGMET-RICS Performance Evaluation Review, 7(3–4):33–38, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Duran:1978:TMP

[108] Joe W. Duran and John J. Wiorkowski. Toward models for probabilistic program correctness. ACM SIGMETRICS Performance Evaluation Review, 7(3– 4):39–44, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yin:1978:EUM

[109] B. H. Yin and J. W. Winchester. The establishment and use of measures to evaluate the quality of software designs. ACM SIGMETRICS Performance Evaluation Review, 7(3– 4):45–52, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pierce:1978:RTT

[110] Robert A. Pierce. A requirements tracing tool. ACM SIGMETRICS Performance Evaluation Review, 7(3– 4):53–60, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Davis:1978:RLP

[111] Alan M. Davis and Walter J. Rataj. Requirements language processing for the effective testing of real-time systems. ACM SIGMETRICS Performance Evaluation Review, 7(3–4):61– 66, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Peters:1978:RSR

[112] Lawrence Peters. Relating software requirements and design. ACM SIGMETRICS Performance Evaluation Review, 7(3–4):67–71, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stavely:1978:DFU

[113] Allan M. Stavely. Design feedback and its use in software design aid systems. ACM SIGMETRICS Performance Evaluation Review, 7(3–4): 72–78, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yoder:1978:NSC

[114] Cornelia M. Yoder and Marilyn L. Schrag. Nassi–Shneiderman charts an alternative to flowcharts for design. ACM SIGMETRICS Performance Evaluation Review, 7(3–4):79– 86, November 1978. CODEN ????

ISSN 0163-5999 (print), 1557-9484 (electronic).

Benson:1978:SQA

[115] J. P. Benson and S. H. Saib. A software quality assurance experiment. ACM SIGMETRICS Performance Evaluation Review, 7(3–4):87–91, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bauer:1978:AGE

[116] Jonathan Bauer, Susan Faasse, Alan Finger, and William Goodhue. The automatic generation and execution of function test plans for electronic switching systems. ACM SIGMET-RICS Performance Evaluation Review, 7(3–4):92–100, November 1978. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Martin:1978:SAT

[117] K. A. Martin. Software acceptance testing that goes beyond the book. ACM SIGMETRICS Performance Evaluation Review, 7(3–4): 101–105, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Drasch:1978:ITP

[118] Frederick J. Drasch and Richard A. Bowen. IDBUG: a tool for program development. ACM SIGMETRICS Performance Evaluation Review, 7(3–4): 106–110, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stickney:1978:AGT

[119] M. E. Stickney. An application of graph theory to software test data selection. ACM SIGMETRICS Performance Evaluation Review, 7(3–4): 111–115, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fischer:1978:SQA

[120] Kurt F. Fischer. Software quality assurance tools: Recent experience and future requirements. ACM SIGMET-RICS Performance Evaluation Review, 7(3–4):116–121, November 1978. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Glasser:1978:ESC

[121] Alan L. Glasser. The evolution of a Source Code Control System. ACM SIGMETRICS Performance Evaluation Review, 7(3–4):122–125, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Josephs:1978:MCB

[122] William H. Josephs. A mini-computer based library control system. ACM SIGMETRICS Performance Evaluation Review, 7(3–4):126–132, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cavano:1978:FMS

[123] Joseph P. Cavano and James A. Mc-Call. A framework for the measurement of software quality. ACM SIGMET-RICS Performance Evaluation Review, 7(3–4):133–139, November 1978. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cobb:1978:MSU

[124] Gary W. Cobb. A measurement of structure for unstructured programming languages. ACM SIGMETRICS Performance Evaluation Review, 7(3– 4):140–147, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bowen:1978:CAS

[125] John B. Bowen. Are current approaches sufficient for measuring software quality? ACM SIGMETRICS Performance Evaluation Review, 7(3– 4):148–155, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lockett:1978:UPM

[126] Joann Lockett. Using performance metrics in system design. ACM SIGMETRICS Performance Evaluation Review, 7(3–4):156–159, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Southworth:1978:RM

[127] Richard N. Southworth. Responding to MIL-S-52779. ACM SIGMETRICS Performance Evaluation Review, 7(3– 4):160–164, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tighe:1978:VPS

[128] Michael F. Tighe. The value of a proper software quality assurance methodology. ACM SIGMETRICS Performance Evaluation Review, 7(3– 4):165–172, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Belford:1978:QEE

[129] Peter Chase Belford and Carlo Broglio. A quantitative evaluation of the effectiveness of quality assurance as experienced on a large-scale software development effort. ACM SIGMETRICS Performance Evaluation Review, 7(3– 4):173–180, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kacik:1978:ESQ

[130] Paul J. Kacik. An example of software quality assurance techniques used in a successful large scale software development. ACM SIGMETRICS Performance Evaluation Review, 7(3–4): 181–186, November 1978. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kreutzer:1979:CSM

[131] Wolfgang Kreutzer. Computer system modelling and simulation. ACM SIG-METRICS Performance Evaluation Review, 8(1–2):9–35, Spring-Summer 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Turner:1979:ISM

[132] Rollins Turner. An investigation of several mathematical models of queueing systems. ACM SIGMETRICS Performance Evaluation Review, 8(1–2): 36–44, Spring-Summer 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sauer:1979:CIQ

[133] Charles H. Sauer. Confidence intervals for queueing simulations of computer systems. ACM SIGMETRICS Performance Evaluation Review, 8(1– 2):45–55, Spring-Summer 1979. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kleijnen:1979:NCS

[134] Jack P. C. Kleijnen. A note on computer system data gathering. ACM SIGMETRICS Performance Evaluation Review, 8(1–2):56, Spring-Summer 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rajaraman:1979:PPV

[135] M. K. Rajaraman. Performance prediction of a virtual machine. ACM SIGMETRICS Performance Evaluation Review, 8(1–2):57–62, Spring-Summer 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jain:1979:GSA

[136] Aridaman K. Jain. A guideline to statistical approaches in computer performance evaluation studies. ACM SIGMETRICS Performance Evaluation Review, 8(1–2):63–77, Spring-Summer 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schwartz:1979:DCC

[137] E. Schwartz. Development of credible computer system simulation models. ACM SIGMETRICS Performance Evaluation Review, 8(1–2):78– 95, Spring-Summer 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Clark:1979:CPE

[138] Jon D. Clark, Thomas J. Reynolds, and Michael J. Intille. Computer performance evaluation: an empirical approach. ACM SIGMETRICS Performance Evaluation Review, 8(1–2):97– 101, Spring-Summer 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Willis:1979:TSW

[139] Ron Willis. Techniques in simulation which enhance software reliability. ACM SIGMETRICS Performance Evaluation Review, 8(1–2):102– 115, Spring-Summer 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Blake:1979:TSM

[140] Russ Blake. Tailor: a simple model that works. ACM SIGMETRICS Performance Evaluation Review, 8(3):1– 11, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Strecker:1979:ACP

[141] William D. Strecker. An analysis of central processor-input-output processor contention. ACM SIGMETRICS Performance Evaluation Review, 8(3): 27–40, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wiecek:1979:PST

[142] Cheryl A. Wiecek and Simon C. Steely, Jr. Performance simulation as a tool in central processing unit design. ACM SIGMETRICS Performance Evaluation Review, 8(3):41–47, Fall 1979. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bennett:1979:SDS

[143] David A. Bennett and Christopher A. Landauer. Simulation of a distributed system for performance modelling. ACM SIGMETRICS Performance Evaluation Review, 8(3):49–56, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lazowska:1979:BTA

[144] Edward D. Lazowska. The benchmarking, tuning and analytic modeling of VAX/VMS. ACM SIGMETRICS Performance Evaluation Review, 8(3):57– 64, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marshall:1979:AMW

[145] William T. Marshall and C. Thomas Nute. Analytic modelling of “working set like” replacement algorithms. ACM SIGMETRICS Performance Evaluation Review, 8(3):65–72, Fall 1979. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Briggs:1979:EBM

[146] Fayé A. Briggs. Effects of buffered memory requests in multiprocessor systems. ACM SIGMETRICS Performance Evaluation Review, 8(3):73–81, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raffi:1979:ECB

[147] Abbas Raffi. Effects of channel blocking on the performance of shared disk pack in a multi-computer system. ACM SIGMETRICS Performance Evaluation Review, 8(3):83–87, Fall 1979. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zahorjan:1979:ESM

[148] John Zahorjan. An exact solution method for the general class of closed separable queueing networks. ACM SIGMETRICS Performance Evaluation Review, 8(3):107–112, Fall 1979.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kienzle:1979:SAQ

[149] Martin G. Kienzle and K. C. Sevcik. Survey of analytic queueing network models of computer systems. ACM SIGMETRICS Performance Evaluation Review, 8(3):113–129, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Landry:1979:SEP

[150] Steve P. Landry and Bruce D. Shriver. A simulation environment for performing dataflow research. ACM SIGMET-RICS Performance Evaluation Review, 8(3):131–139, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Langan:1979:SED

[151] David D. Langan and Bruce D. Shriver. Simulated execution of dataflow programs on processors having finite resources. ACM SIGMETRICS Performance Evaluation Review, 8(3):141– 149, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Unger:1979:OSI

[152] Brian W. Unger and James R. Parker. An operating system implementation and simulation language (OA-SIS). ACM SIGMETRICS Performance Evaluation Review, 8(3):151– 161, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sanguinetti:1979:TIS

[153] John Sanguinetti. A technique for integrating simulation and system design. ACM SIGMETRICS Performance Evaluation Review, 8(3):163– 172, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Razouk:1979:EMS

[154] Rami R. Razouk, Mary Vernon, and Gerald Estrin. Evaluation methods in SARA — the graph model simulator. ACM SIGMETRICS Performance Evaluation Review, 8(3):189–206, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yu:1979:MSD

[155] Stone H. Yu and Tadao Murata. Modeling and simulating data flow computations at machine language level. ACM SIGMETRICS Performance Evaluation Review, 8(3):207– 213, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mattheyses:1979:MSA

[156] R. M. Mattheyses and S. E. Conry. Models for specification and analysis of parallel computing systems. ACM SIGMETRICS Performance Evaluation Review, 8(3):215–224, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gertner:1979:PEC

[157] Ilya Gertner. Performance evaluation of communicating processes. ACM SIGMETRICS Performance Evaluation Review, 8(3):241–248, Fall 1979.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Spooner:1979:BIS

[158] Christopher R. Spooner. Benchmarking interactive systems: Producing the software. ACM SIGMETRICS Performance Evaluation Review, 8(3):249– 257, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dujmovic:1979:CCP

[159] Jozo J. Dujmović. Criteria for computer performance analysis. ACM SIGMETRICS Performance Evaluation Review, 8(3):259–267, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dyal:1979:SBS

[160] James O. Dyal and William DeWald, Jr. Small business system performance analysis. ACM SIGMETRICS Performance Evaluation Review, 8(3):269– 275, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Huff:1979:SCR

[161] Robert W. Huff. System characterization of a Retail Business System. ACM SIGMETRICS Performance Evaluation Review, 8(3):277–284, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stroebel:1979:FPA

[162] Gary Stroebel. Field performance aids for IBM GSD systems. ACM SIGMET-RICS Performance Evaluation Review, 8(3):285–291, Fall 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Clark:1979:FAP

[163] Jon D. Clark. A feature analysis of performance evaluation texts. ACM SIGMETRICS Performance Evaluation Review, 8(4):9–11, December 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dowdy:1979:SWT

[164] Lawrence W. Dowdy. Synopsis of workshop on the theory and application of analytical models to ADP system performance prediction. ACM SIGMET-RICS Performance Evaluation Review, 8(4):13–17, December 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Petrella:1979:SWS

[165] Arthur Petrella and Harold Farrey. Simulating working sets under MVS. ACM SIGMETRICS Performance Evaluation Review, 8(4):24–36, December 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pierson:1979:PEM

[166] Daniel L. Pierson. Performance evaluation of a minicomputer-based data collection system. ACM SIGMET-RICS Performance Evaluation Review, 8(4):37–44, December 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Artis:1979:CPM

[167] H. Pat Artis. Capacity planning for MVS computer systems. ACM SIGMETRICS Performance Evaluation Review, 8(4):45–62, December 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rajaraman:1979:PVM

[168] M. K. Rajaraman. Performance of a virtual memory: some experimental results. ACM SIGMETRICS Performance Evaluation Review, 8(4):63–68, December 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Denning:1980:WWS

[169] Peter J. Denning. What’s a working set? ACM SIGMETRICS Performance Evaluation Review, 9(1):6–10, Spring 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Petrella:1980:SWS

[170] Arthur Petrella and Harold Farrey. Simulating working sets under MVS. ACM SIGMETRICS Performance Evaluation Review, 9(1):11–23, Spring 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Spiegel:1980:MEP

[171] Mitchell G. Spiegel. Measuring and evaluating performance. ACM SIG-METRICS Performance Evaluation Review, 9(1):33–34, Spring 1980. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dixon:1980:PMI

[172] P. J. Dixon. Planning MIS investment and expense levels. ACM SIGMET-RICS Performance Evaluation Review, 9(1):35–37, Spring 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Moran:1980:CPV

[173] Thomas S. Moran. Capacity planning: ‘the volume’. ACM SIGMET-RICS Performance Evaluation Review, 9(1):38–40, Spring 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

DeMarco:1980:BLB

[174] Tom DeMarco. Breaking the language barrier. ACM SIGMETRICS Performance Evaluation Review, 9(1):41–45, Spring 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Giles:1980:CSM

[175] Howard L. Giles. Communications systems management. ACM SIGMET-RICS Performance Evaluation Review, 9(1):46–51, Spring 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Erlandson:1980:SEM

[176] Robert F. Erlandson. System evaluation methodologies: combined multidimensional scaling and ordering techniques. ACM SIGMETRICS Performance Evaluation Review, 9(1):52–58, Spring 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pearson:1980:MCU

[177] Sammy W. Pearson and James E. Bailey. Measurement of computer user satisfaction. ACM SIGMETRICS Performance Evaluation Review, 9(1):59–68, Spring 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chandy:1980:CAP

[178] K. Mani Chandy and Charles H. Sauer. Computational algorithms for product form queueing networks. ACM SIGMETRICS Performance Evaluation Review, 9(2):1, Summer 1980. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gordon:1980:ICP

[179] Karen D. Gordon and Lawrence W. Dowdy. The impact of certain parameter estimation errors in queueing network models. ACM SIGMET-RICS Performance Evaluation Review, 9(2):3–9, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Blake:1980:XIM

[180] Russ Blake. XRAY: Instrumentation for multiple computers. ACM SIGMETRICS Performance Evaluation Review, 9(2):11–25, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hughes:1980:DDA

[181] James H. Hughes. DIAMOND a digital analyzer and monitoring device. ACM SIGMETRICS Performance Evaluation Review, 9(2):27–34, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bashioum:1980:BIS

[182] Douglas L. Bashioum. Benchmarking interactive systems: Calibrating the model. ACM SIGMETRICS Performance Evaluation Review, 9(2):35–41, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lehmann:1980:PEP

[183] Axel Lehmann. Performance evaluation and prediction of storage hierarchies. ACM SIGMETRICS Performance Evaluation Review, 9(2):43–54, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Alanko:1980:MER

[184] Timo O. Alanko, Ilkka J. Haikala, and Petri H. Kutvonen. Methodology and empirical results of program behaviour measurements. ACM SIGMETRICS Performance Evaluation Review, 9(2):55–66, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kumar:1980:PRB

[185] Gopa Kumar and C. Thomas Nute. Program restructuring for block structured languages. ACM SIGMET-RICS Performance Evaluation Review, 9(2):67–79, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vantilborgh:1980:NCD

[186] Hendrik T. Vantilborgh, Richard L. Garner, and Edward D. Lazowska. Near-complete decomposability of queueing networks with clusters of strongly interacting servers. ACM SIGMETRICS Performance Evaluation Review, 9(2):81–92, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brandwajn:1980:FRE

[187] Alexandre Brandwajn. Further results on equivalence and decomposition in queueing network models. ACM SIGMETRICS Performance Evaluation Review, 9(2):93–104, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stewart:1980:ECF

[188] William J. Stewart and Gerald A. Zeiszler. On the existence of composite flow equivalent Markovian servers. ACM SIGMETRICS Performance Evaluation Review, 9(2):105–116, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marie:1980:CEP

[189] Raymond Marie. Calculating equilibrium probabilities for λ(n)/C_{k}/1/N queues. ACM SIGMETRICS Performance Evaluation Review, 9(2):117– 125, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wagner:1980:HCS

[190] Robert A. Wagner and Kishor S. Trivedi. Hardware configuration selection through discretizing a continuous variable solution. ACM SIGMET-RICS Performance Evaluation Review, 9(2):127–142, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bard:1980:MSD

[191] Yonathan Bard. A model of shared DASD and multipathing. ACM SIGMETRICS Performance Evaluation Review, 9(2):143, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lo:1980:CCP

[192] T. L. Lo. Computer capacity planning using queueing network models. ACM SIGMETRICS Performance Evaluation Review, 9(2):145–152, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kurinckx:1980:OVC

[193] A. Kurinckx and G. Pujolle. Overallocation in a virtual circuit computer network. ACM SIGMETRICS Performance Evaluation Review, 9(2): 153–158, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Upton:1980:ADA

[194] Richard A. Upton and Satish K. Tripathi. Analysis of design alternatives for a packet switched I/O system. ACM SIGMETRICS Performance Evaluation Review, 9(2):159–171, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Balkovich:1980:PDS

[195] Edward E. Balkovich and Colin Whitby-Strevens. On the performance of decentralized software. ACM SIG-METRICS Performance Evaluation Review, 9(2):173–180, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Grit:1980:PMA

[196] Dale H. Grit and Rex L. Page. Performance of a multiprocessor for applicative programs. ACM SIGMET-RICS Performance Evaluation Review, 9(2):181–189, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dhas:1980:PEF

[197] C. Retna Dhas. Performance evaluation of a feedback data flow processor using simulation. ACM SIGMET-RICS Performance Evaluation Review, 9(2):191–197, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bryant:1980:HMG

[198] Raymond M. Bryant. On homogeneity in M/G/1 queueing systems. ACM SIGMETRICS Performance Evaluation Review, 9(2):199–208, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coffman:1980:ORP

[199] E. G. Coffman, Jr., Erol Gelenbe, and Roger C. Wood. Optimal replication of parallel-read, sequential-write systems. ACM SIGMETRICS Performance Evaluation Review, 9(2):209– 216, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shore:1980:LRO

[200] John E. Shore. The lazy repairman and other models: Performance collapse due to overhead in simple, single-server queuing systems. ACM SIGMET-RICS Performance Evaluation Review, 9(2):217–224, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lam:1980:RTD

[201] Simon S. Lam and A. Udaya Shankar. Response time distributions for a multi-class queue with feedback. ACM SIGMETRICS Performance Evaluation Review, 9(2):225–234, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:1980:AIO

[202] Y. T. Wang. Analysis of an intrinsic overload control for a class of queueing systems. ACM SIGMET-RICS Performance Evaluation Review, 9(2):235–243, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Smith:1980:ASD

[203] Connie Smith and J. C. Browne. Aspects of software design analysis: Concurrency and blocking. ACM SIGMETRICS Performance Evaluation Review, 9(2):245–253, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Potier:1980:ALP

[204] D. Potier and Ph. Leblanc. Analysis of locking policies in database management systems. ACM SIGMET-RICS Performance Evaluation Review, 9(2):255, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coffman:1980:ONC

[205] E. G. Coffman, Jr., E. Gelenbe, and B. Plateau. Optimization of the number of copies in a distribution data base. ACM SIGMETRICS Performance Evaluation Review, 9(2):257– 263, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ruschitzka:1980:RJC

[206] Manfred Ruschitzka. The response of job classes with distinct policy functions (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 9(2):265, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kim:1980:PTO

[207] K. H. Kim and Mahmoud Naghibzadeh. Prevention of task overruns in realtime non-preemptive multiprogramming systems. ACM SIGMETRICS Performance Evaluation Review, 9(2): 267–276, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

King:1980:NMI

[208] P. J. B. King and I. Mitrani. Numerical methods for infinite Markov processes. ACM SIGMETRICS Performance Evaluation Review, 9(2):277– 282, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fayolle:1980:SCT

[209] G. Fayolle, P. J. B. King, and I. Mitrani. The solution of certain twodimensional Markov models. ACM SIGMETRICS Performance Evaluation Review, 9(2):283–289, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Clark:1980:EIE

[210] Jon D. Clark and Robert M. Golladay. Empirical investigation of the effectiveness of several computer performance evaluation tools. ACM SIGMETRICS Performance Evaluation Review, 9(3): 31–36, Fall 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Estell:1980:BW

[211] Robert G. Estell. Benchmarks and watermarks. ACM SIGMETRICS Performance Evaluation Review, 9(3):39–44, Fall 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kleijnen:1980:SMM

[212] J. P. C. Kleijnen. Scoring methods, multiple criteria, and utility analysis. ACM SIGMETRICS Performance Evaluation Review, 9(3):45–56, Fall 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Denning:1980:TTI

[213] Peter J. Denning. A tale of two islands: a fable. ACM SIGMETRICS Performance Evaluation Review, 9(4):7–10, Winter 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yader:1980:ACP

[214] Mark J. Yader. ADP capacity planning: a case study. ACM SIGMET-RICS Performance Evaluation Review, 9(4):11–25, Winter 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marrevee:1980:HFF

[215] J. Marrevée. How friendly and fast is FAST DUMP RESTORE. ACM SIGMETRICS Performance Evaluation Review, 9(4):28–35, Winter 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bernard:1980:SUM

[216] J. C. Bernard. T-scan: the use of micro computers for response time measurements. ACM SIGMETRICS Performance Evaluation Review, 9(4):39–50, Winter 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bell:1981:SLC

[217] Thomas E. Bell. Structured lifecycle assumptions. ACM SIGMET-RICS Performance Evaluation Review, 10(1):1–3, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coughlin:1981:SDM

[218] Donald T. Coughlin. System development methodology or system research methodology? ACM SIGMETRICS Performance Evaluation Review, 10(1): 5–6, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Earle:1981:AAB

[219] Dennis M. Earle. An alchemical approach to brokerage. ACM SIGMET-RICS Performance Evaluation Review, 10(1):7–8, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Spiegel:1981:PAI

[220] Mitchell G. Spiegel. Prototyping: an approach to information and communication system design. ACM SIGMET-RICS Performance Evaluation Review, 10(1):9–19, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jenkins:1981:APC

[221] C. Wesley Jenkins. Application prototyping: a case study. ACM SIGMET-RICS Performance Evaluation Review, 10(1):21–27, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cox:1981:SRT

[222] Patricia R. Cox. Specification of a regression test for a mini computer operating system. ACM SIGMET-RICS Performance Evaluation Review, 10(1):29–32, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bryan:1981:ASC

[223] William Bryan, Stanley Siegel, and Gary Whiteleather. An approach to software configuration control. ACM SIGMETRICS Performance Evaluation Review, 10(1):33–47, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fredrick:1981:PIS

[224] C. R. Fredrick. Project implementation of Software Configuration Management. ACM SIGMETRICS Performance Evaluation Review, 10(1):49–56, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berlack:1981:ISC

[225] H. Ronald Berlack. Implementing software configuration control in the structured programming environment. ACM SIGMETRICS Performance Evaluation Review, 10(1):57–77, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gross:1981:PCV

[226] Peter Gross. Producers and consumers views of software quality (panel session). ACM SIGMETRICS Performance Evaluation Review, 10(1):79, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Henry:1981:RAT

[227] Sallie Henry, Dennis Kafura, and Kathy Harris. On the relationships among three software metrics. ACM SIGMETRICS Performance Evaluation Review, 10(1):81–88, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Szulewski:1981:MSS

[228] Paul A. Szulewski, Mark H. Whitworth, Philip Buchan, and J. Barton DeWolf. The measurement of software science parameters in software designs. ACM SIGMETRICS Performance Evaluation Review, 10(1):89–94, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Basili:1981:ECS

[229] Victor R. Basili and Tsai-Yun Phillips. Evaluating and comparing software metrics in the software engineering laboratory. ACM SIGMETRICS Performance Evaluation Review, 10(1):95– 106, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ronback:1981:TMS

[230] James Ronback. Test metrics for software quality. ACM SIGMETRICS Performance Evaluation Review, 10(1): 107, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Benson:1981:AST

[231] J. P. Benson. Adaptive search techniques applied to software testing. ACM SIGMETRICS Performance Evaluation Review, 10(1):109– 116, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Paige:1981:DST

[232] Michael Paige. Data space testing. ACM SIGMETRICS Performance Evaluation Review, 10(1):117– 127, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goel:1981:OTP

[233] Amrit L. Goel. Optimal testing policies for software systems. ACM SIGMET-RICS Performance Evaluation Review, 10(1):129–130, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Littlewood:1981:BDD

[234] B. Littlewood. A Bayesian differential debugging model for software reliability. ACM SIGMETRICS Performance Evaluation Review, 10(1):129– 130, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Musa:1981:SRMa

[235] J. D. Musa and A. Iannino. Software reliability modeling accounting for program size variation due to integration or design changes. ACM SIGMET-RICS Performance Evaluation Review, 10(1):129–130, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Musa:1981:SRMb

[236] John D. Musa. Software reliability measurement session. ACM SIGMET-RICS Performance Evaluation Review, 10(1):129–130, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goel:1981:WST

[237] Amrit L. Goel and Kazuhira Okumoto. When to stop testing and start using software? ACM SIGMET-RICS Performance Evaluation Review, 10(1):131–138, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Littlewood:1981:SRG

[238] B. Littlewood. Stochastic reliability growth: a model with applications to computer software faults and hardware design faults. ACM SIGMET-RICS Performance Evaluation Review, 10(1):139–152, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ottenstein:1981:SDS

[239] Linda M. Ottenstein. Software defects — a software science perspective. ACM SIGMETRICS Performance Evaluation Review, 10(1):153– 155, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ottenstein:1981:PNE

[240] Linda Ottenstein. Predicting numbers of errors using software science. ACM SIGMETRICS Performance Evaluation Review, 10(1):157–167, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schneider:1981:SEE

[241] Victor Schneider. Some experimental estimators for developmental and delivered errors in software development projects. ACM SIGMETRICS Performance Evaluation Review, 10 (1):169–172, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sneed:1981:SSA

[242] H. Sneed. SOFTDOC — A system for automated software static analysis and documentation. ACM SIGMET-RICS Performance Evaluation Review, 10(1):173–177, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Crowley:1981:ADP

[243] John D. Crowley. The application development process: What’s wrong with it? ACM SIGMETRICS Performance Evaluation Review, 10(1):179– 187, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bailey:1981:SSU

[244] C. T. Bailey and W. L. Dingee. A software study using Halstead metrics. ACM SIGMETRICS Performance Evaluation Review, 10(1):189– 197, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Esposito:1981:WCT

[245] A. Esposito, A. Mazzeo, and P. Costa. Workload characterization for trend analysis. ACM SIGMETRICS Performance Evaluation Review, 10(2):5–15, Summer 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Musa:1981:SRMc

[246] J. D. Musa and A. Iannino. Software reliability modeling: accounting for program size variation due to integration or design changes. ACM SIGMET-RICS Performance Evaluation Review, 10(2):16–25, Summer 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Comer:1981:CTD

[247] J. R. Comer, J. R. Rinewalt, and M. M. Tanik. A comparison of two different program complexity measures. ACM SIGMETRICS Performance Evaluation Review, 10(2):26–28, Summer 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Abrams:1981:NNM

[248] Marshall D. Abrams and Dorothy C. Neiman. NBS network measurement methodology applied to synchronous communications. ACM SIGMETRICS Performance Evaluation Review, 10 (2):29–36, Summer 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Larsen:1981:CEL

[249] R. L. Larsen, J. R. Agre, and A. K. Agrawala. A comparative evaluation of local area communication technology. ACM SIGMETRICS Performance Evaluation Review, 10(2):37–47, Summer 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hughes:1981:HPT

[250] Herman D. Hughes. A highly parameterized tool for studying performance of computer systems. ACM SIGMET-RICS Performance Evaluation Review, 10(2):48–65, Summer 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Spiegel:1981:RPP

[251] Mitchell G. Spiegel. RTE’s: past is prologue. ACM SIGMETRICS Performance Evaluation Review, 10(2):66–73, Summer 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Browne:1981:DSP

[252] J. C. Browne. Designing systems for performance. ACM SIGMETRICS Performance Evaluation Review, 10(3): 1, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reiner:1981:MAP

[253] David Reiner and Tad Pinkerton. A method for adaptive performance improvement of operating systems. ACM SIGMETRICS Performance Evaluation Review, 10(3):2–10, Fall 1981. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:1981:VVT

[254] Y. T. Wang. On the VAX/VMS time-critical process scheduling. ACM SIGMETRICS Performance Evaluation Review, 10(3):11–18, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Soderlund:1981:ECP

[255] Lars Söderlund. Evaluation of concurrent physical database reorganization through simulation modeling. ACM SIGMETRICS Performance Evaluation Review, 10(3):19–32, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lazowska:1981:AMD

[256] Edward D. Lazowska and John Zahorjan. Analytic modelling of disk I/O subsystems: a tutorial. ACM SIGMETRICS Performance Evaluation Review, 10(3):33–35, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dowdy:1981:MUS

[257] Lawrence W. Dowdy and Hans J. Breitenlohner. A model of Univac 1100/42 swapping. ACM SIGMETRICS Performance Evaluation Review, 10(3):36– 47, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Turner:1981:SFP

[258] Rollins Turner and Henry Levy. Segmented FIFO page replacement. ACM SIGMETRICS Performance Evaluation Review, 10(3):48–51, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ferrari:1981:GMW

[259] Domenico Ferrari. A generative model of working set dynamics. ACM SIGMETRICS Performance Evaluation Review, 10(3):52–57, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zahorjan:1981:BJB

[260] J. Zahorjan, K. C. Sevcik, D. L. Eager, and B. I. Galler. Balanced job bound analysis of queueing networks. ACM SIGMETRICS Performance Evaluation Review, 10(3):58, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Neuse:1981:SHA

[261] D. Neuse and K. Chandy. SCAT: a heuristic algorithm for queueing network models of computing systems. ACM SIGMETRICS Performance Evaluation Review, 10(3):59–79, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zahorjan:1981:SSQ

[262] John Zahorjan and Eugene Wong. The solution of separable queueing network models using mean value analysis. ACM SIGMETRICS Performance Evaluation Review, 10(3):80–85, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thomasian:1981:ASQ

[263] Alexander Thomasian and Behzad Nadji. Aggregation of stations in queueing network models of multiprogrammed computers. ACM SIGMET-RICS Performance Evaluation Review, 10(3):86–104, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schwetman:1981:CSM

[264] Herb Schwetman. Computer system models: an introduction. ACM SIGMETRICS Performance Evaluation Review, 10(3):105, Fall 1981. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Denning:1981:PEE

[265] Peter J. Denning. Performance evaluation: Experimental computer science at its best. ACM SIGMET-RICS Performance Evaluation Review, 10(3):106–109, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rafii:1981:SAM

[266] Abbas Rafii. Structure and application of a measurement tool — SAMPLER/ 3000. ACM SIGMETRICS Performance Evaluation Review, 10(3):110– 120, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tolopka:1981:ETM

[267] Stephen Tolopka. An event trace monitor for the VAX 11/780. ACM SIGMETRICS Performance Evaluation Review, 10(3):121–128, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Artis:1981:LFD

[268] H. Pat Artis. A log file design for analyzing secondary storage occupancy. ACM SIGMETRICS Performance Evaluation Review, 10(3):129– 135, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sanguinetti:1981:ESS

[269] John Sanguinetti. The effects of solid state paging devices in a large time-sharing system. ACM SIGMET-RICS Performance Evaluation Review, 10(3):136–153, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:1981:VMB

[270] Richard T. Wang and J. C. Browne. Virtual machine-based simulation of distributed computing and network computing. ACM SIGMETRICS Performance Evaluation Review, 10(3): 154–156, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Huslende:1981:CEP

[271] Ragnar Huslende. A combined evaluation of performance and reliability for degradable systems. ACM SIGMET-RICS Performance Evaluation Review, 10(3):157–164, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jacobson:1981:MSD

[272] Patricia A. Jacobson and Edward D. Lazowska. The method of surrogate delays: Simultaneous resource possession in analytic models of computer systems. ACM SIGMETRICS Performance Evaluation Review, 10(3):165– 174, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jacobson:1981:AAM

[273] Patricia Jacobson. Approximate analytic models of arbiters. ACM SIGMETRICS Performance Evaluation Review, 10(3):175–180, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Briggs:1981:PCB

[274] Fayé A. Briggs and Michel Dubois. Performance of cache-based multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 10(3):181– 190, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bryant:1981:QNA

[275] R. M. Bryant and J. R. Agre. A queueing network approach to the module allocation problem in distributed systems. ACM SIGMETRICS Performance Evaluation Review, 10(3):191– 204, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marathe:1981:AME

[276] Madhav Marathe and Sujit Kumar. Analytical models for an Ethernetlike local area network link. ACM SIGMETRICS Performance Evaluation Review, 10(3):205–215, Fall 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pechura:1981:PLM

[277] Michael A. Pechura. Page life measurements. ACM SIGMETRICS Performance Evaluation Review, 10(4):10–12, December 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Clark:1981:UES

[278] Jon D. Clark. An update on economiesof-scale in computing systems. ACM SIGMETRICS Performance Evaluation Review, 10(4):13–14, December 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Janusz:1981:GMS

[279] Edward R. Janusz. Getting the most out of a small computer. ACM SIGMETRICS Performance Evaluation Review, 10(4):22–35, December 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cox:1981:DDD

[280] Springer Cox. Data, definition, deduction: an empirical view of operational analysis. ACM SIGMET-RICS Performance Evaluation Review, 10(4):36–44, December 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Muramatsu:1981:SVQ

[281] Hiroshi Muramatsu, Masahiro Date, and Takanori Maki. Structural validation in queueing network models of computer systems. ACM SIGMET-RICS Performance Evaluation Review, 10(4):41–46, December 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sauer:1981:NSS

[282] Charles H. Sauer. Numerical solution of some multiple chain queueing networks. ACM SIGMETRICS Performance Evaluation Review, 10 (4):47–56, December 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nemeth:1981:AIP

[283] Thomas A. Nemeth. An approach to interactive performance analysis in a busy production system (NOS/BE). ACM SIGMETRICS Performance Evaluation Review, 10(4):57–73, December 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Knudson:1981:CPE

[284] Michael E. Knudson. A computer performance evaluation operational methodology. ACM SIGMETRICS Performance Evaluation Review, 10 (4):74–80, December 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Afshari:1981:MNT

[285] P. V. Afshari, S. C. Bruell, and R. Y. Kain. Modeling a new technique for accessing shared buses. ACM SIGMET-RICS Performance Evaluation Review, 11(1):4–13, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lazar:1981:OCM

[286] Aurel A. Lazar. Optimal control of a M/M/m queue. ACM SIGMET-RICS Performance Evaluation Review, 11(1):14–20, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Spirn:1981:NMB

[287] Jeffrey R. Spirn. Network modeling with bursty traffic and finite buffer space. ACM SIGMETRICS Performance Evaluation Review, 11(1):21–28, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lam:1981:ORN

[288] Simon S. Lam and Y. Luke Lien. Optimal routing in networks with flowcontrolled virtual channels. ACM SIGMETRICS Performance Evaluation Review, 11(1):38–46, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Livny:1981:LBH

[289] Miron Livny and Myron Melman. Load balancing in homogeneous broadcast distributed systems. ACM SIGMET-RICS Performance Evaluation Review, 11(1):47–55, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wecker:1981:PGD

[290] Stuart Wecker, Robert Gordon, James Gray, James Herman, Raj Kanodia, and Dan Seligman. Performance of globally distributed networks. ACM SIGMETRICS Performance Evaluation Review, 11(1):58, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gordon:1981:OMH

[291] R. L. Gordon. Operational measurements on a high performance ring. ACM SIGMETRICS Performance Evaluation Review, 11(1):59, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gray:1981:PSL

[292] James P. Gray. Performance of SNA’s LU-LU session protocols. ACM SIGMETRICS Performance Evaluation Review, 11(1):60–61, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Herman:1981:APT

[293] James G. Herman. ARPANET performance tuning techniques. ACM SIGMETRICS Performance Evaluation Review, 11(1):62, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aleh:1981:DUB

[294] Avner Aleh and K. Dan Levin. The determination of upper bounds for economically effective compression in packet switching networks. ACM SIGMETRICS Performance Evaluation Review, 11(1):64–72, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

McGregor:1981:CMP

[295] Patrick V. McGregor. Concentrator modeling with pipelining arrivals compensation. ACM SIGMETRICS Performance Evaluation Review, 11(1):73– 94, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mink:1981:MEC

[296] Alan Mink and Charles B. Silio, Jr. Modular expansion in a class of homogeneous networks. ACM SIGMET-RICS Performance Evaluation Review, 11(1):95–100, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thareja:1981:UBA

[297] Ashok K. Thareja, Satish K. Tripathi, and Richard A. Upton. On updating buffer allocation. ACM SIGMET-RICS Performance Evaluation Review, 11(1):101–110, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Elsanadidi:1981:ATW

[298] M. Y. Elsanadidi and Wesley W. Chu. An analysis of a time window multiaccess protocol with collision size feedback (WCSF). ACM SIGMET-RICS Performance Evaluation Review, 11(1):112–118, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Roehr:1981:PALa

[299] Kuno M. Roehr and Horst Sadlowski. Performance analysis of local communication loops. ACM SIGMETRICS Performance Evaluation Review, 11 (1):119–129, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sherman:1981:DVH

[300] R. H. Sherman, M. G. Gable, and A. W. Chung. Distributed virtual hosts and networks: Measurement and control. ACM SIGMETRICS Performance Evaluation Review, 11(1):130– 136, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brice:1981:NPA

[301] Richard Brice and William Alexander. A network performance analyst’s workbench. ACM SIGMET-RICS Performance Evaluation Review, 11(1):138–146, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

DuBois:1981:HMS

[302] Donald F. DuBois. A Hierarchical Modeling System for computer networks. ACM SIGMETRICS Performance Evaluation Review, 11(1):147– 155, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Terplan:1981:NPR

[303] K. Terplan. Network performance reporting. ACM SIGMETRICS Performance Evaluation Review, 11(1):156– 170, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Spiegel:1981:QLA

[304] Mitchell G. Spiegel. Questions for Local Area Network panelists. ACM SIGMETRICS Performance Evaluation Review, 11(1):172, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Roehr:1981:PALb

[305] Kuno M. Roehr and Horst Sadlowski. Performance analysis of local communication loops. ACM SIGMETRICS Performance Evaluation Review, 11(1): 173, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sternick:1981:SAD

[306] Barbara R. Sternick. Systems aids in determining Local Area Network performance characteristics. ACM SIGMETRICS Performance Evaluation Review, 11(1):174, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anonymous:1981:AI

[307] Anonymous. Authors index. ACM SIGMETRICS Performance Evaluation Review, 11(1):175, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rajaraman:1982:PET

[308] M. K. Rajaraman. Performance evaluation through job scheduler modeling. ACM SIGMETRICS Performance Evaluation Review, 11(2):9–15, Summer 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mager:1982:TPA

[309] Peter S. Mager. Toward a parametric approach for modeling local area network performance. ACM SIGMET-RICS Performance Evaluation Review, 11(2):17–28, Summer 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gaffney:1982:SSI

[310] John E. Gaffney, Jr. Score ‘82 — a summary (at IBM Systems Research Institute, 3/23-3/24/82). ACM SIGMET-RICS Performance Evaluation Review, 11(2):30–32, Summer 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Misek-Falkoff:1982:NFS

[311] Linda D. Misek-Falkoff. The new field of “Software Linguistics”: an early-bird view. ACM SIGMETRICS Performance Evaluation Review, 11 (2):35–51, Summer 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Spiegel:1982:SCR

[312] Mitchell G. Spiegel. Software counting rules: Will history repeat itself? ACM SIGMETRICS Performance Evaluation Review, 11(2):52–56, Summer 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kavi:1982:EDS

[313] Krishna M. Kavi and U. B. Jackson. Effect of declarations on software metrics: an experiment in software science. ACM SIGMETRICS Performance Evaluation Review, 11 (2):57–71, Summer 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gaffney:1982:MIC

[314] John E. Gaffney, Jr. Machine Instruction Count Program. ACM SIGMET-RICS Performance Evaluation Review, 11(2):72–79, Summer 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Misek-Falkoff:1982:UHS

[315] Linda D. Misek-Falkoff. A unification of Halstead’s Software Science counting rules for programs and English text, and a claim space approach to extensions. ACM SIGMET-RICS Performance Evaluation Review, 11(2):80–114, Summer 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Estes:1982:DPO

[316] George E. Estes. Distinguishing the potential operands in FORTRAN programs. ACM SIGMETRICS Performance Evaluation Review, 11(2):115– 117, Summer 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Conte:1982:EDC

[317] S. D. Conte, V. Y. Shen, and K. Dickey. On the effect of different counting rules for control flow operators on Software Science metrics in Fortran. ACM SIGMETRICS Performance Evaluation Review, 11(2):118–126, Summer 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shanthikumar:1982:PCF

[318] J. G. Shanthikumar, P. K. Varshney, and K. Sriram. A priority cutoff flow control scheme for integrated voicedata multiplexers. ACM SIGMET-RICS Performance Evaluation Review, 11(3):8–14, Fall 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cox:1982:DDD

[319] Springer Cox. Data, definition, deduction: an empirical view of operational analysis. ACM SIGMETRICS Performance Evaluation Review, 11(3):15–20, Fall 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Perros:1982:QLD

[320] H. G. Perros. The queue-length distribution of the M/Ck/1 queue. ACM SIGMETRICS Performance Evaluation Review, 11(3):21–24, Fall 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anderson:1982:BMP

[321] Gordon E. Anderson. Bernoulli methods for predicting communication processor performance. ACM SIGMET-RICS Performance Evaluation Review, 11(3):25–29, Fall 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Laurmaa:1982:AHT

[322] Timo Laurmaa and Markku Syrjänen. APL and Halstead’s theory: a measuring tool and some experiments. ACM SIGMETRICS Performance Evaluation Review, 11(3):32–47, Fall 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Beser:1982:FES

[323] Nicholas Beser. Foundations and experiments in software science. ACM SIGMETRICS Performance Evaluation Review, 11(3):48–72, Fall 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schnurer:1982:PAP

[324] Karl Ernst Schnurer. Product Assurance Program Analyzer (P.A.P.A.) a tool for program complexity evaluation. ACM SIGMETRICS Performance Evaluation Review, 11(3):73–74, Fall 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gross:1982:CME

[325] David R. Gross, Mary A. King, Michael R. Murr, and Michael R. Eddy. Complexity measurement of Electronic Switching System (ESS) software. ACM SIGMETRICS Performance Evaluation Review, 11(3):75–85, Fall 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hartman:1982:CTR

[326] Sandra D. Hartman. A counting tool for RPG. ACM SIGMETRICS Performance Evaluation Review, 11(3):86– 100, Fall 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Naib:1982:ASS

[327] Farid A. Naib. An application of software science to the quantitative measurement of code quality. ACM SIGMETRICS Performance Evaluation Review, 11(3):101–128, Fall 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Blake:1982:OCT

[328] Russ Blake. Optimal control of thrashing. ACM SIGMETRICS Performance Evaluation Review, 11(4):1–10, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Babaoglu:1982:HRD

[329] Özalp Babaoğlu. Hierarchical replacement decisions in hierarchical stores. ACM SIGMETRICS Performance Evaluation Review, 11(4):11–19, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hagmann:1982:PPR

[330] Robert B. Hagmann and Robert S. Fabry. Program page reference patterns. ACM SIGMETRICS Performance Evaluation Review, 11(4):20–29, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bunt:1982:EMP

[331] R. B. Bunt, R. S. Harbus, and S. J. Plumb. The effective management of paging storage hierarchies. ACM SIGMETRICS Performance Evaluation Review, 11(4):30–38, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hodges:1982:WCP

[332] Larry F. Hodges and William J. Stewart. Workload characterization and performance evaluation in a research environment. ACM SIGMETRICS Performance Evaluation Review, 11 (4):39–50, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Haring:1982:SDW

[333] Günter Haring. On state-dependent workload characterization by software resources. ACM SIGMETRICS Performance Evaluation Review, 11(4):51–57, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bolzoni:1982:PIS

[334] M. L. Bolzoni, M. C. Calzarossa, P. Mapelli, and G. Serazzi. A package for the implementation of static workload models. ACM SIGMET-RICS Performance Evaluation Review, 11(4):58–67, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

McDaniel:1982:MSI

[335] Gene McDaniel. The Mesa Spy: an interactive tool for performance debugging. ACM SIGMETRICS Performance Evaluation Review, 11(4):68–76, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hercksen:1982:MSE

[336] Uwe Hercksen, Rainer Klar, Wolfgang Kleinöder, and Franz Kneißl. Measuring simultaneous events in a multiprocessor system. ACM SIGMET-RICS Performance Evaluation Review, 11(4):77–88, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gelenbe:1982:SDF

[337] Erol Gelenbe. Stationary deterministic flows in discrete systems: I. ACM SIGMETRICS Performance Evaluation Review, 11(4):89–101, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baccelli:1982:DBR

[338] F. Baccelli and E. G. Coffman. A data base replication analysis using an M/M/m queue with service interruptions. ACM SIGMETRICS Performance Evaluation Review, 11(4): 102–107, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Plateau:1982:MPR

[339] Brigitte Plateau and Andreas Staphylopatis. Modelling of the parallel resolution of a numerical problem on a locally distributed computing system. ACM SIGMETRICS Performance Evaluation Review, 11(4): 108–117, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bard:1982:MSD

[340] Yonathan Bard. Modeling I/O systems with dynamic path selection, and general transmission networks. ACM SIGMETRICS Performance Evaluation Review, 11(4):118–129, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lazowska:1982:MCM

[341] Edward D. Lazowska and John Zahorjan. Multiple class memory constrained queueing networks. ACM SIGMET-RICS Performance Evaluation Review, 11(4):130–140, December 1982. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brandwajn:1982:FAS

[342] Alexandre Brandwajn. Fast approximate solution of multiprogramming models. ACM SIGMETRICS Performance Evaluation Review, 11(4): 141–149, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Agrawal:1982:ASM

[343] Subhash C. Agrawal and Jeffrey P. Buzen. The aggregate server method for analyzing serialization delays in computer systems. ACM SIGMET-RICS Performance Evaluation Review, 11(4):150, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Smith:1982:PAS

[344] Connie U. Smith and David D. Loendorf. Performance analysis of software for an MIMD computer. ACM SIGMETRICS Performance Evaluation Review, 11(4):151–162, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Agre:1982:MRN

[345] Jon R. Agre and Satish K. Tripathi. Modeling reentrant and nonreentrant software. ACM SIGMETRICS Performance Evaluation Review, 11(4): 163–178, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wu:1982:OME

[346] L. T. Wu. Operational models for the evaluation of degradable computing systems. ACM SIGMETRICS Performance Evaluation Review, 11(4): 179–185, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marie:1982:ECA

[347] Raymond A. Marie, Patricia M. Snyder, and William J. Stewart. Extensions and computational aspects of an iterative method. ACM SIGMETRICS Performance Evaluation Review, 11(4): 186–194, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Neuse:1982:HHA

[348] Doug Neuse and K. Mani Chandy. HAM: the heuristic aggregation method for solving general closed queueing network models of computer systems. ACM SIGMETRICS Performance Evaluation Review, 11(4): 195–212, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Eager:1982:PBH

[349] D. L. Eager and K. C. Sevcik. Performance bound hierarchies for queueing networks. ACM SIGMETRICS Performance Evaluation Review, 11(4): 213–214, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brumfield:1982:EAH

[350] Jeffrey A. Brumfield and Peter J. Denning. Error analysis of homogeneous mean queue and response time estimators. ACM SIGMETRICS Performance Evaluation Review, 11(4): 215–221, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harbitter:1982:MTL

[351] Alan Harbitter and Satish K. Tripathi. A model of transport level flow control. ACM SIGMETRICS Performance Evaluation Review, 11(4):222–232, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gelenbe:1982:CPC

[352] Erol Gelenbe and Isi Mitrani. Control policies in CSMA local area networks: Ethernet controls. ACM SIGMET-RICS Performance Evaluation Review, 11(4):233–240, December 1982. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tripathi:1982:ATF

[353] Satish K. Tripathi and Alan Harbitter. An analysis of two flow control techniques. ACM SIGMETRICS Performance Evaluation Review, 11(4): 241–249, December 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

King:1982:MCR

[354] P. J. B. King and I. Mitrani. Modelling the Cambridge Ring. ACM SIGMET-RICS Performance Evaluation Review, 11(4):250–258, December 1982. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marrevee:1982:PRT

[355] J. Marrevee. The power of the read track and the need for a write track command for disk back-up and restore utilities. ACM SIGMETRICS Performance Evaluation Review, 12(1):10– 14, December 1982/1983. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Perros:1982:MPR

[356] H. G. Perros. A model for predicting the response time of an online system for electronic fund transfer. ACM SIGMETRICS Performance Evaluation Review, 12(1):15–21, Winter 1982/1983. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Augustin:1982:CCD

[357] Reinhard Augustin and Klaus-Jürgen Büscher. Characteristics of the COXdistribution. ACM SIGMETRICS Performance Evaluation Review, 12(1):22– 32, December 1982/1983. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Perros:1984:QNB

[358] H. G. Perros. Queueing networks with blocking: a bibliography. ACM SIG-METRICS Performance Evaluation Review, 12(2):8–12, Spring-Summer 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

DeMarco:1984:ASS

[359] Tom DeMarco. An algorithm for sizing software products. ACM SIG-METRICS Performance Evaluation Review, 12(2):13–22, Spring-Summer 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fishwick:1984:PPG

[360] Paul A. Fishwick and Stefan Feyock. PROFGEN: a procedure for generating machine independent high-level language profilers. ACM SIGMETRICS Performance Evaluation Review, 12(2): 27–31, Spring-Summer 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rajaraman:1984:PML

[361] M. K. Rajaraman. Performance measures for a local network. ACM SIG-METRICS Performance Evaluation Review, 12(2):34–37, Spring-Summer 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jones:1984:PEJ

[362] Greg A. Jones. Performance evaluation of a job scheduler. ACM SIGMETRICS Performance Evaluation Review, 12(2): 38–43, Spring-Summer 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Clark:1984:NCP

[363] Jon D. Clark and Thomas C. Richards. A note on the cost-performance ratios of IBM’s 43XX series. ACM SIG-METRICS Performance Evaluation Review, 12(2):44–45, Spring-Summer 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coffman:1984:RPP

[364] E. G. Coffman, Jr. Recent progress in the performance evaluation of fundamental allocation algorithms. ACM SIGMETRICS Performance Evaluation Review, 12(3):2–6, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ferrari:1984:FAW

[365] Domenico Ferrari. On the foundations of artificial workload design. ACM SIGMETRICS Performance Evaluation Review, 12(3):8–14, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Perez-Davila:1984:PIF

[366] Alfredo de J. Perez-Davila and Lawrence W. Dowdy. Parameter interdependencies of file placement models in a Unix system. ACM SIGMET-RICS Performance Evaluation Review, 12(3):15–26, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bunt:1984:MPL

[367] Richard B. Bunt, Jennifer M. Murphy, and Shikharesh Majumdar. A measure of program locality and its application. ACM SIGMETRICS Performance Evaluation Review, 12(3):28–40, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krzesinski:1984:ILM

[368] A. Krzesinski and J. Greyling. Improved lineariser methods for queueing networks with queue dependent centres. ACM SIGMETRICS Performance Evaluation Review, 12(3):41–51, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zahorjan:1984:ILD

[369] John Zahorjan and Edward D. Lazowska. Incorporating load dependent servers in approximate mean value analysis. ACM SIGMETRICS Performance Evaluation Review, 12(3):52–62, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Agrawal:1984:RTP

[370] Subhash C. Agrawal, Jeffrey P. Buzen, and Annie W. Shum. Response Time Preservation: a general technique for developing approximate algorithms for queueing networks. ACM SIGMET-RICS Performance Evaluation Review, 12(3):63–77, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mussi:1984:EPE

[371] Ph. Mussi and Ph. Nain. Evaluation of parallel execution of program tree structures. ACM SIGMET-RICS Performance Evaluation Review, 12(3):78–87, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sanguinetti:1984:POP

[372] John Sanguinetti. Program optimization for a pipelined machine a case study. ACM SIGMETRICS Performance Evaluation Review, 12(3):88–95, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Turner:1984:PDB

[373] Rollins Turner, Jeffrey Schriesheim, and Indrajit Mitra. Performance of a DECnet based disk block server. ACM SIGMETRICS Performance Evaluation Review, 12(3):96–104, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stavenow:1984:TDC

[374] Bengt Stavenow. Throughput-delay characteristics and stability considerations of the access channel in a mobile telephone system. ACM SIGMET-RICS Performance Evaluation Review, 12(3):105–112, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Williams:1984:PQD

[375] Elizabeth Williams. Processor queueing disciplines in distributed systems. ACM SIGMETRICS Performance Evaluation Review, 12(3):113– 119, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stephens:1984:CBH

[376] Lindsey E. Stephens and Lawrence W. Dowdy. Convolutional bound hierarchies. ACM SIGMETRICS Performance Evaluation Review, 12(3):120– 133, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Suri:1984:NBB

[377] Rajan Suri and Gregory W. Diehl. A new ‘building block’ for performance evaluation of queueing networks with finite buffers. ACM SIGMETRICS Performance Evaluation Review, 12 (3):134–142, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lavenberg:1984:SAE

[378] Stephen S. Lavenberg. A simple analysis of exclusive and shared lock contention in a database system. ACM SIGMETRICS Performance Evaluation Review, 12(3):143–148, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Becker:1984:MMS

[379] S. T. Becker, K. M. Rege, and B. Sengupta. A modeling methodology for sizing a computer based system in a netted environment. ACM SIGMET-RICS Performance Evaluation Review, 12(3):149–157, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Peachey:1984:EIS

[380] Darwyn R. Peachey, Richard B. Bunt, Carey L. Williamson, and Tim B. Brecht. An experimental investigation of scheduling strategies for UNIX. ACM SIGMETRICS Performance Evaluation Review, 12(3):158– 166, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menasce:1984:PEI

[381] Daniel A. Menascé and Leonardo Lellis P. Leite. Performance evaluation of isolated and interconnected token bus local area networks. ACM SIGMET-RICS Performance Evaluation Review, 12(3):167–175, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Agrawal:1984:UAS

[382] Subhash C. Agrawal, Jeffrey P. Buzen, and Ashok K. Thareja. A unified approach to scan time analysis of token rings and polling networks. ACM SIGMETRICS Performance Evaluation Review, 12(3):176–185, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brandwajn:1984:EAM

[383] Alexandre Brandwajn and William M. McCormack. Efficient approximation for models of multiprogramming with shared domains. ACM SIGMETRICS Performance Evaluation Review, 12 (3):186–194, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bondi:1984:RTP

[384] André B. Bondi and Jeffrey P. Buzen. The response times of priority classes under preemptive resume in M/G/m queues. ACM SIGMETRICS Performance Evaluation Review, 12(3):195– 201, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thomasian:1984:AQN

[385] Alexander Thomasian and Paul Bay. Analysis of Queueing Network Models with population size constraints and delayed blocked customers. ACM SIGMETRICS Performance Evaluation Review, 12(3):202–216, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gaffney:1984:IEP

[386] John E. Gaffney. Instruction entropy, a possible measure of program/architecture compatibility. ACM SIGMETRICS Performance Evaluation Review, 12(4):13–18, 1984/1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sauer:1984:NSS

[387] Charles H. Sauer. Numerical solution of some multiple chain queueing networks. ACM SIGMETRICS Performance Evaluation Review, 12(4):19– 28, December 1984/1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thomasian:1984:SCS

[388] Alexander Thomasian and Kameshwar Gargeya. Speeding up computer system simulations using hierarchical modeling. ACM SIGMETRICS Performance Evaluation Review, 12(4):34– 39, December 1984/1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Elshoff:1984:PMP

[389] James L. Elshoff. The PEEK measurement program. ACM SIGMET-RICS Performance Evaluation Review, 12(4):40–53, Winter 1984/1985. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hac:1984:STM

[390] Anna Hác. A survey of techniques for the modeling of serialization delays in computer systems. ACM SIGMETRICS Performance Evaluation Review, 12(4):54–56, December 1984/1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mosleh:1985:BPR

[391] Ali Mosleh, E. Richard Hilton, and Peter S. Browne. Bayesian probabilistic risk analysis. ACM SIGMETRICS Performance Evaluation Review, 13(1): 5–12, June 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gong:1985:CMB

[392] Huisheng Gong and Monika Schmidt. A complexity measure based on selection and nesting. ACM SIGMET-RICS Performance Evaluation Review, 13(1):14–19, June 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Knudson:1985:PMS

[393] Michael E. Knudson. A performance measurement and system evaluation project plan proposal. ACM SIGMET-RICS Performance Evaluation Review, 13(1):20–31, June 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ejiogu:1985:SMS

[394] Lem O. Ejiogu. A simple measure of software complexity. ACM SIGMET-RICS Performance Evaluation Review, 13(1):33–47, June 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Eager:1985:CRI

[395] Derek L. Eager, Edward D. Lazowska, and John Zahorjan. A comparison of receiver-initiated and sender-initiated adaptive load sharing (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 13(2):1–3, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gelernter:1985:ACP

[396] David Gelernter, Sunil Podar, and Hussein G. Badr. An adaptive communications protocol for network computers (extended abstract). ACM SIGMET-RICS Performance Evaluation Review, 13(2):4–5, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gelenbe:1985:ADC

[397] Erol Gelenbe, David Finkel, and Satish K. Tripathi. On the availability of a distributed computer system with failing components. ACM SIGMET-RICS Performance Evaluation Review, 13(2):6–13, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Conway:1985:RNE

[398] A. E. Conway and N. D. Georganas. RECAL — a new efficient algorithm for the exact analysis of multiplechain closed queueing networks (abstract). ACM SIGMETRICS Performance Evaluation Review, 13(2):14, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Balbo:1985:MPS

[399] G. Balbo, S. C. Bruell, and S. Ghanta. Modeling priority schemes. ACM SIGMETRICS Performance Evaluation Review, 13(2):15–26, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Walstra:1985:NNQ

[400] Robbe J. Walstra. Nonexponential networks of queues: a maximum entropy analysis. ACM SIGMETRICS Performance Evaluation Review, 13(2):27–37, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Calzarossa:1985:SSC

[401] Maria Calzarossa and Domenico Ferrari. A sensitivity study of the clustering approach to workload modeling (extended abstract). ACM SIGMET-RICS Performance Evaluation Review, 13(2):38–39, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raghavan:1985:CIU

[402] S. V. Raghavan and R. Kalyanakrishnan. On the classification of interactive user behaviour indices. ACM SIGMETRICS Performance Evaluation Review, 13(2):40–48, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Verkamo:1985:ERL

[403] A. Inkeri Verkamo. Empirical results on locality in database referencing. ACM SIGMETRICS Performance Evaluation Review, 13(2):49–58, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Khelalfa:1985:DCS

[404] Halin M. Khelalfa and Anneliese K. von Mayrhauser. Degradable computer systems with dependent subsystems. ACM SIGMETRICS Performance Evaluation Review, 13(2):59–68, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chillarege:1985:ESW

[405] Ram Chillarege and Ravishankar K. Lyer. The effect of system workload on error latency: an experimental study. ACM SIGMETRICS Performance Evaluation Review, 13(2):69–77, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gonsalves:1985:PCT

[406] Timothy A. Gonsalves. Performance characteristics of two Ethernets: an experimental study. ACM SIGMET-RICS Performance Evaluation Review, 13(2):78–86, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chlamtac:1985:PIS

[407] I. Chlamtac and M. Eisinger. Performance of integrated services (voice/ data) CSMA/CD networks. ACM SIGMETRICS Performance Evaluation Review, 13(2):87–93, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chlamtac:1985:AMH

[408] I. Chlamtac and M. Eisinger. An analytic model of the hyperchannel network using multiple channel architecture. ACM SIGMETRICS Performance Evaluation Review, 13(2):94– 104, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bleistein:1985:APM

[409] Sandra Bleistein, Shin-Sun Cho, and Robert T. Goettge. Analytic performance model of the U.S. en route air traffic control computer systems. ACM SIGMETRICS Performance Evaluation Review, 13(2):105–115, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dowdy:1985:AUM

[410] Lawrence W. Dowdy and Manvinder S. Chopra. On the applicability of using multiprogramming level distributions. ACM SIGMETRICS Performance Evaluation Review, 13(2):116– 127, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krzesinski:1985:MQN

[411] A. E. Krzesinski and P. Teunissen. Multiclass queueing networks with population constrainted subnetworks. ACM SIGMETRICS Performance Evaluation Review, 13(2):128– 139, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Branwajn:1985:NSI

[412] Alexandre Branwajn and Yung-Li Lily Jow. A note on service interruptions. ACM SIGMETRICS Performance Evaluation Review, 13(2):140– 148, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Plateau:1985:SSP

[413] Brigitte Plateau. On the stochastic structure of parallelism and synchronization models for distributed algorithms. ACM SIGMETRICS Performance Evaluation Review, 13(2):147– 154, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Snyder:1985:ANS

[414] Patricia M. Snyder and William J. Stewart. An approximate numerical solution for multiclass preemptive priority queues with general service time distributions. ACM SIGMET-RICS Performance Evaluation Review, 13(2):155–165, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hevner:1985:EOD

[415] Alan R. Hevner. Evaluation of optical disk systems for very large database applications. ACM SIGMETRICS Performance Evaluation Review, 13 (2):166–172, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Houtekamer:1985:LDC

[416] Gilbert E. Houtekamer. The local disk controller. ACM SIGMETRICS Performance Evaluation Review, 13 (2):173–182, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yu:1985:MCC

[417] Philip S. Yu, Daniel M. Dias, John T. Robinson, Balakrishna R. Iyer, and Douglas Cornell. Modelling of centralized concurrency control in a multisystem environment. ACM SIGMET-RICS Performance Evaluation Review, 13(2):183–191, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thomasian:1985:ASO

[418] Alexander Thomasian and In Kyung Ryu. Analysis of some optimistic concurrency control schemes based on certification. ACM SIGMETRICS Performance Evaluation Review, 13(2):192– 203, August 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ryu:1985:RPA

[419] In Kyung Ryu. Review of ’OS 1100-of performance algorithms: a guide to the resource allocation algorithms of OS-1100’ by John C. Kelly. ACM SIGMETRICS Performance Evaluation Review, 13(3–4):9, November 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Perros:1985:AMF

[420] H. G. Perros and D. Mirchandani. An analytic model of a file server for bulk file transfers. ACM SIGMETRICS Performance Evaluation Review, 13(3– 4):14–22, November 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Domanski:1985:BIS

[421] Bernard Domanski. Building IMS synthetic workloads. ACM SIGMETRICS Performance Evaluation Review, 13(3– 4):23–28, November 1985. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Buzen:1986:MST

[422] Jeffrey P. Buzen. Modeling I/O subsystems (tutorial). ACM SIGMETRICS Performance Evaluation Review, 14(1): 1, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ferrari:1986:WCT

[423] Domenico Ferrari. Workload characterization (tutorial): issues and approaches. ACM SIGMETRICS Performance Evaluation Review, 14(1):1, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goel:1986:SRM

[424] Amrit L. Goel. Software reliability modeling (tutorial). ACM SIGMET-RICS Performance Evaluation Review, 14(1):2, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hedlund:1986:PMI

[425] Kye Hedlund. Performance modeling in integrated curcuit design (tutorial). ACM SIGMETRICS Performance Evaluation Review, 14(1):2, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Artis:1986:ESP

[426] H. Pat Artis. Expert systems for performance analysis (tutorial). ACM SIGMETRICS Performance Evaluation Review, 14(1):3, May 1986. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tripathi:1986:PIL

[427] Satish K. Tripathi. Performance issues in local area networks (tutorial). ACM SIGMETRICS Performance Evaluation Review, 14(1):3, May 1986. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stone:1986:FC

[428] Harold S. Stone and Dominique Thibaut. Footprints in the cache. ACM SIGMETRICS Performance Evaluation Review, 14(1):4–8, May 1986. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vernon:1986:PAM

[429] Mary K. Vernon and Mark A. Holliday. Performance analysis of multiprocessor cache consistency protocols using generalized timed Petri nets. ACM SIGMETRICS Performance Evaluation Review, 14(1):9–17, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harrison:1986:PMP

[430] P. G. Harrison and A. J. Field. Performance modelling of parallel computer architectures. ACM SIGMET-RICS Performance Evaluation Review, 14(1):18–27, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Madnick:1986:MMC

[431] Stuart Madnick and Y. Richard Wang. Modeling multiprocessor computer systems with unbalanced flows. ACM SIGMETRICS Performance Evaluation Review, 14(1):28–34, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kleeman:1986:APB

[432] Lindsay Kleeman and Antonio Cantoni. The analysis and performance of batching arbiters. ACM SIGMET-RICS Performance Evaluation Review, 14(1):35–43, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lehoczky:1986:PRT

[433] John P. Lehoczky and Lui Sha. Performance of real-time bus scheduling algorithms. ACM SIGMETRICS Performance Evaluation Review, 14(1):44–53, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Leland:1986:LBH

[434] Will Leland and Teunis J. Ott. Loadbalancing heuristics and process behavior. ACM SIGMETRICS Performance Evaluation Review, 14(1):54–69, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:1986:CPB

[435] Kyoo Jeong Lee and Don Towsley. A comparison of priority-based decentralized load balancing policies. ACM SIGMETRICS Performance Evaluation Review, 14(1):70–77, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

LeBoudec:1986:BEM

[436] Jean-Yves Le Boudec. A BCMP extension to multiserver stations with concurrent classes of customers. ACM SIGMETRICS Performance Evaluation Review, 14(1):78–91, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Conway:1986:EAS

[437] A. E. Conway and N. D. Georganas. An efficient algorithm for semihomogeneous queueing network models. ACM SIGMETRICS Performance Evaluation Review, 14(1):92–99, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nain:1986:OMH

[438] Philippe Nain and Keith Ross. Optimal multiplexing of heterogeneous traffic with hard constraint. ACM SIGMET-RICS Performance Evaluation Review, 14(1):100–108, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sevcik:1986:CTP

[439] Kenneth Sevcik and Marjory J. Johnson. Cycle time properties of the FDDI token ring protocol (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 14(1):109– 110, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dallery:1986:ADP

[440] Yves Dallery and Rajan Suri. Approximate disaggregation and performance bounds for queueing networks with multiple-server stations. ACM SIGMETRICS Performance Evaluation Review, 14(1):111–128, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Strelen:1986:GMV

[441] Johann Strelen. A generalization of mean value analysis to higher moments: moment analysis. ACM SIGMET-RICS Performance Evaluation Review, 14(1):129–140, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Massey:1986:PAD

[442] William A. Massey. A probabilistic analysis of a database system. ACM SIGMETRICS Performance Evaluation Review, 14(1):141–146, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Witkowski:1986:PEM

[443] Andrew Witkowski. Performance evaluation of multiversion with the Oracle synchronization. ACM SIGMET-RICS Performance Evaluation Review, 14(1):147–158, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Swinghal:1986:PAA

[444] Mukesh Swinghal and A. K. Agrawala. Performance analysis of an algorithm for concurrency control in replicated database systems. ACM SIGMET-RICS Performance Evaluation Review, 14(1):159–169, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Haikala:1986:AMP

[445] Ilkka Haikala. ARMA models of program behaviour. ACM SIGMET-RICS Performance Evaluation Review, 14(1):170–179, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Majumdar:1986:MAL

[446] Shikharesh Majumdar and Richard B. Bunt. Measurement and analysis of locality phases in file referencing behaviour. ACM SIGMETRICS Performance Evaluation Review, 14(1):180– 192, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Razouk:1986:MOS

[447] Rami R. Razouk, Terri Stewart, and Michael Wilson. Measuring operating system performance on modern micro-processors. ACM SIGMET-RICS Performance Evaluation Review, 14(1):193–202, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nicola:1986:QAF

[448] Victor F. Nicola, V. G. Kulkarni, and Kishor S. Trivedi. Queueing analysis of fault-tolerant computer systems (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 14(1): 203, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coffman:1986:ACQ

[449] E. G. Coffman, Jr., E. Gelenbe, and E. N. Gilbert. Analysis of a conveyor queue in a flexible manufacturing system. ACM SIGMETRICS Performance Evaluation Review, 14(1):204– 223, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kouvatsos:1986:MEQ

[450] Demetres D. Kouvatsos. A maximum entropy queue length distribution for the G/G/1 finite capacity queue. ACM SIGMETRICS Performance Evaluation Review, 14(1):224–236, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Takagi:1986:QAN

[451] Hideaki Takagi and Masayuki Murata. Queueing analysis of nonpreemptive reservation priority discipline. ACM SIGMETRICS Performance Evaluation Review, 14(1):237–244, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hofri:1986:QSP

[452] Micha Hofri. Queueing systems with a procrastinating server. ACM SIGMET-RICS Performance Evaluation Review, 14(1):245–253, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Boxma:1986:WTA

[453] O. J. Boxma and B. Meister. Waitingtime approximations for cyclic-service systems with switch-over times. ACM SIGMETRICS Performance Evaluation Review, 14(1):254–262, May 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hu:1986:MFA

[454] Irene Hu. Measuring file access patterns in UNIX. ACM SIGMET-RICS Performance Evaluation Review, 14(2):15–20, August 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ferrari:1986:CIP

[455] Domenico Ferrari. Considerations on the insularity of performance evaluation. ACM SIGMETRICS Performance Evaluation Review, 14(2):21–32, August 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tripathi:1987:RWD

[456] Satish K. Tripathi, Steve Kaisler, Sharat Chandran, and Ashok K. Agrawala. Report on the Workshop on Design & Performance Issues in Parallel Architectures. ACM SIGMET-RICS Performance Evaluation Review, 14(3–4):16–32, January 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gray:1987:VDS

[457] Jim Gray. A view of database system performance measures. ACM SIGMETRICS Performance Evaluation Review, 15(1):3–4, May 1987. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Heidelberger:1987:PCM

[458] Philip Heidelberger and Seetha Lakshmi. A performance comparison of multi-micro and mainframe database architectures. ACM SIGMETRICS Performance Evaluation Review, 15(1): 5–6, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reed:1987:PRA

[459] Daniel A. Reed and Chong kwon Kim. Packet routing algorithms for integrated switching networks. ACM SIGMETRICS Performance Evaluation Review, 15(1):7–15, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gonsalves:1987:PEV

[460] Timothy A. Gonsalves and Fouad A. Tobagi. Performance of the Expressnet with voice/data traffic. ACM SIGMETRICS Performance Evaluation Review, 15(1):16–26, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Agrawal:1987:ARD

[461] Subhash Agrawal and Ravi Ramaswamy. Analysis of the resequencing delay for M/M/m systems. ACM SIGMETRICS Performance Evaluation Review, 15(1):27–35, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reed:1987:PDE

[462] Daniel A. Reed, Allen D. Malony, and Bradley D. McCredie. Parallel discrete event simulation: a shared memory approach. ACM SIGMETRICS Performance Evaluation Review, 15(1):36–38, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bucher:1987:CLV

[463] Ingrid Y. Bucher and Margaret L. Simmons. A close look at vector performance of register-to-register vector computers and a new model. ACM SIGMETRICS Performance Evaluation Review, 15(1):39–45, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Darema-Rogers:1987:MAP

[464] F. Darema-Rogers, G. F. Pfister, and K. So. Memory access patterns of parallel scientific programs. ACM SIGMETRICS Performance Evaluation Review, 15(1):46–58, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Geist:1987:DSS

[465] Robert Geist, Robert Reynolds, and Eve Pittard. Disk scheduling in System V. ACM SIGMETRICS Performance Evaluation Review, 15(1):59–68, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Livny:1987:MDM

[466] Miron Livny, Setrag Khoshafian, and Haran Boral. Multi-disk management algorithms. ACM SIGMETRICS Performance Evaluation Review, 15(1):69– 77, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Buzen:1987:UOT

[467] Jeffrey P. Buzen and Annie W. Shum. A unified operational treatment of RPS reconnect delays. ACM SIGMET-RICS Performance Evaluation Review, 15(1):78–92, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nelson:1987:PAP

[468] R. Nelson, D. Towsley, and A. N. Tantawi. Performance analysis of parallel processing systems. ACM SIGMETRICS Performance Evaluation Review, 15(1):93–94, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:1987:RDR

[469] Ziao-Nan Tan and Kenneth C. Sevcik. Reduced distance routing in single-state shuffle-exchange interconnection networks. ACM SIGMET-RICS Performance Evaluation Review, 15(1):95–110, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bouras:1987:QDB

[470] Christos Bouras and John Garofalakis. Queueing delays in buffered multistage interconnection networks. ACM SIGMETRICS Performance Evaluation Review, 15(1):111–121, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Garcia-Molina:1987:PTM

[471] Hector Garcia-Molina and Lawrence R. Rogers. Performance through memory. ACM SIGMETRICS Performance Evaluation Review, 15(1):122– 131, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jipping:1987:PPC

[472] Michael J. Jipping and Ray Ford. Predicting performance of concurrency control designs. ACM SIGMETRICS Performance Evaluation Review, 15 (1):132–142, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dahbura:1987:PAF

[473] Anton T. Dahbura, Krishan K. Sabnani, and William J. Hery. Performance analysis of a fault detection scheme in multiprocessor systems. ACM SIGMETRICS Performance Evaluation Review, 15(1):143– 154, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Salsburg:1987:SAC

[474] Michael A. Salsburg. A statistical approach to computer performance modeling. ACM SIGMETRICS Performance Evaluation Review, 15(1):155– 162, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kerola:1987:MPM

[475] Teemu Kerola and Herb Schwetman. Monit: a performance monitoring tool for parallel and pseudo-parallel programs. ACM SIGMETRICS Performance Evaluation Review, 15(1):163– 174, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marsan:1987:MSA

[476] M. Ajmone Marsan, G. Balbo, G. Chiola, and G. Conte. Modeling the software architecture of a prototype parallel machine. ACM SIGMET-RICS Performance Evaluation Review, 15(1):175–185, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Alexander:1987:WCP

[477] William Alexander, Tom W. Keller, and Ellen E. Boughter. A workload characterization pipeline for models of parallel systems. ACM SIGMET-RICS Performance Evaluation Review, 15(1):186–194, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Graf:1987:TBD

[478] Ingrid M. Graf. Transformation between different levels of workload characterization for capacity planning: fundamentals and case study. ACM SIGMETRICS Performance Evaluation Review, 15(1):195–204, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ruan:1987:PAF

[479] Zuwang Ruan and Walter F. Tichy. Performance analysis of file replication schemes in distributed systems. ACM SIGMETRICS Performance Evaluation Review, 15(1):205–215, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cheriton:1987:NMV

[480] David R. Cheriton and Carey L. Williamson. Network measurement of the VMTP request-response protocol in the V distributed system. ACM SIGMETRICS Performance Evaluation Review, 15(1):216–225, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Salehmohamed:1987:PEL

[481] Mohamed Salehmohamed, W. S. Luk, and Joseph G. Peters. Performance evaluation of LAN sorting algorithms. ACM SIGMETRICS Performance Evaluation Review, 15(1):226– 233, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Polyzos:1987:DAW

[482] George C. Polyzos and Mart L. Molle. Delay analysis of a window tree conflict resolution algorithm in a local area network environment. ACM SIGMET-RICS Performance Evaluation Review, 15(1):234–244, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shenker:1987:SCB

[483] Scott Shenker. Some conjectures on the behavior of acknowledgementbased transmission control of random access communication channels. ACM SIGMETRICS Performance Evaluation Review, 15(1):245–255, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mathys:1987:ECE

[484] Peter Mathys and Boi V. Faltings. The effect of channel-exit protocols on the performance of finite population random-access systems. ACM SIGMETRICS Performance Evaluation Review, 15(1):256–267, May 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fisher:1987:IIA

[485] Robert Fisher. The impact of interactive application development with CODESTAR. ACM SIGMETRICS Performance Evaluation Review, 15 (2):13–15, August 1987. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Korner:1988:EED

[486] Ulf Körner, Serge Fdida, Harry Perros, and Gerald Shapiro. End to end delays in a catenet environment. ACM SIGMETRICS Performance Evaluation Review, 15(3–4):20–28, February 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sharma:1988:TSA

[487] Ravi S. Sharma. Three simple algorithms for the N/1/F Problem. ACM SIGMETRICS Performance Evaluation Review, 15(3–4):29–32, February 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Covington:1988:RPP

[488] R. C. Covington, S. Madala, V. Mehta, J. R. Jump, and J. B. Sinclair. The Rice Parallel Processing Testbed. ACM SIGMETRICS Performance Evaluation Review, 16(1):4–11, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lubachevsky:1988:EDE

[489] B. D. Lubachevsky. Efficient distributed event driven simulations of multiple-loop networks. ACM SIG-METRICS Performance Evaluation Review, 16(1):12–24, May 1988. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lucier:1988:PEM

[490] B. J. Lucier. Performance evaluation for multiprocessors programmed using monitors. ACM SIGMETRICS Performance Evaluation Review, 16(1):22–29, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ganz:1988:QAF

[491] A. Ganz and I. Chlamtac. Queueing analysis of finite buffer token networks. ACM SIGMETRICS Performance Evaluation Review, 16(1):30–36, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zafirovic-Vukotic:1988:PMH

[492] M. Zafirovic-Vukotic and I. G. M. M. Niemegeers. Performance modelling of a HSLAN slotted ring protocol. ACM SIGMETRICS Performance Evaluation Review, 16(1):37–46, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chiu:1988:CSD

[493] D.-M. Chiu and R. Sudama. A case study of DECnet applications and protocol performance. ACM SIGMET-RICS Performance Evaluation Review, 16(1):47–55, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shenker:1988:AAL

[494] S. Shenker and A. Weinrib. Asymptotic analysis of large heterogeneous queueing systems. ACM SIGMET-RICS Performance Evaluation Review, 16(1):56–62, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Eager:1988:LPB

[495] D. L. Eager, E. D. Lazowska, and J. Zahorjan. The limited performance benefits of migrating active processes for load sharing. ACM SIGMET-RICS Performance Evaluation Review, 16(1):63–72, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hong:1988:LGA

[496] J. Hong, X. Tan, and M. Chen. From local to global: an analysis of nearest neighbor balancing on hypercube. ACM SIGMETRICS Performance Evaluation Review, 16(1):73–82, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kant:1988:ALM

[497] K. Kant. Application level modeling of parallel machines. ACM SIGMET-RICS Performance Evaluation Review, 16(1):83–93, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Born:1988:ADP

[498] R. G. Born and J. R. Kenevan. Analytic derivation of processor potential utilization in straight line, ring, square mesh, and hypercube networks. ACM SIGMETRICS Performance Evaluation Review, 16(1):94–103, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Majumdar:1988:SMP

[499] S. Majumdar, D. L. Eager, and R. B. Bunt. Scheduling in multiprogrammed parallel systems. ACM SIGMET-RICS Performance Evaluation Review, 16(1):104–113, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Patel:1988:HSC

[500] N. M. Patel and P. G. Harrison. On hot-spot contention in interconnection networks. ACM SIGMET-RICS Performance Evaluation Review, 16(1):114–123, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kothari:1988:PAM

[501] S. C. Kothari, A. Jhunjhunwala, and A. Mukherjee. Performance analysis of multipath multistage interconnection networks. ACM SIGMET-RICS Performance Evaluation Review, 16(1):124–132, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Melus:1988:MPE

[502] J. L. Melús, E. Sanvicente, and J. Magriñá. Modelling and performance evaluation of multiprocessor based packet switches. ACM SIGMETRICS Performance Evaluation Review, 16(1):133– 140, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:1988:MCP

[503] T. P. Lee. A manufacturing capacity planning experiment through functional workload decomposition. ACM SIGMETRICS Performance Evaluation Review, 16(1):141–150, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Irgon:1988:FLS

[504] A. E. Irgon, A. H. Dragoni, Jr., and T. O. Huleatt. FAST: a large scale expert system for application and system software performance tuning. ACM SIGMETRICS Performance Evaluation Review, 16(1):151–156, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Alexander:1988:CDC

[505] W. Alexander and G. Copeland. Comparison of dataflow control techniques in distributed data-intensive systems. ACM SIGMETRICS Performance Evaluation Review, 16(1):157– 166, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Leutenegger:1988:MVP

[506] S. T. Leutenegger and M. K. Vernon. A mean-value performance analysis of a new multiprocessor architecture. ACM SIGMETRICS Performance Evaluation Review, 16(1):167–176, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Blake:1988:SAR

[507] J. T. Blake, A. L. Reibman, and K. S. Trivedi. Sensitivity analysis of reliability and performability measures for multiprocessor systems. ACM SIGMETRICS Performance Evaluation Review, 16(1):177–186, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mukkamala:1988:DPR

[508] R. Mukkamala, S. C. Bruell, and R. K. Shultz. Design of partially replicated distributed database systems: an integrated methodology. ACM SIGMET-RICS Performance Evaluation Review, 16(1):187–196, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wybranietz:1988:MPM

[509] D. Wybranietz and D. Haban. Monitoring and performance measuring distributed systems during operation. ACM SIGMETRICS Performance Evaluation Review, 16(1):197– 206, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Melvin:1988:UMI

[510] S. W. Melvin and Y. N. Patt. The use of microcode instrumentation for development, debugging and tuning of operating system kernels. ACM SIGMET-RICS Performance Evaluation Review, 16(1):207–214, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Agawal:1988:MRC

[511] A. Agawal and A. Gupta. Memoryreference characteristics of multiprocessor applications under MACH. ACM SIGMETRICS Performance Evaluation Review, 16(1):215–225, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Murphy:1988:CPB

[512] J. M. Murphy and R. B. Bunt. Characterising program behaviour with phases and transitions. ACM SIGMETRICS Performance Evaluation Review, 16 (1):226–234, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yoshizawa:1988:ASC

[513] Y. Yoshizawa and T. Arai. Adaptive storage control for page frame supply in large scale computer systems. ACM SIGMETRICS Performance Evaluation Review, 16(1):235–243, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pattipati:1988:PAM

[514] K. R. Pattipati and M. M. Kostreva. On the properties of approximate mean value analysis algorithms for queueing networks. ACM SIGMETRICS Performance Evaluation Review, 16(1):244– 252, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tantawi:1988:OAM

[515] A. N. Tantawi, G. Towsley, and J. Wolf. Optimal allocation of multiple class resources in computer systems. ACM SIGMETRICS Performance Evaluation Review, 16(1):253–260, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hsieh:1988:PNA

[516] C.-H. Hsieh and S. S. Lam. PAM — a noniterative approximate solution method for closed multichain queueing networks. ACM SIGMETRICS Performance Evaluation Review, 16(1):261– 269, May 1988. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hac:1989:LBD

[517] Anna Hać. Load balancing in distributed systems: a summary. ACM SIGMETRICS Performance Evaluation Review, 16(2–4):17–19, February 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hac:1989:KBD

[518] Anna Hać. Knowledge-based distributed system architecture. ACM SIGMETRICS Performance Evaluation Review, 16(2–4):20, February 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hac:1989:DAA

[519] Anna Hać. Design algorithms for asynchronous operations in cache memory. ACM SIGMETRICS Performance Evaluation Review, 16(2–4):21, February 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schneider:1989:AHS

[520] Victor Schneider. Approximations for the Halstead software science software error rate and project effort estimators. ACM SIGMETRICS Performance Evaluation Review, 16(2–4):22– 29, February 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Domanski:1989:PBE

[521] Bernard Domanski. A PROLOGbased expert system for tuning MVS/ XA. ACM SIGMETRICS Performance Evaluation Review, 16(2–4):30– 47, February 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Irvin:1989:QML

[522] David R. Irvin. A queueing model for local area network bridges. ACM SIGMETRICS Performance Evaluation Review, 16(2–4):48–57, February 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wolf:1989:POP

[523] J. Wolf. The placement optimization program: a practical solution to the disk file assignment problem. ACM SIGMETRICS Performance Evaluation Review, 17(1):1–10, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kearns:1989:DDR

[524] J. P. Kearns and S. DeFazio. Diversity in database reference behavior. ACM SIGMETRICS Performance Evaluation Review, 17(1):11–19, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hellerstein:1989:SAD

[525] J. Hellerstein. A statistical approach to diagnosing intermittent performanceproblems using monotone relationships. ACM SIGMETRICS Performance Evaluation Review, 17(1):20–28, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Muntz:1989:BAR

[526] R. R. Muntz, E. de Souza e Silva, and A. Goyal. Bounding availability of repairable computer systems. ACM SIGMETRICS Performance Evaluation Review, 17(1):29–38, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bubenik:1989:POM

[527] R. Bubenik and W. Zwaenepoel. Performance of optimistic make. ACM SIGMETRICS Performance Evaluation Review, 17(1):39–48, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anderson:1989:PIT

[528] T. E. Anderson, D. D. Lazowska, and H. M. Levy. The performance implications of thread management alternatives for shared-memory multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 17(1):49–60, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Carter:1989:OIB

[529] J. B. Carter and W. Zwaenepoel. Optimistic implementation of bulk data transfer protocols. ACM SIGMET-RICS Performance Evaluation Review, 17(1):61–69, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stunkel:1989:TPT

[530] C. B. Stunkel and W. K. Fuchs. TRAPEDS: producing traces for multicomputers via execution driven simulation. ACM SIGMETRICS Performance Evaluation Review, 17(1):70–78, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gallivan:1989:BCM

[531] K. Gallivan, D. Gannon, W. Jalby, A. Malony, and H. Wijshoff. Behavioral characterization of multiprocessor memory systems: a case study. ACM SIGMETRICS Performance Evaluation Review, 17(1):79–88, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Samples:1989:MNL

[532] A. D. Samples. Mache: no-loss trace compaction. ACM SIGMETRICS Performance Evaluation Review, 17(1):89– 97, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mukherjee:1989:ERS

[533] A. Mukherjee, L. H. Landweber, and J. C. Strikwerda. Evaluation of retransmission strategies in a local area network environment. ACM SIGMET-RICS Performance Evaluation Review, 17(1):98–107, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Danzig:1989:FBF

[534] P. B. Danzig. Finite buffers for fast multicast. ACM SIGMETRICS Performance Evaluation Review, 17 (1):108–117, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mukherjee:1989:PDB

[535] B. Mukherjee. Performance of a dualbus unidirectional broadcast network operating under probabilistic scheduling strategy. ACM SIGMETRICS Performance Evaluation Review, 17 (1):118–126, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Greenberg:1989:SCP

[536] A. G. Greenberg and J. McKenna. Solution of closed, product form, queueing networks via the RECAL and tree-RECAL methods on a shared memory multiprocessor. ACM SIGMET-RICS Performance Evaluation Review, 17(1):127–135, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Paterok:1989:FQP

[537] M. Paterok, O. Fischer, and L. Opta. Feedback queues with preemptiondistance priorities. ACM SIGMET-RICS Performance Evaluation Review, 17(1):136–145, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wagner:1989:PSQ

[538] D. B. Wagner and E. D. Lazowska. Parallel simulation of queueing networks: limitations and potentials. ACM SIGMETRICS Performance Evaluation Review, 17(1):146–155, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mitra:1989:CCP

[539] D. Mitra and I. Mitrani. Control and coordination policies for systems with buffers. ACM SIGMET-RICS Performance Evaluation Review, 17(1):156–164, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nicol:1989:AMP

[540] D. M. Nicol and J. C. Townsend. Accurate modeling of parallel scientific computations. ACM SIGMET-RICS Performance Evaluation Review, 17(1):165–170, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sevcik:1989:CPA

[541] K. C. Sevcik. Characterizations of parallelism in applications and their use in scheduling. ACM SIGMET-RICS Performance Evaluation Review, 17(1):171–180, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nelson:1989:ART

[542] R. D. Nelson and T. K. Philips. An approximation to the response time for shortest queue routing. ACM SIGMETRICS Performance Evaluation Review, 17(1):181–189, May 1989.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raatikainen:1989:ART

[543] K. E. E. Raatikainen. Approximating response time distributions. ACM SIGMETRICS Performance Evaluation Review, 17(1):190–199, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mitra:1989:CND

[544] D. Mitra and A. Weiss. A closed network with a discriminatory processorsharing server. ACM SIGMETRICS Performance Evaluation Review, 17 (1):200–208, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Glew:1990:EII

[545] Andy Glew. An empirical investigation of OR indexing. ACM SIGMET-RICS Performance Evaluation Review, 17(2):41–49, January 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gunther:1990:PP

[546] N. J. Gunther. Performance pathways. ACM SIGMETRICS Performance Evaluation Review, 17(2):50–56, January 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gonzales:1990:CHL

[547] Michael G. Gonzales. Correction of the Halstead length estimator skew for small Pascal programs. ACM SIGMETRICS Performance Evaluation Review, 17(2):57–59, January 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Greenberg:1990:UPS

[548] Albert G. Greenberg, Boris D. Lubachevsky, and Isi Mitrani. Unboundedly parallel simulations via recurrence relations. ACM SIGMET-RICS Performance Evaluation Review, 18(1):1–12, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nelson:1990:PEG

[549] Randolph Nelson. A performance evaluation of a general parallel processing model. ACM SIGMETRICS Performance Evaluation Review, 18(1):13–26, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:1990:ETD

[550] Wen-Hann Wang and Jean-Loup Baer. Efficient trace-driven simulation method for cache performance analysis. ACM SIGMETRICS Performance Evaluation Review, 18(1):27–36, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Eggers:1990:TEI

[551] S. J. Eggers, David R. Keppel, Eric J. Koldinger, and Henry M. Levy. Techniques for efficient inline tracing on a shared-memory multiprocessor. ACM SIGMETRICS Performance Evaluation Review, 18(1):37–47, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Agarwal:1990:BES

[552] Anant Agarwal and Minor Huffman. Blocking: exploiting spatial locality for trace compaction. ACM SIGMET-RICS Performance Evaluation Review, 18(1):48–57, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:1990:BAF

[553] Tein-Hsiang Lin and Kang G. Shin. A Bayesian approach to fault classification. ACM SIGMETRICS Performance Evaluation Review, 18(1):58–66, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Moser:1990:PLA

[554] Louise E. Moser, Vikas Kapur, and P. M. Melliar-Smith. Probabilistic language analysis of weighted voting algorithms. ACM SIGMETRICS Performance Evaluation Review, 18(1):67–73, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:1990:ERA

[555] Peter M. Chen, Garth A. Gibson, Randy H. Katz, and David A. Patterson. An evaluation of redundant arrays of disks using an Amdahl 5890. ACM SIGMETRICS Performance Evaluation Review, 18(1):74–85, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mukherjee:1990:SAF

[556] Amarnath Mukherjee, Lawrence H. Landweber, and John C. Strikwerda. Simultaneous analysis of flow and error control strategies with congestiondependent errors. ACM SIGMET-RICS Performance Evaluation Review, 18(1):86–95, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:1990:QAA

[557] Arthur Y. M. Lin and John A. Silvester. Queueing analysis of an ATM switch with multichannel transmission groups. ACM SIGMETRICS Performance Evaluation Review, 18(1):96– 105, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Johnson:1990:AAR

[558] Theodore Johnson. Approximate analysis of reader and writer access to a shared resource. ACM SIGMET-RICS Performance Evaluation Review, 18(1):106–114, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anderson:1990:QTT

[559] Thomas E. Anderson and Edward D. Lazowska. Quartz: a tool for tuning parallel program performance. ACM SIGMETRICS Performance Evaluation Review, 18(1):115–125, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pattipati:1990:CVA

[560] Krishna R. Pattipati, Joel Wolf, and Somnath Deb. A calculus of variations approach to file allocation problems in computer systems. ACM SIGMET-RICS Performance Evaluation Review, 18(1):126–133, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Robinson:1990:DCM

[561] John T. Robinson and Murthy V. Devarakonda. Data cache management using frequency-based replacement. ACM SIGMETRICS Performance Evaluation Review, 18(1):134– 142, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dan:1990:AAL

[562] Asit Dan and Don Towsley. An approximate analysis of the LRU and FIFO buffer replacement schemes. ACM SIGMETRICS Performance Evaluation Review, 18(1):143–152, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Alonso:1990:AFW

[563] Raphael Alonso and Andrew W. Appel. An advisor for flexible working sets. ACM SIGMETRICS Performance Evaluation Review, 18(1):153– 162, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Torrellas:1990:ACA

[564] Joseph Torrellas, John Hennessy, and Thierry Weil. Analysis of critical architectural and programming parameters in a hierarchical. ACM SIGMET-RICS Performance Evaluation Review, 18(1):163–172, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jog:1990:PEC

[565] Rajeev Jog, Philip L. Vitale, and James R. Callister. Performance evaluation of a commercial cache-coherent shared memory multiprocessor. ACM SIGMETRICS Performance Evaluation Review, 18(1):173–182, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gelenbe:1990:PAC

[566] Erol Gelenbe. Performance analysis of the Connection Machine. ACM SIGMETRICS Performance Evaluation Review, 18(1):183–191, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Willick:1990:AMM

[567] Darryl L. Willick and D. L. Eager. An analytic model of multistage interconnection networks. ACM SIGMET-RICS Performance Evaluation Review, 18(1):192–202, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dussa:1990:DPT

[568] K. Dussa, B. Carlson, L. Dowdy, and K.-H. Park. Dynamic partitioning in a transputer environment. ACM SIGMETRICS Performance Evaluation Review, 18(1):203–213, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zahorjan:1990:PSS

[569] John Zahorjan and Cathy McCann. Processor scheduling in shared memory multiprocessors. ACM SIGMET-RICS Performance Evaluation Review, 18(1):214–225, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Leutenegger:1990:PMM

[570] Scott T. Leutenegger and Mary K. Vernon. The performance of multiprogrammed multiprocessor scheduling algorithms. ACM SIGMETRICS Performance Evaluation Review, 18(1):226– 236, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dawkins:1990:ESM

[571] W. P. Dawkins, V. Debbad, J. R. Jump, and J. B. Sinclair. Efficient simulation of multiprogramming. ACM SIGMETRICS Performance Evaluation Review, 18(1):237–238, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shenker:1990:MFC

[572] Scott Shenker. Making flow control work in networks: a controltheoretic analysis of gateway service disciplines. ACM SIGMETRICS Performance Evaluation Review, 18(1): 239–240, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shenker:1990:MGW

[573] Scott Shenker. Making greed work in networks: a game-theoretic analysis of gateway service disciplines. ACM SIGMETRICS Performance Evaluation Review, 18(1):241–242, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghandeharizadeh:1990:FAP

[574] Shahram Ghandeharizadeh and David J. DeWitt. Factors affecting the performance of multiuser database management systems. ACM SIGMET-RICS Performance Evaluation Review, 18(1):243–244, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Englert:1990:BNS

[575] Susanne Englert, Jim Gray, Terrye Kocher, and Praful Shah. A benchmark of NonStop SQL release 2 demonstrating near-linear speedup and scaleup on large databases. ACM SIGMET-RICS Performance Evaluation Review, 18(1):245–246, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Somani:1990:PMR

[576] Arun K. Somani, James A. Ritcey, and Stephen H. L. Au. Phased mission reliability analysis. ACM SIGMET-RICS Performance Evaluation Review, 18(1):247–248, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mitchell:1990:PAF

[577] Lionel C. Mitchell. Performance analysis of a fault tolerant computer system. ACM SIGMETRICS Performance Evaluation Review, 18(1):249– 250, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jensen:1990:RTD

[578] David W. Jensen and Daniel A. Reed. Ray tracing on distributed memory parallel systems. ACM SIGMET-RICS Performance Evaluation Review, 18(1):251–252, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mirchandani:1990:CME

[579] Dinesh Mirchandani and Prabuddha Biswas. Characterizing and modeling Ethernet performance of distributed DECwindows applications. ACM SIGMETRICS Performance Evaluation Review, 18(1):253–254, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

McGehearty:1990:COPa

[580] Patrick F. McGehearty. Challenges in obtaining peak parallel performance with a Convex C240, parallel vector processor. ACM SIGMETRICS Performance Evaluation Review, 18 (1):255–256, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Heimlich:1990:TCN

[581] Steven A. Heimlich. Traffic characterization of the NSFNET national backbone. ACM SIGMETRICS Performance Evaluation Review, 18(1):257– 258, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Davidson:1990:EEA

[582] Jack W. Davidson and David B. Whalley. Ease: an environment for architecture study and experimentation. ACM SIGMETRICS Performance Evaluation Review, 18(1):259–260, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Waclawsky:1990:DQB

[583] John G. Waclawsky and Ashok K. Agrawala. Dynamic queue behavior in networks with window protocols. ACM SIGMETRICS Performance Evaluation Review, 18(1):261–262, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Garofalakis:1990:PMI

[584] John D. Garofalakis and Paul G. Spirakis. The performance of multistage interconnection networks with finite buffers. ACM SIGMETRICS Performance Evaluation Review, 18(1):263– 264, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vasilakos:1990:AWF

[585] Athanasios V. Vasilakos, Christos A. Moschonas, and Constantinos T. Paximadis. Adaptive window flow control and learning algorithms for adaptive routing in data networks. ACM SIGMETRICS Performance Evaluation Review, 18(1):265–266, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nussbaum:1990:MCS

[586] Daniel Nussbaum, Ingmar Vuong-Adlerberg, and Anant Agarwal. Modeling a circuit switched multiprocessor interconnect. ACM SIGMET-RICS Performance Evaluation Review, 18(1):267–269, May 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gaither:1990:ER

[587] Blaine D. Gaither. Editor’s readings. ACM SIGMETRICS Performance Evaluation Review, 18(2):25–26, August 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vance:1990:ARM

[588] R. E. Vance. Article review: ‘A message-based approach to discrete event simulation’ by R. L. Bagrodia, K. M. Chandy, and J. Misra. IEEE Trans. Softw. Eng. SE-13, 6 (June 1987). ACM SIGMETRICS Performance Evaluation Review, 18(2):27, August 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Allen:1990:AMS

[589] Arnold O. Allen and Gary Hynes. Approximate MVA solutions with fixed throughput classes. ACM SIGMET-RICS Performance Evaluation Review, 18(2):31–40, August 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

McGehearty:1990:COPb

[590] Patrick McGehearty. Challenges in obtaining peak parallel performance with a Convex C240, a parallel vector processor. ACM SIGMETRICS Performance Evaluation Review, 18(2):41–47, August 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gaither:1990:SVP

[591] Blaine Gaither. Scientific visualization of performance data: evaluation of DV-Draw. ACM SIGMETRICS Performance Evaluation Review, 18(2):48–53, August 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Taheri:1990:ANN

[592] H. Reza Taheri. An analysis of the Neal Nelson Business Benchmark. ACM SIGMETRICS Performance Evaluation Review, 18(3):13–18, November 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Keller:1990:SBC

[593] Tom Keller. SPEC benchmarks and competitive results. ACM SIGMET-RICS Performance Evaluation Review, 18(3):19–20, November 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1990:BRCa

[594] David Finkel. Book review: ‘Computer Systems Performance Management and Capacity Planning’ by J. Cady and B. Howarth (Prentice-Hall, 1990). ACM SIGMETRICS Performance Evaluation Review, 18(3):21, November 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1990:BRA

[595] David Finkel. Book review: ‘The Art of Computer Systems Performance Analysis’ by R. Jain (Wiley-Interscience, 1991). ACM SIGMETRICS Performance Evaluation Review, 18(3):21–22, November 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1990:BRCb

[596] David Finkel. Book review: ‘Computer and Communication Systems Performance Modelling’ by Peter J. B. King (Prentice Hall 1990). ACM SIGMET-RICS Performance Evaluation Review, 18(3):22, November 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1990:BRQ

[597] David Finkel. Book review: ‘Quantitative Analysis of Computer Systems’ by C. H. C. Leung (Wiley, 1988). ACM SIGMETRICS Performance Evaluation Review, 18(3):22–23, November 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1990:BRF

[598] David Finkel. Book review: ‘Fundamentals of Performance Modeling’ by M. K. Molloy (Macmillan, 1989). ACM SIGMETRICS Performance Evaluation Review, 18(3):23, November 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Saavedra-Barrera:1990:MCB

[599] Rafael Saavedra-Barrera, Alan J. Smith, and Eugene Miya. Machine characterization based on an abstract high-level language machine. ACM SIGMETRICS Performance Evaluation Review, 18(3):24, November 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Panwar:1990:OSP

[600] Shivendra S. Panwar, Don Towsley, and Jack K. Wolf. Optimal scheduling policies for a class of queues with customer deadlines to the beginning of service. ACM SIGMET-RICS Performance Evaluation Review, 18(3):25, November 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tokuda:1990:RTM

[601] Hideyuki Tokuda, Makato Kotera, and Clifford E. Mercer. A real-time monitor for a distributed real-time operating system. ACM SIGMETRICS Performance Evaluation Review, 18(3):26, November 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thiebaut:1990:FDC

[602] Dominique Thiebaut. On the fractal dimension of computer programs and its application to the prediction of the cache miss ratio. ACM SIGMET-RICS Performance Evaluation Review, 18(3):41, November 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ponder:1990:PVA

[603] Carl Ponder. Performance variation across benchmark suites. ACM SIGMETRICS Performance Evaluation Review, 18(3):42–48, November 1990. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1991:BRMa

[604] David Finkel. Book review: ‘Multiprocessor Performance’ by Erol Gelenbe (John Wiley & Sons, 1989). ACM SIGMETRICS Performance Evaluation Review, 18(4):9, April 1991. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1991:BRPb

[605] David Finkel. Book review: ‘Performance Analysis of Transaction Processing Systems’ by Wilbur H. Highleyman (Prentice Hall, 1989). ACM SIGMETRICS Performance Evaluation Review, 18(4):10, April 1991. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1991:BRPa

[606] David Finkel. Book review: ‘Performance Measurement of Computer Systems’ by Phillip McKerrow (Addison-Wesley 1988). ACM SIGMETRICS Performance Evaluation Review, 18(4): 10–11, April 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1991:BRMb

[607] David Finkel. Book review: ‘Multiple Access Protocols: Performance and Analysis’ by Raphael Rom and Moshe Sidi (Springer-Verlag, 1990). ACM SIGMETRICS Performance Evaluation Review, 18(4):11, April 1991. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1991:BRPc

[608] David Finkel. Book review: ‘Performance Engineering of Software Systems’ by Connie U. Smith (Addison-Wesley, 1990). ACM SIGMETRICS Performance Evaluation Review, 18(4): 11–12, April 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Johari:1991:POH

[609] Shyam Johari. Performance objectives — how to define them. ACM SIGMET-RICS Performance Evaluation Review, 18(4):18–19, April 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ponder:1991:BS

[610] Carl G. Ponder. Benchmark semantics. ACM SIGMETRICS Performance Evaluation Review, 18(4):20–24, April 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cabrera:1991:TSS

[611] Luis-Felipe Cabrera. Technical summary of the Second IEEE Workshop on Workstation Operating Systems. ACM SIGMETRICS Performance Evaluation Review, 18(4):25–32, April 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Melliar-Smith:1991:PAB

[612] P. M. Melliar-Smith and Louise E. Moser. Performance analysis of a broadcast communications protocol. ACM SIGMETRICS Performance Evaluation Review, 19(1):1–10, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Danzig:1991:AMO

[613] Peter B. Danzig. An analytical model of operating system protocol processing including effects of multiprogramming. ACM SIGMETRICS Performance Evaluation Review, 19(1):11–20, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harinarayan:1991:LSL

[614] Venkatesh Harinarayan and Leonard Kleinrock. Load sharing in limited access distributed systems. ACM SIGMETRICS Performance Evaluation Review, 19(1):21–30, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:1991:SPA

[615] Tein-Hsiang Lin and Wernhuar Tarng. Scheduling periodic and aperiodic tasks in hard real-time computing systems. ACM SIGMETRICS Performance Evaluation Review, 19(1):31–38, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berry:1991:ADC

[616] Robert Berry and Joseph Hellerstein. An approach to detecting changes in the factors affecting the performance of computer systems. ACM SIGMET-RICS Performance Evaluation Review, 19(1):39–49, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bodnarchuk:1991:SWM

[617] Robert Bodnarchuk and Richard Bunt. A synthetic workload model for a distributed system file server. ACM SIGMETRICS Performance Evaluation Review, 19(1):50–59, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Merchant:1991:MCA

[618] Arif Merchant. A Markov chain approximation for the analysis of Banyan networks. ACM SIGMETRICS Performance Evaluation Review, 19(1):60–67, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:1991:PAF

[619] T. Lin and L. Kleinrock. Performance analysis of finite-buffered multistage interconnection networks with a general traffic pattern. ACM SIGMET-RICS Performance Evaluation Review, 19(1):68–78, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wood:1991:MET

[620] David A. Wood, Mark D. Hill, and R. E. Kessler. A model for estimating trace-sample miss ratios. ACM SIGMETRICS Performance Evaluation Review, 19(1):79–89, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chiang:1991:EMV

[621] Mee-Chow Chiang and Gurindar S. Sohi. Experience with mean value analysis model for evaluating shared bus, throughput-oriented multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 19(1):90– 100, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:1991:PAT

[622] Anurag Gupta, Ian Akyildiz, and Richard M. Fujimoto. Performance analysis of Time Warp with homogeneous processors and exponential task times. ACM SIGMETRICS Performance Evaluation Review, 19(1):101– 110, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kim:1991:SDH

[623] Jong Kim and Chita R. Das. On subcube dependability in a hypercube. ACM SIGMETRICS Performance Evaluation Review, 19(1):111– 119, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:1991:IOS

[624] Anoop Gupta, Andrew Tucker, and Shigeru Urushibara. The impact of operating system scheduling policies and synchronization methods of performance of parallel applications. ACM SIGMETRICS Performance Evaluation Review, 19(1):120–132, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhou:1991:PPB

[625] Songnian Zhou and Timothy Brecht. Processor-pool-based scheduling for large-scale NUMA multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 19(1):133–142, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:1991:ATM

[626] Mark S. Squillante and Randolph D. Nelson. Analysis of task migration in shared-memory multiprocessor scheduling. ACM SIGMETRICS Performance Evaluation Review, 19 (1):143–155, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dan:1991:AMH

[627] Asit Dan, Daniel M. Dias, and Philip S. Yu. Analytical modelling of a hierarchical buffer for a data sharing environment. ACM SIGMETRICS Performance Evaluation Review, 19(1):156– 167, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reiman:1991:PAC

[628] Martin Reiman and Paul E. Wright. Performance analysis of concurrentread exclusive-write. ACM SIGMET-RICS Performance Evaluation Review, 19(1):168–177, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

French:1991:PMP

[629] James C. French, Terrence W. Pratt, and Mriganka Das. Performance measurement of a parallel input/output system for the Intel iPSC/2 Hypercube. ACM SIGMETRICS Performance Evaluation Review, 19(1):178– 187, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chervenak:1991:PDA

[630] Ann L. Chervenak and Randy H. Katz. Performance of a disk array prototype. ACM SIGMETRICS Performance Evaluation Review, 19(1):188– 197, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:1991:PMD

[631] Shenze Chen and Don Towsley. Performance of a mirrored disk in a real-time transaction system. ACM SIGMET-RICS Performance Evaluation Review, 19(1):198–207, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Glenn:1991:IMP

[632] R. R. Glenn and D. V. Pryor. Instrumentation for a massively parallel MIMD application. ACM SIGMET-RICS Performance Evaluation Review, 19(1):208–209, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goldberg:1991:MMD

[633] Aaron Goldberg and John Hennessy. MTOOL: a method for detecting memory bottlenecks. ACM SIGMET-RICS Performance Evaluation Review, 19(1):210–211, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kim:1991:ISS

[634] Yul H. Kim, Mark D. Hill, and David A. Wood. Implementing stack simulation for highly-associative memories. ACM SIGMETRICS Performance Evaluation Review, 19(1):212– 213, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Newman:1991:PAC

[635] Robb Newman. Performance analysis case study (abstract): application of experimental design & statistical data analysis techniques. ACM SIGMET-RICS Performance Evaluation Review, 19(1):214–215, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Park:1991:MPB

[636] Arvin Park and Jeffrey C. Becker. Measurements of the paging behavior of UNIX. ACM SIGMETRICS Performance Evaluation Review, 19(1):216– 217, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pasquale:1991:SDW

[637] Joseph Pasquale, Barbara Bittel, and Daniel Kraiman. A static and dynamic workload characterization study of the San Diego Supercomputer Center Cray X-MP. ACM SIGMET-RICS Performance Evaluation Review, 19(1):218–219, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pu:1991:EMA

[638] Calton Pu, Frederick Korz, and Robert C. Lehman. An experiment on measuring application performance over the Internet. ACM SIGMET-RICS Performance Evaluation Review, 19(1):220–221, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:1991:PBB

[639] Myung K. Yang and Chita R. Das. A parallel branch-and-bound algorithm for MIN-based multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 19(1):222–223, May 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Epema:1991:BRC

[640] Dick H. J. Epema. Book review: ‘Computer and Communication Systems Performance Modelling’ by Peter J. B. King (Prentice Hall, 1990). ACM SIGMETRICS Performance Evaluation Review, 19(2):4–5, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Al-Jaar:1991:BRA

[641] Robert Y. Al-Jaar. Book review: ‘The Art of Computer Systems Performance Analysis: Techniques for Experimental Design, Measurement, Simulation, and Modeling’ by Raj Jain (John Wiley & Sons 1991). ACM SIGMET-RICS Performance Evaluation Review, 19(2):5–11, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1991:BRPd

[642] David Finkel. Brief review: ‘Probability, Statistics and Queueing Theory with Computer Science Applications,’ Second Edition by Arnold O. Allen (Academic Press 1990). ACM SIGMETRICS Performance Evaluation Review, 19(2):11–12, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1991:BRC

[643] David Finkel. Brief review: ‘Computer Networks & Systems: Queueing Theory and Performance Evaluation’ by Thomas Robertazzi (Springer-Verlag, 1990). ACM SIGMETRICS Performance Evaluation Review, 19(2):12, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1991:BRQ

[644] David Finkel. Brief review: “Queueing Networks — Exact Computational Algorithms: A Unified Theory Based on Decomposition and Aggregation” by Adrian E. Conway and Nicholas D. Georganas (MIT Press 1989). ACM SIGMETRICS Performance Evaluation Review, 19(2):12, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1991:BRPe

[645] David Finkel. Brief review: ‘Performance Instrumentation & Visualization’ by Margaret Simmons and Rebecca Koskela (Addison-Wesley & ACM Press, 1989). ACM SIGMET-RICS Performance Evaluation Review, 19(2):12–13, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1991:BRS

[646] David Finkel. Brief review: ‘Stochastic Analysis of Computer and Communication Systems’, Ed. by H. Takagi (Elsevier Science Publishers B.V. 1990). ACM SIGMETRICS Performance Evaluation Review, 19(2):13, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Frankel:1991:BRQ

[647] David Frankel. Brief review: ‘Queueing Analysis: A Foundation of Performance Evaluation. Volume 1: Vacation and Priority Systems, Part 1’ by H. Takagi (North-Holland, 1991). ACM SIGMETRICS Performance Evaluation Review, 19(2):13, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ames:1991:CTP

[648] D. Ames, D. Gibson, and B. Troy. Composite theoretical performance. ACM SIGMETRICS Performance Evaluation Review, 19(2):24–29, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Christianson:1991:ALE

[649] Bruce Christianson. Amdahl’s Law and end of system design. ACM SIGMET-RICS Performance Evaluation Review, 19(2):30–32, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1991:OWB

[650] David Finkel, Robert E. Kinicki, and Jonas A. Lehmann. An overview of the WPI Benchmark Suite. ACM SIGMETRICS Performance Evaluation Review, 19(2):33–35, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Becker:1991:APB

[651] Jeffrey C. Becker and Arvin Park. Analysis of the paging behavior of UNIX. ACM SIGMETRICS Performance Evaluation Review, 19(2):36–43, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fateyev:1991:CEA

[652] A. E. Fateyev, S. M. Porotskiy, and V. I. Drujinin. Comparative evaluation of approximate methods for modelling of network systems. ACM SIGMET-RICS Performance Evaluation Review, 19(2):44–48, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nangia:1992:BRP

[653] Ashvini Nangia. Book review: ‘Performance Analysis of Transaction Processing Systems’ by Wilbur H. Highleyman (Prentice Hall, 1989). ACM SIGMET-RICS Performance Evaluation Review, 19(3):9–11, February 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Meng:1992:BRC

[654] Xiannong Meng. Book review: ‘Computer Networks and Systems: Queueing Theory and Performance Evaluation’ by Thomas G. Robertazzi (Springer-Verlag, 1990). ACM SIG-METRICS Performance Evaluation Review, 19(3):11–12, February 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1992:BRS

[655] ?. Finkel. Brief review: ‘Stochastic Modeling and the Theory of Queues’ by Ronald W. Wolfe (Prentice-Hall, 1989). ACM SIGMETRICS Performance Evaluation Review, 19(3):12, February 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1992:BRMa

[656] ?. Finkel. Brief review: ‘Markovian Queues’ by O. P. Sharma (Ellis Horwood Publishers 1990). ACM SIGMETRICS Performance Evaluation Review, 19(3):12–13, February 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1992:BRB

[657] ?. Finkel. Brief review: ‘The Benchmark Handbook: Database and Transaction Processing Systems,’ Ed. by Jim Gray (Morgan Kaufmann Publishers, Inc., 1991). ACM SIGMET-RICS Performance Evaluation Review, 19(3):13, February 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1992:BRMb

[658] ?. Finkel. Brief review: “Modeling and Analysis of Local Area Networks” by Paul J. Fortier and George Desrochers (CRC Press, 1990). ACM SIGMET-RICS Performance Evaluation Review, 19(3):13, February 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berry:1992:SWC

[659] Michael W. Berry. Scientific workload characterization by loop-based analyses. ACM SIGMETRICS Performance Evaluation Review, 19(3):17–29, February 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Council:1992:CTR

[660] Corporate Transaction Processing Performance Council. Complete TPC results (as of 9/30/91). ACM SIGMET-RICS Performance Evaluation Review, 19(3):32–35, February 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Deike-Glindemann:1992:SPE

[661] Hartmut Deike-Glindemann. SIQUEUE-PET: an environment for queueing network modelling. ACM SIGMETRICS Performance Evaluation Review, 19 (3):36–44, February 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dujmovic:1992:UMS

[662] Jozo J. Dujmović. The use of multiplesubscripted arrays in benchmark programs. ACM SIGMETRICS Performance Evaluation Review, 19(3):45–48, February 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pooley:1992:BRC

[663] Rob Pooley. Book reviews: ‘Computer and Communication Systems Performance Modelling’ by Peter J. B. King (Prentice Hall 1990). ACM SIGMET-RICS Performance Evaluation Review, 19(4):13–14, May 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hac:1992:MDF

[664] Anna Hac. Modeling distributed file systems. ACM SIGMETRICS Performance Evaluation Review, 19(4):22–27, May 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Molloy:1992:ANB

[665] Michael K. Molloy. Anatomy of the NHFSSTONES benchmarks. ACM SIGMETRICS Performance Evaluation Review, 19(4):28–39, May 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Keown:1992:RTP

[666] William F. Keown, Jr., Philip Koopman, Jr., and Aaron Collins. Realtime performance of the HARRIS RTX 2000 stack architecture versus the Sun 4 SPARC and the Sun 3 M68020 architectures with a proposed real-time performance benchmark. ACM SIGMET-RICS Performance Evaluation Review, 19(4):40–48, May 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Martonosi:1992:MAM

[667] Margaret Martonosi, Anoop Gupta, and Thomas Anderson. MemSpy: analyzing memory system bottlenecks in programs. ACM SIGMETRICS Performance Evaluation Review, 20(1):1– 12, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Whalley:1992:FIC

[668] David B. Whalley. Fast instruction cache performance evaluation using compile-time analysis. ACM SIGMET-RICS Performance Evaluation Review, 20(1):13–22, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

LaRowe:1992:ADP

[669] Richard P. LaRowe, Jr., Mark A. Holliday, and Carla Schlatter Ellis. An analysis of dynamic page placement on a NUMA multiprocessor. ACM SIGMETRICS Performance Evaluation Review, 20(1):23–34, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nicola:1992:AGC

[670] Victor F. Nicola, Asit Dan, and Daniel M. Dias. Analysis of the generalized clock buffer replacement scheme for database transaction processing. ACM SIGMETRICS Performance Evaluation Review, 20(1):35–46, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Borst:1992:CCC

[671] S. C. Borst, O. J. Boxma, and M. B. Combé. Collection of customers: a correlated M/G/1 queue. ACM SIGMET-RICS Performance Evaluation Review, 20(1):47–59, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jacquet:1992:STD

[672] Philippe Jacquet. Subexponential tail distribution in LaPalice queues. ACM SIGMETRICS Performance Evaluation Review, 20(1):60–69, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:1992:RBC

[673] Duan-Shin Lee and Bhaskar Sengupta. A reservation based cyclic server queue with limited service. ACM SIGMET-RICS Performance Evaluation Review, 20(1):70–77, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ramakrishnan:1992:AFT

[674] K. K. Ramakrishnan, Prabuddha Biswas, and Ramakrishna Karedla. Analysis of file I/O traces in commercial computing environments. ACM SIGMETRICS Performance Evaluation Review, 20(1):78–90, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sandhu:1992:CBF

[675] Harjinder S. Sandhu and Songnian Zhou. Cluster-based file replication in large-scale distributed systems. ACM SIGMETRICS Performance Evaluation Review, 20(1):91–102, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Merchant:1992:PAD

[676] Arif Merchant, Kun-Lung Wu, Philip S. Yu, and Ming-Syan Chen. Performance analysis of dynamic finite versioning for concurrent transaction and query processing. ACM SIGMET-RICS Performance Evaluation Review, 20(1):103–114, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thomasian:1992:PAL

[677] Alexander Thomasian. Performance analysis of locking policies with limited wait depth. ACM SIGMET-RICS Performance Evaluation Review, 20(1):115–127, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kurose:1992:CPS

[678] Jim Kurose. On computing per-session performance bounds in high-speed multi-hop computer networks. ACM SIGMETRICS Performance Evaluation Review, 20(1):128–139, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lui:1992:AAB

[679] John C. S. Lui and Richard R. Muntz. Algorithmic approach to bounding the mean response time of a minimum expected delay routing system. ACM SIGMETRICS Performance Evaluation Review, 20(1):140–151, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

deSouzaeSilva:1992:SSE

[680] Edmundo de Souza e Silva and Pedro Mejiá Ochoa. State space exploration in Markov models. ACM SIGMETRICS Performance Evaluation Review, 20(1):152–166, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Owicki:1992:FPA

[681] Susan S. Owicki and Anna R. Karlin. Factors in the performance of the AN1 computer network. ACM SIGMET-RICS Performance Evaluation Review, 20(1):167–180, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shankar:1992:PCR

[682] A. Udaya Shankar, Cengiz Alaettinoğlu, Ibrahim Matta, and Klaudia Dussa-Zieger. Performance comparison of routing protocols using MaRS: distance-vector versus link-state. ACM SIGMETRICS Performance Evaluation Review, 20(1):181–192, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Altman:1992:CLC

[683] Eitan Altman and Philippe Nain. Closed-loop control with delayed information. ACM SIGMETRICS Performance Evaluation Review, 20(1):193– 204, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Merchant:1992:AMC

[684] Arif Merchant. Analytical models of combining Banyan networks. ACM SIGMETRICS Performance Evaluation Review, 20(1):205–212, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Akyildiz:1992:PAT

[685] Ian F. Akyildiz, Liang Chen, Samir R. Das, Richard M. Fujimoto, and Richard F. Serfozo. Performance analysis of “Time Warp” with limited memory. ACM SIGMETRICS Performance Evaluation Review, 20(1):213– 224, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Turek:1992:SPT

[686] John Turek, Joel L. Wolf, Krishna R. Pattipati, and Philip S. Yu. Scheduling parallelizable tasks: putting it all on the shelf. ACM SIGMET-RICS Performance Evaluation Review, 20(1):225–236, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bremaud:1992:SLR

[687] P. Brémaud and W.-B. Gong. Stationary likelihood ratios and smoothed perturbation analysis gradient estimates for the routing problem. ACM SIGMETRICS Performance Evaluation Review, 20(1):237–238, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Candlin:1992:SPP

[688] Rosemary Candlin, Peter Fisk, Joe Phillips, and Neil Skilling. Studying the performance properties of concurrent programs by simulation experiments on synthetic programs. ACM SIGMET-RICS Performance Evaluation Review, 20(1):239–240, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berry:1992:CIP

[689] Robert F. Berry and Joseph L. Hellerstein. Characterizing and interpreting periodic behavior in computer systems. ACM SIGMETRICS Performance Evaluation Review, 20(1):241– 242, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rahm:1992:HPC

[690] Erhard Rahm and Donald Ferguson. High performance cache management for sequential data access. ACM SIGMETRICS Performance Evaluation Review, 20(1):243–244, June 1992.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chakka:1992:MSG

[691] Ram Chakka and Isi Mitrani. Multiprocessor systems with general breakdowns and repairs (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 20(1):245– 246, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brewer:1992:PHP

[692] Eric A. Brewer, Chrysanthos N. Dellarocas, Adrian Colbrook, and William E. Weihl. PROTEUS: a highperformance parallel-architecture simulator. ACM SIGMETRICS Performance Evaluation Review, 20(1):247– 248, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Meliksetian:1992:PAC

[693] Dikran S. Meliksetian and C. Y. Roger Chen. Performance analysis of communications in static interconnection networks. ACM SIGMETRICS Performance Evaluation Review, 20(1):249– 250, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dan:1992:CDA

[694] Asit Dan, Philip S. Yu, and Jen-Yao Chung. Characterization of database access skew in a transaction processing environment. ACM SIGMET-RICS Performance Evaluation Review, 20(1):251–252, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:1992:XPE

[695] Aloke Gupta and Wen-Mei W. Hwu. Xprof: profiling the execution of X Window programs. ACM SIGMET-RICS Performance Evaluation Review, 20(1):253–254, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shoham:1992:ETP

[696] Ruth Shoham and Uri Yechiali. Elevator-type polling systems (abstract). ACM SIGMETRICS Performance Evaluation Review, 20(1):255– 257, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baccelli:1992:PSS

[697] François Baccelli and Miguel Canales. Parallel simulation of stochastic Petri nets using recurrence equations. ACM SIGMETRICS Performance Evaluation Review, 20(1):257–258, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jobmann:1992:PAP

[698] Manfred R. Jobmann and Johann Schumann. Performance analysis of a parallel theorem prover. ACM SIGMETRICS Performance Evaluation Review, 20(1):259–260, June 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shanley:1992:TRN

[699] Kim Shanley and Amie Belongia. TPC releases new benchmark: TPC-C. ACM SIGMETRICS Performance Evaluation Review, 20(2):8–22, November 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pooley:1992:BRP

[700] Rob Pooley. Book review: ‘Performance Engineering of Software Systems’ by Connie U. Smith (Addison Wesley 1990). ACM SIGMETRICS Performance Evaluation Review, 20 (2):23–24, November 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Taylor:1992:BRQ

[701] Stephen Taylor. Book review: “Queuing Networks — Exact Computational Algorithms: A Unified Theory Based on Decomposition and Aggregation” by Adrian E. Conway and Nicolas D. Georganas (MIT Press 1989). ACM SIGMETRICS Performance Evaluation Review, 20(2):24–26, November 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kobayashi:1992:CMM

[702] Makoto Kobayashi. A cache multitasking model. ACM SIGMETRICS Performance Evaluation Review, 20 (2):27–37, November 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Porotskiy:1992:DTM

[703] S. M. Porotskiy and A. E. Fateev. Development trends in methods for efficiency evaluation of ES-based computer systems. ACM SIGMETRICS Performance Evaluation Review, 20 (2):38–42, November 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Porotskiy:1992:SRP

[704] S. M. Porotskiy and A. E. Fateev. System and real performance evaluation of computer. ACM SIGMET-RICS Performance Evaluation Review, 20(2):43–46, November 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

vandeLiefvoort:1993:BRM

[705] Apple van de Liefvoort. Book review: “Multiple Access Protocols: Performance and Analysis” by Raphael Rom and Moshe Sidi (Springer-Verlag, 1990). ACM SIGMETRICS Performance Evaluation Review, 20(3):5–6, March 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

TPC:1993:STRa

[706] Corporate TPC. Summary of TPC results (as of December 22, 1992). ACM SIGMETRICS Performance Evaluation Review, 20(3):7–21, March 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Maffeis:1993:FAP

[707] Silvano Maffeis. File access patterns in public FTP archives and an index for locality of reference. ACM SIGMET-RICS Performance Evaluation Review, 20(3):22–35, March 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ulusoy:1993:AAR

[708] Özgür Ulusoy. An approximate analysis of a real-time database concurrency control protocol via Markov modeling. ACM SIGMETRICS Performance Evaluation Review, 20(3):36–48, March 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

IBM:1993:SP

[709] Corporate IBM Systems Analysis Department. Selected publications: 1992. ACM SIGMETRICS Performance Evaluation Review, 20(4):3–9, May 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

TPC:1993:STRb

[710] Corporate TPC. Summary of TPC results (as of March 15, 1993). ACM SIGMETRICS Performance Evaluation Review, 20(4):10–23, May 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raatikainen:1993:CAW

[711] Kimmo E. E. Raatikainen. Cluster analysis and workload classification. ACM SIGMETRICS Performance Evaluation Review, 20(4):24–30, May 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Smith:1993:EPP

[712] Robert B. Smith, James K. Archibald, and Brent E. Nelson. Evaluating performance of prefetching second level caches. ACM SIGMETRICS Performance Evaluation Review, 20(4):31–42, May 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:1993:NAP

[713] Peter M. Chen and David A. Patterson. A new approach to I/O performance evaluation: self-scaling I/O benchmarks, predicted I/O performance. ACM SIGMETRICS Performance Evaluation Review, 21(1):1–12, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Biswas:1993:TDA

[714] Prabuddha Biswas, K. K. Ramakrishnan, and Don Towsley. Trace driven analysis of write caching policies for disks. ACM SIGMETRICS Performance Evaluation Review, 21(1):13–23, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sugumar:1993:ESC

[715] Rabin A. Sugumar and Santosh G. Abraham. Efficient simulation of caches under optimal replacement with applications to miss characterization. ACM SIGMETRICS Performance Evaluation Review, 21(1):24–35, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chame:1993:CIP

[716] Jacqueline Chame and Michel Dubois. Cache inclusion and processor sampling in multiprocessor simulations. ACM SIGMETRICS Performance Evaluation Review, 21(1):36–47, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reinhardt:1993:WWT

[717] Steven K. Reinhardt, Mark D. Hill, James R. Larus, Alvin R. Lebeck, James C. Lewis, and David A. Wood. The Wisconsin Wind Tunnel: virtual prototyping of parallel computers. ACM SIGMETRICS Performance Evaluation Review, 21(1):48–60, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Adve:1993:IRD

[718] Vikram S. Adve and Mary K. Vernon. The influence of random delays on parallel execution times. ACM SIGMET-RICS Performance Evaluation Review, 21(1):61–73, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rosti:1993:KEM

[719] E. Rosti, E. Smirni, T. D. Wagner, A. W. Apon, and L. W. Dowdy. The KSR1: experimentation and modeling of poststore. ACM SIGMET-RICS Performance Evaluation Review, 21(1):74–85, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ganger:1993:PFM

[720] Gregory R. Ganger and Yale N. Patt. The process-flow model: examining I/O performance from the system’s point of view. ACM SIGMETRICS Performance Evaluation Review, 21(1): 86–97, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:1993:APM

[721] Edward K. Lee and Randy H. Katz. An analytic performance model of disk arrays. ACM SIGMETRICS Performance Evaluation Review, 21(1):98– 109, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tang:1993:MMB

[722] Dong Tang and Ravishankar K. Iyer. MEASURE+: a measurement-based dependability analysis package. ACM SIGMETRICS Performance Evaluation Review, 21(1):110–121, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ramesh:1993:STS

[723] A. V. Ramesh and Kishor Trivedi. On the sensitivity of transient solutions of Markov models. ACM SIGMET-RICS Performance Evaluation Review, 21(1):122–134, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nicol:1993:PSM

[724] David M. Nicol and Philip Heidelberger. Parallel simulation of Markovian queueing networks using adaptive uniformization. ACM SIGMET-RICS Performance Evaluation Review, 21(1):135–145, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goldschmidt:1993:ATD

[725] Stephen R. Goldschmidt and John L. Hennessy. The accuracy of trace-driven simulations of multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 21(1):146–157, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Setia:1993:PSM

[726] Sanjeev K. Setia, Mark S. Squillante, and Satish K. Tripathi. Processor scheduling on multiprogrammed, distributed memory parallel computers. ACM SIGMETRICS Performance Evaluation Review, 21(1):158– 170, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wu:1993:PCT

[727] Kun-Lung Wu, Philip S. Yu, and James Z. Teng. Performance comparison of thrashing control policies for concurrent Mergesorts with parallel prefetching. ACM SIGMET-RICS Performance Evaluation Review, 21(1):171–182, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Meliksetian:1993:MMB

[728] Dikran S. Meliksetian and C. Y. Roger Chen. A Markov-modulated Bernoulli process approximation for the analysis of Banyan networks. ACM SIGMET-RICS Performance Evaluation Review, 21(1):183–194, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arakawa:1993:MVR

[729] Hiroshi Arakawa, Daniel I. Katcher, Jay K. Strosnider, and Hideyuki Tokuda. Modeling and validation of the real-time Mach scheduler. ACM SIGMETRICS Performance Evaluation Review, 21(1):195–206, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baruah:1993:RHS

[730] Sanjoy Baruah and Jayant R. Haritsa. ROBUST: a hardware solution to real-time overload. ACM SIGMET-RICS Performance Evaluation Review, 21(1):207–216, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dey:1993:ELP

[731] Jayanta K. Dey, James F. Kurose, Don Towsley, C. M. Krishna, and Mahesh Girkar. Efficient on-line processor scheduling for a class of IRIS (Increasing Reward with Increasing Service) real-time tasks. ACM SIGMET-RICS Performance Evaluation Review, 21(1):217–228, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Morris:1993:ASS

[732] Robert J. T. Morris. Analysis of superposition of streams into a cache buffer. ACM SIGMETRICS Performance Evaluation Review, 21(1):229– 235, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tsai:1993:AMC

[733] Jory Tsai and Anant Agarwal. Analyzing multiprocessor cache behavior through data reference modeling. ACM SIGMETRICS Performance Evaluation Review, 21(1):236–247, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Martonosi:1993:ETS

[734] Margaret Martonosi, Anoop Gupta, and Thomas Anderson. Effectiveness of trace sampling for performance debugging tools. ACM SIGMETRICS Performance Evaluation Review, 21 (1):248–259, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ahn:1993:HTS

[735] Jong-Suk Ahn, Peter B. Danzig, Deborah Estrin, and Brenda Timmerman. Hybrid technique for simulating high bandwidth delay computer networks. ACM SIGMETRICS Performance Evaluation Review, 21(1):260– 261, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Becker:1993:AIC

[736] Jeffrey C. Becker and Arvin Park. An analysis of the information content of address and data reference streams. ACM SIGMETRICS Performance Evaluation Review, 21(1):262– 263, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghandeharizadeh:1993:EAV

[737] Shahram Ghandeharizadeh and Luis Ramos. An evaluation of alternative virtual replication strategies for continuous retrieval of multimedia data. ACM SIGMETRICS Performance Evaluation Review, 21(1):264– 265, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kay:1993:STI

[738] Jonathan Kay and Joseph Pasquale. A summary of TCP/IP networking software performance for the DECstation 5000. ACM SIGMETRICS Performance Evaluation Review, 21(1):266– 267, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lewandowski:1993:AAP

[739] Gary Lewandowski, Anne Condon, and Eric Bach. Asynchronous analysis of parallel dynamic programming. ACM SIGMETRICS Performance Evaluation Review, 21(1):268–269, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shin:1993:ELS

[740] Kang G. Shin and Chao-Ju Hou. Evaluation of load sharing in HARTS while considering message routing and broadcasting. ACM SIGMETRICS Performance Evaluation Review, 21(1):270– 271, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Torrellas:1993:BCA

[741] Josep Torrellas, Andrew Tucker, and Anoop Gupta. Benefits of cache-affinity scheduling in shared-memory multiprocessors: a summary. ACM SIGMET-RICS Performance Evaluation Review, 21(1):272–274, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vetland:1993:CMA

[742] Vidar Vetland, Peter Hughes, and Arne Sølvberg. A composite modelling approach to software performance measurement. ACM SIGMET-RICS Performance Evaluation Review, 21(1):275–276, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wagner:1993:AMV

[743] David B. Wagner. Approximate mean value analysis of interconnection networks with deflection routing. ACM SIGMETRICS Performance Evaluation Review, 21(1):277–278, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Williamson:1993:OFT

[744] Carey L. Williamson. Optimizing file transfer response time using the lossload curve congestion control mechanism. ACM SIGMETRICS Performance Evaluation Review, 21(1):279– 280, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lipsky:1993:BRI

[745] Lester Lipsky. Book review: “Introduction to Computer System Performance Evaluation” by Krishna Kant (McGraw-Hill, 1992). ACM SIGMET-RICS Performance Evaluation Review, 21(2):7–9, December 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kinicki:1993:BRT

[746] Robert E. Kinicki. Book review: “Telecommunications and Networking” by Udo W. Pooch, Denis Machuel and John McCahn (CRC Press, 1991). ACM SIGMETRICS Performance Evaluation Review, 21(2):9–10, December 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cao:1993:SCM

[747] Xiren Cao. Some common misconceptions about performance modeling and validation. ACM SIGMETRICS Performance Evaluation Review, 21 (2):11–15, December 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Maffeis:1993:CMA

[748] Silvano Maffeis. Cache management algorithms for flexible filesystems. ACM SIGMETRICS Performance Evaluation Review, 21(2):16–25, December 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

UI:1993:PMA

[749] UNIX International. Performance management activities within UNIX International. ACM SIGMETRICS Performance Evaluation Review, 21 (2):42, December 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dujmovic:1994:BRB

[750] Jozo J. Dujmović. Book review: “The Benchmarking Handbook for Database and Transaction Processing Systems” Edited by Jim Gray (Morgan Kaufmann Publishers, Inc., 1991). ACM SIGMETRICS Performance Evaluation Review, 21(3–4):4–5, April 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Finkel:1994:BRE

[751] David Finkel. Book review: “Encyclopedia of Computer Science”, Third Edition, edited by Anthony Ralston and Edwin I. Reilly (Van Nostrand Reinhold, 1993). ACM SIGMETRICS Performance Evaluation Review, 21(3– 4):6, April 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schieber:1994:RRT

[752] Colleen D. Schieber and Eric E. Johnson. RATCHET: real-time address trace compression hardware for extended traces. ACM SIGMETRICS Performance Evaluation Review, 21 (3–4):22–32, April 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:1994:SCQ

[753] Surendra M. Gupta and Fikri Karaesmen. Solution to complex queueing systems: a spreadsheet approach. ACM SIGMETRICS Performance Evaluation Review, 21(3–4):33–46, April 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Denning:1994:FLK

[754] Peter J. Denning. The fifteenth level (keynote address). ACM SIGMET-RICS Performance Evaluation Review, 22(1):1–4, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Peris:1994:AIM

[755] Vinod G. J. Peris, Mark S. Squillante, and Vijay K. Naik. Analysis of the impact of memory in distributed parallel processing systems. ACM SIGMETRICS Performance Evaluation Review, 22(1):5–18, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

McCann:1994:PAP

[756] Cathy McCann and John Zahorjan. Processor allocation policies for message-passing parallel computers. ACM SIGMETRICS Performance Evaluation Review, 22(1):19–32, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chiang:1994:UAC

[757] Su-Hui Chiang, Rajesh K. Mansharamani, and Mary K. Vernon. Use of application characteristics and limited preemption for run-to-completion parallel processor scheduling policies. ACM SIGMETRICS Performance Evaluation Review, 22(1):33–44, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wolf:1994:SMQ

[758] Joel L. Wolf, John Turek, Ming-Syan Chen, and Philip S. Yu. Scheduling multiple queries on a parallel machine. ACM SIGMETRICS Performance Evaluation Review, 22(1):45–55, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Patel:1994:AMH

[759] Jignesh M. Patel, Michael J. Carey, and Mary K. Vernon. Accurate modeling of the hybrid hash join algorithm. ACM SIGMETRICS Performance Evaluation Review, 22(1):56–66, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bittan:1994:APB

[760] Avi Bittan, Yaakov Kogan, and Philip S. Yu. Asymptotic performance of a buffer model in a data sharing environment. ACM SIGMETRICS Performance Evaluation Review, 22(1):67–76, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Petriu:1994:AMV

[761] Dorina C. Petriu. Approximate mean value analysis of client-server systems with multi-class requests. ACM SIGMETRICS Performance Evaluation Review, 22(1):77–86, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Balbo:1994:ATP

[762] G. Balbo, S. C. Bruell, and M. Sereno. Arrival theorems for product-form stochastic Petri nets. ACM SIGMET-RICS Performance Evaluation Review, 22(1):87–97, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Somani:1994:PMS

[763] Arun K. Somani and Kishor S. Trivedi. Phased-mission system analysis using Boolean algebraic methods. ACM SIGMETRICS Performance Evaluation Review, 22(1):98–107, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ebling:1994:SEF

[764] Maria R. Ebling and M. Satyanarayanan. SynRGen: an extensible file reference generator. ACM SIGMETRICS Performance Evaluation Review, 22(1):108–117, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raghavan:1994:GNM

[765] S. V. Raghavan, D. Vasukiammaiyar, and Gunter Haring. Generative networkload models for a single server environment. ACM SIGMET-RICS Performance Evaluation Review, 22(1):118–127, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cmelik:1994:SFI

[766] Bob Cmelik and David Keppel. Shade: a fast instruction-set simulator for execution profiling. ACM SIGMET-RICS Performance Evaluation Review, 22(1):128–137, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Noble:1994:ESH

[767] Brian D. Noble and M. Satyanarayanan. An empirical study of a highly available file system. ACM SIGMETRICS Performance Evaluation Review, 22(1):138–149, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dahlin:1994:QAC

[768] Michael D. Dahlin, Clifford J. Mather, Randolph Y. Wang, Thomas E. Anderson, and David A. Patterson. A quantitative analysis of cache policies for scalable network file systems. ACM SIGMETRICS Performance Evaluation Review, 22(1):150–160, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kotz:1994:ELS

[769] David Kotz and Preston Crow. The expected lifetime of “single-address-space” operating systems. ACM SIGMETRICS Performance Evaluation Review, 22(1):161–170, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sivasubramaniam:1994:ASS

[770] Anand Sivasubramaniam, Aman Singla, Umakishore Ramachandran, and H. Venkateswaran. An approach to scalability study of shared memory parallel systems. ACM SIGMET-RICS Performance Evaluation Review, 22(1):171–180, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mehra:1994:CTM

[771] Pankaj Mehra, Catherine H. Schulbach, and Jerry C. Yan. A comparison of two model-based performanceprediction techniques for messagepassing parallel programs. ACM SIGMETRICS Performance Evaluation Review, 22(1):181–190, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Horton:1994:MLS

[772] Graham Horton and Scott T. Leutenegger. A multi-level solution algorithm for steady-state Markov chains. ACM SIGMETRICS Performance Evaluation Review, 22(1):191–200, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Das:1994:AMM

[773] Samir R. Das and Richard M. Fujimoto. An adaptive memory management protocol for Time Warp parallel simulation. ACM SIGMET-RICS Performance Evaluation Review, 22(1):201–210, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:1994:PEE

[774] Hui Zhang and Edward W. Knightly. Providing end-to-end statistical performance guarantees with bounding interval dependent stochastic models. ACM SIGMETRICS Performance Evaluation Review, 22(1):211–220, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pingali:1994:CSI

[775] Sridhar Pingali, Don Towsley, and James F. Kurose. A comparison of sender-initiated and receiver-initiated reliable multicast protocols. ACM SIGMETRICS Performance Evaluation Review, 22(1):221–230, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nikolaidis:1994:TPS

[776] Ioanis Nikolaidis, Richard Fujimoto, and C. Anthony Cooper. Timeparallel simulation of cascaded statistical multiplexers. ACM SIGMET-RICS Performance Evaluation Review, 22(1):231–240, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Worthington:1994:SAM

[777] Bruce L. Worthington, Gregory R. Ganger, and Yale N. Patt. Scheduling algorithms for modern disk drives. ACM SIGMETRICS Performance Evaluation Review, 22(1):241–251, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nicol:1994:OMC

[778] David M. Nicol and Shahid H. Bokhari. Optimal multiphase complete exchange on circuit-switched hypercube architectures. ACM SIGMETRICS Performance Evaluation Review, 22(1):252– 260, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Temam:1994:CIP

[779] O. Temam, C. Fricker, and W. Jalby. Cache interference phenomena. ACM SIGMETRICS Performance Evaluation Review, 22(1):261–271, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Danskin:1994:PXP

[780] John Danskin and Pat Hanrahan. Profiling the X protocol (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 22(1):272– 273, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Drapeau:1994:TWC

[781] Ann L. Drapeau, David A. Patterson, and Randy H. Katz. Toward workload characterization of video server and digital library applications (extended abstract). ACM SIGMET-RICS Performance Evaluation Review, 22(1):274–275, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gill:1994:CSF

[782] Deepinder S. Gill, Songnian Zhou, and Harjinder S. Sandhu. A case study of file system workload in a largescale distributed environment. ACM SIGMETRICS Performance Evaluation Review, 22(1):276–277, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hellerstein:1994:CTD

[783] Joseph L. Hellerstein. A comparison of techniques for diagnosing performance problems in information systems (extended abstract). ACM SIGMET-RICS Performance Evaluation Review, 22(1):278–279, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:1994:EUL

[784] J. William Lee. Efficient userlevel communication on multicomputers with an optimistic flow-control protocol (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 22(1):280–281, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rolia:1994:MRP

[785] J. A. Rolia, M. Starkey, and G. Boersma. Modeling RPC performance. ACM SIGMETRICS Performance Evaluation Review, 22(1):282– 283, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tayyab:1994:SPM

[786] Athar B. Tayyab and Jon G. Kuhl. Stochastic performance models of parallel task systems (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 22(1):284– 285, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Uhlig:1994:KBM

[787] Richard Uhlig, David Nagle, Trevor Mudge, and Stuart Sechrest. Kernelbased memory simulation (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 22(1):286– 287, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wabnig:1994:PPP

[788] Harald Wabnig and Günter Haring. Performance prediction of parallel systems with scalable specifications — methodology and case study. ACM SIGMETRICS Performance Evaluation Review, 22(1):288–289, May 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lavenberg:1995:SPS

[789] S. S. Lavenberg. Selected publications of the Systems Analysis and Systems Applications department of the IBM T. J. Watson Research Center. ACM SIGMETRICS Performance Evaluation Review, 22(2–4):6–17, April 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shanley:1995:TDM

[790] Kim Shanley and Tracy Derossett. TPC-D measures how quickly realworld business questions can be answered. ACM SIGMETRICS Performance Evaluation Review, 22(2–4):18– 45, April 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wabnig:1995:PPP

[791] H. Wabnig and G. Haring. Performance prediction of parallel systems with scalable specifications — methodology and case study. ACM SIGMET-RICS Performance Evaluation Review, 22(2–4):46–62, April 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:1995:QMS

[792] Surendra M. Gupta. Queueing model with state dependent balking and reneging: its complementary and equivalence. ACM SIGMETRICS Performance Evaluation Review, 22(2–4): 63–72, April 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Keehn:1995:VPF

[793] D. G. Keehn. Visualizing performance in the frequency plane. ACM SIGMET-RICS Performance Evaluation Review, 22(2–4):73–81, April 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chapin:1995:MSP

[794] John Chapin, A. Herrod, Mendel Rosenblum, and Anoop Gupta. Memory system performance of UNIX on CC-NUMA multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 23(1):1–13, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bedichek:1995:TFA

[795] Robert C. Bedichek. Talisman: fast and accurate multicomputer simulation. ACM SIGMETRICS Performance Evaluation Review, 23(1):14–24, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Golubchik:1995:RDV

[796] Leana Golubchik, John C. S. Lui, and Richard Muntz. Reducing I/O demand in video-on-demand storage servers. ACM SIGMETRICS Performance Evaluation Review, 23(1):25–36, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghandeharizadeh:1995:CSD

[797] Shahram Ghandeharizadeh, Seon Ho Kim, and Cyrus Shahabi. On configuring a single disk continuous media server. ACM SIGMETRICS Performance Evaluation Review, 23(1):37–46, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krunz:1995:TMC

[798] Marwan Krunz and Herman Hughes. A traffic for MPEG-coded VBR streams. ACM SIGMETRICS Performance Evaluation Review, 23(1):47–55, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Williamson:1995:NTM

[799] Carey L. Williamson. Network traffic measurement and modeling. ACM SIGMETRICS Performance Evaluation Review, 23(1):56–57, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gelenbe:1995:GNN

[800] Erol Gelenbe. G-networks: new queueing models with additional control capabilities. ACM SIGMETRICS Performance Evaluation Review, 23(1):58–59, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tridandapani:1995:FPF

[801] Srini Tridandapani, Anton T. Dahbura, Charles U. Martel, John Matthews, and Arun K. Somani. Free performance and fault tolerance (extended abstract): using system idle capacity efficiently. ACM SIGMETRICS Performance Evaluation Review, 23(1):60– 61, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Malony:1995:DIE

[802] Allen D. Malony. Data interpretation and experiment planning in performance tools. ACM SIGMETRICS Performance Evaluation Review, 23(1): 62–63, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vaidya:1995:CTL

[803] Nitin H. Vaidya. A case for two-level distributed recovery schemes. ACM SIGMETRICS Performance Evaluation Review, 23(1):64–73, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Epema:1995:ADU

[804] D. H. J. Epema. An analysis of decay-usage scheduling in multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 23(1):74–85, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Elwalid:1995:FRP

[805] Anwar Elwalid, Daniel Heyman, T. V. Lakshman, Debasis Mitra, and Alan Weiss. Fundamental results on the performance of ATM multiplexers with applications to video teleconferencing. ACM SIGMETRICS Performance Evaluation Review, 23(1):86–97, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Knightly:1995:FLT

[806] Edward W. Knightly, Dallas E. Wrege, Jörg Liebeherr, and Hui Zhang. Fundamental limits and tradeoffs of providing deterministic guarantees to VBR video traffic. ACM SIGMETRICS Performance Evaluation Review, 23(1):98– 107, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fang:1995:EBW

[807] Youjian Fang, Michael Devetsikiotis, Ioannis Lambadaris, and A. Roger Kaye. Exponential bounds for the waiting time distribution in Markovian queues, with applications to TES/GI/1 systems. ACM SIGMETRICS Performance Evaluation Review, 23(1):108– 115, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Borst:1995:OPA

[808] S. C. Borst. Optimal probabilistic allocation of customer types to servers. ACM SIGMETRICS Performance Evaluation Review, 23(1):116– 125, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Matta:1995:ZIS

[809] Ibrahim Matta and A. Udaya Shankar. Z-iteration: a simple method for throughput estimation in time-dependent multi-class systems. ACM SIGMET-RICS Performance Evaluation Review, 23(1):126–135, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:1995:SRL

[810] Peter M. Chen and Edward K. Lee. Striping in a RAID level 5 disk array. ACM SIGMETRICS Performance Evaluation Review, 23(1):136– 145, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Worthington:1995:LES

[811] Bruce L. Worthington, Gregory R. Ganger, Yale N. Patt, and John Wilkes. On-line extraction of SCSI disk drive parameters. ACM SIGMET-RICS Performance Evaluation Review, 23(1):146–156, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wolf:1995:DDD

[812] Joel L. Wolf, Philip S. Yu, and Hadas Shachnai. DASD dancing: a disk load balancing optimization scheme for video-on-demand computer systems. ACM SIGMETRICS Performance Evaluation Review, 23(1):157– 166, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sandhu:1995:ASD

[813] Harjinder S. Sandhu and Kenneth C. Sevcik. An analytic study of dynamic hardware and software cache coherence strategies. ACM SIGMET-RICS Performance Evaluation Review, 23(1):167–177, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brorsson:1995:SPT

[814] Mats Brorsson. SM-prof: a tool to visualise and find cache coherence performance bottlenecks in multiprocessor programs. ACM SIGMET-RICS Performance Evaluation Review, 23(1):178–187, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cao:1995:SIP

[815] Pei Cao, Edward W. Felten, Anna R. Karlin, and Kai Li. A study of integrated prefetching and caching strategies. ACM SIGMETRICS Performance Evaluation Review, 23(1):188– 197, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sivasubramaniam:1995:CBR

[816] Anand Sivasubramaniam, Aman Singla, Umakishore Ramachandran, and H. Venkateswaran. On characterizing bandwidth requirements of parallel applications. ACM SIGMET-RICS Performance Evaluation Review, 23(1):198–207, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

McCann:1995:SMC

[817] Cathy McCann and John Zahorjan. Scheduling memory constrained jobs on distributed memory parallel computers. ACM SIGMETRICS Performance Evaluation Review, 23(1):208– 219, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lebeck:1995:AMN

[818] Alvin R. Lebeck and David A. Wood. Active memory: a new abstraction for memory-system simulation. ACM SIGMETRICS Performance Evaluation Review, 23(1):220–230, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

deSouzaeSilva:1995:CTD

[819] Edmundo de Souza e Silva, H. Richard Gail, and Reinaldo Vallejos Campos. Calculating transient distributions of cumulative reward. ACM SIGMET-RICS Performance Evaluation Review, 23(1):231–240, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Carrasco:1995:RRT

[820] Juan A. Carrasco and Angel Calderón. Regenerative randomization: theory and application examples. ACM SIGMETRICS Performance Evaluation Review, 23(1):241–252, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Greenberg:1995:CTA

[821] Albert G. Greenberg and R. Srikant. Computational techniques for accurate performance evaluation of multirate, multihop communication networks. ACM SIGMETRICS Performance Evaluation Review, 23(1):253– 260, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ott:1995:IET

[822] Teun Ott. The Internet in evolution, and TCP over ATM (panel session). ACM SIGMETRICS Performance Evaluation Review, 23(1):261– 262, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Trivedi:1995:NMP

[823] Kishor S. Trivedi, A. Bobbio, G. Ciardo, R. German, A. Puliafito, and M. Telek. Non-Markovian Petri nets. ACM SIGMETRICS Performance Evaluation Review, 23(1):263– 264, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Erramilli:1995:PIS

[824] Ashok Erramilli. Performance impacts of self-similarity in traffic. ACM SIGMETRICS Performance Evaluation Review, 23(1):265–266, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arpaci:1995:IPS

[825] Remzi H. Arpaci, Andrea C. Dusseau, Amin M. Vahdat, Lok T. Liu, Thomas E. Anderson, and David A. Patterson. The interaction of parallel and sequential workloads on a network of workstations. ACM SIGMET-RICS Performance Evaluation Review, 23(1):267–278, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Myllymaki:1995:DTJ

[826] Jussi Myllymaki and Miron Livny. Disk-tape joins: synchronizing disk and tape access. ACM SIGMET-RICS Performance Evaluation Review, 23(1):279–290, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Phalke:1995:IRG

[827] Vidyadhar Phalke and Bhaskarpillai Gopinath. An inter-reference gap model for temporal locality in program behavior. ACM SIGMETRICS Performance Evaluation Review, 23(1):291– 300, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Braams:1995:BCP

[828] Jan Braams. Batch class process scheduler for Unix SVR4. ACM SIGMET-RICS Performance Evaluation Review, 23(1):301–302, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Donatelli:1995:SSR

[829] S. Donatelli and G. Franceschinis. State space reductions using stochastic well-formed net simplifications: an application to random polling systems. ACM SIGMETRICS Performance Evaluation Review, 23(1):303– 304, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Balsamo:1995:ART

[830] S. Balsamo and I. Mura. Approximate response time distribution in Fork and Join systems. ACM SIGMET-RICS Performance Evaluation Review, 23(1):305–306, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:1995:SEA

[831] Xiaodong Zhang and Zhichen Xu. A semi-empirical approach to scalability study. ACM SIGMETRICS Performance Evaluation Review, 23(1):307– 308, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hughes:1995:PFP

[832] Eric Hughes and Marianne Winslett. PEDCAD: a framework for performance evaluation of object database applications. ACM SIGMETRICS Performance Evaluation Review, 23 (1):309–310, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Salehi:1995:SCA

[833] James D. Salehi, James F. Kurose, and Don Towsley. Scheduling for cache affinity in parallelized communication protocols. ACM SIGMET-RICS Performance Evaluation Review, 23(1):311–312, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chatterjee:1995:MAM

[834] Amit K. Chatterjee and Vijay K. Konangi. Modeling and analysis of multi channel asymmetric packet switch modules in a bursty and nonuniform traffic environment. ACM SIGMETRICS Performance Evaluation Review, 23(1):313–314, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shah:1995:TNT

[835] Gautam Shah, Umakishore Ramachandran, and Richard Fujimoto. Timepatch: a novel technique for the parallel simulation of multiprocessor caches. ACM SIGMETRICS Performance Evaluation Review, 23(1):315– 316, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sundaram:1995:FAB

[836] C. R. M. Sundaram and Derek L. Eager. Future applicability of bus-based shared memory multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 23(1):317–318, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ciardo:1995:MFC

[837] Gianfranco Ciardo, Ludmila Cherkasova, Vadim Kotov, and Tomas Rokicki. Modeling a Fibre Channel switch with stochastic Petri nets. ACM SIGMET-RICS Performance Evaluation Review, 23(1):319–320, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arunachalam:1995:PPP

[838] Meenakshi Arunachalam and Alok Choudhary. A prefetching prototype for the parallel file systems on the Paragon. ACM SIGMETRICS Performance Evaluation Review, 23(1):321– 322, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gopalakrishnan:1996:BRT

[839] R. Gopalakrishnan and Gurudatta M. Parulkar. Bringing real-time scheduling theory and practice closer for multimedia computing. ACM SIGMETRICS Performance Evaluation Review, 24(1): 1–12, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harchol-Balter:1996:EPL

[840] Mor Harchol-Balter and Allen B. Downey. Exploiting process lifetime distributions for dynamic load balancing. ACM SIGMETRICS Performance Evaluation Review, 24(1):13–24, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dusseau:1996:EDS

[841] Andrea C. Dusseau, Remzi H. Arpaci, and David E. Culler. Effective distributed scheduling of parallel workloads. ACM SIGMETRICS Performance Evaluation Review, 24(1):25–36, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lim:1996:LPB

[842] Beng-Hong Lim and Ricardo Bianchini. Limits on the performance benefits of multithreading and prefetching. ACM SIGMETRICS Performance Evaluation Review, 24(1):37–46, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dinda:1996:FMA

[843] Peter A. Dinda and David R. O’Hallaron. Fast message assembly using compact address relations. ACM SIGMETRICS Performance Evaluation Review, 24(1):47–56, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Parsons:1996:CAM

[844] Eric W. Parsons and Kenneth C. Sevcik. Coordinated allocation of memory and processors in multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 24(1):57–67, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Witchel:1996:EFF

[845] Emmett Witchel and Mendel Rosenblum. Embra: fast and flexible machine simulation. ACM SIGMET-RICS Performance Evaluation Review, 24(1):68–79, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brakmo:1996:ENS

[846] Lawrence S. Brakmo and Larry L. Peterson. Experiences with network simulation. ACM SIGMETRICS Performance Evaluation Review, 24(1):80–90, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Greenberg:1996:AUL

[847] Albert G. Greenberg, S. Shenker, and Alexander L. Stolyar. Asynchronous updates in large parallel systems. ACM SIGMETRICS Performance Evaluation Review, 24(1):91–103, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stiliadis:1996:DAF

[848] Dimitrios Stiliadis and Anujan Varma. Design and analysis of frame-based fair queueing: a new traffic scheduling algorithm for packet-switched networks. ACM SIGMETRICS Performance Evaluation Review, 24(1):104– 115, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yates:1996:NSL

[849] David J. Yates, Erich M. Nahum, James F. Kurose, and Don Towsley. Networking support for large scale multiprocessor servers. ACM SIGMET-RICS Performance Evaluation Review, 24(1):116–125, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arlitt:1996:WSW

[850] Martin F. Arlitt and Carey L. Williamson. Web server workload characterization: the search for invariants. ACM SIGMETRICS Performance Evaluation Review, 24(1):126– 137, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Martonosi:1996:IPM

[851] Margaret Martonosi, David Ofelt, and Mark Heinrich. Integrating performance monitoring and communication in parallel computers. ACM SIGMET-RICS Performance Evaluation Review, 24(1):138–147, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krishnaswamy:1996:MAE

[852] Umesh Krishnaswamy and Isaac D. Scherson. Micro-architecture evaluation using performance vectors. ACM SIGMETRICS Performance Evaluation Review, 24(1):148–159, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Crovella:1996:SSW

[853] Mark E. Crovella and Azer Bestavros. Self-similarity in World Wide Web traffic: evidence and possible causes. ACM SIGMETRICS Performance Evaluation Review, 24(1):160–169, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hillyer:1996:MPC

[854] Bruce K. Hillyer and Avi Silberschatz. On the modeling and performance characteristics of a serpentine tape drive. ACM SIGMET-RICS Performance Evaluation Review, 24(1):170–179, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menasce:1996:AMH

[855] Daniel A. Menascé, Odysseas I. Pentakalos, and Yelena Yesha. An analytic model of hierarchical mass storage systems with network-attached storage devices. ACM SIGMETRICS Performance Evaluation Review, 24(1):180– 189, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:1996:AAW

[856] Ken Chen and Laurent Decreusefond. An approximate analysis of waiting time in multi-class M/G/1/./EDF queues. ACM SIGMETRICS Performance Evaluation Review, 24(1):190– 199, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aggarwal:1996:OPM

[857] Charu Aggarwal, Joel Wolf, and Philip S. Yu. On optimal piggyback merging policies for video-on-demand systems. ACM SIGMET-RICS Performance Evaluation Review, 24(1):200–209, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gerber:1996:EDV

[858] Richard Gerber and Ladan Gharai. Experiments with digital video playback. ACM SIGMETRICS Performance Evaluation Review, 24(1):210– 221, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Salehi:1996:SSV

[859] James D. Salehi, Zhi-Li Zhang, James F. Kurose, and Don Towsley. Supporting stored video: reducing rate variability and end-to-end resource requirements through optimal smoothing. ACM SIGMETRICS Performance Evaluation Review, 24(1):222– 231, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Varki:1996:ABF

[860] Elizabeth Varki and Lawrence W. Dowdy. Analysis of balanced fork-join queueing networks. ACM SIGMET-RICS Performance Evaluation Review, 24(1):232–241, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Carrasco:1996:EEA

[861] Juan A. Carrasco, Javier Escribá, and Angel Calderón. Efficient exploration of availability models guided by failure distances. ACM SIGMET-RICS Performance Evaluation Review, 24(1):242–251, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Garg:1996:MCT

[862] Sachin Garg, Yennun Huang, Chandra Kintala, and Kishor S. Trivedi. Minimizing completion time of a program by checkpointing and rejuvenation. ACM SIGMETRICS Performance Evaluation Review, 24(1):252– 261, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kimbrel:1996:IPP

[863] Tracy Kimbrel, Pei Cao, Edward W. Felten, Anna R. Karlin, and Kai Li. Integrated parallel prefetching and caching. ACM SIGMETRICS Performance Evaluation Review, 24(1):262– 263, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Leutenegger:1996:BME

[864] Scott T. Leutenegger and Mario A. Lopez. A buffer model for evaluating R-tree performance. ACM SIGMET-RICS Performance Evaluation Review, 24(1):264–265, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hellerstein:1996:ASM

[865] Joseph L. Hellerstein. An approach to selecting metrics for detecting performance problems in information systems. ACM SIGMETRICS Performance Evaluation Review, 24(1):266– 267, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Courtright:1996:RRP

[866] William V. Courtright II, Garth Gibson, Mark Holland, and Jim Zelenka. RAIDframe: rapid prototyping for disk arrays. ACM SIGMET-RICS Performance Evaluation Review, 24(1):268–269, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ramany:1996:QAR

[867] Swaminathan Ramany and Derek Eager. Quantifying achievable routing performance in multiprocessor interconnection networks. ACM SIGMET-RICS Performance Evaluation Review, 24(1):270–271, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hotovy:1996:AEW

[868] Steven Hotovy, David Schneider, and Timothy O’Donnell. Analysis of the early workload on the Cornell Theory Center IBM SP2. ACM SIGMET-RICS Performance Evaluation Review, 24(1):272–273, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Braun:1997:APL

[869] Hans-Werner Braun. Architecture and performance of large internets, based on terrestrial and satellite infrastructure. ACM SIGMETRICS Performance Evaluation Review, 25(1):1, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Balakrishnan:1997:ASW

[870] Hari Balakrishnan, Mark Stemm, Srinivasan Seshan, and Randy H. Katz. Analyzing stability in wide-area network performance. ACM SIGMETRICS Performance Evaluation Review, 25(1): 2–12, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Maltzahn:1997:PIE

[871] Carlos Maltzahn, Kathy J. Richardson, and Dirk Grunwald. Performance issues of enterprise level Web proxies. ACM SIGMETRICS Performance Evaluation Review, 25(1):13–23, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Heyman:1997:NMA

[872] D. P. Heyman, T. V. Lakshman, and Arnold L. Neidhardt. A new method for analysing feedback-based protocols with applications to engineering Web traffic over the Internet. ACM SIGMETRICS Performance Evaluation Review, 25(1):24–38, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ma:1997:QME

[873] Qingming Ma and K. K. Ramakrishnan. Queue management for explicit rate based congestion control. ACM SIGMETRICS Performance Evaluation Review, 25(1):39–51, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ott:1997:TAA

[874] Teunis J. Ott and Neil Aggarwal. TCP over ATM: ABR or UBR? ACM SIGMETRICS Performance Evaluation Review, 25(1):52–63, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kasera:1997:SRM

[875] Sneha K. Kasera, Jim Kurose, and Don Towsley. Scalable reliable multicast using multiple multicast groups. ACM SIGMETRICS Performance Evaluation Review, 25(1):64–74, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rajamony:1997:PDS

[876] Ramakrishnan Rajamony and Alan L. Cox. Performance debugging shared memory parallel programs using runtime dependence analysis. ACM SIGMETRICS Performance Evaluation Review, 25(1):75–87, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Herbordt:1997:PSC

[877] Martin C. Herbordt, Owais Kidwai, and Charles C. Weems. Preprototyping SIMD coprocessors using virtual machine emulation and trace compilation. ACM SIGMETRICS Performance Evaluation Review, 25(1):88–99, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tomkins:1997:IMP

[878] Andrew Tomkins, R. Hugo Patterson, and Garth Gibson. Informed multiprocess prefetching and caching. ACM SIGMETRICS Performance Evaluation Review, 25(1):100–114, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Glass:1997:APR

[879] Gideon Glass and Pei Cao. Adaptive page replacement based on memory reference behavior. ACM SIGMET-RICS Performance Evaluation Review, 25(1):115–126, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Voelker:1997:MSL

[880] Geoffrey M. Voelker, Hervé A. Jamrozik, Mary K. Vernon, Henry M. Levy, and Edward D. Lazowska. Managing server load in global memory systems. ACM SIGMETRICS Performance Evaluation Review, 25(1):127– 138, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Woodward:1997:SLB

[881] Michael E. Woodward. Size-limited batch movement in product-form closed discrete-time queueing networks. ACM SIGMETRICS Performance Evaluation Review, 25(1):139– 146, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Golubchik:1997:BPM

[882] Leana Golubchik and John C. S. Lui. Bounding of performance measures for a threshold-based queueing system with hysteresis. ACM SIGMET-RICS Performance Evaluation Review, 25(1):147–157, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lehoczky:1997:URT

[883] John P. Lehoczky. Using real-time queueing theory to control lateness in real-time systems. ACM SIGMET-RICS Performance Evaluation Review, 25(1):158–168, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nahum:1997:CBN

[884] Erich Nahum, David Yates, Jim Kurose, and Don Towsley. Cache behavior of network protocols. ACM SIGMETRICS Performance Evaluation Review, 25(1):169–180, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Knightly:1997:SMR

[885] Edward W. Knightly. Second moment resource allocation in multi-service networks. ACM SIGMETRICS Performance Evaluation Review, 25(1):181– 191, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krunz:1997:CVM

[886] Marwan Krunz and Satish K. Tripathi. On the characterization of VBR MPEG streams. ACM SIGMET-RICS Performance Evaluation Review, 25(1):192–202, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Smith:1997:FSA

[887] Keith A. Smith and Margo I. Seltzer. File system aging — increasing the relevance of file system benchmarks. ACM SIGMETRICS Performance Evaluation Review, 25(1):203–213, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brown:1997:OSB

[888] Aaron B. Brown and Margo I. Seltzer. Operating system benchmarking in the wake of lmbench: a case study of the performance of NetBSD on the Intel x86 architecture. ACM SIGMET-RICS Performance Evaluation Review, 25(1):214–224, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Acharya:1997:UEI

[889] Anurag Acharya, Guy Edjlali, and Joel Saltz. The utility of exploiting idle workstations for parallel computation. ACM SIGMETRICS Performance Evaluation Review, 25(1):225– 234, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Qin:1997:PEC

[890] Xiaohan Qin and Jean-Loup Baer. A performance evaluation of cluster architectures. ACM SIGMETRICS Performance Evaluation Review, 25(1):237– 247, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chiueh:1997:DED

[891] Tzi cker Chiueh and Srinidhi Varadarajan. Design and evaluation of a DRAM-based shared memory ATM switch. ACM SIGMETRICS Performance Evaluation Review, 25(1):248– 259, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Song:1997:ERC

[892] Junehwa Song, Asit Dan, and Dinkar Sitaram. Efficient retrieval of composite multimedia objects in the JINSIL distributed system. ACM SIGMET-RICS Performance Evaluation Review, 25(1):260–271, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gibson:1997:FSS

[893] Garth A. Gibson, David F. Nagle, Khalil Amiri, Fay W. Chang, Eugene M. Feinberg, Howard Gobioff, Chen Lee, Berend Ozceri, Erik Riedel, David Rochberg, and Jim Zelenka. File server scaling with network-attached secure disks. ACM SIGMETRICS Performance Evaluation Review, 25 (1):272–284, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tsiolis:1997:GGC

[894] Athanassios K. Tsiolis and Mary K. Vernon. Group-guaranteed channel capacity in multimedia storage servers. ACM SIGMETRICS Performance Evaluation Review, 25(1):285– 297, June 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Muntz:1997:SIM

[895] Richard Muntz. Special issue on multimedia storage systems. ACM SIGMETRICS Performance Evaluation Review, 25(2):2, September 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ozden:1997:AIM

[896] Banu Özden, Rajeev Rastogi, and Avi Silberschatz. Architecture issues in multimedia storage systems. ACM SIGMETRICS Performance Evaluation Review, 25(2):3–12, September 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shi:1997:BSV

[897] Weifeng Shi and Shahram Ghandeharizadeh. Buffer sharing in videoon-demand servers. ACM SIGMET-RICS Performance Evaluation Review, 25(2):13–20, September 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Golubchik:1997:ITD

[898] Leana Golubchik. On issues and tradeoffs in design of fault tolerant VOD servers. ACM SIGMETRICS Performance Evaluation Review, 25 (2):21–28, September 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Muntz:1997:RRT

[899] Richard Muntz, Jose Renato Santos, and Steve Berson. RIO: a realtime multimedia object server. ACM SIGMETRICS Performance Evaluation Review, 25(2):29–35, September 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Colajanni:1997:ATS

[900] Michele Colajanni and Philip S. Yu. Adaptive TTL schemes for load balancing of distributed Web servers. ACM SIGMETRICS Performance Evaluation Review, 25(2):36–42, September 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kotz:1997:SIP

[901] David Kotz. Special issue on parallel I/O systems. ACM SIGMET-RICS Performance Evaluation Review, 25(3):2, December 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cormen:1997:CFP

[902] Thomas H. Cormen and David M. Nicol. Out-of-core FFTs with parallel disks. ACM SIGMETRICS Performance Evaluation Review, 25(3):3–12, December 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Papadopouli:1997:SVV

[903] Maria Papadopouli and Leana Golubchik. Support of VBR video streams under disk bandwidth limitations. ACM SIGMETRICS Performance Evaluation Review, 25(3):13–20, December 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bordawekar:1997:EEH

[904] Rajesh Bordawekar, Steven Landherr, Don Capps, and Mark Davis. Experimental evaluation of the Hewlett– Packard Exemplar file system. ACM SIGMETRICS Performance Evaluation Review, 25(3):21–28, December 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rochberg:1997:PNE

[905] David Rochberg and Garth Gibson. Prefetching over a network: early experience with CTIP. ACM SIGMET-RICS Performance Evaluation Review, 25(3):29–36, December 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menon:1997:DVD

[906] Jai Menon and Kent Treiber. Daisy: virtual-disk hierarchical storage manager. ACM SIGMETRICS Performance Evaluation Review, 25(3):37–44, December 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nicol:1998:SIT

[907] David M. Nicol. Special issue on the Telecommunications Description Language. ACM SIGMETRICS Performance Evaluation Review, 25(4):3, March 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Perumalla:1998:TLM

[908] Kalyan Perumalla, Richard Fujimoto, and Andrew Ogielski. TED — a language for modeling telecommunication networks. ACM SIGMETRICS Performance Evaluation Review, 25(4):4–11, March 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Perumalla:1998:TMA

[909] Kalyan Perumalla, Matthew Andrews, and Sandeep Bhatt. TED models for ATM internetworks. ACM SIGMET-RICS Performance Evaluation Review, 25(4):12–21, March 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rubenstein:1998:OPS

[910] Dan Rubenstein, Jim Kurose, and Don Towsley. Optimistic parallel simulation of reliable multicast protocols. ACM SIGMETRICS Performance Evaluation Review, 25(4):22–29, March 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Panchal:1998:PSW

[911] Jignesh Panchal, Owen Kelly, Jie Lai, Narayan Mandayam, Andarew T. Ogielski, and Roy Yates. Parallel simulations of wireless networks with TED: radio propagation, mobility and protocols. ACM SIGMETRICS Performance Evaluation Review, 25(4):30–39, March 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Premore:1998:TNT

[912] Brian J. Premore and David M. Nicol. Transformation of ns TCP models to TED. ACM SIGMETRICS Performance Evaluation Review, 25(4):40–48, March 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Srinivasan:1998:FIL

[913] V. Srinivasan and George Varghese. Faster IP lookups using controlled prefix expansion. ACM SIGMETRICS Performance Evaluation Review, 26(1):

1–10, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Paxson:1998:CMP

[914] Vern Paxson. On calibrating measurements of packet transit times. ACM SIGMETRICS Performance Evaluation Review, 26(1):11–21, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:1998:MCP

[915] Randolph Y. Wang, Arvind Krishnamurthy, Richard P. Martin, Thomas E. Anderson, and David E. Culler. Modeling communication pipeline latency. ACM SIGMETRICS Performance Evaluation Review, 26(1):22–32, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Voelker:1998:ICP

[916] Geoffrey M. Voelker, Eric J. Anderson, Tracy Kimbrel, Michael J. Feeley, Jeffrey S. Chase, Anna R. Karlin, and Henry M. Levy. Implementing cooperative prefetching and caching in a globally-managed memory system. ACM SIGMETRICS Performance Evaluation Review, 26(1):33–43, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shenoy:1998:CDS

[917] Prashant J. Shenoy and Harrick M. Vin. Cello: a disk scheduling framework for next generation operating systems. ACM SIGMETRICS Performance Evaluation Review, 26(1):44–55, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rosti:1998:IPB

[918] Emilia Rosti, Giuseppe Serazzi, Evgenia Smirni, and Mark S. Squillante. The impact of I/O on program behavior and parallel scheduling. ACM SIGMETRICS Performance Evaluation Review, 26(1):56–65, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bajaj:1998:SPU

[919] Sandeep Bajaj, Lee Breslau, and Scott Shenker. Is service priority useful in networks? ACM SIGMETRICS Performance Evaluation Review, 26(1):66– 77, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kalampoukas:1998:ITT

[920] Lampros Kalampoukas, Anujan Varma, and K. K. Ramakrishnan. Improving TCP throughput over two-way asymmetric links: analysis and solutions. ACM SIGMETRICS Performance Evaluation Review, 26(1):78–89, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raman:1998:ABG

[921] Suchitra Raman, Steven McCanne, and Scott Shenker. Asymptotic behavior of global recovery in SRM. ACM SIGMETRICS Performance Evaluation Review, 26(1):90–99, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Boxma:1998:BPF

[922] O. J. Boxma and V. Dumas. The busy period in the fluid queue. ACM SIGMETRICS Performance Evaluation Review, 26(1):100–110, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:1998:TLP

[923] Guang-Liang Li, Jun-Hong Cui, Bo Li, and Fang-Ming Li. Transient loss performance of a class of finite buffer queueing systems. ACM SIGMET-RICS Performance Evaluation Review, 26(1):111–120, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

McKinnon:1998:QBA

[924] Martin W. McKinnon, George N. Rouskas, and Harry G. Perros. Queueing-based analysis of broadcast optical networks. ACM SIGMET-RICS Performance Evaluation Review, 26(1):121–130, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bavier:1998:PME

[925] Andy C. Bavier, A. Brady Montz, and Larry L. Peterson. Predicting MPEG execution times. ACM SIGMET-RICS Performance Evaluation Review, 26(1):131–140, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gribble:1998:SSF

[926] Steven D. Gribble, Gurmeet Singh Manku, Drew Roselli, Eric A. Brewer, Timothy J. Gibson, and Ethan L. Miller. Self-similarity in file systems. ACM SIGMETRICS Performance Evaluation Review, 26(1):141– 150, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Barford:1998:GRW

[927] Paul Barford and Mark Crovella. Generating representative Web workloads for network and server performance evaluation. ACM SIGMETRICS Performance Evaluation Review, 26(1): 151–160, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ji:1998:PMM

[928] Minwen Ji, Edward W. Felten, and Kai Li. Performance measurements for multithreaded programs. ACM SIGMET-RICS Performance Evaluation Review, 26(1):161–170, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jiang:1998:MES

[929] Dongming Jiang and Jaswinder Pal Singh. A methodology and an evaluation of the SGI Origin 2000. ACM SIGMETRICS Performance Evaluation Review, 26(1):171–181, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shriver:1998:ABM

[930] Elizabeth Shriver, Arif Merchant, and John Wilkes. An analytic behavior model for disk drives with readahead caches and request reordering. ACM SIGMETRICS Performance Evaluation Review, 26(1):182–191, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fraguela:1998:MSA

[931] Basilio B. Fraguela, Ramón Doallo, and Emilio L. Zapata. Modeling set associative caches behavior for irregular computations. ACM SIGMET-RICS Performance Evaluation Review, 26(1):192–201, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jiang:1998:IRF

[932] Tianji Jiang, Mostafa H. Ammar, and Ellen W. Zegura. Inter-receiver fairness: a novel performance measure for multicast ABR sessions. ACM SIGMETRICS Performance Evaluation Review, 26(1):202–211, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Courcoubetis:1998:AEL

[933] Costas Courcoubetis, Vasilios A. Siris, and George D. Stamoulis. Application and evaluation of large deviation techniques for traffic engineering in broadband networks. ACM SIGMET-RICS Performance Evaluation Review, 26(1):212–221, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Neidhardt:1998:CRT

[934] Arnold L. Neidhardt and Jonathan L. Wang. The concept of relevant time scales and its application to queuing analysis of self-similar traffic (or is Hurst naughty or nice?). ACM SIGMETRICS Performance Evaluation Review, 26(1):222–232, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arpaci-Dusseau:1998:SII

[935] Andrea C. Arpaci-Dusseau, David E. Culler, and Alan M. Mainwaring. Scheduling with implicit information in distributed systems. ACM SIGMET-RICS Performance Evaluation Review, 26(1):233–243, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nguyen:1998:SPS

[936] Thu D. Nguyen and John Zahorjan. Scheduling policies to support distributed 3D multimedia applications. ACM SIGMETRICS Performance Evaluation Review, 26(1):244– 253, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Moritz:1998:LMN

[937] Csaba Andras Moritz and Matthew I. Frank. LoGPC: modeling network contention in message-passing programs. ACM SIGMETRICS Performance Evaluation Review, 26(1):254– 263, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Barve:1998:MOT

[938] Rakesh Barve, Elizabeth Shriver, Phillip B. Gibbons, Bruce K. Hillyer, Yossi Matias, and Jeffrey Scott Vitter. Modeling and optimizing I/O throughput of multiple disks on a bus (summary). ACM SIGMETRICS Performance Evaluation Review, 26(1):264– 265, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Blumofe:1998:PWS

[939] Robert D. Blumofe and Dionisios Papadopoulos. The performance of work stealing in multiprogrammed environments (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 26(1):266–267, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Crovella:1998:TAD

[940] Mark E. Crovella, Mor Harchol-Balter, and Cristina D. Murta. Task assignment in a distributed system (extended abstract): improving performance by unbalancing load. ACM SIGMET-RICS Performance Evaluation Review, 26(1):268–269, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Manley:1998:SSS

[941] Stephen Manley, Margo Seltzer, and Michael Courage. A self-scaling and self-configuring benchmark for Web servers (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 26(1):270–271, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rousskov:1998:PCP

[942] Alex Rousskov and Valery Soloviev. On performance of caching proxies (extended abstract). ACM SIGMET-RICS Performance Evaluation Review, 26(1):272–273, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Waldby:1998:TAE

[943] J. Waldby, U. Madhow, and T. V. Lakshman. Total acknowledgements (extended abstract): a robust feedback mechanism for end-to-end congestion control. ACM SIGMETRICS Performance Evaluation Review, 26 (1):274–275, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Willis:1998:PCR

[944] Thomas E. Willis and George B. Adams III. Portable, continuous recording of complete computer behavior with low overhead (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 26(1):276– 277, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Acharya:1998:UIM

[945] Anurag Acharya and Sanjeev Setia. Using idle memory for data-intensive computations (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 26(1):278–279, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aboutabl:1998:TDD

[946] Mohamed Aboutabl, Ashok Agrawala, and Jean-Dominique Decotignie. Temporally determinate disk access (extended abstract): an experimental approach. ACM SIGMETRICS Performance Evaluation Review, 26(1):280– 281, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marsan:1998:MGS

[947] M. Ajmone Marsan, G. Balbo, G. Conte, S. Donatelli, and G. Franceschinis. Modelling with Generalized Stochastic Petri Nets. ACM SIGMET-RICS Performance Evaluation Review, 26(2):2, August 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bause:1998:SPN

[948] Falko Bause and Pieter S. Kritzinger. Stochastic Petri nets: An introduction to the theory. ACM SIGMET-RICS Performance Evaluation Review, 26(2):2–3, August 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lindemann:1998:PMD

[949] Christoph Lindemann. Performance modelling with deterministic and stochastic Petri nets. ACM SIGMET-RICS Performance Evaluation Review, 26(2):3, August 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lindemann:1998:SIS

[950] Christoph Lindemann. Special issue on stochastic Petri nets. ACM SIGMETRICS Performance Evaluation Review, 26(2):4, August 1998. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Buchholz:1998:GHG

[951] Peter Buchholz and Peter Kemper. On generating a hierarchy for GSPN analysis. ACM SIGMETRICS Performance Evaluation Review, 26(2):5–14, August 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fricks:1998:ANM

[952] Ricardo M. Fricks, Antonio Puliafito, Miklós Telek, and Kishor S.

Trivedi. Applications of non-Markovian stochastic Petri nets. ACM SIGMET-RICS Performance Evaluation Review, 26(2):15–27, August 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marsan:1998:MAS

[953] Marco Ajmone Marsan and Rossano Gaeta. Modeling ATM systems with GSPNs and SWNs. ACM SIGMET-RICS Performance Evaluation Review, 26(2):28–37, August 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ost:1998:AWM

[954] Alexander Ost and Boudewijn R. Haverkort. Analysis of windowing mechanisms with infinite-state stochastic Petri nets. ACM SIGMETRICS Performance Evaluation Review, 26 (2):38–46, August 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dujmovic:1998:EES

[955] Jozo J. Dujmović and Ivo Dujmović. Evolution and evaluation of SPEC benchmarks. ACM SIGMETRICS Performance Evaluation Review, 26(3):2– 9, December 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cao:1998:GEI

[956] Pei Cao and Sekhar Sarukkai. Guest Editors’ introduction. ACM SIGMET-RICS Performance Evaluation Review, 26(3):10, December 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Caceres:1998:WPC

[957] Ramón Cáceres, Fred Douglis, Anja Feldmann, Gideon Glass, and Michael Rabinovich. Web proxy caching: the devil is in the details. ACM SIGMET-RICS Performance Evaluation Review, 26(3):11–15, December 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krishnamurthy:1998:PQE

[958] Diwakar Krishnamurthy and Jerome Rolia. Predicting the QoS of an electronic commerce server: those mean percentiles. ACM SIGMETRICS Performance Evaluation Review, 26(3): 16–22, December 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bangs:1998:BOS

[959] Gaurav Bangs, Peter Druschel, and Jeffrey C. Mogul. Better operating system features for faster network servers. ACM SIGMETRICS Performance Evaluation Review, 26 (3):23–30, December 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mosberger:1998:HTM

[960] David Mosberger and Tai Jin. httperf — a tool for measuring Web server performance. ACM SIGMETRICS Performance Evaluation Review, 26 (3):31–37, December 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ward:1998:ISP

[961] Amy Ward, Peter Glynn, and Kathy Richardson. Internet service performance failure detection. ACM SIGMETRICS Performance Evaluation Review, 26(3):38–43, December 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sayal:1998:SAR

[962] Mehmet Sayal, Yuri Breitbart, Peter Scheuermann, and Radek Vingralek. Selection algorithms for replicated Web servers. ACM SIGMET-RICS Performance Evaluation Review, 26(3):44–50, December 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hillingsworth:1999:SSS

[963] Jeffrey K. Hillingsworth and Barton P. Miller. Summary of the SIGMET-RICS Symposium on Parallel and Distributed Processing. ACM SIGMET-RICS Performance Evaluation Review, 26(4):2–12, March 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sevcik:1999:SIS

[964] Kenneth C. Sevcik. Special issue on scheduling in multiprogrammed parallel systems. ACM SIGMETRICS Performance Evaluation Review, 26(4):13, March 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Downey:1999:EGW

[965] Allen B. Downey and Dror G. Feitelson. The elusive goal of workload characterization. ACM SIGMET-RICS Performance Evaluation Review, 26(4):14–29, March 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Setia:1999:IJM

[966] Sanjeev Setia, Mark S. Squillante, and Vijay K. Naik. The impact of job memory requirements on gang-scheduling performance. ACM SIGMETRICS Performance Evaluation Review, 26 (4):30–39, March 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chan:1999:EPJ

[967] Yuet-Ning Chan, Sivarama P. Dandamudi, and Shikharesh Majumdar. Experiences with parallel job scheduling on a transputer system. ACM SIGMETRICS Performance Evaluation Review, 26(4):40–51, March 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:1999:IJA

[968] Mark S. Squillante, David D. Yao, and Li Zhang. The impact of job arrival patterns on parallel scheduling. ACM SIGMETRICS Performance Evaluation Review, 26(4):52–59, March 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dowdy:1999:SIH

[969] L. W. Dowdy, E. Rosti, G. Serazzi, and E. Smirni. Scheduling issues in high-performance computing. ACM SIGMETRICS Performance Evaluation Review, 26(4):60–69, March 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ribeiro:1999:SNL

[970] Vinay J. Ribeiro, Rudolf H. Riedi, Matthew S. Crouse, and Richard G.

Baraniuk. Simulation of nonGaussian long-range-dependent traffic using wavelets. ACM SIGMETRICS Performance Evaluation Review, 27(1):1–12, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhao:1999:BEC

[971] Wei Zhao and Satish K. Tripathi. Bandwidth-efficient continuous media streaming through optimal multiplexing. ACM SIGMETRICS Performance Evaluation Review, 27(1):13–22, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kumar:1999:ESS

[972] Sanjeev Kumar, Dongming Jiang, Rohit Chandra, and Jaswinder Pal Singh. Evaluating synchronization on shared address space multiprocessors: methodology and performance. ACM SIGMETRICS Performance Evaluation Review, 27(1):23–34, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Acharya:1999:AUI

[973] Anurag Acharya and Sanjeev Setia. Availability and utility of idle memory in workstation clusters. ACM SIGMETRICS Performance Evaluation Review, 27(1):35–46, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kaplan:1999:TRV

[974] Scott F. Kaplan, Yannis Smaragdakis, and Paul R. Wilson. Trace reduction for virtual memory simulations. ACM SIGMETRICS Performance Evaluation Review, 27(1):47–58, June 1999.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Douceur:1999:LSS

[975] John R. Douceur and William J. Bolosky. A large-scale study of filesystem contents. ACM SIGMET-RICS Performance Evaluation Review, 27(1):59–70, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Martin:1999:NSH

[976] Richard P. Martin and David E. Culler. NFS sensitivity to high performance networks. ACM SIGMETRICS Performance Evaluation Review, 27(1):71–82, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Barve:1999:MOT

[977] Rakesh Barve, Elizabeth Shriver, Phillip B. Gibbons, Bruce K. Hillyer, Yossi Matias, and Jeffrey Scott Vitter. Modeling and optimizing I/O throughput of multiple disks on a bus. ACM SIGMETRICS Performance Evaluation Review, 27(1):83–92, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sethuraman:1999:OSS

[978] Jay Sethuraman and Mark S. Squillante. Optimal stochastic scheduling in multiclass parallel queues. ACM SIGMETRICS Performance Evaluation Review, 27(1):93–102, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Varki:1999:MVT

[979] Elizabeth Varki. Mean value technique for closed fork-join networks. ACM SIGMETRICS Performance Evaluation Review, 27(1):103–112, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Franaszek:1999:MFS

[980] Peter A. Franaszek, Philip Heidelberger, and Michael Wazlowski. On management of free space in compressed memory systems. ACM SIGMETRICS Performance Evaluation Review, 27(1):113–121, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Smaragdakis:1999:ESE

[981] Yannis Smaragdakis, Scott Kaplan, and Paul Wilson. EELRU: simple and effective adaptive page replacement. ACM SIGMETRICS Performance Evaluation Review, 27(1):122– 133, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:1999:ESP

[982] Donghee Lee, Jongmoo Choi, Jong-Hun Kim, Sam H. Noh, Sang Lyul Min, Yookun Cho, and Chong Sang Kim. On the existence of a spectrum of policies that subsumes the least recently used (LRU) and least frequently used (LFU) policies. ACM SIGMET-RICS Performance Evaluation Review, 27(1):134–143, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ludwig:1999:MLT

[983] Reiner Ludwig, Bela Rathonyi, Almudena Konrad, Kimberly Oden, and Anthony Joseph. Multi-layer tracing of TCP over a reliable wireless link. ACM SIGMETRICS Performance Evaluation Review, 27(1):144–154, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anjum:1999:BDT

[984] Farooq Anjum and Leandros Tassiulas. On the behavior of different TCP algorithms over a wireless channel with correlated packet losses. ACM SIGMET-RICS Performance Evaluation Review, 27(1):155–165, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sripanidkulchai:1999:TPV

[985] Kunwadee Sripanidkulchai, Andy Myers, and Hui Zhang. A third-party value-added network service approach to reliable multicast. ACM SIGMET-RICS Performance Evaluation Review, 27(1):166–177, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fan:1999:WPB

[986] Li Fan, Pei Cao, Wei Lin, and Quinn Jacobson. Web prefetching between low-bandwidth clients and proxies: potential and performance. ACM SIGMETRICS Performance Evaluation Review, 27(1):178–187, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Barford:1999:PEH

[987] Paul Barford and Mark Crovella. A performance evaluation of hyper text transfer protocols. ACM SIGMET-RICS Performance Evaluation Review, 27(1):188–197, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhu:1999:HRM

[988] Huican Zhu, Ben Smith, and Tao Yang. Hierarchical resource management for Web server clusters with dynamic content. ACM SIGMET-RICS Performance Evaluation Review, 27(1):198–199, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liao:1999:AGS

[989] Cheng Liao, Margaret Martonosi, and Douglas W. Clark. An adaptive globally-synchronizing clock algorithm and its implementation on a Myrinetbased PC cluster. ACM SIGMET-RICS Performance Evaluation Review, 27(1):200–201, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chou:1999:PSD

[990] ChengFu Chou, Leana Golubchik, and John C. S. Lui. A performance study of dynamic replication techniques in continuous media servers. ACM SIGMET-RICS Performance Evaluation Review, 27(1):202–203, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dovrolis:1999:RDS

[991] Constantinos Dovrolis and Dimitrios Stiliadis. Relative differentiated services in the Internet: issues and mechanisms. ACM SIGMETRICS Performance Evaluation Review, 27(1):204– 205, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bartels:1999:PLF

[992] Gretta Bartels, Anna Karlin, Darrell Anderson, Jeffrey Chase, Henry Levy, and Geoffrey Voelker. Potentials and limitations of fault-based Markov prefetching for virtual memory pages. ACM SIGMETRICS Performance Evaluation Review, 27(1):206– 207, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Crowley:1999:UTS

[993] Patrick Crowley and Jean-Loup Baer. On the use of trace sampling for architectural studies of desktop applications. ACM SIGMETRICS Performance Evaluation Review, 27(1):208– 209, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bhola:1999:WMH

[994] Sumeer Bhola and Mustaque Ahamad. Workload modeling for highly interactive applications. ACM SIGMET-RICS Performance Evaluation Review, 27(1):210–211, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Venkitaraman:1999:DEC

[995] Narayanan Venkitaraman, Tae eun Kim, and Kang-Won Lee. Design and evaluation of congestion control algorithms in the future Internet. ACM SIGMETRICS Performance Evaluation Review, 27(1):212–213, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Elnozahy:1999:ATC

[996] E. N. Elnozahy. Address trace compression through loop detection and reduction. ACM SIGMETRICS Performance Evaluation Review, 27(1):214– 215, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nahum:1999:PIW

[997] Erich Nahum, Tsipora Barzilai, and Dilip Kandlur. Performance issues in WWW servers. ACM SIGMET-RICS Performance Evaluation Review, 27(1):216–217, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ng:1999:SBE

[998] T. S. Eugene Ng, Donpaul C. Stephens, Ion Stoica, and Hui Zhang. Supporting best-effort traffic with fair service curve. ACM SIGMETRICS Performance Evaluation Review, 27(1):218– 219, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Padhye:1999:TFR

[999] Jitendra Padhye, Jim Kurose, Don Towsley, and Rajeev Koodli. A TCPfriendly rate adjustment protocol for continuous media flows over best effort networks. ACM SIGMETRICS Performance Evaluation Review, 27 (1):220–221, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Downey:1999:UPE

[1000] Allen B. Downey. Using pathchar to estimate Internet link characteristics. ACM SIGMETRICS Performance Evaluation Review, 27(1):222– 223, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hershko:1999:STS

[1001] Yuval Hershko, Daniel Segal, and Hadas Shachnai. Self-tuning synchronization mechanisms in network operating systems. ACM SIGMET-RICS Performance Evaluation Review, 27(1):224–225, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bose:1999:PEV

[1002] Pradip Bose. Performance evaluation and validation of microprocessors. ACM SIGMETRICS Performance Evaluation Review, 27(1):226– 227, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Majumdar:1999:CMC

[1003] Shikharesh Majumdar, Dale Streibel, Bruce Beninger, Brian Carroll, Neveenta Verma, and Minru Liu. Controlling memory contention on a scalable multiprocessor-based telephone switch. ACM SIGMETRICS Performance Evaluation Review, 27(1):228– 229, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cervetto:1999:MBP

[1004] Eugenio Cervetto. Model-based performance analysis of an EDP/ERPoriented wide area network. ACM SIGMETRICS Performance Evaluation Review, 27(1):230–231, June 1999.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ramanathan:1999:VSA

[1005] Srinivas Ramanathan and Edward H. Perry. The value of a systematic approach to measurement and analysis: an ISP case study. ACM SIGMET-RICS Performance Evaluation Review, 27(1):232–233, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Siebert:1999:IPD

[1006] Janet Siebert. Improving performance of data analysis in data warehouses: a methodology and case study. ACM SIGMETRICS Performance Evaluation Review, 27(1):234–235, June 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Williamson:1999:SIN

[1007] Carey Williamson. Special issue on network traffic measurements and workload characterization. ACM SIGMET-RICS Performance Evaluation Review, 27(2):2, September 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jerkins:1999:MAI

[1008] Judith L. Jerkins, John Monroe, and Jonathan L. Wang. A measurement analysis of Internet traffic over frame relay. ACM SIGMETRICS Performance Evaluation Review, 27(2):3–14, September 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Epsilon:1999:AII

[1009] Raja Epsilon, Jun Ke, and Carey Williamson. Analysis of ISP IP/ATM network traffic measurements. ACM SIGMETRICS Performance Evaluation Review, 27(2):15–24, September 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arlitt:1999:WCW

[1010] Martin Arlitt, Rich Friedrich, and Tai Jin. Workload characterization of a Web proxy in a cable modem environment. ACM SIGMETRICS Performance Evaluation Review, 27(2): 25–36, September 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Barford:1999:MWP

[1011] Paul Barford and Mark Crovella. Measuring Web performance in the wide area. ACM SIGMETRICS Performance Evaluation Review, 27(2):37–48, September 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:1999:SIW

[1012] Mark S. Squillante. Special issue on the Workshop on MAthematical performance Modeling and Analysis (MAMA ‘99). ACM SIGMETRICS Performance Evaluation Review, 27(3):2, December 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coffman:1999:IPP

[1013] E. G. Coffman, Jr., Ph. Robert, and A. L. Stolyar. The interval packing process of linear networks. ACM SIGMETRICS Performance Evaluation Review, 27(3):3–4, December 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Caceres:1999:SII

[1014] R. Cáceres, N. G. Duffield, J. Horowitz, F. Lo Presti, and D. Towsley. Statistical inference of internal network loss and topology. ACM SIGMET-RICS Performance Evaluation Review, 27(3):5–6, December 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Epema:1999:PSS

[1015] D. H. J. Epema and J. F. C. M. de Jongh. Proportional-share scheduling in single-server and multiple-server computing systems. ACM SIGMET-RICS Performance Evaluation Review, 27(3):7–10, December 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bertsimas:1999:PAM

[1016] Dimitris Bertsimas, David Gamarnik, and John N. Tsitsiklis. Performance analysis of multiclass queueing networks. ACM SIGMETRICS Performance Evaluation Review, 27(3):11–14, December 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Herzog:1999:PAG

[1017] Ulrich Herzog. Process algebras are getting mature for performance evaluation?! ACM SIGMETRICS Performance Evaluation Review, 27(3):15–18, December 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gyorfi:1999:DFC

[1018] Laszlo Gyorfi, Andras Racz, Ken Duffy, John T. Lewis, Raymond Russell, and Fergal Toomey. Distribution-free confidence intervals for measurement of effective bandwidths. ACM SIGMET-RICS Performance Evaluation Review, 27(3):19, December 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Juneja:1999:SHT

[1019] Sandeep Juneja and Perwez Shahabuddin. Simulating heavy tailed processes using delayed hazard rate twisting (extended abstract). ACM SIGMET-RICS Performance Evaluation Review, 27(3):20–22, December 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:1999:SBQ

[1020] Zhen Liu and Don Towsley. Stochastic bounds for queueing systems with multiple Markov modulated sources. ACM SIGMETRICS Performance Evaluation Review, 27(3):23, December 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:1999:WTM

[1021] Mark S. Squillante, David D. Yao, and Li Zhang. Web traffic modeling and Web server performance analysis. ACM SIGMETRICS Performance Evaluation Review, 27(3):24–27, December 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bradford:1999:ESH

[1022] Jeffrey P. Bradford and Russell Quong.

An empirical study on how program layout affects cache miss rates. ACM SIGMETRICS Performance Evaluation Review, 27(3):28–42, December 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Moore:1999:ECE

[1023] Andrew Moore and Simon Crosby. An experimental configuration for the evaluation of CAC algorithms. ACM SIGMETRICS Performance Evaluation Review, 27(3):43–54, December 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arlitt:2000:ECM

[1024] Martin Arlitt, Ludmila Cherkasova, John Dilley, Rich Friedrich, and Tai Jin. Evaluating content management techniques for Web proxy caches. ACM SIGMETRICS Performance Evaluation Review, 27(4):3–11, March 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Conti:2000:LDA

[1025] Marco Conti, Enrico Gregori, and Fabio Panzieri. Load distribution among replicated Web servers: QoSbased approach. ACM SIGMET-RICS Performance Evaluation Review, 27(4):12–19, March 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Griwodz:2000:TLP

[1026] Carsten Griwodz, Michael Liepert, Michael Zink, and Ralf Steinmetz. Tune to Lambda patching. ACM SIGMETRICS Performance Evaluation Review, 27(4):20–26, March 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menasec:2000:RMP

[1027] Daniel A. Menaséc, Rodrigo Fonseca, Virgilio A. F. Almeida, and Marco A. Mendes. Resource management policies for e-commerce servers. ACM SIGMETRICS Performance Evaluation Review, 27(4):27–35, March 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Minshall:2000:APP

[1028] Greg Minshall, Yasushi Saito, Jeffrey C. Mogul, and Ben Verghese. Application performance pitfalls and TCP’s Nagle algorithm. ACM SIGMETRICS Performance Evaluation Review, 27(4):36–44, March 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Roadknight:2000:FPC

[1029] Chris Roadknight, Ian Marshall, and Debbie Vearer. File popularity characterisation. ACM SIGMETRICS Performance Evaluation Review, 27(4):45– 50, March 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tomlinson:2000:HCI

[1030] Gary Tomlinson, Drew Major, and Ron Lee. High-capacity Internet middleware: Internet caching system architectural overview. ACM SIGMET-RICS Performance Evaluation Review, 27(4):51–56, March 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

vanderMei:2000:DSS

[1031] R. D. van der Mei, W. K. Ehrlich, P. K.

Reeser, and J. P. Francisco. A decision support system for tuning Web servers in distributed object oriented network architectures. ACM SIGMET-RICS Performance Evaluation Review, 27(4):57–62, March 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chu:2000:CES

[1032] Yang hua Chu, Sanjay G. Rao, and Hui Zhang. A case for end system multicast (keynote address). ACM SIGMETRICS Performance Evaluation Review, 28(1):1–12, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Legout:2000:PFC

[1033] A. Legout and E. W. Biersack. PLM: fast convergence for cumulative layered multicast transmission schemes. ACM SIGMETRICS Performance Evaluation Review, 28(1):13–22, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sahu:2000:ASD

[1034] Sambit Sahu, Philippe Nain, Christophe Diot, Victor Firoiu, Don Towsley, and Don Iowsley. On achievable service differentiation with token bucket marking for TCP. ACM SIGMETRICS Performance Evaluation Review, 28(1):23–33, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bolosky:2000:FSD

[1035] William J. Bolosky, John R. Douceur, David Ely, and Marvin Theimer. Feasibility of a serverless distributed file system deployed on an existing set of desktop PCs. ACM SIGMETRICS Performance Evaluation Review, 28(1):34–43, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Santos:2000:CRD

[1036] Jose Renato Santos, Richard R. Muntz, and Berthier Ribeiro-Neto. Comparing random data allocation and data striping in multimedia servers. ACM SIGMETRICS Performance Evaluation Review, 28(1):44–55, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Griffin:2000:MPM

[1037] John Linwood Griffin, Steven W.

Schlosser, Gregory R. Ganger, and David F. Nagle. Modeling and performance of MEMS-based storage devices. ACM SIGMETRICS Performance Evaluation Review, 28(1):56–65, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raunak:2000:IPC

[1038] Mohammad S. Raunak, Prashant Shenoy, Pawan Goyal, and Krithi Ramamritham. Implications of proxy caching for provisioning networks and servers. ACM SIGMETRICS Performance Evaluation Review, 28(1):66–77, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2000:CWC

[1039] Jiong Yang, Wei Wang, and Richard Muntz. Collaborative Web caching based on proxy affinities. ACM SIGMETRICS Performance Evaluation Review, 28(1):78–89, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aron:2000:CRM

[1040] Mohit Aron, Peter Druschel, and Willy Zwaenepoel. Cluster reserves: a mechanism for resource management in cluster-based network servers. ACM SIGMETRICS Performance Evaluation Review, 28(1):90–101, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Barakat:2000:APS

[1041] Chadi Barakat and Eitan Altman. Analysis of the phenomenon of several slow start phases in TCP (poster session). ACM SIGMETRICS Performance Evaluation Review, 28(1):102– 103, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wong:2000:PGQ

[1042] Wai-Man R. Wong and Richard R. Muntz. Providing guaranteed quality of service for interactive visualization applications (poster session). ACM SIGMETRICS Performance Evaluation Review, 28(1):104–105, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2000:IMF

[1043] Xin Wang, C. Yu, Henning Schulzrinne, Paul Stirpe, and Wei Wu. IP multicast fault recovery in PIM over OSPF (poster session). ACM SIGMET-RICS Performance Evaluation Review, 28(1):106–107, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lety:2000:CBM

[1044] Emmanuel Léty, Thierry Turletti, and François Baccelli. Cell-based multicast grouping in large-scale virtual environments (poster session) (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 28(1):108– 109, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jin:2000:TLW

[1045] Shudong Jin and Azer Bestavros. Temporal locality in Web request streams (poster session) (extended abstract): sources, characteristics, and caching implications. ACM SIGMETRICS Performance Evaluation Review, 28 (1):110–111, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schindler:2000:ADD

[1046] Jiri Schindler and Gregory R. Ganger.

Automated disk drive characterization (poster session). ACM SIGMET-RICS Performance Evaluation Review, 28(1):112–113, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fang:2000:OSP

[1047] Zhen Fang, Lixin Zhang, John Carter, Sally McKee, and Wilson Hsieh. Online superpage promotion revisited (poster session). ACM SIGMETRICS Performance Evaluation Review, 28(1):114– 115, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nikolaidis:2000:ILL

[1048] Ioanis Nikolaidis, Fulu Li, and Ailan Hu. An inherently loss-less and bandwidth-efficient periodic broadcast scheme for VBR video (poster session). ACM SIGMETRICS Performance Evaluation Review, 28(1):116– 117, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Koksal:2000:AST

[1049] Can Emre Koksal, Hisham Kassab, and Hari Balakrishnan. An analysis of short-term fairness in wireless media access protocols (poster session). ACM SIGMETRICS Performance Evaluation Review, 28(1):118–119, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Joshi:2000:RDH

[1050] Srinath R. Joshi and Injong Rhee. RESCU: dynamic hybrid packet-loss recovery for video transmission over the Internet (poster session). ACM SIGMETRICS Performance Evaluation Review, 28(1):120–121, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Padmanabhan:2000:CAD

[1051] Venkata N. Padmanabhan and Lili Qiu. The content and access dynamics of a busy Web server (poster session). ACM SIGMETRICS Performance Evaluation Review, 28(1):122–123, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Altman:2000:TPB

[1052] Eitan Altman, Konstantin Avrachenkov, and Chadi Barakat. TCP in presence of bursty losses. ACM SIGMET-RICS Performance Evaluation Review, 28(1):124–133, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Martin:2000:IDR

[1053] Jim Martin, Arne Nilsson, and Injong Rhee. The incremental deployability of RTT-based congestion avoidance for high speed TCP Internet connections. ACM SIGMETRICS Performance Evaluation Review, 28(1):134– 144, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rubenstein:2000:DSC

[1054] Dan Rubenstein, Jim Kurose, and Don Towsley. Detecting shared congestion of flows via end-to-end measurement. ACM SIGMETRICS Performance Evaluation Review, 28(1):145– 155, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2000:MAL

[1055] Xin Wang, Henning Schulzrinne, Dilip Kandlur, and Dinesh Verma. Measurement and analysis of LDAP performance. ACM SIGMETRICS Performance Evaluation Review, 28(1):156– 165, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cleveland:2000:IPG

[1056] William S. Cleveland, Dong Lin, and Don X. Sun. IP packet generation: statistical models for TCP start times based on connection-rate superposition. ACM SIGMETRICS Performance Evaluation Review, 28(1):166– 177, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hegde:2000:ISH

[1057] Nidhi Hegde and Khosrow Sohraby. On the impact of soft hand-off in cellular systems. ACM SIGMET-RICS Performance Evaluation Review, 28(1):178–187, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shakkottai:2000:DAP

[1058] Sanjay Shakkottai and R. Srikant. Delay asymptotics for a priority queueing system. ACM SIGMETRICS Performance Evaluation Review, 28(1):188– 195, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Golubchik:2000:FAI

[1059] Leana Golubchik and John C. S. Lui. A fast and accurate iterative solution of a multi-class threshold-based queueing system with hysteresis. ACM SIGMETRICS Performance Evaluation Review, 28(1):196–206, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Miner:2000:UES

[1060] Andrew S. Miner, Gianfranco Ciardo, and Susanna Donatelli. Using the exact state space of a Markov model to compute approximate stationary measures. ACM SIGMET-RICS Performance Evaluation Review, 28(1):207–216, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Eager:2000:ATH

[1061] Derek L. Eager, Daniel J. Sorin, and Mary K. Vernon. AMVA techniques for high service time variability. ACM SIGMETRICS Performance Evaluation Review, 28(1):217–228, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ofelt:2000:EPP

[1062] David Ofelt and John L. Hennessy. Efficient performance prediction for modern microprocessors. ACM SIGMET-RICS Performance Evaluation Review, 28(1):229–239, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Endo:2000:IIP

[1063] Yasuhiro Endo and Margo Seltzer. Improving interactive performance using TIPME. ACM SIGMETRICS Performance Evaluation Review, 28(1):240– 251, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Farkas:2000:QEC

[1064] Keith I. Farkas, Jason Flinn, Godmar Back, Dirk Grunwald, and Jennifer M. Anderson. Quantifying the energy consumption of a pocket computer and a Java Virtual Machine. ACM SIGMET-RICS Performance Evaluation Review, 28(1):252–263, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kim:2000:MSB

[1065] Jin-Soo Kim and Yarsun Hsu. Memory system behavior of Java programs: methodology and analysis. ACM SIGMETRICS Performance Evaluation Review, 28(1):264–274, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Karlsson:2000:AMW

[1066] Magnus Karlsson and Per Stenström. An analytical model of the working-set sizes in decision-support systems. ACM SIGMETRICS Performance Evaluation Review, 28(1):275–285, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Choi:2000:TAF

[1067] Jongmoo Choi, Sam H. Noh, Sang Lyul Min, and Yookun Cho. Towards application/file-level characterization of block references: a case for fine-grained buffer management. ACM SIGMET-RICS Performance Evaluation Review, 28(1):286–295, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kodialam:2000:OMR

[1068] Murali S. Kodialam, T. V. Lakshman, and Sudipta Sengupta. Online multicast routing with bandwidth guarantees: a new approach using multicast network flow. ACM SIGMET-RICS Performance Evaluation Review, 28(1):296–306, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gao:2000:SIR

[1069] Lixin Gao and Jennifer Rexford. Stable Internet routing without global coordination. ACM SIGMETRICS Performance Evaluation Review, 28(1):307– 317, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Korkmaz:2000:EAF

[1070] Turgay Korkmaz, Marwan Krunz, and Spyros Tragoudas. An efficient algorithm for finding a path subject to two additive constraints. ACM SIGMET-RICS Performance Evaluation Review, 28(1):318–327, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kant:2000:WPA

[1071] Krishna Kant. Workshop on Performance and Architecture of Web Servers (PAWS-2000). ACM SIGMETRICS Performance Evaluation Review, 28 (2):3–4, September 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kant:2000:SIS

[1072] Krishna Kant and Prasant Mohapatra. Scalable Internet servers: issues and challenges. ACM SIGMET-RICS Performance Evaluation Review, 28(2):5–8, September 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brandman:2000:CFW

[1073] Onn Brandman, Junghoo Cho, Hector Garcia-Molina, and Narayanan Shivakumar. Crawler-friendly Web servers. ACM SIGMETRICS Performance Evaluation Review, 28(2):9–14, September 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Burns:2000:CLD

[1074] Randal C. Burns, Darrell D. E. Long, and Robert M. Rees. Consistency and locking for distributing updates to Web servers using a file system. ACM SIGMETRICS Performance Evaluation Review, 28(2):15–21, September 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vasiliou:2000:PDQ

[1075] Nikolaos Vasiliou and Hanan Lutfiyya. Providing a differentiated quality of service in a World Wide Web server. ACM SIGMETRICS Performance Evaluation Review, 28(2):22–28, September 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bhattacharjee:2000:BFB

[1076] Samrat Bhattacharjee, William C. Cheng, Cheng-Fu Chou, Leana Golubchik, and Samir Khuller. Bistro: a framework for building scalable wide-area upload applications. ACM SIGMETRICS Performance Evaluation Review, 28(2):29–35, September 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kraemer:2000:MIO

[1077] E. Kraemer, G. Paixão, D. Guedes, W. Meira, Jr., and V. Almeida. Minimizing the impact of orphan requests in e-commerce services. ACM SIGMET-RICS Performance Evaluation Review, 28(2):36–42, September 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Davison:2000:PPI

[1078] Brian D. Davison and Vincenzo Liberatore. Pushing politely: improving Web responsiveness one packet at a time. ACM SIGMETRICS Performance Evaluation Review, 28(2):43, September 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arlitt:2000:CWU

[1079] Martin Arlitt. Characterizing Web user sessions. ACM SIGMETRICS Performance Evaluation Review, 28 (2):50–63, September 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hadharan:2000:EEP

[1080] R. Hadharan, W. K. Ehrlich, D. Cura, and P. K. Reeser. End to end performance modeling of Web server architectures. ACM SIGMETRICS Performance Evaluation Review, 28(2): 57–63, September 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhu:2000:AAS

[1081] Weiping Zhu. An approximate analysis of the shortest queue policy on soft real-time scheduling. ACM SIGMET-RICS Performance Evaluation Review, 28(3):3–10, December 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2000:SIP

[1082] Bo Li and Kazem Sohraby. Special issue on performance issues in mobile computing. ACM SIGMET-RICS Performance Evaluation Review, 28(3):11, December 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chang:2000:PWC

[1083] Ming Feng Chang and Yi-Bing Lin. Performance of a weakly consistent wireless Web access mechanism. ACM SIGMETRICS Performance Evaluation Review, 28(3):12–20, December 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Toh:2000:EAH

[1084] C.-K. Toh, Richard Chen, Minar Delwar, and Donald Allen. Experimenting with an Ad Hoc wireless network on campus: insights and experiences. ACM SIGMETRICS Performance Evaluation Review, 28(3):21–29, December 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lang:2000:PED

[1085] Tanja Lang and Daniel Floreani. Performance evaluation of different TCP error detection and congestion control strategies over a wireless link. ACM SIGMETRICS Performance Evaluation Review, 28(3):30–38, December 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chitre:2000:IBS

[1086] Vikrant A. Chitre and John N. Daigle. IP-based services over GPRS. ACM SIGMETRICS Performance Evaluation Review, 28(3):39–47, December 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2001:SIWa

[1087] Mark S. Squillante. Special issue on the Workshop on MAthematical performance Modeling and Analysis (MAMA 2000). ACM SIGMETRICS Performance Evaluation Review, 28(4):2, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harchol-Balter:2001:JPU

[1088] Mor Harchol-Balter. Job placement with unknown duration and no preemption. ACM SIGMETRICS Performance Evaluation Review, 28(4):3–5, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Golubchik:2001:OPT

[1089] Leana Golubchik and John C. S. Lui.

Open problems for threshold-based systems. ACM SIGMETRICS Performance Evaluation Review, 28(4):6–8, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coffman:2001:TPS

[1090] E. G. Coffman, Jr. and Predrag Jelenković. Threshold policies for singleresource reservation systems. ACM SIGMETRICS Performance Evaluation Review, 28(4):9–10, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wolf:2001:LBC

[1091] Joel L. Wolf and Philip S. Yu. Load balancing for clustered Web farms. ACM SIGMETRICS Performance Evaluation Review, 28(4):11–13, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

deSouzaeSilva:2001:TAA

[1092] Edmundo de Souza e Silva, Rosa M. M. Leão, and Morganna C. Diniz. Transient analysis applied to traffic modeling. ACM SIGMETRICS Performance Evaluation Review, 28(4):14–16, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bu:2001:FPAa

[1093] T. Bu and D. Towsley. A fixed point approximation of TCP behavior in a network. ACM SIGMETRICS Performance Evaluation Review, 28(4):17–18, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chang:2001:LDA

[1094] Cheng-Shang Chang, Yuh ming Chiu, and Wheyming Tina Song. Large deviation analysis for multiplexing independent regulated inputs. ACM SIGMET-RICS Performance Evaluation Review, 28(4):19–21, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kuang:2001:CSA

[1095] Lei Kuang and Armand M. Makowski. Convex stability and asymptotic convex ordering for non-stationary arrival processes. ACM SIGMETRICS Performance Evaluation Review, 28(4):22–23, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bachmat:2001:RRM

[1096] Eitan Bachmat. Recent results in mathematical modeling and performance evaluation of disks and disk array. ACM SIGMETRICS Performance Evaluation Review, 28(4):24–26, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hogstedt:2001:GCA

[1097] Karin Hogstedt, Doug Kimelman, V. T. Rajan, Tova Roth, and Mark Wegman. Graph cutting algorithms for distributed applications partitioning. ACM SIGMETRICS Performance Evaluation Review, 28(4):27–29, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fernandes:2001:TSL

[1098] Paulo Fernandes and Brigitte Plateau.

Triangular solution of linear systems in tensor product format. ACM SIGMETRICS Performance Evaluation Review, 28(4):30–32, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Capra:2001:UPS

[1099] L. Capra, C. Dutheillet, G. Franceschinis, and J. M. Ilié. On the use of partial symmetries for lumping Markov chains. ACM SIGMETRICS Performance Evaluation Review, 28(4):33–35, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Haas:2001:EDN

[1100] Peter J. Haas. Estimation of delays in non-regenerative discrete-event stochastic systems. ACM SIGMET-RICS Performance Evaluation Review, 28(4):36–38, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gamarnik:2001:DSC

[1101] David Gamarnik. On deciding stability of constrained random walks and queueing systems. ACM SIGMET-RICS Performance Evaluation Review, 28(4):39–40, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2001:AQU

[1102] Mark S. Squillante, Baffelly Woo, and Li Zhang. Analysis of queues under correlated arrivals with applications to Web server performance. ACM SIGMETRICS Performance Evaluation Review, 28(4):41–43, March 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Narlikar:2001:PMF

[1103] Girija Narlikar and Francis Zane. Performance modeling for fast IP lookups. ACM SIGMETRICS Performance Evaluation Review, 29(1):1–12, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Qie:2001:SCS

[1104] Xiaohu Qie, Andy Bavier, Larry Peterson, and Scott Karlin. Scheduling computations on a software-based router. ACM SIGMETRICS Performance Evaluation Review, 29(1):13–24, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Su:2001:DMP

[1105] Xun Su and Gustavo de Veciana. Dynamic multi-path routing: asymptotic approximation and simulations. ACM SIGMETRICS Performance Evaluation Review, 29(1):25–36, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jones:2001:PRS

[1106] Michael B. Jones and Stefan Saroiu.

Predictability requirements of a soft modem. ACM SIGMETRICS Performance Evaluation Review, 29(1):37–49, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lorch:2001:IDV

[1107] Jacob R. Lorch and Alan Jay Smith.

Improving dynamic voltage scaling algorithms with PACE. ACM SIGMET-RICS Performance Evaluation Review, 29(1):50–61, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vaidyanathan:2001:AIS

[1108] Kalyanaraman Vaidyanathan, Richard E.

Harper, Steven W. Hunter, and Kishor S. Trivedi. Analysis and implementation of software rejuvenation in cluster systems. ACM SIGMETRICS Performance Evaluation Review, 29(1): 62–71, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Loh:2001:TSA

[1109] Gabriel Loh. A time-stamping algorithm for efficient performance estimation of superscalar processors. ACM SIGMETRICS Performance Evaluation Review, 29(1):72–81, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2001:IFI

[1110] Thomas Bonald and Laurent Massoulié. Impact of fairness on Internet performance. ACM SIGMET-RICS Performance Evaluation Review, 29(1):82–91, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Salamatian:2001:HMM

[1111] Kavé Salamatian and Sandrine Vaton. Hidden Markov modeling for network communication channels. ACM SIGMETRICS Performance Evaluation Review, 29(1):92–101, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cao:2001:NIT

[1112] Jin Cao, William S. Cleveland, Dong Lin, and Don X. Sun. On the nonstationarity of Internet traffic. ACM SIGMETRICS Performance Evaluation Review, 29(1):102–112, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hsieh:2001:PCC

[1113] Hung-Yun Hsieh and Raghupathy Sivakumar. Performance comparison of cellular and multi-hop wireless networks: a quantitative study. ACM SIGMETRICS Performance Evaluation Review, 29(1):113–122, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hegde:2001:BLM

[1114] Nidhi Hegde and Khosrow Sohraby. Blocking in large mobile cellular networks with bursty traffic. ACM SIGMETRICS Performance Evaluation Review, 29(1):123–132, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kumar:2001:CEF

[1115] Apurva Kumar and Rajeev Gupta. Capacity evaluation of frequency hopping based ad-hoc systems. ACM SIGMET-RICS Performance Evaluation Review, 29(1):133–142, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Qiu:2001:NPF

[1116] Dongyu Qiu and Ness B. Shroff. A new predictive flow control scheme for efficient network utilization and QoS. ACM SIGMETRICS Performance Evaluation Review, 29(1):143– 153, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Paschalidis:2001:MBE

[1117] Ioannis Ch. Paschalidis and Spyridon Vassilaras. Model-based estimation of buffer overflow probabilities from measurements. ACM SIGMET-RICS Performance Evaluation Review, 29(1):154–163, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dutta:2001:OTG

[1118] Rudra Dutta and George N. Rouskas.

On optimal traffic grooming in WDM rings. ACM SIGMETRICS Performance Evaluation Review, 29(1):164– 174, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

LeBoudec:2001:SPV

[1119] Jean-Yves Le Boudec. Some properties of variable length packet shapers. ACM SIGMETRICS Performance Evaluation Review, 29(1):175–183, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chang:2001:PMI

[1120] Cheng-Shang Chang, Yuh ming Chiu, and Wheyming Tina Song. On the performance of multiplexing independent regulated inputs. ACM SIGMET-RICS Performance Evaluation Review, 29(1):184–193, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shuf:2001:CMB

[1121] Yefim Shuf, Mauricio J. Serrano, Manish Gupta, and Jaswinder Pal Singh. Characterizing the memory behavior of Java workloads: a structured view and opportunities for optimizations. ACM SIGMETRICS Performance Evaluation Review, 29(1):194–205, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sohoni:2001:SMS

[1122] Sohum Sohoni, Rui Min, Zhiyong Xu, and Yiming Hu. A study of memory system performance of multimedia applications. ACM SIGMET-RICS Performance Evaluation Review, 29(1):206–215, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bu:2001:FPAb

[1123] Tian Bu and Don Towsley. Fixed point approximations for TCP behavior in an AQM network. ACM SIGMET-RICS Performance Evaluation Review, 29(1):216–225, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Low:2001:UTV

[1124] Steven H. Low, Larry Peterson, and Limin Wang. Understanding TCP Vegas: a duality model. ACM SIGMET-RICS Performance Evaluation Review, 29(1):226–235, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Talim:2001:CRW

[1125] J. Talim, Z. Liu, Ph. Nain, and E. G.

Coffman, Jr. Controlling the robots of Web search engines. ACM SIGMET-RICS Performance Evaluation Review, 29(1):236–244, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Smith:2001:WTI

[1126] F. Donelson Smith, Félix Hernández Campos, Kevin Jeffay, and David Ott. What TCP/IP protocol headers can tell us about the Web. ACM SIGMET-RICS Performance Evaluation Review, 29(1):245–256, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nahum:2001:EWA

[1127] Erich M. Nahum, Marcel-Catalin Rosu, Srinivasan Seshan, and Jussara Almeida. The effects of wide-area conditions on WWW server performance. ACM SIGMETRICS Performance Evaluation Review, 29(1):257– 267, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nain:2001:MMQ

[1128] Philippe Nain and Redusindo Núñez-Queija. A M/M/1 queue in a semi-Markovian environment. ACM SIGMETRICS Performance Evaluation Review, 29(1):268–278, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bansal:2001:ASS

[1129] Nikhil Bansal and Mor Harchol-Balter. Analysis of SRPT scheduling: investigating unfairness. ACM SIGMET-RICS Performance Evaluation Review, 29(1):279–290, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Luthi:2001:IPC

[1130] Johannes Lüthi and Catalina M. Lladó. Interval parameters for capturing uncertainties in an EJB performance model. ACM SIGMETRICS Performance Evaluation Review, 29(1):291– 300, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

El-Sayed:2001:ASS

[1131] Hesham El-Sayed, Don Cameron, and Murray Woodside. Automation support for software performance engineering. ACM SIGMETRICS Performance Evaluation Review, 29(1):301– 311, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bradshaw:2001:PBP

[1132] Michael K. Bradshaw, Bing Wang, Subhabrata Sen, Lixin Gao, Jim Kurose, Prashant Shenoy, and Don Towsley. Periodic broadcast and patching services: implementation, measurement, and analysis in an Internet streaming video testbed. ACM SIGMETRICS Performance Evaluation Review, 29(1):312–313, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2001:TSR

[1133] Yang Richard Yang, Xiaozhou Li, Simon S. Lam, and Xincheng Zhang. Towards scalable and reliable group key management. ACM SIGMET-RICS Performance Evaluation Review, 29(1):314–315, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bremler-Barr:2001:RPC

[1134] Anat Bremler-Barr, Yehuda Afek, Haim Kaplan, Edith Cohen, and Michael Merritt. Restoration path concatenation: fast recovery of MPLS paths. ACM SIGMETRICS Performance Evaluation Review, 29(1):316– 317, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Savvides:2001:MNW

[1135] Andreas Savvides, Sung Park, and Mani B. Srivastava. On modeling networks of wireless microsensors. ACM SIGMETRICS Performance Evaluation Review, 29(1):318–319, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tsigas:2001:EPN

[1136] Philippas Tsigas and Yi Zhang. Evaluating the performance of non-blocking synchronization on shared-memory multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 29 (1):320–321, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ng:2001:OHP

[1137] Wee Teck Ng and Bruce K. Hillyer. Obtaining high performance for storage outsourcing. ACM SIGMET-RICS Performance Evaluation Review, 29(1):322–323, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Padamanabban:2001:DGL

[1138] Venkata N. Padamanabban and Lealkshminarayanan Subramanian. Determining the geographic location of Internet hosts. ACM SIGMETRICS Performance Evaluation Review, 29 (1):324–325, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mandjes:2001:LCA

[1139] Michel Mandjes, Iraj Saniee, and Alexander Stolyar. Load characterization and anomaly detection for voice over IP traffic. ACM SIGMETRICS Performance Evaluation Review, 29 (1):326–327, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Downey:2001:SCF

[1140] Allen B. Downey. The structural cause of file size distributions. ACM SIGMETRICS Performance Evaluation Review, 29(1):328–329, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bhargava:2001:UAM

[1141] Rishi Bhargava, Ashish Goel, and Adam Meyerson. Using approximate majorization to characterize protocol fairness. ACM SIGMETRICS Performance Evaluation Review, 29(1):330– 331, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mellor-Crummey:2001:PUI

[1142] John Mellor-Crummey, Robert Fowler, and David Whalley. On providing useful information for analyzing and tuning applications. ACM SIGMET-RICS Performance Evaluation Review, 29(1):332–333, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shahabi:2001:ATE

[1143] Cyrus Shahabi, Mohammad R. Kolahdouzan, Greg Barish, Roger Zimmermann, Didi Yao, Kun Fu, and Lingling Zhang. Alternative techniques for the efficient acquisition of haptic data. ACM SIGMETRICS Performance Evaluation Review, 29(1):334– 335, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dinda:2001:OPR

[1144] Peter A. Dinda. Online prediction of the running time of tasks. ACM SIGMETRICS Performance Evaluation Review, 29(1):336–337, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Almeida:2001:ARB

[1145] Virgilío Almeida, Daniel Menascé, Rudolf Riedi, Flávia Peligrinelli, Rodrigo Fonseca, and Wagner Meira, Jr. Analyzing robot behavior in e-business sites. ACM SIGMETRICS Performance Evaluation Review, 29(1):338– 339, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Almeida:2001:CUA

[1146] Jussara M. Almeida, Jeffrey Krueger, and Mary K. Vernon. Characterization of user access to streaming media files. ACM SIGMETRICS Performance Evaluation Review, 29(1):340– 341, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2001:PME

[1147] Thomas Bonald and James Roberts. Performance modeling of elastic traffic in overload. ACM SIGMETRICS Performance Evaluation Review, 29 (1):342–343, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Qiu:2001:FFI

[1148] Lili Qiu, George Varghese, and Subhash Suri. Fast firewall implementations for software-based and hardwarebased routers. ACM SIGMETRICS Performance Evaluation Review, 29 (1):344–345, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kant:2001:CRT

[1149] K. Kant and Prasant Mohapatra. Current research trends in Internet servers. ACM SIGMETRICS Performance Evaluation Review, 29(2):5–7, September 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dalal:2001:OSO

[1150] Amy Csizmar Dalal and Scott Jordan.

An optimal service ordering for a World Wide Web server. ACM SIGMET-RICS Performance Evaluation Review, 29(2):8–13, September 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cardellini:2001:WSS

[1151] Valeria Cardellini, Emiliano Casalicchio, Michele Colajanni, and Marco Mambelli. Web switch support for differentiated services. ACM SIGMET-RICS Performance Evaluation Review, 29(2):14–19, September 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Voigt:2001:KBC

[1152] Thiemo Voigt and Per Gunningberg.

Kernel-based control of persistent Web server connections. ACM SIGMET-RICS Performance Evaluation Review, 29(2):20–25, September 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2001:BPI

[1153] Jun Wang, Rui Min, Zhuying Wu, and Yiming Hu. Boosting I/O performance of Internet servers with user-level custom file systems. ACM SIGMET-RICS Performance Evaluation Review, 29(2):26–31, September 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2001:CDP

[1154] Xin Chen and Xiaodong Zhang. Coordinated data prefetching by utilizing reference information at both proxy and Web servers. ACM SIGMET-RICS Performance Evaluation Review, 29(2):32–38, September 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ardaiz:2001:IST

[1155] Oscar Ardaiz, Felix Freitag, and Leandro Navarro. Improving the service time of Web clients using server redirection. ACM SIGMETRICS Performance Evaluation Review, 29(2): 39–44, September 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jin:2001:GGI

[1156] Shudong Jin and Azer Bestavros. GISMO: a Generator of Internet Streaming Media Objects and workloads. ACM SIGMETRICS Performance Evaluation Review, 29(3):2–10, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2001:SIWb

[1157] Mark S. Squillante. Special issue on the Workshop on MAthematical performance Modeling and Analysis (MAMA 2001). ACM SIGMETRICS Performance Evaluation Review, 29(3):11, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bansal:2001:AMG

[1158] Nikhil Bansal and Mor Harchol-Balter. Analysis of M/G/1/SRPT under transient overload. ACM SIGMETRICS Performance Evaluation Review, 29 (3):12–14, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bachmat:2001:ACA

[1159] E. Bachmat. Average case analysis for batched disk scheduling and increasing subsequences. ACM SIGMET-RICS Performance Evaluation Review, 29(3):15–16, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Riabov:2001:SPT

[1160] Anton Riabov and Jay Sethuraman.

Scheduling periodic task graphs with communication delays. ACM SIG-METRICS Performance Evaluation Review, 29(3):17–18, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fourneau:2001:GNR

[1161] Jean-Michel Fourneau and Erol Gelenbe. G-networks with resets. ACM SIGMETRICS Performance Evaluation Review, 29(3):19–20, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shalmon:2001:QAP

[1162] Michael Shalmon. Queueing analysis for polling and prioritized service of aggregated regenerative variable rate ON-OFF traffic sources. ACM SIGMET-RICS Performance Evaluation Review, 29(3):20, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bain:2001:MPD

[1163] Alan Bain and Peter Key. Modelling the performance of distributed admission control for adaptive applications. ACM SIGMETRICS Performance Evaluation Review, 29(3):21–22, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chang:2001:LBB

[1164] Cheng-Shang Chang, Duan-Shin Lee, and Ching-Ming Lien. Load balanced Birkhoff–von Neumann switches with resequencing. ACM SIGMET-RICS Performance Evaluation Review, 29(3):23–24, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kogan:2001:AEP

[1165] Yaakov Kogan. Asymptotic expansions for probability distributions in large loss and closed queueing networks. ACM SIGMETRICS Performance Evaluation Review, 29(3):25–27, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baryshnikov:2001:KLM

[1166] Yuliy Baryshnikov, E. G. Coffman, Jr., and Predrag Jelenković. Kelly’s LAN model revisited. ACM SIGMET-RICS Performance Evaluation Review, 29(3):28–29, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gamarnik:2001:SOB

[1167] David Gamarnik. Stochastic online binpacking problem: exact conditions for bounded expected queue lengths under the best fit packing heuristic. ACM SIGMETRICS Performance Evaluation Review, 29(3):30–31, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lam:2001:SCS

[1168] S. Lam and Rocky K. C. Chang.

Stability comparison in single-server-multiple-queue systems. ACM SIG-METRICS Performance Evaluation Review, 29(3):32–34, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Szlavik:2001:GGT

[1169] Árpád Szlávik. GI/G/1 type processes: a non-inversive matrix analytical solution. ACM SIGMETRICS Performance Evaluation Review, 29 (3):35–37, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Boots:2001:STP

[1170] Nam Kyoo Boots and Perwez Shahabuddin. Simulating tail probabilities in GI/GI/1 queues and insurance risk processes with subexponential distributions (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 29(3):38–39, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Borst:2001:GPS

[1171] Sem Borst, Michel Mandjes, and Miranda van Uitert. Generalized processor sharing with heterogeneous traffic classes. ACM SIGMETRICS Performance Evaluation Review, 29(3):40–42, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2001:MSL

[1172] Zhen Liu, Mark S. Squillante, and Joel L. Wolf. On maximizing servicelevel-agreement profits. ACM SIG-METRICS Performance Evaluation Review, 29(3):43–44, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2001:PAA

[1173] Yingdong Lu, Jing-Sheng Song, and Weian Zheng. Performance analysis of assemble-to-order systems through strong approximations. ACM SIG-METRICS Performance Evaluation Review, 29(3):45–46, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2001:OSQ

[1174] Mark S. Squillante, Cathy H. Xia, and Li Zhang. Optimal scheduling in queueing network models of high-volume commercial Web sites. ACM SIGMET-RICS Performance Evaluation Review, 29(3):47–48, December 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sevcik:2002:SPC

[1175] Kenneth C. Sevcik and Hai Wang. Solution properties and convergence of an approximate mean value analysis algorithm. ACM SIGMETRICS Performance Evaluation Review, 29(4):3–10, March 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Williamson:2002:CCA

[1176] Carey Williamson and Qian Wu. A case for context-aware TCP/IP. ACM SIGMETRICS Performance Evaluation Review, 29(4):11–23, March 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menasce:2002:SAM

[1177] Daniel A. Menascé. Simple analytic modeling of software contention. ACM SIGMETRICS Performance Evaluation Review, 29(4):24–30, March 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cheng:2002:PSB

[1178] William C. Cheng, Cheng-Fu Chou, Leana Golubchik, and Samir Khuller. A performance study of Bistro, a scalable upload architecture. ACM SIGMETRICS Performance Evaluation Review, 29(4):31–39, March 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lawson:2002:MQB

[1179] Barry G. Lawson and Evgenia Smirni.

Multiple-queue backfilling scheduling with priorities and reservations for parallel systems. ACM SIGMET-RICS Performance Evaluation Review, 29(4):40–47, March 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pasztor:2002:PBP

[1180] Attila Pásztor and Darryl Veitch.

PC based precision timing without GPS. ACM SIGMETRICS Performance Evaluation Review, 30(1):1–10, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coates:2002:MLN

[1181] Mark Coates, Rui Castro, Robert Nowak, Manik Gadhiok, Ryan King, and Yolanda Tsang. Maximum likelihood network topology identification from edge-based unicast measurements. ACM SIGMETRICS Performance Evaluation Review, 30(1):11–20, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bu:2002:NTG

[1182] Tian Bu, Nick Duffield, Francesco Lo Presti, and Don Towsley. Network tomography on general topologies. ACM SIGMETRICS Performance Evaluation Review, 30(1):21–30, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jiang:2002:LEL

[1183] Song Jiang and Xiaodong Zhang. LIRS: an efficient low inter-reference recency set replacement policy to improve buffer cache performance. ACM SIGMETRICS Performance Evaluation Review, 30(1):31–42, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2002:MAD

[1184] Mark S. Squillante, Yanyong Zhang, Anand Sivasubramaniam, Natarajan Gautam, Hubertus Franke, and Jose Moreira. Modeling and analysis of dynamic coscheduling in parallel and distributed environments. ACM SIGMETRICS Performance Evaluation Review, 30(1):43–54, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bachmat:2002:AMS

[1185] Eitan Bachmat and Jiri Schindler.

Analysis of methods for scheduling low priority disk drive tasks. ACM SIGMETRICS Performance Evaluation Review, 30(1):55–65, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Snavely:2002:SJP

[1186] Allan Snavely, Dean M. Tullsen, and Geoff Voelker. Symbiotic jobscheduling with priorities for a simultaneous multithreading processor. ACM SIGMET-RICS Performance Evaluation Review, 30(1):66–76, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harrison:2002:PTD

[1187] Peter G. Harrison and William J. Knottenbelt. Passage time distributions in large Markov chains. ACM SIGMET-RICS Performance Evaluation Review, 30(1):77–85, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Riska:2002:EAS

[1188] Alma Riska and Evgenia Smirni. Exact aggregate solutions for M/G/1-type Markov processes. ACM SIGMET-RICS Performance Evaluation Review, 30(1):86–96, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jin:2002:SMD

[1189] Shudong Jin and Azer Bestavros. Scalability of multicast delivery for nonsequential streaming access. ACM SIGMETRICS Performance Evaluation Review, 30(1):97–107, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mauer:2002:FST

[1190] Carl J. Mauer, Mark D. Hill, and David A. Wood. Full-system timingfirst simulation. ACM SIGMETRICS Performance Evaluation Review, 30 (1):108–116, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jin:2002:PPR

[1191] Ruoming Jin and Gagan Agrawal. Performance prediction for random write reductions: a case study in modeling shared memory programs. ACM SIGMETRICS Performance Evaluation Review, 30(1):117–128, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kandiraju:2002:CTB

[1192] Gokul B. Kandiraju and Anand Sivasubramaniam. Characterizing the d-TLB behavior of SPEC CPU2000 benchmarks. ACM SIGMETRICS Performance Evaluation Review, 30 (1):129–139, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hertz:2002:EFG

[1193] Matthew Hertz, Stephen M. Blackburn, J. Eliot B. Moss, Kathryn S. McKinley, and Darko Stefanović. Error-free garbage collection traces: how to cheat and not get caught. ACM SIGMETRICS Performance Evaluation Review, 30(1):140–151, June 2002.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cameron:2002:HDM

[1194] Craig W. Cameron, Steven H. Low, and David X. Wei. High-density model for server allocation and placement. ACM SIGMETRICS Performance Evaluation Review, 30(1):152–159, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Olshefski:2002:ICR

[1195] David P. Olshefski, Jason Nieh, and Dakshi Agrawal. Inferring client response time at the Web server. ACM SIGMETRICS Performance Evaluation Review, 30(1):160–171, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2002:ACD

[1196] Sam C. M. Lee, John C. S. Lui, and David K. Y. Yau. Admission control and dynamic adaptation for a proportional-delay diffserv-enabled Web server. ACM SIGMETRICS Performance Evaluation Review, 30 (1):172–182, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2002:QSE

[1197] Haonan Tan, Derek L. Eager, Mary K.

Vernon, and Hongfei Guo. Quality of service evaluations of multicast streaming protocols. ACM SIGMET-RICS Performance Evaluation Review, 30(1):183–194, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Balachandran:2002:CUB

[1198] Anand Balachandran, Geoffrey M. Voelker, Paramvir Bahl, and P. Venkat Rangan. Characterizing user behavior and network performance in a public wireless LAN. ACM SIGMET-RICS Performance Evaluation Review, 30(1):195–205, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Singh:2002:ECT

[1199] Harkirat Singh and Suresh Singh. Energy consumption of TCP Reno, Newreno, and SACK in multi-hop wireless networks. ACM SIGMET-RICS Performance Evaluation Review, 30(1):206–216, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Heath:2002:ICA

[1200] Taliver Heath, Richard P. Martin, and Thu D. Nguyen. Improving cluster availability using workstation validation. ACM SIGMETRICS Performance Evaluation Review, 30(1):217– 227, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lai:2002:LWA

[1201] Albert Lai and Jason Nieh. Limits of wide-area thin-client computing. ACM SIGMETRICS Performance Evaluation Review, 30(1):228–239, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vetter:2002:DSP

[1202] Jeffrey Vetter. Dynamic statistical profiling of communication activity in distributed applications. ACM SIGMET-RICS Performance Evaluation Review, 30(1):240–250, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cook:2002:TRP

[1203] Jeanine Cook, Richard L. Oliver, and Eric E. Johnson. Toward reducing processor simulation time via dynamic reduction of microarchitecture complexity. ACM SIGMETRICS Performance Evaluation Review, 30(1):252– 253, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shih:2002:ETC

[1204] Jimmy S. Shih and Randy H. Katz.

Evaluating tradeoffs of congestion pricing for voice calls. ACM SIGMET-RICS Performance Evaluation Review, 30(1):254–255, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sivan-Zimet:2002:WBO

[1205] Miriam Sivan-Zimet and Tara M. Madhyastha. Workload based optimization of probe-based storage. ACM SIGMETRICS Performance Evaluation Review, 30(1):256–257, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lv:2002:SRU

[1206] Qin Lv, Pei Cao, Edith Cohen, Kai Li, and Scott Shenker. Search and replication in unstructured peer-to-peer networks. ACM SIGMETRICS Performance Evaluation Review, 30(1):258– 259, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chandramouli:2002:ALT

[1207] Y. Chandramouli and Arnold Neidhardt. Application level traffic measurements for capacity engineering. ACM SIGMETRICS Performance Evaluation Review, 30(1):260– 261, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Williamson:2002:CAT

[1208] Carey Williamson and Qian Wu. Context-aware TCP/IP. ACM SIG-METRICS Performance Evaluation Review, 30(1):262–263, June 2002. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Barakat:2002:IBT

[1209] Chadi Barakat, Patrick Thiran, Gianluca Iannaccone, and Christophe Diot. On Internet backbone traffic modeling. ACM SIGMETRICS Performance Evaluation Review, 30(1):264– 265, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thomasian:2002:SND

[1210] Alexander Thomasian and Chang Liu. Some new disk scheduling policies and their performance. ACM SIGMET-RICS Performance Evaluation Review, 30(1):266–267, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2002:SCC

[1211] Kang-Won Lee, Khalil Amiri, Sambit Sahu, and Chitra Venkatramani. On the sensitivity of cooperative caching performance to workload and network characteristics. ACM SIGMETRICS Performance Evaluation Review, 30 (1):268–269, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anantharaman:2002:MAT

[1212] Vaidyanathan Anantharaman and Raghupathy Sivakumar. A microscopic analysis of TCP performance over wireless ad-hoc networks. ACM SIGMET-RICS Performance Evaluation Review, 30(1):270–271, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Choi:2002:ARS

[1213] Baek-Young Choi, Jaesung Park, and Zhi-Li Zhang. Adaptive random sampling for load change detection. ACM SIGMETRICS Performance Evaluation Review, 30(1):272–273, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhao:2002:MEN

[1214] Zhili Zhao, Jayesh Ametha, Swaroop Darbha, and A. L. Narasimha Reddy. A method for estimating nonresponsive traffic at a router. ACM SIGMETRICS Performance Evaluation Review, 30(1):274–275, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Guo:2002:SFU

[1215] Liang Guo and Ibrahim Matta.

Scheduling flows with unknown sizes: approximate analysis. ACM SIGMET-RICS Performance Evaluation Review, 30(1):276–277, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Alouf:2002:FVC

[1216] Sara Alouf, Fabrice Huet, and Philippe Nain. Forwarders vs. centralized server: an evaluation of two approaches for locating mobile agents. ACM SIGMET-RICS Performance Evaluation Review, 30(1):278–279, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chang:2002:TCR

[1217] Hyunseok Chang, Ramesh Govindan, Sugih Jamin, Scott J. Shenker, and Walter Willinger. Towards capturing representative AS-level Internet topologies. ACM SIGMETRICS Performance Evaluation Review, 30(1):280– 281, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brownlee:2002:ISS

[1218] Nevil Brownlee and kc claffy. Internet stream size distributions. ACM SIGMETRICS Performance Evaluation Review, 30(1):282–283, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhu:2002:CLD

[1219] Yingwu Zhu and Yiming Hu. Can large disk built-in caches really improve system performance? ACM SIGMET-RICS Performance Evaluation Review, 30(1):284–285, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Simmonds:2002:WSB

[1220] Rob Simmonds, Carey Williamson, Russell Bradford, Martin Arlitt, and Brian Unger. Web server benchmarking using parallel WAN emulation. ACM SIGMETRICS Performance Evaluation Review, 30(1):286–287, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Almeida:2002:AWB

[1221] Virgilio Almeida, Martin Arlitt, and Jerry Rolia. Analyzing a Web-based system’s performance measures at multiple time scales. ACM SIGMET-RICS Performance Evaluation Review, 30(2):3–9, September 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Andreolini:2002:PSD

[1222] Mauro Andreolini, Michele Colajanni, and Ruggero Morselli. Performance study of dispatching algorithms in multi-tier Web architectures. ACM SIGMETRICS Performance Evaluation Review, 30(2):10–20, September 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2002:SND

[1223] Yan Chen, Khian Hao Lim, Randy H.

Katz, and Chris Overton. On the stability of network distance estimation. ACM SIGMETRICS Performance Evaluation Review, 30(2):21–30, September 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thomasian:2002:DSP

[1224] Alexander Thomasian and Chang Liu.

Disk scheduling policies with lookahead. ACM SIGMETRICS Performance Evaluation Review, 30(2):31–40, September 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brandwajn:2002:NSB

[1225] Alexandre Brandwajn. A note on SCSI bus waits. ACM SIGMET-RICS Performance Evaluation Review, 30(2):41–47, September 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menasce:2002:PSP

[1226] Daniel A. Menascé and Lavanya Kanchanapalli. Probabilistic scalable P2P resource location services. ACM SIGMETRICS Performance Evaluation Review, 30(2):48–58, September 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2002:SIW

[1227] Mark S. Squillante. Special issue on the Workshop on MAthematical performance Modeling and Analysis (MAMA 2002). ACM SIGMETRICS Performance Evaluation Review, 30(3):2, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yu:2002:APP

[1228] Shengke Yu, Marianne Winslett, Jonghyun Lee, and Xiaosong Ma. Automatic and portable performance modeling for parallel I/O: a machinelearning approach. ACM SIGMET-RICS Performance Evaluation Review, 30(3):3–5, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Riska:2002:EFL

[1229] Alma Riska, Vesselin Diev, and Evgenia Smirni. Efficient fitting of longtailed data sets into phase-type distributions. ACM SIGMETRICS Performance Evaluation Review, 30(3):6– 8, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harchol-Balter:2002:USL

[1230] Mor Harchol-Balter, Karl Sigman, and Adam Wierman. Understanding the slowdown of large jobs in an M/GI/1 system. ACM SIGMETRICS Performance Evaluation Review, 30(3):9–11, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Golubchik:2002:MPS

[1231] Leana Golubchik and John C. S. Lui.

Multi-path streaming: is it worth the trouble? ACM SIGMETRICS Performance Evaluation Review, 30 (3):12–14, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baryshnikov:2002:FSU

[1232] Y. Baryshnikov, E. Coffman, P. Jelenković, P. Momčilović, and D. Rubenstein. Flood search under the California split strategy. ACM SIGMET-RICS Performance Evaluation Review, 30(3):15–16, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marbukh:2002:RTE

[1233] Vladimir Marbukh. Robust traffic engineering: game theoretic perspective. ACM SIGMETRICS Performance Evaluation Review, 30(3):17–19, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Benaboud:2002:ASC

[1234] H. Benaboud, A. Berqia, and N. Mikou. An analytical study of CANIT algorithm in TCP protocol. ACM SIGMETRICS Performance Evaluation Review, 30(3):20–22, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kamal:2002:MTR

[1235] Ahmed E. Kamal. Modeling TCP Reno with RED-based routers. ACM SIGMETRICS Performance Evaluation Review, 30(3):23–25, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Samios:2002:MTT

[1236] Charalampos (Babis) Samios and Mary K. Vernon. Modeling throughput in TCP Vegas. ACM SIGMET-RICS Performance Evaluation Review, 30(3):26–28, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chandramouli:2002:MAU

[1237] Y. Chandramouli and Arnold L. Neidhardt. Mathematical analysis of the use of application-level traffic measurements for capacity engineering. ACM SIGMETRICS Performance Evaluation Review, 30(3):29–31, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xia:2002:TMP

[1238] Cathy H. Xia, Zhen Liu, Mark S. Squillante, Li Zhang, and Naceur Malouch. Traffic modeling and performance analysis of commercial Web sites. ACM SIGMETRICS Performance Evaluation Review, 30(3):32–34, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Haas:2002:VLR

[1239] Peter J. Haas and Peter W. Glynn.

On the validity of long-run estimation methods for discrete-event systems. ACM SIGMETRICS Performance Evaluation Review, 30(3):35–37, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gamarnik:2002:CSP

[1240] David Gamarnik. Computing stationary probability distributions and large deviation rates for constrained random walks: the undecidability results. ACM SIGMETRICS Performance Evaluation Review, 30(3):38–40, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harrison:2002:PFC

[1241] Peter G. Harrison. Product-forms from a CAT and DOG. ACM SIGMET-RICS Performance Evaluation Review, 30(3):41–43, December 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Allman:2003:EXR

[1242] Mark Allman. An evaluation of XML-RPC. ACM SIGMETRICS Performance Evaluation Review, 30(4):2–11, March 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Weissman:2003:GES

[1243] Jon Weissman. Guest editorial: special issue on grid computing. ACM SIGMETRICS Performance Evaluation Review, 30(4):12, March 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Taylor:2003:PIP

[1244] Valerie Taylor, Xingfu Wu, and Rick Stevens. Prophesy: an infrastructure for performance analysis and modeling of parallel and Grid applications. ACM SIGMETRICS Performance Evaluation Review, 30(4):13–18, March 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lowekamp:2003:CAP

[1245] Bruce B. Lowekamp. Combining active and passive network measurements to build scalable monitoring systems on the Grid. ACM SIGMETRICS Performance Evaluation Review, 30(4):19–26, March 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Snavely:2003:BGC

[1246] Allan Snavely, Greg Chun, Henri Casanova, Rob F. Van der Wijngaart, and Michael A. Frumkin. Benchmarks for Grid computing: a review of ongoing efforts and future directions. ACM SIGMETRICS Performance Evaluation Review, 30(4):27–32, March 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2003:GGR

[1247] Dong Lu and Peter A. Dinda. GridG:

generating realistic computational grids. ACM SIGMETRICS Performance Evaluation Review, 30(4):33–40, March 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wolski:2003:EPR

[1248] Rich Wolski. Experiences with predicting resource performance on-line in computational grid settings. ACM SIGMETRICS Performance Evaluation Review, 30(4):41–49, March 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Girbal:2003:DSR

[1249] Sylvain Girbal, Gilles Mouchard, Albert Cohen, and Olivier Temam. DiST: a simple, reliable and scalable method to significantly reduce processor architecture simulation time. ACM SIGMETRICS Performance Evaluation Review, 31(1):1–12, June 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aamodt:2003:FMO

[1250] Tor M. Aamodt, Pedro Marcuello, Paul Chow, Antonio González, Per Hammarlund, Hong Wang, and John P. Shen. A framework for modeling and optimization of prescient instruction prefetch. ACM SIGMETRICS Performance Evaluation Review, 31(1):13–24, June 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xia:2003:QSL

[1251] Cathy H. Xia and Zhen Liu. Queueing systems with long-range dependent input process and subexponential service times. ACM SIGMETRICS Performance Evaluation Review, 31(1):25–36, June 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Galmes:2003:ACM

[1252] Sebastià Galmés and Ramon Puigjaner. An algorithm for computing the mean response time of a single server queue with generalized on/off traffic arrivals. ACM SIGMETRICS Performance Evaluation Review, 31(1):37–46, June 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Garetto:2003:MSM

[1253] Michele Garetto and Don Towsley. Modeling, simulation and measurements of queuing delay under longtail Internet traffic. ACM SIGMET-RICS Performance Evaluation Review, 31(1):47–57, June 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bohacek:2003:HSM

[1254] Stephan Bohacek, João P. Hespanha, Junsoo Lee, and Katia Obraczka. A hybrid systems modeling framework for fast and accurate simulation of data communication networks. ACM SIGMETRICS Performance Evaluation Review, 31(1):58–69, June 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Samios:2003:MTT

[1255] Charalampos (Babis) Samios and Mary K. Vernon. Modeling the throughput of TCP Vegas. ACM SIGMETRICS Performance Evaluation Review, 31(1):71–81, June 2003.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2003:MAU

[1256] Jiantao Wang, Ao Tang, and Steven H.

Low. Maximum and asymptotic UDP throughput under CHOKe. ACM SIGMETRICS Performance Evaluation Review, 31(1):82–90, June 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2003:FMS

[1257] Yong Liu, Francesco Lo Presti, Vishal Misra, Don Towsley, and Yu Gu. Fluid models and solutions for largescale IP networks. ACM SIGMET-RICS Performance Evaluation Review, 31(1):91–101, June 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harrison:2003:GNP

[1258] P. G. Harrison. G-networks with propagating resets via RCAT. ACM SIGMETRICS Performance Evaluation Review, 31(2):3–5, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wierman:2003:MTV

[1259] Adam Wierman, Takayuki Osogami, and Jörgen Olsén. Modeling TCP-Vegas under on/off traffic. ACM SIGMETRICS Performance Evaluation Review, 31(2):6–8, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gamarnik:2003:WIS

[1260] David Gamarnik and John Hasenbein.

Weak instability in stochastic and fluid queueing networks. ACM SIGMET-RICS Performance Evaluation Review, 31(2):9–10, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Duarte:2003:AFA

[1261] Flávio P. Duarte, Edmundo de Souza e Silva, and Don Towsley. An adaptive FEC algorithm using hidden Markov chains. ACM SIGMETRICS Performance Evaluation Review, 31(2):11–13, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Andrew:2003:AOG

[1262] Lachlan L. H. Andrew, Yuliy Baryshnikov, E. G. Coffman, Stephen V. Hanly, and Jolyon White. An asymptotically optimal greedy algorithm for large optical burst switching systems. ACM SIGMETRICS Performance Evaluation Review, 31 (2):14–16, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marbukh:2003:TMF

[1263] Vladimir Marbukh. Towards mean field theory of wireless networks. ACM SIGMETRICS Performance Evaluation Review, 31(2):17–19, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lam:2003:PQS

[1264] Sum Lam and Rocky K. C. Chang. Per-queue stability analysis of a dynamic quota sharing scheme for wireless networks. ACM SIGMETRICS Performance Evaluation Review, 31 (2):20–22, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ma:2003:IPN

[1265] Richard T. B. Ma, C. M. Lee, John C. S. Lui, and David K. Y. Yau. Incentive P2P networks: a protocol to encourage information sharing and contribution. ACM SIGMET-RICS Performance Evaluation Review, 31(2):23–25, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bachmat:2003:PDR

[1266] Eitan Bachmat. On the performance of D-redundant storage systems. ACM SIGMETRICS Performance Evaluation Review, 31(2):26–27, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Riska:2003:ABM

[1267] Alma Riska, Evgenia Smirni, and Gianfranco Ciardo. An aggregation-based method for the exact analysis of a class of GI/G/1-type processes. ACM SIGMETRICS Performance Evaluation Review, 31(2):28–30, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:2003:HDQ

[1268] Wuqin Lin, Zhen Liu, Harry Stavropoulos, and Cathy H. Xia. Hard deadline queueing system with application to unified messaging service. ACM SIGMETRICS Performance Evaluation Review, 31(2):31–33, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bansal:2003:AST

[1269] Nikhil Bansal. On the average sojourn time under M/M/1/SRPT. ACM SIGMETRICS Performance Evaluation Review, 31(2):34–35, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Feng:2003:MSD

[1270] Hanhua Feng and Vishal Misra. Mixed scheduling disciplines for network flows. ACM SIGMETRICS Performance Evaluation Review, 31(2):36–39, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghosh:2003:RCS

[1271] Soumyadip Ghosh and Mark S. Squillante. Revisiting correlations and scheduling in Web servers. ACM SIGMETRICS Performance Evaluation Review, 31(2):40–42, September 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Titchkosky:2003:PCD

[1272] Lance Titchkosky, Martin Arlitt, and Carey Williamson. A performance comparison of dynamic Web technologies. ACM SIGMETRICS Performance Evaluation Review, 31(3):2–11, December 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Allman:2003:ELR

[1273] Mark Allman, Wesley M. Eddy, and Shawn Ostermann. Estimating loss rates with TCP. ACM SIGMET-RICS Performance Evaluation Review, 31(3):12–24, December 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Douceur:2003:RHA

[1274] John R. Douceur. Is remote host availability governed by a universal law? ACM SIGMETRICS Performance Evaluation Review, 31(3):25–29, December 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brebner:2003:JIS

[1275] Paul Brebner and Jeffrey Gosper. J2EE infrastructure scalability and throughput estimation. ACM SIGMETRICS Performance Evaluation Review, 31 (3):30–36, December 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cui:2003:NHA

[1276] Jike Cui and Mansur. H. Samadzadeh.

A new hybrid approach to exploit localities: LRFU with adaptive prefetching. ACM SIGMETRICS Performance Evaluation Review, 31(3):37–43, December 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Burger:2004:TCA

[1277] Doug Burger and Anand Sivasubramaniam. Tools for computer architecture research. ACM SIGMETRICS Performance Evaluation Review, 31(4):2–3, March 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Burger:2004:RES

[1278] Doug Burger, Todd M. Austin, and Stephen W. Keckler. Recent extensions to the SimpleScalar Tool suite. ACM SIGMETRICS Performance Evaluation Review, 31(4):4–7, March 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bohrer:2004:MFS

[1279] Patrick Bohrer, James Peterson, Mootaz Elnozahy, Ram Rajamony, Ahmed Gheith, Ron Rockhold, Charles Lefurgy, Hazim Shafi, Tarun Nakra, Rick Simpson, Evan Speight, Kartik Sudeep, Eric Van Hensbergen, and Lixin Zhang. Mambo: a full system simulator for the PowerPC architecture. ACM SIGMETRICS Performance Evaluation Review, 31(4):8–12, March 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brooks:2004:PPS

[1280] David Brooks, Pradip Bose, and Margaret Martonosi. Power-performance simulation: design and validation strategies. ACM SIGMETRICS Performance Evaluation Review, 31(4):13– 18, March 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vachharajani:2004:LSE

[1281] Manish Vachharajani, Neil Vachharajani, David A. Penry, Jason A. Blome, and David I. August. The Liberty Simulation Environment, version 1.0. ACM SIGMETRICS Performance Evaluation Review, 31(4):19–24, March 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hamerly:2004:HUS

[1282] Greg Hamerly, Erez Perelman, and Brad Calder. How to use SimPoint to pick simulation points. ACM SIGMET-RICS Performance Evaluation Review, 31(4):25–30, March 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hardavellas:2004:SFA

[1283] Nikolaos Hardavellas, Stephen Somogyi, Thomas F. Wenisch, Roland E. Wunderlich, Shelley Chen, Jangwoo Kim, Babak Falsafi, James C. Hoe, and Andreas G. Nowatzyk. Sim-Flex: a fast, accurate, flexible fullsystem simulation framework for performance evaluation of server architecture. ACM SIGMETRICS Performance Evaluation Review, 31(4):31–34, March 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mitra:2004:STE

[1284] Debasis Mitra. Stochastic traffic engineering for demand uncertainty and risk-aware network revenue management. ACM SIGMETRICS Performance Evaluation Review, 32(1):1, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marin:2004:CAP

[1285] Gabriel Marin and John Mellor-Crummey. Cross-architecture performance predictions for scientific applications using parameterized models. ACM SIGMETRICS Performance Evaluation Review, 32(1):2–13, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Huang:2004:MDS

[1286] Lan Huang, Gang Peng, and Tzi cker Chiueh. Multi-dimensional storage virtualization. ACM SIGMET-RICS Performance Evaluation Review, 32(1):14–24, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Blackburn:2004:MRP

[1287] Stephen M. Blackburn, Perry Cheng, and Kathryn S. McKinley. Myths and realities: the performance impact of garbage collection. ACM SIGMET-RICS Performance Evaluation Review, 32(1):25–36, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jin:2004:IPS

[1288] Wei Jin, Jeffrey S. Chase, and Jasleen Kaur. Interposed proportional sharing for a storage service utility. ACM SIGMETRICS Performance Evaluation Review, 32(1):37–48, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Soule:2004:FCH

[1289] Augustin Soule, Kavé Salamatia, Nina Taft, Richard Emilion, and Konstantina Papagiannaki. Flow classification by histograms: or how to go on safari in the Internet. ACM SIGMET-RICS Performance Evaluation Review, 32(1):49–60, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lakhina:2004:SAN

[1290] Anukool Lakhina, Konstantina Papagiannaki, Mark Crovella, Christophe Diot, Eric D. Kolaczyk, and Nina Taft. Structural analysis of network traffic flows. ACM SIGMETRICS Performance Evaluation Review, 32(1):61–72, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Soule:2004:HIE

[1291] Augustin Soule, Antonio Nucci, Rene Cruz, Emilio Leonardi, and Nina Taft. How to identify and estimate the largest traffic matrix elements in a dynamic environment. ACM SIGMET-RICS Performance Evaluation Review, 32(1):73–84, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Duffield:2004:FSU

[1292] Nick Duffield, Carsten Lund, and Mikkel Thorup. Flow sampling under hard resource constraints. ACM SIGMETRICS Performance Evaluation Review, 32(1):85–96, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aalto:2004:TLP

[1293] Samuli Aalto, Urtzi Ayesta, and Eeva Nyberg-Oksanen. Two-level processorsharing scheduling disciplines: mean delay analysis. ACM SIGMETRICS Performance Evaluation Review, 32(1): 97–105, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rai:2004:PAB

[1294] Idris A. Rai, Guillaume Urvoy-Keller, Mary K. Vernon, and Ernst W. Biersack. Performance analysis of LASbased scheduling disciplines in a packet switched network. ACM SIGMET-RICS Performance Evaluation Review, 32(1):106–117, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Key:2004:ELP

[1295] Peter Key, Laurent Massoulié, and Bing Wang. Emulating low-priority transport at the application layer: a background transfer service. ACM SIGMETRICS Performance Evaluation Review, 32(1):118–129, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raz:2004:RAQ

[1296] David Raz, Hanoch Levy, and Benjamin Avi-Itzhak. A resourceallocation queueing fairness measure. ACM SIGMETRICS Performance Evaluation Review, 32(1):130– 141, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Paxson:2004:MA

[1297] Vern Paxson. Measuring adversaries. ACM SIGMETRICS Performance Evaluation Review, 32(1):142, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kim:2004:FSF

[1298] Hwangnam Kim and Jennifer C. Hou. A fast simulation framework for IEEE 802.11-operated wireless LANs. ACM SIGMETRICS Performance Evaluation Review, 32(1):143–154, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hao:2004:ARF

[1299] Fang Hao, Murali Kodialam, and T. V. Lakshman. ACCEL-RATE: a faster mechanism for memory efficient perflow traffic estimation. ACM SIGMET-RICS Performance Evaluation Review, 32(1):155–166, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Burtscher:2004:VFE

[1300] Martin Burtscher. VPC3: a fast and effective trace-compression algorithm. ACM SIGMETRICS Performance Evaluation Review, 32(1):167– 176, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kumar:2004:DSA

[1301] Abhishek Kumar, Minho Sung, Jun (Jim)

Xu, and Jia Wang. Data streaming algorithms for efficient and accurate estimation of flow size distribution. ACM SIGMETRICS Performance Evaluation Review, 32(1):177–188, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ma:2004:GTA

[1302] Richard T. B. Ma, Sam C. M. Lee, John C. S. Lui, and David K. Y. Yau. A game theoretic approach to provide incentive and service differentiation in P2P networks. ACM SIGMET-RICS Performance Evaluation Review, 32(1):189–198, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lam:2004:FRS

[1303] Simon S. Lam and Huaiyu Liu. Failure recovery for structured P2P networks: protocol design and performance evaluation. ACM SIGMET-RICS Performance Evaluation Review, 32(1):199–210, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2004:ZBP

[1304] Xiaoming Wang, Yueping Zhang, Xiafeng Li, and Dmitri Loguinov. On zone-balancing of peer-to-peer networks: analysis of random node join. ACM SIGMETRICS Performance Evaluation Review, 32(1):211– 222, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kansal:2004:PAT

[1305] Aman Kansal, Dunny Potter, and Mani B. Srivastava. Performance aware tasking for environmentally powered sensor networks. ACM SIGMET-RICS Performance Evaluation Review, 32(1):223–234, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2004:PBI

[1306] Thomas Bonald and Alexandre Proutière. On performance bounds for the integration of elastic and adaptive streaming flows. ACM SIGMETRICS Performance Evaluation Review, 32(1):235– 245, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Deb:2004:RBV

[1307] Supratim Deb and R. Srikant. Ratebased versus queue-based models of congestion control. ACM SIGMET-RICS Performance Evaluation Review, 32(1):246–257, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chandrayana:2004:UCC

[1308] Kartikeya Chandrayana and Shivkumar Kalyanaraman. Uncooperative congestion control. ACM SIGMET-RICS Performance Evaluation Review, 32(1):258–269, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Applegate:2004:CNF

[1309] David Applegate, Lee Breslau, and Edith Cohen. Coping with network failures: routing strategies for optimal demand oblivious restoration. ACM SIGMETRICS Performance Evaluation Review, 32(1):270–281, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sevcik:2004:SSA

[1310] Kenneth C. Sevcik. Some systems, applications and models I have known. ACM SIGMETRICS Performance Evaluation Review, 32(1):282, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tinnakornsrisuphap:2004:CQF

[1311] Peerapol Tinnakornsrisuphap and Richard J. La. Characterization of queue fluctuations in probabilistic AQM mechanisms. ACM SIGMET-RICS Performance Evaluation Review, 32(1):283–294, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vanichpun:2004:OCU

[1312] Sarut Vanichpun and Armand M.

Makowski. The output of a cache under the independent reference model: where did the locality of reference go? ACM SIGMETRICS Performance Evaluation Review, 32(1):295– 306, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Teixeira:2004:DHP

[1313] Renata Teixeira, Aman Shaikh, Tim Griffin, and Jennifer Rexford. Dynamics of hot-potato routing in IP networks. ACM SIGMETRICS Performance Evaluation Review, 32(1):307– 319, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Agarwal:2004:IBD

[1314] Sharad Agarwal, Chen-Nee Chuah, Supratik Bhattacharyya, and Christophe Diot. The impact of BGP dynamics on intra-domain traffic. ACM SIGMET-RICS Performance Evaluation Review, 32(1):319–330, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Feamster:2004:MBR

[1315] Nick Feamster, Jared Winick, and Jennifer Rexford. A model of BGP routing for network engineering. ACM SIGMETRICS Performance Evaluation Review, 32(1):331–342, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baccelli:2004:MFA

[1316] François Baccelli, Augustin Chaintreau, Danny De Vleeschauwer, and David McDonald. A mean-field analysis of short lived interacting TCP flows. ACM SIGMETRICS Performance Evaluation Review, 32(1):343– 354, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hohn:2004:BRP

[1317] N. Hohn, D. Veitch, K. Papagiannaki, and C. Diot. Bridging router performance and queuing theory. ACM SIGMETRICS Performance Evaluation Review, 32(1):355–366, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2004:ILB

[1318] T. Bonald, M. Jonckheere, and A. Proutiére. Insensitive load balancing. ACM SIGMETRICS Performance Evaluation Review, 32(1):367– 377, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2004:WDP

[1319] T. Bonald, S. Borst, N. Hegde, and A. Proutiére. Wireless data performance in multi-cell scenarios. ACM SIGMETRICS Performance Evaluation Review, 32(1):378–380, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kapoor:2004:CSA

[1320] Rohit Kapoor, Ling-Jyh Chen, Alok Nandan, Mario Gerla, and M. Y. Sanadidi. CapProbe: a simple and accurate capacity estimation technique for wired and wireless environments. ACM SIGMETRICS Performance Evaluation Review, 32(1):390–391, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sommers:2004:HFL

[1321] Joel Sommers, Hyungsuk Kim, and Paul Barford. Harpoon: a flow-level traffic generator for router and network tests. ACM SIGMETRICS Performance Evaluation Review, 32(1):392, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ribeiro:2004:STA

[1322] Vinay J. Ribeiro, Rudolf H. Riedi, and Richard G. Baraniuk. Spatiotemporal available bandwidth estimation with STAB. ACM SIGMET-RICS Performance Evaluation Review, 32(1):394–395, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rajendran:2004:OQS

[1323] Raj Kumar Rajendran and Dan Rubenstein. Optimizing the quality of scalable video streams on P2P networks. ACM SIGMETRICS Performance Evaluation Review, 32(1):396– 397, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2004:PAT

[1324] Helen J. Wang, John Platt, Yu Chen, Ruyun Zhang, and Yi-Min Wang. Peer-Pressure for automatic troubleshooting. ACM SIGMETRICS Performance Evaluation Review, 32(1):398– 399, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hahner:2004:QAP

[1325] Jörg Hähner, Dominique Dudkowski, Pedro José Marrón, and Kurt Rothermel. A quantitative analysis of partitioning in mobile ad hoc networks. ACM SIGMETRICS Performance Evaluation Review, 32(1):400– 401, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2004:LTL

[1326] Dalu Zhang, Weili Huang, and Chen Lin. Locating the tightest link of a network path. ACM SIGMET-RICS Performance Evaluation Review, 32(1):402–403, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sullivan:2004:UPR

[1327] David G. Sullivan, Margo I. Seltzer, and Avi Pfeffer. Using probabilistic reasoning to automate software tuning. ACM SIGMETRICS Performance Evaluation Review, 32(1):404– 405, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2004:MST

[1328] Bing Wang, Jim Kurose, Prashant Shenoy, and Don Towsley. Multimedia streaming via TCP: an analytic performance study. ACM SIGMET-RICS Performance Evaluation Review, 32(1):406–407, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wynter:2004:PIQ

[1329] Laura Wynter, Cathy H. Xia, and Fan Zhang. Parameter inference of queueing models for IT systems using endto-end measurements. ACM SIGMET-RICS Performance Evaluation Review, 32(1):408–409, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pfaff:2004:PAB

[1330] Ben Pfaff. Performance analysis of BSTs in system software. ACM SIGMETRICS Performance Evaluation Review, 32(1):410–411, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2004:SDP

[1331] Mengzhi Wang, Kinman Au, Anastassia Ailamaki, Anthony Brockwell, Christos Faloutsos, and Gregory R. Ganger. Storage device performance prediction with CART models. ACM SIGMETRICS Performance Evaluation Review, 32(1):412–413, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kamra:2004:CPT

[1332] Abhinav Kamra, Vishal Misra, and Erich Nahum. Controlling the performance of 3-tiered Web sites: modeling, design and implementation. ACM SIGMETRICS Performance Evaluation Review, 32(1):414–415, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Roughan:2004:CRT

[1333] Matthew Roughan, Tim Griffin, Morley Mao, Albert Greenberg, and Brian Freeman. Combining routing and traffic data for detection of IP forwarding anomalies. ACM SIGMET-RICS Performance Evaluation Review, 32(1):416–417, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tao:2004:EPB

[1334] Shu Tao, Kuai Xu, Ying Xu, Teng Fei, Lixin Gao, Roch Guerin, Jim Kurose, Don Towsley, and Zhi-Li Zhang. Exploring the performance benefits of end-to-end path switching. ACM SIGMETRICS Performance Evaluation Review, 32(1):418–419, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kaplan:2004:CFR

[1335] Scott F. Kaplan. Complete or fast reference trace collection for simulating multiprogrammed workloads: choose one. ACM SIGMETRICS Performance Evaluation Review, 32(1):420– 421, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raghunath:2004:QTO

[1336] Satish Raghunath, Shivkumar Kalyanaraman, and K. K. Ramakrishnan. Quantifying trade-offs in resource allocation for VPNs. ACM SIGMET-RICS Performance Evaluation Review, 32(1):422–423, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ruan:2004:ONS

[1337] Yaoping Ruan and Vivek S. Pai. The origins of network server latency & the myth of connection scheduling. ACM SIGMETRICS Performance Evaluation Review, 32(1):424–425, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anagnostakis:2004:HDI

[1338] K. G. Anagnostakis and M. B. Greenwald. A hybrid direct-indirect estimator of network internal delays. ACM SIGMETRICS Performance Evaluation Review, 32(1):426–427, June 2004.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Carlsson:2004:MPS

[1339] Niklas Carlsson, Derek L. Eager, and Mary K. Vernon. Multicast protocols for scalable on-demand download. ACM SIGMETRICS Performance Evaluation Review, 32(1):428– 429, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pai:2004:IPI

[1340] Vijay S. Pai, Scott Rixner, and Hyong youb Kim. Isolating the performance impacts of network interface cards through microbenchmarks. ACM SIGMETRICS Performance Evaluation Review, 32(1):430–431, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chu:2004:ECU

[1341] Jacky Chu, Kevin Labonte, and Brian Neil Levine. An evaluation of Chord using traces of peer-to-peer file sharing. ACM SIGMETRICS Performance Evaluation Review, 32(1):432– 433, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2004:GEF

[1342] Mark S. Squillante. Guest Editor’s foreword. ACM SIGMETRICS Performance Evaluation Review, 32(2):2, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Osogami:2004:RAT

[1343] Takayuki Osogami, Adam Wierman, Mor Harchol-Balter, and Alan Scheller-Wolf. A recursive analysis technique for multi-dimensionally infinite Markov chains. ACM SIGMETRICS Performance Evaluation Review, 32(2):3–5, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

daSilva:2004:EAT

[1344] Ana Paula Couto da Silva, Rosa M. M.

Leäo, and Edmundo de Souza e Silva. An efficient approximate technique for solving fluid models. ACM SIGMET-RICS Performance Evaluation Review, 32(2):6–8, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kogan:2004:TPI

[1345] Yaakov Kogan and Gagan Choudhury.

Two problems in Internet reliability: new questions for old models. ACM SIGMETRICS Performance Evaluation Review, 32(2):9–11, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wierman:2004:FSS

[1346] Adam Wierman and Mor Harchol-Balter. Formalizing SMART scheduling. ACM SIGMETRICS Performance Evaluation Review, 32(2):12–13, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raz:2004:HFQ

[1347] David Raz, Benjamin Avi-Itzhak, and Hanoch Levy. How fair is queue prioritization? ACM SIGMETRICS Performance Evaluation Review, 32 (2):14–16, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Feng:2004:RBC

[1348] Hanhua Feng and Vishal Misra. On the relationship between coefficient of variation and the performance of M/G/1-FB queues. ACM SIGMETRICS Performance Evaluation Review, 32(2): 17–19, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chang:2004:DSM

[1349] Junxia Chang, Hayriye Ayhan, and Jim Dai. Dynamic scheduling of multiclass open queueing networks in a slowly changing environment. ACM SIGMETRICS Performance Evaluation Review, 32(2):20–21, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marbukh:2004:KPP

[1350] Vladimir Marbukh. A knowledge plane as a pricing mechanism for aggregate, user-centric utility maximization. ACM SIGMETRICS Performance Evaluation Review, 32(2):22–24, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:2004:CMM

[1351] Wuqin Lin, Zhen Liu, Cathy H. Xia, and Li Zhang. Cost minimization of multi-tiered e-business infrastructure with end-to-end delay guarantees. ACM SIGMETRICS Performance Evaluation Review, 32(2):25–27, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Adler:2004:TOP

[1352] Micah Adler, Rakesh Kumar, Keith Ross, Dan Rubenstein, David Turner, and David D. Yao. Two optimal peer selection problems. ACM SIGMET-RICS Performance Evaluation Review, 32(2):28–30, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coffman:2004:CDS

[1353] E. G. Coffman, Andreas Constantinides, Dan Rubenstein, Bruce Shepherd, and Angelos Stavrou. Content distribution for seamless transmission. ACM SIGMETRICS Performance Evaluation Review, 32(2): 31–32, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gamarnik:2004:AOT

[1354] David Gamarnik and Petar Momčilović.

An asymptotic optimality of the transposition rule for linear lists. ACM SIGMETRICS Performance Evaluation Review, 32(2):33–34, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baryshnikov:2004:SAT

[1355] Yuliy Baryshnikov, Ed Coffman, and Petar Momčilović. Self assembly times in DNA-based computation. ACM SIGMETRICS Performance Evaluation Review, 32(2):35–37, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Saniee:2004:PDS

[1356] Iraj Saniee, Indra Widjaja, and John Morrison. Performance of a distributed scheduling protocol for TWIN. ACM SIGMETRICS Performance Evaluation Review, 32(2):38–40, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bekker:2004:ITF

[1357] René Bekker, Sem Borst, and Rudesindo Néñez-Queija. Integration of TCP-friendly streaming sessions and heavy-tailed elastic flows. ACM SIG-METRICS Performance Evaluation Review, 32(2):41–43, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

vanKessel:2004:ARA

[1358] Gijs van Kessel, Rudesindo Núñez-Queija, and Sem Borst. Asymptotic regimes and approximations for discriminatory processor sharing. ACM SIGMETRICS Performance Evaluation Review, 32(2):44–46, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cui:2004:ODM

[1359] Yi Cui, Yuan Xue, and Klara Nahrstedt. Optimal distributed multicast routing using network coding: theory and applications. ACM SIGMET-RICS Performance Evaluation Review, 32(2):47–49, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2004:CPS

[1360] Xuan Li and David D. Yao. Control and pricing in stochastic networks with concurrent resource occupancy. ACM SIGMETRICS Performance Evaluation Review, 32(2):50–52, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Guo:2004:OPR

[1361] Xin Guo, Yingdong Lu, and Mark S.

Squillante. Optimal probabilistic routing in distributed parallel queues. ACM SIGMETRICS Performance Evaluation Review, 32(2):53–54, September 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Neto:2004:CBU

[1362] Humberto T. Marques Neto, Jussara M. Almeida, Leonardo C. D. Rocha, Wagner Meira, Pedro H. C. Guerra, and Virgilio A. F. Almeida. A characterization of broadband user behavior and their e-business activities. ACM SIGMETRICS Performance Evaluation Review, 32(3):3–13, December 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Andreolini:2004:FGP

[1363] Mauro Andreolini, Michele Colajanni, Riccardo Lancellotti, and Francesca Mazzoni. Fine grain performance evaluation of e-commerce sites. ACM SIGMETRICS Performance Evaluation Review, 32(3):14–23, December 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sopitkamol:2004:RCP

[1364] Monchai Sopitkamol. Ranking configuration parameters in multi-tiered e-commerce sites. ACM SIGMET-RICS Performance Evaluation Review, 32(3):24–33, December 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

DAntonio:2004:ASC

[1365] S. D’Antonio, M. Esposito, S. P. Romano, and G. Ventre. Assessing the scalability of component-based frameworks: the CADENUS case study. ACM SIGMETRICS Performance Evaluation Review, 32(3):34–43, December 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ye:2004:RRS

[1366] Tao Ye and Shivkumar Kalyanaraman. A recursive random search algorithm for network parameter optimization. ACM SIGMETRICS Performance Evaluation Review, 32(3):44–53, December 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Haverkort:2005:PV

[1367] Boudewijn R. Haverkort and Joost-Pieter Katoen. Performance and verification. ACM SIGMETRICS Performance Evaluation Review, 32(4):3, March 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ciardo:2005:IDS

[1368] Gianfranco Ciardo and Andrew S. Miner. Implicit data structures for logic and stochastic systems analysis. ACM SIGMETRICS Performance Evaluation Review, 32(4):4–9, March 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baier:2005:MCM

[1369] Christel Baier, Boudewijn R. Haverkort, Holger Hermanns, and Joost-Pieter Katoen. Model checking meets performance evaluation. ACM SIGMET-RICS Performance Evaluation Review, 32(4):10–15, March 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kwiatkowska:2005:PMC

[1370] Marta Kwiatkowska, Gethin Norman, and David Parker. Probabilistic model checking in practice: case studies with PRISM. ACM SIGMETRICS Performance Evaluation Review, 32(4):16–21, March 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baier:2005:PVM

[1371] Christel Baier, Frank Ciesinski, and Marcus Größer. ProbMela and verification of Markov decision processes. ACM SIGMETRICS Performance Evaluation Review, 32(4):22–27, March 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jansen:2005:QMA

[1372] David N. Jansen and Holger Hermanns. QoS modelling and analysis with UML statecharts: the StoCharts approach. ACM SIGMETRICS Performance Evaluation Review, 32(4):28–33, March 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Behrmann:2005:OSU

[1373] Gerd Behrmann, Kim G. Larsen, and Jacob I. Rasmussen. Optimal scheduling using priced timed automata. ACM SIGMETRICS Performance Evaluation Review, 32(4):34–40, March 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

McIver:2005:ARP

[1374] Annabelle McIver and Carroll Morgan. Abstraction and refinement in probabilistic systems. ACM SIGMET-RICS Performance Evaluation Review, 32(4):41–47, March 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hoelzle:2005:GHL

[1375] Urs Hoelzle. Google: or how I learned to love terabytes. ACM SIGMETRICS Performance Evaluation Review, 33(1): 1, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Massoulie:2005:CRS

[1376] Laurent Massoulié and Milan VojnoviĆ. Coupon replication systems. ACM SIGMETRICS Performance Evaluation Review, 33(1):2–13, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tang:2005:LTO

[1377] Chunqiang Tang, Melissa J. Buco, Rong N. Chang, Sandhya Dwarkadas, Laura Z. Luan, Edward So, and Christopher Ward. Low traffic overlay networks with large routing tables. ACM SIGMETRICS Performance Evaluation Review, 33(1):14–25, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Leonard:2005:LBN

[1378] Derek Leonard, Vivek Rai, and Dmitri Loguinov. On lifetime-based node failure and stochastic resilience of decentralized peer-to-peer networks. ACM SIGMETRICS Performance Evaluation Review, 33(1):26–37, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dumitriu:2005:DSR

[1379] D. Dumitriu, E. Knightly, A. Kuzmanovic, I. Stoica, and W. Zwaenepoel. Denial-of-service resilience in peer-to-peer file sharing systems. ACM SIGMETRICS Performance Evaluation Review, 33(1):38–49, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Moore:2005:ITC

[1380] Andrew W. Moore and Denis Zuev.

Internet traffic classification using Bayesian analysis techniques. ACM SIGMETRICS Performance Evaluation Review, 33(1):50–60, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kumar:2005:DSA

[1381] Abhishek Kumar, Minho Sung, Jun (Jim)

Xu, and Ellen W. Zegura. A data streaming algorithm for estimating subpopulation flow size distribution. ACM SIGMETRICS Performance Evaluation Review, 33(1):61–72, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cohen:2005:PCL

[1382] Edith Cohen and Carsten Lund. Packet classification in large ISPs: design and evaluation of decision tree classifiers. ACM SIGMETRICS Performance Evaluation Review, 33(1):73–84, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Keys:2005:RSA

[1383] Ken Keys, David Moore, and Cristian Estan. A robust system for accurate real-time summaries of Internet traffic. ACM SIGMETRICS Performance Evaluation Review, 33(1):85–96, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Choi:2005:PCW

[1384] Sunwoong Choi, Kihong Park, and Chong kwon Kim. On the performance characteristics of WLANs: revisited. ACM SIGMETRICS Performance Evaluation Review, 33(1):97– 108, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ramaiyan:2005:FPA

[1385] Venkatesh Ramaiyan, Anurag Kumar, and Eitan Altman. Fixed point analysis of single cell IEEE 802.11e WLANs: uniqueness, multistability and throughput differentiation. ACM SIGMET-RICS Performance Evaluation Review, 33(1):109–120, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lindemann:2005:MEI

[1386] Christoph Lindemann and Oliver P. Waldhorst. Modeling epidemic information dissemination on mobile devices with finite buffers. ACM SIGMET-RICS Performance Evaluation Review, 33(1):121–132, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kumar:2005:AAC

[1387] V. S. Anil Kumar, Madhav V.

Marathe, Srinivasan Parthasarathy, and Aravind Srinivasan. Algorithmic aspects of capacity in wireless networks. ACM SIGMETRICS Performance Evaluation Review, 33(1):133– 144, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2005:EEM

[1388] Zhifeng Chen, Yan Zhang, Yuanyuan Zhou, Heidi Scott, and Berni Schiefer. Empirical evaluation of multi-level buffer cache collaboration for storage systems. ACM SIGMETRICS Performance Evaluation Review, 33(1):145– 156, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Butt:2005:PIK

[1389] Ali R. Butt, Chris Gniady, and Y. Charlie Hu. The performance impact of kernel prefetching on buffer cache replacement algorithms. ACM SIGMETRICS Performance Evaluation Review, 33(1):157–168, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berg:2005:FDL

[1390] Erik Berg and Erik Hagersten. Fast data-locality profiling of native execution. ACM SIGMETRICS Performance Evaluation Review, 33(1):169– 180, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yotov:2005:AMM

[1391] Kamen Yotov, Keshav Pingali, and Paul Stodghill. Automatic measurement of memory hierarchy parameters. ACM SIGMETRICS Performance Evaluation Review, 33(1):181– 192, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jonckheere:2005:OIR

[1392] M. Jonckheere and J. Virtamo. Optimal insensitive routing and bandwidth sharing in simple data networks. ACM SIGMETRICS Performance Evaluation Review, 33(1):193–204, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wierman:2005:NIB

[1393] Adam Wierman, Mor Harchol-Balter, and Takayuki Osogami. Nearly insensitive bounds on SMART scheduling. ACM SIGMETRICS Performance Evaluation Review, 33(1):205– 216, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kortebi:2005:ENA

[1394] A. Kortebi, L. Muscariello, S. Oueslati, and J. Roberts. Evaluating the number of active flows in a scheduler realizing fair statistical bandwidth sharing. ACM SIGMETRICS Performance Evaluation Review, 33(1):217– 228, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wierman:2005:CSP

[1395] Adam Wierman and Mor Harchol-Balter. Classifying scheduling policies with respect to higher moments of conditional response time. ACM SIGMET-RICS Performance Evaluation Review, 33(1):229–240, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jiang:2005:WIT

[1396] Hao Jiang and Constantinos Dovrolis.

Why is the Internet traffic bursty in short time scales? ACM SIGMET-RICS Performance Evaluation Review, 33(1):241–252, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Roughan:2005:FBA

[1397] Matthew Roughan. Fundamental bounds on the accuracy of network performance measurements. ACM SIGMETRICS Performance Evaluation Review, 33(1):253–264, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jain:2005:EEE

[1398] Manish Jain and Constantinos Dovrolis. End-to-end estimation of the available bandwidth variation range. ACM SIGMETRICS Performance Evaluation Review, 33(1):265–276, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chiang:2005:NUM

[1399] Mung Chiang, J. W. Lee, R. Calderbank, D. Palomar, and M. Fazel. Network utility maximization with nonconcave, coupled, and reliability-based utilities. ACM SIGMETRICS Performance Evaluation Review, 33(1):277, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chiang:2005:OCC

[1400] Mung Chiang and Steven Low. Optimization and control of communication networks. ACM SIGMETRICS Performance Evaluation Review, 33(1):277, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Low:2005:OMI

[1401] Steven Low, J. Doyle, L. Li, A. Tang, and J. Wang. Optimization model of Internet protocols. ACM SIGMET-RICS Performance Evaluation Review, 33(1):277, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mitra:2005:JPN

[1402] Debasis Mitra. Joint pricing-network design and stochastic traffic engineering to manage demand uncertainty. ACM SIGMETRICS Performance Evaluation Review, 33(1):278, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Musacchio:2005:AFR

[1403] John Musacchio and Jean Walrand. Achieving fair rates with ingress policing. ACM SIGMETRICS Performance Evaluation Review, 33(1):278, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shroff:2005:OBA

[1404] Ness Shroff and Xiaojun Lin. An optimization based approach for cross-layer design in wireless communication networks. ACM SIGMETRICS Performance Evaluation Review, 33(1):278, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ciucu:2005:NSC

[1405] Florin Ciucu, Almut Burchard, and Jörg Liebeherr. A network service curve approach for the stochastic analysis of networks. ACM SIGMET-RICS Performance Evaluation Review, 33(1):279–290, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Urgaonkar:2005:AMM

[1406] Bhuvan Urgaonkar, Giovanni Pacifici, Prashant Shenoy, Mike Spreitzer, and Asser Tantawi. An analytical model for multi-tier Internet services and its applications. ACM SIGMET-RICS Performance Evaluation Review, 33(1):291–302, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2005:MSE

[1407] Yiyu Chen, Amitayu Das, Wubi Qin, Anand Sivasubramaniam, Qian Wang, and Natarajan Gautam. Managing server energy and operational costs in hosting centers. ACM SIGMET-RICS Performance Evaluation Review, 33(1):303–314, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ruan:2005:EIS

[1408] Yaoping Ruan, Vivek S. Pai, Erich Nahum, and John M. Tracey. Evaluating the impact of simultaneous multithreading on network servers using real hardware. ACM SIGMET-RICS Performance Evaluation Review, 33(1):315–326, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Donnet:2005:EAL

[1409] Benoit Donnet, Philippe Raoult, Timur Friedman, and Mark Crovella. Efficient algorithms for large-scale topology discovery. ACM SIGMETRICS Performance Evaluation Review, 33 (1):327–338, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mao:2005:LPI

[1410] Z. Morley Mao, Lili Qiu, Jia Wang, and Yin Zhang. On AS-level path inference. ACM SIGMETRICS Performance Evaluation Review, 33(1):339– 349, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhao:2005:DSA

[1411] Qi (George) Zhao, Abhishek Kumar, Jia Wang, and Jun (Jim) Xu. Data streaming algorithms for accurate and efficient measurement of traffic and flow matrices. ACM SIGMETRICS Performance Evaluation Review, 33 (1):350–361, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Soule:2005:TMB

[1412] Augustin Soule, Anukool Lakhina, Nina Taft, Konstantina Papagiannaki, Kave Salamatian, Antonio Nucci, Mark Crovella, and Christophe Diot. Traffic matrices: balancing measurements, inference and modeling. ACM SIGMET-RICS Performance Evaluation Review, 33(1):362–373, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ganeriwal:2005:RAT

[1413] Saurabh Ganeriwal, Deepak Ganesanl, Mark Hansen, Mani B. Srivastava, and Deborah Estrin. Rate-adaptive time synchronization for long-lived sensor networks. ACM SIGMETRICS Performance Evaluation Review, 33(1):374– 375, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2005:IPS

[1414] An-I A. Wang, Peter Reiher, and Geoff Kuenning. Introducing permuted states for analyzing conflict rates in optimistic replication. ACM SIGMET-RICS Performance Evaluation Review, 33(1):376–377, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mickens:2005:PNA

[1415] James W. Mickens and Brian D. Noble. Predicting node availability in peer-to-peer networks. ACM SIGMET-RICS Performance Evaluation Review, 33(1):378–379, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Qiu:2005:TMW

[1416] Lili Qiu, Paramvir Bahl, Ananth Rao, and Lidong Zhou. Troubleshooting multihop wireless networks. ACM SIGMETRICS Performance Evaluation Review, 33(1):380–381, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raz:2005:FOM

[1417] David Raz, Benjamin Avi-Itzhak, and Hanoch Levy. Fair operation of multiserver and multi-queue systems. ACM SIGMETRICS Performance Evaluation Review, 33(1):382–383, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anderson:2005:DSA

[1418] Eric Anderson, Dirk Beyer, Kamalika Chaudhuri, Terence Kelly, Norman Salazar, Cipriano Santos, Ram Swaminathan, Robert Tarjan, Janet Wiener, and Yunhong Zhou. Deadline scheduling for animation rendering. ACM SIGMETRICS Performance Evaluation Review, 33(1):384–385, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

He:2005:SSP

[1419] Simin He, Shutao Sun, Wei Zhao, Yanfeng Zheng, and Wen Gao. Smooth switching problem in buffered crossbar switches. ACM SIGMETRICS Performance Evaluation Review, 33 (1):386–387, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

He:2005:PTT

[1420] Qi He, Constantinos Dovrolis, and Mostafa Ammar. Prediction of TCP throughput: formula-based and history-based methods. ACM SIG-METRICS Performance Evaluation Review, 33(1):388–389, June 2005. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chua:2005:SFE

[1421] David Chua, Eric D. Kolaczyk, and Mark Crovella. A statistical framework for efficient monitoring of end-to-end network properties. ACM SIGMET-RICS Performance Evaluation Review, 33(1):390–391, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhu:2005:TSA

[1422] Ningning Zhu, Jiawu Chen, Tzi cker Chiueh, and Daniel Ellard. TBBT: scalable and accurate trace replay for file server evaluation. ACM SIGMET-RICS Performance Evaluation Review, 33(1):392–393, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sarat:2005:UAD

[1423] Sandeep Sarat, Vasileios Pappas, and Andreas Terzis. On the use of anycast in DNS. ACM SIGMETRICS Performance Evaluation Review, 33 (1):394–395, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mudigonda:2005:MMA

[1424] Jayaram Mudigonda, Harrick M. Vin, and Raj Yavatkar. Managing memory access latency in packet processing. ACM SIGMETRICS Performance Evaluation Review, 33(1):396– 397, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bharambe:2005:SOB

[1425] Ashwin R. Bharambe, Cormac Herley, and Venkata N. Padmanabhan. Some observations on BitTorrent performance. ACM SIGMETRICS Performance Evaluation Review, 33(1):398– 399, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Machiraju:2005:TPC

[1426] Sridhar Machiraju, Darryl Veitch, François Baccelli, Antonio Nucci, and Jean C. Bolot. Theory and practice of cross-traffic estimation. ACM SIGMETRICS Performance Evaluation Review, 33(1):400–401, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stutzbach:2005:CTT

[1427] Daniel Stutzbach and Reza Rejaie. Characterizing the two-tier Gnutella topology. ACM SIGMETRICS Performance Evaluation Review, 33(1):402– 403, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tewari:2005:ASR

[1428] Saurabh Tewari and Leonard Kleinrock. Analysis of search and replication in unstructured peer-to-peer networks. ACM SIGMETRICS Performance Evaluation Review, 33(1):404– 405, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2005:ILS

[1429] Jianyong Zhang, Anand Sivasubramaniam, Alma Riska, Qian Wang, and Erik Riedel. An interposed 2-level I/O scheduling framework for performance virtualization. ACM SIGMET-RICS Performance Evaluation Review, 33(1):406–407, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wenisch:2005:TAM

[1430] Thomas F. Wenisch, Roland E. Wunderlich, Babak Falsafi, and James C.

Hoe. TurboSMARTS: accurate microarchitecture simulation sampling in minutes. ACM SIGMETRICS Performance Evaluation Review, 33(1):408– 409, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hu:2005:RCM

[1431] Chunyu Hu and Jennifer C. Hou. A reactive channel model for expediting wireless network simulation. ACM SIGMETRICS Performance Evaluation Review, 33(1):410–411, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Groenevelt:2005:MDM

[1432] Robin Groenevelt, Philippe Nain, and Ger Koole. Message delay in MANET. ACM SIGMETRICS Performance Evaluation Review, 33(1):412– 413, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2005:SIW

[1433] Mark S. Squillante. Special issue on the workshop on MAthematical performance Modeling And Analysis (MAMA 2005): Guest Editor’s foreword. ACM SIGMETRICS Performance Evaluation Review, 33(2):2, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Carofiglio:2005:SPA

[1434] Giovanna Carofiglio, Rossano Gaeta, Michele Garetto, Paolo Giaccone, Emilio Leonardi, and Matteo Sereno. A statistical physics approach for modelling P2P systems. ACM SIGMET-RICS Performance Evaluation Review, 33(2):3–5, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sundararaj:2005:OPA

[1435] Ananth I. Sundararaj, Manan Sanghi, John R. Lange, and Peter A. Dinda. An optimization problem in adaptive virtual environments. ACM SIGMET-RICS Performance Evaluation Review, 33(2):6–8, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nicol:2005:OPC

[1436] David M. Nicol. Optimized pre-copy calibration of hard drives. ACM SIGMETRICS Performance Evaluation Review, 33(2):9–11, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kumaran:2005:SAC

[1437] J. Kumaran, K. Mitchell, and A. van de Liefvoort. A spectral approach to compute performance measures in a correlated single server queue. ACM SIGMETRICS Performance Evaluation Review, 33(2):12–14, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fiorini:2005:UCS

[1438] Pierre M. Fiorini, Robert Sheahan, and Lester Lipsky. On unreliable computing systems when heavy-tails appear as a result of the recovery procedure. ACM SIGMETRICS Performance Evaluation Review, 33(2): 15–17, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2005:MDP

[1439] Qi Zhang, Armin Heindl, and Evgenia Smirni. Models of the departure process of a BMAP/MAP/1 queue. ACM SIGMETRICS Performance Evaluation Review, 33(2):18–20, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ramachandran:2005:PBA

[1440] Krishna K. Ramachandran and Biplab Sikdar. A population based approach to model network lifetime in wireless sensor networks. ACM SIGMET-RICS Performance Evaluation Review, 33(2):21–23, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kamra:2005:DPS

[1441] Abhinav Kamra, Jon Feldman, Vishal Misra, and Dan Rubenstein. Data persistence in sensor networks: towards optimal encoding for data recovery in partial network failures. ACM SIGMETRICS Performance Evaluation Review, 33(2):24–26, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jiang:2005:ION

[1442] Wenjie Jiang, John C. S. Lui, and Dah-Ming Chiu. Interaction of overlay networks: properties and implications. ACM SIGMETRICS Performance Evaluation Review, 33(2): 27–29, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ma:2005:CNC

[1443] Richard T. B. Ma, Vishal Misra, and Dan Rubenstein. Cooperative and noncooperative models for slotted-Aloha type MAC protocols. ACM SIGMET-RICS Performance Evaluation Review, 33(2):30–32, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Covell:2005:PMS

[1444] Michele Covell, Sumit Roy, and Beomjoo Seo. Predictive modeling of streaming servers. ACM SIGMET-RICS Performance Evaluation Review, 33(2):33–35, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harchol-Balter:2005:RTP

[1445] Mor Harchol-Balter, Takayuki Osogami, and Alan Scheller-Wolf. Robustness of threshold policies in beneficiary-donor model. ACM SIG-METRICS Performance Evaluation Review, 33(2):36–38, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raz:2005:LRU

[1446] David Raz, Benjamin Avi-Itzhak, and Hanoch Levy. Locality of reference and the use of sojourn time variance for measuring queue unfairness: extended abstract. ACM SIGMET-RICS Performance Evaluation Review, 33(2):39–41, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2005:DSO

[1447] Yingdong Lu and Mark S. Squillante. Dynamic scheduling to optimize utility functions of sojourn time moments in queueing systems. ACM SIGMET-RICS Performance Evaluation Review, 33(2):42–44, September 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Papagiannaki:2005:GEF

[1448] Konstantina Papagiannaki and Yin Zhang. Guest Editor’s foreword. ACM SIGMETRICS Performance Evaluation Review, 33(3):2, December 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chandramouli:2005:ANC

[1449] Y. Chandramouli and Arnold Neidhardt. Analysis of network congestion inference techniques. ACM SIGMET-RICS Performance Evaluation Review, 33(3):3–9, December 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Burch:2005:MLD

[1450] Hal Burch and Chris Chase. Monitoring link delays with one measurement host. ACM SIGMETRICS Performance Evaluation Review, 33 (3):10–17, December 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Choi:2005:OCS

[1451] Baek-Young Choi and Supratik Bhattacharyya. Observations on Cisco sampled NetFlow. ACM SIGMETRICS Performance Evaluation Review, 33 (3):18–23, December 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Soule:2005:TMT

[1452] Augustin Soule, Kavé Salamatian, Antonio Nucci, and Nina Taft. Traffic matrix tracking using Kalman filters. ACM SIGMETRICS Performance Evaluation Review, 33(3):24–31, December 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lance:2005:RTT

[1453] Ryan Lance and Ian Frommer. Roundtrip time inference via passive monitoring. ACM SIGMETRICS Performance Evaluation Review, 33(3):32–38, December 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lawrence:2005:LAN

[1454] Earl Lawrence, George Michailidis, and Vijay N. Nair. Local area network analysis using end-to-end delay tomography. ACM SIGMETRICS Performance Evaluation Review, 33(3):39–45, December 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tian:2005:TAL

[1455] Wenhong Tian. The transient analysis of loss networks. ACM SIGMET-RICS Performance Evaluation Review, 33(3):46–50, December 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fiedler:2005:TMT

[1456] Daniel Fiedler, Kristen Walcott, Thomas Richardson, Gregory M. Kapfhammer, Ahmed Amer, and Panos K. Chrysanthis. Towards the measurement of tuple space performance. ACM SIGMETRICS Performance Evaluation Review, 33(3):51–62, December 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Riska:2006:GEF

[1457] Alma Riska and Erik Riedel. Guest Editor’s foreword: bigger and faster and smaller. ACM SIGMETRICS Performance Evaluation Review, 33(4):2–3, March 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Keeton:2006:CMD

[1458] Kimberly Keeton and Arif Merchant.

Challenges in managing dependable data systems. ACM SIGMETRICS Performance Evaluation Review, 33(4): 4–10, March 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2006:ACT

[1459] Jianyong Zhang, Prasenjit Sarkar, and Anand Sivasubramaniam. Achieving completion time guarantees in an opportunistic data migration scheme. ACM SIGMETRICS Performance Evaluation Review, 33(4):11–16, March 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thomasian:2006:MLR

[1460] Alexander Thomasian. Multi-level RAID for very large disk arrays. ACM SIGMETRICS Performance Evaluation Review, 33(4):17–22, March 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mesnier:2006:RFM

[1461] Michael Mesnier, Matthew Wachs, Brandon Salmon, and Gregory R. Ganger. Relative fitness models for storage. ACM SIGMETRICS Performance Evaluation Review, 33(4):23–28, March 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arpaci-Dusseau:2006:SSD

[1462] Andrea C. Arpaci-Dusseau, Remzi H. Arpaci-Dusseau, Lakshmi N. Bairavasundaram, Timothy E. Denehy, Florentina I. Popovici, Vijayan Prabhakaran, and Muthian Sivathanu. Semantically-smart disk systems: past, present, and future. ACM SIGMET-RICS Performance Evaluation Review, 33(4):29–35, March 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bachmat:2006:BDS

[1463] Eitan Bachmat and Vladimir Braverman. Batched disk scheduling with delays. ACM SIGMETRICS Performance Evaluation Review, 33(4):36–41, March 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zarandioon:2006:OOD

[1464] Saman Zarandioon and Alexander Thomasian. Optimization of online disk scheduling algorithms. ACM SIGMETRICS Performance Evaluation Review, 33(4):42–46, March 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reed:2006:PRU

[1465] Daniel A. Reed. Performance and reliability: the ubiquitous challenge. ACM SIGMETRICS Performance Evaluation Review, 34(1):1–2, June 2006. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thereska:2006:STA

[1466] Eno Thereska, Brandon Salmon, John Strunk, Matthew Wachs, Michael Abd-El-Malek, Julio Lopez, and Gregory R.

Ganger. Stardust: tracking activity in a distributed storage system. ACM SIGMETRICS Performance Evaluation Review, 34(1):3–14, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pinheiro:2006:ERC

[1467] Eduardo Pinheiro, Ricardo Bianchini, and Cezary Dubnicki. Exploiting redundancy to conserve energy in storage systems. ACM SIGMETRICS Performance Evaluation Review, 34(1):15–26, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Modiano:2006:MTW

[1468] Eytan Modiano, Devavrat Shah, and Gil Zussman. Maximizing throughput in wireless networks via gossiping. ACM SIGMETRICS Performance Evaluation Review, 34(1):27–38, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gao:2006:DEE

[1469] Yan Gao, Dah-Ming Chiu, and John C. S. Lui. Determining the end-to-end throughput capacity in multi-hop networks: methodology and applications. ACM SIGMETRICS Performance Evaluation Review, 34(1):39–50, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Koksal:2006:ICV

[1470] Can Emre Koksal, Kyle Jamieson, Emre Telatar, and Patrick Thiran. Impacts of channel variability on link-level throughput in wireless networks. ACM SIGMETRICS Performance Evaluation Review, 34(1):51–62, June 2006.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mishra:2006:POC

[1471] Arunesh Mishra, Vivek Shrivastava, Suman Banerjee, and William Arbaugh. Partially overlapped channels not considered harmful. ACM SIGMETRICS Performance Evaluation Review, 34(1):63–74, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lieshout:2006:GSS

[1472] P. Lieshout, M. Mandjes, and S. Borst. GPS scheduling: selection of optimal weights and comparison with strict priorities. ACM SIGMETRICS Performance Evaluation Review, 34(1):75–86, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gromoll:2006:IRP

[1473] H. Christian Gromoll, Philippe Robert, Bert Zwart, and Richard Bakker. The impact of reneging in processor sharing queues. ACM SIGMETRICS Performance Evaluation Review, 34(1):87–96, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2006:TAP

[1474] Chang-Woo Yang, Adam Wierman, Sanjay Shakkottai, and Mor Harchol-Balter. Tail asymptotics for policies favoring short jobs in a many-flows regime. ACM SIGMETRICS Performance Evaluation Review, 34(1):97– 108, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2006:LHT

[1475] Thomas Bonald, Aleksi Penttinen, and Jorma Virtamo. On light and heavy traffic approximations of balanced fairness. ACM SIGMETRICS Performance Evaluation Review, 34(1):109– 120, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Song:2006:NFF

[1476] Han Hee Song, Lili Qiu, and Yin Zhang. NetQuest: a flexible framework for large-scale network measurement. ACM SIGMETRICS Performance Evaluation Review, 34(1):121– 132, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhao:2006:RTM

[1477] Qi Zhao, Zihui Ge, Jia Wang, and Jun Xu. Robust traffic matrix estimation with imperfect information: making use of multiple data sources. ACM SIGMETRICS Performance Evaluation Review, 34(1):133–144, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lall:2006:DSA

[1478] Ashwin Lall, Vyas Sekar, Mitsunori Ogihara, Jun Xu, and Hui Zhang. Data streaming algorithms for estimating entropy of network traffic. ACM SIGMETRICS Performance Evaluation Review, 34(1):145–156, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2006:SEE

[1479] Sanghwan Lee, Zhi-Li Zhang, Sambit Sahu, and Debanjan Saha. On suitability of Euclidean embedding of Internet hosts. ACM SIGMETRICS Performance Evaluation Review, 34 (1):157–168, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2006:EAE

[1480] Giuliano Casale. An efficient algorithm for the exact analysis of multiclass queueing networks with large population sizes. ACM SIGMET-RICS Performance Evaluation Review, 34(1):169–180, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

VanVelthoven:2006:TAT

[1481] J. Van Velthoven, B. Van Houdt, and C. Blondia. Transient analysis of tree-like processes and its application to random access systems. ACM SIGMETRICS Performance Evaluation Review, 34(1):181–190, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Buchholz:2006:BSR

[1482] Peter Buchholz. Bounding stationary results of Tandem networks with MAP input and PH service time distributions. ACM SIGMETRICS Performance Evaluation Review, 34(1):191– 202, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:2006:FCQ

[1483] Varun Gupta, Mor Harchol-Balter, Alan Scheller Wolf, and Uri Yechiali. Fundamental characteristics of queues with fluctuating load. ACM SIGMET-RICS Performance Evaluation Review, 34(1):203–215, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Narayanasamy:2006:ALO

[1484] Satish Narayanasamy, Cristiano Pereira, Harish Patil, Robert Cohn, and Brad Calder. Automatic logging of operating system effects to guide applicationlevel architecture simulation. ACM SIGMETRICS Performance Evaluation Review, 34(1):216–227, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Guo:2006:AMC

[1485] Fei Guo and Yan Solihin. An analytical model for cache replacement policy performance. ACM SIGMET-RICS Performance Evaluation Review, 34(1):228–239, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Olshefski:2006:UMC

[1486] David Olshefski and Jason Nieh. Understanding the management of client perceived response time. ACM SIG-METRICS Performance Evaluation Review, 34(1):240–251, June 2006. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thorup:2006:CIP

[1487] Mikkel Thorup. Confidence intervals for priority sampling. ACM SIGMET-RICS Performance Evaluation Review, 34(1):252–263, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Osogami:2006:FPBa

[1488] Takayuki Osogami and Toshinari Itoko. Finding probably better system configurations quickly. ACM SIGMET-RICS Performance Evaluation Review, 34(1):264–275, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2006:EMN

[1489] Thomas Bonald. The Erlang model with non-Poisson call arrivals. ACM SIGMETRICS Performance Evaluation Review, 34(1):276–286, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fidler:2006:WDS

[1490] Markus Fidler and Jens B. Schmitt. On the way to a distributed systems calculus: an end-to-end network calculus with data scaling. ACM SIGMET-RICS Performance Evaluation Review, 34(1):287–298, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Peserico:2006:RNC

[1491] Enoch Peserico and Larry Rudolph. Robust network connectivity: when it’s the big picture that matters. ACM SIGMETRICS Performance Evaluation Review, 34(1):299–310, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dong:2006:PCT

[1492] Qunfeng Dong, Suman Banerjee, Jia Wang, Dheeraj Agrawal, and Ashutosh Shukla. Packet classifiers in ternary CAMs can be smaller. ACM SIGMET-RICS Performance Evaluation Review, 34(1):311–322, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhao:2006:DNS

[1493] Qi Zhao, Jun Xu, and Zhen Liu. Design of a novel statistics counter architecture with optimal space and time efficiency. ACM SIGMETRICS Performance Evaluation Review, 34(1):323– 334, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kumar:2006:FMP

[1494] Rakesh Kumar, David D. Yao, Amitabha Bagchi, Keith W. Ross, and Dan Rubenstein. Fluid modeling of pollution proliferation in P2P networks. ACM SIGMETRICS Performance Evaluation Review, 34(1):335– 346, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2006:FSS

[1495] Kang Li and Zhenyu Zhong. Fast statistical spam filter by approximate classifications. ACM SIGMETRICS Performance Evaluation Review, 34 (1):347–358, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kola:2006:QAB

[1496] George Kola and Mary K. Vernon.

QuickProbe: available bandwidth estimation in two roundtrips. ACM SIGMETRICS Performance Evaluation Review, 34(1):359–360, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kaushik:2006:FTW

[1497] Neena R. Kaushik, Silvia M. Figueira, and Stephen A. Chiappari. Flexible time-windows for advance reservation in LambdaGrids. ACM SIGMET-RICS Performance Evaluation Review, 34(1):361–362, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Verbowski:2006:APS

[1498] Chad Verbowski, Emre Kiciman, Brad Daniels, Yi-Min Wang, Roussi Roussev, Shan Lu, and Juhan Lee. Analyzing persistent state interactions to improve state management. ACM SIGMETRICS Performance Evaluation Review, 34(1):363–364, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Verloop:2006:DOS

[1499] Maaike Verloop, Rudesindo Núñez-Queija, and Sem Borst. Delay-optimal scheduling in bandwidth-sharing networks. ACM SIGMETRICS Performance Evaluation Review, 34(1):365– 366, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menth:2006:TPP

[1500] Michael Menth, Robert Henjes, Christian Zepfel, and Sebastian Gehrsitz. Throughput performance of popular JMS servers. ACM SIGMETRICS Performance Evaluation Review, 34 (1):367–368, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Garg:2006:OHR

[1501] Rahul Garg and Yogish Sabharwal.

Optimizing the HPCC randomaccess benchmark on Blue Gene/L supercomputer. ACM SIGMETRICS Performance Evaluation Review, 34(1):369– 370, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Piotrowski:2006:PPS

[1502] Tadeusz Piotrowski, Suman Banerjee, Sudeept Bhatnagar, Samrat Ganguly, and Rauf Izmailov. Peer-to-peer streaming of stored media: the indirect approach. ACM SIGMET-RICS Performance Evaluation Review, 34(1):371–372, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dholakia:2006:ANI

[1503] Ajay Dholakia, Evangelos Eleftheriou, Xiao-Yu Hu, Ilias Iliadis, Jai Menon, and KK Rao. Analysis of a new intra-disk redundancy scheme for highreliability RAID storage systems in the presence of unrecoverable errors. ACM SIGMETRICS Performance Evaluation Review, 34(1):373–374, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bower:2006:AAV

[1504] Fred A. Bower, Derek Hower, Mahmut Yilmaz, Daniel J. Sorin, and Sule Ozev. Applying architectural vulnerability Analysis to hard faults in the microprocessor. ACM SIGMET-RICS Performance Evaluation Review, 34(1):375–376, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Broberg:2006:MFM

[1505] James A. Broberg, Zhen Liu, Cathy H. Xia, and Li Zhang. A multicommodity flow model for distributed stream processing. ACM SIGMETRICS Performance Evaluation Review, 34(1):377– 378, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2006:GEF

[1506] T. Bonald. Guest Editor’s foreword. ACM SIGMETRICS Performance Evaluation Review, 34(2):2, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hautphenne:2006:EPP

[1507] Sophie Hautphenne, Kenji Leibnitz, and Marie-Ange Remiche. Extinction probability in peer-to-peer file diffusion. ACM SIGMETRICS Performance Evaluation Review, 34(2):3–4, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mundinger:2006:APPa

[1508] J. Mundinger, R. R. Weber, and G. Weiss. Analysis of peer-to-peer file dissemination amongst users of different upload capacities. ACM SIGMET-RICS Performance Evaluation Review, 34(2):5–6, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Besson:2006:GSE

[1509] Emmanuel Besson, Aline Gouget, and Hervé Sibert. The GAIA sensor: an early DDoS detection tool. ACM SIGMETRICS Performance Evaluation Review, 34(2):7–8, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hardy:2006:PCR

[1510] G. Hardy, C. Lucet, and N. Limnios.

Probability of connection in regular stochastic networks. ACM SIGMET-RICS Performance Evaluation Review, 34(2):9–10, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Busic:2006:BTS

[1511] Ana Bušić and Jean-Michel Fourneau.

Bounding transient and steady-state dependability measures through algorithmic stochastic comparison. ACM SIGMETRICS Performance Evaluation Review, 34(2):11–12, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bossie:2006:CHT

[1512] Craig Bossie and Pierre M. Fiorini. On checkpointing and heavy-tails in unreliable computing environments. ACM SIGMETRICS Performance Evaluation Review, 34(2):13–15, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mickens:2006:IDS

[1513] James W. Mickens and Brian D. Noble. Improving distributed system performance using machine availability prediction. ACM SIGMETRICS Performance Evaluation Review, 34 (2):16–18, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chydzinski:2006:BOC

[1514] Andrzej Chydzinski. Buffer overflow calculations in a batch arrival queue. ACM SIGMETRICS Performance Evaluation Review, 34(2):19–21, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menasce:2006:ECP

[1515] Daniel A. Menascé and Vasudeva Akula. Evaluating caching policies for online auctions. ACM SIGMET-RICS Performance Evaluation Review, 34(2):22–23, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vincent:2006:PSI

[1516] Jean-Marc Vincent and Jérôme Vienne. Perfect simulation of index based routing queueing networks. ACM SIGMETRICS Performance Evaluation Review, 34(2):24–25, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chang:2006:STQ

[1517] Cheng-Shang Chang, Yi-Ting Chen, Jay Cheng, Po-Kai Huang, and Duan-Shin Lee. From switching theory to ‘queueing’ theory. ACM SIGMET-RICS Performance Evaluation Review, 34(2):26–28, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Giannoulis:2006:CLP

[1518] Anastasios Giannoulis, Konstantinos P. Tsoukatos, and Leandros Tassiulas. Cross-layer power control in wireless networks. ACM SIGMETRICS Performance Evaluation Review, 34(2): 29–31, September 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2006:F

[1519] Mark S. Squillante. Foreword. ACM SIGMETRICS Performance Evaluation Review, 34(3):2, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nakassis:2006:TPQ

[1520] Anastase Nakassis and Vladimir Marbukh. Towards power and QoS aware wireless networks. ACM SIGMET-RICS Performance Evaluation Review, 34(3):3–5, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yazici:2006:EPD

[1521] Emine Şule Yazici, Selda Küçükçifçi, Öznur Özkasap, and Mine Çağlar. Exact probability distributions for peerto-peer epidemic information diffusion. ACM SIGMETRICS Performance Evaluation Review, 34(3):6–8, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Luan:2006:MOC

[1522] Hao Luan, Danny H. K. Tsang, and Kin Wah Kwong. Media overlay construction via a Markov chain Monte Carlo method. ACM SIGMETRICS Performance Evaluation Review, 34 (3):9–11, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mundinger:2006:APPb

[1523] Jochen Mundinger, Richard Weber, and Gideon Weiss. Analysis of peerto-peer file dissemination. ACM SIGMETRICS Performance Evaluation Review, 34(3):12–14, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raz:2006:TMS

[1524] David Raz, Hanoch Levy, and Benjamin Avi-Itzhak. On the twin measure and system predictability and fairness. ACM SIGMETRICS Performance Evaluation Review, 34(3):15–17, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brown:2006:CFP

[1525] Patrick Brown. Comparing FB and PS scheduling policies. ACM SIGMET-RICS Performance Evaluation Review, 34(3):18–20, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wierman:2006:EIS

[1526] Adam Wierman. On the effect of inexact size information in size based policies. ACM SIGMETRICS Performance Evaluation Review, 34(3):21–23, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sheahan:2006:CTD

[1527] Robert Sheahan, Lester Lipsky, Pierre M. Fiorini, and Søren Asmussen. On the completion time distribution for tasks that must restart from the beginning if a failure occurs. ACM SIGMET-RICS Performance Evaluation Review, 34(3):24–26, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Silveira:2006:MST

[1528] Fernando Silveira and Edmundo de Souza e Silva. Modeling the shortterm dynamics of packet losses. ACM SIGMETRICS Performance Evaluation Review, 34(3):27–29, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ott:2006:SSP

[1529] Teunis J. Ott and Jason Swanson. Stationarity of some processes in transport protocols. ACM SIGMETRICS Performance Evaluation Review, 34 (3):30–32, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baryshnikov:2006:FDT

[1530] Yuliy Baryshnikov, Ed Coffman, Jing Feng, and Vishal Misra. Free-drop TCP. ACM SIGMETRICS Performance Evaluation Review, 34(3):33–35, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Carofiglio:2006:ARS

[1531] G. Carofiglio, C. Chiasserini, M. Garetto, and E. Leonardi. Analysis of route stability under the random direction mobility model. ACM SIGMETRICS Performance Evaluation Review, 34 (3):36–38, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Osogami:2006:FPBb

[1532] Takayuki Osogami. Finding probably best system configurations quickly. ACM SIGMETRICS Performance Evaluation Review, 34(3):39–41, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yao:2006:AOT

[1533] David D. Yao and Heng-Qing Ye. Asymptotic optimality of threshold control in a stochastic network based on a fixed-point approximation. ACM SIGMETRICS Performance Evaluation Review, 34(3):42–44, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bayati:2006:OSM

[1534] Mohsen Bayati, Mayank Sharma, and Mark S. Squillante. Optimal scheduling in a multiserver stochastic network. ACM SIGMETRICS Performance Evaluation Review, 34(3):45–47, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Elhaddad:2006:ATS

[1535] Mahmoud Elhaddad, Rami Melhem, and Taieb Znati. Analysis of a transmission scheduling algorithm for supporting bandwidth guarantees in bufferless networks. ACM SIGMET-RICS Performance Evaluation Review, 34(3):48–63, December 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harchol-Balter:2007:F

[1536] Mor Harchol-Balter. Foreword. ACM SIGMETRICS Performance Evaluation Review, 34(4):2–3, March 2007.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wierman:2007:FC

[1537] Adam Wierman. Fairness and classifications. ACM SIGMETRICS Performance Evaluation Review, 34(4):4–12, March 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Boxma:2007:TS

[1538] Onno Boxma and Bert Zwart. Tails in scheduling. ACM SIGMETRICS Performance Evaluation Review, 34(4):13– 20, March 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Biersack:2007:SP

[1539] Ernst W. Biersack, Bianca Schroeder, and Guillaume Urvoy-Keller. Scheduling in practice. ACM SIGMET-RICS Performance Evaluation Review, 34(4):21–28, March 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2007:SNT

[1540] Thomas Bonald and James Roberts.

Scheduling network traffic. ACM SIGMETRICS Performance Evaluation Review, 34(4):29–35, March 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aalto:2007:BPS

[1541] Samuli Aalto, Urtzi Ayesta, Sem Borst, Vishal Misra, and Rudesindo Núñez-Queija. Beyond processor sharing. ACM SIGMETRICS Performance Evaluation Review, 34(4):36–43, March 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2007:SAM

[1542] Mark S. Squillante. Stochastic analysis of multiserver systems. ACM SIGMETRICS Performance Evaluation Review, 34(4):44–51, March 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pruhs:2007:COS

[1543] Kirk Pruhs. Competitive online scheduling for server systems. ACM SIGMETRICS Performance Evaluation Review, 34(4):52–58, March 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2007:AMJ

[1544] Hui Li and Michael Muskulus. Analysis and modeling of job arrivals in a production grid. ACM SIGMET-RICS Performance Evaluation Review, 34(4):59–70, March 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kadayif:2007:MID

[1545] Ismail Kadayif and Mahmut Kandemir. Modeling and improving data cache reliability: 1. ACM SIGMETRICS Performance Evaluation Review, 35(1):12, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gulati:2007:PAC

[1546] Ajay Gulati, Arif Merchant, and Peter J. Varman. pClock: an arrival curve based approach for QoS guarantees in shared storage systems. ACM SIGMETRICS Performance Evaluation Review, 35(1):13–24, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Iyer:2007:QPA

[1547] Ravi Iyer, Li Zhao, Fei Guo, Ramesh Illikkal, Srihari Makineni, Don Newell, Yan Solihin, Lisa Hsu, and Steve Reinhardt. QoS policies and architecture for cache/memory in CMP platforms. ACM SIGMETRICS Performance Evaluation Review, 35(1):25–36, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mesnier:2007:MRF

[1548] Michael P. Mesnier, Matthew Wachs, Raja R. Sambasivan, Alice X. Zheng, and Gregory R. Ganger. Modeling the relative fitness of storage. ACM SIGMETRICS Performance Evaluation Review, 35(1):37–48, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wen:2007:FFI

[1549] Zhihua Wen, Sipat Triukose, and Michael Rabinovich. Facilitating focused Internet measurements. ACM SIGMETRICS Performance Evaluation Review, 35(1):49–60, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Huang:2007:DND

[1550] Yiyi Huang, Nick Feamster, Anukool Lakhina, and Jim (Jun) Xu. Diagnosing network disruptions with networkwide analysis. ACM SIGMETRICS Performance Evaluation Review, 35(1): 61–72, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pucha:2007:UND

[1551] Himabindu Pucha, Ying Zhang, Z. Morley Mao, and Y. Charlie Hu.

Understanding network delay changes caused by routing events. ACM SIG-METRICS Performance Evaluation Review, 35(1):73–84, June 2007. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kashyap:2007:TPR

[1552] Abhishek Kashyap, Sudipta Sengupta, Randeep Bhatia, and M. Kodialam. Two-phase routing, scheduling and power control for wireless mesh networks with variable traffic. ACM SIGMETRICS Performance Evaluation Review, 35(1):85–96, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mirza:2007:MLA

[1553] Mariyam Mirza, Joel Sommers, Paul Barford, and Xiaojin Zhu. A machine learning approach to TCP throughput prediction. ACM SIGMETRICS Performance Evaluation Review, 35(1):97– 108, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ringberg:2007:SPT

[1554] Haakon Ringberg, Augustin Soule, Jennifer Rexford, and Christophe Diot. Sensitivity of PCA for traffic anomaly detection. ACM SIGMETRICS Performance Evaluation Review, 35(1):109– 120, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2007:BCS

[1555] Seungjoon Lee, Dave Levin, Vijay Gopalakrishnan, and Bobby Bhattacharjee. Backbone construction in selfish wireless networks. ACM SIGMETRICS Performance Evaluation Review, 35(1):121–132, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xia:2007:SFJ

[1556] Cathy H. Xia, Zhen Liu, Don Towsley, and Marc Lelarge. Scalability of fork/ join queueing networks with blocking. ACM SIGMETRICS Performance Evaluation Review, 35(1):133– 144, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Osogami:2007:OSC

[1557] Takayuki Osogami and Sei Kato. Optimizing system configurations quickly by guessing at the performance. ACM SIGMETRICS Performance Evaluation Review, 35(1):145–156, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2007:SSR

[1558] Zhe Wang, Wei Dong, William Josephson, Qin Lv, Moses Charikar, and Kai Li. Sizing sketches: a rank-based analysis for similarity search. ACM SIGMETRICS Performance Evaluation Review, 35(1):157–168, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Park:2007:MEP

[1559] Soyeon Park, Weihang Jiang, Yuanyuan Zhou, and Sarita Adve. Managing energy-performance tradeoffs for multithreaded applications on multiprocessor architectures. ACM SIGMET-RICS Performance Evaluation Review, 35(1):169–180, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cvetkovski:2007:AAC

[1560] Andrej Cvetkovski. An algorithm for approximate counting using limited memory resources. ACM SIGMET-RICS Performance Evaluation Review, 35(1):181–190, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2007:SDN

[1561] Eric S. Lee and Thom Whalen. Synthetic designs: a new form of true experimental design for use in information systems development. ACM SIGMETRICS Performance Evaluation Review, 35(1):191–202, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Feng:2007:PUP

[1562] Hanhua Feng, Vishal Misra, and Dan Rubenstein. PBS: a unified prioritybased scheduler. ACM SIGMET-RICS Performance Evaluation Review, 35(1):203–214, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jelenkovic:2007:ASC

[1563] Predrag R. Jelenkovic, Xiaozhu Kang, and Jian Tan. Adaptive and scalable comparison scheduling. ACM SIGMETRICS Performance Evaluation Review, 35(1):215–226, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bhadra:2007:OCP

[1564] Sandeep Bhadra, Yingdong Lu, and Mark S. Squillante. Optimal capacity planning in stochastic loss networks with time-varying workloads. ACM SIGMETRICS Performance Evaluation Review, 35(1):227–238, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2007:FLS

[1565] Jiaping Liu, Alexandre Proutière, Yung Yi, Mung Chiang, and H. Vincent Poor. Flow-level stability of data networks with non-convex and timevarying rate regions. ACM SIGMET-RICS Performance Evaluation Review, 35(1):239–250, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Smirni:2007:FDP

[1566] Evgenia Smirni, Frederica Darema, Albert Greenberg, Adolfy Hoisie, and Don Towsley. Future directions in performance evaluation research. ACM SIGMETRICS Performance Evaluation Review, 35(1):251–252, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dong:2007:WSP

[1567] Qunfeng Dong, Suman Banerjee, Jia Wang, and Dheeraj Agrawal. Wire speed packet classification without TCAMs: a few more registers (and a bit of logic) are enough. ACM SIGMETRICS Performance Evaluation Review, 35(1):253–264, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hirzel:2007:DLO

[1568] Martin Hirzel. Data layouts for objectoriented programs. ACM SIGMET-RICS Performance Evaluation Review, 35(1):265–276, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hao:2007:BHA

[1569] Fang Hao, Murali Kodialam, and T. V. Lakshman. Building high accuracy Bloom filters using partitioned hashing. ACM SIGMETRICS Performance Evaluation Review, 35(1):277– 288, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bairavasundaram:2007:ALS

[1570] Lakshmi N. Bairavasundaram, Garth R. Goodson, Shankar Pasupathy, and Jiri Schindler. An analysis of latent sector errors in disk drives. ACM SIGMET-RICS Performance Evaluation Review, 35(1):289–300, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Legout:2007:CSI

[1571] Arnaud Legout, Nikitas Liogkas, Eddie Kohler, and Lixia Zhang. Clustering and sharing incentives in Bit-Torrent systems. ACM SIGMET-RICS Performance Evaluation Review, 35(1):301–312, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sanghavi:2007:DLS

[1572] Sujay Sanghavi, Loc Bui, and R. Srikant. Distributed link scheduling with constant overhead. ACM SIGMETRICS Performance Evaluation Review, 35(1):313–324, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rajendran:2007:TBC

[1573] Raj Kumar Rajendran, Vishal Misra, and Dan Rubenstein. Theoretical bounds on control-plane selfmonitoring in routing protocols. ACM SIGMETRICS Performance Evaluation Review, 35(1):325–336, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yuan:2007:ORF

[1574] Xin Yuan, Wickus Nienaber, Zhenhai Duan, and Rami Melhem. Oblivious routing for fat-tree based system area networks with uncertain traffic demands. ACM SIGMETRICS Performance Evaluation Review, 35(1):337– 348, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nahum:2007:ESS

[1575] Erich M. Nahum, John Tracey, and Charles P. Wright. Evaluating SIP server performance. ACM SIGMET-RICS Performance Evaluation Review, 35(1):349–350, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Puzak:2007:PS

[1576] Thomas R. Puzak, Allan Hartstein, Viji Srinivasan, Philip Emma, and Arthur Nadas. Pipeline spectroscopy. ACM SIGMETRICS Performance Evaluation Review, 35(1):351– 352, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cohen:2007:BSB

[1577] Edith Cohen and Haim Kaplan.

Bottom-k sketches: better and more efficient estimation of aggregates. ACM SIGMETRICS Performance Evaluation Review, 35(1):353–354, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gu:2007:GEM

[1578] Yu Gu, Lee Breslau, Nick G. Duffield, and Subhabrata Sen. GRE encapsulated multicast probing: a scalable technique for measuring one-way loss. ACM SIGMETRICS Performance Evaluation Review, 35(1):355– 356, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mirkovic:2007:WSR

[1579] Jelena Mirkovic, Alefiya Hussain, Brett Willson, Sonia Fahmy, Wei-Min Yao, Peter Reiher, Stephen Schwab, and Roshan Thomas. When is service really denied?: a user-centric DoS metric. ACM SIGMETRICS Performance Evaluation Review, 35(1):357– 358, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Guo:2007:DIM

[1580] Lei Guo, Enhua Tan, Songqing Chen, Zhen Xiao, and Xiaodong Zhang. Does Internet media traffic really follow Zipflike distribution? ACM SIGMET-RICS Performance Evaluation Review, 35(1):359–360, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hoflehner:2007:CCS

[1581] Gerolf F. Hoflehner, Darshan Desai, Daniel M. Lavery, Alexandru Nicolau, and Alexander V. Veidenbaum.

Comparative characterization of SPEC CPU2000 and CPU2006 on Itanium(R) architecture. ACM SIGMETRICS Performance Evaluation Review, 35 (1):361–362, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:2007:PRT

[1582] Bin Lin, Arindam Mallik, Peter A.

Dinda, Gokhan Memik, and Robert P. Dick. Power reduction through measurement and modeling of users and CPUs: summary. ACM SIGMET-RICS Performance Evaluation Review, 35(1):363–364, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2007:GRI

[1583] Chong Wang and John W. Byers. Generating representative ISP topologies from first-principles. ACM SIGMET-RICS Performance Evaluation Review, 35(1):365–366, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bissias:2007:BDL

[1584] George Dean Bissias, Brian Neil Levine, and Arnold Rosenberg. Bounding damage from link destruction, with application to the Internet. ACM SIGMETRICS Performance Evaluation Review, 35(1):367–368, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Erman:2007:SSN

[1585] Jeffrey Erman, Anirban Mahanti, Martin Arlitt, Ira Cohen, and Carey Williamson. Semi-supervised network traffic classification. ACM SIGMET-RICS Performance Evaluation Review, 35(1):369–370, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mi:2007:EMI

[1586] Ningfang Mi, Alma Riska, Qi Zhang, Evgenia Smirni, and Erik Riedel. Efficient management of idleness in systems. ACM SIGMETRICS Performance Evaluation Review, 35(1):371– 372, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

deJager:2007:AIS

[1587] Douglas V. de Jager and Jeremy T. Bradley. Asynchronous iterative solution for state-based performance metrics. ACM SIGMETRICS Performance Evaluation Review, 35(1):373– 374, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hoste:2007:ACP

[1588] Kenneth Hoste, Lieven Eeckhout, and Hendrik Blockeel. Analyzing commercial processor performance numbers for predicting performance of applications of interest. ACM SIGMET-RICS Performance Evaluation Review, 35(1):375–376, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

He:2007:BSS

[1589] Jiayue He and Augustin Chaintreau. BRADO: scalable streaming through reconfigurable trees. ACM SIGMET-RICS Performance Evaluation Review, 35(1):377–378, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nurmi:2007:QQB

[1590] Daniel Charles Nurmi, John Brevik, and Rich Wolski. QBETS: queue bounds estimation from time series. ACM SIGMETRICS Performance Evaluation Review, 35(1):379– 380, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Deng:2007:PDS

[1591] Leiwen Deng and Aleksandar Kuzmanovic. Pong: diagnosing spatiotemporal Internet congestion properties. ACM SIGMETRICS Performance Evaluation Review, 35(1):381– 382, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aalto:2007:MDO

[1592] Samuli Aalto and Urtzi Ayesta. Mean delay optimization for the M/G/1 queue with Pareto type service times. ACM SIGMETRICS Performance Evaluation Review, 35(1):383–384, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2007:F

[1593] Mark S. Squillante. Foreword. ACM SIGMETRICS Performance Evaluation Review, 35(2):2, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gianini:2007:PNR

[1594] Gabriele Gianini and Ernesto Damiani. Poisson-noise removal in selfsimilarity studies based on packetcounting: factorial-moment/strip-integral approach. ACM SIGMET-RICS Performance Evaluation Review, 35(2):3–5, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marbukh:2007:FBS

[1595] Vladimir Marbukh. Fair bandwidth sharing under flow arrivals/departures: effect of retransmissions on stability and performance. ACM SIGMET-RICS Performance Evaluation Review, 35(2):6–8, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Osogami:2007:AMT

[1596] Takayuki Osogami. Accuracy of measured throughputs and mean response times. ACM SIGMETRICS Performance Evaluation Review, 35(2):9–11, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:2007:EHM

[1597] Varun Gupta, Jim Dai, Mor Harchol-Balter, and Bert Zwart. The effect of higher moments of job size distribution on the performance of an m/g/s queueing system. ACM SIGMET-RICS Performance Evaluation Review, 35(2):12–14, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hossfeld:2007:MOT

[1598] Tobias Hoßfeld, Kenji Leibnitz, and Marie-Ange Remiche. Modeling of an online TV recording service. ACM SIGMETRICS Performance Evaluation Review, 35(2):15–17, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2007:OTC

[1599] Peng Wang and Stephan Bohacek.

An overview of tractable computation of optimal scheduling and routing in mesh networks. ACM SIGMET-RICS Performance Evaluation Review, 35(2):18–20, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ciucu:2007:ESE

[1600] Florin Ciucu. Exponential supermartingales for evaluating end-to-end backlog bounds. ACM SIGMETRICS Performance Evaluation Review, 35 (2):21–23, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:2007:IPS

[1601] Varun Gupta, Karl Sigman, Mor Harchol-Balter, and Ward Whitt. Insensitivity for PS server farms with JSQ routing. ACM SIGMETRICS Performance Evaluation Review, 35 (2):24–26, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2007:CMA

[1602] Giuliano Casale, Eddy Z. Zhang, and Evgenia Smirni. Characterization of moments and autocorrelation in MAPs. ACM SIGMETRICS Performance Evaluation Review, 35(2):27–29, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Field:2007:AAN

[1603] Tony Field and Peter Harrison. Approximate analysis of a network of fluid queues. ACM SIGMETRICS Performance Evaluation Review, 35 (2):30–32, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reich:2007:TCU

[1604] Joshua Reich, Vishal Misra, and Dan Rubenstein. The time-correlated update problem. ACM SIGMETRICS Performance Evaluation Review, 35 (2):33–35, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kang:2007:PFS

[1605] W. N. Kang, F. P. Kelly, N. H. Lee, and R. J. Williams. Product form stationary distributions for diffusion approximations to a flow-level model operating under a proportional fair sharing policy. ACM SIGMET-RICS Performance Evaluation Review, 35(2):36–38, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2007:OCP

[1606] Yingdong Lu, Ana Radovanović, and Mark S. Squillante. Optimal capacity planning in stochastic loss networks. ACM SIGMETRICS Performance Evaluation Review, 35(2):39–41, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cherkasova:2007:CTC

[1607] Ludmila Cherkasova, Diwaker Gupta, and Amin Vahdat. Comparison of the three CPU schedulers in Xen. ACM SIGMETRICS Performance Evaluation Review, 35(2):42–51, September 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marsan:2007:F

[1608] Marco Ajmone Marsan and Prashant Shenoy. Foreword. ACM SIGMET-RICS Performance Evaluation Review, 35(3):2–3, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cesana:2007:EPC

[1609] M. Cesana, L. Campelli, and F. Borgonovo. Efficiency of physical carrier sensing in wireless access networks. ACM SIGMETRICS Performance Evaluation Review, 35(3):4–6, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cano:2007:HDE

[1610] Juan-Carlos Cano, José-Manuel Cano, Eva González, Carlos Calafate, and Pietro Manzoni. How does energy consumption impact performance in Bluetooth? ACM SIGMETRICS Performance Evaluation Review, 35 (3):7–9, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lukas:2007:IBL

[1611] Georg Lukas, André Herms, and Daniel Mahrenholz. Interval based off-line clock synchronization for wireless mesh networks. ACM SIGMETRICS Performance Evaluation Review, 35(3):10–12, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chydzinski:2007:SFB

[1612] Andrzej Chydzinski. Solving finitebuffer queues with Markovian arrivals. ACM SIGMETRICS Performance Evaluation Review, 35(3):13–15, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ciardo:2007:ASM

[1613] Gianfranco Ciardo, Andrew S. Miner, Min Wan, and Andy Jinqing Yu. Approximating stationary measures of structured continuous-time Markov models using matrix diagrams. ACM SIGMETRICS Performance Evaluation Review, 35(3):16–18, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Silveira:2007:PPL

[1614] Fernando Silveira and Edmundo de Souza e Silva. Predicting packet loss statistics with hidden Markov models. ACM SIGMETRICS Performance Evaluation Review, 35(3):19–21, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menth:2007:NSM

[1615] Michael Menth, Andreas Binzenhöfer, and Stefan Mühleck. A note on source models for speech traffic. ACM SIGMETRICS Performance Evaluation Review, 35(3):22–24, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bracciale:2007:OOP

[1616] Lorenzo Bracciale, Francesca Lo Piccolo, Dario Luzzi, and Stefano Salsano. OPSS: an overlay peer-to-peer streaming simulator for large-scale networks. ACM SIGMETRICS Performance Evaluation Review, 35(3):25–27, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Engels:2007:ETS

[1617] Kai Engels, Ralf Heidger, Reinhold Kroeger, Morris Milekovic, Jan Schaefer, Markus Schmid, and Marcus Thoss. eMIVA: tool support for the instrumentation of critical distributed applications. ACM SIGMETRICS Performance Evaluation Review, 35 (3):28–30, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dube:2007:CPQ

[1618] Parijat Dube, Corinne Touati, and Laura Wynter. Capacity planning, quality of service and price wars. ACM SIGMETRICS Performance Evaluation Review, 35(3):31–33, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Begin:2007:HLA

[1619] Thomas Begin, Alexandre Brandwajn, Bruno Baynat, Bernd E. Wolfinger, and Serge Fdida. High-level approach to modeling of observed system behavior. ACM SIGMETRICS Performance Evaluation Review, 35(3):34–36, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Misra:2007:F

[1620] Vishal Misra. Foreword. ACM SIGMETRICS Performance Evaluation Review, 35(3):37, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhu:2007:LWA

[1621] Wenbin Zhu, Patrick G. Bridges, and Arthur B. Maccabe. Light-weight application monitoring and tuning with embedded gossip. ACM SIGMET-RICS Performance Evaluation Review, 35(3):38–39, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kortebi:2007:IAS

[1622] Riadh M. Kortebi, Yvon Gourhant, and Nazim Agoulmine. Interferenceaware SINR-based routing: algorithms and evaluation. ACM SIGMETRICS Performance Evaluation Review, 35 (3):40–42, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bui:2007:ORA

[1623] Loc Bui, R. Srikant, and Alexander Stolyar. Optimal resource allocation for multicast flows in multihop wireless networks. ACM SIGMET-RICS Performance Evaluation Review, 35(3):43, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mi:2007:PIA

[1624] Ningfang Mi. Performance impacts of autocorrelated flows in multi-tiered systems. ACM SIGMETRICS Performance Evaluation Review, 35(3):44–45, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kaushik:2007:RCA

[1625] Neena Kaushik, Silvia Figueira, and Stephen A. Chiappari. Resource coallocation using advance reservations with flexible time-windows. ACM SIGMETRICS Performance Evaluation Review, 35(3):46–48, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Verloop:2007:ERA

[1626] Maaike Verloop and Rudesindo Núñez-Queija. Efficient resource allocation in bandwidth-sharing networks. ACM SIGMETRICS Performance Evaluation Review, 35(3):49–50, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Miretskiy:2007:TQS

[1627] D. I. Miretskiy, W. R. W. Scheinhardt, and M. R. H. Mandjes. Tandem queue with server slow-down. ACM SIGMETRICS Performance Evaluation Review, 35(3):51–52, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Volkovich:2007:SMW

[1628] Y. Volkovich, D. Donato, and N. Litvak. Stochastic models for Web ranking. ACM SIGMETRICS Performance Evaluation Review, 35(3):53, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hylick:2007:HDP

[1629] Anthony Hylick, Andrew Rice, Brian Jones, and Ripduman Sohan. Hard drive power consumption uncovered. ACM SIGMETRICS Performance Evaluation Review, 35(3):54–55, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gulati:2007:TFE

[1630] Ajay Gulati, Peter Varman, Arif Merchant, and Mustafa Uysal. Towards fairness and efficiency in storage systems. ACM SIGMETRICS Performance Evaluation Review, 35(3):56–58, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Heimlicher:2007:EEV

[1631] Simon Heimlicher, Pavan Nuggehalli, and Martin May. End-to-end vs. hopby-hop transport. ACM SIGMET-RICS Performance Evaluation Review, 35(3):59–60, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Balakrichenan:2007:SPT

[1632] Sandoche Balakrichenan, Thomas Bugnazet, and Monique Becker. A simulation platform: for testing and optimization of ENUM architecture. ACM SIGMETRICS Performance Evaluation Review, 35(3):61–63, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mohror:2007:SEB

[1633] Kathryn Mohror and Karen L. Karavanic. Scalable event-based performance measurement in high-end environments. ACM SIGMETRICS Performance Evaluation Review, 35(3): 64–65, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vicari:2007:DRP

[1634] Claudio Vicari, Chiara Petrioli, and Francesco Lo Presti. Dynamic replica placement and traffic redirection in content delivery networks. ACM SIGMETRICS Performance Evaluation Review, 35(3):66–68, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Papadopoulos:2007:PPI

[1635] Fragkiskos Papadopoulos and Konstantinos Psounis. Predicting the performance of Internet-like networks using scaled-down replicas. ACM SIGMETRICS Performance Evaluation Review, 35(3):69–71, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shamsi:2007:PPS

[1636] Jawwad Shamsi and Monica Brockmeyer. PSON: predictable service overlay networks. ACM SIGMETRICS Performance Evaluation Review, 35 (3):72–74, December 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gilmore:2008:F

[1637] Stephen Gilmore and Jane Hillston.

Foreword. ACM SIGMETRICS Performance Evaluation Review, 35(4):2, March 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gilmore:2008:PEC

[1638] Stephen Gilmore and Jane Hillston.

Performance evaluation comes to life: quantitative methods applied to biological systems. ACM SIGMET-RICS Performance Evaluation Review, 35(4):3–13, March 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kwiatkowska:2008:UPM

[1639] Marta Kwiatkowska, Gethin Norman, and David Parker. Using probabilistic model checking in systems biology. ACM SIGMETRICS Performance Evaluation Review, 35(4):14–21, March 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jeschke:2008:PDD

[1640] Matthias Jeschke, Roland Ewald, Alfred Park, Richard Fujimoto, and Adelinde M. Uhrmacher. A parallel and distributed discrete event approach for spatial cell-biological simulations. ACM SIGMETRICS Performance Evaluation Review, 35(4):22–31, March 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dematte:2008:MSB

[1641] Lorenzo Dematté, Corrado Priami, and Alessandro Romanel. Modelling and simulation of biological processes in BlenX. ACM SIGMETRICS Performance Evaluation Review, 35(4):32–39, March 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sommers:2008:SPR

[1642] Joel Sommers, Paul Barford, Albert Greenberg, and Walter Willinger. An SLA perspective on the router buffer sizing problem. ACM SIGMETRICS Performance Evaluation Review, 35 (4):40–51, March 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Korzun:2008:DMR

[1643] Dmitry Korzun and Andrei Gurtov. A Diophantine model of routes in structured P2P overlays. ACM SIGMET-RICS Performance Evaluation Review, 35(4):52–61, March 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sibai:2008:EPS

[1644] Fadi N. Sibai. Evaluating the performance of single and multiple core processors with PCMARK(R)05 and benchmark analysis. ACM SIGMET-RICS Performance Evaluation Review, 35(4):62–71, March 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bordenave:2008:PRM

[1645] Charles Bordenave, David McDonald, and Alexandre Proutiere. Performance of random medium access control, an asymptotic approach. ACM SIGMET-RICS Performance Evaluation Review, 36(1):1–12, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2008:BAC

[1646] Giuliano Casale, Ningfang Mi, and Evgenia Smirni. Bound analysis of closed queueing networks with workload burstiness. ACM SIGMETRICS Performance Evaluation Review, 36(1): 13–24, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wierman:2008:SDI

[1647] Adam Wierman and Misja Nuyens. Scheduling despite inexact job-size information. ACM SIGMETRICS Performance Evaluation Review, 36(1):25– 36, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lelarge:2008:NED

[1648] Marc Lelarge and Jean Bolot. Network externalities and the deployment of security features and protocols in the Internet. ACM SIGMETRICS Performance Evaluation Review, 36(1):37–48, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brosh:2008:DFT

[1649] Eli Brosh, Salman Abdul Baset, Dan Rubenstein, and Henning Schulzrinne. The delay-friendliness of TCP. ACM SIGMETRICS Performance Evaluation Review, 36(1):49–60, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kim:2008:SVR

[1650] Changhoon Kim, Alexandre Gerber, Carsten Lund, Dan Pei, and Subhabrata Sen. Scalable VPN routing via relaying. ACM SIGMETRICS Performance Evaluation Review, 36(1):61–72, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tschopp:2008:HRD

[1651] Dominique Tschopp, Suhas Diggavi, and Matthias Grossglauser. Hierarchical routing over dynamic wireless networks. ACM SIGMETRICS Performance Evaluation Review, 36(1):73–84, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rayanchu:2008:LAN

[1652] Shravan Rayanchu, Sayandeep Sen, Jianming Wu, Suman Banerjee, and Sudipta Sengupta. Loss-aware network coding for unicast wireless sessions: design, implementation, and performance evaluation. ACM SIGMET-RICS Performance Evaluation Review, 36(1):85–96, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schmid:2008:EMV

[1653] Thomas Schmid, Zainul Charbiwala, Jonathan Friedman, Young H. Cho, and Mani B. Srivastava. Exploiting manufacturing variations for compensating environment-induced clock drift in time synchronization. ACM SIGMETRICS Performance Evaluation Review, 36(1):97–108, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cohen:2008:CEM

[1654] Edith Cohen, Nick Duffield, Carsten Lund, and Mikkel Thorup. Confident estimation for multistage measurement sampling and aggregation. ACM SIGMETRICS Performance Evaluation Review, 36(1):109–120, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2008:CBN

[1655] Yi Lu, Andrea Montanari, Balaji Prabhakar, Sarang Dharmapurikar, and Abdul Kabbani. Counter braids: a novel counter architecture for perflow measurement. ACM SIGMET-RICS Performance Evaluation Review, 36(1):121–132, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anandkumar:2008:TSB

[1656] Animashree Anandkumar, Chatschik Bisdikian, and Dakshi Agrawal. Tracking in a spaghetti bowl: monitoring transactions using footprints. ACM SIGMETRICS Performance Evaluation Review, 36(1):133–144, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Singhal:2008:OSS

[1657] Harsh Singhal and George Michailidis. Optimal sampling in state space models with applications to network monitoring. ACM SIGMETRICS Performance Evaluation Review, 36(1):145– 156, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ioannidis:2008:DHP

[1658] Stratis Ioannidis and Peter Marbach. On the design of hybrid peer-to-peer systems. ACM SIGMETRICS Performance Evaluation Review, 36(1):157– 168, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2008:UMP

[1659] Minghua Chen, Miroslav Ponec, Sudipta Sengupta, Jin Li, and Philip A. Chou. Utility maximization in peerto-peer systems. ACM SIGMET-RICS Performance Evaluation Review, 36(1):169–180, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Simatos:2008:QSM

[1660] Florian Simatos, Philippe Robert, and Fabrice Guillemin. A queueing system for modeling a file sharing principle. ACM SIGMETRICS Performance Evaluation Review, 36(1):181– 192, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goldberg:2008:PQM

[1661] Sharon Goldberg, David Xiao, Eran Tromer, Boaz Barak, and Jennifer Rexford. Path-quality monitoring in the presence of adversaries. ACM SIGMETRICS Performance Evaluation Review, 36(1):193–204, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pedarsani:2008:DAS

[1662] Pedram Pedarsani, Daniel R. Figueiredo, and Matthias Grossglauser. Densification arising from sampling fixed graphs. ACM SIGMETRICS Performance Evaluation Review, 36(1):205– 216, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Oliveira:2008:SEG

[1663] Ricardo V. Oliveira, Dan Pei, Walter Willinger, Beichuan Zhang, and Lixia Zhang. In search of the elusive ground truth: the Internet’s as-level connectivity structure. ACM SIGMET-RICS Performance Evaluation Review, 36(1):217–228, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bao:2008:HPI

[1664] Yungang Bao, Mingyu Chen, Yuan Ruan, Li Liu, Jianping Fan, Qingbo Yuan, Bo Song, and Jianwei Xu. HMTT: a platform independent fullsystem memory trace monitoring system. ACM SIGMETRICS Performance Evaluation Review, 36(1):229– 240, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Iliadis:2008:DSV

[1665] Ilias Iliadis, Robert Haas, Xiao-Yu Hu, and Evangelos Eleftheriou. Disk scrubbing versus intra-disk redundancy for high-reliability raid storage systems. ACM SIGMETRICS Performance Evaluation Review, 36(1):241– 252, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thereska:2008:IRP

[1666] Eno Thereska and Gregory R. Ganger. Ironmodel: robust performance models in the wild. ACM SIGMET-RICS Performance Evaluation Review, 36(1):253–264, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2008:XFS

[1667] Alex X. Liu, Fei Chen, JeeHyun Hwang, and Tao Xie. Xengine: a fast and scalable XACML policy evaluation engine. ACM SIGMETRICS Performance Evaluation Review, 36 (1):265–276, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Traeger:2008:DDA

[1668] Avishay Traeger, Ivan Deras, and Erez Zadok. DARC: dynamic analysis of root causes of latency distributions. ACM SIGMETRICS Performance Evaluation Review, 36(1):277– 288, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chaitanya:2008:QQM

[1669] Shiva Chaitanya, Bhuvan Urgaonkar, and Anand Sivasubramaniam. QDSL: a queuing model for systems with differential service levels. ACM SIGMET-RICS Performance Evaluation Review, 36(1):289–300, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Parvez:2008:ABL

[1670] Nadim Parvez, Carey Williamson, Anirban Mahanti, and Niklas Carlsson. Analysis of BitTorrent-like protocols for on-demand stored media streaming. ACM SIGMETRICS Performance Evaluation Review, 36(1):301– 312, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2008:PBP

[1671] Shao Liu, Rui Zhang-Shen, Wenjie Jiang, Jennifer Rexford, and Mung Chiang. Performance bounds for peer-assisted live streaming. ACM SIGMETRICS Performance Evaluation Review, 36(1):313–324, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2008:ELS

[1672] Thomas Bonald, Laurent Massoulié, Fabien Mathieu, Diego Perino, and Andrew Twigg. Epidemic live streaming: optimal performance trade-offs. ACM SIGMETRICS Performance Evaluation Review, 36(1):325–336, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:2008:STM

[1673] Jiang Lin, Hongzhong Zheng, Zhichun Zhu, Eugene Gorbatov, Howard David, and Zhao Zhang. Software thermal management of DRAM memory for multicore systems. ACM SIGMET-RICS Performance Evaluation Review, 36(1):337–348, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menache:2008:NPC

[1674] Ishai Menache and Nahum Shimkin. Noncooperative power control and transmission scheduling in wireless collision channels. ACM SIGMETRICS Performance Evaluation Review, 36 (1):349–358, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kandemir:2008:SDC

[1675] Mahmut Kandemir and Ozcan Ozturk. Software-directed combined CPU/link voltage scaling for NoCbased CMPs. ACM SIGMETRICS Performance Evaluation Review, 36 (1):359–370, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Crk:2008:IAE

[1676] Igor Crk, Mingsong Bi, and Chris Gniady. Interaction-aware energy management for wireless network cards. ACM SIGMETRICS Performance Evaluation Review, 36(1):371–382, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stanojevi:2008:FDE

[1677] Rade Stanojevi and Robert Shorten.

Fully decentralized emulation of besteffort and processor sharing queues. ACM SIGMETRICS Performance Evaluation Review, 36(1):383–394, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jagabathula:2008:ODS

[1678] Srikanth Jagabathula and Devavrat Shah. Optimal delay scheduling in networks with arbitrary constraints. ACM SIGMETRICS Performance Evaluation Review, 36(1):395–406, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jung:2008:RSL

[1679] Kyomin Jung, Yingdong Lu, Devavrat Shah, Mayank Sharma, and Mark S. Squillante. Revisiting stochastic loss networks: structures and algorithms. ACM SIGMETRICS Performance Evaluation Review, 36(1):407– 418, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2008:TCM

[1680] Thomas Bonald, Ali Ibrahim, and James Roberts. Traffic capacity of multi-cell WLANS. ACM SIGMET-RICS Performance Evaluation Review, 36(1):419–430, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reineke:2008:RCC

[1681] Jan Reineke and Daniel Grund. Relative competitiveness of cache replacement policies. ACM SIGMETRICS Performance Evaluation Review, 36 (1):431–432, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wen:2008:NDE

[1682] Zhihua Wen and Michael Rabinovich. Network distance estimation with dynamic landmark triangles. ACM SIGMETRICS Performance Evaluation Review, 36(1):433–434, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yuksel:2008:CSI

[1683] Murat Yuksel, Kadangode K. Ramakrishnan, Shivkumar Kalyanaraman, Joseph D. Houle, and Rita Sadhvani. Class-of-service in IP backbones: informing the network neutrality debate. ACM SIGMETRICS Performance Evaluation Review, 36 (1):435–436, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dreger:2008:PRC

[1684] Holger Dreger, Anja Feldmann, Vern Paxson, and Robin Sommer. Predicting the resource consumption of network intrusion detection systems. ACM SIGMETRICS Performance Evaluation Review, 36(1):437–438, June 2008.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2008:EMA

[1685] Bin Li, Lu Peng, and Balachandran Ramadass. Efficient MART-aided modeling for microarchitecture design space exploration and performance prediction. ACM SIGMETRICS Performance Evaluation Review, 36(1):439– 440, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Balon:2008:CII

[1686] Simon Balon and Guy Leduc. Combined intra- and inter-domain traffic engineering using hot-potato aware link weights optimization. ACM SIGMET-RICS Performance Evaluation Review, 36(1):441–442, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anderson:2008:MDW

[1687] Eric W. Anderson, Caleb T. Phillips, Kevin S. Bauer, Dirk C. Grunwald, and Douglas C. Sicker. Modeling directionality in wireless networks: extended abstract. ACM SIGMET-RICS Performance Evaluation Review, 36(1):443–444, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bremler-Barr:2008:LIC

[1688] Anat Bremler-Barr, David Hay, Danny Hendler, and Boris Farber. Layered interval codes for TCAM-based classification. ACM SIGMETRICS Performance Evaluation Review, 36(1):445– 446, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ramabhadran:2008:DRD

[1689] Sriram Ramabhadran and Joseph Pasquale. Durability of replicated distributed storage systems. ACM SIGMETRICS Performance Evaluation Review, 36(1):447–448, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2008:IEM

[1690] Feihui Li, Mahmut Kandemir, and Mary J. Irwin. Implementation and evaluation of a migration-based NUCA design for chip multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 36(1):449–450, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Alouf:2008:MGQ

[1691] Sara Alouf, Eitan Altman, and Amar Prakash Azad. M/G/1 queue with repeated inhomogeneous vacations applied to IEEE 802.16e power saving. ACM SIGMETRICS Performance Evaluation Review, 36(1):451– 452, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Seetharaman:2008:MID

[1692] Srinivasan Seetharaman and Mostafa H. Ammar. Managing inter-domain traffic in the presence of BitTorrent filesharing. ACM SIGMETRICS Performance Evaluation Review, 36(1):453– 454, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mota-Garcia:2008:COE

[1693] Edmar Mota-Garcia and Rogelio Hasimoto-Beltran. Clock offset estimation using collaborative one-way transit time. ACM SIGMETRICS Performance Evaluation Review, 36(1):455– 456, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:2008:SQL

[1694] Gagan R. Gupta and Ness B. Shroff.

Scheduling with queue length guarantees for shared resource systems. ACM SIGMETRICS Performance Evaluation Review, 36(1):457–458, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2008:ECD

[1695] Aiyou Chen, Li Li, and Jin Cao. Estimating cardinality distributions in network traffic: extended abstract. ACM SIGMETRICS Performance Evaluation Review, 36(1):459–460, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Grit:2008:WFS

[1696] Laura E. Grit and Jeffrey S. Chase.

Weighted fair sharing for dynamic virtual clusters. ACM SIGMETRICS Performance Evaluation Review, 36 (1):461–462, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sundaram:2008:ETF

[1697] Vasumathi Sundaram, Abhishek Chandra, and Jon Weissman. Exploring the throughput-fairness tradeoff of deadline scheduling in heterogeneous computing environments. ACM SIGMET-RICS Performance Evaluation Review, 36(1):463–464, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Papp:2008:CMV

[1698] Gabor Papp and Chris GauthierDickey. Characterizing multiparty voice communication for multiplayer games. ACM SIGMETRICS Performance Evaluation Review, 36(1):465–466, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Meiners:2008:AAR

[1699] Chad R. Meiners, Alex X. Liu, and Eric Torng. Algorithmic approaches to redesigning TCAM-based systems. ACM SIGMETRICS Performance Evaluation Review, 36(1):467–468, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Douceur:2008:PAR

[1700] John R. Douceur. Performance analysis in the real world. ACM SIGMET-RICS Performance Evaluation Review, 36(1):469–470, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2008:IMV

[1701] Tingxi Tan, Rob Simmonds, Bradley Arlt, Martin Arlitt, and Bruce Walker. Image management in a virtualized data center. ACM SIGMETRICS Performance Evaluation Review, 36 (2):4–9, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chandra:2008:CDF

[1702] Abhishek Chandra, Rohini Prinja, Sourabh Jain, and ZhiLi Zhang. Codesigning the failure analysis and monitoring of large-scale systems. ACM SIGMETRICS Performance Evaluation Review, 36(2):10–15, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sharma:2008:ARC

[1703] Abhishek B. Sharma, Ranjita Bhagwan, Monojit Choudhury, Leana Golubchik, Ramesh Govindan, and Geoffrey M. Voelker. Automatic request categorization in Internet services. ACM SIGMETRICS Performance Evaluation Review, 36(2):16–25, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kansal:2008:FGE

[1704] Aman Kansal and Feng Zhao. Finegrained energy profiling for poweraware application design. ACM SIGMETRICS Performance Evaluation Review, 36(2):26–31, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fonseca:2008:LRM

[1705] Nahur Fonseca, Mark Crovella, and Kavé Salamatian. Long range mutual information. ACM SIGMET-RICS Performance Evaluation Review, 36(2):32–37, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2008:HPM

[1706] Giuliano Casale, Ningfang Mi, Ludmila Cherkasova, and Evgenia Smirni. How to parameterize models with bursty workloads. ACM SIGMETRICS Performance Evaluation Review, 36(2): 38–44, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:2008:DPF

[1707] Bill Lin and Jun (Jim) Xu. DRAM is plenty fast for wirespeed statistics counting. ACM SIGMETRICS Performance Evaluation Review, 36 (2):45–51, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Agrawal:2008:TRF

[1708] Nitin Agrawal, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Towards realistic file-system benchmarks with CodeMRI. ACM SIGMETRICS Performance Evaluation Review, 36(2):52–57, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Weingartner:2008:SNE

[1709] Elias Weingärtner, Florian Schmidt, Tobias Heer, and Klaus Wehrle. Synchronized network emulation: matching prototypes with complex simulations. ACM SIGMETRICS Performance Evaluation Review, 36(2):58–63, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krishnamurthy:2008:WOS

[1710] Balachander Krishnamurthy and Walter Willinger. What are our standards for validation of measurement-based networking research? ACM SIGMET-RICS Performance Evaluation Review, 36(2):64–69, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Curry:2008:RAE

[1711] Roger Curry, Cameron Kiddle, Nayden Markatchev, Rob Simmonds, Tingxi Tan, Martin Arlitt, and Bruce Walker. Running applications efficiently in online social networks. ACM SIGMET-RICS Performance Evaluation Review, 36(2):71–74, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2008:KTB

[1712] Eddy Zheng Zhang, Giuliano Casale, and Evgenia Smirni. KPC-Toolbox: best recipes toward automatization of workload fitting. ACM SIGMETRICS Performance Evaluation Review, 36 (2):75–78, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

DeVera:2008:AQE

[1713] Daniel De Vera, Pablo Rodríguez-Bocca, and Gerardo Rubino. Automatic quality of experience measuring on video delivering networks. ACM SIGMETRICS Performance Evaluation Review, 36(2):79–82, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rossi:2008:PS

[1714] Dario Rossi, Silvio Valenti, Paolo Veglia, Dario Bonfiglio, Marco Mellia, and Michela Meo. Pictures from the Skype. ACM SIGMETRICS Performance Evaluation Review, 36(2):83–86, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ormont:2008:CMW

[1715] Justin Ormont, Jordan Walker, and Suman Banerjee. Continuous monitoring of wide-area wireless networks: data collection and visualization. ACM SIGMETRICS Performance Evaluation Review, 36(2):87–89, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anouar:2008:OOW

[1716] Hicham Anouar, Christian Bonnet, Daniel Câmara, Fethi Filali, and Raymond Knopp. An overview of OpenAirInterface wireless network emulation methodology. ACM SIGMET-RICS Performance Evaluation Review, 36(2):90–94, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jiang:2008:NPN

[1717] Xiaoyue Jiang. New perspectives on network calculus. ACM SIGMET-RICS Performance Evaluation Review, 36(2):95–97, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Garikiparthi:2008:BPA

[1718] Chaitanya Garikiparthi, Appie van de Liefvoort, and Ken Mitchell. Busy period analysis of finite QBD processes. ACM SIGMETRICS Performance Evaluation Review, 36(2): 98–100, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jelenkovic:2008:FRS

[1719] Predrag R. Jelenković and Xiaozhu Kang. Is fair resource sharing responsible for spreading long delays? ACM SIGMETRICS Performance Evaluation Review, 36(2):101–103, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:2008:FOQ

[1720] Varun Gupta. Finding the optimal quantum size: Sensitivity analysis of the m/g/1 round-robin queue. ACM SIGMETRICS Performance Evaluation Review, 36(2):104–106, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bachmat:2008:ASI

[1721] Eitan Bachmat and Hagit Sarfati.

Analysis of size interval task assignment policies. ACM SIGMETRICS Performance Evaluation Review, 36(2): 107–109, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2008:ELS

[1722] Ho-Lin Chen, Jason R. Marden, and Adam Wierman. The effect of local scheduling in load balancing designs. ACM SIGMETRICS Performance Evaluation Review, 36(2):110– 112, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wu:2008:JRP

[1723] Yuan Wu and Danny H. K. Tsang.

Joint rate-and-power allocation for multi-channel spectrum sharing networks. ACM SIGMETRICS Performance Evaluation Review, 36(2):

113–115, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2008:SMB

[1724] Pei Li, John C. S. Lui, and Yinlong Xu. A stochastic model for BitTorrentlike systems. ACM SIGMETRICS Performance Evaluation Review, 36(2): 116–118, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jelenkovic:2008:CMS

[1725] Predrag R. Jelenković and Xiaozhu Kang. Characterizing the miss sequence of the LRU cache. ACM SIGMETRICS Performance Evaluation Review, 36(2):119–121, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Simatos:2008:SSM

[1726] Florian Simatos and Danielle Tibi. Study of a stochastic model for mobile networks. ACM SIGMETRICS Performance Evaluation Review, 36(2): 122–124, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Momcilovic:2008:TSL

[1727] Petar Momčilović and Mark S. Squillante. On throughput in stochastic linear loss networks. ACM SIGMET-RICS Performance Evaluation Review, 36(2):125–127, September 2008. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:2008:FLR

[1728] Varun Gupta and Peter G. Harrison. Fluid level in a reservoir with an on-off source. ACM SIGMETRICS Performance Evaluation Review, 36(2): 128–130, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kwak:2008:SAS

[1729] K. J. Kwak, Y. M. Baryshnikov, and E. G. Coffman. Self-assembling sweep-and-sleep sensor systems. ACM SIGMETRICS Performance Evaluation Review, 36(2):131–133, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2008:CCO

[1730] Giuliano Casale. CoMoM: classoriented evaluation of multiclass models. ACM SIGMETRICS Performance Evaluation Review, 36(2):134– 136, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dieker:2008:COF

[1731] A. B. Dieker, S. Ghosh, and M. S.

Squillante. Capacity optimization in feedforward Brownian networks. ACM SIGMETRICS Performance Evaluation Review, 36(2):137–139, September 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Haverkort:2008:QAG

[1732] Boudewijn R. Haverkort, Markus Siegle, and Maarten van Steen. Quantitative analysis of gossiping protocols. ACM SIGMETRICS Performance Evaluation Review, 36(3):2, December 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Katoen:2008:HMA

[1733] Joost-Pieter Katoen. How to model and analyze gossiping protocols? ACM SIGMETRICS Performance Evaluation Review, 36(3):3–6, December 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Crouzen:2008:AFM

[1734] Pepijn Crouzen, Jaco van de Pol, and Arend Rensink. Applying formal methods to gossiping networks with mCRL and groove. ACM SIGMET-RICS Performance Evaluation Review, 36(3):7–16, December 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kwiatkowska:2008:AGP

[1735] Marta Kwiatkowska, Gethin Norman, and David Parker. Analysis of a gossip protocol in PRISM. ACM SIGMET-RICS Performance Evaluation Review, 36(3):17–22, December 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krieger:2008:VPM

[1736] Thomas Krieger, Martin Riedl, Johann Schuster, and Markus Siegle. A viewprobability-matrix approach to the modelling of gossiping protocols. ACM SIGMETRICS Performance Evaluation Review, 36(3):23–30, December 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bakhshi:2008:MAE

[1737] Rena Bakhshi, Lucia Cloth, Wan Fokkink, and Boudewijn R. Haverkort. MeanField analysis for the evaluation of gossip protocols. ACM SIGMET-RICS Performance Evaluation Review, 36(3):31–39, December 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Estrada:2008:DEM

[1738] Trilce Estrada, Olac Fuentes, and Michela Taufer. A distributed evolutionary method to design scheduling policies for volunteer computing. ACM SIGMETRICS Performance Evaluation Review, 36(3):40–49, December 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Eddy:2008:BPI

[1739] Wesley M. Eddy. Basic properties of the IPv6 AS-level topology. ACM SIGMETRICS Performance Evaluation Review, 36(3):50–57, December 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2009:SIT

[1740] Giuliano Casale, Richard R. Muntz, and Giuseppe Serazzi. Special issue on tools for computer performance modeling and reliability analysis. ACM SIGMETRICS Performance Evaluation Review, 36(4):2–3, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baarir:2009:GTR

[1741] Soheib Baarir, Marco Beccuti, Davide Cerotti, Massimiliano De Pierro, Susanna Donatelli, and Giuliana Franceschinis. The GreatSPN tool: recent enhancements. ACM SIGMETRICS Performance Evaluation Review, 36(4):4– 9, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bertoli:2009:JPE

[1742] Marco Bertoli, Giuliano Casale, and Giuseppe Serazzi. JMT: performance engineering tools for system modeling. ACM SIGMETRICS Performance Evaluation Review, 36(4):10–15, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gaonkar:2009:PDM

[1743] Shravan Gaonkar, Ken Keefe, Ruth Lamprecht, Eric Rozier, Peter Kemper, and William H. Sanders. Performance and dependability modeling with Möbius. ACM SIGMET-RICS Performance Evaluation Review, 36(4):16–21, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arns:2009:OTO

[1744] Markus Arns, Peter Buchholz, and Dennis Müller. OPEDo: a tool for the optimization of performance and dependability models. ACM SIGMET-RICS Performance Evaluation Review, 36(4):22–27, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tribastone:2009:PEP

[1745] Mirco Tribastone, Adam Duguid, and Stephen Gilmore. The PEPA Eclipse plugin. ACM SIGMETRICS Performance Evaluation Review, 36(4):28–33, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dingle:2009:PTP

[1746] Nicholas J. Dingle, William J. Knottenbelt, and Tamas Suto. PIPE2: a tool for the performance evaluation of generalised stochastic Petri Nets. ACM SIGMETRICS Performance Evaluation Review, 36(4):34–39, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kwiatkowska:2009:PPM

[1747] Marta Kwiatkowska, Gethin Norman, and David Parker. PRISM: probabilistic model checking for performance and reliability analysis. ACM SIGMET-RICS Performance Evaluation Review, 36(4):40–45, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kounev:2009:QPM

[1748] Samuel Kounev and Christofer Dutz.

QPME: a performance modeling tool based on queueing Petri Nets. ACM SIGMETRICS Performance Evaluation Review, 36(4):46–51, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Trivedi:2009:SAT

[1749] Kisho S. Trivedi and Robin Sahner. SHARPE at the age of twenty two. ACM SIGMETRICS Performance Evaluation Review, 36(4):52–57, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ciardo:2009:AFS

[1750] Gianfranco Ciardo, Andrew S. Miner, and Min Wan. Advanced features in SMART: the stochastic model checking analyzer for reliability and timing. ACM SIGMETRICS Performance Evaluation Review, 36(4):58–63, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

deSouzaeSilva:2009:TIM

[1751] Edmundo de Souza e Silva, Daniel R. Figueiredo, and Rosa M. M. Leão. The TANGRAMII integrated modeling environment for computer systems and networks. ACM SIGMETRICS Performance Evaluation Review, 36(4):64–69, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lelarge:2009:ECE

[1752] Marc Lelarge. Efficient control of epidemics over random networks. ACM SIGMETRICS Performance Evaluation Review, 37(1):1–12, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/2492101.
1555351.

Pathak:2009:BSC

[1753] Abhinav Pathak, Feng Qian, Y. Charlie Hu, Z. Morley Mao, and Supranamaya Ranjan. Botnet spam campaigns can be long lasting: evidence, implications, and analysis. ACM SIGMETRICS Performance Evaluation Review, 37(1): 13–24, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/2492101.1555352.

Torres:2009:IUB

[1754] Ruben D. Torres, Mohammad Y. Hajjat, Sanjay G. Rao, Marco Mellia, and Maurizio M. Munafo. Inferring undesirable behavior from P2P traffic analysis. ACM SIGMETRICS Performance Evaluation Review, 37(1):25– 36, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/2492101.1555353.

Anand:2009:RNT

[1755] Ashok Anand, Chitra Muthukrishnan, Aditya Akella, and Ramachandran Ramjee. Redundancy in network traffic: findings and implications. ACM SIGMETRICS Performance Evaluation Review, 37(1):37–48, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/2492101.1555355.

Jin:2009:UCN

[1756] Yu Jin, Esam Sharafuddin, and Zhi-Li Zhang. Unveiling core network-wide communication patterns through application traffic activity graph decomposition. ACM SIGMETRICS Performance Evaluation Review, 37(1):49– 60, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/2492101.1555356.

Ramasubramanian:2009:TIL

[1757] Venugopalan Ramasubramanian, Dahlia Malkhi, Fabian Kuhn, Mahesh Balakrishnan, Archit Gupta, and Aditya Akella. On the treeness of Internet latency and bandwidth. ACM SIGMETRICS Performance Evaluation Review, 37(1):61–72, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/2492101.
1555357.

Meiners:2009:TTA

[1758] Chad R. Meiners, Alex X. Liu, and Eric Torng. Topological transformation approaches to optimizing TCAM-based packet classification systems. ACM SIGMETRICS Performance Evaluation Review, 37(1):73–84, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/2492101.
                                        1555359.

Shen:2009:RDP

[1759] Kai Shen, Christopher Stewart, Chuanpeng Li, and Xin Li. Referencedriven performance anomaly identification. ACM SIGMETRICS Performance Evaluation Review, 37(1):85–96, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/2492101.1555360.

Gupta:2009:NWS

[1760] Gagan R. Gupta, Sujay Sanghavi, and Ness B. Shroff. Node weighted scheduling. ACM SIGMETRICS Performance Evaluation Review, 37(1):97– 108, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/2492101.1555361.

Chaintreau:2009:AGS

[1761] Augustin Chaintreau, Jean-Yves Le Boudec, and Nikodin Ristanovic. The age of gossip: spatial mean field regime. ACM SIGMETRICS Performance Evaluation Review, 37(1):109– 120, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/2492101.1555363.

Bodas:2009:SMC

[1762] Shreeshankar Bodas, Sanjay Shakkottai, Lei Ying, and R. Srikant. Scheduling in multi-channel wireless networks: rate function optimality in the smallbuffer regime. ACM SIGMETRICS Performance Evaluation Review, 37(1): 121–132, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/2492101.1555364.

Rajagopalan:2009:NAT

[1763] Shreevatsa Rajagopalan, Devavrat Shah, and Jinwoo Shin. Network adiabatic theorem: an efficient randomized protocol for contention resolution. ACM SIGMETRICS Performance Evaluation Review, 37(1):133– 144, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/2492101.1555365.

Sharma:2009:DDC

[1764] Abhishek B. Sharma, Leana Golubchik, Ramesh Govindan, and Michael J. Neely. Dynamic data compression in multi-hop wireless networks. ACM SIGMETRICS Performance Evaluation Review, 37(1):145– 156, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/2492101.1555367.

Gandhi:2009:OPA

[1765] Anshul Gandhi, Mor Harchol-Balter, Rajarshi Das, and Charles Lefurgy. Optimal power allocation in server farms. ACM SIGMETRICS Performance Evaluation Review, 37(1):157– 168, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/2492101.1555368.

Coskun:2009:EIJ

[1766] Ayse K. Coskun, Richard Strong, Dean M. Tullsen, and Tajana Simunic Rosing. Evaluating the impact of job scheduling and power management on processor lifetime for chip multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 37(1):169– 180, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/2492101.1555369.

Chen:2009:UIC

[1767] Feng Chen, David A. Koufaty, and Xiaodong Zhang. Understanding intrinsic characteristics and system implications of flash memory based solid state drives. ACM SIGMETRICS Performance Evaluation Review, 37(1):181– 192, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/2492101.1555371.

Schroeder:2009:DEW

[1768] Bianca Schroeder, Eduardo Pinheiro, and Wolf-Dietrich Weber. DRAM errors in the wild: a large-scale field study. ACM SIGMETRICS Performance Evaluation Review, 37(1):193– 204, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/2492101.1555372.

Mi:2009:RUI

[1769] Ningfang Mi, Alma Riska, Xin Li, Evgenia Smirni, and Erik Riedel. Restrained utilization of idleness for transparent scheduling of background tasks. ACM SIGMETRICS Performance Evaluation Review, 37(1):205– 216, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/2492101.1555373.

Wang:2009:NSB

[1770] Yi Wang, Michael Schapira, and Jennifer Rexford. Neighbor-specific BGP: more flexible routing policies while improving global stability. ACM SIG-METRICS Performance Evaluation Review, 37(1):217–228, June 2009. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/2492101.
1555375.

Laoutaris:2009:DTB

[1771] Nikolaos Laoutaris, Georgios Smaragdakis, Pablo Rodriguez, and Ravi Sundaram. Delay tolerant bulk data transfers on the Internet. ACM SIGMET-RICS Performance Evaluation Review, 37(1):229–238, June 2009. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/2492101.
1555376.

Jiang:2009:CCD

[1772] Wenjie Jiang, Rui Zhang-Shen, Jennifer Rexford, and Mung Chiang. Cooperative content distribution and traffic engineering in an ISP network. ACM SIGMETRICS Performance Evaluation Review, 37(1):239– 250, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/2492101.1555377.

Cohen:2009:LDS

[1773] Edith Cohen and Haim Kaplan. Leveraging discarded samples for tighter estimation of multiple-set aggregates. ACM SIGMETRICS Performance Evaluation Review, 37(1):251–262, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/2492101.1555379.

Loiseau:2009:MLE

[1774] Patrick Loiseau, Paulo Gonçalves, Stéphane Girard, Florence Forbes, and Pascale Vicat-Blanc Primet. Maximum likelihood estimation of the flow size distribution tail index from sampled packet data. ACM SIGMETRICS Performance Evaluation Review, 37(1): 263–274, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/2492101.1555380.

Qiu:2009:MCP

[1775] Tongqing Qiu, Zihui Ge, Seungjoon Lee, Jia Wang, Qi Zhao, and Jun Xu. Modeling channel popularity dynamics in a large IPTV system. ACM SIGMETRICS Performance Evaluation Review, 37(1):275– 286, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/2492101.1555381.

Harchol-Balter:2009:SRT

[1776] Mor Harchol-Balter, Alan Scheller-Wolf, and Andrew R. Young. Surprising results on task assignment in server farms with high-variability workloads. ACM SIGMETRICS Performance Evaluation Review, 37(1):287– 298, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/2492101.1555383.

Sandholm:2009:MOU

[1777] Thomas Sandholm and Kevin Lai.

MapReduce optimization using regulated dynamic prioritization. ACM SIGMETRICS Performance Evaluation Review, 37(1):299–310, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/2492101.
1555384.

Gupta:2009:SAA

[1778] Varun Gupta and Mor Harchol-Balter.

Self-adaptive admission control policies for resource-sharing systems. ACM SIGMETRICS Performance Evaluation Review, 37(1):311–322, June 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/2492101.
1555385.

Scheuermann:2009:WSS

[1779] Björn Scheuermann and Wolfgang Kiess. Who said that?: the sendreceive correlation problem in network log analysis. ACM SIGMET-RICS Performance Evaluation Review, 37(2):3–5, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anandkumar:2009:SRM

[1780] Animashree Anandkumar, Chatschik Bisdikian, Ting He, and Dakshi Agrawal. Selectively retrofitting monitoring in distributed systems. ACM SIGMETRICS Performance Evaluation Review, 37(2):6–8, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dubey:2009:PMD

[1781] Abhishek Dubey, Rajat Mehrotra, Sherif Abdelwahed, and Asser Tantawi. Performance modeling of distributed multi-tier enterprise systems. ACM SIGMETRICS Performance Evaluation Review, 37(2):9–11, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2009:DCR

[1782] Chao Wang and Xiaoli Ma. Deriving Cramér–Rao bounds and maximum likelihood estimators for traffic matrix inference. ACM SIGMET-RICS Performance Evaluation Review, 37(2):12–14, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krioukov:2009:GFS

[1783] Dmitri Krioukov, Fragkiskos Papadopoulos, Marián Boguñá, and Amin Vahdat. Greedy forwarding in scalefree networks embedded in hyperbolic metric spaces. ACM SIGMETRICS Performance Evaluation Review, 37 (2):15–17, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cho:2009:BTB

[1784] Jeong woo Cho and Yuming Jiang. Basic theorems on the backoff process in 802.11. ACM SIGMETRICS Performance Evaluation Review, 37 (2):18–20, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nair:2009:OJF

[1785] Jayakrishnan Nair and Steven H. Low. Optimal job fragmentation. ACM SIGMETRICS Performance Evaluation Review, 37(2):21–23, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yao:2009:EAL

[1786] Erlin Yao, Yungang Bao, Guangming Tan, and Mingyu Chen. Extending Amdahl’s Law in the multicore era. ACM SIGMETRICS Performance Evaluation Review, 37(2):24–26, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Korzun:2009:LEM

[1787] Dmitry Korzun and Andrei Gurtov.

A local equilibrium model for P2P resource ranking. ACM SIGMET-RICS Performance Evaluation Review, 37(2):27–29, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menasche:2009:MCAa

[1788] Daniel Sadoc Menasché, Antonio A. Aragão Rocha, Edmundo de Souza e Silva, Rosa M. Meri Leão, Don Towsley, and Arun Venkataramani. Modeling chunk availability in P2P swarming systems. ACM SIGMETRICS Performance Evaluation Review, 37(2):30–32, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hohlfeld:2009:VIV

[1789] Oliver Hohlfeld and Florin Ciucu.

Viewing impaired video transmissions from a modeling perspective. ACM SIGMETRICS Performance Evaluation Review, 37(2):33–35, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:2009:WOS

[1790] Gagan R. Gupta, Sujay Sanghavi, and Ness B. Shroff. Workload optimality in switches without arrivals. ACM SIGMETRICS Performance Evaluation Review, 37(2):36–38, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Andrew:2009:OSS

[1791] Lachlan L. H. Andrew, Adam Wierman, and Ao Tang. Optimal speed scaling under arbitrary power functions. ACM SIGMETRICS Performance Evaluation Review, 37(2):39–41, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Verloop:2009:HTA

[1792] I. M. Verloop, U. Ayesta, and R. Núñez-Queija. Heavy-traffic analysis of the M/PH/1 discriminatory processor sharing queue with phasedependent weights. ACM SIGMET-RICS Performance Evaluation Review, 37(2):42–44, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anselmi:2009:IAS

[1793] J. Anselmi, Y. Lu, M. Sharma, and M. S. Squillante. Improved approximations for stochastic loss networks. ACM SIGMETRICS Performance Evaluation Review, 37(2):45–47, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Weingartner:2009:TAI

[1794] Elias Weingärtner, Florian Schmidt, Tobias Heer, and Klaus Wehrle. Time accurate integration of software prototypes with event-based network simulations. ACM SIGMETRICS Performance Evaluation Review, 37(2): 49–50, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2009:ETC

[1795] Haifeng Chen, Wenxuan Zhang, and Guofei Jiang. Experience transfer for the configuration tuning in large scale computing systems. ACM SIGMET-RICS Performance Evaluation Review, 37(2):51–52, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:2009:RID

[1796] Bill Lin, Jun (Jim) Xu, Nan Hua, Hao Wang, and Haiquan (Chuck) Zhao. A randomized interleaved DRAM architecture for the maintenance of exact statistics counters. ACM SIGMET-RICS Performance Evaluation Review, 37(2):53–54, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhao:2009:MPA

[1797] Bridge Zhao, y. K. Li, John C. S.

Lui, and Dah-Ming Chiu. On modeling product advertisement in social networks. ACM SIGMETRICS Performance Evaluation Review, 37(2): 55–56, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zahn:2009:ESF

[1798] Thomas Zahn, Greg O’Shea, and Antony Rowstron. An empirical study of flooding in mesh networks. ACM SIGMETRICS Performance Evaluation Review, 37(2):57–58, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Triukose:2009:CDN

[1799] Sipat Triukose, Zhihua Wen, and Michael Rabinovich. Content delivery networks: how big is big enough? ACM SIGMETRICS Performance Evaluation Review, 37(2):59–60, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yu:2009:SFM

[1800] Zhibin Yu and Hai Jin. Simple and fast micro-architecture simulation: a trisection Cantor fractal approach. ACM SIGMETRICS Performance Evaluation Review, 37(2):61–62, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Key:2009:RGE

[1801] Peter Key and Alexandre Proutiere. Routing games with elastic traffic. ACM SIGMETRICS Performance Evaluation Review, 37(2):63–64, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lange:2009:ESI

[1802] John R. Lange, J. Scott Miller, and Peter A. Dinda. EmNet: satisfying the individual user through empathic home networks: summary. ACM SIGMET-RICS Performance Evaluation Review, 37(2):65–66, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Riska:2009:EDL

[1803] Alma Riska and Erik Riedel. Evaluation of disk-level workloads at different time scales. ACM SIGMET-RICS Performance Evaluation Review, 37(2):67–68, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reddy:2009:MDC

[1804] Vinith Reddy, Younghoon Kim, Srinivas Shakkottai, and A. L. Narasimha Reddy. MultiTrack: a delay and cost aware P2P overlay architecture. ACM SIGMETRICS Performance Evaluation Review, 37(2):69–70, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Borst:2009:SOA

[1805] Sem Borst, Varun Gupta, and Anwar Walid. Self-organizing algorithms for cache cooperation in content distribution networks. ACM SIGMET-RICS Performance Evaluation Review, 37(2):71–72, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rubinstein:2009:SPA

[1806] Benjamin I. P. Rubinstein, Blaine Nelson, Ling Huang, Anthony D. Joseph, Shing hon Lau, Satish Rao, Nina Taft, and J. D. Tygar. Stealthy poisoning attacks on PCA-based anomaly detectors. ACM SIGMETRICS Performance Evaluation Review, 37(2): 73–74, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Down:2009:SDR

[1807] Douglas G. Down, H. Christian Gromoll, and Amber L. Puha. Statedependent response times via fluid limits in shortest remaining processing time queues. ACM SIGMET-RICS Performance Evaluation Review, 37(2):75–76, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2009:SPP

[1808] Jianwei Chen, Murali Annavaram, and Michel Dubois. SlackSim: a platform for parallel simulations of CMPs on CMPs. ACM SIGMETRICS Performance Evaluation Review, 37(2): 77–78, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gulati:2009:EAP

[1809] Ajay Gulati, Arif Merchant, Mustafa Uysal, Pradeep Padala, and Peter Varman. Efficient and adaptive proportional share I/O scheduling. ACM SIGMETRICS Performance Evaluation Review, 37(2):79–80, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2009:DDS

[1810] Yang Liu, Linfeng Zhang, and Yong Guan. A distributed data streaming algorithm for network-wide traffic anomaly detection. ACM SIGMET-RICS Performance Evaluation Review, 37(2):81–82, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baccelli:2009:TMA

[1811] François Baccelli, Bruno Kauffmann, and Darryl Veitch. Towards multihop available bandwidth estimation. ACM SIGMETRICS Performance Evaluation Review, 37(2):83–84, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nandi:2009:WMU

[1812] Animesh Nandi, Bobby Bhattacharjee, and Peter Druschel. What a mesh: understanding the design tradeoffs for streaming multicast. ACM SIGMET-RICS Performance Evaluation Review, 37(2):85–86, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menasche:2009:MCAb

[1813] Daniel Sadoc Menasche, Antonio A. Aragao Rocha, Bin Li, Don Towsley, and Arun Venkataramani. Modeling content availability in peer-to-peer swarming systems. ACM SIGMET-RICS Performance Evaluation Review, 37(2):87–88, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Iyer:2009:VPA

[1814] Ravi Iyer, Ramesh Illikkal, Li Zhao, Don Newell, and Jaideep Moses. Virtual platform architectures for resource metering in datacenters. ACM SIGMETRICS Performance Evaluation Review, 37(2):89–90, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kant:2009:CDE

[1815] Krishna Kant. Challenges in distributed energy adaptive computing. ACM SIGMETRICS Performance Evaluation Review, 37(3):3–7, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pan:2009:GBB

[1816] Xinghao Pan, Jiaqi Tan, Soila Kavulya, Rajeev Gandhi, and Priya Narasimhan. Ganesha: black-box diagnosis of MapReduce systems. ACM SIGMET-RICS Performance Evaluation Review, 37(3):8–13, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anand:2009:NNN

[1817] Ashok Anand and Aditya Akella. NetReplay: a new network primitive. ACM SIGMETRICS Performance Evaluation Review, 37(3):14–19, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Crocey:2009:QBE

[1818] Daniele Crocey, Marco Melliay, and Emilio Leonardiy. The quest for bandwidth estimation techniques for large-scale distributed systems. ACM SIGMETRICS Performance Evaluation Review, 37(3):20–25, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Keeton:2009:DYK

[1819] Kimberly Keeton, Pankaj Mehra, and John Wilkes. Do you know your IQ?: a research agenda for information quality in systems. ACM SIGMET-RICS Performance Evaluation Review, 37(3):26–31, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2009:AGB

[1820] Giuliano Casale, Amir Kalbasi, Diwakar Krishnamurthy, and Jerry Rolia. Automatically generating bursty benchmarks for multitier systems. ACM SIGMETRICS Performance Evaluation Review, 37(3):32–37, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hellerstein:2009:ACT

[1821] Joseph L. Hellerstein, Vance Morrison, and Eric Eilebrecht. Applying control theory in the real world: experience with building a controller for the .NET thread pool. ACM SIGMET-RICS Performance Evaluation Review, 37(3):38–42, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Riska:2009:FRE

[1822] Alma Riska, Ningfang Mi, Evgenia Smirni, and Giuliano Casale. Feasibility regions: exploiting tradeoffs between power and performance in disk drives. ACM SIGMETRICS Performance Evaluation Review, 37(3):43–48, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Willinger:2009:ROS

[1823] Walter Willinger, Reza Rejaie, Mojtaba Torkjazi, Masoud Valafar, and Mauro Maggioni. Research on online social networks: time to face the real challenges. ACM SIGMET-RICS Performance Evaluation Review, 37(3):49–54, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tickoo:2009:MVM

[1824] Omesh Tickoo, Ravi Iyer, Ramesh Illikkal, and Don Newell. Modeling virtual machine performance: challenges and approaches. ACM SIGMETRICS Performance Evaluation Review, 37 (3):55–60, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gulati:2009:MWD

[1825] Ajay Gulati, Chethan Kumar, and Irfan Ahmad. Modeling workloads and devices for IO load balancing in virtualized environments. ACM SIGMET-RICS Performance Evaluation Review, 37(3):61–66, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fay:2009:WSM

[1826] Damien Fay, Hamed Haddadi, Andrew W. Moore, Richard Mortier, Steve Uhlig, and Almerima Jamakovic. A weighted spectrum metric for comparison of Internet topologies. ACM SIGMETRICS Performance Evaluation Review, 37(3):67–72, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Illikkal:2010:PQP

[1827] Ramesh Illikkal, Vineet Chadha, Andrew Herdrich, Ravi Iyer, and Donald Newell. PIRATE: QoS and performance management in CMP architectures. ACM SIGMETRICS Performance Evaluation Review, 37(4):3–10, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dube:2010:PLL

[1828] Parijat Dube, Li Zhang, David Daly, and Alan Bivens. Performance of large low-associativity caches. ACM SIGMETRICS Performance Evaluation Review, 37(4):11–18, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhu:2010:ROW

[1829] Yaping Zhu, Jennifer Rexford, Subhabrata Sen, and Aman Shaikh. Route Oracle: where have all the packets gone? ACM SIGMETRICS Performance Evaluation Review, 37(4):19–25, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Doebel:2010:TVP

[1830] Bjoern Doebel, Peter Nobel, Eno Thereska, and Alice Zheng. Towards versatile performance models for complex, popular applications. ACM SIGMETRICS Performance Evaluation Review, 37(4):26–33, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mishra:2010:TCC

[1831] Asit K. Mishra, Joseph L. Hellerstein, Walfredo Cirne, and Chita R. Das. Towards characterizing cloud backend workloads: insights from Google compute clusters. ACM SIGMET-RICS Performance Evaluation Review, 37(4):34–41, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arlitt:2010:SIQ

[1832] Martin Arlitt, Keith Farkas, Subu Iyer, Preethi Kumaresan, and Sandro Rafaeli. Systematically improving the quality of IT utilization data. ACM SIGMETRICS Performance Evaluation Review, 37(4):42–49, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hu:2010:PMI

[1833] Jianying Hu, Yingdong Lu, Aleksandra Mojsilović, Mayank Sharma, and Mark S. Squillante. Performance management of IT services delivery. ACM SIGMETRICS Performance Evaluation Review, 37(4):50–57, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2010:BPI

[1834] Shuyi Chen, Kaustubh R. Joshi, Matti A. Hiltunen, Richard D. Schlichting, and William H. Sanders. Blackbox prediction of the impact of DVFS on end-to-end performance of multitier systems. ACM SIGMETRICS Performance Evaluation Review, 37(4):59–63, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marwah:2010:QSI

[1835] Manish Marwah, Paulo Maciel, Amip Shah, Ratnesh Sharma, Tom Christian, Virgilio Almeida, Carlos Araújo, Erica Souza, Gustavo Callou, Bruno Silva, Sérgio Galdino, and Jose Pires. Quantifying the sustainability impact of data center availability. ACM SIGMET-RICS Performance Evaluation Review, 37(4):64–68, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marsan:2010:EEM

[1836] Marco Ajmone Marsan and Michela Meo. Energy efficient management of two cellular access networks. ACM SIGMETRICS Performance Evaluation Review, 37(4):69–73, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tsiaflakis:2010:FGD

[1837] Paschalis Tsiaflakis, Yung Yi, Mung Chiang, and Marc Moonen. Fair greening for DSL broadband access. ACM SIGMETRICS Performance Evaluation Review, 37(4):74–78, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ord:2010:PEM

[1838] Jason Ord, Ellen Chappell, Scott Canonico, and Tim Strecker. Product environmental metrics for printers. ACM SIGMETRICS Performance Evaluation Review, 37(4):79–83, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cayzer:2010:SHI

[1839] Steve Cayzer and Chris Preist. The sustainability hub: an information management tool for analysis and decision making. ACM SIGMETRICS Performance Evaluation Review, 37(4):84–88, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thereska:2010:PPM

[1840] Eno Thereska, Bjoern Doebel, Alice X.

Zheng, and Peter Nobel. Practical performance models for complex, popular applications. ACM SIGMETRICS Performance Evaluation Review, 38(1):1– 12, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gast:2010:MFM

[1841] Nicolas Gast and Gaujal Bruno. A mean field model of work stealing in large-scale systems. ACM SIGMET-RICS Performance Evaluation Review, 38(1):13–24, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Balsamo:2010:UAP

[1842] Simonetta Balsamo, Peter G. Harrison, and Andrea Marin. A unifying approach to product-forms in networks with finite capacity constraints. ACM SIGMETRICS Performance Evaluation Review, 38(1):25–36, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Andrew:2010:OFR

[1843] Lachlan L. H. Andrew, Minghong Lin, and Adam Wierman. Optimality, fairness, and robustness in speed scaling designs. ACM SIGMETRICS Performance Evaluation Review, 38(1):37–48, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dong:2010:EEE

[1844] Wei Dong, Yunhao Liu, Xiaofan Wu, Lin Gu, and Chun Chen. Elon: enabling efficient and long-term reprogramming for wireless sensor networks. ACM SIGMETRICS Performance Evaluation Review, 38(1):49–60, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Karbasi:2010:DSN

[1845] Amin Karbasi and Sewoong Oh. Distributed sensor network localization from local connectivity: performance analysis for the HOP-TERRAIN algorithm. ACM SIGMETRICS Performance Evaluation Review, 38(1):61–70, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xu:2010:SSP

[1846] Kuang Xu, Olivier Dousse, and Patrick Thiran. Self-synchronizing properties of CSMA wireless multi-hop networks. ACM SIGMETRICS Performance Evaluation Review, 38(1):71–82, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Moallemi:2010:FLD

[1847] Ciamac Moallemi and Devavrat Shah.

On the flow-level dynamics of a packetswitched network. ACM SIGMET-RICS Performance Evaluation Review, 38(1):83–94, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Godfrey:2010:ICD

[1848] P. Brighten Godfrey, Michael Schapira, Aviv Zohar, and Scott Shenker. Incentive compatibility and dynamics of congestion control. ACM SIGMET-RICS Performance Evaluation Review, 38(1):95–106, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shah:2010:DCG

[1849] Devavrat Shah and Jinwoo Shin. Dynamics in congestion games. ACM SIGMETRICS Performance Evaluation Review, 38(1):107–118, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xiang:2010:ORS

[1850] Liping Xiang, Yinlong Xu, John C. S. Lui, and Qian Chang. Optimal recovery of single disk failure in RDP code storage systems. ACM SIGMET-RICS Performance Evaluation Review, 38(1):119–130, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghanbari:2010:QLR

[1851] Saeed Ghanbari, Gokul Soundararajan, and Cristiana Amza. A query language and runtime tool for evaluating behavior of multi-tier servers. ACM SIGMETRICS Performance Evaluation Review, 38(1):131–142, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goel:2010:SSQ

[1852] Ashish Goel and Pankaj Gupta. Small subset queries and bloom filters using ternary associative memories, with applications. ACM SIGMETRICS Performance Evaluation Review, 38 (1):143–154, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Laadan:2010:TLA

[1853] Oren Laadan, Nicolas Viennot, and Jason Nieh. Transparent, lightweight application execution replay on commodity multiprocessor operating systems. ACM SIGMETRICS Performance Evaluation Review, 38(1):155– 166, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ni:2010:CSP

[1854] Jian Ni, R. Srikant, and Xinzhou Wu.

Coloring spatial point processes with applications to peer discovery in large wireless networks. ACM SIGMET-RICS Performance Evaluation Review, 38(1):167–178, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

vandeVen:2010:OTB

[1855] Peter M. van de Ven, Augustus J.

E. M. Janssen, and Johan S. H. van Leeuwaarden. Optimal tradeoff between exposed and hidden nodes in large wireless networks. ACM SIGMETRICS Performance Evaluation Review, 38(1):179–190, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2010:SMW

[1856] Shihuan Liu, Lei Ying, and R. Srikant.

Scheduling in multichannel wireless networks with flow-level dynamics. ACM SIGMETRICS Performance Evaluation Review, 38(1):191– 202, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shah:2010:DSC

[1857] Devavrat Shah and Tauhid Zaman.

Detecting sources of computer viruses in networks: theory and experiment. ACM SIGMETRICS Performance Evaluation Review, 38(1):203– 214, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Misra:2010:IPA

[1858] Vishal Misra, Stratis Ioannidis, Augustin Chaintreau, and Laurent Massoulié. Incentivizing peer-assisted services: a fluid Shapley value approach. ACM SIGMETRICS Performance Evaluation Review, 38(1):215– 226, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ma:2010:LPM

[1859] Yadi Ma, Suman Banerjee, Shan Lu, and Cristian Estan. Leveraging parallelism for multi-dimensional packet classification on software routers. ACM SIGMETRICS Performance Evaluation Review, 38(1):227–238, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shah:2010:QPW

[1860] Devavrat Shah, John N. Tsitsiklis, and Yuan Zhong. Qualitative properties of α-weighted scheduling policies. ACM SIGMETRICS Performance Evaluation Review, 38(1):239–250, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2010:CMD

[1861] Giuliano Casale, Ningfang Mi, and Evgenia Smirni. CWS: a model-driven scheduling policy for correlated workloads. ACM SIGMETRICS Performance Evaluation Review, 38(1):251– 262, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zheng:2010:RAU

[1862] Haoqiang Zheng and Jason Nieh. RSIO: automatic user interaction detection and scheduling. ACM SIG-METRICS Performance Evaluation Review, 38(1):263–274, June 2010. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bramson:2010:RLB

[1863] Maury Bramson, Yi Lu, and Balaji Prabhakar. Randomized load balancing with general service time distributions. ACM SIGMETRICS Performance Evaluation Review, 38(1):275– 286, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ganesh:2010:LBR

[1864] Ayalvadi Ganesh, Sarah Lilienthal, D. Manjunath, Alexandre Proutiere, and Florian Simatos. Load balancing via random local search in closed and open systems. ACM SIGMET-RICS Performance Evaluation Review, 38(1):287–298, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhao:2010:UMF

[1865] Haiquan (Chuck) Zhao, Cathy H. Xia, Zhen Liu, and Don Towsley. A unified modeling framework for distributed resource allocation of general fork and join processing networks. ACM SIGMETRICS Performance Evaluation Review, 38(1):299–310, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ioannidis:2010:DCH

[1866] Stratis Ioannidis, Laurent Massoulie, and Augustin Chaintreau. Distributed caching over heterogeneous mobile networks. ACM SIGMETRICS Performance Evaluation Review, 38(1):311– 322, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Antunes:2010:AFI

[1867] Nelson Antunes, Gonçalo Jacinto, and António Pacheco. An analytical framework to infer multihop path reliability in MANETs. ACM SIGMET-RICS Performance Evaluation Review, 38(1):323–332, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coffman:2010:CFD

[1868] Ed Coffman, Philippe Robert, Florian Simatos, Shuzo Tarumi, and Gil Zussman. Channel fragmentation in dynamic spectrum access systems: a theoretical study. ACM SIGMET-RICS Performance Evaluation Review, 38(1):333–344, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bermond:2010:DSA

[1869] Jean-Claude Bermond, Dorian Mazauric, Vishal Misra, and Philippe Nain. A distributed scheduling algorithm for wireless networks with constant overhead and arbitrary binary interference. ACM SIGMETRICS Performance Evaluation Review, 38(1):345– 346, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sagnol:2010:SOD

[1870] Guillaume Sagnol, Mustapha Bouhtou, and Stéphane Gaubert. Successive c-optimal designs: a scalable technique to optimize the measurements on large networks. ACM SIGMET-RICS Performance Evaluation Review, 38(1):347–348, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cuevas:2010:DDB

[1871] Rubén Cuevas, Nikolaos Laoutaris, Xiaoyuan Yang, Georgos Siganos, and Pablo Rodriguez. Deep diving into BitTorrent locality. ACM SIGMET-RICS Performance Evaluation Review, 38(1):349–350, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jin:2010:IAN

[1872] Yu Jin, Nick Duffield, Patrick Haffner, Subhabrata Sen, and Zhi-Li Zhang. Inferring applications at the network layer using collective traffic statistics. ACM SIGMETRICS Performance Evaluation Review, 38(1):351– 352, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anselmi:2010:PAP

[1873] Jonatha Anselmi and Bruno Gaujal. The price of anarchy in parallel queues revisited. ACM SIGMET-RICS Performance Evaluation Review, 38(1):353–354, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Khouzani:2010:OPS

[1874] M. H. R. Khouzani, Saswati Sarkar, and Eitan Altman. Optimal propagation of security patches in mobile wireless networks: extended abstract. ACM SIGMETRICS Performance Evaluation Review, 38(1):355–356, June 2010.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Le:2010:MCE

[1875] Kien Le, Ozlem Bilgir, Ricardo Bianchini, Margaret Martonosi, and Thu D. Nguyen. Managing the cost, energy consumption, and carbon footprint of Internet services. ACM SIGMET-RICS Performance Evaluation Review, 38(1):357–358, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mishra:2010:CPM

[1876] Asit K. Mishra, Shekhar Srikantaiah, Mahmut Kandemir, and Chita R. Das. Coordinated power management of voltage islands in CMPs. ACM SIGMETRICS Performance Evaluation Review, 38(1):359–360, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nguyen:2010:RSA

[1877] Hung X. Nguyen and Matthew Roughan. Rigorous statistical analysis of Internet loss measurements. ACM SIGMETRICS Performance Evaluation Review, 38(1):361–362, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Osogami:2010:SOT

[1878] Takayuki Osogami and Rudy Raymond. Semidefinite optimization for transient analysis of queues. ACM SIGMETRICS Performance Evaluation Review, 38(1):363–364, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Park:2010:CCF

[1879] Dongchul Park, Biplob Debnath, and David Du. CFTL: a convertible flash translation layer adaptive to data access patterns. ACM SIGMETRICS Performance Evaluation Review, 38 (1):365–366, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Qian:2010:CUL

[1880] Feng Qian, Abhinav Pathak, Yu Charlie Hu, Zhuoqing Morley Mao, and Yinglian Xie. A case for unsupervisedlearning-based spam filtering. ACM SIGMETRICS Performance Evaluation Review, 38(1):367–368, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rajagopalan:2010:DAD

[1881] Shreevatsa Rajagopalan and Devavrat Shah. Distributed averaging in dynamic networks. ACM SIGMET-RICS Performance Evaluation Review, 38(1):369–370, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sarikaya:2010:PBP

[1882] Ruhi Sarikaya, Canturk Isci, and Alper Buyuktosunoglu. Program behavior prediction using a statistical metric model. ACM SIGMETRICS Performance Evaluation Review, 38(1):371– 372, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shah:2010:DOQ

[1883] Devavrat Shah and Jinwoo Shin. Delay optimal queue-based CSMA. ACM SIGMETRICS Performance Evaluation Review, 38(1):373–374, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shye:2010:CMU

[1884] Alex Shye, Benjamin Scholbrock, Gokhan Memik, and Peter A. Dinda. Characterizing and modeling user activity on smartphones: summary. ACM SIGMETRICS Performance Evaluation Review, 38(1):375–376, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Silveira:2010:DTA

[1885] Fernando Silveira, Christophe Diot, Nina Taft, and Ramesh Govindan. Detecting traffic anomalies using an equilibrium property. ACM SIGMET-RICS Performance Evaluation Review, 38(1):377–378, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Soundararajan:2010:CSE

[1886] Niranjan Soundararajan, Anand Sivasubramaniam, and Vijay Narayanan. Characterizing the soft error vulnerability of multicores running multithreaded applications. ACM SIGMET-RICS Performance Evaluation Review, 38(1):379–380, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2010:CMM

[1887] Jian Tan, Wei Wei, Bo Jiang, Ness Shroff, and Don Towsley. Can multipath mitigate power law delays?: effects of parallelism on tail performance. ACM SIGMETRICS Performance Evaluation Review, 38(1):381– 382, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tomozei:2010:DUP

[1888] Dan-Cristian Tomozei and Laurent Massoulié. Distributed user profiling via spectral methods. ACM SIGMET-RICS Performance Evaluation Review, 38(1):383–384, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

George:2010:AAC

[1889] David K. George and Cathy H. Xia.

Asymptotic analysis of closed queueing networks and its implications to achievable service levels. ACM SIGMET-RICS Performance Evaluation Review, 38(2):3–5, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Buaic:2010:SBM

[1890] Ana Buaic, Varun Gupta, and Jean Mairesse. Stability of the bipartite matching model. ACM SIGMET-RICS Performance Evaluation Review, 38(2):6–8, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tizghadam:2010:RWD

[1891] Ali Tizghadam and Alberto Leon-Garcia. On random walks in directionaware network problems. ACM SIG-METRICS Performance Evaluation Review, 38(2):9–11, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:2010:ART

[1892] Minghong Lin, Adam Wierman, and Bert Zwart. The average response time in a heavy-traffic SRPT queue. ACM SIGMETRICS Performance Evaluation Review, 38(2):12–14, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sigman:2010:HTL

[1893] Karl Sigman and Ward Whitt. Heavytraffic limits for nearly deterministic queues. ACM SIGMETRICS Performance Evaluation Review, 38(2): 15–17, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ye:2010:DLT

[1894] Heng-Qing Ye and David D. Yao. Diffusion limit of a two-class network: stationary distributions and interchange of limits. ACM SIGMETRICS Performance Evaluation Review, 38(2): 18–20, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nino-Mora:2010:IPA

[1895] José Niño-Mora. Index policies for admission and routing of soft realtime traffic to parallel queues. ACM SIGMETRICS Performance Evaluation Review, 38(2):21–23, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Robert:2010:MFA

[1896] Philippe Robert and Jim Roberts. A mean field approximation for the capacity of server-limited, gate-limited multi-server polling systems. ACM SIGMETRICS Performance Evaluation Review, 38(2):24–26, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2010:FAL

[1897] Yunan Liu and Ward Whitt. A fluid approximation for large-scale service systems. ACM SIGMETRICS Performance Evaluation Review, 38 (2):27–29, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gast:2010:MFL

[1898] Nicolas Gast and Bruno Gaujal. Mean field limit of non-smooth systems and differential inclusions. ACM SIGMET-RICS Performance Evaluation Review, 38(2):30–32, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Radovanovic:2010:RMT

[1899] Ana Radovanović and Assaf Zeevi.

Revenue maximization through “smart” inventory management in reservationbased online advertising. ACM SIG-METRICS Performance Evaluation Review, 38(2):33–35, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cho:2010:VFP

[1900] Jeong woo Cho, Jean-Yves Le Boudec, and Yuming Jiang. On the validity of the fixed point equation and decoupling assumption for analyzing the 802.11 MAC protocol. ACM SIGMET-RICS Performance Evaluation Review, 38(2):36–38, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

vandeVen:2010:ETR

[1901] P. M. van de Ven, S. C. Borst, D. Denteneer, A. J. E. M. Janssen, and J. S. H. van Leeuwaarden. Equalizing throughputs in random-access networks. ACM SIGMETRICS Performance Evaluation Review, 38(2):39–41, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marot:2010:RCP

[1902] Michel Marot and Vincent Gauthier. Reducing collision probability on a shared medium using a variational method. ACM SIGMETRICS Performance Evaluation Review, 38(2): 42–44, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2010:AMM

[1903] Yingdong Lu and Mark S. Squillante. On approximations for multiple multidimensional stochastic knapsacks. ACM SIGMETRICS Performance Evaluation Review, 38(2):45–47, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gandhi:2010:DRM

[1904] Anshul Gandhi, Mor Harchol-Balter, and Ivo Adan. Decomposition results for an M/M/k with staggered setup. ACM SIGMETRICS Performance Evaluation Review, 38(2):48–50, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pal:2010:EIS

[1905] Ranjan Pal and Leana Golubchik. On the economics of information security: the problem of designing optimal cyber-insurance contracts. ACM SIGMETRICS Performance Evaluation Review, 38(2):51–53, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dube:2010:RDC

[1906] Parijat Dube and Li Zhang. Resiliency of distributed clock synchronization networks. ACM SIGMET-RICS Performance Evaluation Review, 38(2):54–56, September 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2010:RAD

[1907] Xiaozhou Li, Mark Lillibridge, and Mustafa Uysal. Reliability analysis of deduplicated and erasure-coded storage. ACM SIGMETRICS Performance Evaluation Review, 38(3):4–9, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kulkarni:2010:TAI

[1908] Milind Kulkarni, Vijay Pai, and Derek Schuff. Towards architecture independent metrics for multicore performance analysis. ACM SIGMETRICS Performance Evaluation Review, 38 (3):10–14, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shepard:2010:LMW

[1909] Clayton Shepard, Ahmad Rahmati, Chad Tossell, Lin Zhong, and Phillip Kortum. LiveLab: measuring wireless networks and smartphone users in the field. ACM SIGMETRICS Performance Evaluation Review, 38 (3):15–20, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hahn:2010:UVL

[1910] Dongwoon Hahn, Ginnah Lee, Brenton Walker, Matt Beecher, and Padma Mundur. Using virtualization and live migration in a scalable mobile wireless testbed. ACM SIGMETRICS Performance Evaluation Review, 38 (3):21–25, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shakkottai:2010:TCD

[1911] Srinivas Shakkottai, Lei Ying, and Sankalp Sah. Targeted coupon distribution using social networks. ACM SIGMETRICS Performance Evaluation Review, 38(3):26–30, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gopalakrishnan:2010:AVG

[1912] Ragavendran Gopalakrishnan, Jason R. Marden, and Adam Wierman. An architectural view of game theoretic control. ACM SIGMETRICS Performance Evaluation Review, 38 (3):31–36, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yao:2010:DDL

[1913] Zhongmei Yao, Daren B. H. Cline, and Dmitri Loguinov. In-degree dynamics of large-scale P2P systems. ACM SIGMETRICS Performance Evaluation Review, 38(3):37–42, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Joumblatt:2010:HAE

[1914] Diana Joumblatt, Renata Teixeira, Jaideep Chandrashekar, and Nina Taft. HostView: annotating end-host performance measurements with user feedback. ACM SIGMETRICS Performance Evaluation Review, 38(3):43–48, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Adhikari:2010:TMR

[1915] Vijay Kumar Adhikari, Sourabh Jain, and Zhi-Li Zhang. From traffic matrix to routing matrix: PoP level traffic characteristics for a Tier-1 ISP. ACM SIGMETRICS Performance Evaluation Review, 38(3):49–54, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arlitt:2010:SIG

[1916] Martin Arlitt, Niklas Carlsson, and Jerry Rolia. Special issue on the 2010 GreenMetrics workshop. ACM SIGMETRICS Performance Evaluation Review, 38(3):??, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krishnan:2010:VPM

[1917] Bhavani Krishnan, Hrishikesh Amur, Ada Gavrilovska, and Karsten Schwan. VM power metering: feasibility and challenges. ACM SIGMETRICS Performance Evaluation Review, 38(3): 56–60, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Phillips:2010:RAI

[1918] Steven Phillips, Sheryl L. Woodward, Mark D. Feuer, and Peter D. Magill. A regression approach to infer electricity consumption of legacy telecom equipment. ACM SIGMET-RICS Performance Evaluation Review, 38(3):61–65, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sikdar:2010:EII

[1919] Biplab Sikdar. Environmental impact of IEEE 802.11 access points: a case study. ACM SIGMETRICS Performance Evaluation Review, 38 (3):66–70, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

V:2010:NDB

[1920] Prabhakar T.V., Akshay Uttama Nambi S.N., Jamadagni H.S., Krishna Swaroop, R. Venkatesha Prasad, and I. G. M. M. Niemegeers. A novel DTN based energy neutral transfer scheme for energy harvested WSN Gateways. ACM SIGMETRICS Performance Evaluation Review, 38(3):71– 75, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lombardo:2010:AES

[1921] Alfio Lombardo, Carla Panarello, and Giovanni Schembra. Achieving energy savings and QoS in Internet access routers. ACM SIGMETRICS Performance Evaluation Review, 38 (3):76–80, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bianzino:2010:AAF

[1922] Aruna Prem Bianzino, Anand Kishore Raju, and Dario Rossi. Applesto-apples: a framework analysis for energy-efficiency in networks. ACM SIGMETRICS Performance Evaluation Review, 38(3):81–85, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Janssen:2011:USD

[1923] Curtis L. Janssen, Helgi Adalsteinsson, and Joseph P. Kenny. Using simulation to design extremescale applications and architectures: programming model exploration. ACM SIGMETRICS Performance Evaluation Review, 38(4):4–8, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Giles:2011:PAO

[1924] M. B. Giles, G. R. Mudalige, Z. Sharif, G. Markall, and P. H. J. Kelly. Performance analysis of the OP2 framework on many-core architectures. ACM SIGMETRICS Performance Evaluation Review, 38(4):9–15, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Herdman:2011:BMP

[1925] J. A. Herdman, W. P. Gaudin, D. Turland, and S. D. Hammond. Benchmarking and modelling of POWER7, Westmere, BG/P, and GPUs: an industry case study. ACM SIGMET-RICS Performance Evaluation Review, 38(4):16–22, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pennycook:2011:PAH

[1926] S. J. Pennycook, S. D. Hammond, S. A.

Jarvis, and G. R. Mudalige. Performance analysis of a hybrid MPI/CUDA implementation of the NASLU benchmark. ACM SIGMETRICS Performance Evaluation Review, 38(4):23–29, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Budanur:2011:MTC

[1927] Sandeep Budanur, Frank Mueller, and Todd Gamblin. Memory trace compression and replay for SPMD systems using extended PRSDs? ACM SIGMET-RICS Performance Evaluation Review, 38(4):30–36, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rodrigues:2011:SST

[1928] A. F. Rodrigues, K. S. Hemmert, B. W. Barrett, C. Kersey, R. Oldfield, M. Weston, R. Risen, J. Cook, P. Rosenfeld, E. CooperBalls, and B. Jacob. The Structural Simulation Toolkit. ACM SIGMETRICS Performance Evaluation Review, 38(4):37–42, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Karlin:2011:PMP

[1929] Ian Karlin, Elizabeth Jessup, Geoffrey Belter, and Jeremy G. Siek. Parallel memory prediction for fused linear algebra kernels. ACM SIGMET-RICS Performance Evaluation Review, 38(4):43–49, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nakasato:2011:FGI

[1930] Naohito Nakasato. A fast GEMM implementation on the Cypress GPU. ACM SIGMETRICS Performance Evaluation Review, 38(4):50–55, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wu:2011:PCH

[1931] Xingfu Wu and Valerie Taylor. Performance characteristics of hybrid MPI/ OpenMP implementations of NAS parallel benchmarks SP and BT on largescale multicore supercomputers. ACM SIGMETRICS Performance Evaluation Review, 38(4):56–62, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hsieh:2011:FAL

[1932] Ming yu Hsieh, Arun Rodrigues, Rolf Riesen, Kevin Thompson, and William Song. A framework for architecturelevel power, area, and thermal simulation and its application to networkon-chip design exploration. ACM SIGMETRICS Performance Evaluation Review, 38(4):63–68, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Perks:2011:SWW

[1933] O. Perks, S. D. Hammond, S. J. Pennycook, and S. A. Jarvis. Should we worry about memory loss? ACM SIGMET-RICS Performance Evaluation Review, 38(4):69–74, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cook:2011:SPM

[1934] Jeanine Cook, Jonathan Cook, and Waleed Alkohlani. A statistical performance model of the Opteron processor. ACM SIGMETRICS Performance Evaluation Review, 38(4):75–80, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tabbal:2011:PDE

[1935] Alexandre Tabbal, Matthew Anderson, Maciej Brodowicz, Hartmut Kaiser, and Thomas Sterling. Preliminary design examination of the ParalleX system from a software and hardware perspective. ACM SIGMETRICS Performance Evaluation Review, 38(4):81–87, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

McIntosh-Smith:2011:EAM

[1936] Simon McIntosh-Smith, Terry Wilson, Jon Crisp, Amaurys Ávila Ibarra, and Richard B. Sessions. Energyaware metrics for benchmarking heterogeneous systems. ACM SIGMET-RICS Performance Evaluation Review, 38(4):88–94, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2011:MPR

[1937] Jian Chen, Lizy Kurian John, and Dimitris Kaseridis. Modeling program resource demand using inherent program characteristics. ACM SIGMETRICS Performance Evaluation Review, 39(1): 1–12, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sharifi:2011:MME

[1938] Akbar Sharifi, Shekhar Srikantaiah, Asit K. Mishra, Mahmut Kandemir, and Chita R. Das. METE: meeting end-to-end QoS in multicores through system-wide resource management. ACM SIGMETRICS Performance Evaluation Review, 39(1):13–24, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2011:SIC

[1939] Yuanrui Zhang, Mahmut Kandemir, and Taylan Yemliha. Studying intercore data reuse in multicores. ACM SIGMETRICS Performance Evaluation Review, 39(1):25–36, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2011:SIH

[1940] Fang Liu and Yan Solihin. Studying the impact of hardware prefetching and bandwidth partitioning in chipmultiprocessors. ACM SIGMETRICS Performance Evaluation Review, 39(1): 37–48, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Alizadeh:2011:SAQ

[1941] Mohammad Alizadeh, Abdul Kabbani, Berk Atikoglu, and Balaji Prabhakar. Stability analysis of QCN: the averaging principle. ACM SIGMET-RICS Performance Evaluation Review, 39(1):49–60, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Joseph:2011:SNM

[1942] Vinay Joseph and Gustavo de Veciana. Stochastic networks with multipath flow control: impact of resource pools on flow-level performance and network congestion. ACM SIGMET-RICS Performance Evaluation Review, 39(1):61–72, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Alizadeh:2011:ADS

[1943] Mohammad Alizadeh, Adel Javanmard, and Balaji Prabhakar. Analysis of DCTCP: stability, convergence, and fairness. ACM SIGMETRICS Performance Evaluation Review, 39(1):73–84, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Suh:2011:SEB

[1944] Jinho Suh, Mehrtash Manoochehri, Murali Annavaram, and Michel Dubois. Soft error benchmarking of L2 caches with PARMA. ACM SIGMET-RICS Performance Evaluation Review, 39(1):85–96, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Suchara:2011:NAJ

[1945] Martin Suchara, Dahai Xu, Robert Doverspike, David Johnson, and Jennifer Rexford. Network architecture for joint failure recovery and traffic engineering. ACM SIGMETRICS Performance Evaluation Review, 39(1):97– 108, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Subhraveti:2011:RTP

[1946] Dinesh Subhraveti and Jason Nieh. Record and transplay: partial checkpointing for replay debugging across heterogeneous systems. ACM SIG-METRICS Performance Evaluation Review, 39(1):109–120, June 2011. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tsitsiklis:2011:PEL

[1947] John N. Tsitsiklis and Kuang Xu. On the power of (even a little) centralization in distributed processing. ACM SIGMETRICS Performance Evaluation Review, 39(1):121–132, June 2011.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nguyen:2011:WPA

[1948] Thanh Nguyen and Milan Vojnovic. Weighted proportional allocation. ACM SIGMETRICS Performance Evaluation Review, 39(1):133– 144, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aalto:2011:OTB

[1949] Samuli Aalto, Aleksi Penttinen, Pasi Lassila, and Prajwal Osti. On the optimal trade-off between SRPT and opportunistic scheduling. ACM SIGMETRICS Performance Evaluation Review, 39(1):145–155, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cohen:2011:SAS

[1950] Edith Cohen, Graham Cormode, and Nick Duffield. Structure-aware sampling on data streams. ACM SIGMET-RICS Performance Evaluation Review, 39(1):157–168, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Korada:2011:GP

[1951] Satish Babu Korada, Andrea Montanari, and Sewoong Oh. Gossip PCA. ACM SIGMETRICS Performance Evaluation Review, 39(1):169– 180, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Urgaonkar:2011:OPC

[1952] Rahul Urgaonkar, Bhuvan Urgaonkar, Michael J. Neely, and Anand Sivasubramaniam. Optimal power cost management using stored energy in data centers. ACM SIGMETRICS Performance Evaluation Review, 39(1):181– 192, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2011:GGL

[1953] Zhenhua Liu, Minghong Lin, Adam Wierman, Steven H. Low, and Lachlan L. H. Andrew. Greening geographical load balancing. ACM SIGMET-RICS Performance Evaluation Review, 39(1):193–204, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nguyen:2011:SP

[1954] Giang T. K. Nguyen, Rachit Agarwal, Junda Liu, Matthew Caesar, P. Brighten Godfrey, and Scott Shenker. Slick packets. ACM SIG-METRICS Performance Evaluation Review, 39(1):205–216, June 2011. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lam:2011:GRD

[1955] Simon S. Lam and Chen Qian. Geographic routing in d-dimensional spaces with guaranteed delivery and low stretch. ACM SIGMETRICS Performance Evaluation Review, 39(1):217– 228, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rozner:2011:MDO

[1956] Eric Rozner, Mi Kyung Han, Lili Qiu, and Yin Zhang. Model-driven optimization of opportunistic routing. ACM SIGMETRICS Performance Evaluation Review, 39(1):229– 240, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kurant:2011:WGM

[1957] Maciej Kurant, Minas Gjoka, Carter T.

Butts, and Athina Markopoulou. Walking on a graph with a magnifying glass: stratified sampling via weighted random walks. ACM SIGMETRICS Performance Evaluation Review, 39 (1):241–252, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anandkumar:2011:TDS

[1958] Animashree Anandkumar, Avinatan Hassidim, and Jonathan Kelner. Topology discovery of sparse random graphs with few participants. ACM SIGMET-RICS Performance Evaluation Review, 39(1):253–264, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shafiq:2011:CMI

[1959] M. Zubair Shafiq, Lusheng Ji, Alex X.

Liu, and Jia Wang. Characterizing and modeling Internet traffic dynamics of cellular devices. ACM SIGMET-RICS Performance Evaluation Review, 39(1):265–276, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xu:2011:CDN

[1960] Qiang Xu, Junxian Huang, Zhaoguang Wang, Feng Qian, Alexandre Gerber, and Zhuoqing Morley Mao. Cellular data network infrastructure characterization and implication on mobile content placement. ACM SIGMET-RICS Performance Evaluation Review, 39(1):277–288, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2011:FGL

[1961] Myungjin Lee, Sharon Goldberg, Ramana Rao Kompella, and George Varghese. Fine-grained latency and loss measurements in the presence of reordering. ACM SIGMETRICS Performance Evaluation Review, 39(1):289– 300, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhou:2011:SOU

[1962] Xia Zhou, Stratis Ioannidis, and Laurent Massoulie. On the stability and optimality of universal swarms. ACM SIGMETRICS Performance Evaluation Review, 39(1):301–312, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Eibl:2011:FBE

[1963] Patrick J. Eibl, Albert Meixner, and Daniel J. Sorin. An FPGA-based experimental evaluation of microprocessor core error detection with Argus-2. ACM SIGMETRICS Performance Evaluation Review, 39(1):313– 314, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2011:RKD

[1964] Lele Zhang, Darryl Veitch, and Kotagiri Ramamohanarao. The role of KL divergence in anomaly detection. ACM SIGMETRICS Performance Evaluation Review, 39(1):315–316, June 2011.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krevat:2011:AIL

[1965] Elie Krevat, Tomer Shiran, Eric Anderson, Joseph Tucek, Jay J. Wylie, and Gregory R. Ganger. Applying idealized lower-bound runtime models to understand inefficiencies in dataintensive computing. ACM SIGMET-RICS Performance Evaluation Review, 39(1):317–318, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Han:2011:HPC

[1966] Jinyoung Han, Taejoong Chung, Seungbae Kim, Ted Taekyoung Kwon, Hyun chul Kim, and Yanghee Choi. How prevalent is content bundling in BitTorrent? ACM SIGMET-RICS Performance Evaluation Review, 39(1):319–320, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rao:2011:SAP

[1967] Jia Rao, Xiangping Bu, Kun Wang, and Cheng-Zhong Xu. Self-adaptive provisioning of virtualized resources in cloud computing. ACM SIGMET-RICS Performance Evaluation Review, 39(1):321–322, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2011:CAR

[1968] Chao Li, Amer Qouneh, and Tao Li.

Characterizing and analyzing renewable energy driven data centers. ACM SIGMETRICS Performance Evaluation Review, 39(1):323–324, June 2011.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:2011:TMB

[1969] Varun Gupta and Takayuki Osogami. Tight moments-based bounds for queueing systems. ACM SIGMET-RICS Performance Evaluation Review, 39(1):325–326, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2011:SMT

[1970] Suk-Bok Lee, Dan Pei, Mohammad-Taghi Hajiaghayi, Ioannis Pefkianakis, Songwu Lu, He Yan, Zihui Ge, Jennifer Yates, and Mario Kosseifi. Scalable monitoring via threshold compression in a large operational 3G network. ACM SIGMETRICS Performance Evaluation Review, 39(1):327– 328, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Adhikari:2011:HDY

[1971] Vijay Kumar Adhikari, Sourabh Jain, Yingying Chen, and Zhi-Li Zhang. How do you ’Tube’? ACM SIGMET-RICS Performance Evaluation Review, 39(1):329–330, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kant:2011:CSB

[1972] Krishna Kant. A control scheme for batching DRAM requests to improve power efficiency. ACM SIGMET-RICS Performance Evaluation Review, 39(1):331–332, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2011:ONS

[1973] Hao Zhang, Ziyu Shao, Minghua Chen, and Kannan Ramchandran. Optimal neighbor selection in BitTorrent-like peer-to-peer networks. ACM SIGMET-RICS Performance Evaluation Review, 39(1):333–334, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ihm:2011:TUM

[1974] Sunghwan Ihm and Vivek S. Pai.

Towards understanding modern Web traffic. ACM SIGMETRICS Performance Evaluation Review, 39(1):335– 336, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Akella:2011:OIR

[1975] Aditya Akella, Shuchi Chawla, Holly Esquivel, and Chitra Muthukrishnan. De-ossifying Internet routing through intrinsic support for end-network and ISP selfishness. ACM SIGMETRICS Performance Evaluation Review, 39 (1):337–338, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hong:2011:DSP

[1976] Yu-Ju Hong, Jiachen Xue, and Mithuna Thottethodi. Dynamic server provisioning to minimize cost in an IaaS cloud. ACM SIGMETRICS Performance Evaluation Review, 39(1): 339–340, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Srinivasan:2011:HHA

[1977] Sadagopan Srinivasan, Ravishankar Iyer, Li Zhao, and Ramesh Illikkal. HeteroScouts: hardware assist for OS scheduling in heterogeneous CMPs. ACM SIGMETRICS Performance Evaluation Review, 39(1):341– 342, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ribeiro:2011:CCT

[1978] Bruno Ribeiro, Daniel Figueiredo, Edmundo de Souza e Silva, and Don Towsley. Characterizing continuoustime random walks on dynamic networks. ACM SIGMETRICS Performance Evaluation Review, 39(1):343– 344, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2011:AAN

[1979] Jian Chen and Lizy Kurian John. Autocorrelation analysis: a new and improved method for measuring branch predictability. ACM SIGMETRICS Performance Evaluation Review, 39 (1):345–346, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Singh:2011:IGM

[1980] Satinder Pal Singh, Randolph Baden, Choon Lee, Bobby Bhattacharjee, Richard La, and Mark Shayman. IP geolocation in metropolitan areas. ACM SIGMETRICS Performance Evaluation Review, 39(1):347–348, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2011:TBS

[1981] Jay Chen, Janardhan Iyengar, Lakshminarayanan Subramanian, and Bryan Ford. TCP behavior in sub packet regimes. ACM SIGMETRICS Performance Evaluation Review, 39(1):349– 350, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bowden:2011:NLT

[1982] Rhys Alistair Bowden, Matthew Roughan, and Nigel Bean. Network link tomography and compressive sensing. ACM SIGMETRICS Performance Evaluation Review, 39(1):351– 352, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gulati:2011:STM

[1983] Ajay Gulati and Irfan Ahmad. Storage technologies, management and troubleshooting in virtualized datacenters. ACM SIGMETRICS Performance Evaluation Review, 39(1):353– 354, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sengupta:2011:CDC

[1984] Sudipta Sengupta. Cloud data center networks: technologies, trends, and challenges. ACM SIGMETRICS Performance Evaluation Review, 39 (1):355–356, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2011:BAW

[1985] Giuliano Casale. Building accurate workload models using Markovian arrival processes. ACM SIGMETRICS Performance Evaluation Review, 39 (1):357–358, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ciucu:2011:NAC

[1986] Florin Ciucu. Non-asymptotic capacity and delay analysis of mobile wireless networks. ACM SIGMET-RICS Performance Evaluation Review, 39(1):359–360, June 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Elmokashfi:2011:SSI

[1987] Ahmed Elmokashfi, Amund Kvalbein, and Constantine Dovrolis. SIMROT: a scalable inter-domain routing toolbox. ACM SIGMETRICS Performance Evaluation Review, 39(2):4–13, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Sen:2011:CIH

[1988] Aritra Sen, Ankit Garg, Akshat Verma, and Tapan Nayak. CloudBridge: on integrated hardware-software consolidation. ACM SIGMETRICS Performance Evaluation Review, 39(2):14– 25, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Nair:2011:ENE

[1989] Jayakrishnan Nair, Adam Wierman, and Bert Zwart. Exploiting network effects in the provisioning of large scale systems. ACM SIGMETRICS Performance Evaluation Review, 39(2):26– 28, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Nightingale:2011:PES

[1990] James Nightingale, Qi Wang, and Christos Grecos. Performance evaluation of scalable video streaming in multihomed mobile networks. ACM SIGMETRICS Performance Evaluation Review, 39(2):29–31, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Bouman:2011:DPB

[1991] N. Bouman, S. C. Borst, and J. S. H.

van Leeuwaarden. Delay performance of backlog based random access. ACM SIGMETRICS Performance Evaluation Review, 39(2):32–34, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Shneer:2011:CSC

[1992] Seva Shneer and Peter M. van de Ven. Comparing slotted and continuous CSMA: throughputs and fairness. ACM SIGMETRICS Performance Evaluation Review, 39(2):35– 37, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Shvets:2011:AMI

[1993] Evgeny Shvets, Andrey Lyakhov, Alexander Safonov, and Evgeny Khorov. Analytical model of IEEE 802.11s MCCAbased streaming in the presence of noise. ACM SIGMET-RICS Performance Evaluation Review, 39(2):38–40, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PER-FORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Ayesta:2011:HTA

[1994] U. Ayesta, A. Izagirre, and I. M. Verloop. Heavy traffic analysis of the discriminatory random-order-of-service discipline. ACM SIGMETRICS Performance Evaluation Review, 39(2):41– 43, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Boon:2011:QNS

[1995] M. A. A. Boon, R. D. van der Mei, and E. M. M. Winands. Queueing networks with a single shared server: light and heavy traffic. ACM SIGMET-RICS Performance Evaluation Review, 39(2):44–46, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PER-FORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Frolkova:2011:FPA

[1996] Maria Frolkova, Josh Reed, and Bert Zwart. Fixed-point approximations of bandwidth sharing networks with rate constraints. ACM SIGMETRICS Performance Evaluation Review, 39(2):47– 49, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Cano:2011:IPF

[1997] Maria Dolores Cano. Improving path failure detection in SCTP using adaptive heartbeat time intervals. ACM SIGMETRICS Performance Evaluation Review, 39(2):50–52, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Varis:2011:NSB

[1998] Nuutti Varis and Jukka Manner. In the network: Sandy Bridge versus Nehalem. ACM SIGMETRICS Performance Evaluation Review, 39(2):53– 55, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Anselmi:2011:EPS

[1999] Jonatha Anselmi Anselmi and Bruno Gaujal. On the efficiency of perfect simulation in monotone queueing networks. ACM SIGMETRICS Performance Evaluation Review, 39(2):56– 58, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Baryshnikov:2011:CLD

[2000] Y. M. Baryshnikov, E. G. Coffman, and K. J. Kwak. CAUCHY localization: a distributed computation of WSNs. ACM SIGMETRICS Performance Evaluation Review, 39(2):59– 61, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Goga:2011:IFS

[2001] Oana Goga, Patrick Loiseau, and Paulo Gonçalves. On the impact of the flow size distribution’s tail index on network performance with TCP connections. ACM SIGMETRICS Performance Evaluation Review, 39(2):62– 64, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

VanHoudt:2011:LBP

[2002] B. Van Houdt. Load balancing and the power of preventive probing. ACM SIGMETRICS Performance Evaluation Review, 39(2):65–67, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Altman:2011:PAC

[2003] Eitan Altman, Rachid El Azouzi, Daniel S. Menasché, and Yuedong Xu. Poster: Aging control for smartphones in hybrid networks. ACM SIGMET-RICS Performance Evaluation Review, 39(2):68, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PER-FORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Bokharaei:2011:PTN

[2004] Hossein Kaffash Bokharaei, Yashar Ganjali, Ram Keralapura, and Antonio Nucci. Poster: Telephony network characterization for spammer identification. ACM SIGMETRICS Performance Evaluation Review, 39(2): 69, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Bosman:2011:POD

[2005] Joost Bosman, Rob van der Mei, and Gerard Hoekstra. Poster: Optimal dispatching policies for parallel processor sharing nodes with partial information. ACM SIGMETRICS Performance Evaluation Review, 39(2): 70, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Dong:2011:PPS

[2006] Ke Dong, Diptanil DebBarma, R. Venkatesha, and Prasad Cheng Guo. Poster: Performance study of clustering of Zigbee devices in OPNET. ACM SIGMETRICS Performance Evaluation Review, 39(2):71, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Lubben:2011:PCD

[2007] Ralf Lübben and Markus Fidler. Poster: On the capacity delay error tradeoff of source coding. ACM SIGMETRICS Performance Evaluation Review, 39(2):72, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Marbukh:2011:PTE

[2008] Vladimir Marbukh. Poster: Tcp effective bandwidth and Internet performance. ACM SIGMETRICS Performance Evaluation Review, 39(2): 73, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Massey:2011:PSV

[2009] William A. Massey and Jamol Pender. Poster: Skewness variance approximation for dynamic rate MultiServer queues with abandonment. ACM SIGMETRICS Performance Evaluation Review, 39(2):74, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Rahman:2011:PGF

[2010] Ashikur Rahman and Carey Williamson.

Poster: Δ-graphs: flexible topology control in wireless ad hoc networks. ACM SIGMETRICS Performance Evaluation Review, 39(2):75, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Rahman:2011:PCM

[2011] Ashikur Rahman, Hanan Shpungin, and Carey Williamson. Poster: On capacity maximization in wireless relay networks. ACM SIGMETRICS Performance Evaluation Review, 39(2): 76, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Romano:2011:PSB

[2012] Paolo Romano and Matteo Leonetti.

Poster: Selftuning batching in total order broadcast via analytical modelling and reinforcement learning. ACM SIGMETRICS Performance Evaluation Review, 39(2):77, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Yan:2011:PDV

[2013] Zhichao Yan, Dan Feng, and Yujuan Tan. Poster: Dissection the version management schemes in hardware transactional memory systems. ACM SIGMETRICS Performance Evaluation Review, 39(2):78, September 2011. ISSN 0163-5999 (print), 1557-9484 (electronic). Special Issue on IFIP PERFORMANCE 2011- 29th International Symposium on Computer Performance, Modeling, Measurement and Evaluation.

Czekster:2011:EVD

[2014] Ricardo M. Czekster, Paulo Fernandes, and Thais Webber. Efficient vectordescriptor product exploiting timememory trade-offs! ACM SIGMET-RICS Performance Evaluation Review, 39(3):2–9, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lilja:2011:PAS

[2015] David J. Lilja, Raffaela Mirandola, and Kai Sachs. Paper abstracts of the second international conference on performance engineering (ICPE 2011). ACM SIGMETRICS Performance Evaluation Review, 39(3):2–9, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2011:IBT

[2016] Mark S. Squillante. Instrumentationbased tool for latency measurements (abstracts only). ACM SIGMET-RICS Performance Evaluation Review, 39(3):20, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Papadimitriou:2011:PVR

[2017] Dimitri Papadimitriou, Florin Coras, and Albert Cabellos. Path-vector routing stability analysis. ACM SIGMET-RICS Performance Evaluation Review, 39(3):22–24, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhao:2011:DAS

[2018] Haotian Zhao and Yinlong Xu. A deterministic algorithm of single failed node recovery in MSR-based distributed storage systems. ACM SIGMETRICS Performance Evaluation Review, 39 (3):25–27, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Garg:2011:RHD

[2019] Siddharth Garg, Shreyas Sundaram, and Hiren D. Patel. Robust heterogeneous data center design: a principled approach. ACM SIGMETRICS Performance Evaluation Review, 39 (3):28–30, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tizghadam:2011:RWN

[2020] Ali Tizghadam, Alberto Leon-Garcia, and Hassan Naser. On robust wireless network optimization using network criticality. ACM SIGMETRICS Performance Evaluation Review, 39 (3):31–33, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lelarge:2011:DCB

[2021] Marc Lelarge. Diffusion and cascading behavior in random networks. ACM SIGMETRICS Performance Evaluation Review, 39(3):34–36, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Abdelrahman:2011:SNH

[2022] Omer H. Abdelrahman and Erol Gelenbe. Search in non-homogeneous random environments? ACM SIGMET-RICS Performance Evaluation Review, 39(3):37–39, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Feng:2011:EPQ

[2023] Hanhua Feng, Parijat Dube, and Li Zhang. On estimation problems for the G/G/∞ queue. ACM SIGMET-RICS Performance Evaluation Review, 39(3):40–42, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Doroudi:2011:DIF

[2024] Sherwin Doroudi, Ragavendran Gopalakrishnan, and Adam Wierman. Dispatching to incentivize fast service in multi-server queues. ACM SIGMET-RICS Performance Evaluation Review, 39(3):43–45, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Akgun:2011:PPP

[2025] Osman T. Akgun, Rhonda Righter, and Ronald Wolff. The power of partial power of two choices. ACM SIGMET-RICS Performance Evaluation Review, 39(3):46–48, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pal:2011:SLQ

[2026] Ranjan Pal, Sokol Kosta, and Pan Hui. Settling for less: a QoS compromise mechanism for opportunistic mobile networks. ACM SIGMET-RICS Performance Evaluation Review, 39(3):49–51, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2011:IEN

[2027] Zichao Yang and John C. S. Lui. Investigating the effect of node heterogeneity and network externality on security adoption. ACM SIGMET-RICS Performance Evaluation Review, 39(3):52–54, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menasche:2011:IPS

[2028] Daniel Sadoc Menasché, Antonio A.

de A. Rocha, Edmundo A. de Souza e Silva, Don Towsley, and Rosa M. Meri Leäo. Implications of peer selection strategies by publishers on the performance of P2P swarming systems. ACM SIGMETRICS Performance Evaluation Review, 39(3):55–57, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aalto:2011:HIA

[2029] Samuli Aalto, Pasi Lassila, Petri Savolainen, and Sasu Tarkoma. How impatience affects the performance and scalability of P2P video-on-demand systems. ACM SIGMETRICS Performance Evaluation Review, 39(3):58–60, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arlitt:2011:PGW

[2030] Martin Arlitt, Niklas Carlsson, and Jerry Rolia. Proceedings of the 2011 GreenMetrics workshop. ACM SIGMETRICS Performance Evaluation Review, 39(3):58–60, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2011:GLB

[2031] Zhenhua Liu, Minghong Lin, Adam Wierman, Steven H. Low, and Lachlan L. H. Andrew. Geographical load balancing with renewables. ACM SIGMETRICS Performance Evaluation Review, 39(3):62–66, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Altman:2011:TGC

[2032] Eitan Altman, Manjesh K. Hanawal, Rachid ElAzouzi, and Sholomo Shamai. Tradeoffs in green cellular networks. ACM SIGMETRICS Performance Evaluation Review, 39(3):67–71, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sucevic:2011:PEE

[2033] Andrew Sucevic, Lachlan L. H. Andrew, and Thuy T. T. Nguyen. Powering down for energy efficient peer-to-peer file distribution. ACM SIGMET-RICS Performance Evaluation Review, 39(3):72–76, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brown:2011:RPS

[2034] Michael Brown and Jose Renau.

ReRack: power simulation for data centers with renewable energy generation. ACM SIGMETRICS Performance Evaluation Review, 39(3):77–81, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yan:2011:CRS

[2035] Feng Yan, Xenia Mountrouidou, Alma Riska, and Evgenia Smirni. Copy rate synchronization with performance guarantees for work consolidation in storage clusters. ACM SIGMETRICS Performance Evaluation Review, 39 (3):82–86, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:2011:APR

[2036] Vishal Gupta, Ripal Nathuji, and Karsten Schwan. An analysis of power reduction in datacenters using heterogeneous chip multiprocessors. ACM SIGMETRICS Performance Evaluation Review, 39(3):87–91, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2011:HSS

[2037] Giuliano Casale and Ioan Raicu.

HPDC/SIGMETRICS student research posters. ACM SIGMETRICS Performance Evaluation Review, 39 (3):92–96, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2011:UCG

[2038] Doron Chen, Ealan Henis, Ronen I.

Kat, Dmitry Sotnikov, Cinzia Cappiello, Alexandre Mello Ferreira, Barbara Pernici, Monica Vitali, Tao Jiang, Jia Liu, and Alexander Kipp. Usage centric green performance indicators. ACM SIGMETRICS Performance Evaluation Review, 39(3):92–96, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2011:BBH

[2039] Yuanrui Zhang, Jun Liu, Sai Prashanth Muralidhara, and Mahmut Kandemir. BrickX: building hybrid systems for recursive computations. ACM SIGMET-RICS Performance Evaluation Review, 39(3):98–100, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Blackburn:2011:CGS

[2040] Jeremy Blackburn, Ramanuja Simha, Clayton Long, Xiang Zuo, Nicolas Kourtellis, John Skvoretz, and Adriana Iamnitchi. Cheaters in a gaming social network. ACM SIGMET-RICS Performance Evaluation Review, 39(3):101–103, December 2011. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stefanek:2011:FCP

[2041] Anton Stefanek, Richard A. Hayden, and Jeremy T. Bradley. Fluid computation of the performance: energy tradeoff in large scale Markov models. ACM SIGMETRICS Performance Evaluation Review, 39(3):104–106, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kim:2011:IHP

[2042] Shingyu Kim, Junghee Won, Hyuck Han, Hyeonsang Eom, and Heon Y. Yeom. Improving Hadoop performance in intercloud environments. ACM SIGMETRICS Performance Evaluation Review, 39(3):107–109, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2011:IPE

[2043] Yong Oh Lee. Improving performance and energy savings through alternative forwarding. ACM SIGMETRICS Performance Evaluation Review, 39(3): 110–112, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Choi:2011:IPM

[2044] Seungmi Choi, Shingyu Kim, Hyuck Han, and Heon Y. Yeom. Improving performance of MapReduce framework on InterCloud by avoiding transmission of unnecessary data. ACM SIGMET-RICS Performance Evaluation Review, 39(3):113–115, December 2011. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gadre:2011:IMF

[2045] Hrishikesh Gadre, Ivan Rodero, and Manish Parashar. Investigating MapReduce framework extensions for efficient processing of geographically scattered datasets. ACM SIGMET-RICS Performance Evaluation Review, 39(3):116–118, December 2011. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hayden:2011:MFA

[2046] Richard A. Hayden. Mean-field approximations for performance models with generally-timed transitions. ACM SIGMETRICS Performance Evaluation Review, 39(3):119–121, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gandhi:2011:MMV

[2047] Rohan Gandhi, Dimitrios Koutsonikolas, and Y. Charlie Hu. Multicasting MDC videos to receivers with different screen resolution. ACM SIGMET-RICS Performance Evaluation Review, 39(3):122–124, December 2011. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sawalha:2011:TSH

[2048] Lina Sawalha, Monte P. Tull, and Ronald D. Barnes. Thread scheduling for heterogeneous multicore processors using phase identification. ACM SIGMETRICS Performance Evaluation Review, 39(3):125–127, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2011:EDH

[2049] Tonglin Li, Raman Verma, Xi Duan, Hui Jin, and Ioan Raicu. Exploring distributed hash tables in High-End computing. ACM SIGMETRICS Performance Evaluation Review, 39(3): 128–130, December 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Burdette:2012:ECJ

[2050] Philip F. Burdette, William F. Jones, Brian C. Blose, and Gregory M.

Kapfhammer. An empirical comparison of Java remote communication primitives for intra-node data transmission. ACM SIGMETRICS Performance Evaluation Review, 39(4):2–11, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gopalakrishnan:2012:SUT

[2051] Sathish Gopalakrishnan. Sharp utilization thresholds for some realtime scheduling problems. ACM SIGMET-RICS Performance Evaluation Review, 39(4):12–22, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coffman:2012:SLR

[2052] Edward G. Coffman. Synthesis of localrule processes: successes and challenges (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4): 24, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kou:2012:FPT

[2053] Steven S. G. Kou. First passage times and option pricing under a mixedexponential jump diffusion model (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):24, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Neuts:2012:AMS

[2054] Marcel F. Neuts. The algorithmization of mathematics: the story of stochastic models (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):24, April 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shah:2012:PFD

[2055] Devavrat Shah. Product-form distributions and network algorithms (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):24, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baek:2012:FPM

[2056] Jung Woo Baek, Ho Woo Lee, Se Won Lee, and Soohan Ahn. Factorization properties for a MAP-modulated fluid flow model under server vacation policies (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):25, April 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bladt:2012:BME

[2057] Mogens Bladt, Luz Judith R. Esparza, and Bo Friis Nielsen. Bilateral matrix-exponential distributions (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):25, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bladt:2012:MDP

[2058] Mogens Bladt and Bo Friis Nielsen. Moment distributions of phase-type (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4): 25–26, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Drekic:2012:SPP

[2059] Steve Drekic, David Stanford, and Douglas Woolford. A self-promoting priority model for transplant queues (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4): 26, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fackrell:2012:CME

[2060] Mark Fackrell. Characterizing matrixexponential distributions of order 4 (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4): 26, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hautphenne:2012:EAM

[2061] Sophie Hautphenne. An EM algorithm for the model fitting of Markovian binary trees (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):26–27, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hautphenne:2012:MTS

[2062] Sophie Hautphenne, Guy Latouche, and Giang T. Nguyen. Markovian trees subject to catastrophes: would they survive forever? (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):27, April 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

He:2012:DMV

[2063] Qi-Ming He and Jiandong Ren. On a discrete multi-variate phase-type distribution and its applications (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):27, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

He:2012:MEP

[2064] Qi-Ming He, Hanqin Zhang, and Juan Vera. Majorization and Extremal PHdistributions (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):27, April 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Horvath:2012:ARM

[2065] Gábor Horváth and Miklós Telek. Acceptance-rejection methods for generating random variates from matrix exponential distributions and rational arrival processes (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):27, April 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kobayashi:2012:TAS

[2066] Masahiro Kobayashi, Yutaka Sakuma, and Masakiyo Miyazawa. Tail asymptotics of the stationary distribution for M/M-JSQ with k parallel queues (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):28, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krishnamoorthy:2012:SDP

[2067] A. Krishnamoorthy and Viswanath C. Narayanan. Stochastic decomposition in production inventory with service time (abstract only). ACM SIGMET-RICS Performance Evaluation Review, 39(4):28, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Latouche:2012:TDF

[2068] Guy Latouche, Giang T. Nguyen, and Zbigniew Palmowski. Two-dimensional fluid queues with temporary assistance (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4): 28, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ramaswami:2012:FIB

[2069] V. Ramaswami. A fluid introduction to Brownian motion & stochastic integration (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):29, April 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sonenberg:2012:NFM

[2070] Nikki Sonenberg and Peter G. Taylor. A network of fluid models and its application in MANETs (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):29, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stanford:2012:NPP

[2071] David Stanford, Peter G. Taylor, and Ilze Ziedins. A new paradigm for priority patient selection (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):29, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Toyoizumi:2012:ADS

[2072] Hiroshi Toyoizumi and Jeremy Field.

Analysis of the dynamics of social queues by quasi-birth-and-death processes (abstract only). ACM SIGMET-RICS Performance Evaluation Review, 39(4):29–30, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

VanHoudt:2012:IDD

[2073] B. Van Houdt and J. F. Pérez. The impact of dampening demand variability in a production/inventory system with multiple retailers (abstract only). ACM SIGMETRICS Performance Evaluation Review, 39(4):30, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bean:2012:AQR

[2074] Nigel G. Bean and Bo Friis Nielsen. Analysis of queues with rational arrival process components: a general approach. ACM SIGMETRICS Performance Evaluation Review, 39(4):31, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bean:2012:SFM

[2075] Nigel G. Bean and Małgorzata M. O’Reilly. A stochastic fluid model driven by an uncountable-state process, which is a stochastic fluid model itself: the stochastic fluid-fluid model. ACM SIGMETRICS Performance Evaluation Review, 39(4):32, April 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bini:2012:CCR

[2076] Dario A. Bini, Paola Favati, and Beatrice Meini. A compressed cyclic reduction for QBDs with low rank upper and lower transitions. ACM SIGMETRICS Performance Evaluation Review, 39(4): 33, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bladt:2012:OMG

[2077] Mogens Bladt and Bo Friis Nielsen. An overview of multivariate gamma distributions as seen from a (multivariate) matrix exponential perspective. ACM SIGMETRICS Performance Evaluation Review, 39(4):34, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Blanchet:2012:RES

[2078] Jose Blanchet and Jing Dong. Rareevent simulation for multi-server queues in the Halfin–Whitt regime. ACM SIGMETRICS Performance Evaluation Review, 39(4):35, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2012:PFA

[2079] Giuliano Casale, Peter G. Harrison, and Maria Grazia Vigliotti. Productform approximation of queueing networks with phase-type service. ACM SIGMETRICS Performance Evaluation Review, 39(4):36, April 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dai:2012:NAD

[2080] J. G. Dai and Shuangchi He. Numerical analysis for diffusion models of GI/Ph/n + GI queues. ACM SIGMETRICS Performance Evaluation Review, 39(4):37, April 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Denardo:2012:SFM

[2081] Eric V. Denardo, Eugene A. Feinberg, and Uriel G. Rothblum. Splitting in a finite Markov decision problem. ACM SIGMETRICS Performance Evaluation Review, 39(4):38, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Denardo:2012:MAB

[2082] Eric V. Denardo, Eugene A. Feinberg, and Uriel G. Rothblum. The multiarmed bandit, with constraints. ACM SIGMETRICS Performance Evaluation Review, 39(4):39, April 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dendievel:2012:SDP

[2083] S. Dendievel, G. Latouche, and M-A. Remiche. Stationary distribution of a perturbed QBD process. ACM SIGMETRICS Performance Evaluation Review, 39(4):40, April 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Govorun:2012:PRP

[2084] Maria Govorun, Guy Latouche, and Marie-Ange Remiche. Profits and risks of pension plans. ACM SIGMETRICS Performance Evaluation Review, 39(4): 41, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kobayashi:2012:RTA

[2085] Masahiro Kobayashi and Masakiyo Miyazawa. Revisit to the tail asymptotics of the double QBD process by the analytic function method. ACM SIGMETRICS Performance Evaluation Review, 39(4):42, April 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2012:FMM

[2086] Yunan Liu and Ward Whitt. A fluid model for many-server queues with time-varying arrivals and phase-type service distribution. ACM SIGMET-RICS Performance Evaluation Review, 39(4):43, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Margolius:2012:NSM

[2087] Barbara Margolius. Numerical study of Markovian arrival processes (MAP) with time-varying periodic arrival rates. ACM SIGMETRICS Performance Evaluation Review, 39(4):44, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

OReilly:2012:SDF

[2088] Małgorzata M. O’Reilly and Nigel G.

Bean. Stochastic 2-dimensional fluid model. ACM SIGMETRICS Performance Evaluation Review, 39(4):45, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bini:2012:SQM

[2089] D. Bini, B. Meini, S. Steffé, J. F.

Pérez, and B. Van Houdt. SMCSolver and Q-MAM: tools for matrix-analytic methods. ACM SIGMETRICS Performance Evaluation Review, 39(4):46, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2012:KTF

[2090] Giuliano Casale and Evgenia Smirni.

KPC-toolbox: fitting Markovian arrival processes and phase-type distributions with MATLAB. ACM SIGMET-RICS Performance Evaluation Review, 39(4):47, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cote:2012:JPS

[2091] Marco Cote, German Riano, Raha Akhavan-Tabatabaei, Juan Fernando Perez, Andres Sarmiento, and Julio Goez. jMarkov package: a stochastic modeling tool. ACM SIGMETRICS Performance Evaluation Review, 39(4): 48, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casolari:2012:SRC

[2092] Sara Casolari, Michele Colajanni, and Stefania Tosi. Selective resource characterization for evaluation of system dynamics. ACM SIGMETRICS Performance Evaluation Review, 39(4):51– 60, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aceto:2012:RUE

[2093] Giuseppe Aceto and Antonio Pescapè. On the recent use of email through traffic and network analysis: the impact of OSNs, new trends, and other communication platforms. ACM SIGMET-RICS Performance Evaluation Review, 39(4):61–70, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Distefano:2012:DAB

[2094] Salvatore Distefano, Antonio Puliafito, and Kishor S. Trivedi. Dynamic aspects and behaviors of complex systems in performance and reliability assessment. ACM SIGMETRICS Performance Evaluation Review, 39(4):71–78, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mahmud:2012:CST

[2095] Nidhal Mahmud, Martin Walker, and Yiannis Papadopoulos. Compositional synthesis of temporal fault trees from state machines. ACM SIGMETRICS Performance Evaluation Review, 39(4): 79–88, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Abundo:2012:ACP

[2096] Marco Abundo, Valeria Cardellini, and Francesco Lo Presti. Admission control policies for a multi-class QoS-aware service oriented architecture. ACM SIGMETRICS Performance Evaluation Review, 39(4):89–98, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Persona:2012:HQM

[2097] Vittoria de Nitto Personà. Heuristics for QoS maintenance: adaptive policies in differentiated services wireless networks. ACM SIGMETRICS Performance Evaluation Review, 39(4):99– 107, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anceaume:2012:PEL

[2098] Emmanuelle Anceaume, Romaric Ludinard, and Bruno Sericola. Performance evaluation of large-scale dynamic systems. ACM SIGMETRICS Performance Evaluation Review, 39 (4):108–117, April 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Patel:2012:PIF

[2099] Naresh M. Patel. Performance implications of flash and storage class memories. ACM SIGMETRICS Performance Evaluation Review, 40(1):1– 2, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2012:HPC

[2100] Zhen Liu. High-performance computing in mobile services. ACM SIGMET-RICS Performance Evaluation Review, 40(1):3–4, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2012:DTM

[2101] Jian Tan, Xiaoqiao Meng, and Li Zhang. Delay tails in MapReduce scheduling. ACM SIGMETRICS Performance Evaluation Review, 40(1):5– 16, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shah:2012:OQS

[2102] Devavrat Shah, Neil Walton, and Yuan Zhong. Optimal queue-size scaling in switched networks. ACM SIGMET-RICS Performance Evaluation Review, 40(1):17–28, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hyytia:2012:MSH

[2103] Esa Hyytiä, Samuli Aalto, and Aleksi Penttinen. Minimizing slowdown in heterogeneous size-aware dispatching systems. ACM SIGMETRICS Performance Evaluation Review, 40(1):29–40, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Leconte:2012:BGS

[2104] Mathieu Leconte, Marc Lelarge, and Laurent Massoulié. Bipartite graph structures for efficient balancing of heterogeneous loads. ACM SIGMET-RICS Performance Evaluation Review, 40(1):41–52, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Atikoglu:2012:WAL

[2105] Berk Atikoglu, Yuehai Xu, Eitan Frachtenberg, Song Jiang, and Mike Paleczny. Workload analysis of a largescale key-value store. ACM SIGMET-RICS Performance Evaluation Review, 40(1):53–64, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shafiq:2012:FLC

[2106] Muhammad Zubair Shafiq, Lusheng Ji, Alex X. Liu, Jeffrey Pang, and Jia Wang. A first look at cellular machine-to-machine traffic: large scale measurement and characterization. ACM SIGMETRICS Performance Evaluation Review, 40(1):65–76, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Han:2012:BPB

[2107] Jinyoung Han, Seungbae Kim, Taejoong Chung, Ted Taekyoung Kwon, Hyun chul Kim, and Yanghee Choi. Bundling practice in BitTorrent: what, how, and why. ACM SIGMETRICS Performance Evaluation Review, 40(1): 77–88, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gan:2012:EEC

[2108] Lingwen Gan, Anwar Walid, and Steven Low. Energy-efficient congestion control. ACM SIGMETRICS Performance Evaluation Review, 40(1):89– 100, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jelenkovic:2012:UAD

[2109] Predrag R. Jelenkovic and Evangelia D. Skiani. Uniform approximation of the distribution for the number of retransmissions of bounded documents. ACM SIGMETRICS Performance Evaluation Review, 40(1):101–112, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

VanHoudt:2012:FLA

[2110] Benny Van Houdt and Luca Bortolussi. Fluid limit of an asynchronous optical packet switch with shared per link full range wavelength conversion. ACM SIGMETRICS Performance Evaluation Review, 40(1):113–124, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hua:2012:TOE

[2111] Nan Hua, Ashwin Lall, Baochun Li, and Jun Xu. Towards optimal errorestimating codes through the lens of Fisher information analysis. ACM SIGMETRICS Performance Evaluation Review, 40(1):125–136, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vulimiri:2012:HWC

[2112] Ashish Vulimiri, Gul A. Agha, Philip Brighten Godfrey, and Karthik Lakshminarayanan. How well can congestion pricing neutralize denial of service attacks? ACM SIGMET-RICS Performance Evaluation Review, 40(1):137–150, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Niu:2012:PCB

[2113] Di Niu, Chen Feng, and Baochun Li. Pricing cloud bandwidth reservations under demand uncertainty. ACM SIGMETRICS Performance Evaluation Review, 40(1):151–162, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

El-Sayed:2012:TMD

[2114] Nosayba El-Sayed, Ioan A. Stefanovici, George Amvrosiadis, Andy A. Hwang, and Bianca Schroeder. Temperature management in data centers: why some (might) like it hot. ACM SIGMET-RICS Performance Evaluation Review, 40(1):163–174, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2012:RCA

[2115] Zhenhua Liu, Yuan Chen, Cullen Bash, Adam Wierman, Daniel Gmach, Zhikui Wang, Manish Marwah, and Chris Hyser. Renewable and cooling aware workload management for sustainable data centers. ACM SIGMETRICS Performance Evaluation Review, 40 (1):175–186, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2012:ESD

[2116] Di Wang, Chuangang Ren, Anand Sivasubramaniam, Bhuvan Urgaonkar, and Hosam Fathy. Energy storage in datacenters: what, where, and how much? ACM SIGMETRICS Performance Evaluation Review, 40(1):187– 198, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shah:2012:RCU

[2117] Devavrat Shah and Tauhid Zaman. Rumor centrality: a universal source detector. ACM SIGMETRICS Performance Evaluation Review, 40(1):199– 210, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Netrapalli:2012:LGE

[2118] Praneeth Netrapalli and Sujay Sanghavi. Learning the graph of epidemic cascades. ACM SIGMETRICS Performance Evaluation Review, 40 (1):211–222, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Milling:2012:NFR

[2119] Chris Milling, Constantine Caramanis, Shie Mannor, and Sanjay Shakkottai. Network forensics: random infection vs spreading epidemic. ACM SIGMET-RICS Performance Evaluation Review, 40(1):223–234, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kim:2012:WGB

[2120] Hyojun Kim, Moonkyung Ryu, and Umakishore Ramachandran. What is a good buffer cache replacement scheme for mobile flash storage? ACM SIGMETRICS Performance Evaluation Review, 40(1):235–246, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Alizadeh:2012:VRL

[2121] Mohammad Alizadeh, Adel Javanmard, Shang-Tse Chuang, Sundar Iyer, and Yi Lu. Versatile refresh: low complexity refresh scheduling for high-throughput multi-banked eDRAM. ACM SIGMETRICS Performance Evaluation Review, 40(1):247– 258, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bhattacharya:2012:DLI

[2122] Suparna Bhattacharya, Karthick Rajamani, K. Gopinath, and Manish Gupta. Does lean imply green?: a study of the power performance implications of Java runtime bloat. ACM SIGMET-RICS Performance Evaluation Review, 40(1):259–270, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lim:2012:DFQ

[2123] Seung-Hwan Lim, Jae-Seok Huh, Youngjae Kim, Galen M. Shipman, and Chita R. Das. D-factor: a quantitative model of application slow-down in multi-resource shared systems. ACM SIGMETRICS Performance Evaluation Review, 40(1):271–282, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yoo:2012:AAD

[2124] Wucherl Yoo, Kevin Larson, Lee Baugh, Sangkyum Kim, and Roy H. Campbell. ADP: automated diagnosis of performance pathologies using hardware events. ACM SIGMET-RICS Performance Evaluation Review, 40(1):283–294, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xu:2012:PFS

[2125] Di Xu, Chenggang Wu, Pen-Chung Yew, Jianjun Li, and Zhenjiang Wang. Providing fairness on shared-memory multiprocessors via process scheduling. ACM SIGMETRICS Performance Evaluation Review, 40(1):295– 306, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Figueiredo:2012:CCT

[2126] Daniel Figueiredo, Philippe Nain, Bruno Ribeiro, Edmundo de Souza e Silva, and Don Towsley. Characterizing continuous time random walks on time varying graphs. ACM SIGMET-RICS Performance Evaluation Review, 40(1):307–318, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2012:BRW

[2127] Chul-Ho Lee, Xin Xu, and Do Young Eun. Beyond random walk and Metropolis–Hastings samplers: why you should not backtrack for unbiased graph sampling. ACM SIGMET-RICS Performance Evaluation Review, 40(1):319–330, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Song:2012:CEM

[2128] Han Hee Song, Berkant Savas, Tae Won Cho, Vacha Dave, Zhengdong Lu, Inderjit S. Dhillon, Yin Zhang, and Lili Qiu. Clustered embedding of massive social networks. ACM SIGMET-RICS Performance Evaluation Review, 40(1):331–342, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cohen:2012:DLN

[2129] Edith Cohen, Graham Cormode, and Nick Duffield. Don’t let the negatives bring you down: sampling from streams of signed updates. ACM SIGMETRICS Performance Evaluation Review, 40(1):343–354, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ammar:2012:ERA

[2130] Ammar Ammar and Devavrat Shah.

Efficient rank aggregation using partial data. ACM SIGMETRICS Performance Evaluation Review, 40(1):355– 366, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Duffield:2012:FSA

[2131] Nick Duffield. Fair sampling across network flow measurements. ACM SIGMETRICS Performance Evaluation Review, 40(1):367–378, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Peng:2012:TBN

[2132] Kunyang Peng and Qunfeng Dong.

TCAM-based NFA implementation. ACM SIGMETRICS Performance Evaluation Review, 40(1):379–380, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anshelevich:2012:SEP

[2133] Elliot Anshelevich, Ameya Hate, Koushik Kar, and Michael Usher. Stable and efficient pricing for interdomain traffic forwarding. ACM SIG-METRICS Performance Evaluation Review, 40(1):381–382, June 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

DiCioccio:2012:MCH

[2134] Lucas DiCioccio, Renata Teixeira, and Catherine Rosenberg. Measuring and characterizing home networks. ACM SIGMETRICS Performance Evaluation Review, 40(1):383–384, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sommers:2012:CMA

[2135] Joel Sommers and Paul Barford. Comparing metro-area cellular and WiFi performance: extended abstract. ACM SIGMETRICS Performance Evaluation Review, 40(1):385–386, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nemeth:2012:TSC

[2136] Gábor Németh and Gábor Rétvári. Towards a statistical characterization of the competitiveness of oblivious routing. ACM SIGMETRICS Performance Evaluation Review, 40(1):387– 388, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zarifzadeh:2012:RT

[2137] Sajjad Zarifzadeh, Madhwaraj G. K., and Constantine Dovrolis. Range tomography. ACM SIGMETRICS Performance Evaluation Review, 40(1): 389–390, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2012:SAM

[2138] Myungjin Lee, Nick Duffield, and Ramana Rao Kompella. A scalable architecture for maintaining packet latency measurements. ACM SIGMET-RICS Performance Evaluation Review, 40(1):391–392, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Laner:2012:MRN

[2139] Markus Laner, Philipp Svoboda, and Markus Rupp. Modeling randomness in network traffic. ACM SIGMET-RICS Performance Evaluation Review, 40(1):393–394, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gallo:2012:PER

[2140] Massimo Gallo, Bruno Kauffmann, Luca Muscariello, Alain Simonian, and Christian Tanguy. Performance evaluation of the random replacement policy for networks of caches. ACM SIGMET-RICS Performance Evaluation Review, 40(1):395–396, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mukherjee:2012:SCT

[2141] Koyel Mukherjee, Samir Khuller, and Amol Deshpande. Saving on cooling: the thermal scheduling problem. ACM SIGMETRICS Performance Evaluation Review, 40(1):397–398, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bodas:2012:CCM

[2142] Shreeshankar Bodas, Devavrat Shah, and Damon Wischik. Congestion control meets medium access: throughput, delay, and complexity. ACM SIGMET-RICS Performance Evaluation Review, 40(1):399–400, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tantawi:2012:OCP

[2143] Asser N. Tantawi. Optimized cloud placement of virtual clusters using biased importance sampling. ACM SIGMETRICS Performance Evaluation Review, 40(1):401–402, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shen:2012:PEC

[2144] Kai Shen, Arrvindh Shriraman, Sandhya Dwarkadas, and Xiao Zhang. Power and energy containers for multicore servers. ACM SIGMETRICS Performance Evaluation Review, 40 (1):403–404, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2012:CIW

[2145] Kai Wang, Minghong Lin, Florin Ciucu, Adam Wierman, and Chuang Lin. Characterizing the impact of the workload on the value of dynamic resizing in data centers. ACM SIGMET-RICS Performance Evaluation Review, 40(1):405–406, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2012:PLSa

[2146] Yue Tan, Yingdong Lu, and Cathy H. Xia. Provisioning for large scale cloud computing services. ACM SIGMET-RICS Performance Evaluation Review, 40(1):407–408, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Narayana:2012:DWA

[2147] Srinivas Narayana, Joe Wenjie Jiang, Jennifer Rexford, and Mung Chiang. Distributed wide-area traffic management for cloud services. ACM SIGMETRICS Performance Evaluation Review, 40(1):409–410, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dixit:2012:EFG

[2148] Advait Abhay Dixit, Pawan Prakash, Ramana Rao Kompella, and Charlie Hu. On the efficacy of finegrained traffic splitting protocols in data center networks. ACM SIGMET-RICS Performance Evaluation Review, 40(1):411–412, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Frank:2012:CAT

[2149] Benjamin Frank, Ingmar Poese, Georgios Smaragdakis, Steve Uhlig, and Anja Feldmann. Content-aware traffic engineering. ACM SIGMETRICS Performance Evaluation Review, 40 (1):413–414, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hu:2012:UPA

[2150] Jian Hu, Hong Jiang, and Prakash Manden. Understanding performance anomalies of SSDs and their impact in enterprise application environment. ACM SIGMETRICS Performance Evaluation Review, 40(1):415– 416, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Glatz:2012:CIO

[2151] Eduard Glatz and Xenofontas Dimitropoulos. Classifying Internet oneway traffic. ACM SIGMETRICS Performance Evaluation Review, 40 (1):417–418, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arora:2012:FCE

[2152] Manish Arora, Feng Wang, Bob Rychlik, and Dean Tullsen. Fast cost efficient designs by building upon the Plackett and Burman method. ACM SIGMETRICS Performance Evaluation Review, 40(1):419–420, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Keller:2012:MHN

[2153] Matthias Keller, Jan Beutel, and Lothar Thiele. Multi-hop network tomography: path reconstruction and per-hop arrival time estimation from partial information. ACM SIGMET-RICS Performance Evaluation Review, 40(1):421–422, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Papapanagiotou:2012:SVL

[2154] Ioannis Papapanagiotou, Erich M. Nahum, and Vasileios Pappas. Smartphones vs. laptops: comparing Web browsing behavior and the implications for caching. ACM SIGMET-RICS Performance Evaluation Review, 40(1):423–424, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reinecke:2012:MMV

[2155] Philipp Reinecke, Miklós Telek, and Katinka Wolter. Micro and macro views of discrete-state Markov models and their application to efficient simulation with phase-type distributions. ACM SIGMETRICS Performance Evaluation Review, 40(1):425– 426, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bertran:2012:PFB

[2156] Ramon Bertran, Marc Gonzàlez, Xavier Martorell, Nacho Navarro, and Eduard Ayguadé. POTRA: a framework for building power models for next generation multicore architectures. ACM SIGMETRICS Performance Evaluation Review, 40(1):427– 428, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hayden:2012:BTS

[2157] Richard A. Hayden. Basic theory and some applications of martingales. ACM SIGMETRICS Performance Evaluation Review, 40(1):429–430, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

deSouzaeSilva:2012:AML

[2158] Edmundo de Souza e Silva and Daniel Sadoc Menasche. Applications of machine learning to performance evaluation. ACM SIGMET-RICS Performance Evaluation Review, 40(1):431–432, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aikat:2012:INE

[2159] Jay Aikat and Kevin Jeffay. Introduction to network experiments using the GENI cyberinfrastructure. ACM SIGMETRICS Performance Evaluation Review, 40(1):433–434, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Eriksson:2012:PLA

[2160] Brian Eriksson, Paul Barford, Bruce Maggs, and Robert Nowak. Posit: a lightweight approach for IP geolocation. ACM SIGMETRICS Performance Evaluation Review, 40(2):2–11, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coucheney:2012:CSE

[2161] Pierre Coucheney, Patrick Maillé, and Bruno Tuffin. Comparison of search engines non-neutral and neutral behaviors. ACM SIGMETRICS Performance Evaluation Review, 40(2): 14–17, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hanawal:2012:GTA

[2162] Manjesh K. Hanawal, Eitan Altman, and Rajesh Sundaresan. Game theoretic analysis of collusions in nonneutral networks. ACM SIGMET-RICS Performance Evaluation Review, 40(2):18–21, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yu:2012:GUW

[2163] Seung Min Yu and Seong-Lyun Kim. Guaranteeing user welfare in network service: comparison of two subsidy schemes. ACM SIGMETRICS Performance Evaluation Review, 40(2): 22–25, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berry:2012:NMC

[2164] R. Berry, M. Honig, T. Nguyen, V. Subramanian, H. Zhou, and R. Vohra. Newsvendor model of capacity sharing. ACM SIGMETRICS Performance Evaluation Review, 40 (2):26–29, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ma:2012:PDK

[2165] Richard T. B. Ma, Dah Ming Chiu, John C. S. Lui, Vishal Misra, and Dan Rubenstein. Price differentiation in the Kelly mechanism. ACM SIGMET-RICS Performance Evaluation Review, 40(2):30–33, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Houidi:2012:PTB

[2166] Zied Ben Houidi and Helia Pouyllau. The price of tussles: bankrupt in cyberspace? ACM SIGMET-RICS Performance Evaluation Review, 40(2):34–37, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lodhi:2012:PSA

[2167] Aemen Lodhi, Amogh Dhamdhere, and Constantine Dovrolis. Peering strategy adoption by transit providers in the Internet: a game theoretic approach? ACM SIGMETRICS Performance Evaluation Review, 40(2):

38–41, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mastroeni:2012:PIP

[2168] Loretta Mastroeni and Maurizio Naldi. Pricing of insurance policies against cloud storage price rises. ACM SIGMETRICS Performance Evaluation Review, 40(2):42–45, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2012:IVI

[2169] Dongmyung Lee, Jeonghoon Mo, and Jinwoo Park. ISP vs. ISP + CDN: can ISPs in duopoly profit by introducing CDN services? ACM SIGMET-RICS Performance Evaluation Review, 40(2):46–48, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gulyas:2012:GNF

[2170] András Gulyás, Attila Korösi, Dávid Szabó, and Gergely Biczók. On greedy network formation. ACM SIGMET-RICS Performance Evaluation Review, 40(2):49–52, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ramakrishnan:2012:EIV

[2171] Lavanya Ramakrishnan, R. Shane Canon, Krishna Muriki, Iwona Sakrejda, and Nicholas J. Wright. Evaluating interconnect and virtualization performance for high performance computing. ACM SIGMETRICS Performance Evaluation Review, 40(2): 55–60, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mudalige:2012:PMA

[2172] G. R. Mudalige, M. B. Giles, C. Bertolli, and P. H. J. Kelly. Predictive modeling and analysis of OP2 on distributed memory GPU clusters. ACM SIGMETRICS Performance Evaluation Review, 40(2):61–67, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mateescu:2012:OMT

[2173] Gabriel Mateescu, Gregory H. Bauer, and Robert A. Fiedler. Optimizing matrix transposes using a POWER7 cache model and explicit prefetching. ACM SIGMETRICS Performance Evaluation Review, 40(2):68–73, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Danalis:2012:BPH

[2174] Anthony Danalis, Piotr Luszczek, Gabriel Marin, Jeffrey S. Vetter, and Jack Dongarra. Blackjack-Bench: portable hardware characterization. ACM SIGMETRICS Performance Evaluation Review, 40(2):74–79, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tineo:2012:TAA

[2175] Adrian Tineo, Sadaf R. Alam, and Thomas C. Schulthess. Towards autotuning by alternating communication methods. ACM SIGMETRICS Performance Evaluation Review, 40 (2):80–85, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Iakymchuk:2012:MPT

[2176] Roman Iakymchuk and Paolo Bientinesi. Modeling performance through memory-stalls. ACM SIGMETRICS Performance Evaluation Review, 40 (2):86–91, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shan:2012:PEH

[2177] Hongzhang Shan, Nicholas J. Wright, John Shalf, Katherine Yelick, Marcus Wagner, and Nathan Wichmann. A preliminary evaluation of the hardware acceleration of the Cray Gemini interconnect for PGAS languages and comparison with MPI. ACM SIGMET-RICS Performance Evaluation Review, 40(2):92–98, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Deshpande:2012:AGC

[2178] Vivek Deshpande, Xing Wu, and Frank Mueller. Auto-generation of communication benchmark traces. ACM SIGMETRICS Performance Evaluation Review, 40(2):99–105, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Su:2012:CPB

[2179] ChunYi Su, Dong Li, Dimitrios S. Nikolopoulos, Matthew Grove, Kirk Cameron, and Bronis R. de Supinski. Critical path-based thread placement for NUMA systems. ACM SIGMET-RICS Performance Evaluation Review, 40(2):106–112, September 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2012:BMD

[2180] DongJin Lee, Michael O’Sullivan, and Cameron Walker. Benchmarking and modeling disk-based storage tiers for practical storage design. ACM SIGMETRICS Performance Evaluation Review, 40(2):113–118, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2012:TEG

[2181] Lingyuan Wang, Miaoqing Huang, and Tarek El-Ghazawi. Towards efficient GPU sharing on multicore processors. ACM SIGMETRICS Performance Evaluation Review, 40(2): 119–124, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sun:2012:APM

[2182] Xian-He Sun and Dawei Wang. APC:

a performance metric of memory systems. ACM SIGMETRICS Performance Evaluation Review, 40(2):125– 130, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vitali:2012:LSO

[2183] Roberto Vitali, Alessandro Pellegrini, and Francesco Quaglia. Load sharing for optimistic parallel simulations on multi core machines. ACM SIGMET-RICS Performance Evaluation Review, 40(3):2–11, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hahnel:2012:MEC

[2184] Marcus Hähnel, Björn Döbel, Marcus Völp, and Hermann Härtig. Measuring energy consumption for short code paths using RAPL. ACM SIGMET-RICS Performance Evaluation Review, 40(3):13–17, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mazzucco:2012:EEP

[2185] Michele Mazzucco and Isi Mitrani. Empirical evaluation of power saving policies for data centers. ACM SIGMET-RICS Performance Evaluation Review, 40(3):18–22, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghumre:2012:ENC

[2186] Pooja Ghumre, Junwei Li, Mukil Kesavan, Ada Gavrilovska, and Karsten Schwan. Evaluating the need for complexity in energy-aware management for cloud platforms. ACM SIGMET-RICS Performance Evaluation Review, 40(3):23–27, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gast:2012:OSP

[2187] Nicolas Gast, Dan-Cristian Tomozei, and Jean-Yves Le Boudec. Optimal storage policies with wind forecast uncertainties. ACM SIGMET-RICS Performance Evaluation Review, 40(3):28–32, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bernstein:2012:SAP

[2188] Andrey Bernstein, Daniel Bienstock, David Hay, Meric Uzunoğlu, and Gil Zussman. Sensitivity analysis of the power grid vulnerability to large-scale cascading failures. ACM SIGMET-RICS Performance Evaluation Review, 40(3):33–37, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ardakanian:2012:RDC

[2189] O. Ardakanian, C. Rosenberg, and S. Keshav. RealTime distributed congestion control for electrical vehicle charging. ACM SIGMETRICS Performance Evaluation Review, 40(3):38–42, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ardakanian:2012:ISR

[2190] Omid Ardakanian, Catherine Rosenberg, and S. Keshav. On the impact of storage in residential power distribution systems. ACM SIGMET-RICS Performance Evaluation Review, 40(3):43–47, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chiu:2012:EGB

[2191] David Chiu, Christopher Stewart, and Bart McManus. Electric grid balancing through low-cost workload migration. ACM SIGMETRICS Performance Evaluation Review, 40(3):48–52, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Menasche:2012:SAP

[2192] Daniel S. Menasché, Rosa Maria Meri Leäo, Edmundo de Souza e Silva, Alberto Avritzer, Sindhu Suresh, Kishor Trivedi, Raymond A. Marie, Lucia Happe, and Anne Koziolek. Survivability analysis of power distribution in smart grids with active and reactive power modeling. ACM SIGMET-RICS Performance Evaluation Review, 40(3):53–57, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coffman:2012:UDA

[2193] E. G. Coffman, Jr., Y. Kogan, W. Lai, and V. Ramaswami. Uptime and downtime analysis for hierarchical redundant systems in telecommunications. ACM SIGMETRICS Performance Evaluation Review, 40(3):59–61, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Avrachenkov:2012:OCC

[2194] K. Avrachenkov, U. Ayesta, J. Doncel, and P. Jacko. Optimal congestion control of TCP flows for Internet routers. ACM SIGMETRICS Performance Evaluation Review, 40 (3):62–64, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schorgendorfer:2012:TLB

[2195] Angela Schörgendorfer, Peter M. van de Ven, and Bo Zhang. Temporal load balancing for distributed backup scheduling. ACM SIGMET-RICS Performance Evaluation Review, 40(3):65–67, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rochman:2012:ERM

[2196] Yuval Rochman, Hanoch Levy, and Eli Brosh. Efficient replication in multi-regional peer-supported VoD systems. ACM SIGMETRICS Performance Evaluation Review, 40(3):68–70, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Borgs:2012:PQ

[2197] Christian Borgs, Jennifer T. Chayes, Sherwin Doroudi, Mor Harchol-Balter, and Kuang Xu. Pricing and queueing. ACM SIGMETRICS Performance Evaluation Review, 40(3):71–73, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Godtschalk:2012:SBR

[2198] Antonie S. Godtschalk and Florin Ciucu. Stochastic bounds for randomized load balancing. ACM SIGMET-RICS Performance Evaluation Review, 40(3):74–76, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Myers:2012:EQL

[2199] Daniel S. Myers and Mary K. Vernon.

Estimating queue length distributions for queues with random arrivals. ACM SIGMETRICS Performance Evaluation Review, 40(3):77–79, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cremonesi:2012:MRT

[2200] Paolo Cremonesi and Andrea Sansottera. Modeling response times in the Google ROADEF/EURO challenge. ACM SIGMETRICS Performance Evaluation Review, 40(3):80–82, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2012:PLSb

[2201] Yue Tan, Yingdong Lu, and Cathy H.

Xia. Provisioning for large scale loss network systems with applications in cloud computing. ACM SIGMET-RICS Performance Evaluation Review, 40(3):83–85, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pal:2012:CCT

[2202] Ranjan Pal and Pan Hui. CyberInsurance for cybersecurity a topological take on modulating insurance premiums. ACM SIGMETRICS Performance Evaluation Review, 40(3):86–88, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Elahi:2012:MFD

[2203] Maryam Elahi, Carey Williamson, and Philipp Woelfel. Meeting the fairness deadline in speed scaling systems: is turbocharging enough? ACM SIGMETRICS Performance Evaluation Review, 40(3):89–91, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bachmat:2012:ASQ

[2204] Eitan Bachmat and Assaf Natanzon. Analysis of SITA queues with many servers and spacetime geometry. ACM SIGMETRICS Performance Evaluation Review, 40(3):92–94, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2012:RSS

[2205] Thomas Bonald and Davide Cuda. RateOptimal scheduling schemes for asynchronous InputQueued packet switches. ACM SIGMETRICS Performance Evaluation Review, 40(3):95–97, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:2012:OOS

[2206] Minghong Lin, Adam Wierman, Alan Roytman, Adam Meyerson, and Lachlan L. H. Andrew. Online optimization with switching cost. ACM SIGMET-RICS Performance Evaluation Review, 40(3):98–100, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Blaszczyszyn:2012:FVW

[2207] B. Blaszczyszyn and K. Gaurav. Farout vertices in weighted repeated configuration model. ACM SIGMETRICS Performance Evaluation Review, 40(3): 100–103, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Papadopoulos:2012:RGG

[2208] Fragkiskos Papadopoulos, Constantinos Psomas, and Dmitri Krioukov. Replaying the geometric growth of complex networks and application to the AS Internet. ACM SIGMETRICS Performance Evaluation Review, 40(3): 104–106, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tizghadam:2012:NCV

[2209] Ali Tizghadam, Weiwei Li, and Alberto Leon-Garcia. Network criticality in vehicular networks. ACM SIGMET-RICS Performance Evaluation Review, 40(3):107–109, December 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lui:2013:SPC

[2210] John C. S. Lui and Li Zhang. A study of pricing for cloud resources. ACM SIGMETRICS Performance Evaluation Review, 40(4):3–12, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2013:SCI

[2211] Zhizhong Zhang, Chuan Wu, and David W. L. Cheung. A survey on cloud interoperability: taxonomies, standards, and practice. ACM SIGMETRICS Performance Evaluation Review, 40(4):13–22, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2013:FPE

[2212] Lei Yang, Jiannong Cao, Yin Yuan, Tao Li, Andy Han, and Alvin Chan. A framework for partitioning and execution of data stream applications in mobile cloud computing. ACM SIGMET-RICS Performance Evaluation Review, 40(4):23–32, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2013:TOA

[2213] Weina Wang, Kai Zhu, Lei Ying, Jian Tan, and Li Zhang. A throughput optimal algorithm for map task scheduling in MapReduce with data locality. ACM SIGMETRICS Performance Evaluation Review, 40(4):33–42, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Huang:2013:ESC

[2214] Qun Huang and Patrick P. C. Lee. An experimental study of cascading performance interference in a virtualized environment. ACM SIGMET-RICS Performance Evaluation Review, 40(4):43–52, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Singh:2013:AMW

[2215] Rahul Singh, Prashant Shenoy, Maitreya Natu, Vaishali Sadaphal, and Harrick Vin. Analytical modeling for what-if analysis in complex cloud computing applications. ACM SIGMET-RICS Performance Evaluation Review, 40(4):53–62, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2013:DCR

[2216] Jia Liu, Cathy H. Xia, Ness B. Shroff, and Xiaodong Zhang. On distributed computation rate optimization for deploying cloud computing programming frameworks. ACM SIGMETRICS Performance Evaluation Review, 40(4):63– 72, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2013:MEV

[2217] Giuliano Casale and Mirco Tribastone. Modelling exogenous variability in cloud deployments. ACM SIGMET-RICS Performance Evaluation Review, 40(4):73–82, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mahmood:2013:TNE

[2218] Shah Mahmood and Yvo Desmedt.

Two new economic models for privacy. ACM SIGMETRICS Performance Evaluation Review, 40(4):84–89, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hutton:2013:AEP

[2219] Luke Hutton and Tristan Henderson. An architecture for ethical and privacy-sensitive social network experiments. ACM SIGMETRICS Performance Evaluation Review, 40(4):90–95, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:2013:LCI

[2220] Minaxi Gupta, Yuqing (Melanie) Wu, Swapnil S. Joshi, Aparna Tiwari, Ashish Nair, and Ezhilan Ilangovan. On the linkability of complementary information from free versions of people databases. ACM SIGMETRICS Performance Evaluation Review, 40 (4):96–100, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tschorsch:2013:HBT

[2221] Florian Tschorsch and Björn Scheuermann. How (not) to build a transport layer for anonymity overlays. ACM SIGMETRICS Performance Evaluation Review, 40(4):101–106, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Prabhakar:2013:DLS

[2222] Balaji Prabhakar. Designing large-scale nudge engines. ACM SIGMETRICS Performance Evaluation Review, 41(1): 1–2, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Maltz:2013:CCS

[2223] David A. Maltz. Challenges in cloud scale data centers. ACM SIGMET-RICS Performance Evaluation Review, 41(1):3–4, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhou:2013:PCG

[2224] Xia Zhou, Zengbin Zhang, Gang Wang, Xiaoxiao Yu, Ben Y. Zhao, and Haitao Zheng. Practical conflict graphs for dynamic spectrum distribution. ACM SIGMETRICS Performance Evaluation Review, 41(1):5–16, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shafiq:2013:FLC

[2225] Muhammad Zubair Shafiq, Lusheng Ji, Alex X. Liu, Jeffrey Pang, Shobha Venkataraman, and Jia Wang. A first look at cellular network performance during crowded events. ACM SIGMETRICS Performance Evaluation Review, 41(1):17–28, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ding:2013:CMI

[2226] Ning Ding, Daniel Wagner, Xiaomeng Chen, Abhinav Pathak, Y. Charlie Hu, and Andrew Rice. Characterizing and modeling the impact of wireless signal strength on smartphone battery drain. ACM SIGMETRICS Performance Evaluation Review, 41(1):29–40, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stolyar:2013:LSS

[2227] Alexander L. Stolyar and Yuan Zhong.

A large-scale service system with packing constraints: minimizing the number of occupied servers. ACM SIGMETRICS Performance Evaluation Review, 41(1):41–52, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2013:OEG

[2228] Lian Lu, Jinlong Tu, Chi-Kin Chau, Minghua Chen, and Xiaojun Lin. Online energy generation scheduling for microgrids with intermittent energy sources and co-generation. ACM SIGMETRICS Performance Evaluation Review, 41(1):53–66, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shanmuganathan:2013:DCU

[2229] Ganesha Shanmuganathan, Ajay Gulati, and Peter Varman. Defragmenting the cloud using demand-based resource allocation. ACM SIGMET-RICS Performance Evaluation Review, 41(1):67–80, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Karger:2013:ECM

[2230] David R. Karger, Sewoong Oh, and Devavrat Shah. Efficient crowdsourcing for multi-class labeling. ACM SIGMETRICS Performance Evaluation Review, 41(1):81–92, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kim:2013:RCD

[2231] Myunghwan Kim, Roshan Sumbaly, and Sam Shah. Root cause detection in a service-oriented architecture. ACM SIGMETRICS Performance Evaluation Review, 41(1):93–104, June 2013.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jaggard:2013:DSP

[2232] Aaron D. Jaggard, Swara Kopparty, Vijay Ramachandran, and Rebecca N. Wright. The design space of probing algorithms for network-performance measurement. ACM SIGMETRICS Performance Evaluation Review, 41 (1):105–116, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bouman:2013:DMT

[2233] Niek Bouman, Sem Borst, and Johan van Leeuwaarden. Delays and mixing times in random-access networks. ACM SIGMETRICS Performance Evaluation Review, 41(1):117–128, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cecchi:2013:SUM

[2234] Fabio Cecchi and Peter Jacko. Scheduling of users with Markovian timevarying transmission rates. ACM SIGMETRICS Performance Evaluation Review, 41(1):129–140, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Simatos:2013:LID

[2235] Florian Simatos, Niek Bouman, and Sem Borst. Lingering issues in distributed scheduling. ACM SIGMET-RICS Performance Evaluation Review, 41(1):141–152, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gandhi:2013:EAM

[2236] Anshul Gandhi, Sherwin Doroudi, Mor Harchol-Balter, and Alan Scheller-Wolf. Exact analysis of the M/M/k/ setup class of Markov chains via recursive renewal reward. ACM SIGMET-RICS Performance Evaluation Review, 41(1):153–166, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tsitsiklis:2013:QST

[2237] John N. Tsitsiklis and Kuang Xu. Queueing system topologies with limited flexibility. ACM SIGMETRICS Performance Evaluation Review, 41 (1):167–178, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2013:SML

[2238] Yongkun Li, Patrick P. C. Lee, and John C. S. Lui. Stochastic modeling of large-scale solid-state storage systems: analysis, design tradeoffs and optimization. ACM SIGMET-RICS Performance Evaluation Review, 41(1):179–190, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

VanHoudt:2013:MFM

[2239] Benny Van Houdt. A mean field model for a class of garbage collection algorithms in flash-based solid state drives. ACM SIGMETRICS Performance Evaluation Review, 41(1):191– 202, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jung:2013:RWH

[2240] Myoungsoo Jung and Mahmut Kandemir. Revisiting widely held SSD expectations and rethinking systemlevel implications. ACM SIGMET-RICS Performance Evaluation Review, 41(1):203–216, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cintra:2013:CIP

[2241] Marcelo Cintra and Niklas Linkewitsch. Characterizing the impact of process variation on write endurance enhancing techniques for non-volatile memory systems. ACM SIGMET-RICS Performance Evaluation Review, 41(1):217–228, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sharma:2013:DCS

[2242] Abhigyan Sharma, Arun Venkataramani, and Ramesh K. Sitaraman. Distributing content simplifies ISP traffic engineering. ACM SIGMETRICS Performance Evaluation Review, 41 (1):229–242, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Valancius:2013:QBJ

[2243] Vytautas Valancius, Bharath Ravi, Nick Feamster, and Alex C. Snoeren. Quantifying the benefits of joint content and network routing. ACM SIGMETRICS Performance Evaluation Review, 41(1):243–254, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Simha:2013:HTL

[2244] Dilip Nijagal Simha, Tzi cker Chiueh, Ganesh Karuppur Rajagopalan, and Pallav Bose. High-throughput lowlatency fine-grained disk logging. ACM SIGMETRICS Performance Evaluation Review, 41(1):255–266, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tudor:2013:UEC

[2245] Bogdan Marius Tudor and Yong Meng Teo. On understanding the energy consumption of ARM-based multicore servers. ACM SIGMETRICS Performance Evaluation Review, 41(1):267– 278, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sen:2013:RBO

[2246] Rathijit Sen and David A. Wood. Reuse-based online models for caches. ACM SIGMETRICS Performance Evaluation Review, 41(1):279–292, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shahzad:2013:POT

[2247] Muhammad Shahzad and Alex X. Liu. Probabilistic optimal tree hopping for RFID identification. ACM SIGMET-RICS Performance Evaluation Review, 41(1):293–304, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Peng:2013:MTA

[2248] Qiuyu Peng, Anwar Walid, and Steven H. Low. Multipath TCP algorithms: theory and design. ACM SIGMETRICS Performance Evaluation Review, 41(1):305–316, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2013:TAU

[2249] Guang Tan, Zhimeng Yin, and Hongbo Jiang. Trap array: a unified model for scalability evaluation of geometric routing. ACM SIGMETRICS Performance Evaluation Review, 41(1):317– 328, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Andrew:2013:TTM

[2250] Lachlan Andrew, Siddharth Barman, Katrina Ligett, Minghong Lin, Adam Meyerson, Alan Roytman, and Adam Wierman. A tale of two metrics: simultaneous bounds on competitiveness and regret. ACM SIGMET-RICS Performance Evaluation Review, 41(1):329–330, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yu:2013:AGA

[2251] Zhibin Yu, Lieven Eeckhout, Nilanjan Goswami, Tao Li, Lizy John, Hai Jin, and Chengzhong Xu. Accelerating GPGPU architecture simulation. ACM SIGMETRICS Performance Evaluation Review, 41(1):331–332, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2013:AAC

[2252] Di Wang, Chuangang Ren, Sriram Govindan, Anand Sivasubramaniam, Bhuvan Urgaonkar, Aman Kansal, and Kushagra Vaid. ACE: abstracting, characterizing and exploiting peaks and valleys in datacenter power consumption. ACM SIGMETRICS Performance Evaluation Review, 41(1): 333–334, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Potharaju:2013:EAI

[2253] Rahul Potharaju and Navendu Jain. An empirical analysis of intra- and inter-datacenter network failures for geo-distributed services. ACM SIG-METRICS Performance Evaluation Review, 41(1):335–336, June 2013. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mazauric:2013:CAC

[2254] Dorian Mazauric, Saleh Soltan, and Gil Zussman. Computational analysis of cascading failures in power networks. ACM SIGMETRICS Performance Evaluation Review, 41(1):337– 338, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nelson:2013:DCA

[2255] John C. Nelson, Jonathan Connell, Canturk Isci, and Jonathan Lenchner. Data center asset tracking using a mobile robot. ACM SIGMET-RICS Performance Evaluation Review, 41(1):339–340, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2013:DCD

[2256] Zhenhua Liu, Adam Wierman, Yuan Chen, Benjamin Razon, and Niangjun Chen. Data center demand response: avoiding the coincident peak via workload shifting and local generation. ACM SIGMETRICS Performance Evaluation Review, 41(1):341– 342, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Saez:2013:DFP

[2257] Juan Carlos Sáez, Fernando Castro, Daniel Chaver, and Manuel Prieto. Delivering fairness and priority enforcement on asymmetric multicore systems via OS scheduling. ACM SIGMET-RICS Performance Evaluation Review, 41(1):343–344, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arvidsson:2013:DUD

[2258] Ake Arvidsson and Ying Zhang. Detecting user dissatisfaction and understanding the underlying reasons. ACM SIGMETRICS Performance Evaluation Review, 41(1):345–346, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kong:2013:DMD

[2259] Deguang Kong and Guanhua Yan. Discriminant malware distance learning on structural information for automated malware classification. ACM SIGMET-RICS Performance Evaluation Review, 41(1):347–348, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Peserico:2013:EP

[2260] Enoch Peserico. Elastic paging. ACM SIGMETRICS Performance Evaluation Review, 41(1):349–350, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gan:2013:ECR

[2261] Lingwen Gan, Na Li, Steven Low, and Ufuk Topcu. Exact convex relaxation for optimal power flow in distribution networks. ACM SIGMET-RICS Performance Evaluation Review, 41(1):351–352, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kwak:2013:EPR

[2262] Jaewook Kwak, Chul-Ho Lee, and Do Young Eun. Exploiting the past to reduce delay in CSMA scheduling: a high-order Markov chain approach. ACM SIGMETRICS Performance Evaluation Review, 41(1):353– 354, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Paredes-Oliva:2013:FFR

[2263] Ignasi Paredes-Oliva, Pere Barlet-Ros, and Xenofontas Dimitropoulos. FaRNet: fast recognition of high multi-dimensional network traffic patterns. ACM SIGMETRICS Performance Evaluation Review, 41(1):355– 356, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghiassi-Farrokhfal:2013:FSP

[2264] Yashar Ghiassi-Farrokhfal, Srinivasan Keshav, Catherine Rosenberg, and Florin Ciucu. Firming solar power. ACM SIGMETRICS Performance Evaluation Review, 41(1):357–358, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2013:GNL

[2265] Yi Wang, Dongzhe Tai, Ting Zhang, Jianyuan Lu, Boyang Xu, Huichen Dai, and Bin Liu. Greedy name lookup for named data networking. ACM SIGMETRICS Performance Evaluation Review, 41(1):359–360, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dong:2013:HDE

[2266] Mian Dong, Tian Lan, and Lin Zhong.

How does energy accounting matter for energy management? ACM SIGMET-RICS Performance Evaluation Review, 41(1):361–362, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Moharir:2013:OLB

[2267] Sharayu Moharir, Sujay Sanghavi, and Sanjay Shakkottai. Online load balancing under graph constraints. ACM SIGMETRICS Performance Evaluation Review, 41(1):363–364, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kambadur:2013:PSP

[2268] Melanie Kambadur, Kui Tang, Joshua Lopez, and Martha A. Kim. Parallel scaling properties from a basic block view. ACM SIGMETRICS Performance Evaluation Review, 41(1):365– 366, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ciucu:2013:SBS

[2269] Florin Ciucu, Felix Poloczek, and Jens Schmitt. Sharp bounds in stochastic network calculus. ACM SIGMET-RICS Performance Evaluation Review, 41(1):367–368, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhu:2013:SSU

[2270] Ji Zhu, Stratis Ioannidis, Nidhi Hegde, and Laurent Massoulie. Stable and scalable universal swarms. ACM SIGMETRICS Performance Evaluation Review, 41(1):369–370, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Paschos:2013:SSP

[2271] Georgios S. Paschos and Leandros Tassiulas. Sustainability of service provisioning systems under attack. ACM SIGMETRICS Performance Evaluation Review, 41(1):371–372, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xu:2013:TAW

[2272] Hong Xu, Chen Feng, and Baochun Li. Temperature aware workload management in geo-distributed datacenters. ACM SIGMETRICS Performance Evaluation Review, 41(1):373– 374, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2013:TPH

[2273] Ming Li, Andrey Lukyanenko, Sasu Tarkoma, Yong Cui, and Antti Ylä-Jääski. Tolerating path heterogeneity in multipath TCP with bounded receive buffers. ACM SIGMETRICS Performance Evaluation Review, 41 (1):375–376, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dai:2013:UAC

[2274] Chen Dai, Chao Lv, Jiaxin Li, Weihua Zhang, and Binyu Zang. Understanding architectural characteristics of multimedia retrieval workloads. ACM SIGMETRICS Performance Evaluation Review, 41(1):377–378, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Balachandran:2013:UIV

[2275] Athula Balachandran, Vyas Sekar, Aditya Akella, and Srinivasan Seshan. Understanding Internet video viewing behavior in the wild. ACM SIGMET-RICS Performance Evaluation Review, 41(1):379–380, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jiang:2013:USS

[2276] Nan Jiang, Yu Jin, Ann Skudlark, and Zhi-Li Zhang. Understanding SMS spam in a large cellular network. ACM SIGMETRICS Performance Evaluation Review, 41(1):381–382, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sundaresan:2013:WPB

[2277] Srikanth Sundaresan, Nazanin Magharei, Nick Feamster, Renata Teixeira, and Sam Crawford. Web performance bottlenecks in broadband access networks. ACM SIGMETRICS Performance Evaluation Review, 41(1):383– 384, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aguilera:2013:TGR

[2278] Marcos K. Aguilera. Tutorial on geo-replication in data center applications. ACM SIGMETRICS Performance Evaluation Review, 41(1):385– 386, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nair:2013:FHT

[2279] Jayakrishnan Nair, Adam Wierman, and Bert Zwart. The fundamentals of heavy-tails: properties, emergence, and identification. ACM SIGMET-RICS Performance Evaluation Review, 41(1):387–388, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schindler:2013:PAP

[2280] Jiri Schindler. Profiling and analyzing the I/O performance of NoSQL DBs. ACM SIGMETRICS Performance Evaluation Review, 41(1):389– 390, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gao:2013:SOC

[2281] X. Gao, y. Lu, M. Sharma, M. S. Squillante, and J. W. Bosman. Stochastic optimal control for a general class of dynamic resource allocation problems. ACM SIGMETRICS Performance Evaluation Review, 41(2):3–14, September 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jelenkovic:2013:RCC

[2282] Predrag R. Jelenković and Evangelia D. Skiani. Retransmissions over correlated channels. ACM SIGMETRICS Performance Evaluation Review, 41 (2):15–25, September 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mahmud:2013:OCP

[2283] A. Hasan Mahmud and Shaolei Ren. Online capacity provisioning for carbon-neutral data center with demand-responsive electricity prices. ACM SIGMETRICS Performance Evaluation Review, 41(2):26–37, September 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Asghari:2013:OEM

[2284] Naser M. Asghari, Michel Mandjes, and Anwar Walid. Optimizing energy management in multi-core servers. ACM SIGMETRICS Performance Evaluation Review, 41(2):38–40, September 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bekker:2013:SPS

[2285] R. Bekker, J. L. Dorsman, R. D.

van der Mei, P. Vis, and E. M. M. Winands. Scheduling in polling systems in heavy traffic. ACM SIGMET-RICS Performance Evaluation Review, 41(2):41–43, September 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casas:2013:YSL

[2286] Pedro Casas, Michael Seufert, and Raimund Schatz. YOUQMON: a system for on-line monitoring of YouTube QoE in operational 3G networks. ACM SIGMETRICS Performance Evaluation Review, 41(2):44–46, September 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dorsman:2013:PQN

[2287] Jan-Pieter Dorsman, Maria Vlasiou, and Bert Zwart. Parallel queueing networks with Markov-modulated service speeds in heavy traffic. ACM SIGMET-RICS Performance Evaluation Review, 41(2):47–49, September 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fiems:2013:SRE

[2288] Dieter Fiems and Koen De Turck. Spreading rumours in Euclidean space. ACM SIGMETRICS Performance Evaluation Review, 41(2):50–52, September 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vatamidou:2013:CPT

[2289] Eleni Vatamidou, Ivo Adan, Maria Vlasiou, and Bert Zwart. Corrected phasetype approximations for the workload of the MAP/G/1 queue with heavytailed service times. ACM SIGMET-RICS Performance Evaluation Review, 41(2):53–55, September 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Koziolek:2013:TSP

[2290] Heiko Koziolek, Steffen Becker, Jens Happe, Petr Tuma, and Thijmen de Gooijer. Towards software performance engineering for multicore and manycore systems. ACM SIGMET-RICS Performance Evaluation Review, 41(3):2–11, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bachmat:2013:AGD

[2291] Eitan Bachmat and Ilan Elhanan.

Analysis of the GSTF disk scheduling algorithm. ACM SIGMETRICS Performance Evaluation Review, 41 (3):13–15, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:2013:JOO

[2292] Minghong Lin, Li Zhang, Adam Wierman, and Jian Tan. Joint optimization of overlapping phases in MapReduce. ACM SIGMETRICS Performance Evaluation Review, 41(3):16–18, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghaderi:2013:RAW

[2293] Javad Ghaderi, Sem Borst, and Phil Whiting. Random access in wireless networks: how much aggressiveness can cause instability? ACM SIGMET-RICS Performance Evaluation Review, 41(3):19–21, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Adan:2013:QSB

[2294] Ivo Adan, Marko Boon, Ana Busić, Jean Mairesse, and Gideon Weiss. Queues with skill based parallel servers and a FCFS infinite matching model. ACM SIGMETRICS Performance Evaluation Review, 41(3):22–24, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Feinber:2013:DPO

[2295] Eugene A. Feinber and Fenghsu Yang.

Dynamic price optimization for an M/M/k/N queue with several customer types. ACM SIGMETRICS Performance Evaluation Review, 41 (3):25–27, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Urgaonkar:2013:PSC

[2296] Bhuvan Urgaonkar, George Kesidis, Uday V. Shanbhag, and Cheng Wang. Pricing of service in clouds: optimal response and strategic interactions. ACM SIGMETRICS Performance Evaluation Review, 41(3):28–30, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lim:2013:PTM

[2297] Sungsu Lim, Kyomin Jung, and John C. S. Lui. Phase transition of multi-state diffusion process in networks. ACM SIGMETRICS Performance Evaluation Review, 41(3):31–33, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Antunes:2013:PMG

[2298] Nelson Antunes, Gonçalo Jacinto, and António Pacheco. Probing a M/G/1 queue with general input and service times. ACM SIGMETRICS Performance Evaluation Review, 41(3):34–36, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harrison:2013:STD

[2299] Peter Harrison. Sojourn time distributions in tandem batch-networks. ACM SIGMETRICS Performance Evaluation Review, 41(3):37–39, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Spencer:2013:QFI

[2300] Joel Spencer, Madhu Sudan, and Kuang Xu. Queueing with future information. ACM SIGMETRICS Performance Evaluation Review, 41(3):40–42, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Poloczek:2013:MEA

[2301] Felix Poloczek and Florin Ciucu.

A martingale-envelope and applications. ACM SIGMETRICS Performance Evaluation Review, 41(3):43–45, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kang:2013:FAM

[2302] Weining Kang, Hongyuan Lu, and Guodong Pang. Fluid approximations of many-server queues with delayed feedback and heterogeneous service and patience times. ACM SIGMETRICS Performance Evaluation Review, 41 (3):46–48, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dieker:2013:DAL

[2303] A. B. Dieker and T. Suk. Diffusion approximations for large-scale buffered systems. ACM SIGMET-RICS Performance Evaluation Review, 41(3):49–51, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2013:ACR

[2304] Yingdong Lu, Mark S. Squillante, and David D. Yao. Asymptotics of a class of resource planning problems. ACM SIGMETRICS Performance Evaluation Review, 41(3):52–54, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ye:2013:ILH

[2305] Heng-Qing Ye and David D. Yao. Interchange of limits in heavy traffic analysis under a moment condition. ACM SIGMETRICS Performance Evaluation Review, 41(3):55–57, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Coffman:2013:PPU

[2306] Edward G. Coffman and Petar Momcilović. A particle process underlying SSD storage structures. ACM SIGMETRICS Performance Evaluation Review, 41(3):58–60, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gao:2013:RCF

[2307] X. Gao, Y. Lu, M. Sharma, M. S. Squillante, and J. W. Bosman. Rewards, costs and flexibility in dynamic resource allocation: a stochastic optimal control approach. ACM SIGMET-RICS Performance Evaluation Review, 41(3):61–63, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Temple:2013:MMP

[2308] William G. Temple and Richard T. B. Ma. Monotonic marginal pricing: demand response with price certainty. ACM SIGMETRICS Performance Evaluation Review, 41(3):65–70, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Singla:2013:BPS

[2309] Sahil Singla, Yashar Ghiassi-Farrokhfal, and Srinivasan Keshav. Battery provisioning and scheduling for a hybrid battery-diesel generator system. ACM SIGMETRICS Performance Evaluation Review, 41(3):71–76, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gan:2013:RTD

[2310] Lingwen Gan, Adam Wierman, Ufuk Topcu, Niangjun Chen, and Steven H. Low. Real-time deferrable load control: handling the uncertainties of renewable generation. ACM SIGMET-RICS Performance Evaluation Review, 41(3):77–79, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2013:OCT

[2311] Kai Yang and Anwar Walid. Outagecapacity tradeoff for smart grid with renewables. ACM SIGMETRICS Performance Evaluation Review, 41(3):80–82, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chan:2013:CVI

[2312] Christine S. Chan, Boxiang Pan, Kenny Gross, Kalyan Vaidyanathan, and Tajana Simunić Rosing. Correcting vibration-induced performance degradation in enterprise servers. ACM SIGMETRICS Performance Evaluation Review, 41(3):83–88, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2013:ESG

[2313] Hao Wang, Jianwei Huang, Xiaojun Lin, and Hamed Mohsenian-Rad. Exploring smart grid and data center interactions for electric power load balancing. ACM SIGMETRICS Performance Evaluation Review, 41(3):89–94, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pervila:2013:HHU

[2314] Mikko Pervilä, Lassi Remes, and Jussi Kangasharju. Harvesting heat in an urban greenhouse. ACM SIGMET-RICS Performance Evaluation Review, 41(3):95–97, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Widjaja:2013:SSE

[2315] Indra Widjaja, Anwar Walid, Yanbin Luo, Yang Xu, and H. Jonathan Chao. Switch sizing for energy-efficient datacenter networks. ACM SIGMET-RICS Performance Evaluation Review, 41(3):98–100, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hou:2013:HHE

[2316] Chenying Hou, Fa Zhang, Antonio Fernández Anta, Lin Wang, and Zhiyong Liu. A hop-by-hop energy efficient distributed routing scheme. ACM SIGMETRICS Performance Evaluation Review, 41(3):101–106, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2013:JVM

[2317] Lin Wang, Fa Zhang, Athanasios V.

Vasilakos, Chenying Hou, and Zhiyong Liu. Joint virtual machine assignment and traffic engineering for green data center networks. ACM SIGMET-RICS Performance Evaluation Review, 41(3):107–112, December 2013. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Loiseau:2014:MSG

[2318] Patrick Loiseau, David C. Parkes, and Jean Walrand. MultiDefender security games on networks. ACM SIGMET-RICS Performance Evaluation Review, 41(4):4–7, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Laszka:2014:QAO

[2319] Aron Laszka and Assane Gueye.

Quantifying all-to-one network topology robustness under budget constraints. ACM SIGMETRICS Performance Evaluation Review, 41(4):8–11, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dritsoula:2014:GCE

[2320] Lemonia Dritsoula and John Musacchio. A game of clicks: economic incentives to fight click fraud in ad networks. ACM SIGMETRICS Performance Evaluation Review, 41(4):12–15, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kavurmacioglu:2014:DIP

[2321] Emir Kavurmacioğlu, Murat Alanyali, and David Starobinski. Demandinsensitive price relationships in competitive private commons. ACM SIGMETRICS Performance Evaluation Review, 41(4):16–19, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Courcoubetis:2014:RMP

[2322] Costas Courcoubetis, Kostas Sdrolias, and Richard Weber. Revenue models, price differentiation and network neutrality implications in the Internet. ACM SIGMETRICS Performance Evaluation Review, 41(4):20–23, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Park:2014:ICR

[2323] Jiwon Park and Jeonghoon Mo. ISP and CP revenue sharing and content piracy. ACM SIGMETRICS Performance Evaluation Review, 41(4):24–27, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ifrach:2014:BSL

[2324] Bar Ifrach, Costis Maglaras, and Marco Scarsini. Bayesian social learning with consumer reviews. ACM SIGMET-RICS Performance Evaluation Review, 41(4):28, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dahleh:2014:CLI

[2325] Munther Dahleh, Alireza TahbazSalehi, John N. Tsitsiklis, and Spyros I. Zoumpoulis. Coordination with local information. ACM SIGMETRICS Performance Evaluation Review, 41(4):29, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Abbassi:2014:DCC

[2326] Zeinab Abbassi, Nidhi Hegde, and Laurent Massoulié. Distributed content curation on the Web. ACM SIGMET-RICS Performance Evaluation Review, 41(4):30–33, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xu:2014:IDH

[2327] Jie Xu and Mihaela van der Schaar. Incentive design for heterogeneous usergenerated content networks. ACM SIGMETRICS Performance Evaluation Review, 41(4):34–37, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jiang:2014:BLS

[2328] Chong Jiang, Carolyn L. Beck, and R. Srikant. Bidding with limited statistical knowledge in online auctions. ACM SIGMETRICS Performance Evaluation Review, 41(4):38–41, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rallapalli:2014:MVI

[2329] Swati Rallapalli, Qiang Ma, Han Hee Song, Mario Baldi, S. Muthukrishnan, and Lili Qiu. Modeling the value of information granularity in targeted advertising. ACM SIGMETRICS Performance Evaluation Review, 41(4):42–45, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chiang:2014:SSD

[2330] Mung Chiang. SDP: Smart Data Pricing, from theorems to trials. ACM SIGMETRICS Performance Evaluation Review, 41(4):46, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anselmi:2014:ECP

[2331] Jonatha Anselmi, Danilo Ardagna, John C. S. Lui, Adam Wierman, Yunjian Xu, and Zichao Yang. The economics of the cloud: price competition and congestion. ACM SIGMET-RICS Performance Evaluation Review, 41(4):47–49, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berbeglia:2014:PMD

[2332] Gerardo Berbeglia, Peter Sloan, and Adrian Vetta. Pricing mechanisms for a durable good monopolist. ACM SIGMETRICS Performance Evaluation Review, 41(4):50, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ifrach:2014:PBS

[2333] Bar Ifrach and Rameh Johari. Pricing a bestseller: sales and visibility in the marketplace for mobile apps. ACM SIGMETRICS Performance Evaluation Review, 41(4):51, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wagner:2014:DAL

[2334] Daniel T. Wagner, Andrew Rice, and Alastair R. Beresford. Device analyzer: largescale mobile data collection. ACM SIGMETRICS Performance Evaluation Review, 41(4):53–56, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Feinberg:2014:OCU

[2335] Eugene Feinberg and Xiaoxuan Zhang.

Optimizing cloud utilization via switching decisions. ACM SIGMET-RICS Performance Evaluation Review, 41(4):57–60, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yilmaz:2014:FDK

[2336] Onur Yilmaz, Mustafa U. Torun, and Ali N. Akansu. A fast derivation of Karhunen–Loeve transform kernel for first-order autoregressive discrete process. ACM SIGMETRICS Performance Evaluation Review, 41(4):61–64, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Madan:2014:ATA

[2337] Bharat B. Madan and Manoj Banik.

Attack tolerant architecture for big data file systems. ACM SIGMET-RICS Performance Evaluation Review, 41(4):65–69, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Suthaharan:2014:BDC

[2338] Shan Suthaharan. Big data classification: problems and challenges in network intrusion prediction with machine learning. ACM SIGMETRICS Performance Evaluation Review, 41(4):70–73, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sharma:2014:MAC

[2339] Abhishek B. Sharma, Franjo Ivancić, Alexandru Niculescu-Mizil, Haifeng Chen, and Guofei Jiang. Modeling and analytics for cyber-physical systems in the age of big data. ACM SIGMET-RICS Performance Evaluation Review, 41(4):74–77, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hu:2014:AIM

[2340] Jie Hu, Kun Meng, Xiaomin Chen, Chuang Lin, and Jiwei Huang. Analysis of influence maximization in largescale social networks. ACM SIGMET-RICS Performance Evaluation Review, 41(4):78–81, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Whitworth:2014:SPC

[2341] Jeff Whitworth and Shan Suthaharan. Security problems and challenges in a machine learning-based hybrid big data processing network systems. ACM SIGMETRICS Performance Evaluation Review, 41(4):82–85, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Savas:2014:TBD

[2342] Onur Savas, Yalin Sagduyu, Julia Deng, and Jason Li. Tactical big data analytics: challenges, use cases, and solutions. ACM SIGMETRICS Performance Evaluation Review, 41(4):86–89, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2014:FOL

[2343] Yu Zhang, Daby Sow, Deepak Turaga, and Mihaela van der Schaar. A fast online learning algorithm for distributed mining of BigData. ACM SIGMET-RICS Performance Evaluation Review, 41(4):90–93, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Heintz:2014:BGT

[2344] Benjamin Heintz and Abhishek Chandra. Beyond graphs: toward scalable hypergraph analysis systems. ACM SIGMETRICS Performance Evaluation Review, 41(4):94–97, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Al-Jaroodi:2014:DDB

[2345] Jameela Al-Jaroodi, Nader Mohamed, and Abdulla Eid. Dual direction big data download and analysis. ACM SIGMETRICS Performance Evaluation Review, 41(4):98–101, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brock:2014:LAN

[2346] Jacob Brock, Hao Luo, and Chen Ding. Locality analysis: a nonillion time window problem. ACM SIGMET-RICS Performance Evaluation Review, 41(4):102–105, March 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2014:RSD

[2347] Zhaoxu Wang, Wenxiang Dong, Wenyi Zhang, and Chee Wei Tan. Rumor source detection with multiple observations: fundamental limits and algorithms. ACM SIGMETRICS Performance Evaluation Review, 42(1):1–13, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Han:2014:COS

[2348] Jinyoung Han, Daejin Choi, Byung-Gon Chun, Ted Kwon, Hyun chul Kim, and Yanghee Choi. Collecting, organizing, and sharing pins in pinterest: interest-driven or socialdriven? ACM SIGMETRICS Performance Evaluation Review, 42(1):15–27, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xu:2014:JCR

[2349] Jiaming Xu, Rui Wu, Kai Zhu, Bruce Hajek, R. Srikant, and Lei Ying. Jointly clustering rows and columns of binary matrices: algorithms and tradeoffs. ACM SIGMETRICS Performance Evaluation Review, 42(1):29–41, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

May:2014:FFH

[2350] Avner May, Augustin Chaintreau, Nitish Korula, and Silvio Lattanzi. Filter & follow: how social media foster content curation. ACM SIGMET-RICS Performance Evaluation Review, 42(1):43–55, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ioannidis:2014:PTP

[2351] Stratis Ioannidis, Andrea Montanari, Udi Weinsberg, Smriti Bhagat, Nadia Fawaz, and Nina Taft. Privacy tradeoffs in predictive analytics. ACM SIGMETRICS Performance Evaluation Review, 42(1):57–69, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shi:2014:OAF

[2352] Weijie Shi, Linquan Zhang, Chuan Wu, Zongpeng Li, and Francis C. M. Lau. An online auction framework for dynamic resource provisioning in cloud computing. ACM SIGMET-RICS Performance Evaluation Review, 42(1):71–83, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nair:2014:EPS

[2353] Jayakrishnan Nair, Sachin Adlakha, and Adam Wierman. Energy procurement strategies in the presence of intermittent sources. ACM SIGMET-RICS Performance Evaluation Review, 42(1):85–97, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2014:RAD

[2354] Linquan Zhang, Zongpeng Li, and Chuan Wu. Randomized auction design for electricity markets between grids and microgrids. ACM SIGMET-RICS Performance Evaluation Review, 42(1):99–110, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2014:PDC

[2355] Zhenhua Liu, Iris Liu, Steven Low, and Adam Wierman. Pricing data center demand response. ACM SIGMET-RICS Performance Evaluation Review, 42(1):111–123, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Larranaga:2014:IPM

[2356] Maialen Larrañaga, Urtzi Ayesta, and Ina Maria Verloop. Index policies for a multi-class queue with convex holding cost and abandonments. ACM SIGMETRICS Performance Evaluation Review, 42(1):125–137, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Walton:2014:CSS

[2357] Neil Stuart Walton. Concave switching in single and multihop networks. ACM SIGMETRICS Performance Evaluation Review, 42(1):139–151, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Huang:2014:POL

[2358] Longbo Huang, Xin Liu, and Xiaohong Hao. The power of online learning in stochastic network optimization. ACM SIGMETRICS Performance Evaluation Review, 42(1):153–165, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jelenkovic:2014:SRC

[2359] Predrag R. Jelenkovic and Evangelia D. Skiani. Is sharing with retransmissions causing instabilities? ACM SIGMET-RICS Performance Evaluation Review, 42(1):167–179, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2014:NWC

[2360] Jian Tan, Yandong Wang, Weikuan Yu, and Li Zhang. Non-work-conserving effects in MapReduce: diffusion limit and criticality. ACM SIGMETRICS Performance Evaluation Review, 42 (1):181–192, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stoica:2014:CBD

[2361] Ion Stoica. Conquering big data with spark and BDAS. ACM SIGMETRICS Performance Evaluation Review, 42(1): 193, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shamsi:2014:HSP

[2362] Zain Shamsi, Ankur Nandwani, Derek Leonard, and Dmitri Loguinov. Hershel: single-packet OS fingerprinting. ACM SIGMETRICS Performance Evaluation Review, 42(1):195– 206, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shahzad:2014:NCH

[2363] Muhammad Shahzad and Alex X. Liu.

Noise can help: accurate and efficient per-flow latency measurement without packet probing and time stamping. ACM SIGMETRICS Performance Evaluation Review, 42(1):207– 219, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Viennot:2014:MSG

[2364] Nicolas Viennot, Edward Garcia, and Jason Nieh. A measurement study of Google Play. ACM SIGMET-RICS Performance Evaluation Review, 42(1):221–233, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kim:2014:ITC

[2365] Chung Hwan Kim, Junghwan Rhee, Hui Zhang, Nipun Arora, Guofei Jiang, Xiangyu Zhang, and Dongyan Xu. IntroPerf: transparent context-sensitive multi-layer performance inference using system stack traces. ACM SIGMET-RICS Performance Evaluation Review, 42(1):235–247, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Suneja:2014:NIB

[2366] Sahil Suneja, Canturk Isci, Vasanth Bala, Eyal de Lara, and Todd Mummert. Non-intrusive, out-of-band and out-of-the-box systems monitoring in the cloud. ACM SIGMETRICS Performance Evaluation Review, 42(1):249– 261, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krishnasamy:2014:BEU

[2367] Subhashini Krishnasamy, Siddhartha Banerjee, and Sanjay Shakkottai. The behavior of epidemics under bounded susceptibility. ACM SIGMETRICS Performance Evaluation Review, 42 (1):263–275, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gabielkov:2014:SSN

[2368] Maksym Gabielkov, Ashwin Rao, and Arnaud Legout. Studying social networks at scale: macroscopic anatomy of the Twitter social graph. ACM SIGMETRICS Performance Evaluation Review, 42(1):277–288, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Buccapatnam:2014:SBS

[2369] Swapna Buccapatnam, Atilla Eryilmaz, and Ness B. Shroff. Stochastic bandits with side observations on networks. ACM SIGMETRICS Performance Evaluation Review, 42(1):289– 300, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ok:2014:MDS

[2370] Jungseul Ok, Youngmi Jin, Jinwoo Shin, and Yung Yi. On maximizing diffusion speed in social networks: impact of random seeding and clustering. ACM SIGMETRICS Performance Evaluation Review, 42(1):301–313, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yallouz:2014:TSS

[2371] Jose Yallouz, Ori Rottenstreich, and Ariel Orda. Tunable survivable spanning trees. ACM SIGMETRICS Performance Evaluation Review, 42(1): 315–327, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghit:2014:BRA

[2372] Bogdan Ghit, Nezih Yigitbasi, Alexandru Iosup, and Dick Epema. Balanced resource allocations across multiple dynamic MapReduce clusters. ACM SIGMETRICS Performance Evaluation Review, 42(1):329–341, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berger:2014:RAQ

[2373] Daniel S. Berger, Martin Karsten, and Jens Schmitt. On the relevance of adversarial queueing theory in practice. ACM SIGMETRICS Performance Evaluation Review, 42(1):343– 354, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nachiappan:2014:GFE

[2374] Nachiappan Chidambaram Nachiappan, Praveen Yedlapalli, Niranjan Soundararajan, Mahmut Taylan Kandemir, Anand Sivasubramaniam, and Chita R. Das. GemDroid: a framework to evaluate mobile platforms. ACM SIGMETRICS Performance Evaluation Review, 42(1):355–366, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shafiq:2014:UIN

[2375] Muhammad Zubair Shafiq, Jeffrey Erman, Lusheng Ji, Alex X. Liu, Jeffrey Pang, and Jia Wang. Understanding the impact of network dynamics on mobile video user engagement. ACM SIGMETRICS Performance Evaluation Review, 42(1):367–379, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Huang:2014:EEC

[2376] Jiwei Huang, Sen Yang, Ashwin Lall, Justin Romberg, Jun Xu, and Chuang Lin. Error estimating codes for insertion and deletion channels. ACM SIGMETRICS Performance Evaluation Review, 42(1):381–393, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Meyfroyt:2014:DDP

[2377] Thomas M. M. Meyfroyt, Sem C. Borst, Onno J. Boxma, and Dee Denteneer. Data dissemination performance in large-scale sensor networks. ACM SIGMETRICS Performance Evaluation Review, 42(1):395–406, June 2014.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gorlatova:2014:MSK

[2378] Maria Gorlatova, John Sarik, Guy Grebla, Mina Cong, Ioannis Kymissis, and Gil Zussman. Movers and shakers: kinetic energy harvesting for the Internet of Things. ACM SIGMET-RICS Performance Evaluation Review, 42(1):407–419, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lai:2014:PLT

[2379] Chengdi Lai, Steven H. Low, Ka-Cheong Leung, and Victor O. K. Li. Pricing link by time. ACM SIGMET-RICS Performance Evaluation Review, 42(1):421–433, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Moharir:2014:SCU

[2380] Sharayu Moharir, Javad Ghaderi, Sujay Sanghavi, and Sanjay Shakkottai. Serving content with unknown demand: the high-dimensional regime. ACM SIGMETRICS Performance Evaluation Review, 42(1):435– 447, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tune:2014:NDS

[2381] Paul Tune and Matthew Roughan.

Network-design sensitivity analysis. ACM SIGMETRICS Performance Evaluation Review, 42(1):449–461, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ai:2014:MSS

[2382] Lingqing Ai, Xian Wu, Lingxiao Huang, Longbo Huang, Pingzhong Tang, and Jian Li. The multi-shop ski rental problem. ACM SIGMET-RICS Performance Evaluation Review, 42(1):463–475, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ding:2014:CCC

[2383] Wei Ding and Mahmut Kandemir. CApRI: CAche-conscious data reordering for irregular codes. ACM SIGMET-RICS Performance Evaluation Review, 42(1):477–489, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cai:2014:NCA

[2384] Yu Cai, Gulay Yalcin, Onur Mutlu, Erich F. Haratsch, Osman Unsal, Adrian Cristal, and Ken Mai. Neighbor-cell assisted error correction for MLC NAND flash memories. ACM SIGMETRICS Performance Evaluation Review, 42(1):491–504, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gulur:2014:AAM

[2385] Nagendra Gulur, Mahesh Mehendale, Raman Manikantan, and Ramaswamy Govindarajan. ANATOMY: an analytical model of memory system performance. ACM SIGMETRICS Performance Evaluation Review, 42(1):505– 517, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Khan:2014:EEM

[2386] Samira Khan, Donghyuk Lee, Yoongu Kim, Alaa R. Alameldeen, Chris Wilkerson, and Onur Mutlu. The efficacy of error mitigation techniques for DRAM retention failures: a comparative experimental study. ACM SIGMET-RICS Performance Evaluation Review, 42(1):519–532, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2014:GDM

[2387] Kaibo Wang, Xiaoning Ding, Rubao Lee, Shinpei Kato, and Xiaodong Zhang. GDM: device memory management for GPGPU computing. ACM SIGMETRICS Performance Evaluation Review, 42(1):533–545, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Diegues:2014:EPC

[2388] Nuno Diegues, Paolo Romano, and Luís Rodrigues. On the energy and performance of commodity hardware transactional memory. ACM SIGMET-RICS Performance Evaluation Review, 42(1):547–548, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2014:ICM

[2389] Qi Wang, Liang Liu, Jinbei Zhang, Xinyu Wang, Xinbing Wang, and Songwu Lu. Impact of correlated mobility and cluster scalability on connectivity of wireless networks. ACM SIGMETRICS Performance Evaluation Review, 42(1):549–550, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tavakkol:2014:UPD

[2390] Arash Tavakkol, Mohammad Arjomand, and Hamid Sarbazi-Azad. Unleashing the potentials of dynamism for page allocation strategies in SSDs. ACM SIGMETRICS Performance Evaluation Review, 42(1):551– 552, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mandayam:2014:TCM

[2391] Chinmoy V. Mandayam and Balaji Prabhakar. Traffic congestion: models, costs and optimal transport. ACM SIGMETRICS Performance Evaluation Review, 42(1):553–554, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mukhopadhyay:2014:RRS

[2392] Arpan Mukhopadhyay and Ravi R. Mazumdar. Randomized routing schemes for large processor sharing systems with multiple service rates. ACM SIGMETRICS Performance Evaluation Review, 42(1):555–556, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tarvo:2014:AAM

[2393] Alexander Tarvo and Steven P. Reiss. Automated analysis of multithreaded programs for performance modeling. ACM SIGMETRICS Performance Evaluation Review, 42(1):557– 558, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arora:2014:CCP

[2394] Manish Arora, Srilatha Manne, Yasuko Eckert, Indrani Paul, Nuwan Jayasena, and Dean Tullsen. A comparison of core power gating strategies implemented in modern hardware. ACM SIGMETRICS Performance Evaluation Review, 42(1):559–560, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ray:2014:TMN

[2395] Avik Ray, Sujay Sanghavi, and Sanjay Shakkottai. Topic modeling from network spread. ACM SIGMETRICS Performance Evaluation Review, 42 (1):561–562, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mahmud:2014:BBC

[2396] A. Hasan Mahmud, Yuxiong He, and Shaolei Ren. BATS: budgetconstrained autoscaling for cloud performance optimization. ACM SIG-METRICS Performance Evaluation Review, 42(1):563–564, June 2014. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ammar:2014:WYC

[2397] Ammar Ammar, Sewoong Oh, Devavrat Shah, and Luis Filipe Voloch. What’s your choice?: learning the mixed multi-nomial. ACM SIGMET-RICS Performance Evaluation Review, 42(1):565–566, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shafiq:2014:RCC

[2398] Muhammad Zubair Shafiq, Alex X.

Liu, and Amir R. Khakpour. Revisiting caching in content delivery networks. ACM SIGMETRICS Performance Evaluation Review, 42(1):567– 568, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xu:2014:FSL

[2399] Qiang Xu, Thomas Andrews, Yong Liao, Stanislav Miskovic, Z. Morley Mao, Mario Baldi, and Antonio Nucci. FLOWR: a self-learning system for classifying mobile-application traffic. ACM SIGMETRICS Performance Evaluation Review, 42(1):569– 570, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dong:2014:ART

[2400] Wei Dong, Xuefeng Zhang, Jiliang Wang, Yi Gao, Chun Chen, and Jiajun Bu. Accurate and robust time reconstruction for deployed sensor networks. ACM SIGMETRICS Performance Evaluation Review, 42(1):571– 572, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2014:EPS

[2401] Shaoquan Zhang, Longbo Huang, Minghua Chen, and Xin Liu. Effect of proactive serving on user delay reduction in service systems. ACM SIGMETRICS Performance Evaluation Review, 42(1):573–574, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kong:2014:OES

[2402] Fanxin Kong, Xue Liu, and Lei Rao. Optimal energy source selection and capacity planning for green datacenters. ACM SIGMETRICS Performance Evaluation Review, 42(1):575– 576, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shin:2014:SUI

[2403] Jinwoo Shin and Tonghoon Suk.

Scheduling using interactive oracles: connection between iterative optimization and low-complexity scheduling. ACM SIGMETRICS Performance Evaluation Review, 42(1):577– 578, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rallapalli:2014:ULF

[2404] Swati Rallapalli, Wei Dong, Lili Qiu, and Yin Zhang. Unified localization framework using trajectory signatures. ACM SIGMETRICS Performance Evaluation Review, 42(1):579– 580, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kang:2014:TCT

[2405] Dong Hyun Kang, Changwoo Min, and Young Ik Eom. TS-CLOCK: temporal and spatial locality aware buffer replacement algorithm for NAND flash storages. ACM SIGMETRICS Performance Evaluation Review, 42(1):581– 582, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kim:2014:MSM

[2406] Juhoon Kim, Yung-Chih Chen, Ramin Khalili, Don Towsley, and Anja Feldmann. Multi-source multipath HTTP (mHTTP): a proposal. ACM SIGMET-RICS Performance Evaluation Review, 42(1):583–584, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vlachou:2014:PAM

[2407] Christina Vlachou, Albert Banchs, Julien Herzen, and Patrick Thiran. Performance analysis of MAC for power-line communications. ACM SIGMETRICS Performance Evaluation Review, 42(1):585–586, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vu:2014:IDC

[2408] Long Vu, Deepak Turaga, and Srinivasan Parthasarathy. Impact of DHCP churn on network characterization. ACM SIGMETRICS Performance Evaluation Review, 42(1):587– 588, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Guo:2014:OAJ

[2409] Yang Guo, Alexander L. Stolyar, and Anwar Walid. Online algorithms for joint application-VM-physical-machine auto-scaling in a cloud. ACM SIG-METRICS Performance Evaluation Review, 42(1):589–590, June 2014. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2014:DOL

[2410] Jia Liu, Cathy H. Xia, Ness B. Shroff, and Hanif D. Sherali. Distributed optimal load shedding for disaster recovery in smart electric power grids: a secondorder approach. ACM SIGMETRICS Performance Evaluation Review, 42 (1):591–592, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Clegg:2014:TSS

[2411] Richard G. Clegg, Raul Landa, João Taveira Araújo, Eleni Mykoniati, David Griffin, and Miguel Rio. TARDIS: stably shifting traffic in space and time. ACM SIGMETRICS Performance Evaluation Review, 42(1):593– 594, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berger:2014:EAT

[2412] Daniel S. Berger, Philipp Gland, Sahil Singla, and Florin Ciucu. Exact analysis of TTL cache networks: the case of caching policies driven by stopping times. ACM SIGMETRICS Performance Evaluation Review, 42(1):595– 596, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jyothi:2014:MTD

[2413] Sangeetha Abdu Jyothi, Ankit Singla, P. Brighten Godfrey, and Alexandra Kolla. Measuring throughput of data center network topologies. ACM SIGMETRICS Performance Evaluation Review, 42(1):597–598, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2014:ETR

[2414] Da Wang, Gauri Joshi, and Gregory Wornell. Efficient task replication for fast response times in parallel computation. ACM SIGMET-RICS Performance Evaluation Review, 42(1):599–600, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Buchholz:2014:JLC

[2415] Peter Buchholz and Benny Van Houdt. Joint latency and cost optimization for erasure-coded data center storage. ACM SIGMETRICS Performance Evaluation Review, 42(2):3–14, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2014:RPS

[2416] Bo Zhang, Guodong Pang, and Bert Zwart. Refining piecewise stationary approximation for a Markov-regulated fluid queue. ACM SIGMETRICS Performance Evaluation Review, 42(2): 15–17, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Izagirre:2014:LTP

[2417] A. Izagirre and A. M. Makowski. Light traffic performance under the power of two load balancing strategy: the case of server heterogeneity. ACM SIGMET-RICS Performance Evaluation Review, 42(2):18–20, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shioda:2014:RWB

[2418] Shigeo Shioda. Random walk based biased sampling for data collection on communication networks. ACM SIGMETRICS Performance Evaluation Review, 42(2):21–23, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Haddad:2014:SEE

[2419] Majed Haddad, Oussama Habachi, Piotr Wiecek, and Yezekael Hayel. Spectral efficiency of energy efficient multicarrier systems. ACM SIGMET-RICS Performance Evaluation Review, 42(2):24–26, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2014:MCI

[2420] Tao Zhang, Guangshuo Chen, Wei Shu, and Min-You Wu. Microarchitectural characterization of irregular applications on GPGPUs. ACM SIGMET-RICS Performance Evaluation Review, 42(2):27–29, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nair:2014:CPC

[2421] Jayakrishnan Nair, Vijay G. Subramanian, and Adam Wierman. On competitive provisioning of cloud services. ACM SIGMETRICS Performance Evaluation Review, 42(2):30–32, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bosman:2014:PCT

[2422] Joost Bosman, Jayakrishnan Nair, and Bert Zwart. On the probability of current and temperature overloading in power grids: a large deviations approach. ACM SIGMETRICS Performance Evaluation Review, 42(2): 33–35, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gelenbe:2014:SNE

[2423] Erol Gelenbe. A sensor node with energy harvesting. ACM SIGMET-RICS Performance Evaluation Review, 42(2):37–39, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Meyfroyt:2014:CSA

[2424] Thomas M. M. Meyfroyt. A cooperative sequential adsorption model for wireless gossiping. ACM SIGMET-RICS Performance Evaluation Review, 42(2):40–42, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tune:2014:MET

[2425] Paul Tune and Matthew Roughan. Maximum entropy traffic matrix synthesis. ACM SIGMETRICS Performance Evaluation Review, 42(2):43–45, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bradonjic:2014:SCR

[2426] Milan Bradonjić, Iraj Saniee, and Indra Widjaja. Scaling of capacity and reliability in data center networks. ACM SIGMETRICS Performance Evaluation Review, 42(2):46–48, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rochman:2014:ERP

[2427] Yuval Rochman, Hanoch Levy, and Eli Brosh. Efficient resource placement in cloud computing and network applications. ACM SIGMET-RICS Performance Evaluation Review, 42(2):49–51, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xie:2014:MCS

[2428] Hong Xie and John C. S. Lui. Modeling crowdsourcing systems: design and analysis of incentive mechanism and rating system. ACM SIGMET-RICS Performance Evaluation Review, 42(2):52–54, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Asadi:2014:MDC

[2429] Arash Asadi, Peter Jacko, and Vincenzo Mancuso. Modeling D2D communications with LTE and WiFi. ACM SIGMETRICS Performance Evaluation Review, 42(2):55–57, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zheng:2014:EFF

[2430] Liang Zheng and Chee Wei Tan. Egalitarian fairness framework for joint rate and power optimization in wireless networks. ACM SIGMETRICS Performance Evaluation Review, 42 (2):58–60, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goldberg:2014:AOC

[2431] D. A. Goldberg, D. A. Katz, Y. Lu, M. Sharma, and M. S. Squillante. Asymptotic optimality of constant capacity allocation policies for dynamic resource planning. ACM SIGMET-RICS Performance Evaluation Review, 42(2):61–63, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghaderi:2014:AOB

[2432] Javad Ghaderi, Yuan Zhong, and R. Srikant. Asymptotic optimality of BestFit for stochastic bin packing. ACM SIGMETRICS Performance Evaluation Review, 42(2):64–66, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tizghadam:2014:ISI

[2433] Ali Tizghadam and Alberto Leon-Garcia. On the impact of subgraph insertion or removal on Moore– Penrose Laplacian and resistance distance. ACM SIGMETRICS Performance Evaluation Review, 42(2):67–69, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Miyazawa:2014:TAS

[2434] Masakiyo Miyazawa. Tail asymptotics of the stationary distribution for a two-node generalized Jackson network. ACM SIGMETRICS Performance Evaluation Review, 42(2):70–72, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2014:ISS

[2435] Mark S. Squillante, Don Towsley, and Sean Barker. Improving the scalability of search in networks through multiple random walks. ACM SIGMET-RICS Performance Evaluation Review, 42(2):73–75, September 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chuang:2014:JWP

[2436] John Chuang and Patrick Loiseau. The Joint Workshop on Pricing and Incentives in Networks and Systems (WPIN+NetEcon 2014). ACM SIG-METRICS Performance Evaluation Review, 42(3):2–3, December 2014.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kamble:2014:SMP

[2437] Vijay Kamble and Jean Walrand.

Strategy-proof mechanisms for purchasing a shared resource. ACM SIGMETRICS Performance Evaluation Review, 42(3):4–7, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Manickam:2014:ITM

[2438] Saravana Manickam, Mahesh K. Marina, Sofia Pediaditaki, and Maziar Nekovee. An iterative and truthful multi-unit auction scheme for coordinated sharing of spectrum white spaces. ACM SIGMETRICS Performance Evaluation Review, 42(3):8–11, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sinha:2014:GMD

[2439] Abhinav Sinha and Achilleas Anastasopoulos. A general mechanism design methodology for social utility maximization with linear constraints. ACM SIGMETRICS Performance Evaluation Review, 42(3):12–15, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Weber:2014:FAS

[2440] Steven Weber and Roch Guérin. Facilitating adoption of services with positive externalities via subsidies. ACM SIGMETRICS Performance Evaluation Review, 42(3):16–19, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ajorlou:2014:SID

[2441] Amir Ajorlou, Ali Jadbabaie, and Ali Kakhbod. Strategic information diffusion: Spread vs. exploit. ACM SIGMETRICS Performance Evaluation Review, 42(3):20–23, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Acemoglu:2014:HIL

[2442] Daron Acemoglu, Giacomo Como, Fabio Fagnani, and Asuman Ozdaglar. Harmonic influence in largescale networks. ACM SIGMETRICS Performance Evaluation Review, 42(3):24, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raja:2014:FFF

[2443] Vamseedhar Reddyvari Raja, Srinivas Shakkottai, Amogh Dhamdhere, and kc claffy. Fair, flexible and feasible ISP billing. ACM SIGMETRICS Performance Evaluation Review, 42 (3):25–28, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gyarmati:2014:APB

[2444] Laszlo Gyarmati, Nikolaos Laoutaris, Kostas Sdrolias, Pablo Rodriguez, and Costas Courcoubetis. From advertising profits to bandwidth prices: a quantitative methodology for negotiating premium peering. ACM SIGMET-RICS Performance Evaluation Review, 42(3):29–32, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Simhon:2014:ARG

[2445] Eran Simhon and David Starobinski.

Advance reservation games and the price of conservatism. ACM SIGMET-RICS Performance Evaluation Review, 42(3):33, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bentov:2014:PAE

[2446] Iddo Bentov, Charles Lee, Alex Mizrahi, and Meni Rosenfeld. Proof of activity: Extending Bitcoin’s proof of work via proof of stake [extended abstract]y. ACM SIGMETRICS Performance Evaluation Review, 42(3):34–37, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Acemoglu:2014:NSC

[2447] Daron Acemoglu, Azarakhsh Malekian, and Asu Ozdaglar. Network security and contagion. ACM SIGMET-RICS Performance Evaluation Review, 42(3):38, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Roth:2014:DPT

[2448] Aaron Roth. Differential privacy as a tool for mechanism design in large systems. ACM SIGMETRICS Performance Evaluation Review, 42(3):39, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Georgiadis:2014:DEC

[2449] Leonidas Georgiadis, George Iosifidis, and Leandros Tassiulas. Dynamic exchange of communication services. ACM SIGMETRICS Performance Evaluation Review, 42(3):40, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kazumori:2014:GDA

[2450] Eiichiro Kazumori. Generalizing deferred acceptance auctions to allow multiple relinquishment options. ACM SIGMETRICS Performance Evaluation Review, 42(3):41, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

AlDaoud:2014:GUS

[2451] Ashraf Al Daoud, George Kesidis, and Jörg Liebeherr. A game of uncoordinated sharing of private spectrum commons. ACM SIGMETRICS Performance Evaluation Review, 42(3):42, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Poularakis:2014:QPQ

[2452] Konstantinos Poularakis, Ioannis Pefkianakis, Jaideep Chandrashekar, and Leandros Tassiulas. Quid pro quo: Reducing peak traffic costs with (subscriber) price incentives. ACM SIG-METRICS Performance Evaluation Review, 42(3):43, December 2014. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lotfi:2014:NNI

[2453] Mohammad Hassan Lotfi, George Kesidis, and Saswati Sarkar. Network NonNeutrality on the Internet: Content provision under a subscription revenue model. ACM SIGMETRICS Performance Evaluation Review, 42(3):44, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Joseph:2014:MFT

[2454] Siny Joseph, Vinod Namboodiri, and Vishnu C. Dev. A market-driven framework towards environmentally sustainable mobile computing. ACM SIGMETRICS Performance Evaluation Review, 42(3):46–48, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jalali:2014:ECC

[2455] Fatemeh Jalali, Rob Ayre, Arun Vishwanath, Kerry Hinton, Tansu Alpcan, and Rod Tucker. Energy consumption of content distribution from nano data centers versus centralized data centers. ACM SIGMETRICS Performance Evaluation Review, 42 (3):49–54, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Miwa:2014:ECH

[2456] Shinobu Miwa and Charles R. Lefurgy. Evaluation of core hopping on POWER7. ACM SIGMETRICS Performance Evaluation Review, 42 (3):55–60, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Debele:2014:ERS

[2457] Fikru Getachew Debele, Nanfang Li, Michela Meo, Marco Ricca, and Yi Zhang. Experimenting resource-on-demand strategies for green WLANs. ACM SIGMETRICS Performance Evaluation Review, 42(3):61–66, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yi:2014:MEC

[2458] Qing Yi and Suresh Singh. Minimizing energy consumption of FatTree data center networks. ACM SIGMET-RICS Performance Evaluation Review, 42(3):67–72, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ren:2014:FLC

[2459] Shaolei Ren and Mohammad A. Islam. A first look at colocation demand response. ACM SIGMETRICS Performance Evaluation Review, 42 (3):73–75, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cavdar:2014:QBS

[2460] Derya Çavdar, Andrea Rosà, Lydia Y. Chen, Walter Binder, and Fatih Alagöz. Quantifying the brown side of priority schedulers: Lessons from big clusters. ACM SIGMET-RICS Performance Evaluation Review, 42(3):76–81, December 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ardagna:2015:SIP

[2461] Danilo Ardagna and Mark S. Squillante. Special issue on performance and resource management in big data applications. ACM SIGMETRICS Performance Evaluation Review, 42(4):2, March 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2015:ALA

[2462] Yue Tan and Cathy H. Xia. An adaptive learning approach for efficient resource provisioning in cloud services. ACM SIGMETRICS Performance Evaluation Review, 42(4):3–11, March 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rosa:2015:DCE

[2463] Andrea Rosà, Lydia Y. Chen, Robert Birke, and Walter Binder. Demystifying casualties of evictions in big data priority scheduling. ACM SIGMET-RICS Performance Evaluation Review, 42(4):12–21, March 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ying:2015:EAE

[2464] Yijun Ying, Robert Birke, Cheng Wang, Lydia Y. Chen, and Gautam Natarajan. On energyaware allocation and execution for batch and interactive MapReduce. ACM SIGMET-RICS Performance Evaluation Review, 42(4):22–30, March 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2015:MRF

[2465] Jian Tan, Li Zhang, Min Li, and Yandong Wang. Multi-resource fair sharing for multiclass workflows. ACM SIGMETRICS Performance Evaluation Review, 42(4):31–37, March 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2015:ECH

[2466] Zhuoyao Zhang, Ludmila Cherkasova, and Boon Thau Loo. Exploiting cloud heterogeneity to optimize performance and cost of MapReduce processing. ACM SIGMETRICS Performance Evaluation Review, 42(4):38–50, March 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Malekimajd:2015:OMR

[2467] Marzieh Malekimajd, Danilo Ardagna, Michele Ciavotta, Alessandro Maria Rizzi, and Mauro Passacantando. Optimal map reduce job capacity allocation in cloud systems. ACM SIGMET-RICS Performance Evaluation Review, 42(4):51–61, March 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2015:MIM

[2468] Wei Zhang, Sundaresan Rajasekaran, Shaohua Duan, Timothy Wood, and Mingfa Zhuy. Minimizing interference and maximizing progress for Hadoop virtual machines. ACM SIGMET-RICS Performance Evaluation Review, 42(4):62–71, March 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hajek:2015:BID

[2469] Bruce Hajek. Bounds implied by drift with applications. ACM SIGMETRICS Performance Evaluation Review, 43(1): 1, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2015:OAI

[2470] Xiaoxi Zhang, Zhiyi Huang, Chuan Wu, Zongpeng Li, and Francis C. M. Lau. Online auctions in IaaS clouds: Welfare and profit maximization with server costs. ACM SIGMETRICS Performance Evaluation Review, 43(1):3– 15, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yun:2015:DPF

[2471] Se-Young Yun and Alexandre Proutiere.

Distributed proportional fair load balancing in heterogeneous systems. ACM SIGMETRICS Performance Evaluation Review, 43(1):17–30, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2015:MRF

[2472] Thomas Bonald and James Roberts.

Multi-resource fairness: Objectives, algorithms and performance. ACM SIGMETRICS Performance Evaluation Review, 43(1):31–42, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Georgiadis:2015:ESN

[2473] Leonidas Georgiadis, George Iosifidis, and Leandros Tassiulas. Exchange of services in networks: Competition, cooperation, and fairness. ACM SIGMETRICS Performance Evaluation Review, 43(1):43–56, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aalto:2015:WIA

[2474] Samuli Aalto, Pasi Lassila, and Prajwal Osti. Whittle index approach to sizeaware scheduling with time-varying channels. ACM SIGMETRICS Performance Evaluation Review, 43(1):57–69, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sur:2015:GIN

[2475] Sanjib Sur, Vignesh Venkateswaran, Xinyu Zhang, and Parmesh Ramanathan. 60 GHz indoor networking through flexible beams: a link-level profiling. ACM SIGMETRICS Performance Evaluation Review, 43(1):71–84, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2015:SDP

[2476] Liang Zhang, Weijie Wu, and Dan Wang. Sponsored data plan: a twoclass service model in wireless data networks. ACM SIGMETRICS Performance Evaluation Review, 43(1):85–96, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2015:QPR

[2477] Bin Li and Rayadurgam Srikant. Queue-proportional rate allocation with per-link information in multihop networks. ACM SIGMETRICS Performance Evaluation Review, 43(1):97– 108, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marasevic:2015:RAR

[2478] Jelena Marasevic, Jin Zhou, Harish Krishnaswamy, Yuan Zhong, and Gil Zussman. Resource allocation and rate gains in practical full-duplex systems. ACM SIGMETRICS Performance Evaluation Review, 43(1):109– 122, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gast:2015:TSS

[2479] Nicolas Gast and Benny Van Houdt. Transient and steady-state regime of a family of list-based cache replacement algorithms. ACM SIGMET-RICS Performance Evaluation Review, 43(1):123–136, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kandemir:2015:MRR

[2480] Mahmut Kandemir, Hui Zhao, Xulong Tang, and Mustafa Karakoy. Memory row reuse distance and its role in optimizing application performance. ACM SIGMETRICS Performance Evaluation Review, 43(1):137–149, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2015:SED

[2481] Xiaomeng Chen, Ning Ding, Abhilash Jindal, Y. Charlie Hu, Maruti Gupta, and Rath Vannithamby. Smartphone energy drain in the wild: Analysis and implications. ACM SIGMET-RICS Performance Evaluation Review, 43(1):151–164, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2015:NSB

[2482] Ming Chen, Dean Hildebrand, Geoff Kuenning, Soujanya Shankaranarayana, Bharat Singh, and Erez Zadok. Newer is sometimes better: an evaluation of NFSv4.1. ACM SIGMETRICS Performance Evaluation Review, 43(1):165–176, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Meza:2015:LSS

[2483] Justin Meza, Qiang Wu, Sanjev Kumar, and Onur Mutlu. A large-scale study of flash memory failures in the field. ACM SIGMETRICS Performance Evaluation Review, 43(1):177– 190, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2015:OCO

[2484] Niangjun Chen, Anish Agarwal, Adam Wierman, Siddharth Barman, and Lachlan L. H. Andrew. Online convex optimization using predictions. ACM SIGMETRICS Performance Evaluation Review, 43(1):191–204, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2015:RMC

[2485] Donghyeon Lee, Joonyoung Kim, Hyunmin Lee, and Kyomin Jung. Reliable multiple-choice iterative algorithm for crowdsourcing systems. ACM SIGMETRICS Performance Evaluation Review, 43(1):205–216, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2015:OLA

[2486] Yang Liu and Mingyan Liu. An online learning approach to improving the quality of crowd-sourcing. ACM SIGMETRICS Performance Evaluation Review, 43(1):217–230, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Combes:2015:LRR

[2487] Richard Combes, Stefan Magureanu, Alexandre Proutiere, and Cyrille Laroche. Learning to rank: Regret lower bounds and efficient algorithms. ACM SIGMETRICS Performance Evaluation Review, 43(1):231– 244, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Combes:2015:BBR

[2488] Richard Combes, Chong Jiang, and Rayadurgam Srikant. Bandits with budgets: Regret lower bounds and optimal algorithms. ACM SIGMET-RICS Performance Evaluation Review, 43(1):245–257, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chalermsook:2015:SNM

[2489] Parinya Chalermsook, Atish Das Sarma, Ashwin Lall, and Danupon Nanongkai. Social network monetization via sponsored viral marketing. ACM SIGMETRICS Performance Evaluation Review, 43(1):259– 270, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fanti:2015:SVS

[2490] Giulia Fanti, Peter Kairouz, Sewoong Oh, and Pramod Viswanath. Spy vs. spy: Rumor source obfuscation. ACM SIGMETRICS Performance Evaluation Review, 43(1):271–284, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Massoulie:2015:GBT

[2491] Laurent Massoulié, Mesrob I. Ohannessian, and Alexandre Proutière. Greedy-Bayes for targeted news dissemination. ACM SIGMETRICS Performance Evaluation Review, 43(1):285– 296, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tarihi:2015:DAD

[2492] Mojtaba Tarihi, Hossein Asadi, and Hamid Sarbazi-Azad. DiskAccel: Accelerating disk-based experiments by representative sampling. ACM SIGMETRICS Performance Evaluation Review, 43(1):297–308, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jin:2015:CPI

[2493] Ye Jin, Xiaosong Ma, Mingliang Liu, Qing Liu, Jeremy Logan, Norbert Podhorszki, Jong Youl Choi, and Scott Klasky. Combining phase identification and statistic modeling for automated parallel benchmark generation. ACM SIGMETRICS Performance Evaluation Review, 43(1):309–320, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xie:2015:PDC

[2494] Qiaomin Xie, Xiaobo Dong, Yi Lu, and Rayadurgam Srikant. Power of d choices for large-scale bin packing: a loss model. ACM SIGMET-RICS Performance Evaluation Review, 43(1):321–334, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rizk:2015:CBF

[2495] Amr Rizk, Felix Poloczek, and Florin Ciucu. Computable bounds in forkjoin queueing systems. ACM SIGMET-RICS Performance Evaluation Review, 43(1):335–346, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gardner:2015:RLR

[2496] Kristen Gardner, Samuel Zbarsky, Sherwin Doroudi, Mor Harchol-Balter, and Esa Hyytia. Reducing latency via redundant requests: Exact analysis. ACM SIGMETRICS Performance Evaluation Review, 43(1):347– 360, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Soltan:2015:JCP

[2497] Saleh Soltan, Mihalis Yannakakis, and Gil Zussman. Joint cyber and physical attacks on power grids: Graph theoretical approaches for information recovery. ACM SIGMETRICS Performance Evaluation Review, 43(1):361– 374, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shah:2015:IFH

[2498] Virag Shah and Gustavo de Veciana.

Impact of fairness and heterogeneity on delays in large-scale content delivery networks. ACM SIGMET-RICS Performance Evaluation Review, 43(1):375–387, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

He:2015:FIB

[2499] Ting He, Chang Liu, Ananthram Swami, Don Towsley, Theodoros Salonidis, Andrei Iu. Bejan, and Paul Yu. Fisher information-based experiment design for network tomography. ACM SIGMETRICS Performance Evaluation Review, 43(1):389–402, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Birke:2015:WVM

[2500] Robert Birke, Mathias Björkqvist, Cyriel Minkenberg, Martin Schmatz, and Lydia Y. Chen. When virtual meets physical at the edge: a field study on datacenters’ virtual traffic. ACM SIGMETRICS Performance Evaluation Review, 43(1):403– 415, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xiao:2015:HCV

[2501] Qingjun Xiao, Shigang Chen, Min Chen, and Yibei Ling. Hyper-compact virtual estimators for big network data based on register sharing. ACM SIGMETRICS Performance Evaluation Review, 43(1):417–428, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kotronis:2015:IPI

[2502] Vasileios Kotronis, Rowan Klöti, Matthias Rost, Panagiotis Georgopoulos, Bernhard Ager, Stefan Schmid, and Xenofontas Dimitropoulos. Investigating the potential of the inter-IXP multigraph for the provisioning of guaranteed end-to-end services. ACM SIGMETRICS Performance Evaluation Review, 43(1):429–430, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Singh:2015:MSA

[2503] Rahul Singh and Alexander Stolyar. MaxWeight scheduling: Asymptotic behavior of unscaled queue-differentials in heavy traffic. ACM SIGMET-RICS Performance Evaluation Review, 43(1):431–432, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fuerst:2015:KTE

[2504] Carlo Fuerst, Stefan Schmid, Lalith Suresh, and Paolo Costa. Kraken: Towards elastic performance guarantees in multi-tenant data centers. ACM SIGMETRICS Performance Evaluation Review, 43(1):433–434, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

He:2015:LSD

[2505] Keqiang He, Junaid Khalid, Sourav Das, Aaron Gember-Jacobson, Chaithan Prakash, Aditya Akella, Li Erran Li, and Marina Thottan. Latency in software defined networks: Measurements and mitigation techniques. ACM SIGMETRICS Performance Evaluation Review, 43(1):435–436, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fu:2015:TSB

[2506] Yongquan Fu and Ernst Biersack. Treestructured Bloom filters for joint optimization of false positive probability and transmission bandwidth. ACM SIGMETRICS Performance Evaluation Review, 43(1):437–438, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghaderi:2015:SSS

[2507] Javad Ghaderi, Sanjay Shakkottai, and Rayadurgam Srikant. Scheduling storms and streams in the cloud. ACM SIGMETRICS Performance Evaluation Review, 43(1):439–440, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Meirom:2015:LED

[2508] Eli A. Meirom, Chris Milling, Constantine Caramanis, Shie Mannor, Sanjay Shakkottai, and Ariel Orda. Localized epidemic detection in networks with overwhelming noise. ACM SIGMET-RICS Performance Evaluation Review, 43(1):441–442, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhou:2015:PBE

[2509] Zhi Zhou, Fangming Liu, and Zongpeng Li. Pricing bilateral electricity trade between smart grids and hybrid green datacenters. ACM SIGMET-RICS Performance Evaluation Review, 43(1):443–444, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krishnasamy:2015:DSR

[2510] Subhashini Krishnasamy, Rajat Sen, Sewoong Oh, and Sanjay Shakkottai. Detecting sponsored recommendations. ACM SIGMETRICS Performance Evaluation Review, 43(1):445– 446, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhao:2015:UPP

[2511] Yong Zhao, Jia Rao, Xiaobo Zhou, and Qing Yi. Understanding parallel performance under interferences in multi-tenant clouds. ACM SIGMET-RICS Performance Evaluation Review, 43(1):447–448, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wu:2015:CIP

[2512] Rui Wu, Jiaming Xu, Rayadurgam Srikant, Laurent Massoulie, Marc Lelarge, and Bruce Hajek. Clustering and inference from pairwise comparisons. ACM SIGMETRICS Performance Evaluation Review, 43(1):449– 450, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Venkatakrishnan:2015:DNO

[2513] Shaileshh Bojja Venkatakrishnan and Pramod Viswanath. Deterministic near-optimal P2P streaming. ACM SIGMETRICS Performance Evaluation Review, 43(1):451–452, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mirhoseini:2015:FTL

[2514] Azalia Mirhoseini, Ebrahim M. Songhori, Bita Darvish Rouhani, and Farinaz Koushanfar. Flexible transformations for learning big data. ACM SIGMET-RICS Performance Evaluation Review, 43(1):453–454, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2015:ECM

[2515] Jian Li, Bainan Xia, Xinbo Geng, Hao Ming, Srinivas Shakkottai, Vijay Subramanian, and Le Xie. Energy coupon: a mean field game perspective on demand response in smart grids. ACM SIGMETRICS Performance Evaluation Review, 43(1):455–456, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Umar:2015:DLA

[2516] Ibrahim Umar, Otto Johan Anshus, and Phuong Hoai Ha. DeltaTree: a locality-aware concurrent search tree. ACM SIGMETRICS Performance Evaluation Review, 43(1):457– 458, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ahmed:2015:DLE

[2517] Faraz Ahmed, Jeffrey Erman, Zihui Ge, Alex X. Liu, Jia Wang, and He Yan. Detecting and localizing endto-end performance degradation for cellular data services. ACM SIGMET-RICS Performance Evaluation Review, 43(1):459–460, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Varloot:2015:SGD

[2518] Rémi Varloot, Ana Busíü’c, and Anne Bouillard. Speeding up Glauber dynamics for random generation of independent sets. ACM SIGMETRICS Performance Evaluation Review, 43 (1):461–462, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2015:OEC

[2519] Linquan Zhang, Zongpeng Li, Chuan Wu, and Shaolei Ren. Online electricity cost saving algorithms for colocation data centers. ACM SIGMET-RICS Performance Evaluation Review, 43(1):463–464, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ducoffe:2015:WTC

[2520] Guillaume Ducoffe, Mathias Lécuyer, Augustin Chaintreau, and Roxana Geambasu. Web transparency for complex targeting: Algorithms, limits, and tradeoffs. ACM SIGMET-RICS Performance Evaluation Review, 43(1):465–466, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:2015:LBO

[2521] Varun Gupta and Ana Radovanovic.

Lagrangian-based online stochastic bin packing. ACM SIGMETRICS Performance Evaluation Review, 43(1):467– 468, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gupta:2015:TCI

[2522] Vani Gupta, Stephen Lee, Prashant Shenoy, Ramesh Sitaraman, and Rahul Urgaonkar. Towards cooling Internetscale distributed networks on the cheap. ACM SIGMETRICS Performance Evaluation Review, 43(1):469– 470, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Clapp:2015:SMQ

[2523] Russell Clapp, Martin Dimitrov, Karthik Kumar, Vish Viswanathan, and Thomas Willhalm. A simple model to quantify the impact of memory latency and bandwidth on performance. ACM SIGMETRICS Performance Evaluation Review, 43(1):471– 472, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xu:2015:PCH

[2524] Qiumin Xu, Huzefa Siyamwala, Mrinmoy Ghosh, Manu Awasthi, Tameesh Suri, Zvika Guz, Anahita Shayesteh, and Vijay Balakrishnan. Performance characterization of hyperscale applicationson on NVMe SSDs. ACM SIGMETRICS Performance Evaluation Review, 43(1):473–474, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Golubchik:2015:SLS

[2525] Leana Golubchik and Bert Zwart. Spatial loss systems: Exact simulation and rare event behavior. ACM SIGMET-RICS Performance Evaluation Review, 43(2):3–6, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Onderwater:2015:LOP

[2526] Martijn Onderwater, Sandjai Bhulai, and Rob van der Mei. Learning optimal policies in Markov decision processes with value function discovery? ACM SIGMETRICS Performance Evaluation Review, 43(2):7–9, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

vanLeeuwaarden:2015:DWS

[2527] Johan S. H. van Leeuwaarden, Britt W. J. Mathijsen, and Fiona Sloothaak. Delayed workload shifting in manyserver systems. ACM SIGMETRICS Performance Evaluation Review, 43 (2):10–12, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cecchi:2015:MFA

[2528] F. Cecchi, S. C. Borst, and J. S. H.

van Leeuwaardena. Mean-field analysis of ultra-dense CSMA networks. ACM SIGMETRICS Performance Evaluation Review, 43(2):13–15, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Patch:2015:PFL

[2529] Brendan Patch, Thomas Taimre, and Yoni Nazarathy. Performance of faulty loss systems with persistent connections. ACM SIGMETRICS Performance Evaluation Review, 43(2):16–18, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shneer:2015:SII

[2530] Seva Shneer and Peter M. van de Ven. Stability and instability of individual nodes in multi-hop wireless CSMA/CA networks. ACM SIGMET-RICS Performance Evaluation Review, 43(2):19–21, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Brun:2015:FMB

[2531] O. Brun, H. Ben Cheikh, and B. J.

Prabhu. A fluid model based heuristic for optimal speed-scaling of multi-class single server queues. ACM SIGMET-RICS Performance Evaluation Review, 43(2):22–23, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jonckheere:2015:GBA

[2532] Matthieu Jonckheere and Seva Shneer.

Gradient bandwidth allocations. ACM SIGMETRICS Performance Evaluation Review, 43(2):24–25, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kleinrouweler:2015:MES

[2533] Jan Willem Kleinrouweler, Sergio Cabrero, Rob van der Mei, and Pablo Cesar. Modeling the effect of sharing policies for network-assisted HTTP adaptive video streaming. ACM SIG-METRICS Performance Evaluation Review, 43(2):26–27, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Patel:2015:HLR

[2534] Naresh M. Patel. Half-latency rule for finding the knee of the latency curve. ACM SIGMETRICS Performance Evaluation Review, 43(2):28–29, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Touati:2015:AJS

[2535] Mikael Touati, Rachid Elazouzi, Marceau Coupechoux, Eitan Altman, and Jean-Marc Kelif. About joint stable user association and resource allocation in multi-rate IEEE 802.11 WLANs. ACM SIGMETRICS Performance Evaluation Review, 43(2):30–31, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wu:2015:AER

[2536] Huaming Wu, Yi Sun, and Katinka Wolter. Analysis of the energy-response time tradeoff for delayed mobile cloud offloading. ACM SIGMETRICS Performance Evaluation Review, 43(2):

33–35, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2015:GMT

[2537] Niangjun Chen, Xiaoqi Ren, Shaolei Ren, and Adam Wierman. Greening multi-tenant data center demand response. ACM SIGMETRICS Performance Evaluation Review, 43(2): 36–38, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2015:PSD

[2538] Shaoquan Zhang, Longbo Huang, Minghua Chen, and Xin Liu. Proactive serving decreases user delay exponentially. ACM SIGMETRICS Performance Evaluation Review, 43(2):39–41, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ren:2015:SAC

[2539] Xiaoqi Ren, Ganesh Ananthanarayanan, Adam Wierman, and Minlan Yu. Speculation-aware cluster scheduling. ACM SIGMETRICS Performance Evaluation Review, 43(2):42–44, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2015:MLE

[2540] Weikun Wang and Giuliano Casale.

Maximum likelihood estimation of closed queueing network demands from queue length data. ACM SIGMET-RICS Performance Evaluation Review, 43(2):45–47, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kesidis:2015:NCP

[2541] G. Kesidis, Y. Shan, B. Urgaonkar, and J. Liebeherr. Network calculus for parallel processing. ACM SIGMET-RICS Performance Evaluation Review, 43(2):48–50, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fiorini:2015:EAS

[2542] Pierre M. Fiorini and Lester Lipsky. Exact analysis of some splitmerge queues. ACM SIGMETRICS Performance Evaluation Review, 43 (2):51–53, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Joshi:2015:QRL

[2543] Gauri Joshi, Emina Soljanin, and Gregory Wornell. Queues with redundancy: Latency-cost analysis. ACM SIGMETRICS Performance Evaluation Review, 43(2):54–56, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berger:2015:MCH

[2544] Daniel S. Berger, Sebastian Henningsen, Florin Ciucu, and Jens B. Schmitt. Maximizing cache hit ratios by variance reduction. ACM SIGMET-RICS Performance Evaluation Review, 43(2):57–59, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2015:MBC

[2545] Jian Tan, Li Zhang, and Yandong Wang. Miss behavior for caching with lease. ACM SIGMETRICS Performance Evaluation Review, 43(2):

60–62, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2015:OGG

[2546] Yudong Yang, Vishal Misra, and Dan Rubenstein. On the optimality of greedy garbage collection for SSDs. ACM SIGMETRICS Performance Evaluation Review, 43(2):63–65, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Spencer:2015:ILM

[2547] Sam Spencer and R. Srikant. On the impossibility of localizing multiple rumor sources in a line graph. ACM SIGMETRICS Performance Evaluation Review, 43(2):66–68, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gast:2015:PTC

[2548] Nicolas Gast. The power of two choices on graphs: the pair-approximation is accurate? ACM SIGMETRICS Performance Evaluation Review, 43 (2):69–71, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Maguluri:2015:HTB

[2549] Siva Theja Maguluri and R. Srikant.

Heavy-traffic behavior of the MaxWeight algorithm in a switch with uniform traffic. ACM SIGMETRICS Performance Evaluation Review, 43(2):72–74, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Busic:2015:AOB

[2550] Ana Bušić and Sean Meyn. Approximate optimality with bounded regret in dynamic matching models. ACM SIGMETRICS Performance Evaluation Review, 43(2):75–77, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2015:CEL

[2551] Yingdong Lu, Mark S. Squillante, Chai Wah Wu, and Bo Zhang. On the control of epidemic-like stochastic processes with time-varying behavior? ACM SIGMETRICS Performance Evaluation Review, 43(2): 78–80, September 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Canini:2015:HMP

[2552] Marco Canini, James Kempf, and Stefan Schmid. How many planet-wide leaders should there be? ACM SIGMETRICS Performance Evaluation Review, 43(3):3–6, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2015:USR

[2553] Da Wang, Gauri Joshi, and Gregory Wornell. Using straggler replication to reduce latency in large-scale parallel computing. ACM SIGMETRICS Performance Evaluation Review, 43 (3):7–11, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gandhi:2015:ANA

[2554] Anshul Gandhi and Justin Chan. Analyzing the network for AWS distributed cloud computing. ACM SIGMET-RICS Performance Evaluation Review, 43(3):12–15, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jia:2015:PCA

[2555] Rui Jia, Sherif Abdelwahed, and Abdelkarim Erradi. A predictive control approach for fault management of computing systems. ACM SIGMET-RICS Performance Evaluation Review, 43(3):16–20, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Netto:2015:ARI

[2556] Hylson Vescovi Netto, Lau Cheuk Lung, Tulio Alberton Ribeiro, Miguel Correia, and Aldelir Fernando Luiz. Anticipating requests to improve performance and reduce costs in cloud storage. ACM SIGMETRICS Performance Evaluation Review, 43(3):21–24, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lavi:2015:ARP

[2557] Nadav Lavi and Hanoch Levy. Admit or reject? Preserve or drop?: Operational dilemmas upon server failures on the cloud. ACM SIGMET-RICS Performance Evaluation Review, 43(3):25–29, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2015:TDT

[2558] Tianrong Zhang and Yufeng Xin.

Towards designing a truthful online auction framework for deadline-aware cloud resource allocation. ACM SIGMETRICS Performance Evaluation Review, 43(3):30–33, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tran:2015:CCD

[2559] Nguyen H. Tran, Cuong T. Do, Choong Seon Hong, Shaolei Ren, and Zhu Han. Coordinated colocation datacenters for economic demand responce. ACM SIGMETRICS Performance Evaluation Review, 43(3):34–37, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ludwig:2015:DCM

[2560] Arne Ludwig and Stefan Schmid. Distributed cloud market: Who benefits from specification flexibilities? ACM SIGMETRICS Performance Evaluation Review, 43(3):38–41, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mao:2015:DAD

[2561] Bo Mao and Suzhen Wu. Deduplicationassisted data reduction and distribution in cloud-of-clouds. ACM SIG-METRICS Performance Evaluation Review, 43(3):42, December 2015. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gandhi:2015:OLB

[2562] Anshul Gandhi, Naman Mittal, and Xi Zhang. Optimal load-balancing for heterogeneous clusters. ACM SIGMETRICS Performance Evaluation Review, 43(3):43, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Le:2015:ECA

[2563] Tan N. Le, Bong Jun Choi, and Pradipta De. Energy cost aware scheduling of MapReduce jobs across geographically distributed nodes. ACM SIGMETRICS Performance Evaluation Review, 43(3):44, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bhojwani:2015:IDC

[2564] Sushil Bhojwani, Matt Hemmings, Dan Ingalls, Jens Lincke, Robert Krahn, David Lary, Rick McGeer, Glenn Ricart, Marko Roder, Yvonne Coady, and Ulrike Stege. The ignite distributed collaborative visualization system. ACM SIGMETRICS Performance Evaluation Review, 43(3):45–46, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Maille:2015:ICD

[2565] Patrick Maillé and Bruno Tuffin. Impact of content delivery networks on service and content innovation. ACM SIGMETRICS Performance Evaluation Review, 43(3):49–52, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ahuja:2015:PDW

[2566] Kartik Ahuja, Simpson Zhang, and Mihaela van der Schaar. The population dynamics of Websites: [extended abstract]. ACM SIGMET-RICS Performance Evaluation Review, 43(3):53–56, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Luo:2015:PPP

[2567] Yuan Luo, Nihar B. Shah, Jianwei Huang, and Jean Walrand. Parametric prediction from parametric agents? ACM SIGMETRICS Performance Evaluation Review, 43(3):57, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Acemoglu:2015:PCN

[2568] Daron Acemoglu, Ali Makhdoumi, Azarakhsh Malekian, and Asu Ozdaglar. Privacy-constrained network formation. ACM SIGMETRICS Performance Evaluation Review, 43(3):58, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ramachandran:2015:NEP

[2569] Arthi Ramachandran and Augustin Chaintreau. The network effect of privacy choices. ACM SIGMETRICS Performance Evaluation Review, 43 (3):59–62, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Afrasiabi:2015:CBP

[2570] M. H. Afrasiabi and R. Guérin. Choicebased pricing for user-provided connectivity? ACM SIGMETRICS Performance Evaluation Review, 43(3):63–66, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Meir:2015:PWG

[2571] Reshef Meir and David Parkes. Playing the wrong game: Smoothness bounds for congestion games with behavioral biases. ACM SIGMETRICS Performance Evaluation Review, 43 (3):67–70, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Feldman:2015:CSE

[2572] Michal Feldman and Ophir Friedler.

Convergence to strong equilibrium in network design games. ACM SIGMET-RICS Performance Evaluation Review, 43(3):71, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Touati:2015:CSA

[2573] Mikael Touati, Rachid El-Azouzi, Marceau Coupechoux, Eitan Altman, and Jean-Marc Kelif. Core stable algorithms for coalition games with complementarities and peer effects. ACM SIGMETRICS Performance Evaluation Review, 43(3):72–75, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kilcioglu:2015:RMC

[2574] Cinar Kilcioglu and Costis Maglaras.

Revenue maximization for cloud computing services. ACM SIGMETRICS Performance Evaluation Review, 43 (3):76, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kulkarni:2015:DCM

[2575] Janardhan Kulkarni and Vahab Mirrokni. Dynamic coordination mechanisms: [extended abstract]. ACM SIGMETRICS Performance Evaluation Review, 43(3):77, December 2015.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tavafoghi:2015:SCU

[2576] Hamidreza Tavafoghi and Demosthenis Teneketzis. Sequential contracts for uncertain electricity resources. ACM SIGMETRICS Performance Evaluation Review, 43(3):78–81, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Simhon:2015:ISI

[2577] Eran Simhon and David Starobinski. On the impact of sharing information in advance reservation systems. ACM SIGMETRICS Performance Evaluation Review, 43(3):82, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ceppi:2015:PPS

[2578] Sofia Ceppi and Ian Kash. Personalized payments for storage-as-a-service. ACM SIGMETRICS Performance Evaluation Review, 43(3):83–86, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Benjaafar:2015:MAC

[2579] Saif Benjaafar, Guangwen Kong, and Xiang Li. Modeling and analysis of collaborative consumption in peer-to-peer car sharing. ACM SIGMET-RICS Performance Evaluation Review, 43(3):87–90, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krishnamurthy:2016:PCC

[2580] Diwakar Krishnamurthy and Anne Koziolek. Performance challenges, current bad practices, and hints in PaaS cloud application design. ACM SIGMETRICS Performance Evaluation Review, 43(4):3–12, March 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Heinrich:2016:ART

[2581] Robert Heinrich. Architectural runtime models for performance and privacy analysis in dynamic cloud applications? ACM SIGMETRICS Performance Evaluation Review, 43(4):13–22, March 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2016:DTD

[2582] Zheng Li, Liam O’Brien, and Maria Kihl. DoKnowMe: Towards a domain knowledge-driven methodology for performance evaluation. ACM SIGMET-RICS Performance Evaluation Review, 43(4):23–32, March 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Grottke:2016:ESC

[2583] Michael Grottke, Alberto Avritzer, Daniel S. Menasché, Leandro P. de Aguiar, and Eitan Altman. On the efficiency of sampling and countermeasures to critical-infrastructure-targeted malware campaigns. ACM SIGMET-RICS Performance Evaluation Review, 43(4):33–42, March 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rehmann:2016:PMS

[2584] Kim-Thomas Rehmann, Changyun Seo, Dongwon Hwang, Binh Than Truong, Alexander Boehm, and Dong Hun Lee. Performance monitoring in SAP HANA’s continuous integration process. ACM SIGMET-RICS Performance Evaluation Review, 43(4):43–52, March 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nambiar:2016:MDS

[2585] Manoj Nambiar, Ajay Kattepur, Gopal Bhaskaran, Rekha Singhal, and Subhasri Duttagupta. Model driven software performance engineering: Current challenges and way ahead. ACM SIGMETRICS Performance Evaluation Review, 43(4):53–62, March 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gamarnik:2016:DMM

[2586] David Gamarnik, John N. Tsitsiklis, and Martin Zubeldia. Delay, memory, and messaging tradeoffs in distributed service systems. ACM SIGMETRICS Performance Evaluation Review, 44(1): 1–12, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Maguluri:2016:OHT

[2587] Siva Theja Maguluri, Sai Kiran Burle, and R. Srikant. Optimal heavy-traffic queue length scaling in an incompletely saturated switch. ACM SIGMET-RICS Performance Evaluation Review, 44(1):13–24, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zeng:2016:NSC

[2588] Yun Zeng, Augustin Chaintreau, Don Towsley, and Cathy H. Xia. A necessary and sufficient condition for throughput scalability of fork and join networks with blocking. ACM SIGMETRICS Performance Evaluation Review, 44(1):25–36, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Garetto:2016:GTB

[2589] Michele Garetto, Emilio Leonardi, and Giovanni Luca Torrisi. Generalized threshold-based epidemics in random graphs: The power of extreme values. ACM SIGMETRICS Performance Evaluation Review, 44(1):37–50, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Buchnik:2016:RRG

[2590] Eliav Buchnik and Edith Cohen. Reverse ranking by graph structure: Model and scalable algorithms. ACM SIGMETRICS Performance Evaluation Review, 44(1):51–62, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cullina:2016:IAC

[2591] Daniel Cullina and Negar Kiyavash.

Improved achievability and converse bounds for Erdős–Rényi graph matching. ACM SIGMETRICS Performance Evaluation Review, 44(1):63–72, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harchol-Balter:2016:BMT

[2592] Mor Harchol-Balter. A better model for task assignment in server farms: How replication can help. ACM SIGMET-RICS Performance Evaluation Review, 44(1):73, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Venkatakrishnan:2016:CCS

[2593] Shaileshh Bojja Venkatakrishnan, Mohammad Alizadeh, and Pramod Viswanath. Costly circuits, submodular schedules and approximate Carathéodory theorems. ACM SIG-METRICS Performance Evaluation Review, 44(1):75–88, June 2016. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Narayanan:2016:RLT

[2594] Shankaranarayanan Puzhavakath Narayanan, Yun Seong Nam, Ashiwan Sivakumar, Balakrishnan Chandrasekaran, Bruce Maggs, and Sanjay Rao. Reducing latency through page-aware management of Web objects by content delivery networks. ACM SIGMETRICS Performance Evaluation Review, 44(1): 89–100, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ferragut:2016:OTC

[2595] Andrés Ferragut, Ismael Rodriguez, and Fernando Paganini. Optimizing TTL caches under heavy-tailed demands. ACM SIGMETRICS Performance Evaluation Review, 44(1):101– 112, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ioannidis:2016:ACN

[2596] Stratis Ioannidis and Edmund Yeh. Adaptive caching networks with optimality guarantees. ACM SIGMET-RICS Performance Evaluation Review, 44(1):113–124, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jacquet:2016:BMT

[2597] Philippe Jacquet. Breathing mankind thoughts. ACM SIGMETRICS Performance Evaluation Review, 44(1):125, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shamsi:2016:UCU

[2598] Zain Shamsi and Dmitri Loguinov.

Unsupervised clustering under temporal feature volatility in network stack fingerprinting. ACM SIGMETRICS Performance Evaluation Review, 44 (1):127–138, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dai:2016:NBF

[2599] Haipeng Dai, Yuankun Zhong, Alex X.

Liu, Wei Wang, and Meng Li. Noisy Bloom filters for multi-set membership testing. ACM SIGMETRICS Performance Evaluation Review, 44(1):139– 151, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fanti:2016:RSO

[2600] Giulia Fanti, Peter Kairouz, Sewoong Oh, Kannan Ramchandran, and Pramod Viswanath. Rumor source obfuscation on irregular trees. ACM SIGMETRICS Performance Evaluation Review, 44(1):153–164, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Avrachenkov:2016:IOL

[2601] Konstantin Avrachenkov, Bruno Ribeiro, and Jithin K. Sreedharan. Inference in OSNs via lightweight partial crawls. ACM SIGMETRICS Performance Evaluation Review, 44(1):165– 177, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gabielkov:2016:SCW

[2602] Maksym Gabielkov, Arthi Ramachandran, Augustin Chaintreau, and Arnaud Legout. Social clicks: What and who gets read on Twitter? ACM SIGMETRICS Performance Evaluation Review, 44(1):179–192, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2016:UPO

[2603] Niangjun Chen, Joshua Comden, Zhenhua Liu, Anshul Gandhi, and Adam Wierman. Using predictions in online optimization: Looking forward with an eye on the past. ACM SIGMET-RICS Performance Evaluation Review, 44(1):193–206, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bresler:2016:CFL

[2604] Guy Bresler, Devavrat Shah, and Luis Filipe Voloch. Collaborative filtering with low regret. ACM SIGMET-RICS Performance Evaluation Review, 44(1):207–220, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2016:ALD

[2605] Jia Liu. Achieving low-delay and fast-convergence in stochastic network optimization: a Nesterovian approach. ACM SIGMETRICS Performance Evaluation Review, 44(1):221– 234, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zheng:2016:VCV

[2606] Liang Zheng, Carlee Joe-Wong, Christopher G. Brinton, Chee Wei Tan, Sangtae Ha, and Mung Chiang. On the viability of a cloud virtual service provider. ACM SIGMETRICS Performance Evaluation Review, 44(1):235– 248, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2016:VPS

[2607] Weina Wang, Lei Ying, and Junshan Zhang. The value of privacy: Strategic data subjects, incentive mechanisms and fundamental limits. ACM SIGMETRICS Performance Evaluation Review, 44(1):249–260, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2016:IDM

[2608] Yuanjie Li, Haotian Deng, Jiayao Li, Chunyi Peng, and Songwu Lu. Instability in distributed mobility management: Revisiting configuration management in 3G/4G mobile networks. ACM SIGMETRICS Performance Evaluation Review, 44(1):261– 272, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ludwig:2016:TSN

[2609] Arne Ludwig, Szymon Dudyzc, Matthias Rost, and Stefan Schmid. Transiently secure network updates. ACM SIG-METRICS Performance Evaluation Review, 44(1):273–284, June 2016. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ying:2016:AEM

[2610] Lei Ying. On the approximation error of mean-field models. ACM SIGMET-RICS Performance Evaluation Review, 44(1):285–297, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jiang:2016:DIC

[2611] Bo Jiang, Daniel R. Figueiredo, Bruno Ribeiro, and Don Towsley. On the duration and intensity of competitions in nonlinear Pólya urn processes with fitness. ACM SIGMETRICS Performance Evaluation Review, 44(1):299– 310, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jonckheere:2016:AIL

[2612] Matthieu Jonckheere and Balakrishna J. Prabhu. Asymptotics of insensitive load balancing and blocking phases. ACM SIGMETRICS Performance Evaluation Review, 44(1):311– 322, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chang:2016:ULV

[2613] Kevin K. Chang, Abhijith Kashyap, Hasan Hassan, Saugata Ghose, Kevin Hsieh, Donghyuk Lee, Tianshi Li, Gennady Pekhimenko, Samira Khan, and Onur Mutlu. Understanding latency variation in modern DRAM chips: Experimental characterization, analysis, and optimization. ACM SIGMET-RICS Performance Evaluation Review, 44(1):323–336, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yaniv:2016:HDC

[2614] Idan Yaniv and Dan Tsafrir. Hash, don’t cache (the page table). ACM SIGMETRICS Performance Evaluation Review, 44(1):337–350, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jog:2016:ECC

[2615] Adwait Jog, Onur Kayiran, Ashutosh Pattnaik, Mahmut T. Kandemir, Onur Mutlu, Ravishankar Iyer, and Chita R. Das. Exploiting core criticality for enhanced GPU performance. ACM SIGMETRICS Performance Evaluation Review, 44(1):351–363, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nguyen:2016:SSR

[2616] Lam M. Nguyen and Alexander L. Stolyar. A service system with randomly behaving on-demand agents. ACM SIGMETRICS Performance Evaluation Review, 44(1):365–366, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Novakovic:2016:ALI

[2617] Stanko Novakovic, Alexandros Daglis, Edouard Bugnion, Babak Falsafi, and Boris Grot. An analysis of load imbalance in scale-out data serving. ACM SIGMETRICS Performance Evaluation Review, 44(1):367–368, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cao:2016:APC

[2618] Yi Cao, Javad Nejati, Pavan Maguluri, Aruna Balasubramanian, and Anshul Gandhi. Analyzing the power consumption of the mobile page load. ACM SIGMETRICS Performance Evaluation Review, 44(1):369– 370, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Qureshi:2016:ATL

[2619] Mubashir Adnan Qureshi, Ajay Mahimkar, Lili Qiu, Zihui Ge, Sarat Puthenpura, Nabeel Mir, and Sanjeev Ahuja. Automated test location selection for cellular network upgrades. ACM SIGMETRICS Performance Evaluation Review, 44(1):371– 372, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2016:CCA

[2620] Wenjie Liu, Ping Huang, Kun Tang, Ke Zhou, and Xubin He. CAR: a compression-aware refresh approach to improve memory performance and energy efficiency. ACM SIGMET-RICS Performance Evaluation Review, 44(1):373–374, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Poloczek:2016:CER

[2621] Felix Poloczek and Florin Ciucu. Contrasting effects of replication in parallel systems: From overload to underload and back. ACM SIGMET-RICS Performance Evaluation Review, 44(1):375–376, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

VanHoudt:2016:EBR

[2622] Benny Van Houdt. Explicit backoff rates for achieving target throughputs in CSMA/CA networks. ACM SIGMETRICS Performance Evaluation Review, 44(1):377–379, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2016:FDR

[2623] Liang Liu, Yating Wang, Lance Fortnow, Jin Li, and Jun Xu. Freestyle dancing: Randomized algorithms for dynamic storage load-balancing. ACM SIGMETRICS Performance Evaluation Review, 44(1):381–382, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ren:2016:JDP

[2624] Xiaoqi Ren, Palma London, Juba Ziani, and Adam Wierman. Joint data purchasing and data placement in a geo-distributed data market. ACM SIGMETRICS Performance Evaluation Review, 44(1):383–384, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mukhopadhyay:2016:MRB

[2625] Arpan Mukhopadhyay, Ravi R.

Mazumdar, and Rahul Roy. Majority rule based opinion dynamics with biased and stubborn agents. ACM SIGMETRICS Performance Evaluation Review, 44(1):385–386, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raja:2016:MFE

[2626] Vamseedhar Reddyvari Raja, Vinod Ramaswamy, Srinivas Shakkottai, and Vijay Subramanian. Mean field equilibria of pricing games in Internet marketplaces. ACM SIGMETRICS Performance Evaluation Review, 44(1):387– 388, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shafaei:2016:MSD

[2627] Mansour Shafaei, Mohammad Hossein Hajkazemi, Peter Desnoyers, and Abutalib Aghayev. Modeling SMR drive performance. ACM SIGMET-RICS Performance Evaluation Review, 44(1):389–390, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Combes:2016:MSF

[2628] Richard Combes, Habib Sidi, and Salah Elayoubi. Multipath streaming: Fundamental limits and efficient algorithms. ACM SIGMETRICS Performance Evaluation Review, 44(1):391– 392, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shekaramiz:2016:NCA

[2629] Alireza Shekaramiz, Jorg Liebeherr, and Almut Burchard. Network calculus analysis of a feedback system with random service. ACM SIGMET-RICS Performance Evaluation Review, 44(1):393–394, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ahmed:2016:QAL

[2630] Adnan Ahmed, Zubair Shafiq, and Amir Khakpour. QoE analysis of a large-scale live video streaming event. ACM SIGMETRICS Performance Evaluation Review, 44(1):395– 396, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2016:SRL

[2631] Sen Yang, Bill Lin, and Jun Xu. Safe randomized load-balanced switching by diffusing extra loads. ACM SIGMET-RICS Performance Evaluation Review, 44(1):397–398, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ray:2016:SSC

[2632] Avik Ray, Sujay Sanghavi, and Sanjay Shakkottai. Searching for a single community in a graph. ACM SIGMET-RICS Performance Evaluation Review, 44(1):399–400, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2016:SMY

[2633] Daiping Liu, Xing Gao, Mingwei Zhang, and Haining Wang. Shoot for the Moon and you will never miss: Characterizing and detecting aimbots in online games. ACM SIGMET-RICS Performance Evaluation Review, 44(1):401–402, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Giovanidis:2016:SML

[2634] Anastasios Giovanidis and Apostolos Avranas. Spatial multi-LRU caching for wireless networks with coverage overlaps. ACM SIGMETRICS Performance Evaluation Review, 44(1):403– 405, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Narayanan:2016:SFD

[2635] Iyswarya Narayanan, Di Wang, Myeongjae Jeon, Bikash Sharma, Laura Caulfield, Anand Sivasubramaniam, Ben Cutler, Jie Liu, Badriddine Khessib, and Kushagra Vaid. SSD failures in datacenters: What, when and why? ACM SIGMETRICS Performance Evaluation Review, 44(1):407– 408, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gardner:2016:PCR

[2636] Kristen Gardner, Samuel Zbarsky, Mor Harchol-Balter, and Alan Scheller-Wolf. The power of d choices for redundancy. ACM SIGMETRICS Performance Evaluation Review, 44(1):409– 410, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2016:TBB

[2637] Hui Wang and Peter Varman. Timebased bandwidth allocation for heterogeneous storage. ACM SIGMET-RICS Performance Evaluation Review, 44(1):411–413, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2016:TMR

[2638] Wei Wang, Baochun Li, Ben Liang, and Jun Li. Towards multi-resource fair allocation with placement constraints. ACM SIGMETRICS Performance Evaluation Review, 44(1):415– 416, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xie:2016:TDR

[2639] Hong Xie, Richard T. B. Ma, and John C. S. Lui. Trading discount for reputation?: On the design and analysis of Ecommerce discount mechanisms. ACM SIGMETRICS Performance Evaluation Review, 44(1):417–418, June 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2016:ETI

[2640] Mark S. Squillante. Estimation of the traffic intensity in a piecewisestationary Mt/Gt/1 queue with probing. ACM SIGMETRICS Performance Evaluation Review, 44(2):3–5, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gast:2016:CLF

[2641] Nicolas Gast. Construction of Lyapunov functions via relative entropy with application to caching. ACM SIGMETRICS Performance Evaluation Review, 44(2):6–8, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Domingues:2016:SPT

[2642] Guilherme Domingues, Edmundo de Souza e Silva, Rosa Leão, Daniel Menasché, and Don Towsley. Search and placement in tiered cache networks. ACM SIGMETRICS Performance Evaluation Review, 44(2):9–11, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cai:2016:GIS

[2643] Kechao Cai, Hong Xie, and John C. S. Lui. Generalization of information spreading forensics via sequential dependent snapshots. ACM SIGMET-RICS Performance Evaluation Review, 44(2):12–14, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gelenbe:2016:ASS

[2644] Erol Gelenbe. Agreement in spins and social networks. ACM SIGMET-RICS Performance Evaluation Review, 44(2):15–17, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Doncel:2016:MFG

[2645] Josu Doncel, Nicolas Gast, and Bruno Gaujal. Are mean-field games the limits of finite stochastic games? ACM SIGMETRICS Performance Evaluation Review, 44(2):18–20, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Feinberg:2016:SOS

[2646] Eugene A. Feinberg and Yan Liang. Structure of optimal solutions to periodic-review total-cost stochastic inventory control problems. ACM SIG-METRICS Performance Evaluation Review, 44(2):21–23, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fricker:2016:ADR

[2647] Christine Fricker, Fabrice Guillemin, and Philippe Robert. Analysis of downgrading for resource allocation. ACM SIGMETRICS Performance Evaluation Review, 44(2):24–26, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2016:RBD

[2648] Yingdong Lu, Siva Theja Maguluri, Mark S. Squillante, and Chai Wah Wu. Risk-based dynamic allocation of computing resources. ACM SIGMET-RICS Performance Evaluation Review, 44(2):27–29, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sermpezis:2016:IDS

[2649] Pavlos Sermpezis and Xenofontas Dimitropoulos. Inter-domain SDN: Analysing the effects of routing centralization on BGP convergence time. ACM SIGMETRICS Performance Evaluation Review, 44(2):30–32, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gardner:2016:URT

[2650] Kristen Gardner, Samuel Zbarsky, Mark Velednitsky, Mor Harchol-Balter, and Alan Scheller-Wolf. Understanding response time in the redundancyd system. ACM SIGMETRICS Performance Evaluation Review, 44(2): 33–35, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mukherjee:2016:UPD

[2651] Debankur Mukherjee, Sem Borst, Johan van Leeuwaarden, and Phil Whiting. Universality of power-of-d load balancing schemes. ACM SIGMET-RICS Performance Evaluation Review, 44(2):36–38, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Juneja:2016:LQ

[2652] S. Juneja and D. Manjunath. To lounge or to queue up. ACM SIGMET-RICS Performance Evaluation Review, 44(2):39–41, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stolyar:2016:LSS

[2653] Alexander L. Stolyar. Large-scale service systems with packing constraints and heterogeneous servers. ACM SIGMETRICS Performance Evaluation Review, 44(2):42–44, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2016:DOS

[2654] Y. Lu, S. T. Maguluri, M. S. Squillante, and T. Suk. Delay-optimal scheduling for some input-queued switches. ACM SIGMETRICS Performance Evaluation Review, 44(2):45–47, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ruhi:2016:OPM

[2655] Navid Azizan Ruhi, Niangjun Chen, Krishnamurthy Dvijotham, and Adam Wierman. Opportunities for price manipulation by aggregators in electricity markets. ACM SIGMETRICS Performance Evaluation Review, 44 (2):49–51, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Comden:2016:OLC

[2656] Joshua Comden, Zhenhua Liu, and Yue Zhao. Optimizing the level of commitment in demand response. ACM SIGMETRICS Performance Evaluation Review, 44(2):52–67, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhou:2016:EDR

[2657] Ruiting Zhou, Zongpeng Li, and Chuan Wu. An emergency demand response mechanism for cloud computing. ACM SIGMETRICS Performance Evaluation Review, 44(2):58–63, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Neglia:2016:GLB

[2658] Giovanni Neglia, Matteo Sereno, and Giuseppe Bianchi. Geographical load balancing across green datacenters: a mean field analysis. ACM SIGMET-RICS Performance Evaluation Review, 44(2):64–69, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Facchini:2016:ESB

[2659] Alessandro Facchini, Cristina Rottondi, and Giacomo Verticale. Emergence of shared behaviour in distributed scheduling systems for domestic appliances. ACM SIGMET-RICS Performance Evaluation Review, 44(2):70–75, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rossi:2016:AEE

[2660] Claudio Rossi, Manuel Gaetani, and Antonio Defina. AURORA: an energy efficient public lighting IoT system for smart cities. ACM SIGMET-RICS Performance Evaluation Review, 44(2):76–81, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dalmasso:2016:RRM

[2661] Mattia Dalmasso, Michela Meo, and Daniela Renga. Radio resource management for improving energy self-sufficiency of green mobile networks. ACM SIGMETRICS Performance Evaluation Review, 44(2):82–87, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fan:2016:BSA

[2662] Xiaoyi Fan, Feng Wang, and Jiangchuan Liu. Boosting service availability for base stations of cellular networks by event-driven battery profiling. ACM SIGMETRICS Performance Evaluation Review, 44(2):88–93, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2016:TPE

[2663] Jianyuan Lu, Liang Liu, Jun “Jim” Xu, and Bin Liu. Toward power-efficient backbone routers. ACM SIGMET-RICS Performance Evaluation Review, 44(2):94–99, September 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vaze:2016:OBT

[2664] Rahul Vaze and Marceau Coupechoux. Online budgeted truthful matching. ACM SIGMETRICS Performance Evaluation Review, 44(3):3– 6, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lim:2016:CRS

[2665] Yongwhan Lim, Asuman Ozdaglar, and Alexander Teytelboym. Competitive rumor spread in social networks. ACM SIGMETRICS Performance Evaluation Review, 44(3):7–14, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goel:2016:NFC

[2666] Arpit Goel, Vijay Kamble, Siddhartha Banerjee, and Ashish Goel. Network formation of coalition loyalty programs. ACM SIGMETRICS Performance Evaluation Review, 44(3):15–20, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harder:2016:TSG

[2667] Reed Harder and Vikrant Vaze. Twostage game theoretic modelling of airline frequency and fare competition. ACM SIGMETRICS Performance Evaluation Review, 44(3):21, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hota:2016:STG

[2668] Ashish R. Hota, Harsha Honnappa, and Shreyas Sundaram. The superpositiontraffic game. ACM SIGMETRICS Performance Evaluation Review, 44 (3):22–25, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Reiffers-Masson:2016:TPD

[2669] Alexandre Reiffers-Masson, Eduardo Hargreaves, Eitan Altman, Wouter Caarls, and Daniel S. Menasché. Timelines are publisher-driven caches: Analyzing and shaping timeline networks. ACM SIGMETRICS Performance Evaluation Review, 44(3):26–29, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shan:2016:SFU

[2670] Y. Shan, C. Lo Prete, G. Kesidis, and D. J. Miller. A simulation framework for uneconomic virtual bidding in day-ahead electricity markets: Short talk. ACM SIGMETRICS Performance Evaluation Review, 44(3):30, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

LEcuyer:2016:SNN

[2671] Pierre L’Ecuyer, Patrick Maillé, Nicolás Stier-Moses, and Bruno Tuffin. Search (non-)Neutrality and impact on innovation: Short talk. ACM SIGMETRICS Performance Evaluation Review, 44(3):31, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ma:2016:PSE

[2672] Qian Ma, Jianwei Huang, Tamer Basar, Ji Liu, and Xudong Chen. Pricing for sharing economy with reputation. ACM SIGMETRICS Performance Evaluation Review, 44(3):32, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gregoire:2016:PHD

[2673] J.-Ch. Grégoire, Angèle M. Hamel, and D. Marc Kilgour. Pricing for a hybrid delivery model of video streaming. ACM SIGMETRICS Performance Evaluation Review, 44(3):33–36, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Antonopoulos:2016:ISP

[2674] Angelos Antonopoulos, Chiara Perillo, and Christos Verikoukis. Internet service providers vs. over-the-top companies: Friends or foes? — short talk. ACM SIGMETRICS Performance Evaluation Review, 44(3):37, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xia:2016:HMY

[2675] Chaolun Xia, Saikat Guha, and S. Muthukrishnan. How much is your attention worth?: Analysis of prices in LinkedIn advertising network — short talk. ACM SIGMETRICS Performance Evaluation Review, 44(3):38, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nguyen:2016:PFR

[2676] Thi Thu Han Nguyen, Olivier Brun, and Balakrishna J. Prabhu. Performance of a fixed reward incentive scheme for two-hop DTNs with competing relays: Short talk. ACM SIGMETRICS Performance Evaluation Review, 44(3):39, December 2016. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Onderwater:2017:TMI

[2677] Martijn Onderwater, Gerard Hoekstra, and Rob van der Mei. Throughput modeling of the IEEE MAC for sensor networks. ACM SIGMETRICS Performance Evaluation Review, 44(4):2–9, March 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cardellini:2017:OOR

[2678] Valeria Cardellini, Vincenzo Grassi, Francesco Lo Presti, and Matteo Nardelli. Optimal operator replication and placement for distributed stream processing systems. ACM SIGMET-RICS Performance Evaluation Review, 44(4):11–22, March 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gianniti:2017:FPN

[2679] Eugenio Gianniti, Alessandro Maria Rizzi, Enrico Barbierato, Marco Gribaudo, and Danilo Ardagna. Fluid Petri nets for the performance evaluation of MapReduce and Spark applications. ACM SIGMETRICS Performance Evaluation Review, 44(4):23–36, March 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Longo:2017:ARQ

[2680] Francesco Longo, Rahul Ghosh, Vijay K. Naik, Andrew J. Rindos, and Kishor S. Trivedi. An approach for resiliency quantification of large scale systems. ACM SIGMETRICS Performance Evaluation Review, 44(4):37–48, March 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Canali:2017:ICP

[2681] Claudia Canali and Riccardo Lancellotti. Identifying communication patterns between virtual machines in software-defined data centers. ACM SIGMETRICS Performance Evaluation Review, 44(4):49–56, March 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bianchi:2017:MRB

[2682] Francesco Bianchi and Francesco Lo Presti. A Markov reward based resource-latency aware heuristic for the virtual network embedding problem. ACM SIGMETRICS Performance Evaluation Review, 44(4):57–68, March 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Donatiello:2017:ASL

[2683] Lorenzo Donatiello and Gustavo Marfia. Analyzing and shaping the lifetime and the performance of barrier coverage sensor networks. ACM SIGMETRICS Performance Evaluation Review, 44(4):69–79, March 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pinciroli:2017:CEM

[2684] Riccardo Pinciroli and Salvatore Distefano. Characterization and evaluation of Mobile Crowdsensing performance and energy indicators. ACM SIGMET-RICS Performance Evaluation Review, 44(4):80–90, March 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Totis:2017:OLK

[2685] Niccoló Totis, Laura Follia, Chiara Riganti, Francesco Novelli, Francesca Cordero, and Marco Beccuti. Overcoming the lack of kinetic information in biochemical reactions networks. ACM SIGMETRICS Performance Evaluation Review, 44(4):91–102, March 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Golubchik:2017:DSM

[2686] Leana Golubchik. Delay scalings and mean-field limits in networked systems. ACM SIGMETRICS Performance Evaluation Review, 45(1):1, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Avrachenkov:2017:LCA

[2687] Konstantin Avrachenkov, Jasper Goseling, and Berksan Serbetci. A lowcomplexity approach to distributed cooperative caching with geographic constraints. ACM SIGMETRICS Performance Evaluation Review, 45(1):2, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mukherjee:2017:OSE

[2688] Debankur Mukherjee, Souvik Dhara, Sem C. Borst, and Johan S. H. van Leeuwaarden. Optimal service elasticity in large-scale distributed systems. ACM SIGMETRICS Performance Evaluation Review, 45(1):3, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gong:2017:QPS

[2689] Long Gong, Paul Tune, Liang Liu, Sen Yang, and Jun (Jim) Xu. Queueproportional sampling: a better approach to crossbar scheduling for inputqueued switches. ACM SIGMETRICS Performance Evaluation Review, 45(1):

4, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ju:2017:HLS

[2690] Xiaoen Ju, Hani Jamjoom, and Kang G. Shin. Hieroglyph: Locallysufficient graph processing via computesync-merge. ACM SIGMETRICS Performance Evaluation Review, 45(1):5, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2017:SYE

[2691] Lingda Li, Robel Geda, Ari B. Hayes, Yanhao Chen, Pranav Chaudhari, Eddy Z. Zhang, and Mario Szegedy. A simple yet effective balanced edge partition model for parallel computing. ACM SIGMETRICS Performance Evaluation Review, 45(1):6, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cohen:2017:OCS

[2692] Maxime C. Cohen, Philipp Keller, Vahab Mirrokni, and Morteza Zadimoghadddam. Overcommitment in cloud services bin packing with chance constraints. ACM SIGMETRICS Performance Evaluation Review, 45(1):7, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Quach:2017:ILT

[2693] Alan Quach, Zhongjie Wang, and Zhiyun Qian. Investigation of the 2016 Linux TCP stack vulnerability at scale. ACM SIGMETRICS Performance Evaluation Review, 45(1):8, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2017:CMP

[2694] Brandon Wang, Xiaoye Li, Leandro P. de Aguiar, Daniel S. Menasche, and Zubair Shafiq. Characterizing and modeling patching practices of industrial control systems. ACM SIGMET-RICS Performance Evaluation Review, 45(1):9, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2017:SGN

[2695] Sinong Wang and Ness Shroff. Security game with non-additive utilities and multiple attacker resources. ACM SIGMETRICS Performance Evaluation Review, 45(1):10, June 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Braverman:2017:FMB

[2696] Anton Braverman, J. G. Dai, Xin Liu, and Lei Ying. Fluid-model-based car routing for modern ridesharing systems. ACM SIGMETRICS Performance Evaluation Review, 45(1):11–12, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kuhnle:2017:PSA

[2697] Alan Kuhnle, Tianyi Pan, Victoria G. Crawford, Md Abdul Alim, and My T. Thai. Pseudo-separation for assessment of structural vulnerability of a network. ACM SIGMETRICS Performance Evaluation Review, 45(1):13–14, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Deng:2017:CRA

[2698] Han Deng and I-Hong Hou. On the capacity requirement for arbitrary endto-end deadline and reliability guarantees in multi-hop networks. ACM SIGMETRICS Performance Evaluation Review, 45(1):15–16, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xu:2017:OSE

[2699] Maotong Xu, Sultan Alamro, Tian Lan, and Suresh Subramaniam. Optimizing speculative execution of deadline-sensitive jobs in cloud. ACM SIGMETRICS Performance Evaluation Review, 45(1):17–18, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Islam:2017:SCM

[2700] Mohammad A. Islam, Xiaoqi Ren, Shaolei Ren, and Adam Wierman. A spot capacity market to increase power infrastructure utilization in multitenant data centers. ACM SIGMET-RICS Performance Evaluation Review, 45(1):19–20, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2017:HAO

[2701] Lin Yang, Mohammad H. Hajiesmaili, Hanling Yi, and Minghua Chen. Hourahead offering strategies in electricity market for power producers with storage and intermittent supply. ACM SIGMETRICS Performance Evaluation Review, 45(1):21–22, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gao:2017:WSL

[2702] Xing Gao, Zhang Xu, Haining Wang, Li Li, and Xiaorui Wang. Why “Some” like it hot too: Thermal attack on data centers. ACM SIGMETRICS Performance Evaluation Review, 45(1):23–24, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Comden:2017:IRD

[2703] Joshua Comden, Zhenhua Liu, and Yue Zhao. Incentivizing reliable demand response with customers’ uncertainties and capacity planning. ACM SIGMET-RICS Performance Evaluation Review, 45(1):25–26, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jadidi:2017:SPP

[2704] Amin Jadidi, Mohammad Arjomand, Mahmut Kandemir, and Chita Das. A study on performance and power efficiency of dense non-volatile caches in multi-core systems. ACM SIGMET-RICS Performance Evaluation Review, 45(1):27–28, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shafiee:2017:SCD

[2705] Mehrnoosh Shafiee and Javad Ghaderi. Scheduling coflows in datacenter networks: Improved bound for total weighted completion time. ACM SIGMETRICS Performance Evaluation Review, 45(1):29–30, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xiong:2017:CFG

[2706] Qin Xiong, Fei Wu, Zhonghai Lu, Yue Zhu, You Zhou, Yibing Chu, Changsheng Xie, and Ping Huang. Characterizing 3D floating gate NAND flash. ACM SIGMETRICS Performance Evaluation Review, 45(1):31–32, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lim:2017:EMP

[2707] Yeon sup Lim, Erich M. Nahum, Don Towsley, and Richard J. Gibbens. ECF: an MPTCP path scheduler to manage heterogeneous paths. ACM SIGMET-RICS Performance Evaluation Review, 45(1):33–34, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aktas:2017:SQH

[2708] Mehmet Fatih Aktas, Elie Najm, and Emina Soljanin. Simplex queues for hot-data download. ACM SIGMET-RICS Performance Evaluation Review, 45(1):35–36, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Singh:2017:EAF

[2709] Siddharth Singh, Vedant Nanda, Rijurekha Sen, Sohaib Ahmad, Satadal Sengupta, Amreesh Phokeer, Zaid Ahmed Farooq, Taslim Arefin Khan, Ponnurangam Kumaragaguru, Ihsan Ayyub Qazi, David Choffnes, and Krishna P. Gummadi. An empirical analysis of Facebook’s free basics. ACM SIGMETRICS Performance Evaluation Review, 45(1):37–38, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mena:2017:MTV

[2710] Jorge Mena, Peter Bankole, and Mario Gerla. Multipath TCP on a VANET: a performance study. ACM SIGMET-RICS Performance Evaluation Review, 45(1):39–40, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yu:2017:FSD

[2711] Ye Yu, Djamal Belazzougui, Chen Qian, and Qin Zhang. A fast, small, and dynamic forwarding information base. ACM SIGMETRICS Performance Evaluation Review, 45(1):41–42, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wu:2017:HHF

[2712] Ning Wu, Yingjie Bi, Nithin Michael, Ao Tang, John Doyle, and Nikolai Matni. HFTraC: High-frequency traffic control. ACM SIGMETRICS Performance Evaluation Review, 45(1):43–44, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Basu:2017:ATB

[2713] Soumya Basu, Aditya Sundarrajan, Javad Ghaderi, Sanjay Shakkottai, and Ramesh Sitaraman. Adaptive TTL-based caching for content delivery. ACM SIGMETRICS Performance Evaluation Review, 45(1):45–46, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mirrokni:2017:OOM

[2714] Vahab Mirrokni. Online optimization for markets and the cloud: Theory and practice. ACM SIGMETRICS Performance Evaluation Review, 45(1):47–48, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ying:2017:SMM

[2715] Lei Ying. Stein’s method for mean-field approximations in light and heavy traffic regimes. ACM SIGMETRICS Performance Evaluation Review, 45(1):49, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gast:2017:EVE

[2716] Nicolas Gast. Expected values estimated via mean-field approximation are 1/N-accurate: Extended abstract. ACM SIGMETRICS Performance Evaluation Review, 45(1):50, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sun:2017:ASM

[2717] Wen Sun, Véronique Simon, Sébastien Monnet, Philippe Robert, and Pierre Sens. Analysis of a stochastic model of replication in large distributed storage systems: a mean-field approach. ACM SIGMETRICS Performance Evaluation Review, 45(1):51, June 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chang:2017:URV

[2718] Kevin K. Chang, Abdullah Giray Yaglikçi, Saugata Ghose, Aditya Agrawal, Niladrish Chatterjee, Abhijith Kashyap, Donghyuk Lee, Mike O’Connor, Hasan Hassan, and Onur Mutlu. Understanding reduced-voltage operation in modern DRAM devices: Experimental characterization, analysis, and mechanisms. ACM SIGMET-RICS Performance Evaluation Review, 45(1):52, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Choi:2017:EDL

[2719] Wonil Choi, Mohammad Arjomand, Myoungsoo Jung, and Mahmut T. Kandemir. Exploiting data longevity for enhancing the lifetime of flashbased storage class memory. ACM SIGMETRICS Performance Evaluation Review, 45(1):53, June 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lee:2017:DIL

[2720] Donghyuk Lee, Samira Khan, Lavanya Subramanian, Saugata Ghose, Rachata Ausavarungnirun, Gennady Pekhimenko, Vivek Seshadri, and Onur Mutlu. Design-induced latency variation in modern DRAM chips: Characterization, analysis, and latency reduction mechanisms. ACM SIGMETRICS Performance Evaluation Review, 45(1): 54, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gibbens:2017:HND

[2721] Mathias Gibbens, Chris Gniady, Lei Ye, and Beichuan Zhang. Hadoop on named data networking: Experience and results. ACM SIGMETRICS Performance Evaluation Review, 45(1):55, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2017:UBI

[2722] Cheng Wang, Bhuvan Urgaonkar, Neda Nasiriani, and George Kesidis. Using burstable instances in the public cloud: Why, when and how? ACM SIGMETRICS Performance Evaluation Review, 45(1):56, June 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Venkatakrishnan:2017:DRB

[2723] Shaileshh Bojja Venkatakrishnan, Giulia Fanti, and Pramod Viswanath. Dandelion: Redesigning the Bitcoin network for anonymity. ACM SIGMET-RICS Performance Evaluation Review, 45(1):57, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jordan:2017:GBO

[2724] Michael Jordan. On gradient-based optimization: Accelerated, distributed, asynchronous and stochastic. ACM SIGMETRICS Performance Evaluation Review, 45(1):58, June 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sharma:2017:PDR

[2725] Prateek Sharma, David Irwin, and Prashant Shenoy. Portfolio-driven resource management for transient cloud servers. ACM SIGMETRICS Performance Evaluation Review, 45(1):59, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2017:OPP

[2726] Zijun Zhang, Zongpeng Li, and Chuan Wu. Optimal posted prices for online cloud resource allocation. ACM SIGMETRICS Performance Evaluation Review, 45(1):60, June 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2017:OTS

[2727] Xin Wang, Richard T. B. Ma, and Yinlong Xu. On optimal two-sided pricing of congested networks. ACM SIGMETRICS Performance Evaluation Review, 45(1):61, June 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Oh:2017:MFF

[2728] Sewoong Oh. Matrix factorization at the frontier of non-convex optimizations: Abstract for SIGMETRICS 2017 Rising Star Award talk. ACM SIGMETRICS Performance Evaluation Review, 45(1):62, June 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nguyen:2017:OIC

[2729] Hung T. Nguyen, Tri P. Nguyen, Tam N. Vu, and Thang N. Dinh. Outward influence and cascade size estimation in billion-scale networks. ACM SIGMETRICS Performance Evaluation Review, 45(1):63, June 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2017:API

[2730] Giuliano Casale. Accelerating performance inference over closed systems by asymptotic methods. ACM SIGMETRICS Performance Evaluation Review, 45(1):64, June 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bondorf:2017:QCD

[2731] Steffen Bondorf, Paul Nikolaus, and Jens B. Schmitt. Quality and cost of deterministic network calculus: Design and evaluation of an accurate and fast analysis. ACM SIGMETRICS Performance Evaluation Review, 45(1):65, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Formby:2017:CSP

[2732] David Formby, Anwar Walid, and Raheem Beyah. A case study in power substation network dynamics. ACM SIGMETRICS Performance Evaluation Review, 45(1):66, June 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhou:2017:PSM

[2733] You Zhou, Yian Zhou, Min Chen, and Shigang Chen. Persistent spread measurement for big network data based on register intersection. ACM SIGMETRICS Performance Evaluation Review, 45(1):67, June 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cao:2017:DEC

[2734] Yi Cao, Javad Nejati, Muhammad Wajahat, Aruna Balasubramanian, and Anshul Gandhi. Deconstructing the energy consumption of the mobile page load. ACM SIGMETRICS Performance Evaluation Review, 45(1):68, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ma:2017:RMP

[2735] Richard T. B. Ma and Vishal Misra.

Routing money, not packets: a tutorial on Internet economics. ACM SIGMET-RICS Performance Evaluation Review, 45(1):69–70, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2017:ORC

[2736] Mark S. Squillante. On the optimality of reflection control, with productioninventory applications. ACM SIG-METRICS Performance Evaluation Review, 45(2):3–5, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Feinberg:2017:SPA

[2737] Eugene A. Feinberg and Jefferson Huang. Strongly polynomial algorithms for transient and average-cost MDPs. ACM SIGMETRICS Performance Evaluation Review, 45(2):6–8, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Joshi:2017:BSC

[2738] Gauri Joshi. Boosting service capacity via adaptive task replication. ACM SIGMETRICS Performance Evaluation Review, 45(2):9–11, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aktas:2017:ESM

[2739] Mehmet Fatih Aktas, Pei Peng, and Emina Soljanin. Effective straggler mitigation: Which clones should attack and when? ACM SIGMETRICS Performance Evaluation Review, 45(2):12– 14, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lin:2017:NCC

[2740] Weixuan Lin, John Z. F. Pang, Eilyan Bitar, and Adam Wierman. Networked Cournot competition in platform markets: Access control and efficiency loss. ACM SIGMETRICS Performance Evaluation Review, 45(2):15– 17, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Maxey:2017:WAB

[2741] Tyler Maxey, Hakjin Chung, Hyun-Soo Ahn, and Rhonda Righter. When is anarchy beneficial? ACM SIGMETRICS Performance Evaluation Review, 45(2): 18–20, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Panigrahy:2017:HRV

[2742] Nitish K. Panigrahy, Jian Li, and Don Towsley. Hit rate vs. hit probability based cache utility maximization. ACM SIGMETRICS Performance Evaluation Review, 45(2):21–23, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jiang:2017:LCU

[2743] Bo Jiang, Philippe Nain, and Don Towsley. LRU cache under stationary requests. ACM SIGMETRICS Performance Evaluation Review, 45(2):24–26, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goel:2017:TFS

[2744] Gautam Goel, Niangjun Chen, and Adam Wierman. Thinking fast and slow: Optimization decomposition across timescales. ACM SIGMETRICS Performance Evaluation Review, 45(2): 27–29, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

London:2017:DOL

[2745] Palma London, Niangjun Chen, Shai Vardi, and Adam Wierman. Distributed optimization via local computation algorithms. ACM SIGMET-RICS Performance Evaluation Review, 45(2):30–32, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aveklouris:2017:EVC

[2746] Angelos Aveklouris, Yorie Nakahira, Maria Vlasiou, and Bert Zwart. Electric vehicle charging: a queueing approach. ACM SIGMETRICS Performance Evaluation Review, 45(2):33–35, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Scully:2017:OSJ

[2747] Ziv Scully, Guy Blelloch, Mor Harchol-Balter, and Alan Scheller-Wolf. Optimally scheduling jobs with multiple tasks. ACM SIGMETRICS Performance Evaluation Review, 45(2):36–38, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baryshnikov:2017:LDIa

[2748] Yuliy Baryshnikov and Abram Magner. Large deviations for increasing subsequences of permutations and a concurrency application. ACM SIGMET-RICS Performance Evaluation Review, 45(2):39–41, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ganguly:2017:LSN

[2749] Arnab Ganguly, Kavita Ramanan, Philippe Robert, and Wen Sun. A large-scale network with moving servers. ACM SIGMETRICS Performance Evaluation Review, 45(2):42–44, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Foss:2017:JIQ

[2750] Sergey Foss and Alexander L. Stolyar. Join-idle-queue system with general service times: Large-scale limit of stationary distributions. ACM SIGMET-RICS Performance Evaluation Review, 45(2):45–47, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2017:NCD

[2751] Yingdong Lu, Mark S. Squillante, and Chai Wah Wu. Nearly completely decomposable epidemic-like stochastic processes with time-varying behavior. ACM SIGMETRICS Performance Evaluation Review, 45(2):48–50, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Comden:2017:DAD

[2752] Joshua Comden, Zhenhua Liu, and Yue Zhao. Distributed algorithm design for probabilistic demand response. ACM SIGMETRICS Performance Evaluation Review, 45(2):52–54, 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hajiesmaili:2017:SRR

[2753] Mohammad H. Hajiesmaili, Minghua Chen, Enrique Mallada, and Chi-Kin Chau. Summary of recent results: Crowd-sourced storage-assisted demand response in microgrids. ACM SIGMETRICS Performance Evaluation Review, 45(2):55–57, 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Le:2017:OEPa

[2754] Tan N. Le, Jie Liang, Zhenhua Liu, Ramesh K. Sitaraman, Jayakrishnan Nair, and Bong Jun Choi. Optimal energy procurement for geo-distributed data centers in multi-timescale electricity markets. ACM SIGMETRICS Performance Evaluation Review, 45(2):58– 63, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2017:PBP

[2755] Wei Wang and Nanpeng Yu. Phase balancing in power distribution network with data center. ACM SIGMETRICS Performance Evaluation Review, 45(2): 64–69, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Oi:2017:CSE

[2756] Hitoshi Oi. A case study of energy efficiency on a heterogeneous multiprocessor. ACM SIGMETRICS Performance Evaluation Review, 45(2):70– 72, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2017:GPE

[2757] Qiang Wang and Xiaowen Chu. GPGPU power estimation with core and memory frequency scaling. ACM SIGMETRICS Performance Evaluation Review, 45(2):73–78, 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2017:BBS

[2758] Dong Chen and David Irwin. Black-box solar performance modeling: Comparing physical, machine learning, and hybrid approaches. ACM SIGMETRICS Performance Evaluation Review, 45(2): 79–84, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

You:2017:BSA

[2759] Pengcheng You, Youxian Sun, John Pang, Steven Low, and Minghua Chen. Battery swapping assignment for electric vehicles: a bipartite matching approach. ACM SIGMETRICS Performance Evaluation Review, 45(2):85–87, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Doan:2017:DLM

[2760] Thinh T. Doan, Subhonmesh Bose, and Carolyn L. Beck. Distributed Lagrangian method for tie-line scheduling in power grids under uncertainty. ACM SIGMETRICS Performance Evaluation Review, 45(2):88–90, 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Islam:2017:FLP

[2761] Mohammad A. Islam, Shaolei Ren, and Adam Wierman. A first look at power attacks in multi-tenant data centers. ACM SIGMETRICS Performance Evaluation Review, 45(2):91–93, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pang:2017:LSF

[2762] John Z. F. Pang, Linqi Guo, and Steven H. Low. Load-side frequency regulation with limited control coverage. ACM SIGMETRICS Performance Evaluation Review, 45(2):94–96, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kelic:2017:ICI

[2763] Andjelka Kelic. Interdependencies in critical infrastructure modeling. ACM SIGMETRICS Performance Evaluation Review, 45(2):99–102, 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Guo:2017:MPS

[2764] Linqi Guo, Chen Liang, and Steven H.

Low. Monotonicity properties and spectral characterization of power redistribution in cascading failures. ACM SIGMETRICS Performance Evaluation Review, 45(2):103–106, 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Oostenbrink:2017:CID

[2765] Jorik Oostenbrink and Fernando Kuipers. Computing the impact of disasters on networks. ACM SIGMET-RICS Performance Evaluation Review, 45(2):107–110, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Soltan:2017:APG

[2766] Saleh Soltan and Gil Zussman. Algorithms for power grid state estimation after cyber-physical attacks. ACM SIGMETRICS Performance Evaluation Review, 45(2):111–114, 2017. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bienstock:2017:CUA

[2767] Daniel Bienstock and Mauro Escobar. Computing undetectable attacks on power grids. ACM SIGMETRICS Performance Evaluation Review, 45(2): 115–118, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stergiopoulos:2017:IAJ

[2768] George Stergiopoulos, Evangelos Valvis, Foivos Anagnou-Misyris, Nick Bozovic, and Dimitris Gritzalis. Interdependency analysis of junctions for congestion mitigation in transportation infrastructures. ACM SIGMETRICS Performance Evaluation Review, 45(2): 119–124, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2017:DGA

[2769] Juntao Chen, Corinne Touati, and Quanyan Zhu. A dynamic game analysis and design of infrastructure network protection and recovery: 125. ACM SIGMETRICS Performance Evaluation Review, 45(2):128, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ding:2017:CBT

[2770] Jianguo Ding, Yacine Atif, Sten F.

Andler, Birgitta Lindström, and Manfred Jeusfeld. CPS-based threat modeling for critical infrastructure protection. ACM SIGMETRICS Performance Evaluation Review, 45(2):129– 132, 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhou:2017:WIC

[2771] Xunyu Zhou. Who are I: Intrapersonal conflicts in performance measure and control. ACM SIGMETRICS Performance Evaluation Review, 45(3):1, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yekkehkhany:2017:GPT

[2772] Ali Yekkehkhany, Avesta Hojjati, and Mohammad H. Hajiesmaili. GB-PANDAS: Throughput and heavytraffic optimality analysis for affinity scheduling. ACM SIGMETRICS Performance Evaluation Review, 45 (3):2–14, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goldsztajn:2017:CNA

[2773] Diego Goldsztajn, Andres Ferragut, Fernando Paganini, and Matthieu Jonckheere. Controlling the number of active instances in a cloud environment. ACM SIGMETRICS Performance Evaluation Review, 45(3):15–20, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Joshi:2017:SRB

[2774] Gauri Joshi. Synergy via redundancy: Boosting service capacity with adaptive replication. ACM SIGMETRICS Performance Evaluation Review, 45 (3):21–28, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

He:2017:DLA

[2775] Ting He. Distributed link anomaly detection via partial network tomography. ACM SIGMETRICS Performance Evaluation Review, 45(3):29–42, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tootaghaj:2017:PTO

[2776] Diman Zad Tootaghaj, Ting He, and Thomas La Porta. Parsimonious tomography: Optimizing costidentifiability trade-off for probingbased network monitoring. ACM SIGMETRICS Performance Evaluation Review, 45(3):43–55, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jansen:2017:PEW

[2777] Bart Jansen, Timothy Goodwin, Varun Gupta, Fernando Kuipers, and Gil Zussman. Performance evaluation of WebRTC-based video conferencing. ACM SIGMETRICS Performance Evaluation Review, 45(3):56–68, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Moka:2017:APS

[2778] S. B. Moka, S. Juneja, and M. R. H.

Mandjes. Analysis of perfect sampling methods for hard-sphere models. ACM SIGMETRICS Performance Evaluation Review, 45(3):69–75, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hollocou:2017:MLC

[2779] Alexandre Hollocou, Thomas Bonald, and Marc Lelarge. Multiple local community detection. ACM SIGMET-RICS Performance Evaluation Review, 45(3):76–83, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Baryshnikov:2017:LDIb

[2780] Yuli Baryshnikov and Abram Magner.

Large deviations for increasing subsequences of permutations and a concurrency application. ACM SIGMET-RICS Performance Evaluation Review, 45(3):84–89, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bhatt:2017:IIF

[2781] Sujay Bhatt and Vikram Krishnamurthy. Incentivized information fusion with social sensors. ACM SIGMETRICS Performance Evaluation Review, 45(3):90–95, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Abbe:2017:LGD

[2782] Emmanuel Abbe. Learning from graphical data. ACM SIGMETRICS Performance Evaluation Review, 45(3):96, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sun:2017:SPW

[2783] Fengyou Sun and Yuming Jiang. A statistical property of wireless channel capacity: Theory and application. ACM SIGMETRICS Performance Evaluation Review, 45(3):97–108, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cecchi:2017:MFL

[2784] F. Cecchi, P. M. Van de Ven, and S. Shneer. Mean-field limits for multihop random-access networks. ACM SIGMETRICS Performance Evaluation Review, 45(3):109–122, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cecchi:2017:SMF

[2785] F. Cecchi, S. C. Borst, J. S. H. van Leeuwaarden, and P. A. Whiting. Spatial mean-field limits for ultradense random-access networks. ACM SIGMETRICS Performance Evaluation Review, 45(3):123–136, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Juneja:2017:CDU

[2786] S. Juneja and N. Shimkin. On the computation of dynamic user equilibrium in the multiclass transient fluid queue. ACM SIGMETRICS Performance Evaluation Review, 45(3): 137–142, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Telek:2017:RTD

[2787] Miklos Telek and Benny Van Houdt.

Response time distribution of a class of limited processor sharing queues. ACM SIGMETRICS Performance Evaluation Review, 45(3):143–155, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tay:2017:TES

[2788] Y. C. Tay. A technique to estimate a system’s asymptotic delay and throughput. ACM SIGMETRICS Performance Evaluation Review, 45(3): 156–159, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2017:ELS

[2789] Yingdong Lu, Mark S. Squillante, and Chai Wah Wu. Epidemic-like stochastic processes with time-varying behavior: Structural properties and asymptotic limits. ACM SIGMETRICS Performance Evaluation Review, 45(3): 160–166, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Allybokus:2017:LBF

[2790] Zaid Allybokus, Konstantin Avrachenkov, Jérémie Leguay, and Lorenzo Maggi. Lower bounds for the fair resource allocation problem. ACM SIGMETRICS Performance Evaluation Review, 45(3): 167–173, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2017:ODU

[2791] Ruidi Chen and Ioannis Paschalidis.

Outlier detection using robust optimization with uncertainty sets constructed from risk measures. ACM SIGMETRICS Performance Evaluation Review, 45(3):174–179, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2017:ORC

[2792] Jiankui Yang, David D. Yao, and Heng-Qing Ye. On the optimality of reflection control, with production-inventory applications. ACM SIGMETRICS Performance Evaluation Review, 45(3): 180–183, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Greenberg:2017:AN

[2793] Albert Greenberg. Azure networking. ACM SIGMETRICS Performance Evaluation Review, 45(3):184, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Le:2017:OEPb

[2794] Tan N. Le, Jie Liang, Zhenhua Liu, Ramesh K. Sitaraman, Jayakrishnan Nair, and Bong Jun Choi. Optimal energy procurement for geo-distributed data centers in multi-timescale electricity markets. ACM SIGMETRICS Performance Evaluation Review, 45(3): 185–197, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Cetinay:2017:ACF

[2795] Hale Cetinay, Saleh Soltan, Fernando A. Kuipers, Gil Zussman, and Piet Van Mieghem. Analyzing cascading failures in power grids under the AC and DC power flow models. ACM SIGMETRICS Performance Evaluation Review, 45(3):198–203, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Deiana:2017:FFM

[2796] Eleonora Deiana, Guy Latouche, and Marie-Ange Remiche. Fluid flow model for energy-aware server performance evaluation. ACM SIGMETRICS Performance Evaluation Review, 45(3): 204–209, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mitra:2017:MSI

[2797] Debasis Mitra and Qiong Wang. Management strategies for industrial laboratories with knowledge memory. ACM SIGMETRICS Performance Evaluation Review, 45(3):210–216, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lu:2017:OEP

[2798] Y. Lu, S. T. Maguluri, M. S. Squillante, T. Suk, and X. Wu. Optimal energy procurement for geo-distributed data centers in multi-timescale electricity markets. ACM SIGMETRICS Performance Evaluation Review, 45(3): 217–223, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aktas:2017:SMD

[2799] Mehmet Fatih Aktas, Pei Peng, and Emina Soljanin. Straggler mitigation by delayed relaunch of tasks. ACM SIGMETRICS Performance Evaluation Review, 45(3):224–231, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2017:HTD

[2800] Weina Wang, Siva Theja Maguluri, R. Srikant, and Lei Ying. Heavy-traffic delay insensitivity in connection-level models of data transfer with proportionally fair bandwidth sharing. ACM SIGMETRICS Performance Evaluation Review, 45(3):232–245, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2017:PEJ

[2801] Giuliano Casale, Giuseppe Serazzi, and Lulai Zhu. Performance evaluation with Java modelling tools: a handson introduction. ACM SIGMETRICS Performance Evaluation Review, 45(3): 246–247, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nannicini:2017:SMD

[2802] Giacomo Nannicini. Straggler mitigation by delayed relaunch of tasks. ACM SIGMETRICS Performance Evaluation Review, 45(3):248, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Maguluri:2017:DMH

[2803] Siva Theja Maguluri, R. Srikant, and Weina Wang. The drift method for heavy traffic limits, with applications in data centers and networks. ACM SIGMETRICS Performance Evaluation Review, 45(3):249, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Braverman:2017:SMS

[2804] Anton Braverman and Jim Dai. Stein’s method for steady-state approximations: Error bounds and engineering solutions. ACM SIGMETRICS Performance Evaluation Review, 45 (3):250, December 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Avrachenkov:2018:EBM

[2805] Konstantin Avrachenkov and Tejas Bodas. On the equivalence between multiclass processor sharing and random order scheduling policies. ACM SIGMETRICS Performance Evaluation Review, 45(4):2–6, March 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pal:2018:ICS

[2806] Ranjan Pal, Leana Golubchik, Konstantinos Psounis, and Pan Hui. Improving cyber-security via profitable insurance markets. ACM SIGMET-RICS Performance Evaluation Review, 45(4):7–15, March 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Misra:2018:SDP

[2807] Vishal Misra. Session details: Performance evaluation review. ACM SIGMETRICS Performance Evaluation Review, 45(4), March 2018. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dai:2018:SSA

[2808] Jim Dai. Steady-state approximations:

Achievement lecture. ACM SIGMET-RICS Performance Evaluation Review, 46(1):1, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Banerjee:2018:SDC

[2809] Siddhartha Banerjee, Yash Kanoria, and Pengyu Qian. State dependent control of closed queueing networks. ACM SIGMETRICS Performance Evaluation Review, 46(1):2–4, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fanti:2018:DLC

[2810] Giulia Fanti, Shaileshh Bojja Venkatakrishnan, Surya Bakshi, Bradley Denby, Shruti Bhargava, Andrew Miller, and Pramod Viswanath. Dandelion++: Lightweight cryptocurrency networking with formal anonymity guarantees. ACM SIGMETRICS Performance Evaluation Review, 46(1):5–7, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Buchnik:2018:BGD

[2811] Eliav Buchnik and Edith Cohen. Bootstrapped graph diffusions: Exposing the power of nonlinearity. ACM SIGMETRICS Performance Evaluation Review, 46(1):8–10, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hoffmann:2018:CUC

[2812] Jessica Hoffmann and Constantine Caramanis. The cost of uncertainty in curing epidemics. ACM SIGMET-RICS Performance Evaluation Review, 46(1):11–13, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sejourne:2018:PFM

[2813] Thibault Séjourné, Samitha Samaranayake, and Siddhartha Banerjee.

The price of fragmentation in mobilityon-demand services. ACM SIGMET-RICS Performance Evaluation Review, 46(1):14–16, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Amjad:2018:CDE

[2814] Muhammad J. Amjad and Devavrat Shah. Censored demand estimation in retail. ACM SIGMETRICS Performance Evaluation Review, 46(1):17–19, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Martonosi:2018:NMM

[2815] Margaret Martonosi. New metrics and models for a post-ISA era: Managing complexity and scaling performance in heterogeneous parallelism and Internet-of-Things. ACM SIG-METRICS Performance Evaluation Review, 46(1):20, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Borst:2018:DSM

[2816] Sem Borst and Martin Zubeldia. Delay scaling in many-sources wireless networks without queue state information. ACM SIGMETRICS Performance Evaluation Review, 46(1):21–23, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berger:2018:PBO

[2817] Daniel S. Berger, Nathan Beckmann, and Mor Harchol-Balter. Practical bounds on optimal caching with variable object sizes. ACM SIGMET-RICS Performance Evaluation Review, 46(1):24–26, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2018:RPS

[2818] Jian Tan, Guocong Quan, Kaiyi Ji, and Ness Shroff. On resource pooling and separation for LRU caching. ACM SIGMETRICS Performance Evaluation Review, 46(1):27, June 2018. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2018:ORO

[2819] Lin Yang, Wing Shing Wong, and Mohammad H. Hajiesmaili. An optimal randomized online algorithm for QoS buffer management. ACM SIGMET-RICS Performance Evaluation Review, 46(1):28–30, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liang:2018:MQL

[2820] Qingkai Liang and Eytan Modiano. Minimizing queue length regret under adversarial network models. ACM SIGMETRICS Performance Evaluation Review, 46(1):31–32, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Freeman:2018:DPS

[2821] Rupert Freeman, Seyed Majid Zahedi, Vincent Conitzer, and Benjamin C. Lee. Dynamic proportional sharing: a game-theoretic approach. ACM SIGMETRICS Performance Evaluation Review, 46(1):33–35, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Scully:2018:SOC

[2822] Ziv Scully, Mor Harchol-Balter, and Alan Scheller-Wolf. SOAP: One clean analysis of all age-based scheduling policies. ACM SIGMETRICS Performance Evaluation Review, 46(1):36–38, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anand:2018:WIB

[2823] Arjun Anand and Gustavo de Veciana. A Whittle’s index based approach for QoE optimization in wireless networks. ACM SIGMETRICS Performance Evaluation Review, 46(1):39, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kleinberg:2018:ITO

[2824] Jon Kleinberg. Inherent trade-offs in algorithmic fairness. ACM SIGMET-RICS Performance Evaluation Review, 46(1):40, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2018:OAO

[2825] Lin Yang, Lei Deng, Mohammad H.

Hajiesmaili, Cheng Tan, and Wing Shing Wong. An optimal algorithm for online non-convex learning. ACM SIGMET-RICS Performance Evaluation Review, 46(1):41–43, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Duran:2018:AOC

[2826] Santiago Duran and Ina Maria Verloop.

Asymptotic optimal control of Markovmodulated restless bandits. ACM SIGMETRICS Performance Evaluation Review, 46(1):44–46, June 2018.

CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Magureanu:2018:OLO

[2827] Stefan Magureanu, Alexandre Proutiere, Marcus Isaksson, and Boxun Zhang. Online learning of optimally diverse rankings. ACM SIGMETRICS Performance Evaluation Review, 46(1):47–49, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Talebi:2018:LPF

[2828] Mohammad Sadegh Talebi and Alexandre Proutiere. Learning proportionally fair allocations with low regret. ACM SIGMETRICS Performance Evaluation Review, 46(1):50–52, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yun:2018:MAB

[2829] Donggyu Yun, Sumyeong Ahn, Alexandre Proutiere, Jinwoo Shin, and Yung Yi. Multi-armed bandit with additional observations. ACM SIGMET-RICS Performance Evaluation Review, 46(1):53–55, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wei:2018:OLW

[2830] Xiaohan Wei, Hao Yu, and Michael J. Neely. Online learning in weakly coupled Markov decision processes: a convergence time study. ACM SIGMET-RICS Performance Evaluation Review, 46(1):56–58, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zheng:2018:HCL

[2831] Pengfei Zheng and Benjamin C. Lee.

Hound: Causal learning for datacenterscale straggler diagnosis. ACM SIG-METRICS Performance Evaluation Review, 46(1):59–61, June 2018. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nitu:2018:WSS

[2832] Vlad Nitu, Aram Kocharyan, Hannas Yaya, Alain Tchana, Daniel Hagimont, and Hrachya Astsatryan. Working set size estimation techniques in virtualized environments: One size does not fit all. ACM SIGMETRICS Performance Evaluation Review, 46(1):62–63, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2018:PSF

[2833] Shenglin Zhang, Ying Liu, Weibin Meng, Zhiling Luo, Jiahao Bu, Sen Yang, Peixian Liang, Dan Pei, Jun Xu, Yuzhi Zhang, Yu Chen, Hui Dong, Xianping Qu, and Lei Song. PreFix: Switch failure prediction in datacenter networks. ACM SIGMETRICS Performance Evaluation Review, 46(1):64–66, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Psychas:2018:NPV

[2834] Konstantinos Psychas and Javad Ghaderi. On non-preemptive VM scheduling in the cloud. ACM SIG-METRICS Performance Evaluation Review, 46(1):67–69, June 2018. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Islam:2018:WSL

[2835] Mohammad A. Islam, Luting Yang, Kiran Ranganath, and Shaolei Ren. Why some like it loud: Timing power attacks in multi-tenant data centers using an acoustic side channel. ACM SIGMET-RICS Performance Evaluation Review, 46(1):70–72, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ahmadian:2018:ECH

[2836] Saba Ahmadian, Onur Mutlu, and Hossein Asadi. ECI-Cache: a high-endurance and cost-efficient I/O caching scheme for virtualized platforms. ACM SIGMETRICS Performance Evaluation Review, 46(1):73, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2018:SMV

[2837] Zhaowei Tan, Yuanjie Li, Qianru Li, Zhehui Zhang, Zhehan Li, and Songwu Lu. Supporting mobile VR in LTE networks: How close are we? ACM SIGMETRICS Performance Evaluation Review, 46(1):74, June 2018. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pignolet:2018:TNP

[2838] Yvonne-Anne Pignolet, Stefan Schmid, and Gilles Tredan. Tomographic node placement strategies and the impact of the routing model. ACM SIGMET-RICS Performance Evaluation Review, 46(1):75–77, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vlachou:2018:LTL

[2839] Christina Vlachou, Ioannis Pefkianakis, and Kyu-Han Kim. LTERadar: Towards LTE-aware Wi-Fi access points. ACM SIGMETRICS Performance Evaluation Review, 46(1):78–80, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kuhnle:2018:NRL

[2840] Alan Kuhnle, Victoria G. Crawford, and My T. Thai. Network resilience and the length-bounded multicut problem: Reaching the dynamic billionscale with guarantees. ACM SIGMET-RICS Performance Evaluation Review, 46(1):81–83, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2018:PIA

[2841] Sen Yang, He Yan, Zihui Ge, Dongmei Wang, and Jun Xu. Predictive impact analysis for designing a resilient cellular backhaul network. ACM SIGMET-RICS Performance Evaluation Review, 46(1):84–86, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Subramanian:2018:SFT

[2842] Kausik Subramanian, Loris D’Antoni, and Aditya Akella. Synthesis of faulttolerant distributed router configurations. ACM SIGMETRICS Performance Evaluation Review, 46(1):87–89, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xu:2018:RFM

[2843] Kuang Xu and Se-Young Yun. Reinforcement with fading memories. ACM SIGMETRICS Performance Evaluation Review, 46(1):90–92, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Doan:2018:CRD

[2844] Thinh T. Doan, Carolyn L. Beck, and R. Srikant. On the convergence rate of distributed gradient methods for finite-sum optimization under communication delays. ACM SIGMET-RICS Performance Evaluation Review, 46(1):93–95, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2018:DSM

[2845] Yudong Chen, Lili Su, and Jiaming Xu. Distributed statistical machine learning in adversarial settings: Byzantine gradient descent. ACM SIGMETRICS Performance Evaluation Review, 46(1): 96, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2018:NNM

[2846] Mowei Wang, Yong Cui, Shihan Xiao, Xin Wang, Dan Yang, Kai Chen, and Jun Zhu. Neural network meets DCN: Traffic-driven topology adaptation with deep learning. ACM SIGMETRICS Performance Evaluation Review, 46(1):97–99, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schardl:2018:CFC

[2847] Tao B. Schardl, Tyler Denniston, Damon Doucet, Bradley C. Kuszmaul, I-Ting Angelina Lee, and Charles E. Leiserson. The CSI framework for compiler-inserted program instrumentation. ACM SIGMETRICS Performance Evaluation Review, 46(1):100– 102, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jain:2018:QEC

[2848] Akshay Jain, Mahmoud Khairy, and Timothy G. Rogers. A quantitative evaluation of contemporary GPU simulation methodology. ACM SIGMET-RICS Performance Evaluation Review, 46(1):103–105, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Luo:2018:INF

[2849] Yixin Luo, Saugata Ghose, Yu Cai, Erich F. Haratsch, and Onur Mutlu. Improving 3D NAND flash memory lifetime by tolerating early retention loss and process variation. ACM SIGMETRICS Performance Evaluation Review, 46(1):106, June 2018. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chen:2018:FGE

[2850] Xiaomeng Chen, Jiayi Meng, Y. Charlie Hu, Maruti Gupta, Ralph Hasholzner, Venkatesan Nallampatti Ekambaram, Ashish Singh, and Srikathyayani Srikanteswara. A fine-grained eventbased modem power model for enabling in-depth modem energy drain analysis. ACM SIGMETRICS Performance Evaluation Review, 46(1):107– 109, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghose:2018:WYD

[2851] Saugata Ghose, Abdullah Giray Yaglikçi, Raghav Gupta, Donghyuk Lee, Kais Kudrolli, William X. Liu, Hasan Hassan, Kevin K. Chang, Niladrish Chatterjee, Aditya Agrawal, Mike O’Connor, and Onur Mutlu. What your DRAM power models are not telling you: Lessons from a detailed experimental study. ACM SIG-METRICS Performance Evaluation Review, 46(1):110, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Oleksenko:2018:IME

[2852] Oleksii Oleksenko, Dmitrii Kuvaiskii, Pramod Bhatotia, Pascal Felber, and Christof Fetzer. Intel MPX explained: a cross-layer analysis of the Intel MPX system stack. ACM SIGMETRICS Performance Evaluation Review, 46 (1):111–112, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gast:2018:RMFa

[2853] Nicolas Gast and Benny Van Houdt. A refined mean field approximation. ACM SIGMETRICS Performance Evaluation Review, 46(1):113, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hellemans:2018:PDC

[2854] Tim Hellemans and Benny Van Houdt. On the power-of-d-choices with least loaded server selection. ACM SIG-METRICS Performance Evaluation Review, 46(1):114, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhou:2018:DQI

[2855] Xingyu Zhou, Fei Wu, Jian Tan, Kannan Srinivasan, and Ness Shroff. Degree of queue imbalance: Overcoming the limitation of heavy-traffic delay optimality in load balancing systems. ACM SIGMETRICS Performance Evaluation Review, 46(1):115, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berg:2018:TOP

[2856] Benjamin Berg, Jan-Pieter Dorsman, and Mor Harchol-Balter. Towards optimality in parallel job scheduling. ACM SIGMETRICS Performance Evaluation Review, 46(1):116–118, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jiang:2018:CSM

[2857] Bo Jiang, Philippe Nain, Don Towsley, and Saikat Guha. On a class of stochastic multilayer networks. ACM SIGMETRICS Performance Evaluation Review, 46(1):119–121, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zeng:2018:FJQ

[2858] Yun Zeng, Jian Tan, and Cathy Honghui Xia. Fork and join queueing networks with heavy tails: Scaling dimension and throughput limit. ACM SIGMET-RICS Performance Evaluation Review, 46(1):122–124, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bonald:2018:PBF

[2859] Thomas Bonald, Céline Comte, and Fabien Mathieu. Performance of balanced fairness in resource pools: a recursive approach. ACM SIGMET-RICS Performance Evaluation Review, 46(1):125–127, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhou:2018:DLC

[2860] Xingyu Zhou, Fei Wu, Jian Tan, Yin Sun, and Ness Shroff. Designing lowcomplexity heavy-traffic delay-optimal load balancing schemes: Theory to algorithms. ACM SIGMETRICS Performance Evaluation Review, 46(1):128, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2018:TFC

[2861] Sinong Wang and Ness Shroff. Towards fast-convergence, low-delay and low-complexity network optimization. ACM SIGMETRICS Performance Evaluation Review, 46(1):129– 131, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aghajani:2018:PMA

[2862] Reza Aghajani, Xingjie Li, and Kavita Ramanan. The PDE method for the analysis of randomized load balancing networks. ACM SIGMET-RICS Performance Evaluation Review, 46(1):132–134, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2018:SRL

[2863] Sen Yang, Bill Lin, and Jun Xu. Safe randomized load-balanced switching by diffusing extra loads. ACM SIGMET-RICS Performance Evaluation Review, 46(1):135–137, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mukherjee:2018:AOL

[2864] Debankur Mukherjee, Sem Borst, and Johan S. H. van Leeuwaarden. Asymptotically optimal load balancing topologies. ACM SIGMETRICS Performance Evaluation Review, 46(1):138, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hegde:2018:ASP

[2865] Nidhi Hegde. ACM Sigmetrics Performance Evaluation Review: a new series on diversity. ACM SIGMETRICS Performance Evaluation Review, 46(1): 139, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Golubchik:2018:DFR

[2866] Leana Golubchik and Mallory Redel. Diversity in faculty recruiting: a WiSE approach. ACM SIGMET-RICS Performance Evaluation Review, 46(1):140–142, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fanti:2018:SDL

[2867] Giulia Fanti. Session details: Load balancing. ACM SIGMETRICS Performance Evaluation Review, 46(1), June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gast:2018:SDR

[2868] Nicolas Gast. Session details: Resource management II. ACM SIGMETRICS Performance Evaluation Review, 46(1), June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Golubchik:2018:SDL

[2869] Leana Golubchik. Session details: Learning I. ACM SIGMETRICS Performance Evaluation Review, 46(1), June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Harchol-Balter:2018:SDS

[2870] Mor Harchol-Balter. Session details: Scheduling I. ACM SIGMETRICS Performance Evaluation Review, 46(1), June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Houdt:2018:SDN

[2871] Benny Van Houdt. Session details: Networking. ACM SIGMETRICS Performance Evaluation Review, 46(1), June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Maguluri:2018:SDE

[2872] Siva Theja Maguluri. Session details: Emerging areas. ACM SIGMETRICS Performance Evaluation Review, 46(1), June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Misra:2018:SDR

[2873] Vishal Misra. Session details: Resource management I. ACM SIGMETRICS Performance Evaluation Review, 46(1), June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ren:2018:SDS

[2874] Shaolei Ren. Session details: Systems. ACM SIGMETRICS Performance Evaluation Review, 46(1), June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shah:2018:SDL

[2875] Devavrat Shah. Session details: Learning II. ACM SIGMETRICS Performance Evaluation Review, 46(1), June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Squillante:2018:SIW

[2876] Mark S. Squillante. Special issue on The Workshop on MAthematical performance Modeling and Analysis (MAMA 2018). ACM SIGMET-RICS Performance Evaluation Review, 46(2):2, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xie:2018:MDP

[2877] Hong Xie and John C. S Lui. A Markov decision process approach to analyze discount & reputation trade-offs in ecommerce systems. ACM SIGMET-RICS Performance Evaluation Review, 46(2):3–5, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yang:2018:MAC

[2878] Yudong Yang, Vishal Misra, and Dan Rubenstein. A modeling approach to classifying malicious cloud users via shuffling. ACM SIGMETRICS Performance Evaluation Review, 46 (2):6–8, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Grosof:2018:SMSa

[2879] Isaac Grosof, Ziv Scully, and Mor Harchol-Balter. SRPT for multiserver systems. ACM SIGMETRICS Performance Evaluation Review, 46(2):9–11, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nakahira:2018:MVDa

[2880] Yorie Nakahira, Andres Ferragut, and Adam Wierman. Minimal-variance distributed scheduling under strict demands and deadlines. ACM SIGMET-RICS Performance Evaluation Review, 46(2):12–14, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2018:SSS

[2881] Xin Liu and Lei Ying. A simple steady-state analysis of load balancing algorithms in the sub-Halfin– Whitt regime. ACM SIGMETRICS Performance Evaluation Review, 46 (2):15–17, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Mukherjee:2018:JIQ

[2882] Debankur Mukherjee and Alexander Stolyar. Join-idle-queue with service elasticity. ACM SIGMETRICS Performance Evaluation Review, 46 (2):18–20, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sun:2018:FAH

[2883] Xiao Sun, Tan N. Le, Mosharaf Chowdhury, and Zhenhua Liu. Fair allocation of heterogeneous and interchangeable resources. ACM SIGMET-RICS Performance Evaluation Review, 46(2):21–23, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ayesta:2018:RDC

[2884] Urtzi Ayesta. On redundancy-d with cancel-on-start a.k.a join-shortest-work (d). ACM SIGMETRICS Performance Evaluation Review, 46(2):24–26, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Panigrahy:2018:QTM

[2885] Nitish K. Panigrahy, Prithwish Basu, Don Towsley, Ananthram Swami, Kevin S. Chan, and Kin K. Leung. A queueing-theoretic model for resource allocation in one-dimensional distributed analytics network? ACM SIGMETRICS Performance Evaluation Review, 46(2):27–29, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gast:2018:RMFb

[2886] Nicolas Gast, Diego Latella, and Mieke Massink. A refined mean field approximation for synchronous population processes. ACM SIGMETRICS Performance Evaluation Review, 46 (2):30–32, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shneer:2018:SSD

[2887] Seva Shneer and Alexander Stolyar.

Stability of a standard decentralised medium access. ACM SIGMETRICS Performance Evaluation Review, 46 (2):33–35, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sabnis:2018:OOB

[2888] Anirudh Anirudh Sabnis, Ramesh K.

Sitaraman, and Donald Towsley. OC-CAM: an optimization based approach to network inference. ACM SIGMET-RICS Performance Evaluation Review, 46(2):36–38, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Qin:2018:CPIa

[2889] Junjie Qin, Ram Rajagopal, Shai Vardi, and Adam Wierman. Convex prophet inequalities. ACM SIGMET-RICS Performance Evaluation Review, 46(2):39–41, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goel:2018:SOC

[2890] Gautam Goel. Smoothed online convex optimization via online balanced descent. ACM SIGMETRICS Performance Evaluation Review, 46(2): 42–44, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ghosh:2018:MMO

[2891] Soumyadip Ghosh, Mark S. Squillante, and Ebisa D. Wollega. On min-max optimization over large data sets. ACM SIGMETRICS Performance Evaluation Review, 46(2):45–47, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thai:2018:ASI

[2892] My T. Thai, Arun Sen, and Arun Das. ACM SIGMETRICS international workshop on critical infrastructure network security. ACM SIGMET-RICS Performance Evaluation Review, 46(2):48–49, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Soltan:2018:RCP

[2893] Saleh Soltan, Mihalis Yannakakis, and Gil Zussman. REACT to cyberphysical attacks on power grids (extended abstract). ACM SIGMET-RICS Performance Evaluation Review, 46(2):50–51, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Huang:2018:ASC

[2894] Linan Huang and Quanyan Zhu. Adaptive strategic cyber defense for advanced persistent threats in critical infrastructure networks. ACM SIGMET-RICS Performance Evaluation Review, 46(2):52–56, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Guo:2018:FLPa

[2895] Linqi Guo, Chen Liang, Alessandro Zocca, Steven H. Low, and Adam Wierman. Failure localization in power systems via tree partitions. ACM SIGMETRICS Performance Evaluation Review, 46(2):57–61, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Oostenbrink:2018:ELD

[2896] Jorik Oostenbrink, Fernando A.

Kuipers, Poul E. Heegaard, and Bjarne E. Helvik. Evaluating local disaster recovery strategies. ACM SIGMETRICS Performance Evaluation Review, 46(2):62–66, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Khamfroush:2018:VII

[2897] Hana Khamfroush, Samuel Lofumbwa Iloo, and Mahshid Rahnamay-Naeini.

Vulnerability of interdependent infrastructures under random attacks. ACM SIGMETRICS Performance Evaluation Review, 46(2):67–71, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kelic:2018:CRC

[2898] Andjelka Kelic. Cyber risk in critical infrastructure. ACM SIGMET-RICS Performance Evaluation Review, 46(2):72–75, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Liu:2018:SIW

[2899] Zhenhua Liu and Ganesh Ananthanarayanan. Special issue on the workin-progress (WIP) session at the ACM SIGMETRICS 2018 Conference. ACM SIGMETRICS Performance Evaluation Review, 46(2):76, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Le:2018:BMB

[2900] Tan N. Le, Xiao Sun, Mosharaf Chowdhury, and Zhenhua Liu. BoPF: Mitigating the burstiness-fairness tradeoff in multi-resource clusters. ACM SIGMETRICS Performance Evaluation Review, 46(2):77–78, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Guo:2018:FLPb

[2901] Linqi Guo, Chen Liang, Alessandro Zocca, Steven H. Low, and Adam Wierman. Failure localization in power systems via tree partitions. ACM SIGMETRICS Performance Evaluation Review, 46(2):79–80, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hurtado-Lange:2018:NVD

[2902] Daniela Hurtado-Lange and Siva Theja Maguluri. A novel view of the drift: Method for heavy traffic limits of queueing systems. ACM SIGMET-RICS Performance Evaluation Review, 46(2):81–82, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ruan:2018:EVV

[2903] Yichen Ruan and Carlee Joe-Wong.

On the economic value of vehicular caching. ACM SIGMETRICS Performance Evaluation Review, 46(2): 83–84, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Qin:2018:CPIb

[2904] Junjie Qin, Ram Rajagopal, Shai Vardi, and Adam Wierman. Convex prophet inequalities. ACM SIGMET-RICS Performance Evaluation Review, 46(2):85–86, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Le:2018:AAA

[2905] Tan N. Le, Xiao Sun, Mosharaf Chowdhury, and Zhenhua Liu. AlloX: Allocation across computing resources for hybrid CPU/GPU clusters. ACM SIGMETRICS Performance Evaluation Review, 46(2):87–88, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zuo:2018:OBP

[2906] Jinhang Zuo, Xiaoxi Zhang, and Carlee Joe-Wong. Observe before play: Multi-armed bandit with preobservations. ACM SIGMETRICS Performance Evaluation Review, 46 (2):89–90, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2018:MPL

[2907] Justin Wang, Benjamin Berg, Daniel S. Berger, and Siddhartha Sen. Maximizing page-level cache hit ratios in LargeWeb services. ACM SIGMET-RICS Performance Evaluation Review, 46(2):91–92, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shao:2018:FLT

[2908] Zhihui Shao, Mohammad A. Islam, and Shaolei Ren. A first look at thermal attacks in multi-tenant data centers. ACM SIGMETRICS Performance Evaluation Review, 46(2): 93–94, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shen:2018:ENM

[2909] Shan-Hsiang Shen. An efficient network monitor for SDN networks. ACM SIGMETRICS Performance Evaluation Review, 46(2):95–96, September 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tassiulas:2018:ONE

[2910] Leandros Tassiulas. Optimizing the network edge for flexible service provisioning. ACM SIGMETRICS Performance Evaluation Review, 46(3):1, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2018:DAB

[2911] Weina Wang, Mor Harchol-Balter, Haotian Jiang, Alan Scheller-Wolf, and R. Srikant. Delay asymptotics and bounds for multi-task parallel jobs. ACM SIGMETRICS Performance Evaluation Review, 46(3):2–7, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Grosof:2018:SMSb

[2912] Isaac Grosof, Ziv Scully, and Mor Harchol-Balter. SRPT for multiserver systems. ACM SIGMETRICS Performance Evaluation Review, 46(3):8–9, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhou:2018:FLB

[2913] Xingyu Zhou, Jian Tan, and Ness Shroff. Flexible load balancing with multi-dimensional state-space collapse: Throughput and heavy-traffic delay optimality. ACM SIGMETRICS Performance Evaluation Review, 46(3):10–11, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chatzieleftheriou:2018:JUA

[2914] L. E. Chatzieleftheriou, G. Darzanos, M. Karaliopoulos, and I. Koutsopoulos. Joint user association, content caching and recommendations in wireless edge networks. ACM SIGMET-RICS Performance Evaluation Review, 46(3):12–17, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wong:2018:HSM

[2915] Yung Fei Wong, Lachlan L. H. Andrew, and Y. Ahmet Sekercioglu. Hidden semi-Markov models for electricity load disaggregation. ACM SIGMET-RICS Performance Evaluation Review, 46(3):18–23, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kakhki:2018:IMW

[2916] Arash Molavi Kakhki, Vijay Erramilli, Phillipa Gill, Augustin Chaintreau, and Balachander Krishnamurthy. Information market for Web browsing: Design, usability and incremental adoption. ACM SIGMETRICS Performance Evaluation Review, 46(3):24, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gast:2018:SEM

[2917] Nicolas Gast, Luca Bortolussi, and Mirco Tribastone. Size expansions of mean field approximation: Transient and steady-state analysis. ACM SIGMETRICS Performance Evaluation Review, 46(3):25–26, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bermolen:2018:DGA

[2918] Paola Bermolen, Matthieu Jonckheere, Federico Larroca, and Manuel Saenz. Degree-greedy algorithms on large random graphs. ACM SIGMETRICS Performance Evaluation Review, 46 (3):27–32, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yajima:2018:CLT

[2919] M. Yajima and T. Phung-Duc. A central limit theorem for a Markovmodulated infinite-server queue with batch Poisson arrivals and binomial catastrophes. ACM SIGMETRICS Performance Evaluation Review, 46 (3):33–34, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shi:2018:WFC

[2920] Lianjie Shi, Xin Wang, Richard T. B.

Ma, and Y. C. Tay. Weighted fair caching: Occupancy-centric allocation for space-shared resources. ACM SIGMETRICS Performance Evaluation Review, 46(3):35–36, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Carlsson:2018:WCB

[2921] Niklas Carlsson and Derek Eager.

Worst-case bounds and optimized cache on M-th request cache insertion policies under elastic conditions. ACM SIGMETRICS Performance Evaluation Review, 46(3):37–38, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ni:2018:WEW

[2922] Fan Ni, Xingbo Wu, Weijun Li, Lei Wang, and Song Jiang. WOJ: Enabling write-once full-data journaling in SSDs by using weak-hashing-based deduplication. ACM SIGMETRICS Performance Evaluation Review, 46 (3):39–40, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Durand:2018:EBR

[2923] Stephane Durand, Federica Garin, and Bruno Gaujal. Efficiency of best response dynamics with high playing rates in potential games. ACM SIGMETRICS Performance Evaluation Review, 46(3):41–42, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chouayakh:2018:AML

[2924] Ayman Chouayakh, Aurelien Bechler, Isabel Amigo, Loutfi Nuaymi, and Patrick Maillé. Auction mechanisms for licensed shared access: reserve prices and revenue-fairness trade offs. ACM SIGMETRICS Performance Evaluation Review, 46(3):43–48, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zou:2018:TEM

[2925] Mao Zou, Richard T. B. Ma, and Yinlong Xu. Towards an efficient market mediator for divisible resources. ACM SIGMETRICS Performance Evaluation Review, 46(3):49–50, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thiran:2018:LSD

[2926] Patrick Thiran. Locating the source of diffusion in large-scale and random networks. ACM SIGMETRICS Performance Evaluation Review, 46(3):51, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zocca:2018:TSM

[2927] Alessandro Zocca. Temporal starvation in multi-channel CSMA networks: an analytical framework. ACM SIGMET-RICS Performance Evaluation Review, 46(3):52–53, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vasantam:2018:MFB

[2928] Thirupathaiah Vasantam, Arpan Mukhopadhyay, and Ravi R. Mazumdar. The mean-field behavior of processor sharing systems with general job lengths under the SQ(d) policy. ACM SIGMETRICS Performance Evaluation Review, 46(3):54–55, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nakahira:2018:MVDb

[2929] Yorie Nakahira, Andres Ferragut, and Adam Wierman. Minimal-variance distributed deadline scheduling in a stationary environment. ACM SIGMET-RICS Performance Evaluation Review, 46(3):56–61, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zeballos:2018:AFE

[2930] Martin Zeballos, Andres Ferragut, and Fernando Paganini. Achieving fairness for EV charging in overload: a fluid approach. ACM SIGMETRICS Performance Evaluation Review, 46 (3):62–67, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hargreaves:2018:FOS

[2931] Eduardo Hargreaves, Claudio Agosti, Daniel Menasche, Giovanni Neglia, Alexandre Reiffers-Masson, and Eitan Altman. Fairness in online social network timelines: Measurements, models and mechanism design. ACM SIGMETRICS Performance Evaluation Review, 46(3):68–69, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Floquet:2018:HBR

[2932] Julien Floquet, Richard Combes, and Zwi Altman. Hierarchical beamforming: Resource allocation, fairness and flow level performance. ACM SIGMETRICS Performance Evaluation Review, 46(3):70–71, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Raaijmakers:2018:DPP

[2933] Y. Raaijmakers, S. C. Borst, and O. J. Boxma. Delta probing policies for redundancy. ACM SIGMET-RICS Performance Evaluation Review, 46(3):72–73, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hellemans:2018:ARD

[2934] T. Hellemans and B. Vanhoudt. Analysis of redundancy(d) with identical replicas. ACM SIGMETRICS Performance Evaluation Review, 46(3):74–79, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ayesta:2018:UPF

[2935] Urtzi Ayesta, Tejas Bodas, and Ina Maria Verloop. On a unifying product form framework for redundancy models. ACM SIGMETRICS Performance Evaluation Review, 46 (3):80–81, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rosenberg:2018:HTB

[2936] Catherine Rosenberg. Highlight talk on battery modeling: Trade-offs between accuracy and complexity. ACM SIGMETRICS Performance Evaluation Review, 46(3):82–83, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ardakanian:2018:LSD

[2937] Omid Ardakanian. Leveraging sparsity in distribution grids: System identification and harmonic state estimation. ACM SIGMETRICS Performance Evaluation Review, 46(3):84–85, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Danner:2018:SEP

[2938] Dominik Danner and Hermann de Meer.

State estimation in the power distribution system. ACM SIGMETRICS Performance Evaluation Review, 46 (3):86–88, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vinot:2018:CAL

[2939] Benoit Vinot, Florent Cadoux, and Nicolas Gast. Congestion avoidance in low-voltage networks by using the advanced metering infrastructure. ACM SIGMETRICS Performance Evaluation Review, 46(3):89–91, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tang:2018:PDG

[2940] Yujie Tang, Emiliano Dall’Anese, Andrey Bernstein, and S. H. Low. A primal-dual gradient method for timevarying optimization with application to power systems. ACM SIGMET-RICS Performance Evaluation Review, 46(3):92, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Alagha:2018:SAI

[2941] Nader Alagha. Satellite air interface evolutions in the 5G and IoT era. ACM SIGMETRICS Performance Evaluation Review, 46(3):93–95, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Palattella:2018:AMT

[2942] Maria Rita Palattella, Ridha Soua, André Stemper, and Thomas Engel. Aggregation of MQTT topics over integrated satellite-terrestrial networks. ACM SIGMETRICS Performance Evaluation Review, 46(3):96–97, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Luong:2018:SHV

[2943] Doanh Kim Luong, Muhammad Ali, Fouad Benamrane, Ibrahim Ammar, and Yim-Fun Hu. Seamless handover for video streaming over an SDN-based aeronautical communications network. ACM SIGMETRICS Performance Evaluation Review, 46 (3):98–99, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bas:2018:IMS

[2944] J. Bas, M. Caus, A. Perez, R. Soleymani, and N. A. K. Beigi. Interference management schemes for high spectral efficiency satellite communications. ACM SIGMETRICS Performance Evaluation Review, 46(3): 100–103, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Medina-Caballero:2018:LQO

[2945] Julio A. Medina-Caballero and M. Angeles Vazquez-Castro. Link quality optimization for hybrid LEO–GSO systems. ACM SIGMETRICS Performance Evaluation Review, 46(3): 104–107, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Asuquo:2018:SEC

[2946] P. Asuquo, H. Cruickshank, and C. P. A. Ogah. Securing emergency communications for the nextgeneration networks: Performance analysis of distributed trust computing for future emergency communications. ACM SIGMETRICS Performance Evaluation Review, 46(3): 108–111, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kuhn:2018:RTS

[2947] Nicolas Kuhn. Research trends in SAT-COM. ACM SIGMETRICS Performance Evaluation Review, 46(3):112, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

vanMoorsel:2018:BMB

[2948] Aad van Moorsel. Benchmarks and models for blockchain: Consensus algorithms. ACM SIGMETRICS Performance Evaluation Review, 46(3):113, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hellemans:2018:MCM

[2949] Tim Hellemans, Benny Van Houdt, Daniel S. Menasche, Mandar Datar, Swapnil Dhamal, and Corinne Touati. Mining competition in a multicryptocurrency ecosystem at the network edge: a congestion game approach. ACM SIGMETRICS Performance Evaluation Review, 46(3): 114–117, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zander:2018:DSD

[2950] Manuel Zander, Tom Waite, and Dominik Harz. DAGsim: Simulation of DAG-based distributed ledger protocols. ACM SIGMETRICS Performance Evaluation Review, 46(3): 118–121, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ricci:2018:LBD

[2951] Saulo Ricci, Eduardo Ferreira, Daniel Sadoc Menasche, Artur Ziviani, Jose Eduardo Souza, and Alex Borges Vieira. Learning blockchain delays: a queueing theory approach. ACM SIGMETRICS Performance Evaluation Review, 46(3): 122–125, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Grunspan:2018:PBW

[2952] C. Grunspan and R. Perez-Marco.

On profitability of block withholding strategies. ACM SIGMETRICS Performance Evaluation Review, 46 (3):126, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bruschi:2018:MIS

[2953] Francesco Bruschi, Vincenzo Rana, Lorenzo Gentile, and Donatella Sciuto. Mine with it or sell it: the superhashing power dilemma. ACM SIGMET-RICS Performance Evaluation Review, 46(3):127–130, December 2018. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Smuts:2018:WDC

[2954] Nico Smuts. What drives cryptocurrency prices?: an investigation of Google Trends and Telegram sentiment. ACM SIGMETRICS Performance Evaluation Review, 46(3): 131–134, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Alharby:2018:BSF

[2955] Maher Alharby and Aad van Moorsel.

BlockSim: a simulation framework for blockchain systems. ACM SIGMET-RICS Performance Evaluation Review, 46(3):135–138, December 2018. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fedchenko:2018:FNN

[2956] Vladyslav Fedchenko, Giovanni Neglia, and Bruno Ribeiro. Feedforward neural networks for caching: Enough or too much? ACM SIGMETRICS Performance Evaluation Review, 46(3): 139–142, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Trevisan:2018:RUC

[2957] Martino Trevisan and Idilio Drago. Robust URL classification with generative adversarial networks. ACM SIGMET-RICS Performance Evaluation Review, 46(3):143–146, December 2018. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Marin:2018:DMR

[2958] Gonzalo Marín, Pedro Casas, and Germán Capdehourat. DeepSec meets RawPower — deep learning for detection of network attacks using raw representations. ACM SIGMETRICS Performance Evaluation Review, 46(3): 147–150, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Piskozub:2018:MDM

[2959] Michal Piskozub, Riccardo Spolaor, and Ivan Martinovic. MalAlert: Detecting malware in large-scale network traffic using statistical features. ACM SIGMETRICS Performance Evaluation Review, 46(3):151–154, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wassermann:2018:MLM

[2960] Sarah Wassermann, Nikolas Wehner, and Pedro Casas. Machine learning models for YouTube QoE and user engagement prediction in Smartphones. ACM SIGMETRICS Performance Evaluation Review, 46(3):

155–158, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Manzo:2018:DLS

[2961] Gaetano Manzo, Juan Sebastian Otalora, Marco Ajmone Marsan, and Gianluca Rizzo. A deep learning strategy for vehicular floating content management. ACM SIGMETRICS Performance Evaluation Review, 46(3): 159–162, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Li:2018:LDM

[2962] Rui Li, Chaoyun Zhang, Paul Patras, Razvan Stanica, and Fabrice Valois. Learning driven mobility control of airborne base stations in emergency networks. ACM SIGMETRICS Performance Evaluation Review, 46(3): 163–166, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Garcia:2018:RCP

[2963] Johan Garcia and Topi Korhonen.

On runtime and classification performance of the discretize–optimize (DISCO) classification approach. ACM SIGMETRICS Performance Evaluation Review, 46(3):167–170, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hanawal:2018:DLA

[2964] Manjesh K. Hanawal and Sumit J.

Darak. Distributed learning in adhoc networks with unknown number of players. ACM SIGMETRICS Performance Evaluation Review, 46(3): 171–174, December 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pal:2019:ESC

[2965] Ranjan Pal, Aditya Ahuja, Sung-Han Lin, Abhishek Kumar, Leana Golubchik, and Nachikethas A. Jagadeesan. On the economic sustainability of cloud sharing systems are dynamic single resource sharing markets stable? ACM SIGMETRICS Performance Evaluation Review, 46(4):2–10, March 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Su:2019:CLB

[2966] Lili Su, Martin Zubeldia, and Nancy Lynch. Collaboratively learning the best option on graphs, using bounded local memory. ACM SIGMETRICS Performance Evaluation Review, 47 (1):1–2, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/3376930.
                                        3376932.

Hoffmann:2019:LGN

[2967] Jessica Hoffmann and Constantine Caramanis. Learning graphs from noisy epidemic cascades. ACM SIGMET-RICS Performance Evaluation Review, 47(1):3–4, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/3376930.
                                        3376933.

Zhou:2019:HTD

[2968] Xingyu Zhou, Jian Tan, and Ness Shroff. Heavy-traffic delay optimality in pull-based load balancing systems: Necessary and sufficient conditions. ACM SIGMETRICS Performance Evaluation Review, 47(1):5–6, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3376930.3376935.

Hellemans:2019:PAW

[2969] Tim Hellemans, Tejas Bodas, and Benny Van Houdt. Performance analysis of workload dependent load balancing policies. ACM SIGMETRICS Performance Evaluation Review, 47(1):7– 8, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3376930.3376936.

Grosof:2019:LBG

[2970] Isaac Grosof, Ziv Scully, and Mor Harchol-Balter. Load balancing guardrails: Keeping your heavy traffic on the road to low response times. ACM SIGMETRICS Performance Evaluation Review, 47(1):9–10, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3376930.3376937.

Tang:2019:RWB

[2971] Dengwang Tang and Vijay G. Subramanian. Random walk based sampling for load balancing in multi-server systems. ACM SIGMETRICS Performance Evaluation Review, 47(1):11–12, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3376930.3376938.

Nguyen:2019:NRA

[2972] Lan N. Nguyen and My T. Thai. Network resilience assessment via QoS degradation metrics: an algorithmic approach. ACM SIGMETRICS Performance Evaluation Review, 47(1):13–14, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3376930.3376940.

Kumar:2019:TBA

[2973] Dhruv Kumar, Jian Li, Ramesh Sitaraman, and Abhishek Chandra. A TTLbased approach for data aggregation in geo-distributed streaming analytics. ACM SIGMETRICS Performance Evaluation Review, 47(1):15–16, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3376930.3376941.

Nikolopoulos:2019:RPS

[2974] Pavlos Nikolopoulos, Christos Pappas, Katerina Argyraki, and Adrian Perrig. Retroactive packet sampling for traffic receipts. ACM SIGMETRICS Performance Evaluation Review, 47(1):17–18, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3376930.3376942.

Sermpezis:2019:ICI

[2975] Pavlos Sermpezis and Vasileios Kotronis. Inferring catchment in Internet routing. ACM SIGMETRICS Performance Evaluation Review, 47(1):19–20, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3376930.3376943.

Akram:2019:CGP

[2976] Shoaib Akram, Jennifer B. Sartor, Kathryn S. McKinley, and Lieven Eeckhout. Crystal Gazer: Profiledriven write-rationing garbage collection for hybrid memories. ACM SIGMETRICS Performance Evaluation Review, 47(1):21–22, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376945.

Karakoy:2019:AAA

[2977] Mustafa Karakoy, Orhan Kislal, Xulong Tang, Mahmut Taylan Kandemir, and Meenakshi Arunachalam. Architecture-aware approximate computing. ACM SIGMETRICS Performance Evaluation Review, 47(1):23–24, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3376930.3376946.

Tang:2019:QDL

[2978] Xulong Tang, Ashutosh Pattnaik, Onur Kayiran, Adwait Jog, Mahmut Taylan Kandemir, and Chita Das. Quantifying data locality in dynamic parallelism in GPUs. ACM SIGMETRICS Performance Evaluation Review, 47(1):25–26, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3376930.3376947.

Tang:2019:CND

[2979] Xulong Tang, Mahmut Taylan Kandemir, Hui Zhao, Myoungsoo Jung, and Mustafa Karakoy. Computing with near data. ACM SIGMETRICS Performance Evaluation Review, 47 (1):27–28, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/3376930.
                                        3376948.

Balseiro:2019:DPR

[2980] Santiago R. Balseiro, David B. Brown, and Chen Chen. Dynamic pricing of relocating resources in large networks. ACM SIGMETRICS Performance Evaluation Review, 47(1):29–30, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3376930.3376950.

Alijani:2019:STT

[2981] Reza Alijani, Siddhartha Banerjee, Sreenivas Gollapudi, Kostas Kollias, and Kamesh Munagala. The segmentation-thickness tradeoff in online marketplaces. ACM SIGMET-RICS Performance Evaluation Review, 47(1):31–32, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/3376930.
                                        3376951.

Shi:2019:VLA

[2982] Ming Shi, Xiaojun Lin, and Lei Jiao. On the value of look-ahead in competitive online convex optimization. ACM SIGMETRICS Performance Evaluation Review, 47(1):33–34, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3376930.3376952.

Lin:2019:COO

[2983] Qiulin Lin, Hanling Yi, John Pang, Minghua Chen, Adam Wierman, Michael Honig, and Yuanzhang Xiao. Competitive online optimization under inventory constraints. ACM SIGMET-RICS Performance Evaluation Review, 47(1):35–36, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3376930.
3376953.

Yu:2019:ALB

[2984] Haoran Yu, Ermin Wei, and Randall A.

Berry. Analyzing location-based advertising for vehicle service providers using effective resistances. ACM SIGMETRICS Performance Evaluation Review, 47(1):37–38, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376955.

Vial:2019:SRP

[2985] Daniel Vial and Vijay Subramanian.

A structural result for personalized PageRank and its algorithmic consequences. ACM SIGMETRICS Performance Evaluation Review, 47(1):39–40, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3376930.3376956.

Cayci:2019:LCR

[2986] Semih Cayci, Atilla Eryilmaz, and R. Srikant. Learning to control renewal processes with bandit feedback. ACM SIGMETRICS Performance Evaluation Review, 47(1):41–42, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376957.

Henzinger:2019:EDR

[2987] Monika Henzinger, Stefan Neumann, and Stefan Schmid. Efficient distributed workload (re-)Embedding. ACM SIGMETRICS Performance Evaluation Review, 47(1):43–44, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3376930.3376959.

Ambati:2019:OCE

[2988] Pradeep Ambati and David Irwin. Optimizing the cost of executing mixed interactive and batch workloads on transient VMs. ACM SIGMETRICS Performance Evaluation Review, 47 (1):45–46, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/3376930.
                                        3376960.

Comden:2019:OOC

[2989] Joshua Comden, Sijie Yao, Niangjun Chen, Haipeng Xing, and Zhenhua Liu. Online optimization in cloud resource provisioning: Predictions, regrets, and algorithms. ACM SIGMET-RICS Performance Evaluation Review, 47(1):47–48, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/3376930.
                                        3376961.

Quan:2019:NFM

[2990] Guocong Quan, Jian Tan, Atilla Eryilmaz, and Ness Shroff. A new flexible multi-flow LRU cache management paradigm for minimizing misses. ACM SIGMETRICS Performance Evaluation Review, 47(1):49–50, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376962.

Zarchy:2019:ACC

[2991] Doron Zarchy, Radhika Mittal, Michael Schapira, and Scott Shenker. Axiomatizing congestion control. ACM SIGMETRICS Performance Evaluation Review, 47(1):51–52, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376964.

Xu:2019:IMC

[2992] Kuang Xu and Yuan Zhong. Information, memory and capacity in dynamic resource allocation. ACM SIGMETRICS Performance Evaluation Review, 47(1):53–54, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376965.

Amjad:2019:MMD

[2993] Muhammad Jehangir Amjad, Vishal Misra, Devavrat Shah, and Dennis Shen. mRSC: Multi-dimensional robust synthetic control. ACM SIGMET-RICS Performance Evaluation Review, 47(1):55–56, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3376930.
3376966.

Jose:2019:DAC

[2994] Lavanya Jose, Stephen Ibanez, Mohammad Alizadeh, and Nick McKeown. A distributed algorithm to calculate max-min fair rates without per-flow state. ACM SIGMETRICS Performance Evaluation Review, 47(1):57–58, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3376930.3376967.

VanHoudt:2019:GAO

[2995] Benny Van Houdt. Global attraction of ODE-based mean field models with hyperexponential job sizes. ACM SIGMETRICS Performance Evaluation Review, 47(1):59–60, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3376930.3376969.

vanderBoor:2019:HSJ

[2996] Mark van der Boor, Sem Borst, and Johan van Leeuwaarden. Hyper-scalable JSQ with sparse feedback. ACM SIGMETRICS Performance Evaluation Review, 47(1):61–62, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3376930.3376970.

Ciucu:2019:TEK

[2997] Florin Ciucu and Felix Poloczek. Two extensions of Kingman’s GI/G/1 bound. ACM SIGMETRICS Performance Evaluation Review, 47(1):63–64, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3376930.3376971.

Ciucu:2019:QLD

[2998] Florin Ciucu, Felix Poloczek, and Amr Rizk. Queue and loss distributions in finite-buffer queues. ACM SIGMETRICS Performance Evaluation Review, 47(1):65–66, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376972.

Xu:2019:IQS

[2999] Jiaming Xu and Yuan Zhong. Improved queue-size scaling for inputqueued switches via graph factorization. ACM SIGMETRICS Performance Evaluation Review, 47(1):67–68, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3376930.3376973.

Han:2019:TWD

[3000] Youil Han, Bryan S. Kim, Jeseong Yeon, Sungjin Lee, and Eunji Lee. TeksDB: Weaving data structures for a high-performance key–value store. ACM SIGMETRICS Performance Evaluation Review, 47(1):69–70, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376975.

Radulovic:2019:PMS

[3001] Milan Radulovic, Rommel Sánchez Verdejo, Paul Carpenter, Petar Radojković, Bruce Jacob, and Eduard Ayguadé. PROFET: Modeling system performance and energy without simulating the CPU. ACM SIGMETRICS Performance Evaluation Review, 47(1):71–72, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376976.

Wei:2019:HBS

[3002] Song Wei, Kun Zhang, and Bibo Tu. HyperBench: a benchmark suite for virtualization capabilities. ACM SIGMETRICS Performance Evaluation Review, 47(1):73–74, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3376930.3376977.

Zhang:2019:AMD

[3003] Lei Zhang, Zhemin Yang, Yuyu He, Mingqi Li, Sen Yang, Min Yang, Yuan Zhang, and Zhiyun Qian. App in the middle: Demystify application virtualization in Android and its security threats. ACM SIGMETRICS Performance Evaluation Review, 47(1):75–76, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3376930.3376978.

Ngoc:2019:EYS

[3004] Tu Dinh Ngoc, Bao Bui, Stella Bitchebe, Alain Tchana, Valerio Schiavoni, Pascal Felber, and Daniel Hagimont. Everything you should know about Intel SGX performance on virtualized systems. ACM SIGMETRICS Performance Evaluation Review, 47 (1):77–78, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/3376930.
                                        3376979.

Wei:2019:QMO

[3005] Honghao Wei, Xiaohan Kang, Weina Wang, and Lei Ying. QuickStop: a Markov optimal stopping approach for quickest misinformation detection. ACM SIGMETRICS Performance Evaluation Review, 47(1):79–80, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3376930.3376981.

Vera:2019:BPL

[3006] Alberto Vera and Siddhartha Banerjee. The Bayesian prophet: a low-regret framework for online decision making. ACM SIGMETRICS Performance Evaluation Review, 47(1):81–82, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376982.

Su:2019:SDG

[3007] Lili Su and Jiaming Xu. Securing distributed gradient descent in high dimensional statistical learning. ACM SIGMETRICS Performance Evaluation Review, 47(1):83–84, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376983.

Agarwal:2019:MAT

[3008] Anish Agarwal, Muhammad Jehangir Amjad, Devavrat Shah, and Dennis Shen. Model agnostic time series analysis via matrix estimation. ACM SIGMETRICS Performance Evaluation Review, 47(1):85–86, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376984.

Pourghassemi:2019:WIA

[3009] Behnam Pourghassemi, Ardalan Amiri Sani, and Aparna Chandramowlishwaran. What-if analysis of page load time in Web browsers using causal profiling. ACM SIGMETRICS Performance Evaluation Review, 47(1):87–88, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3376930.3376986.

Liu:2019:PCL

[3010] Ran Liu, Edmund Yeh, and Atilla Eryilmaz. Proactive caching for low accessdelay services under uncertain predictions. ACM SIGMETRICS Performance Evaluation Review, 47(1):89–90, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3376930.3376987.

Zhu:2019:UNP

[3011] Xiao Zhu, Yihua Ethan Guo, Ashkan Nikravesh, Feng Qian, and Z. Morley Mao. Understanding the networking performance of Wear OS. ACM SIGMETRICS Performance Evaluation Review, 47(1):91–92, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3376930.3376988.

Ghose:2019:DCD

[3012] Saugata Ghose, Tianshi Li, Nastaran Hajinazar, Damla Senol Cali, and Onur Mutlu. Demystifying complex workload-DRAM interactions: an experimental study. ACM SIGMET-RICS Performance Evaluation Review, 47(1):93, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3376930.
3376989.

Lee:2019:NMM

[3013] Chul-Ho Lee, Min Kang, and Do Young Eun. Non-Markovian Monte Carlo on directed graphs. ACM SIGMET-RICS Performance Evaluation Review, 47(1):94–95, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3376930.
3376991.

Dai:2019:ACL

[3014] Osman Emre Dai, Daniel Cullina, Negar Kiyavash, and Matthias Grossglauser. Analysis of a canonical labeling algorithm for the alignment of correlated Erdős–Rényi graphs. ACM SIGMETRICS Performance Evaluation Review, 47(1):96–97, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3376930.3376992.

Combes:2019:CEE

[3015] Richard Combes and Mikael Touati.

Computationally efficient estimation of the spectral gap of a Markov chain. ACM SIGMETRICS Performance Evaluation Review, 47(1): 98–100, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3376930.
3376993.

Squillante:2019:SIW

[3016] Mark S. Squillante. Special issue on The Workshop on MAthematical performance Modeling and Analysis (MAMA 2019). ACM SIGMET-RICS Performance Evaluation Review, 47(2):2, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Abuthahir:2019:DWN

[3017] Abuthahir, Gaurav Raina, and Thomas Voice. Do we need two forms of feedback in the Rate Control Protocol (RCP)? ACM SIGMETRICS Performance Evaluation Review, 47 (2):3–5, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Goel:2019:OAS

[3018] Gautam Goel and Adam Wierman. An online algorithm for smoothed online convex optimization. ACM SIGMET-RICS Performance Evaluation Review, 47(2):6–8, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Tan:2019:OPP

[3019] Xiaoqi Tan, Alberto Leon-Garcia, and Danny H. K. Tsang. Optimal posted prices for online resource allocation with supply costs. ACM SIGMET-RICS Performance Evaluation Review, 47(2):9–11, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gardner:2019:SDH

[3020] Kristen Gardner and Cole Stephens. Smart dispatching in heterogeneous systems. ACM SIGMETRICS Performance Evaluation Review, 47(2): 12–14, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Anton:2019:RPS

[3021] Elene Anton, Urtzi Ayesta, Matthieu Jonckheere, and Ina Maria Verloop. Redundancy with processor sharing servers. ACM SIGMETRICS Performance Evaluation Review, 47(2): 15–17, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Berg:2019:HOS

[3022] Benjamin Berg, Rein Vesilo, and Mor Harchol-Balter. heSRPT: Optimal scheduling of parallel jobs with known sizes. ACM SIGMETRICS Performance Evaluation Review, 47(2):18–20, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Su:2019:CAS

[3023] Yu Su, Xiaoqi Ren, Shai Vardi, Adam Wierman, and Yuxiong He. Communication-aware scheduling of precedence-constrained tasks. ACM SIGMETRICS Performance Evaluation Review, 47(2):21–23, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Scully:2019:SNO

[3024] Ziv Scully, Mor Harchol-Balter, and Alan Scheller-Wolf. Simple near-optimal scheduling for the M/G/1. ACM SIGMETRICS Performance Evaluation Review, 47(2):24–26, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vardoyan:2019:SAQ

[3025] Gayane Vardoyan, Saikat Guha, Philippe Nain, and Don Towsley. On the stochastic analysis of a quantum entanglement switch. ACM SIGMET-RICS Performance Evaluation Review, 47(2):27–29, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Casale:2019:NSC

[3026] Giuliano Casale, Peter G. Harrison, and Ong Wai Hong. Novel solutions for closed queueing networks with loaddependent stations. ACM SIGMET-RICS Performance Evaluation Review, 47(2):30–32, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Somashekar:2019:TLT

[3027] Gagan Somashekar, Mohammad Delasay, and Anshul Gandhi. Tighter Lyapunov truncation for multi-dimensional continuous time Markov chains with known moments. ACM SIGMET-RICS Performance Evaluation Review, 47(2):33–35, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lange:2019:HTA

[3028] Daniela Hurtado Lange and Siva Theja Maguluri. Heavy-traffic analysis of the generalized switch under multidimensional state space collapse. ACM SIGMETRICS Performance Evaluation Review, 47(2):36–38, September 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Azizan:2020:OAL

[3029] Navid Azizan. Optimization algorithms for large-scale systems: From deep learning to energy markets. ACM SIGMETRICS Performance Evaluation Review, 47(3):2–5, January 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3380908.3380910.

Comden:2020:AOD

[3030] Joshua Comden. Algorithms for online and distributed optimization and their applications. ACM SIGMETRICS Performance Evaluation Review, 47(3):6– 9, January 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3380908.3380911.

Dipietro:2020:PMO

[3031] Salvatore Dipietro. Performance modelling and optimisation of NoSQL database systems. ACM SIGMET-RICS Performance Evaluation Review, 47(3):10–13, January 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3380908.
3380912.

Javadi:2020:AAD

[3032] Seyyed Ahmad Javadi. Analytical approaches for dynamic scheduling in cloud environments. ACM SIGMETRICS Performance Evaluation Review, 47(3):14–16, January 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3380908.3380913.

Li:2020:VSC

[3033] Qian Li. Vision-based sensor coverage in uncertain geometric domains. ACM SIGMETRICS Performance Evaluation Review, 47(3):17–19, January 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3380908.3380914.

Pourghassemi:2020:SDA

[3034] Behnam Pourghassemi. Scalable dynamic analysis of browsers for privacy and performance. ACM SIGMET-RICS Performance Evaluation Review, 47(3):20–23, January 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/3380908.
                                        3380915.

Su:2020:DDS

[3035] Lili Su. Defending distributed systems against adversarial attacks: Consensus, consensus-based learning, and statistical learning. ACM SIGMETRICS Performance Evaluation Review, 47(3):24– 27, January 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/abs/10.1145/3380908.3380916.

Wajahat:2020:CDM

[3036] Muhammad Wajahat. Cost-efficient dynamic management of cloud resources through supervised learning. ACM SIGMETRICS Performance Evaluation Review, 47(3):28–30, January 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3380908.3380917.

YU:2020:NCS

[3037] Pei-Duo YU. Network centralities as statistical inference for large networks: Combinatorics, probability and efficient graph algorithms. ACM SIGMETRICS Performance Evaluation Review, 47(3):31–33, January 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3380908.3380918.

Zhou:2020:AOLa

[3038] Xingyu Zhou. Asymptotically optimal load balancing: Theory and algorithms. ACM SIGMETRICS Performance Evaluation Review, 47(3):34–37, January 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3380908.3380919.

Palani:2020:OMS

[3039] Kartik Palani and David M. Nicol. Optimal monitoring strategies for critical infrastructure networks. ACM SIG-METRICS Performance Evaluation Review, 47(4):4–7, April 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3397776.
3397778.

Padhee:2020:IUP

[3040] Malhar Padhee, Reetam Sen Biswas, Anamitra Pal, Kaustav Basu, and Arunabha Sen. Identifying unique power system signatures for determining vulnerability of critical power system assets. ACM SIGMETRICS Performance Evaluation Review, 47(4):8– 11, April 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3397776.3397779.

Oostenbrink:2020:MWP

[3041] Jorik Oostenbrink and Fernando A. Kuipers. A moment of weakness: Protecting against targeted attacks following a natural disaster. ACM SIGMET-RICS Performance Evaluation Review, 47(4):12–15, April 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3397776.
3397780.

Pournaras:2020:CFI

[3042] Evangelos Pournaras, Riccardo Taormina, Manish Thapa, Stefano Galelli, Venkata Palleti, and Robert Kooij. Cascading failures in interconnected power-to-water networks. ACM SIGMETRICS Performance Evaluation Review, 47(4):16–20, April 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3397776.3397781.

Tan:2020:SPD

[3043] Yushi Tan, Arindam K. Das, Mareldi Ahumada-Paras, Payman Arabshahi, and Daniel S. Kirschen. Scheduling post-disaster repairs in electricity distribution networks with uncertain repair times. ACM SIGMETRICS Performance Evaluation Review, 47(4):21– 24, April 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3397776.3397782.

Kant:2020:APT

[3044] Krishna Kant. Advanced persistent threats in autonomous driving. ACM SIGMETRICS Performance Evaluation Review, 47(4):25–28, April 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3397776.3397783.

Banerjee:2020:ULA

[3045] Siddhartha Banerjee and Daniel Freund. Uniform loss algorithms for online stochastic decision-making with applications to bin packing. ACM SIG-METRICS Performance Evaluation Review, 48(1):1–2, July 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3410048.
3410050.

Wei:2020:OPD

[3046] Xiaohan Wei, Hao Yu, and Michael J.

Neely. Online primal-dual mirror descent under stochastic constraints. ACM SIGMETRICS Performance Evaluation Review, 48(1):3–4, July 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3410048.3410051.

Im:2020:DWF

[3047] Sungjin Im, Benjamin Moseley, Kamesh Munagala, and Kirk Pruhs. Dynamic weighted fairness with minimal disruptions. ACM SIGMETRICS Performance Evaluation Review, 48(1): 5–6, July 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3410048.3410052.

Yang:2020:OLO

[3048] Lin Yang, Mohammad H. Hajiesmaili, Ramesh Sitaraman, Adam Wierman, Enrique Mallada, and Wing S. Wong. Online linear optimization with inventory management constraints. ACM SIGMETRICS Performance Evaluation Review, 48(1):7–8, July 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3410048.3410053.

Lin:2020:OOP

[3049] Yiheng Lin, Gautam Goel, and Adam Wierman. Online optimization with predictions and non-convex losses. ACM SIGMETRICS Performance Evaluation Review, 48(1):9–10, July 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3410048.3410054.

Tan:2020:MDO

[3050] Xiaoqi Tan, Bo Sun, Alberto Leon-Garcia, Yuan Wu, and Danny H. K. Tsang. Mechanism design for online resource allocation: a unified approach. ACM SIGMETRICS Performance Evaluation Review, 48(1):11–12, July 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/abs/
                                        10.1145/3410048.3410055.

Alijani:2020:PMP

[3051] Reza Alijani, Siddhartha Banerjee, Sreenivas Gollapudi, Kamesh Munagala, and Kangning Wang. Predict and match: Prophet inequalities with uncertain supply. ACM SIGMETRICS Performance Evaluation Review, 48(1):

13–14, July 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3410048.3410056.

Bhattacharjee:2020:FLR

[3052] Rajarshi Bhattacharjee, Subhankar Banerjee, and Abhishek Sinha. Fundamental limits on the regret of online network-caching. ACM SIGMETRICS Performance Evaluation Review, 48(1): 15–16, July 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3410048.3410057.

Squillante:2020:SIW

[3053] Mark S. Squillante. Special issue on The Workshop on MAthematical performance Modeling and Analysis (MAMA 2020). ACM SIGMETRICS Performance Evaluation Review, 48(2): 2, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3439602.3439604.

Lu:2020:UGC

[3054] Ping-En Lu and Cheng-Shang Chang.

Using graph convolutional networks to compute approximations of dominant eigenvectors. ACM SIGMETRICS Performance Evaluation Review, 48(2): 3–5, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3439602.3439605.

Bhattacharjee:2020:CAM

[3055] Rajarshi Bhattacharjee and Abhishek Sinha. Competitive algorithms for minimizing the maximum age-of-information. ACM SIGMETRICS Performance Evaluation Review, 48(2):6– 8, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3439602.3439606.

Levy:2020:WCA

[3056] Hanoch Levy and Jhonatan Tavori. Worst case attacks on distributed resources systems. ACM SIGMETRICS Performance Evaluation Review, 48(2): 9–11, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3439602.3439607.

Bachmat:2020:PAO

[3057] Eitan Bachmat and Sveinung Erland. Performance analysis, optimization and optics. ACM SIGMETRICS Performance Evaluation Review, 48(2): 12–14, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3439602.3439608.

Datar:2020:RPC

[3058] Mandar Datar, Eitan Altman, and Ghilas Ferrat. Routing into parallel collision channels. ACM SIGMETRICS Performance Evaluation Review, 48(2): 15–17, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3439602.3439609.

Fricker:2020:MFA

[3059] Christine Fricker, Hanene Mohamed, and Cedric Bourdais. A mean field analysis of a stochastic model for reservation in car-sharing systems. ACM SIGMETRICS Performance Evaluation Review, 48(2):18–20, November

2020.  CODEN ????  ISSN

0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3439602.3439610.

Haverkort:2020:MLD

[3060] Boudewijn R. Haverkort, Felix Finkbeiner, and Pieter-Tjerk de Boer. Machine learning data center workloads using generative adversarial networks. ACM SIGMETRICS Performance Evaluation Review, 48(2):21–23, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3439602.3439611.

Menasche:2020:CTO

[3061] Daniel S. Menasché, Mark Shifrin, and Eduardo Hargreaves. Caches and timelines operate under heavy traffic. ACM SIGMETRICS Performance Evaluation Review, 48(2):24–26, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3439602.3439612.

Kesidis:2020:TGQ

[3062] George Kesidis and Takis Konstantopoulos. TB/GI/1 queues with arrival traffic envelopes. ACM SIGMETRICS Performance Evaluation Review, 48(2): 27–29, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3439602.3439613.

Jaleel:2020:GPD

[3063] Jazeem Abdul Jaleel, Alexander Wickeham, Sherwin Doroudi, and Kristen Gardner. A general “Power-of-d” dispatching framework for heterogeneous systems. ACM SIGMETRICS Performance Evaluation Review, 48(2):30–32, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3439602.3439614.

Scully:2020:OMS

[3064] Ziv Scully, Isaac Grosof, and Mor Harchol-Balter. Optimal multiserver scheduling with unknown job sizes in heavy traffic. ACM SIGMETRICS Performance Evaluation Review, 48(2): 33–35, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3439602.3439615.

Lu:2020:ODC

[3065] Yingdong Lu, Mark S. Squillante, and Tonghoon Suk. Optimal delaycost scheduling control in fluid models of general n × n input-queued switches. ACM SIGMETRICS Performance Evaluation Review, 48(2):36–38, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3439602.3439616.

Araldo:2020:ASI

[3066] Andrea Araldo, Ivona Brandic, and Stefan Schmid. ACM SIGMETRICS international workshop on distributed cloud computing (DCC). ACM SIG-METRICS Performance Evaluation Review, 48(2):39–40, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3439602.
3439617.

Balouek-Thomert:2020:HCC

[3067] Daniel Balouek-Thomert, Ivan Rodero, and Manish Parashar. Harnessing the computing continuum for urgent science. ACM SIGMETRICS Performance Evaluation Review, 48(2):41–46, November 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3439602.3439618.

Shi:2020:MRP

[3068] Lianjie Shi, Xin Wang, and Richard T. B. Ma. On multi-resource procurement in Internet access markets: Optimal strategies and market equilibrium. ACM SIGMETRICS Performance Evaluation Review, 48(3):4–5, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453953.3453954.

Hoque:2020:NAP

[3069] Mohammad A. Hoque, Ashwin Rao, and Sasu Tarkoma. Network and application performance measurement challenges on Android devices. ACM SIGMETRICS Performance Evaluation Review, 48(3):6–11, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453953.3453955.

Scherrer:2020:ISP

[3070] Simon Scherrer, Markus Legner, Adrian Perrig, and Stefan Schmid. Incentivizing stable path selection in future Internet architectures. ACM SIG-METRICS Performance Evaluation Review, 48(3):12–13, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3453953.
                                        3453956.

Christensen:2020:LIL

[3071] Niels Christensen, Mark Glavind, Stefan Schmid, and Jir’ Srba. Latte: Improving the latency of transiently consistent network update schedules. ACM SIGMETRICS Performance Evaluation Review, 48(3):14–26, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3453953.3453957.

Bronzino:2020:ISV

[3072] Francesco Bronzino, Paul Schmitt, Sara Ayoubi, Guilherme Martins, Renata Teixeira, and Nick Feamster. Inferring streaming video quality from encrypted traffic: Practical models and deployment experience. ACM SIGMETRICS Performance Evaluation Review, 48(3):27–32, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3453953.3453958.

Hurtado-Lange:2020:HTA

[3073] Daniela Hurtado-Lange and Siva Theja Maguluri. Heavy-traffic analysis of the generalized switch under multidimensional state space collapse. ACM SIGMETRICS Performance Evaluation Review, 48(3):33–34, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3453953.3453959.

Berg:2020:HPS

[3074] Benjamin Berg, Rein Vesilo, and Mor Harchol-Balter. heSRPT: Parallel scheduling to minimize mean slowdown. ACM SIGMETRICS Performance Evaluation Review, 48(3):35–36, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453953.3453960.

Gardner:2020:SLB

[3075] Kristen Gardner, Jazeem Abdul Jaleel, Alexander Wickeham, and Sherwin Doroudi. Scalable load balancing in the presence of heterogeneous servers. ACM SIGMETRICS Performance Evaluation Review, 48(3):37–38, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453953.3453961.

Dai:2020:LOR

[3076] Wenkai Dai, Klaus-Tycho Foerster, David Fuchssteiner, and Stefan Schmid. Load-optimization in reconfigurable networks: Algorithms and complexity of flow routing. ACM SIGMET-RICS Performance Evaluation Review, 48(3):39–44, December 2020. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3453953.
3453962.

Vardoyan:2020:CRB

[3077] Gayane Vardoyan, Saikat Guha, Philippe Nain, and Don Towsley. On the capacity region of bipartite and tripartite entanglement switching. ACM SIGMETRICS Performance Evaluation Review, 48(3):45–50, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/10.
                                        1145/3453953.3453963. See erratum

[3208].

Panigrahy:2020:ASC

[3078] Nitish K. Panigrahy, Prithwish Basu, Don Towsley, Ananthram Swami, and Kin K. Leung. On the analysis of spatially constrained power of two choice policies. ACM SIGMETRICS Performance Evaluation Review, 48(3):51–56, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3453953.3453964.

Zhou:2020:AOLb

[3079] Xingyu Zhou, Ness Shroff, and Adam Wierman. Asymptotically optimal load balancing in large-scale heterogeneous systems with multiple dispatchers. ACM SIGMETRICS Performance Evaluation Review, 48(3):57–58, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3453953.3453965.

Niu:2020:IAB

[3080] Jianyu Niu, Ziyu Wang, Fangyu Gai, and Chen Feng. Incentive analysis of Bitcoin-NG, revisited. ACM SIGMETRICS Performance Evaluation Review, 48(3):59–60, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3453953.3453966.

Vaze:2020:NSS

[3081] Rahul Vaze and Jayakrishnan Nair. Network speed scaling. ACM SIG-METRICS Performance Evaluation Review, 48(3):61–62, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3453953.
3453967.

Pokhrel:2020:RSF

[3082] Shiva Raj Pokhrel and Carey Williamson.

A rent-seeking framework for multipath TCP. ACM SIGMETRICS Performance Evaluation Review, 48(3):63– 70, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3453953.3453968.

Meng:2020:SWQ

[3083] Jingfan Meng, Long Gong, and Jun (Jim) Xu. Sliding-window QPS (SW-QPS): a perfect parallel iterative switching algorithm for input-queued switches. ACM SIGMETRICS Performance Evaluation Review, 48(3):71–76, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453953.3453969.

Quan:2020:PCM

[3084] Guocong Quan, Atilla Eryilmaz, Jian Tan, and Ness Shroff. Prefetching and caching for minimizing service costs: Optimal and approximation strategies. ACM SIGMETRICS Performance Evaluation Review, 48(3):77–78, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453953.3453970.

Vardoyan:2020:EAI

[3085] Gayane Vardoyan, Saikat Guha, Philippe Nain, and Don Towsley. On the exact analysis of an idealized quantum switch. ACM SIGMETRICS Performance Evaluation Review, 48(3):79– 80, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3453953.3453971.

Gilman:2020:DPP

[3086] Guin Gilman, Samuel S. Ogden, Tian Guo, and Robert J. Walls. Demystifying the placement policies of the NVIDIA GPU thread block scheduler for concurrent kernels. ACM SIGMETRICS Performance Evaluation Review, 48(3):81–88, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453953.3453972.

Liu:2020:FCN

[3087] Yuezhou Liu, Yuanyuan Li, Qian Ma, Stratis Ioannidis, and Edmund Yeh. Fair caching networks. ACM SIGMETRICS Performance Evaluation Review, 48(3):89–90, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453953.3453973.

Shioda:2020:DCB

[3088] Shigeo Shioda. Distribution of consensus in a broadcasting-based consensusforming algorithm. ACM SIGMET-RICS Performance Evaluation Review, 48(3):91–96, December 2020. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3453953.
3453974.

Elahi:2020:FSM

[3089] Maryam Elahi, Andrea Marin, Sabina Rossi, and Carey Williamson. Frequency scaling in multilevel queues. ACM SIGMETRICS Performance Evaluation Review, 48(3):97–98, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3453953.3453975.

Bienkowski:2020:ODB

[3090] Marcin Bienkowski, David Fuchssteiner, Jan Marcinkowski, and Stefan Schmid. Online dynamic B-matching: With applications to reconfigurable datacenter networks. ACM SIGMETRICS Performance Evaluation Review, 48(3):99– 108, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3453953.3453976.

Huang:2020:HTA

[3091] Yu Huang and Longbo Huang. Heavy traffic analysis of approximate maxweight matching algorithms for inputqueued switches. ACM SIGMETRICS Performance Evaluation Review, 48(3): 109–110, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3453953.3453977.

Malik:2020:RSI

[3092] Fehmina Malik, Manjesh K. Hanawal, Yezekael Hayel, and Jayakrishnan Nair. Revenue sharing on the Internet: a case for going soft on neutrality regulations. ACM SIGMETRICS Performance Evaluation Review, 48(3):111– 112, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3453953.3453978.

Pourghassemi:2020:ORS

[3093] Behnam Pourghassemi, Ardalan Amiri Sani, and Aparna Chandramowlishwaran. Only relative speed matters: Virtual causal profiling. ACM SIGMETRICS Performance Evaluation Review, 48(3):113–119, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3453953.3453979.

Kim:2020:RIP

[3094] Myungsuk Kim, Myoungjun Chun, Duwon Hong, Yoona Kim, Geonhee Cho, Dusol Lee, and Jihong Kim. Real-Wear: Improving performance and lifetime of SSDs using a NAND aging marker. ACM SIGMETRICS Performance Evaluation Review, 48(3):120– 121, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3453953.3453980.

Lu:2020:OCF

[3095] Yingdong Lu, Mark S. Squillante, and Tonghoon Suk. Optimal control in fluid models of n × n input-queued switches under linear fluid-flow costs. ACM SIGMETRICS Performance Evaluation Review, 48(3):122–127, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3453953.3453981.

Kar:2020:TOL

[3096] Sounak Kar, Robin Rehrmann, Arpan Mukhopadhyay, Bastian Alt, Florin Ciucu, Heinz Koeppl, Carsten Binnig, and Amr Rizk. On the throughput optimization in large-scale batchprocessing systems. ACM SIGMET-RICS Performance Evaluation Review, 48(3):128–129, December 2020. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3453953.
3453982.

Bayat:2020:ZRN

[3097] Niloofar Bayat, Richard Ma, Vishal Misra, and Dan Rubenstein. Zerorating and net neutrality: Who wins, who loses? ACM SIGMETRICS Performance Evaluation Review, 48(3): 130–135, December 2020. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3453953.3453983.

Knottenbelt:2021:MC

[3098] William Knottenbelt and Katinka Wolter. Message from the chairs. ACM SIGMETRICS Performance Evaluation Review, 48(4):2, May 2021. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3466826.
3466828.

Weber:2021:KAD

[3099] Ingo Weber. Keynote: Analysing data from blockchains. ACM SIGMETRICS Performance Evaluation Review, 48(4): 3, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3466826.3466829.

Coutinho:2021:CHI

[3100] Felipe Ribas Coutinho, Victor Pires, Claudio Miceli, and Daniel S. Menasche.

Crypto-hotwire: Illegal blockchain mining at zero cost using public infrastructures. ACM SIGMETRICS Performance Evaluation Review, 48(4):4– 7, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3466826.3466830.

Simoes:2021:BPT

[3101] Jefferson E. Simoes, Eduardo Ferreira, Daniel S. Menasché, and Carlos A. V. Campos. Blockchain privacy through merge avoidance and mixing services: a hardness and an impossibility result. ACM SIGMETRICS Performance Evaluation Review, 48(4):8–11, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3466826.3466831.

Oliveira:2021:ATC

[3102] Vinicius C. Oliveira, Julia Almeida Valadares, Jose Eduardo A. Sousa, Alex Borges Vieira, Heder Soares Bernardino, Saulo Moraes Villela, and Glauber Dias Goncalves. Analyzing transaction confirmation in Ethereum using machine learning techniques. ACM SIGMETRICS Performance Evaluation Review, 48(4):12–15, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3466826.3466832.

Gundlach:2021:PCT

[3103] Rowel Gundlach, Martijn Gijsbers, David Koops, and Jacques Resing. Predicting confirmation times of bitcoin transactions. ACM SIGMETRICS Performance Evaluation Review, 48(4): 16–19, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3466826.3466833.

Stoepker:2021:RAB

[3104] Ivo Stoepker, Rowel Gundlach, and Stella Kapodistria. Robustness analysis of bitcoin confirmation times. ACM SIGMETRICS Performance Evaluation Review, 48(4):20–23, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3466826.
3466834.

Sousa:2021:FUP

[3105] Jose Eduardo A. Sousa, Vinicius C.

Oliveira, Julia Almeida Valadares, Alex Borges Vieira, Heder S. Bernardino, Saulo Moraes Villela, and Glauber Dias Goncalves. Fighting under-price DoS attack in Ethereum with machine learning techniques. ACM SIGMET-RICS Performance Evaluation Review, 48(4):24–27, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3466826.3466835.

Vassio:2021:MOW

[3106] Luca Vassio, Zhi-Li Zhang, Danilo Giordano, and Abhishek Chandra. Message from the organizers of WAIN. ACM SIGMETRICS Performance Evaluation Review, 48(4):28, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3466826.3466837.

Hossen:2021:MTO

[3107] Md Rajib Hossen and Mohammad A.

Islam. Mobile task offloading under unreliable edge performance. ACM SIGMETRICS Performance Evaluation Review, 48(4):29–32, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3466826.
3466838.

Celenk:2021:MLB

[3108] Özge Celenk, Thomas Bauschert, and Marcus Eckert. Machine learning based KPI monitoring of video streaming traffic for QoE estimation. ACM SIGMETRICS Performance Evaluation Review, 48(4):33–36, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3466826.
3466839.

Wehner:2021:IWQ

[3109] Nikolas Wehner, Michael Seufert, Joshua Schuler, Sarah Wassermann, Pedro Casas, and Tobias Hossfeld. Improving Web QoE monitoring for encrypted network traffic through time series modeling. ACM SIGMETRICS Performance Evaluation Review, 48(4): 37–40, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3466826.3466840.

Markudova:2021:WMA

[3110] Dena Markudova, Martino Trevisan, Paolo Garza, Michela Meo, Maurizio M. Munafo, and Giovanna Carofiglio. What’s my app?: MLbased classification of RTC applications. ACM SIGMETRICS Performance Evaluation Review, 48(4):41–44, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3466826.3466841.

Higuchi:2021:FLP

[3111] Shunsuke Higuchi, Junji Takemasa, Yuki Koizumi, Atsushi Tagami, and Toru Hasegawa. Feasibility of Longest Prefix Matching using Learned Index Structures. ACM SIGMETRICS Performance Evaluation Review, 48(4):45– 48, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3466826.3466842.

Gonzalez:2021:UGM

[3112] Gastón García González, Pedro Casas, Alicia Fernández, and Gabriel Gómez. On the usage of generative models for network anomaly detection in multivariate time-series. ACM SIGMET-RICS Performance Evaluation Review, 48(4):49–52, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3466826.3466843.

Marin:2021:CCQ

[3113] Andrea Marin and Carey Williamson.

Cheating at craps: a quantitative analysis. ACM SIGMETRICS Performance Evaluation Review, 48(4):53–61, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3466826.3466845.

Masetti:2021:EMM

[3114] Giulio Masetti, Silvano Chiaradonna, Felicita Di Giandomenico, William H. Sanders, and Brett Feddersen. Extending the Möbius modeling environment with the advanced replication operator. ACM SIGMETRICS Performance Evaluation Review, 48(4):62–67, May 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3466826.3466846.

Arora:2021:OBB

[3115] Sanjeev Arora. Opening the Black Box of deep learning: Some lessons and take-aways. ACM SIGMETRICS Performance Evaluation Review, 49(1): 1, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3453910.

Zhu:2021:FBG

[3116] Zhaowei Zhu, Jingxuan Zhu, Ji Liu, and Yang Liu. Federated bandit: a gossiping approach. ACM SIGMETRICS Performance Evaluation Review, 49(1): 3–4, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3453919.

Cuvelier:2021:SEP

[3117] Thibaut Cuvelier, Richard Combes, and Eric Gourdin. Statistically efficient, polynomial-time algorithms for combinatorial semi-bandits. ACM SIG-METRICS Performance Evaluation Review, 49(1):5–6, June 2021. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3543516.
3453926.

Li:2021:IAC

[3118] Tongxin Li, Yue Chen, Bo Sun, Adam Wierman, and Steven Low. Information aggregation for constrained online control. ACM SIGMETRICS Performance Evaluation Review, 49(1):7– 8, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3461737.

Buchbinder:2021:OVM

[3119] Niv Buchbinder, Yaron Fairstein, Konstantina Mellou, Ishai Menache, and Joseph (Seffi) Naor. Online virtual machine allocation with lifetime and load predictions. ACM SIGMETRICS Performance Evaluation Review, 49(1):9– 10, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3456278.

Grosof:2021:NSI

[3120] Isaac Grosof, Kunhe Yang, Ziv Scully, and Mor Harchol-Balter. Nudge: Stochastically improving upon FCFS. ACM SIGMETRICS Performance Evaluation Review, 49(1):11–12, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3543516.3460102.

Wang:2021:ZQM

[3121] Weina Wang, Qiaomin Xie, and Mor Harchol-Balter. Zero queueing for multi-server jobs. ACM SIGMETRICS Performance Evaluation Review, 49(1): 13–14, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3453924.

Scully:2021:GPN

[3122] Ziv Scully, Isaac Grosof, and Mor Harchol-Balter. The Gittins Policy is nearly optimal in the M/G/k under extremely general conditions. ACM SIGMETRICS Performance Evaluation Review, 49(1):15–16, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3543516.
3456281.

Bijlani:2021:WDM

[3123] Ashish Bijlani, Umakishore Ramachandran, and Roy Campbell. Where did my 256 GB go? A measurement analysis of storage consumption on smart mobile devices. ACM SIGMETRICS Performance Evaluation Review, 49(1): 17–18, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3460108.

Zhang:2021:MSW

[3124] Yue Zhang, Bayan Turkistani, Allen Yuqing Yang, Chaoshun Zuo, and Zhiqiang Lin. A measurement study of Wechat mini-apps. ACM SIGMETRICS Performance Evaluation Review, 49(1):19– 20, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3460106.

Singh:2021:PNP

[3125] Rachee Singh, David Tench, Phillipa Gill, and Andrew McGregor. PredictRoute: a network path prediction toolkit. ACM SIGMETRICS Performance Evaluation Review, 49(1):21–22, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3460107.

Akbari:2021:LBC

[3126] Iman Akbari, Mohammad A. Salahuddin, Leni Ven, Noura Limam, Raouf Boutaba, Bertrand Mathieu, Stephanie Moteau, and Stephane Tuffin. A look behind the curtain: Traffic classification in an increasingly encrypted Web. ACM SIGMETRICS Performance Evaluation Review, 49(1):23–24, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3453921.

Weng:2021:AZA

[3127] Wentao Weng and Weina Wang.

Achieving zero asymptotic queueing delay for parallel jobs. ACM SIG-METRICS Performance Evaluation Review, 49(1):25–26, June 2021. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3543516.
3456268.

Raaijmakers:2021:ASR

[3128] Youri Raaijmakers and Sem Borst.

Achievable stability in redundancy systems. ACM SIGMETRICS Performance Evaluation Review, 49(1):27–28, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3456267.

Kielanski:2021:AIS

[3129] Grzegorz Kielanski and Benny Van Houdt. On the asymptotic insensitivity of the supermarket model in processor sharing systems. ACM SIGMETRICS Performance Evaluation Review, 49(1): 29–30, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3460100.

Randone:2021:RMF

[3130] Francesca Randone, Luca Bortolussi, and Mirco Tribastone. Refining meanfield approximations by dynamic state truncation. ACM SIGMETRICS Performance Evaluation Review, 49(1):31– 32, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3460099.

Gao:2021:TCC

[3131] Bingyu Gao, Haoyu Wang, Pengcheng Xia, Siwei Wu, Yajin Zhou, Xiapu Luo, and Gareth Tyson. Tracking counterfeit cryptocurrency end-to-end. ACM SIGMETRICS Performance Evaluation Review, 49(1):33–34, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3456282.

Chen:2021:SDC

[3132] Weimin Chen, Xinran Li, Yuting Sui, Ningyu He, Haoyu Wang, Lei Wu, and Xiapu Luo. SADPonzi: Detecting and characterizing Ponzi schemes in Ethereum smart contracts. ACM SIGMETRICS Performance Evaluation Review, 49(1):35–36, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3543516.
3460105.

Pourghassemi:2021:ACP

[3133] Behnam Pourghassemi, Jordan Bonecutter, Zhou Li, and Aparna Chandramowlishwaran. adPerf: Characterizing the performance of third-party ads. ACM SIGMETRICS Performance Evaluation Review, 49(1):37–38, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3453920.

Hon:2021:ASI

[3134] Hsiao-Wuen Hon. AI for system — infusing AI into cloud computing systems. ACM SIGMETRICS Performance Evaluation Review, 49(1):39–40, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3453911.

Zhang:2021:SFI

[3135] Qingzhao Zhang, David Ke Hong, Ze Zhang, Qi Alfred Chen, Scott Mahlke, and Z. Morley Mao. A systematic framework to identify violations of scenario-dependent driving rules in autonomous vehicle software. ACM SIGMETRICS Performance Evaluation Review, 49(1):43–44, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3460101.

Yang:2021:SSG

[3136] Lishan Yang, Bin Nie, Adwait Jog, and Evgenia Smirni. SUGAR: Speeding up GPGPU application resilience estimation with input sizing. ACM SIGMET-RICS Performance Evaluation Review, 49(1):45–46, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3543516.3453917.

Tang:2021:MMR

[3137] Xulong Tang, Mahmut Taylan Kandemir, and Mustafa Karakoy. Mix and match: Reorganizing tasks for enhancing data locality. ACM SIGMETRICS Performance Evaluation Review, 49(1): 47–48, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3543516.3460103.

Weng:2021:OLB

[3138] Wentao Weng, Xingyu Zhou, and R. Srikant. Optimal load balancing with locality constraints. ACM SIGMETRICS Performance Evaluation Review, 49(1):49–50, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3543516.
                                        3456279.

Rutten:2021:LBU

[3139] Daan Rutten and Debankur Mukherjee. Load balancing under strict compatibility constraints. ACM SIGMET-RICS Performance Evaluation Review, 49(1):51–52, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3543516.3456275.

Hellemans:2021:MWT

[3140] Tim Hellemans and Benny Van Houdt. Mean waiting time in large-scale and critically loaded power of d load balancing systems. ACM SIGMETRICS Performance Evaluation Review, 49(1): 53–54, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3543516.3460097.

Anton:2021:IPH

[3141] Elene Anton, Urtzi Ayesta, Matthieu Jonckheere, and Ina Maria Verloop. Improving the performance of heterogeneous data centers through redundancy. ACM SIGMETRICS Performance Evaluation Review, 49(1):55–56, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3456274.

Wang:2021:PPA

[3142] Xin Wang and Richard T. B. Ma. On private peering agreements between content and access providers: a contractual equilibrium analysis. ACM SIGMETRICS Performance Evaluation Review, 49(1):57–58, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3543516.
3456277.

Fleder:2021:KWY

[3143] Michael Fleder and Devavrat Shah. I know what you bought at Chipotle for $9.81 by solving a linear inverse problem. ACM SIGMETRICS Performance Evaluation Review, 49(1):59– 60, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3456273.

Varma:2021:DPM

[3144] Sushil Mahavir Varma, Francisco Castro, and Siva Theja Maguluri. Dynamic pricing and matching for two-sided markets with strategic servers. ACM SIGMETRICS Performance Evaluation Review, 49(1):61–62, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3543516.
3456272.

Tang:2021:RMG

[3145] Jing Tang, Xueyan Tang, Andrew Lim, Kai Han, Chongshou Li, and Junsong Yuan. Revisiting modified greedy algorithm for monotone submodular maximization with a knapsack constraint. ACM SIGMETRICS Performance Evaluation Review, 49(1):63–64, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3453925.

Han:2021:AAS

[3146] Kai Han, Shuang Cui, Tianshuai Zhu, Enpei Zhang, Benwei Wu, Zhizhuo Yin, Tong Xu, Shaojie Tang, and He Huang. Approximation algorithms for submodular data summarization with a knapsack constraint. ACM SIGMETRICS Performance Evaluation Review, 49(1): 65–66, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3453922.

Sun:2021:CAO

[3147] Bo Sun, Ali Zeynali, Tongxin Li, Mohammad Hajiesmaili, Adam Wierman, and Danny H. K. Tsang. Competitive algorithms for the online multiple knapsack problem with application to electric vehicle charging. ACM SIGMETRICS Performance Evaluation Review, 49(1):67–68, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3456271.

Tassiulas:2021:EIS

[3148] Leandros Tassiulas. Enabling intelligent services at the network edge. ACM SIGMETRICS Performance Evaluation Review, 49(1):69–70, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3543516.
3453912.

Foerster:2021:IDD

[3149] Klaus-Tycho Foerster, Janne H. Korhonen, Ami Paz, Joel Rybicki, and Stefan Schmid. Input-dynamic distributed algorithms for communication networks. ACM SIGMETRICS Performance Evaluation Review, 49(1):71– 72, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3453923.

Perivier:2021:RTA

[3150] Noémie Périvier, Chamsi Hssaine, Samitha Samaranayake, and Siddhartha Banerjee. Real-time approximate routing for smart transit systems. ACM SIGMETRICS Performance Evaluation Review, 49(1):73–74, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3543516.3460096.

Asgari:2021:BSO

[3151] Kamiar Asgari and Michael J. Neely.

Bregman-style online convex optimization with energy harvesting constraints. ACM SIGMETRICS Performance Evaluation Review, 49(1):75–76, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/10.
                                        1145/3543516.3456270.

Yu:2021:PDH

[3152] Liren Yu, Jiaming Xu, and Xiaojun Lin. The power of D-hops in matching power-law graphs. ACM SIGMET-RICS Performance Evaluation Review, 49(1):77–78, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3543516.3460098.

Zhang:2021:CHE

[3153] Yiguang Zhang, Jessy Xinyi Han, Ilica Mahajan, Priyanjana Bengani, and Augustin Chaintreau. Chasm in hegemony: Explaining and reproducing disparities in homophilous networks. ACM SIGMETRICS Performance Evaluation Review, 49(1):79–80, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3543516.3460109.

Hazimeh:2021:MGT

[3154] Ahmad Hazimeh, Adrian Herrera, and Mathias Payer. Magma: a groundtruth fuzzing benchmark. ACM SIG-METRICS Performance Evaluation Review, 49(1):81–82, June 2021. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3543516.
                                        3456276.

Abanto-Leon:2021:SCL

[3155] Luis F. Abanto-Leon, Andreas Bäuml, Gek Hong (Allyson) Sim, Matthias Hollick, and Arash Asadi. Stay connected, leave no trace: Enhancing security and privacy in WiFi via obfuscating radiometric fingerprints. ACM SIGMETRICS Performance Evaluation Review, 49(1):83–84, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543516.3456280.

Squillante:2021:SIW

[3156] Mark S. Squillante. Special issue on The Workshop on MAthematical performance Modeling and Analysis (MAMA 2021). ACM SIGMETRICS Performance Evaluation Review, 49(2): 2, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512798.3512800.

Su:2021:LAE

[3157] Yu Su, Jannie Yu, Vivek Anand, and Adam Wierman. Learningaugmented energy-aware scheduling of precedence-constrained tasks. ACM SIGMETRICS Performance Evaluation Review, 49(2):3–5, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512798.3512801.

Liao:2021:PPO

[3158] Guocheng Liao, Yu Su, Juba Ziani, Adam Wierman, and Jianwei Huang. The privacy paradox and optimal biasvariance trade-offs in data acquisition. ACM SIGMETRICS Performance Evaluation Review, 49(2):6–8, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512798.3512802.

Tavori:2021:CVO

[3159] Jhonatan Tavori and Hanoch Levy. Continual versus occasional spreading in networks: Modeling spreading thresholds in epidemic processes. ACM SIGMETRICS Performance Evaluation Review, 49(2):9–11, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3512798.3512803.

Ancel:2021:MFA

[3160] Julien Ancel, Christine Fricker, and Hanene Mohamed. Mean field analysis for bike and e-bike sharing systems. ACM SIGMETRICS Performance Evaluation Review, 49(2):12–14, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3512798.3512804.

Miguelez:2021:RSG

[3161] Fernando Miguélez, Josu Doncel, and Urtzi Ayesta. A resource sharing game for the freshness of status updates. ACM SIGMETRICS Performance Evaluation Review, 49(2):15–17, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3512798.3512805.

Scully:2021:WDG

[3162] Ziv Scully and Lucas van Kreveld. When does the Gittins policy have asymptotically optimal response time tail? ACM SIGMETRICS Performance Evaluation Review, 49(2):18–20, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512798.3512806.

Lu:2021:PAQ

[3163] Yingdong Lu. Performance analysis of a queueing system with server arrival and departure. ACM SIGMETRICS Performance Evaluation Review, 49(2): 21–23, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3512798.3512807.

Rutten:2021:CSA

[3164] Daan Rutten and Debankur Mukherjee. Capacity scaling augmented with unreliable machine learning predictions. ACM SIGMETRICS Performance Evaluation Review, 49(2):24–26, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512798.3512808.

Ferragut:2021:EVC

[3165] Andres Ferragut, Lucas Narbondo, and Fernando Paganini. EDF vehicle charging under deadline uncertainty. ACM SIGMETRICS Performance Evaluation Review, 49(2):27–29, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512798.3512809.

Bayat:2021:REA

[3166] Niloofar Bayat, Cody Morrin, Yuheng Wang, and Vishal Misra. Rank estimation for (approximately) low-rank matrices. ACM SIGMETRICS Performance Evaluation Review, 49(2):30–32, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3512798.3512810.

Ramtin:2021:CDA

[3167] Amir Reza Ramtin, Don Towsley, Philippe Nain, Edmundo de Souza e Silva, and Daniel S. Menasche. Are covert DDoS attacks facing multifeature detectors feasible? ACM SIGMETRICS Performance Evaluation Review, 49(2):33–35, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3512798.3512811.

Scully:2021:BMS

[3168] Ziv Scully. Bounding mean slowdown in multiserver systems. ACM SIGMETRICS Performance Evaluation Review, 49(2):36–38, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3512798.3512812.

Ghosh:2021:UGE

[3169] Soumyadip Ghosh and Mark S. Squillante. Unbiased gradient estimation for robust optimization. ACM SIGMETRICS Performance Evaluation Review, 49(2):39–41, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3512798.3512813.

Gast:2021:SIW

[3170] Nicolas Gast and Neil Walton. Special issue on the workshop about reinforcement learning in networks and queues (RLNQ 2021). ACM SIGMETRICS Performance Evaluation Review, 49(2):

42, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3512798.3512814.

Tessler:2021:RLD

[3171] Chen Tessler, Yuval Shpigelman, Gal Dalal, Amit Mandelbaum, Doron Haritan Kazakov, Benjamin Fuhrer, Gal Chechik, and Shie Mannor. Reinforcement learning for Datacenter Congestion Control. ACM SIGMETRICS Performance Evaluation Review, 49(2): 43–46, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3512798.3512815.

Robledo:2021:QQL

[3172] Francisco Robledo, Vivek Borkar, Urtzi Ayesta, and Konstantin Avrachenkov. QWI: Q-learning with Whittle Index. ACM SIGMETRICS Performance Evaluation Review, 49(2):47–50, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512798.3512816.

Newton:2021:AOD

[3173] Conor Newton, Ayalvadi Ganesh, and Henry Reeve. Asymptotic optimality for decentralised bandits. ACM SIGMETRICS Performance Evaluation Review, 49(2):51–53, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512798.3512817.

Liu:2021:CBA

[3174] Xin Liu, Bin Li, Pengyi Shi, and Lei Ying. A constrained bandit approach for online dispatching. ACM SIGMETRICS Performance Evaluation Review, 49(2):54–56, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3512798.3512818.

Archer:2021:OBS

[3175] Christopher Archer, Siddhartha Banerjee, Mayleen Cortez, Carrie Rucker, Sean R. Sinclair, Max Solberg, Qiaomin Xie, and Christina Lee Yu. ORSuite: Benchmarking suite for sequential operations models. ACM SIG-METRICS Performance Evaluation Review, 49(2):57–61, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3512798.
                                        3512819.

Fanti:2021:ASS

[3176] Giulia Fanti. ACM SIGMETRICS 2021 Student Research Competition. ACM SIGMETRICS Performance Evaluation Review, 49(2):62, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3512798.
                                        3512820.

Tuli:2021:SIP

[3177] Shreshth Tuli. SplitPlace: Intelligent placement of split neural nets in mobile edge environments. ACM SIGMETRICS Performance Evaluation Review, 49(2):63–65, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3512798.3512821.

Hong:2021:SZQ

[3178] Yige Hong. Sharp zero-queueing bounds for multi-server jobs. ACM SIGMETRICS Performance Evaluation Review, 49(2):66–68, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512798.3512822.

Ghasemi:2021:ASA

[3179] Mahshid Ghasemi. Auto-SDA: Automated video-based social distancing analyzer. ACM SIGMETRICS Performance Evaluation Review, 49(2):69–71, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512798.3512823.

Peng:2021:ERT

[3180] Edwin Peng. Exact response time analysis of preemptive priority scheduling with switching overhead. ACM SIGMETRICS Performance Evaluation Review, 49(2):72–74, September 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3512798.3512824.

Liu:2021:SAS

[3181] Qingsong Liu, Wenfei Wu, Longbo Huang, and Zhixuan Fang. Simultaneously achieving sublinear regret and constraint violations for online convex optimization with time-varying constraints. ACM SIGMETRICS Performance Evaluation Review, 49(3):4–5, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529114.

Lee:2021:CBS

[3182] Russell Lee, Yutao Zhou, Lin Yang, Mohammad Hajiesmaili, and Ramesh Sitaraman. Competitive bidding strategies for online linear optimization with inventory management constraints. ACM SIGMETRICS Performance Evaluation Review, 49(3):6–7, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3529113.3529115.

Song:2021:OLH

[3183] Jianhan Song, Gustavo de Veciana, and Sanjay Shakkottai. Online learning for hierarchical scheduling to support network slicing in cellular networks. ACM SIGMETRICS Performance Evaluation Review, 49(3):8–9, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3529113.3529116.

Ferragut:2021:SEC

[3184] Andres Ferragut, Lucas Narbondo, and Fernando Paganini. Scheduling EV charging with uncertain departure times. ACM SIGMETRICS Performance Evaluation Review, 49(3):10–15, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3529113.3529117.

Scherrer:2021:APP

[3185] Simon Scherrer, Markus Legner, Adrian Perrig, and Stefan Schmid. An axiomatic perspective on the performance effects of end-host path selection. ACM SIGMETRICS Performance Evaluation Review, 49(3):16–17, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529118.

Liu:2021:OCN

[3186] Bai Liu and Eytan Modiano. Optimal control for networks with unobservable malicious nodes. ACM SIGMETRICS Performance Evaluation Review, 49(3): 18–19, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3529113.3529119.

Ramtin:2021:FSL

[3187] Amir Reza Ramtin, Philippe Nain, Daniel S. Menasche, Don Towsley, and Edmundo de Souza e Silva. Fundamental scaling laws of covert DDoS attacks. ACM SIGMETRICS Performance Evaluation Review, 49(3):20–21, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529120.

Chaturvedi:2021:ITA

[3188] Anya Chaturvedi, Chandra Chekuri, Andréa W. Richa, Matthias Rost, Stefan Schmid, and Jamison Weber. Improved throughput for all-or-nothing multicommodity flows with arbitrary demands. ACM SIGMETRICS Performance Evaluation Review, 49(3):22–27, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529121.

Huang:2021:EFB

[3189] Xiandong Huang, Qinglin Wang, Shuyu Lu, Ruochen Hao, Songzhu Mei, and Jie Liu. Evaluating FFT-based algorithms for strided convolutions on ARMv8 architectures? ACM SIG-METRICS Performance Evaluation Review, 49(3):28–29, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3529113.
3529122.

Vardoyan:2021:QPE

[3190] Gayane Vardoyan, Matthew Skrzypczyk, and Stephanie Wehner. On the quantum performance evaluation of two distributed quantum architectures. ACM SIGMETRICS Performance Evaluation Review, 49(3):30–31, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3529113.3529123.

Gilman:2021:CCM

[3191] Guin Gilman and Robert J. Walls. Characterizing concurrency mechanisms for NVIDIA GPUs under deep learning workloads (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 49(3):32–34, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529124.

Pacut:2021:ISD

[3192] Maciej Pacut, Wenkai Dai, Alexandre Labbe, Klaus-Tycho Foerster, and Stefan Schmid. Improved scalability of demand-aware datacenter topologies with minimal route lengths and congestion. ACM SIGMETRICS Performance Evaluation Review, 49(3):35–36, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529125.

Cadas:2021:FCH

[3193] Arnaud Cadas, Josu Doncel, Jean-Michel Fourneau, and Ana Busic. Flexibility can hurt dynamic matching system performance. ACM SIGMETRICS Performance Evaluation Review, 49(3): 37–42, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3529113.3529126.

Varma:2021:HTT

[3194] Sushil Mahavir Varma and Siva Theja Maguluri. A heavy traffic theory of two-sided queues. ACM SIGMETRICS Performance Evaluation Review, 49(3): 43–44, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3529113.3529127.

Vaze:2021:SSM

[3195] Rahul Vaze and Jayakrishnan Nair.

Speed scaling with multiple servers under a sum–power constraint. ACM SIGMETRICS Performance Evaluation Review, 49(3):45–50, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529128.

Casale:2021:FLD

[3196] Giuliano Casale, Peter G. Harrison, and Wai Hong Ong. Facilitating load-dependent queueing analysis through factorization (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 49(3):51–52, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529129.

Zubeldia:2021:LTC

[3197] Martin Zubeldia and Michel Mandjes. Learning traffic correlations in multi-class queueing systems by sampling queue lengths, with routing applications. ACM SIGMETRICS Performance Evaluation Review, 49(3):53–54, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529130.

Spang:2021:UTB

[3198] Bruce Spang. Updating the theory of buffer sizing. ACM SIGMETRICS Performance Evaluation Review, 49(3): 55–56, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3529113.3529131.

Singhal:2021:CFR

[3199] Shiksha Singhal and Veeraruna Kavitha. Coalition formation resource sharing games in networks. ACM SIGMET-RICS Performance Evaluation Review, 49(3):57–58, December 2021. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3529113.
3529132.

Sivaraman:2021:ENT

[3200] Vibhaalakshmi Sivaraman, Weizhao Tang, Shaileshh Bojja Venkatakrishnan, Giulia Fanti, and Mohammad Alizadeh. The effect of network topology on credit network throughput. ACM SIGMETRICS Performance Evaluation Review, 49(3):59–60, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529133.

Chen:2021:ERC

[3201] Xusheng Chen, Shixiong Zhao, Ji Qi, Jianyu Jiang, Haoze Song, Cheng Wang, Tsz On Li, T-H. Hubert Chan, Fengwei Zhang, Xiapu Luo, Sen Wang, Gong Zhang, and Heming Cuih. Efficient and DoS-resistant consensus for permissioned blockchains. ACM SIGMETRICS Performance Evaluation Review, 49(3):61–62, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529134.

Jain:2021:SCM

[3202] Shubham Anand Jain, Shreyas Goenka, Divyam Bapna, Nikhil Karamchandani, and Jayakrishnan Nair. Sequential community mode estimation. ACM SIGMETRICS Performance Evaluation Review, 49(3):63–64, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529135.

Berg:2021:CPA

[3203] Benjamin Berg, Justin Whitehouse, Benjamin Moseley, Weina Wang, and Mor Harchol-Balter. The case for phase-aware scheduling of parallelizable jobs. ACM SIGMETRICS Performance Evaluation Review, 49(3):65–66, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529136.

Fu:2021:EJS

[3204] Xinzhe Fu and Eytan Modiano. Elastic job scheduling with unknown utility functions. ACM SIGMETRICS Performance Evaluation Review, 49(3):67–68, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3529113.3529137.

Ruuskanen:2021:IMF

[3205] Johan Ruuskanen, Tommi Berner, Karl-Erik Arzen, and Anton Cervin. Improving the mean-field fluid model of processor sharing queueing networks for dynamic performance models in cloud computing. ACM SIGMETRICS Performance Evaluation Review, 49(3): 69–70, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3529113.3529138.

Anselmi:2021:OSP

[3206] Jonatha Anselmi, Bruno Gaujal, and Louis-Sébastien Rebuffi. Optimal speed profile of a DVFS processor under soft deadlines. ACM SIGMETRICS Performance Evaluation Review, 49(3):71–72, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3529113.3529139.

Buchholz:2021:RCE

[3207] Peter Buchholz. On the representation of correlated exponential distributions by phase type distributions. ACM SIGMETRICS Performance Evaluation Review, 49(3):73–78, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529140.

Vardoyan:2021:CRB

[3208] Gayane Vardoyan, Saikat Guha, Philippe Nain, and Don Towsley. On the capacity region of bipartite and tripartite entanglement switching: Erratum. ACM SIGMETRICS Performance Evaluation Review, 49(3):79–80, December 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3529113.3529141. See

[3077].

Harchol-Balter:2022:MCQ

[3209] Mor Harchol-Balter and Ziv Scully.

The most common queueing theory questions asked by computer systems practitioners. ACM SIGMETRICS Performance Evaluation Review, 49(4): 3–7, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543146.3543148.

Tan:2022:VCA

[3210] Chee Wei Tan. The value of cooperation: From AIMD to flipped classroom teaching. ACM SIGMETRICS Performance Evaluation Review, 49(4):8–13, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543146.3543149.

Xia:2022:TPM

[3211] Cathy H. Xia, Nanshan Chen, and Priya Natarajan. Teaching performance modeling via software and instructional technology. ACM SIG-METRICS Performance Evaluation Review, 49(4):14–19, March 2022. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3543146.
                                        3543150.

Boudec:2022:PEP

[3212] Jean-Yves Le Boudec. Performance evaluation: a preparation for statistics and data science? ACM SIGMET-RICS Performance Evaluation Review, 49(4):20–23, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3543146.3543151.

Serazzi:2022:UCP

[3213] Giuseppe Serazzi. Updating the content of performance analysis textbooks. ACM SIGMETRICS Performance Evaluation Review, 49(4):24–27, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3543146.3543152.

Donatelli:2022:SIW

[3214] Susanna Donatelli and Giuliana Franceschinis. Special issue on the workshop on TOols for Stochastic Modeling and Evaluation (TOSME 2021). ACM SIGMETRICS Performance Evaluation Review, 49(4):28, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/10.
                                        1145/3543146.3543154.

Almousa:2022:CME

[3215] Salah Al-Deen Almousa, G´abor Horv´ath, Ill ´es Horv´ath, Andr´as M´esz´aros, and Mikl ´os Telek. The CME method: Efficient numerical inverse Laplace transformation with Concentrated Matrix Exponential distribution. ACM SIGMETRICS Performance Evaluation Review, 49(4):29–34, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3543146.3543155.

Allmeier:2022:RTL

[3216] Sebastian Allmeier and Nicolas Gast.

rmf tool — a library to compute (refined) mean field approximation(s). ACM SIGMETRICS Performance Evaluation Review, 49(4):35–40, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3543146.3543156.

Masetti:2022:TTS

[3217] Giulio Masetti, Leonardo Robol, Silvano Chiaradonna, and Felicita Di Giandomenico. TAPAS: a tool for stochastic evaluation of large interdependent composed models with absorbing states. ACM SIGMETRICS Performance Evaluation Review, 49(4):41– 46, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543146.3543157.

Marzolla:2022:QNM

[3218] Moreno Marzolla. Queueing networks and Markov chains analysis with the Octave queueing package. ACM SIG-METRICS Performance Evaluation Review, 49(4):47–52, March 2022. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3543146.
3543158.

Cortellessa:2022:SMR

[3219] Vittorio Cortellessa, Daniele Di Pompeo, Vincenzo Stoico, and Michele Tucci. Software model refactoring driven by performance antipattern detection. ACM SIGMETRICS Performance Evaluation Review, 49(4):53–58, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3543146.3543159.

Budde:2022:FFI

[3220] Carlos E. Budde. FIG: the Finite Improbability Generator v1.3. ACM SIGMETRICS Performance Evaluation Review, 49(4):59–64, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3543146.
                                        3543160.

Ballarini:2022:CES

[3221] Paolo Ballarini and Benoît Barbot. Cosmos: Evolution of a statistical model checking platform. ACM SIG-METRICS Performance Evaluation Review, 49(4):65–69, March 2022. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3543146.
                                        3543161.

Sheldon:2022:TSM

[3222] Matthew Sheldon and Giuliano Casale. TauSSA: Simulating Markovian queueing networks with Tau leaping. ACM SIGMETRICS Performance Evaluation Review, 49(4):70–75, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3543146.
                                        3543162.

Llado:2022:POP

[3223] Catalina M. Llado. PIPE 2.7 overview:

a Petri net tool for performance modeling and evaluation. ACM SIGMET-RICS Performance Evaluation Review, 49(4):76–80, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3543146.3543163.

Carnevali:2022:OTA

[3224] Laura Carnevali, Marco Paolieri, and Enrico Vicario. The ORIS tool: app, library, and toolkit for quantitative evaluation of non-Markovian systems. ACM SIGMETRICS Performance Evaluation Review, 49(4):81–86, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543146.3543164.

Amparore:2022:SME

[3225] Elvio G. Amparore. Stochastic modelling and evaluation using Great-SPN. ACM SIGMETRICS Performance Evaluation Review, 49(4):87–91, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543146.3543165.

Vassio:2022:MOO

[3226] Luca Vassio, Danilo Giordano, Jinoh Kim, and Jon Crowcroft. Message from the organizers of WAIN. ACM SIGMETRICS Performance Evaluation Review, 49(4):92, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3543146.
3543167.

Tao:2022:LAO

[3227] Shimin Tao, Weibin Meng, Yimeng Cheng, Yichen Zhu, Ying Liu, Chunning Du, Tao Han, Yongpeng Zhao, Xiangguang Wang, and Hao Yang. LogStamp: Automatic online log parsing based on sequence labelling. ACM SIGMETRICS Performance Evaluation Review, 49(4):93–98, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3543146.
3543168.

Hao:2022:IAA

[3228] Wenwen Hao, Ben Niu, Yin Luo, Kangkang Liu, and Na Liu. Improving accuracy and adaptability of SSD failure prediction in hyper-scale data centers. ACM SIGMETRICS Performance Evaluation Review, 49(4):99– 104, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543146.3543169.

Ketabi:2022:CAF

[3229] Shiva Ketabi, Matthew Buckley, Parsa Pazhooheshy, Faraz Farahvash, and Yashar Ganjali. Correlation-aware flow consolidation for load balancing and beyond. ACM SIGMETRICS Performance Evaluation Review, 49(4):105– 110, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543146.3543170.

Pujol-Perich:2022:UPG

[3230] David Pujol-Perich, Jose Suarez-Varela, Albert Cabellos-Aparicio, and Pere Barlet-Ros. Unveiling the potential of graph neural networks for robust intrusion detection. ACM SIG-METRICS Performance Evaluation Review, 49(4):111–117, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3543146.
3543171.

Bertoli:2022:IDS

[3231] Gustavo de Carvalho Bertoli, Lourenço Alves Pereira Júnior, and Osamu Saotome. Improving detection of scanning attacks on heterogeneous networks with Federated Learning. ACM SIGMETRICS Performance Evaluation Review, 49(4):118–123, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543146.3543172.

Barros:2022:UMN

[3232] Matheus F. C. Barros, Carlos H. G.

Ferreira, Lourenco A. P. Junior, Marco Mellia, Jussara M. Almeida, and Bruno Pereira dos Santos. Understanding mobility in networks: a node embedding approach. ACM SIGMETRICS Performance Evaluation Review, 49(4):124–130, March 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3543146.3543173.

Roberts:2022:SBS

[3233] James Roberts and Dario Rossi. Sizebased scheduling vs fairness for datacenter flows: a queuing perspective. ACM SIGMETRICS Performance Evaluation Review, 50(2):2–10, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3561074.3561076.

Squillante:2022:SIW

[3234] Mark S. Squillante. Special issue on The Workshop on MAthematical performance Modeling and Analysis (MAMA 2022). ACM SIGMETRICS Performance Evaluation Review, 50(2): 11, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3561074.3561078.

Shioda:2022:ETR

[3235] Shigeo Shioda and Kenta Takehara. Ergodicity of time reversal process of stochastic consensus formation and its application. ACM SIGMETRICS Performance Evaluation Review, 50(2):12– 14, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3561074.3561079.

Tavori:2022:QNB

[3236] Jhonatan Tavori and Hanoch Levy. Queueing-network based applications under worst-case attacks. ACM SIGMETRICS Performance Evaluation Review, 50(2):15–17, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3561074.3561080.

Kalvit:2022:DLL

[3237] Anand Kalvit and Assaf Zeevi. Dynamic learning in large matching markets. ACM SIGMETRICS Performance Evaluation Review, 50(2):18–20, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3561074.3561081.

Jiang:2022:DDO

[3238] Harry Jiang, Xiaoxi Zhang, and Carlee Joe-Wong. DOLL: Distributed On-Line Learning using preemptible cloud instances. ACM SIGMETRICS Performance Evaluation Review, 50(2):21–23, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3561074.3561082.

Yao:2022:STM

[3239] Yuan Yao, Marco Paolieri, and Leana Golubchik. Sojourn time minimization of successful jobs. ACM SIGMETRICS Performance Evaluation Review, 50(2): 24–26, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3561074.3561083.

Ye:2022:ORR

[3240] Heng-Qing Ye. Optimal round-robin routing to parallel servers in heavy traffic. ACM SIGMETRICS Performance Evaluation Review, 50(2):27–29, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3561074.3561084.

Anton:2022:SUR

[3241] E. Anton, R. Righter, and I. M.

Verloop. Scheduling under redundancy. ACM SIGMETRICS Performance Evaluation Review, 50(2):30–32, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3561074.3561085.

Tan:2022:OSC

[3242] Xiaoqi Tan, Siyuan Yu, and Raouf Boutaba. Online selection with convex costs. ACM SIGMETRICS Performance Evaluation Review, 50(2):33–35, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3561074.3561086.

Horvath:2022:ONI

[3243] Illes Horvath, Andras Meszaros, and Miklos Telek. Optimized numerical inverse Laplace transformation. ACM SIGMETRICS Performance Evaluation Review, 50(2):36–38, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3561074.3561087.

Ding:2022:COM

[3244] Rui Ding and Eugene Feinberg. CVaR optimization for MDPs: Existence and computation of optimal policies. ACM SIGMETRICS Performance Evaluation Review, 50(2):39–41, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3561074.3561088.

Kalantzis:2022:QAR

[3245] Vasileios Kalantzis, Mark S. Squillante, Shashanka Ubaru, and Lior Horesh. On quantum algorithms for random walks in the nonnegative quarter plane. ACM SIGMETRICS Performance Evaluation Review, 50(2):42–44, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3561074.3561089.

Kannan:2022:SIW

[3246] Pravein Govindan Kannan, Priyanka Naik, Praveen Tammana, and Mythili Vutukuru. Special issue on The Workshop on Performance of hostbased Network Applications (PerfNA 2022). ACM SIGMETRICS Performance Evaluation Review, 50(2):45, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3561074.3561091.

Wu:2022:GGD

[3247] Ziyan Wu, Tianming Cui, Arvind Narayanan, Yang Zhang, Kangjie Lu, Antonia Zhai, and Zhi-Li Zhang. GranularNF: Granular decomposition of stateful NFV at 100 Gbps line speed and beyond. ACM SIGMETRICS Performance Evaluation Review, 50(2):46– 51, September 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3561074.3561092.

Scully:2022:NTS

[3248] Ziv Scully. A new toolbox for scheduling theory. ACM SIGMETRICS Performance Evaluation Review, 50(3):3– 6, December 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3579342.3579344.

Agarwal:2022:CIS

[3249] Anish Agarwal. Causal inference for social and engineering systems. ACM SIGMETRICS Performance Evaluation Review, 50(3):7–11, December 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3579342.3579345.

Chen:2022:ULF

[3250] Zaiwei Chen. A unified Lyapunov framework for finite-sample analysis of reinforcement learning algorithms. ACM SIGMETRICS Performance Evaluation Review, 50(3):12–15, December 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3579342.3579346.

Tuli:2022:ACS

[3251] Shreshth Tuli. AI and co-simulation driven resource management in fog computing environments. ACM SIG-METRICS Performance Evaluation Review, 50(3):16–19, December 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3579342.
3579347.

Salem:2022:OLN

[3252] Tareq Si Salem. Online learning for network resource allocation. ACM SIGMETRICS Performance Evaluation Review, 50(3):20–23, December 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3579342.3579348.

Maghakian:2022:ORA

[3253] Jessica Maghakian. Online resource allocation with noisy predictions. ACM SIGMETRICS Performance Evaluation Review, 50(3):24–27, December 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3579342.3579349.

UlGias:2022:MBR

[3254] Alim Ul Gias. Model-based resource management for fine-grained services. ACM SIGMETRICS Performance Evaluation Review, 50(3):28–31, December 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3579342.3579350.

Sun:2022:RAS

[3255] Xiao Sun. Resource allocation and scheduling in modern cloud computing. ACM SIGMETRICS Performance Evaluation Review, 50(3):32–35, December 2022. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3579342.3579351.

Shang:2022:EDI

[3256] Xiaojun Shang. Enabling dataintensive workflows in heterogeneous edge-cloud networks. ACM SIGMET-RICS Performance Evaluation Review, 50(3):36–38, December 2022. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3579342.
3579352.

Marin:2023:PFN

[3257] Andrea Marin, Sabina Rossi, and Diletta Olliaro. A product-form network for systems with job stealing policies. ACM SIGMETRICS Performance Evaluation Review, 50(4):2–4, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595246.

Kumar:2023:ADC

[3258] Nitesh Kumar, Gaurav S. Kasbekar, and D. Manjunath. Application of data collected by endpoint detection and response systems for implementation of a network security system based on zero trust principles and the EigenTrust algorithm. ACM SIGMETRICS Performance Evaluation Review, 50(4):5–7, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595247.

Jinan:2023:AAP

[3259] Rooji Jinan, Gaurav Gautam, Parimal Parag, and Vaneet Aggarwal. Asymptotic analysis of probabilistic scheduling for erasure-coded heterogeneous systems. ACM SIGMETRICS Performance Evaluation Review, 50(4):8–10, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595248.

Hwang:2023:AEM

[3260] Wonjun Hwang, Yoora Kim, and Kyunghan Lee. Augmenting epidemic models with graph neural networks. ACM SIGMETRICS Performance Evaluation Review, 50(4):11–13, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595249.

Biswas:2023:EDT

[3261] Sudeshna Biswas, Himanshu, Sushmita Ghosh, Payali Das, Kaushik Saha, and Swades De. Efficient data transfer mechanism for DLMS/COSEM enabled smart energy metering platform. ACM SIGMETRICS Performance Evaluation Review, 50(4):14–16, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595250.

Umrawal:2023:LCS

[3262] Abhishek K. Umrawal and Vaneet Aggarwal. Leveraging the community structure of a social network for maximizing the spread of influence. ACM SIGMETRICS Performance Evaluation Review, 50(4):17–19, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595251.

Gupta:2023:CCB

[3263] Samarth Gupta, Jinhang Zuo, Carlee Joe-Wong, Gauri Joshi, and Osman Yagan. Correlated combinatorial bandits for online resource allocation. ACM SIGMETRICS Performance Evaluation Review, 50(4):20–22, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595252.

Talukder:2023:EFL

[3264] Zahidur Talukder and Mohammad A.

Islam. Efficient federated learning with self-regulating clients. ACM SIGMET-RICS Performance Evaluation Review, 50(4):23–25, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3595244.3595253.

Jhunjhunwala:2023:HTQ

[3265] Prakirt Raj Jhunjhunwala and Siva Theja Maguluri. Heavy traffic queue length distribution without resource pooling in an input-queued switch. ACM SIGMETRICS Performance Evaluation Review, 50(4):26–28, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3595244.
3595254.

Tuli:2023:LDS

[3266] Shreshth Tuli, Giuliano Casale, and Nicholas R. Jennings. Learning to dynamically select cost optimal schedulers in cloud computing environments. ACM SIGMETRICS Performance Evaluation Review, 50(4):29–31, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595255.

Saurav:2023:MAI

[3267] Kumar Saurav and Rahul Vaze. Minimizing age of information under arbitrary arrival model with arbitrary packet size. ACM SIGMETRICS Performance Evaluation Review, 50(4):32– 34, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595256.

Prakash:2023:ROE

[3268] R Sri Prakash, Nikhil Karamchandani, and Sharayu Moharir. On the regret of online edge service hosting. ACM SIGMETRICS Performance Evaluation Review, 50(4):35–37, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595257.

Reiffers-Masson:2023:OMA

[3269] Alexandre Reiffers-Masson and Isabel Amigo. Online multi-agent decentralized Byzantine-robust gradient estimation. ACM SIGMETRICS Performance Evaluation Review, 50(4):38–40, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595258.

Sam:2023:OLHa

[3270] Tyler Sam, Yudong Chen, and Christina Lee Yu. Overcoming the long horizon barrier for sample-efficient reinforcement learning with latent lowrank structure. ACM SIGMETRICS Performance Evaluation Review, 50(4): 41–43, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3595244.3595259.

Li:2023:PPD

[3271] Qi Li and Dong Chen. Peer to peer distributed solar energy trading. ACM SIGMETRICS Performance Evaluation Review, 50(4):44–46, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595260.

Li:2023:PAM

[3272] Wenxin Li. Performance analysis of modified SRPT in multiple-processor multitask scheduling. ACM SIGMET-RICS Performance Evaluation Review, 50(4):47–49, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3595244.3595261.

Hossen:2023:PEM

[3273] Md Rajib Hossen and Mohammad A. Islam. Practical efficient microservice autoscaling. ACM SIGMETRICS Performance Evaluation Review, 50(4):50– 52, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3595244.3595262.

Zhu:2023:R

[3274] Yinan Zhu, Chunhui Duan, and Xuan Ding. RoSense. ACM SIGMETRICS Performance Evaluation Review, 50(4): 53–55, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3595244.3595263.

Mittal:2023:SSR

[3275] Daksh Mittal, Sandeep Juneja, and Shubhada Agrawal. Shift, scale and restart smaller models to estimate larger ones: Agent based simulators in epidemiology. ACM SIGMETRICS Performance Evaluation Review, 50(4): 56–58, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3595244.3595264.

Choudhury:2023:THT

[3276] Tuhinangshu Choudhury, Weina Wang, and Gauri Joshi. Tackling heterogeneous traffic in multi-access systems via erasure coded servers. ACM SIGMET-RICS Performance Evaluation Review, 50(4):59–61, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3595244.3595265.

Regmi:2023:TDL

[3277] Hem Regmi and Sanjib Sur. Towards deep learning augmented robust Dband millimeter-wave picocell deployment. ACM SIGMETRICS Performance Evaluation Review, 50(4):62–64, March 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3595244.3595266.

Maruf:2023:MMD

[3278] Hasan Al Maruf, Yuhong Zhong, Hongyi Wang, Mosharaf Chowdhury, Asaf Cidon, and Carl Waldspurger. Memtrade: Marketplace for disaggregated memory clouds. ACM SIGMET-RICS Performance Evaluation Review, 51(1):1–2, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3606376.3593553.

Addanki:2023:MNO

[3279] Vamsi Addanki, Chen Avin, and Stefan Schmid. Mars: Near-optimal throughput with shallow buffers in reconfigurable datacenter networks. ACM SIGMETRICS Performance Evaluation Review, 51(1):3–4, June 2023. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606376.
3593551.

Giannoula:2023:ASE

[3280] Christina Giannoula, Kailong Huang, Jonathan Tang, Nectarios Koziris, Georgios Goumas, Zeshan Chishti, and Nandita Vijaykumar. Architectural support for efficient data movement in fully disaggregated systems. ACM SIGMETRICS Performance Evaluation Review, 51(1):5–6, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593533.

Zerwas:2023:DHT

[3281] Johannes Zerwas, Csaba Györgyi, Andreas Blenk, Stefan Schmid, and Chen Avin. Duo: a high-throughput reconfigurable datacenter network using local routing and control. ACM SIGMETRICS Performance Evaluation Review, 51(1):7–8, June 2023. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606376.
3593537.

Lin:2023:TAD

[3282] Jiaxin Lin, Tao Ji, Xiangpeng Hao, Hokeun Cha, Yanfang Le, Xiangyao Yu, and Aditya Akella. Towards accelerating data intensive Application’s shuffle process using Smart-NICs. ACM SIGMETRICS Performance Evaluation Review, 51(1):9–10, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593577.

Zhu:2023:DBF

[3283] Wenzhe Zhu, Yongkun Li, Erci Xu, Fei Li, Yinlong Xu, and John C. S. Lui. DiffForward: On balancing forwarding traffic for modern cloud block services via differentiated forwarding. ACM SIGMETRICS Performance Evaluation Review, 51(1):11– 12, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593536.

Kumar:2023:SCD

[3284] Adithya Kumar, Anand Sivasubramaniam, and Timothy Zhu. SplitRPC: a control + data path splitting RPC stack for ML inference serving. ACM SIGMETRICS Performance Evaluation Review, 51(1):13–14, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606376.
3593571.

Gozlan:2023:GCB

[3285] Itamar Gozlan, Chen Avin, Gil Einziger, and Gabriel Scalosub. Go-to-controller is better: Efficient and optimal LPM caching with splicing. ACM SIGMETRICS Performance Evaluation Review, 51(1):15–16, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606376.
3593546.

DeSensi:2023:NCI

[3286] Daniele De Sensi, Tiziano De Matteis, Konstantin Taranov, Salvatore Di Girolamo, Tobias Rahn, and Torsten Hoefler. Noise in the clouds: Influence of network performance variability on application scalability. ACM SIGMET-RICS Performance Evaluation Review, 51(1):17–18, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3606376.3593555.

Liu:2023:SFF

[3287] Yi Liu, Shouqian Shi, Minghao Xie, Heiner Litz, and Chen Qian. Smash: Flexible, fast, and resource-efficient placement and lookup of distributed storage. ACM SIGMETRICS Performance Evaluation Review, 51(1):19–20, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3606376.3593569.

Jin:2023:SSL

[3288] Wangkai Jin and Xiangjun Peng. SLITS: Sparsity-lightened intelligent thread scheduling. ACM SIGMET-RICS Performance Evaluation Review, 51(1):21–22, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3606376.3593568.

Liu:2023:AAP

[3289] Hongyuan Liu, Sreepathi Pai, and Adwait Jog. Asynchronous automata processing on GPUs. ACM SIGMETRICS Performance Evaluation Review, 51(1): 23–24, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3606376.3593524.

Chen:2023:GGA

[3290] Canhui Chen and Zhixuan Fang. Gacha game analysis and design. ACM SIG-METRICS Performance Evaluation Review, 51(1):25–26, June 2023. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3606376.
                                        3593544.

Rutten:2023:MFA

[3291] Daan Rutten and Debankur Mukherjee. Mean-field analysis for load balancing on spatial graphs. ACM SIGMET-RICS Performance Evaluation Review, 51(1):27–28, June 2023. CODEN ????

ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3606376.3593552.

Allmeier:2023:BRM

[3292] Sebastian Allmeier and Nicolas Gast.

Bias and refinement of multiscale mean field models. ACM SIGMETRICS Performance Evaluation Review, 51(1):29– 30, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593527.

Salem:2023:ELT

[3293] Tareq Si Salem, Georgios Iosifidis, and Giovanni Neglia. Enabling long-term fairness in dynamic resource allocation. ACM SIGMETRICS Performance Evaluation Review, 51(1):31–32, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593541.

Ayvasik:2023:PPE

[3294] Serkut Ayvasik, Fidan Mehmeti, Edwin Babaians, and Wolfgang Kellerer. PEACH: Proactive and environment aware channel state information prediction with depth images. ACM SIGMETRICS Performance Evaluation Review, 51(1):33–34, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606376.
3593563.

Liu:2023:FPS

[3295] Yunzhuo Liu, Bo Jiang, Tian Guo, Zimeng Huang, Wenhao Ma, Xinbing Wang, and Chenghu Zhou. FuncPipe: a pipelined serverless framework for fast and cost-efficient training of deep learning models. ACM SIGMETRICS Performance Evaluation Review, 51(1): 35–36, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593543.

K:2023:CPA

[3296] Prashanthi S. K, Sai Anuroop Kesanapalli, and Yogesh Simmhan. Characterizing the performance of accelerated Jetson edge devices for training deep learning models. ACM SIGMETRICS Performance Evaluation Review, 51(1): 37–38, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593530.

Abyaneh:2023:MMA

[3297] Ali Hossein Abbasi Abyaneh, Maizi Liao, and Seyed Majid Zahedi. Malcolm: Multi-agent learning for cooperative load management at rack scale. ACM SIGMETRICS Performance Evaluation Review, 51(1):39–40, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593550.

MacMillan:2023:CAO

[3298] Kyle MacMillan, Tarun Mangla, James Saxon, Nicole P. Marwell, and Nick Feamster. A comparative analysis of Ookla Speedtest and Measurement Labs Network Diagnostic Test (NDT7). ACM SIGMETRICS Performance Evaluation Review, 51(1):41– 42, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593522.

Kumar:2023:EOP

[3299] Rashna Kumar, Sana Asif, Elise Lee, and Fabián E. Bustamante. Each at its own pace: Third-party dependency and centralization around the world. ACM SIGMETRICS Performance Evaluation Review, 51(1):43–44, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593539.

Lu:2023:DMA

[3300] Haoran Lu, Qingchuan Zhao, Yongliang Chen, Xiaojing Liao, and Zhiqiang Lin. Detecting and measuring aggressive location harvesting in mobile apps via data-flow path embedding. ACM SIGMETRICS Performance Evaluation Review, 51(1):45–46, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606376.
3593535.

Hsu:2023:FLI

[3301] Amanda Hsu, Frank Li, and Paul Pearce. Fiat Lux: Illuminating IPv6 apportionment with different datasets. ACM SIGMETRICS Performance Evaluation Review, 51(1):47– 48, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593542.

Chiesa:2023:NMM

[3302] Marco Chiesa and Fábio L. Verdi.

Network monitoring on multi-pipe switches. ACM SIGMETRICS Performance Evaluation Review, 51(1):49– 50, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593554.

Wang:2023:RTS

[3303] Haibo Wang, Dimitrios Melissourgos, Chaoyi Ma, and Shigang Chen. Realtime spread burst detection in data streaming. ACM SIGMETRICS Performance Evaluation Review, 51(1):51– 52, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593566.

Naseer:2023:JCF

[3304] Usama Naseer and Theophilus A. Benson. JS capsules: a framework for capturing fine-grained JavaScript memory measurements for the mobile Web. ACM SIGMETRICS Performance Evaluation Review, 51(1):53–54, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593548.

Banerjee:2023:OFA

[3305] Siddhartha Banerjee, Chamsi Hssaine, and Sean R. Sinclair. Online fair allocation of perishable resources. ACM SIGMETRICS Performance Evaluation Review, 51(1):55–56, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606376.
3593558.

Liu:2023:DPP

[3306] Mozhengfu Liu and Xueyan Tang. Dynamic bin packing with predictions. ACM SIGMETRICS Performance Evaluation Review, 51(1):57–58, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593538.

Sun:2023:OKP

[3307] Bo Sun, Lin Yang, Mohammad Hajiesmaili, Adam Wierman, John C. S. Lui, Don Towsley, and Danny H. K. Tsang. The online knapsack problem with departures. ACM SIGMETRICS Performance Evaluation Review, 51(1): 59–60, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593576.

Li:2023:PKB

[3308] Fengjiao Li, Xingyu Zhou, and Bo Ji.

(Private) kernelized bandits with distributed biased feedback. ACM SIG-METRICS Performance Evaluation Review, 51(1):61–62, June 2023. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606376.
3593565.

Balseiro:2023:ORA

[3309] Santiago Balseiro, Christian Kroer, and Rachitesh Kumar. Online resource allocation under horizon uncertainty. ACM SIGMETRICS Performance Evaluation Review, 51(1):63–64, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593559.

Cui:2023:SAC

[3310] Shuang Cui, Kai Han, Jing Tang, He Huang, Xueying Li, and Zhiyu Li. Streaming algorithms for constrained submodular maximization. ACM SIG-METRICS Performance Evaluation Review, 51(1):65–66, June 2023. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3606376.
                                        3593573.

Vial:2023:RMA

[3311] Daniel Vial, Sanjay Shakkottai, and R. Srikant. Robust multi-agent bandits over undirected graphs. ACM SIGMETRICS Performance Evaluation Review, 51(1):67–68, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3606376.
                                        3593567.

Mhaisen:2023:ONR

[3312] Naram Mhaisen, Abhishek Sinha, Georgios Paschos, and George Iosifidis. Optimistic no-regret algorithms for discrete caching. ACM SIGMETRICS Performance Evaluation Review, 51(1): 69–70, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3606376.3593561.

Rutten:2023:SOO

[3313] Daan Rutten, Nicolas Christianson, Debankur Mukherjee, and Adam Wierman. Smoothed online optimization with unreliable predictions. ACM SIGMETRICS Performance Evaluation Review, 51(1):71–72, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3606376.
                                        3593570.

Yu:2023:OAS

[3314] Jing Yu, Dimitar Ho, and Adam Wierman. Online adversarial stabilization of unknown networked systems. ACM SIGMETRICS Performance Evaluation Review, 51(1):73–74, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593557.

Williams:2023:MMK

[3315] Jalani K. Williams, Mor Harchol-Balter, and Weina Wang. The M/M/k with deterministic setup times. ACM SIGMETRICS Performance Evaluation Review, 51(1):75–76, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606376.
3593575.

Fu:2023:JLC

[3316] Xinzhe Fu and Eytan Modiano. Joint learning and control in stochastic queueing networks with unknown utilities. ACM SIGMETRICS Performance Evaluation Review, 51(1):77–78, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593547.

Wei:2023:CRP

[3317] Yehua Wei, Jiaming Xu, and Sophie H. Yu. Constant regret primaldual policy for multi-way dynamic matching. ACM SIGMETRICS Performance Evaluation Review, 51(1):79– 80, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593532.

Huo:2023:BEM

[3318] Dongyan (Lucy) Huo, Yudong Chen, and Qiaomin Xie. Bias and extrapolation in Markovian linear stochastic approximation with constant stepsizes. ACM SIGMETRICS Performance Evaluation Review, 51(1):81–82, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593526.

Zhang:2023:GCL

[3319] Yizhou Zhang, Guannan Qu, Pan Xu, Yiheng Lin, Zaiwei Chen, and Adam Wierman. Global convergence of localized policy iteration in networked multi-agent reinforcement learning. ACM SIGMETRICS Performance Evaluation Review, 51(1):83– 84, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593545.

Sam:2023:OLHb

[3320] Tyler Sam, Yudong Chen, and Christina Lee Yu. Overcoming the long horizon barrier for sample-efficient reinforcement learning with latent lowrank structure. ACM SIGMETRICS Performance Evaluation Review, 51(1): 85–86, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593562.

Zhang:2023:DMS

[3321] Mingming Zhang, Xiang Li, Baojun Liu, JianYu Lu, Yiming Zhang, Jianjun Chen, Haixin Duan, Shuang Hao, and Xiaofeng Zheng. Detecting and measuring security risks of hosting-based dangling domains. ACM SIGMET-RICS Performance Evaluation Review, 51(1):87–88, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3606376.3593534.

Goldberg:2023:BTU

[3322] Alexander Goldberg, Giulia Fanti, and Nihar B. Shah. Batching of tasks by users of pseudonymous forums: Anonymity compromise and protection. ACM SIGMETRICS Performance Evaluation Review, 51(1):89–90, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593525.

Wang:2023:CCT

[3323] Kailong Wang, Yuxi Ling, Yanjun Zhang, Zhou Yu, Haoyu Wang, Guangdong Bai, Beng Chin Ooi, and Jin Song Dong. Characterizing cryptocurrencythemed malicious browser extensions. ACM SIGMETRICS Performance Evaluation Review, 51(1):91–92, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3606376.3593529.

Tang:2023:SLR

[3324] Weizhao Tang, Lucianna Kiffer, Giulia Fanti, and Ari Juels. Strategic latency reduction in blockchain peer-to-peer networks. ACM SIGMETRICS Performance Evaluation Review, 51(1): 93–94, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593572.

Varma:2023:PDC

[3325] Sushil Mahavir Varma, Francisco Castro, and Siva Theja Maguluri. Powerof-d choices load balancing in the sub-Halfin Whitt regime. ACM SIGMET-RICS Performance Evaluation Review, 51(1):95–96, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3606376.3593564.

VanHoudt:2023:SAI

[3326] Benny Van Houdt. On the stochastic and asymptotic improvement of firstcome first-served and nudge scheduling. ACM SIGMETRICS Performance Evaluation Review, 51(1):97–98, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593556.

Grosof:2023:OSMa

[3327] Isaac Grosof, Ziv Scully, Mor Harchol-Balter, and Alan Scheller-Wolf. Optimal scheduling in the multiserverjob model under heavy traffic. ACM SIGMETRICS Performance Evaluation Review, 51(1):99–100, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606376.
3593560.

Kadota:2023:SRP

[3328] Igor Kadota, Dror Jacoby, Hagit Messer, Gil Zussman, and Jonatan Ostrometzky. Switching in the rain: Predictive wireless x-haul network reconfiguration. ACM SIGMETRICS Performance Evaluation Review, 51(1): 101–102, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3606376.3593574.

Liu:2023:FLW

[3329] Ruofeng Liu and Nakjung Choi. A first look at Wi-Fi 6 in action: Throughput, latency, energy efficiency, and security. ACM SIGMETRICS Performance Evaluation Review, 51(1):103– 104, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3606376.3593523.

Zhang:2023:CCB

[3330] Ding Zhang, Panneer Selvam Santhalingam, Parth Pathak, and Zizhan Zheng. CoBF: Coordinated beamforming in dense mmWave networks. ACM SIGMETRICS Performance Evaluation Review, 51(1):105–106, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606376.
3593531.

Liu:2023:LPM

[3331] Yilin Liu, Shijia Zhang, Mahanth Gowda, and Srihari Nelakuditi. Leveraging the properties of mmWave signals for 3D finger motion tracking for interactive IoT applications. ACM SIGMETRICS Performance Evaluation Review, 51(1):107–108, June 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606376.
3593549.

Squillante:2023:SIW

[3332] Mark S. Squillante. Special issue on The Workshop on MAthematical performance Modeling and Analysis (MAMA 2023). ACM SIGMETRICS Performance Evaluation Review, 51(2): 2, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3626570.3626572.

Pan:2023:SCO

[3333] Weici Pan and Zhenhua Liu. Switching constrained online convex optimization with predictions and feedback delays. ACM SIGMETRICS Performance Evaluation Review, 51(2):3–5, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3626570.3626573.

Grosof:2023:NSR

[3334] Isaac Grosof, Mor Harchol-Balter, and Alan Scheller-Wolf. New stability results for multiserver-job models via product-form saturated systems. ACM SIGMETRICS Performance Evaluation Review, 51(2):6–8, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3626570.3626574.

Anton:2023:SDR

[3335] E. Anton and K. Gardner. The stationary distribution of the redundancyd model with random order of service. ACM SIGMETRICS Performance Evaluation Review, 51(2):9–11, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3626570.3626575.

Xie:2023:ILS

[3336] Runhan Xie, Kristen Gardner, and Rhonda Righter. Insensitivity for loss systems with compatibilities. ACM SIGMETRICS Performance Evaluation Review, 51(2):12–14, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3626570.3626576.

Bor:2023:FME

[3337] Julianna Bor, Giuliano Casale, William Knottenbelt, Evgenia Smirni, and Andreas Stathopoulos. Fitting with matrix exponential mixtures generated by discrete probabilistic scaling. ACM SIGMETRICS Performance Evaluation Review, 51(2):15–17, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3626570.3626577.

Zhao:2023:ORM

[3338] Zhisheng Zhao and Debankur Mukherjee. Optimal rate-matrix pruning for heterogeneous systems. ACM SIGMETRICS Performance Evaluation Review, 51(2):18–20, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3626570.3626578.

Rutten:2023:DRS

[3339] Daan Rutten, Martin Zubeldia, and Debankur Mukherjee. Distributed rate scaling in large-scale service systems. ACM SIGMETRICS Performance Evaluation Review, 51(2):21–23, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3626570.3626579.

Jhunjhunwala:2023:ETB

[3340] Prakirt Jhunjhunwala, Daniela Hurtado-Lange, and Siva Theja Maguluri. Exponential tail bounds on queues. ACM SIGMETRICS Performance Evaluation Review, 51(2):24–26, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/10.
                                        1145/3626570.3626580.

Ciucu:2023:USQ

[3341] Florin Ciucu, Sima Mehri, and Amr Rizk. On ultra-sharp queueing bounds. ACM SIGMETRICS Performance Evaluation Review, 51(2):27–29, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3626570.3626581.

Gast:2023:WOP

[3342] Nicolas Gast, Bruno Gaujal, and Kimang Khun. What is an optimal policy in time-average MDP? ACM SIGMETRICS Performance Evaluation Review, 51(2):30–32, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3626570.3626582.

Hong:2023:PGP

[3343] Yige Hong and Ziv Scully. Performance of the Gittins Policy in the G/G/1 and G/G/k, with and without setup times. ACM SIGMETRICS Performance Evaluation Review, 51(2):33–35, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3626570.3626583.

Xie:2023:RHT

[3344] Runhan Xie and Ziv Scully. Reducing heavy-traffic response time with asymmetric dispatching. ACM SIGMET-RICS Performance Evaluation Review, 51(2):36–38, September 2023. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3626570.
3626584.

Lu:2023:MDP

[3345] Yingdong Lu, Mark S. Squillante, and Chai Wah Wu. Markov decision process framework for control-based reinforcement learning. ACM SIGMETRICS Performance Evaluation Review, 51(2): 39–41, September 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3626570.3626585.

Alouf:2023:ASS

[3346] Sara Alouf. ACM SIGMETRICS 2023 Student Research Competition. ACM SIGMETRICS Performance Evaluation Review, 51(3):2, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3639830.
3639832.

NovaesdeOliveira:2023:LOR

[3347] Amanda Camacho Novaes de Oliveira.

Learning the optimal representation dimension for restricted Boltzmann machines. ACM SIGMETRICS Performance Evaluation Review, 51(3):3–5, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3639830.3639833.

Niu:2023:GLB

[3348] Zifeng Niu. Graph learning based performance analysis for queueing networks. ACM SIGMETRICS Performance Evaluation Review, 51(3):6–7, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3639830.3639834.

Grosof:2023:RTM

[3349] Isaac Grosof. The RESET technique for multiserver-job analysis. ACM SIGMETRICS Performance Evaluation Review, 51(3):8–9, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3639830.
                                        3639835.

Hossen:2023:PCR

[3350] Md Rajib Hossen. PEMA+: a comprehensive resource manager for microservices. ACM SIGMETRICS Performance Evaluation Review, 51(3):10–12, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3639830.3639836.

Li:2023:DED

[3351] Yuanyuan Li. Distributed experimental design networks. ACM SIGMETRICS Performance Evaluation Review, 51(3): 13–15, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3639830.3639837.

He:2023:ZCE

[3352] Xiangan He. Zephyr: a costeffective, zero-knowledge light client for enhanced blockchain interoperability. ACM SIGMETRICS Performance Evaluation Review, 51(3):16–18, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3639830.3639838.

deSouza:2023:EEW

[3353] Beatriz Pereira de Souza, Marcio Nunes de Miranda, and Luiz Maltar Castello Branco. An energyefficient wireless sensor network applied to greenhouse cultivation. ACM SIGMETRICS Performance Evaluation Review, 51(3):19–21, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3639830.3639839.

Ahmed:2023:BPE

[3354] Nishat Ahmed, Amaan Rahman, and Lucia Rhode. Best practices for exoskeleton evaluation using DeepLab-Cut. ACM SIGMETRICS Performance Evaluation Review, 51(3):22–24, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3639830.3639840.

Huang:2023:EEB

[3355] Tianhao Huang, Xiaozhi Zhu, and Mo Niu. An end-to-end benchmarking tool for analyzing the hardwaresoftware implications of multi-modal DNNs. ACM SIGMETRICS Performance Evaluation Review, 51(3):25–27, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3639830.3639841.

Ayesta:2023:FCA

[3356] Urtzi Ayesta. Foreword from Chair of 2023 ACM SIGMETRICS Doctoral Dissertation Award Committee. ACM SIGMETRICS Performance Evaluation Review, 51(3):28, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3639830.3639843.

Grosof:2023:OSMb

[3357] Isaac Grosof. Optimal scheduling in multiserver queues. ACM SIGMET-RICS Performance Evaluation Review, 51(3):29–32, December 2023. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3639830.
                                        3639844.

Jhunjhunwala:2023:DAS

[3358] Prakirt Raj Jhunjhunwala. Design and analysis of stochastic processing and matching networks. ACM SIGMETRICS Performance Evaluation Review, 51(3):33–37, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3639830.3639845.

Sinclair:2023:ASO

[3359] Sean R. Sinclair. Adaptivity, structure, and objectives in sequential decisionmaking. ACM SIGMETRICS Performance Evaluation Review, 51(3):38–41, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3639830.3639846.

Corbett:2023:ESP

[3360] Matthew Corbett. Enhancing security and privacy in head-mounted augmented reality systems using eye gaze. ACM SIGMETRICS Performance Evaluation Review, 51(3):42–45, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3639830.3639848.

Liu:2023:DFI

[3361] Zhongdong Liu. Data freshness in information-update systems: Modeling, scheduling, and tradeoffs. ACM SIGMETRICS Performance Evaluation Review, 51(3):46–49, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3639830.3639849.

Ornee:2023:TTR

[3362] Tasmeen Zaman Ornee. Theory of timely remote estimation and application to situational awareness. ACM SIGMETRICS Performance Evaluation Review, 51(3):50–53, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3639830.3639850.

Shah:2023:DRC

[3363] Abhin Shah. Data-rich causal inference. ACM SIGMETRICS Performance Evaluation Review, 51(3):54–57, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3639830.3639851.

Shisher:2023:TIN

[3364] Md Kamran Chowdhury Shisher.

Timely inference over networks. ACM SIGMETRICS Performance Evaluation Review, 51(3):58–61, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3639830.3639852.

Tripathi:2023:IFM

[3365] Vishrant Tripathi. Information freshness for monitoring and control over wireless networks. ACM SIGMET-RICS Performance Evaluation Review, 51(3):62–65, December 2023. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3639830.
                                        3639853.

Zuo:2023:LEN

[3366] Jinhang Zuo. Learning-enabled networked systems. ACM SIGMETRICS Performance Evaluation Review, 51(3): 66–69, December 2023. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3639830.3639854.

Braverman:2024:FGC

[3367] Anton Braverman and Varun Gupta. Foreword from the general chairs. ACM SIGMETRICS Performance Evaluation Review, 51(4):2, March 2024. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3649477.
                                        3649479.

Barford:2024:MTP

[3368] Paul Barford, Maria Vlasiou, and Lei Ying. Message from the Technical Program Committee Chairs. ACM SIGMETRICS Performance Evaluation Review, 51(4):3, March 2024. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3649477.
                                        3649480.

Scherrer:2024:QCA

[3369] Simon Scherrer, Seyedali Tabaeiaghdaei, and Adrian Perrig. Quality competition among Internet service providers. ACM SIGMETRICS Performance Evaluation Review, 51(4):4– 5, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3649477.3649481.

Grosof:2024:RMT

[3370] Isaac Grosof, Yige Hong, Mor Harchol-Balter, and Alan Scheller-Wolf. The RESET and MARC techniques, with application to multiserver-job analysis. ACM SIGMETRICS Performance Evaluation Review, 51(4):6–7, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3649477.3649482.

Pai:2024:FCC

[3371] Meera Pai, Nikhil Karamchandani, and Jayakrishnan Nair. Fixed confidence community mode estimation. ACM SIGMETRICS Performance Evaluation Review, 51(4):8–9, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3649477.
3649483.

Saurav:2024:MAI

[3372] Kumar Saurav and Rahul Vaze. Minimizing age of information under arbitrary arrival model with arbitrary packet size. ACM SIGMETRICS Performance Evaluation Review, 51(4):10– 11, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3649477.3649484.

Hong:2024:PGP

[3373] Yige Hong and Ziv Scully. Performance of the Gittins policy in the G/G/1 and G/G/k, with and without setup times. ACM SIGMETRICS Performance Evaluation Review, 51(4):12–13, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3649477.3649485.

Yeger:2024:ATQ

[3374] Yaron Yeger, Onno Boxma, Jacques Resing, and Maria Vlasiou. ASIP tandem queues with consumption. ACM SIGMETRICS Performance Evaluation Review, 51(4):14–15, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3649477.
3649486.

Jhunjhunwala:2024:HTJ

[3375] Prakirt Raj Jhunjhunwala and Siva Theja Maguluri. Heavy traffic joint queue length distribution without resource pooling. ACM SIGMETRICS Performance Evaluation Review, 51(4):16–17, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3649477.3649487.

Jhunjhunwala:2024:ETB

[3376] Prakirt Raj Jhunjhunwala, Daniela Hurtado-Lange, and Siva Theja Maguluri. Exponential tail bounds on queues: a confluence of non- asymptotic heavy traffic and large deviations. ACM SIGMETRICS Performance Evaluation Review, 51(4):18–19, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3649477.3649488.

Guo:2024:PSO

[3377] Hengquan Guo, Hongchen Cao, Jingzhu He, Xin Liu, and Yuanming Shi. POBO: Safe and optimal resource management for cloud microservices. ACM SIGMETRICS Performance Evaluation Review, 51(4):20–21, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3649477.3649489.

Senapati:2024:OCO

[3378] Spandan Senapati and Rahul Vaze. Online convex optimization with switching cost and delayed gradients. ACM SIGMETRICS Performance Evaluation Review, 51(4):22–23, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3649477.
3649490.

Guan:2024:WNI

[3379] Xiaoding Guan, Noman Bashir, David Irwin, and Prashant Shenoy. WattScope: Non-intrusive applicationlevel power disaggregation in datacenters. ACM SIGMETRICS Performance Evaluation Review, 51(4):24–25, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3649477.3649491.

Zhao:2024:ORM

[3380] Zhisheng Zhao and Debankur Mukherjee. Optimal rate-matrix pruning for heterogeneous systems. ACM SIG-METRICS Performance Evaluation Review, 51(4):26–27, March 2024. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3649477.
3649492.

Olliaro:2024:SMJ

[3381] Diletta Olliaro, Marco Ajmone Marsan, Simonetta Balsamo, and Andrea Marin. The saturated multiserver job queuing model with two classes of jobs: Exact and approximate results. ACM SIGMETRICS Performance Evaluation Review, 51(4):28–29, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3649477.
3649493.

Agarwal:2024:RFP

[3382] Khushboo Agarwal and Veeraruna Kavitha. Robust fake-post detection against real-coloring adversaries. ACM SIGMETRICS Performance Evaluation Review, 51(4):30–31, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3649477.
3649494.

Zhou:2024:BCD

[3383] Boyang Zhou, Isaac Howenstine, Liang Cheng, and Steffen Bondorf. Breaking cyclic dependencies for network calculus using service partitioning. ACM SIGMETRICS Performance Evaluation Review, 51(4):32–42, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3649477.
3649495.

Ferragut:2024:QAI

[3384] Andres Ferragut and Fernando Paganini. Queueing analysis of imbalance between multiple server pools with an application to 3-phase EV charging. ACM SIGMETRICS Performance Evaluation Review, 51(4):43–53, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3649477.3649496.

Nguyen:2024:FRT

[3385] Hai Duc Nguyen and Andrew A. Chien.

A foundation for real-time applications on function-as-a-service. ACM SIGMETRICS Performance Evaluation Review, 51(4):54–65, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3649477.
3649497.

Cao:2024:COP

[3386] Ying Cao, Siyuan Yu, Xiaoqi Tan, and Danny H. K. Tsang. Competitive online path-aware path selection. ACM SIGMETRICS Performance Evaluation Review, 51(4):66–72, March 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3649477.3649498.

Gast:2024:ASI

[3387] Nicolas Gast and Benny Van Houdt.

Approximations to study the impact of the service discipline in systems with redundancy. ACM SIGMETRICS Performance Evaluation Review, 52(1):1– 2, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655045.

Cardinaels:2024:MDS

[3388] Ellen Cardinaels, Sem Borst, and Johan S. H. van Leeuwaarden. Multidimensional state space collapse in non-complete resource pooling scenarios. ACM SIGMETRICS Performance Evaluation Review, 52(1):3–4, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655067.

Yu:2024:STO

[3389] George Yu and Ziv Scully. Strongly tail-optimal scheduling in the lighttailed M/G/1. ACM SIGMETRICS Performance Evaluation Review, 52(1): 5–6, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655084.

Xie:2024:HTO

[3390] Runhan Xie, Isaac Grosof, and Ziv Scully. Heavy-traffic optimal sizeand state-aware dispatching. ACM SIGMETRICS Performance Evaluation Review, 52(1):7–8, June 2024. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3673660.
3655059.

Moura:2024:DDN

[3391] Giovane C. M. Moura, Marco Davids, Caspar Schutijser, Cristian Hesselman, John Heidemann, and Georgios Smaragdakis. Deep dive into NTP Pool’s popularity and mapping. ACM SIGMETRICS Performance Evaluation Review, 52(1):9–10, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655051.

Lyu:2024:MMM

[3392] Minzhao Lyu, Rahul Dev Tripathi, and Vijay Sivaraman. MetaVRadar: Measuring metaverse virtual reality network activity. ACM SIGMETRICS Performance Evaluation Review, 52(1): 11–12, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655065.

Ortiz:2024:SWE

[3393] Neil Ortiz, Alvaro A. Cardenas, and Avishai Wool. SCADA world: an exploration of the diversity in power grid networks. ACM SIGMETRICS Performance Evaluation Review, 52(1):13– 14, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655078.

Izhikevich:2024:DLS

[3394] Liz Izhikevich, Manda Tran, Katherine Izhikevich, Gautam Akiwate, and Zakir Durumeric. Democratizing LEO satellite network measurement. ACM SIGMETRICS Performance Evaluation Review, 52(1):15–16, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3673660.
3655052.

Dozier:2024:AFN

[3395] Kahlil Dozier, Loqman Salamatian, and Dan Rubenstein. Analysis of false negative rates for recycling Bloom filters (yes, they happen!). ACM SIG-METRICS Performance Evaluation Review, 52(1):17–18, June 2024. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3673660.
                                        3655044.

Mizrahi:2024:IBL

[3396] Avi Mizrahi, Daniella Bar-Lev, Eitan Yaakobi, and Ori Rottenstreich. Invertible Bloom lookup tables with listing guarantees. ACM SIGMETRICS Performance Evaluation Review, 52(1):19– 20, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3673660.3655060.

Monterubbiano:2024:LAR

[3397] Andrea Monterubbiano, Jonatan Langlet, Stefan Walzer, Gianni Antichi, Pedro Reviriego, and Salvatore Pontarelli. Lightweight acquisition and ranging of flows in the data plane. ACM SIGMETRICS Performance Evaluation Review, 52(1):21–22, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3673660.
                                        3655063.

Pandey:2024:TRT

[3398] Santosh Pandey, Amir Yazdanbakhsh, and Hang Liu. TAO: Re-thinking DL-based microarchitecture simulation. ACM SIGMETRICS Performance Evaluation Review, 52(1):23–24, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

                                        URL https://dl.acm.org/doi/10.
                                        1145/3673660.3655085.

Bakhshalipour:2024:AAS

[3399] Mohammad Bakhshalipour and Phillip B.

Gibbons. Agents of autonomy: a systematic study of robotics on modern hardware. ACM SIGMETRICS Performance Evaluation Review, 52(1):25– 26, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655043.

Iyer:2024:ABA

[3400] Venkatraman Iyer, Sungho Lee, Semun Lee, Juitem Joonwoo Kim, Hyunjun Kim, and Youngjae Shin. Automated backend allocation for multi-model, ondevice AI inference. ACM SIGMET-RICS Performance Evaluation Review, 52(1):27–28, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3673660.3655046.

Liu:2024:LOC

[3401] Qingsong Liu and Zhixuan Fang.

Learning the optimal control for evolving systems with converging dynamics. ACM SIGMETRICS Performance Evaluation Review, 52(1):29–30, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655062.

Pan:2024:SRE

[3402] Jiayu Pan, Yin Sun, and Ness B.

Shroff. Sampling for remote estimation of the Wiener process over an unreliable channel. ACM SIGMETRICS Performance Evaluation Review, 52(1): 31–32, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655077.

Tsanikidis:2024:NOP

[3403] Christos Tsanikidis and Javad Ghaderi. Near-optimal packet scheduling in multihop networks with end-to-end deadline constraints. ACM SIGMETRICS Performance Evaluation Review, 52(1): 33–34, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655069.

Zhang:2024:PCS

[3404] Yixuan Zhang, Dongyan (Lucy) Huo, Yudong Chen, and Qiaomin Xie. Prelimit coupling and steady-state convergence of constant-stepsize nonsmooth contractive SA. ACM SIGMETRICS Performance Evaluation Review, 52(1): 35–36, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655076.

Zhang:2024:MLS

[3405] Huaifeng Zhang, Mohannad Alhanahnah, Fahmi Abdulqadir Ahmed, Dyako Fatih, Philipp Leitner, and Ahmed Ali-Eldin. Machine learning systems are bloated and vulnerable. ACM SIGMETRICS Performance Evaluation Review, 52(1):37–38, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3673660.
3655064.

Cheng:2024:TCA

[3406] Scott Cheng, Jun-Liang Lin, Murali Emani, Siddhisanket Raskar, Sam Foreman, Zhen Xie, Venkatram Vishwanath, and Mahmut T. Kandemir.

Thorough characterization and analysis of large transformer model training at-scale. ACM SIGMETRICS Performance Evaluation Review, 52(1):39– 40, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655087.

Pan:2024:HPT

[3407] Yicheng Pan, Yang Zhang, Tingzhu Bi, Linlin Han, Yu Zhang, Meng Ma, Xiangzhuang Shen, Xinrui Jiang, Feng Wang, Xian Liu, and Ping Wang. HEAL: Performance troubleshooting deep inside data center hosts. ACM SIGMETRICS Performance Evaluation Review, 52(1):41–42, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655058.

Cai:2024:KVU

[3408] Peter Cai and Martin Karsten. Kernel vs. user-level networking: Don’t throw out the stack with the interrupts. ACM SIGMETRICS Performance Evaluation Review, 52(1):43–44, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655061.

Lechowicz:2024:OCS

[3409] Adam Lechowicz, Nicolas Christianson, Bo Sun, Noman Bashir, Mohammad Hajiesmaili, Adam Wierman, and Prashant Shenoy. Online conversion with switching costs: Robust and learning-augmented algorithms. ACM SIGMETRICS Performance Evaluation Review, 52(1):45–46, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3673660.3655074.

Lechowicz:2024:OPR

[3410] Adam Lechowicz, Nicolas Christianson, Jinhang Zuo, Noman Bashir, Mohammad Hajiesmaili, Adam Wierman, and Prashant Shenoy. The online pause and resume problem: Optimal algorithms and an application to carbonaware load shifting. ACM SIGMET-RICS Performance Evaluation Review, 52(1):47–48, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3673660.3655086.

Hanafy:2024:CLC

[3411] Walid A. Hanafy, Qianlin Liang, Noman Bashir, David Irwin, and Prashant Shenoy. CarbonScaler: Leveraging cloud workload elasticity for optimizing carbon-efficiency. ACM SIGMETRICS Performance Evaluation Review, 52(1): 49–50, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3673660.3655048.

Huang:2024:OCP

[3412] Yudi Huang, Yilei Lin, and Ting He. Optimized cross-path attacks via adversarial reconnaissance. ACM SIG-METRICS Performance Evaluation Review, 52(1):51–52, June 2024. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3673660.
                                        3655075.

Anghel:2024:WGM

[3413] Radu Anghel, Yury Zhauniarovich, and Carlos Gañán. Who’s got my back? Measuring the adoption of an Internet-wide BGP RTBH service. ACM SIGMETRICS Performance Evaluation Review, 52(1):53–54, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655090.

Choo:2024:LSS

[3414] Euijin Choo, Mohamed Nabeel, Doowon Kim, Ravindu De Silva, Ting Yu, and Issa Khalil. A large scale study and classification of VirusTotal reports on phishing and malware URLs. ACM SIGMETRICS Performance Evaluation Review, 52(1):55–56, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3673660.
3655042.

Yang:2024:OAR

[3415] Jianyi Yang, Pengfei Li, Mohammad J.

Islam, and Shaolei Ren. Online allocation with replenishable budgets: Worst case and beyond. ACM SIGMETRICS Performance Evaluation Review, 52(1): 57–58, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655073.

Aslan:2024:FRA

[3416] Fatih Aslan, George Iosifidis, Jose A.

Ayala-Romero, Andres Garcia-Saavedra, and Xavier Costa-Perez. Fair resource allocation in virtualized O-RAN platforms. ACM SIGMETRICS Performance Evaluation Review, 52(1):59–60, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3673660.3655054.

Wang:2024:BNO

[3417] Lingdong Wang, Simran Singh, Jacob Chakareski, Mohammad Hajiesmaili, and Ramesh K. Sitaraman. BONES: Near-optimal neural-enhanced video streaming. ACM SIGMETRICS Performance Evaluation Review, 52(1):61– 62, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3673660.3655047.

Jonatan:2024:SLP

[3418] Gilbert Jonatan, Haeyoon Cho, Hyojun Son, Xiangyu Wu, Neal Livesay, Evelio Mora, Kaustubh Shivdikar, José L. Abellán, Ajay Joshi, David Kaeli, and John Kim. Scalability limitations of processing-in-memory using real system evaluations. ACM SIGMETRICS Performance Evaluation Review, 52(1): 63–64, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3673660.3655079.

Bitchebe:2024:GEB

[3419] Stella Bitchebe, Yves Kone, Pierre Olivier, Jalil Boukhobza, Yérom-David Bromberg, Daniel Hagimont, and Alain Tchana. GuaNary: Efficient buffer overflow detection in virtualized clouds using Intel EPT-based sub-page write protection support. ACM SIGMET-RICS Performance Evaluation Review, 52(1):65–66, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3673660.3655056.

Akbarzadeh:2024:HBH

[3420] Negar Akbarzadeh, Sina Darabi, Atiyeh Gheibi-Fetrat, Amir Mirzaei, Mohammad Sadrosadati, and Hamid Sarbazi-Azad. A high-bandwidth high-capacity hybrid 3D memory for GPUs. ACM SIGMETRICS Performance Evaluation Review, 52(1):67–68, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3673660.3655057.

Harris:2024:SDM

[3421] Keegan Harris, Anish Agarwal, Chara Podimata, and Zhiwei Steven Wu. Strategyproof decision-making in panel data settings and beyond. ACM SIG-METRICS Performance Evaluation Review, 52(1):69–70, June 2024. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3673660.
3655083.

Dierks:2024:WSP

[3422] Ludwig Dierks and Makoto Yokoo.

When should prices stay fixed? On the chances and limitations of spot pricing in larger markets. ACM SIGMETRICS Performance Evaluation Review, 52(1): 71–72, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655089.

Cheng:2024:CQB

[3423] Jin Cheng, Ningning Ding, John C. S.

Lui, and Jianwei Huang. Continuous query-based data trading. ACM SIGMETRICS Performance Evaluation Review, 52(1):73–74, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3673660.
3655050.

Lo:2024:SVT

[3424] Chi-Jen (Roger) Lo, Mahesh K. Marina, Nishanth Sastry, Kai Xu, Saeed Fadaei, and Yong Li. Shrinking VOD traffic via Rényi–entropic optimal transport. ACM SIGMETRICS Performance Evaluation Review, 52(1):75– 76, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3673660.3655081.

Shi:2024:CLI

[3425] Ruizhe Shi, Ruizhi Cheng, Bo Han, Yue Cheng, and Songqing Chen. A closer look into IPFS: Accessibility, content, and performance. ACM SIGMETRICS Performance Evaluation Review, 52(1): 77–78, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3673660.3655040.

Roy:2024:SMB

[3426] Rohan Basu Roy and Devesh Tiwari. StarShip: Mitigating I/O bottlenecks in serverless computing for scientific workflows. ACM SIGMETRICS Performance Evaluation Review, 52(1):79– 80, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3673660.3655082.

Zhang:2024:NFA

[3427] Yiguang Zhang, Reetahan Mukhopadhyay, and Augustin Chaintreau. Network fairness ambivalence: When does social network capital mitigate or amplify unfairness? ACM SIGMETRICS Performance Evaluation Review, 52(1): 81–82, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655072.

Lindstaahl:2024:CPD

[3428] Simon Lindståhl, Alexandre Proutiere, and Andreas Johnsson. Change point detection with adaptive measurement schedules for network performance verification. ACM SIGMETRICS Performance Evaluation Review, 52(1):83– 84, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655049.

Jiang:2024:NND

[3429] Xi Jiang, Shinan Liu, Aaron Gember-Jacobson, Arjun Nitin Bhagoji, Paul Schmitt, Francesco Bronzino, and Nick Feamster. NetDiffusion: Network data augmentation through protocolconstrained traffic generation. ACM SIGMETRICS Performance Evaluation Review, 52(1):85–86, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3673660.
3655071.

Huang:2024:MMM

[3430] Jintao Huang, Ningyu He, Kai Ma, Jiang Xiao, and Haoyu Wang. Miracle or mirage? A measurement study of NFT rug pulls. ACM SIGMETRICS Performance Evaluation Review, 52(1): 87–88, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655066.

Li:2024:TUC

[3431] Kai Li, Shixuan Guan, and Darren Lee. Towards understanding and characterizing the arbitrage bot scam in the wild. ACM SIGMETRICS Performance Evaluation Review, 52(1):89– 90, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3673660.3655088.

Chu:2024:FLQ

[3432] Tianyue Chu and Nikolaos Laoutaris. FedQV: Leveraging quadratic voting in federated learning. ACM SIGMET-RICS Performance Evaluation Review, 52(1):91–92, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3673660.3655055.

Hong:2024:NOS

[3433] Yige Hong, Qiaomin Xie, and Weina Wang. Near-optimal stochastic binpacking in large service systems with time-varying item sizes. ACM SIG-METRICS Performance Evaluation Review, 52(1):93–94, June 2024. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3673660.
                                        3655070.

Rutten:2024:DSS

[3434] Daan Rutten, Martin Zubeldia, and Debankur Mukherjee. Distributed speed scaling in large-scale service systems. ACM SIGMETRICS Performance Evaluation Review, 52(1):95–96, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3673660.3655053.

Goldsztajn:2024:SSS

[3435] Diego Goldsztajn, Sem C. Borst, and Johan S. H. van Leeuwaarden. Server saturation in skewed networks. ACM SIGMETRICS Performance Evaluation Review, 52(1):97–98, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3673660.3655080.

Ramanathan:2024:XIC

[3436] Alagappan Ramanathan, Rishika Sankaran, and Sangeetha Abdu Jyothi. Xaminer: an Internet cross-layer resilience analysis tool. ACM SIGMET-RICS Performance Evaluation Review, 52(1):99–100, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3673660.3655091.

Ramanathan:2024:NFC

[3437] Alagappan Ramanathan and Sangeetha Abdu Jyothi. Nautilus: a framework for cross-layer cartography of submarine cables and IP links. ACM SIGMET-RICS Performance Evaluation Review, 52(1):101–102, June 2024. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3673660.
3655068.

Carisimo:2024:HEV

[3438] Esteban Carisimo, Caleb J. Wang, Mia Weaver, Fabián Bustamante, and Paul Barford. A hop away from everywhere: a view of the intercontinental long-haul infrastructure. ACM SIGMETRICS Performance Evaluation Review, 52(1): 103–104, June 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3673660.3655041.

Squillante:2024:SIW

[3439] Mark S. Squillante. Special issue on the workshop on MAthematical performance Modeling and Analysis (MAMA 2024). ACM SIGMETRICS Performance Evaluation Review, 52(2): 2, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3695411.3695413.

Jazi:2024:OCG

[3440] Hossein Nekouyan Jazi and Faraz Zargari. Online conversion with group fairness constraints. ACM SIGMETRICS Performance Evaluation Review, 52(2):3–5, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3695411.3695414.

Christianson:2024:RSO

[3441] Nicolas Christianson, Bo Sun, Steven Low, and Adam Wierman. Risksensitive online algorithms. ACM SIGMETRICS Performance Evaluation Review, 52(2):6–8, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3695411.3695415.

Ferragut:2024:TBP

[3442] Andres Ferragut, Matias Carrasco, and Fernando Paganini. Timerbased pre-fetching for increasing hazard rates. ACM SIGMETRICS Performance Evaluation Review, 52(2):9–11, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3695411.3695416.

Li:2024:HPQ

[3443] Zhouzi Li, Mor Harchol-Balter, and Alan Scheller-Wolf. Hybrid priority queue and its applications. ACM SIGMETRICS Performance Evaluation Review, 52(2):12–14, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3695411.3695417.

Harlev:2024:GPO

[3444] Amit Harlev, George Yu, and Ziv Scully. A gittins policy for optimizing tail latency. ACM SIGMETRICS Performance Evaluation Review, 52(2): 15–17, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

org/doi/10.1145/3695411.3695418.

Ramakrishna:2024:TAP

[3445] Shefali Ramakrishna and Ziv Scully.

Transform analysis of preemption overhead in the M/G/1. ACM SIGMET-RICS Performance Evaluation Review, 52(2):18–20, September 2024. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3695411.
3695419.

Charlet:2024:TON

[3446] Nils Charlet and Benny Van Houdt.

Tail optimality of the Nudge-M scheduling algorithm. ACM SIGMET-RICS Performance Evaluation Review, 52(2):21–23, September 2024. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3695411.
3695420.

Grosof:2024:BMG

[3447] Isaac Grosof and Ziyuan Wang. Bounds on M/G/k scheduling under moderate load improving on SRPT-k and tightening lower bounds. ACM SIGMETRICS Performance Evaluation Review, 52(2):24–26, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3695411.3695421.

Chen:2024:SPM

[3448] Zhongrui Chen, Isaac Grosof, and Benjamin Berg. Simple policies for multiresource job scheduling. ACM SIGMETRICS Performance Evaluation Review, 52(2):27–29, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3695411.3695422.

Liu:2024:OSC

[3449] Larkin Liu, Shiqi Liu, and Matej Jusup. Optimizing stochastic control through state transition separability and resource-utility exchange. ACM SIGMETRICS Performance Evaluation Review, 52(2):30–32, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3695411.3695423.

Pan:2024:NSB

[3450] Weici Pan and Zhenhua Liu. Nonstationary bandits with heavy tail. ACM SIGMETRICS Performance Evaluation Review, 52(2):33–35, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

URL https://dl.acm.org/doi/10.
1145/3695411.3695424.

Sankagiri:2024:PRF

[3451] Suryanarayana Sankagiri and Bruce Hajek. Pricing for routing and flow-control in payment channel networks. ACM SIGMETRICS Performance Evaluation Review, 52(2):36–38, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3695411.3695425.

Dai:2024:SSC

[3452] J. G. Dai, Jin Guang, and Yaosheng Xu. Steady-state convergence of the continuous-time JSQ system with general distributions in heavy traffic. ACM SIGMETRICS Performance Evaluation Review, 52(2):39–41, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3695411.3695426.

Kalantzis:2024:QAE

[3453] Vasileios Kalantzis, Mark S. Squillante, and Shashanka Ubaru. On quantum algorithms for efficient solutions of general classes of structured Markov processes. ACM SIGMETRICS Performance Evaluation Review, 52(2):42–44, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3695411.3695427.

Persone:2024:RTI

[3454] Vittoria de Nitto Persone and Y. C.

Tay. Report on the Third International Workshop on Teaching Performance Analysis of Computer Systems 2024. ACM SIGMETRICS Performance Evaluation Review, 52(2):45–48, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3695411.3695429.

Fiems:2024:TPA

[3455] Dieter Fiems. Teaching performance analysis: essential skills andlearning outcomes. ACM SIGMETRICS Performance Evaluation Review, 52(2):49– 52, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3695411.3695430.

Meo:2024:WST

[3456] Michela Meo. Why should I teach performance evaluation to students in networking? ACM SIGMETRICS Performance Evaluation Review, 52(2):53–57, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3695411.3695431.

Abad:2024:HCW

[3457] Cristina L. Abad. How can we teach workload modeling in CS systems classes? ACM SIGMETRICS Performance Evaluation Review, 52(2):58–62, September 2024. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3695411.3695432.